A Complete Bibliography of Publications in SIAM Journal on Numerical Analysis: 1964–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 3.30

Title word cross-reference

#11058 [Jel79].

(−1, 1) [DE79]. (ln z)^{2} [RW89b]. (t, m, s) [LLNS96]. 1 [BM96, BK97a, Pet98b]. 1 + \sqrt{2} [TW88]. 2 [AL96a, AC98, BH93, HMM98, Kuz99]. 2r [Die90]. 2 × 2 [Vil86]. 3 [Gui87]. 9 [ALY88]. > 1 [GO81]. A [Bar99, Hil97, Bog71, BC89, Buz77, Car72, CH95, Kam78, Lin83, You72a, Zir82]. A(0) [Jel76]. A_{0} [Jel76, Fei82]. Av = b [Bor76]. Ax = b [DT73, Was82]. Ax = λBx [PW70, Ste72]. B [FSU81, GC98, BN74, DM82, Far85, Hak82, Mey87b]. β [Hu96]. C^{0} [Arc78, Whe77]. C^{1} [HS97, APS87, ASW93, AD92, Her85, KW85, PS92, Pet95, Zha96, PW80]. C^{2} [LS97a, Del93]. C^{m} [Bub78]. cosh μz [RW89a]. D [van81c]. δ [Por73]. Δ^{2}u = f [Sig76]. d ≥ 8 [ASW93]. e^{−x} [Sch82]. E^{3} [Wix78]. E_{n}(f) [EL73]. ϵ [Han93b]. exp(−x) [Ise81b]. exp(tA) [Kam78]. G [MF97, BC72, RSK76, GAM71]. GBQ [VH93]. H [LB97, Ain96, BG88, GHS96, GC98, IB97, MS97b, SS97, TD82, VSM80]. H^{−1} [KW76, Día79, FKD84, Whe75]. H^{1} [Pan98, BC94a]. H^{1}(0, s) [Mar83]. H^{m,∞} [CS74b]. H^{p} [Sin73]. H^{p, p/2} [Hac81]. H_{2} [GMP93]. hp [MS98, OPF97]. k [KRD78, Par87a, Rag85, PS80, Pet65]. K(a_{n}/1) [JTW83]. K × K [PS80]. L [Cha94, Koc84, ST76]. L^{1} [STW97, hTZ97, DS68, DS84]. L^{2} [CL90, Den74a, Dup73b, HL87, Pat76, Sch82, Sch69b, Sch69c, LP78, Sch71b]. L^{2}(E_{ρ}) [Bar67a, Bar67b]. L^{∞} [FR78, KM88, LK74, Sch69d, Dob80, LP78]. l^{p} [Shr73]. L_{0} [KT86]. l_{1} [BR73, BR78b, BCS78, AH77, EAVD79, HM85, Kam79a, MS76, Uso67a, Uso67b]. L_{2} [Joe99, GV74, Bra94, GMP93, MS76, Sha85b, Whe73a]. l_{∞} [BCC78, Cad73, Cad74, BCL89a, BCL89b, CL92, MT75, Whe73b, Whe73c, BPR72]. L_{∞}L_{2} [EJ95a]. L_{∞}L_{∞} [EJ95a]. L_{p} [Asa89, Bre77, FGH74]. L_{q} [Kro81]. λ [JS83, Kub70]. [−1, 1] [Bru91]. LU [PP73a]. LZ [Kau74]. m [GW90, El 90, Por71, Prü82]. MGR[ν] [KP87]. n [EL73, Geo77, GPV78, GM78c, KS99, Str69a, Str72, Str73, Str75b, GK96]. n + 1 [BG71]. n + 2 [MS65c]. ∇^{4}u = f [Smi73]. O(h^{4}) [Arc77, HVR88]. O(n^{2}) [BR75]. p [Ain96, BSK81, BS87a, BG88, BCMP91, BL94, Dor84, Dor86, Eri86, GHS96, HR99, IB97, KW85, LB96a, MS97b, SS96a, SS97]. P1 [Pie95]. P^{1} [JNV99, JN98]. P^{4} [EGLP85]. P DL_{g} [dC90]. Q [Sor80]. QR [EH75, Nan85, PP73a, Ste77]. QZ [War75]. r [Die90, FM90b]. R^{1} [Ste75b]. R^{2} [Mae82]. R_{∞}^{2} [Gov86]. R^{3} [Gue92, Ned78, BvA94]. R^{n} [Bac80, Bus85, Can83]. s [Vin86, KD83]. SOR [KL89]. t [ST76, SL83a, Jef69]. T_{n+1}(X) [RJ74]. Θ [KR94a]. u_{tt} = u_{xx} [RW89a, RW89b]. V [Neu98, BH83, McC85]. W [BSW95]. x′ = f(t, x) [Ste70]. y = f(t, y) [Cha85b]. y′′ = F (x, y, y′) [Hen70b]. Z_{2} [JSC91].

-Algorithm [Kau74, VH93]. -and [Jel76]. -Approximation [BPR72, FGH74, Sch69b, EAVD79, Kam79a]. -approximations [Kro81]. -Box [Str72]. -Collocation-Finite [Whe77]. -continuity [KW85]. -contractive [Lin83]. -contractivity [CH95]. -convergence [Dob80, FSU81]. -convex [KRD78]. -cycle [BH83, McC85, Neu98]. -D [Gui87]. -Dependent [GW90]. -Error [FR78, LK74, MT75, KM88, STW97]. -Estimates [Bre77, Dup73b]. -Estimation [MS76]. -Estimator [MS76]. -Fractions [Jef69]. -free [Han93b]. -Function [Prü82]. -functional [Rag85]. -Functions [El 90, Por71, MF97]. -Galerkin [FKD84, Día79, Pan98]. -Inverses [RSK76]. -Laplacian [BL94]. -Line [PS80]. -Matrices [Kub70, JS83, VSM80]. -methods [BSW95]. -Multivariate [Sch69d, Sch69c]. -Nets [LLNS96]. -Norm [Bra94, Cad74, BC94a, EAVD79]. -Nyström [vS87]. -order [FM90b]. -orthogonal [Vin86]. -P.D [Pet65]. -part [Par87a]. -Perfect [ST76]. -point [ALY88]. -polynomial [Hu96]. -Polynomials [Por73]. -Problems [Sch82]. -Rate [hTZ97]. -Scaling [GV74, Sha85b]. -scheme [KR94a]. -sets [TD82]. -Simplex [Str69a, GM78c]. -Solutions [DS68]. -Sphere [Str73, Str75b]. -spline [Hak82]. -Splines [BC72, CS74b, Koc84, ST76, DM82, Far85, Mey87b]. -Stability [Jel76, van81c, Bar99]. -Stable [Bog71, Buz77, Car72, KT86, BC89, Cha94, Fei82, Zir82, Hil97]. -superlinear [Sor80]. -surface [Pet95]. -symmetry [JSC91]. -System [HR99]. -Transformation [SL83a]. -Transformations [GAM71]. -Version [BSK81, BS87a, Dor84, Dor86, Eri86, KW85, LB96a, SS96a, LB97, BCMP91]. -Widths [KS99].

1 [But65, Ost65]. 10 [But65]. 12 [DF76a]. 13 [Cra78, Hil78]. 130 [Ost65]. 14 [Jel79]. 15 [HL82, de 78a]. 16 [Sob79a]. 17 [Sym81]. 18 [DW82, Laz82]. 1983 [Bir83].

2 [But65]. 2-10 [But65]. 20 [GN83a]. 21 [MR85]. 22 [Gup85b]. 25 [CGT89]. 26 [BCL89a].

31 [LL95]. 33 [Hic97].

57 [Jel79].

6 [Pie71]. 60th [Bir83]. 65th [MSW94].

7 [Wei95].

80f [Sob79a]. 81k [Sym81]. 82i [Laz82]. 89h [CGT89].

9 [HEFS74]. 95g [LL95].

A-BDF [Fre98]. A-Convergence [Gek75]. A-Stable [SW94a]. A.D.I [FM67]. Abel [Atk74, Egg88, ML69, SP68]. Abel-type [Egg88]. Abramov [Pet98b]. Abscissa [Hen70a]. Abscissas [AF68]. Absorbing [BJR90, BDH98, CDW96, JT97]. Absorption [MMD94]. Abstract [CN88, DD93, GP98, JS68a, Pet65, Var67, VW95]. Accelerated [AP73, DF84, Del88, Del89, DF91, de 68, DGB82, GHY83]. Accelerating [Per67a, Wim77]. Acceleration [ABDH95, Bro95, Cor83, HLY80, Lei98, LB90, Osa90, SFS86, SF79, BD85, BDGB83, DK82, KS83, ST78, SL83b]. Accepting [WF97]. Accuracy [CM90a, DeL94, DMW83, GS96, Hig89b, JS87, RR96, Sni72, Ten94, War77, WES80, And83b, BX89, DN89, HK79, LeV88, MK80, Sch80, Ske82, Ske86, Tad86]. Accurate [BD90, Byr90, CJR97, Del93, Drm98a, HMMR95, Huy93, Huy95, Kel69, Kel74, KC93, LO96, SW99, Var71b, VWI97, AF89, Ber86, HO87, SO86, Vil89]. Achieve [CH97]. Achieving [Den74a]. Acoustic [BGT97, CJ96, Hol78, Wal88, FN80]. Active [BM88, Bur90]. Actual [CS74a]. Adams [CLM89, Fei82, Hal74, Ste79a]. Adaptation [VAF98]. Adapted [BNP91]. Adaptive [AE94, AK98, BL98, CW96, CS96, Dar91, Dör96, EJ91, EJ95a, EJ95b, EJ95c, EJL98, HK94, HW97b, Jen92, JNT90, LP77, Liu93, McC94, MS99b, MS99c, NPV93, OPF97, Rüd93, RC78, BR78a, Gup85a, Gup85b, Joh88, Luc85b, SC87, Sty89]. Adaptively [JNV99]. Adaptivity [SB84]. Addition [KSC99]. Additional [Heg92]. Additive [GC98, SA73]. ADI [CW97, DM79, Dyk87]. Adjoint [Don70, DKR68, EW96, JS98, Osb67, OS92]. Admissible [OPW65]. Admitting [CY77]. Adsorption [BK97a, BK97b, BK98, Bou98, DVW94]. Advection [AW95, Bec92, BHM^{+}99, Daw93, EW96, FF98, Fun93, GSY92, LeV96, Mat98, RS92, VS97, Cha84b]. Advection-Diffusion [Bec92, BHM^{+}99, Daw93, Fun93, Mat98, VS97, Cha84b]. Advection-Dominated [AW95]. Advection-Reaction [EW96]. Advective [Daw91]. Affine [BD99, DH79]. Aggregation [CM84, Hav87]. Aggregation/Disaggregation [CM84, Hav87]. aided [JP92]. AIDS [ILM^{+}96]. Airfoil [MS97a, Scu98, Mos83]. Algebra [Bab72, BL80]. Algebraic [AC90, AP91, AP95b, ABDH95, AV90, BGM96, BP89, Bro95, BP90, Cas82, DTW76, Eic93, FF70, GGM99, Gay83, GP84, HMM98, Her75, JK96, KG91, Kno92, Kun75, KM96, Lee91, LPG91, LM90b, LM90a, McC82, Ost64, Ost65, Pet98b, Pet86, RW95, SW94a, Sch90a, Asc89, BH93, Gea90, HS87, MMR88, MP78a, MP85, Mur82, Roc89]. Algebraically [Bur82, HW81, Bur87]. Algorithm [AP83, Alo97, AH77, Bac80, BT99, BPR72, BR73, BR78b, BS79, Ber95, BB71, Bog77, Bör90, BPV97, Bre92, Cad73, CF88, CW76, CGT91, DM95, Dör96, Drm98b, DL73, Dun74a, DT73, EH75, EA91, Erd65, FH72, GM78a, GPS76, GGK98, GCB74, Gir95, Gri74, Hel76, Hen66, HZ90, HY96, JT70, KS97a, Kau74, Kel95, KLM94, KW93, Lan77, Li88, Li98, Lin72, LM86, Mat72, MS73d, Nan85, PRS90, Prü82, PA73, RR87, Ree90, Rit90, Swe74, Swe77, TMV98, TH98, Tse96, Van75, Ver84, War75, Wat75, Wat79, Wol80, Wyn66, YR99, de 68, AS85, AG87, Bea82a, BT89, BSS87, CF89, CO80, Die82, DS89a, EAVD79, EGLP85, FS87, Fra87, GL78b, Geo80, GL80b, GM89, HB86, HP78b, Jer88, Kar78, Ker86, Ker88, Mar78, Mey87b]. algorithm [Saa82, SW82, Sha78, Sor80, Toi79, Var85, VH93, Wan93b, Whi86a]. Algorithmic [Cad74]. Algorithms [AH74, Asc77, BR77, Ban77, BL98, Bre73, Car91, CHL91, Coh72, CGT89, DTW78, DMP75, DM92, DM88, DD93, DSW94, Dus95, DGR96, El 90, EG94, Gar78, GL78a, GM78b, GHK94, Gra65, Gut93, Hal74, HYS94, HKT92, Hel74, Hin72, Hop87, Jac75, Kro73, Lam71, LeV96, Les91, LVQ95, LS76, MW95, Mey76, Mil76, Per67a, PR98, RSK76, Ruh73, Sch71a, Ste65a, Sun96, Ver88, VAF98, Wal96, Wan92, Wan93a, WM67, de 78a, dR81, CW93, CG84, CGT88, Dou84, HHS78, Hor83, LM79, MP78b, Naz79, NO85, Pet89, Que89, ST78, Sch85, SC87, SSB85, Sid86]. Aligned [Ott96a]. alloy [Whi85]. Almost [GS91b, Hes98a, Kah69, Kee71, Mah82a, Sch86]. Almost-Linear [Mah82a]. Almost-Triangular [Kee71]. Along [ELW91, EYL89, MP85, Rus85]. Alternate [Var76]. Alternating [Bad91, BF99, BEL89, CJZ83, CP91a, DF75, DF76a, Den77b, Den77a, DGP66, Dry78, EZB83, EW91, FF91, GM69b, Hub65, Mat98, MH98, Prü82, Rac68a, Sta91, Sto73, Dun80a, Hay81a, KHR89, RBC79]. Alternating-Direction [CJZ83, DF75, DF76a, Hay81a, KHR89]. Alternative [Dut90, Enr93]. Always [KC74]. Always-Convergent [KC74]. ambiguities [CLNP89]. Anal [BCL89a, CGT89, Cra78, DF76a, DW82, GN83a, Gup85b, Hic97, HL82, Hil78, HEFS74, Jel79, Laz82, LL95, MR85, Pie71, Sob79a, Sym81, Wei95, de 78a, But65, Ost65]. Analog [BH65]. Analogue [Cry67, Esp94, GB89]. Analogues [BHZ68, Naz86, Ste71a]. Analyse [Cia75].

Analyses [GYF86, Mei83]. Analysis [ALY88, Arb89, Asa86, ABDH95, BR81, BDR92, Bai91, BS97b, Ban82, BS93a, BW68, BW69, BSW85, Ben92, BDM^{+}95, BL92, Bia98, Boc97, BCMM98, BMM99, BFGP99, BPV97, Bra94, BQ86, BHM^{+}99, BC99, Bun71, Bun74a, Cas88, Cha87, Che95, CE97a, Cia75, CJ96, COP98, DG98a, Daw98, DF84, Den75, DSW94, DN91, DNW98, Dub90, Eis67b, Enr89, EH98, FF98, FR87a, Fer86a, For78, FR68, Gol89, GV80, Gre84, GLBN97, Gre94, GO83, Hal74, HDS97, HV96, Hei94, HR90, HR97, ILM^{+}96, IIMPL82, JP96, JS96, JPRT96, KS97a, KG91, KK98a, Kir93, KSC99, Lam67, LB97, LL98, LSZ97, LS76, Liu94, Lor84, LL99, LM92, MT89a, MG72, MS99a, Mil76, Mon92, MS94, Mor98, Nic92a, Nie99, NS93, Obi90, Oli74, Ott96b, Pal96, Pao99, PR98, Pet99, Rei99, Rul96, Saa92]. Analysis [SM96, Ske76, SS97, Sun99, Tan98, Tar96, TV96, VMSB97, Var67, VS97, VTG91, VIA96, VF93, Wan92, Wan93a, Wei92a, Wet97a, Wet97b, Woz74b, Zha97, vS89b, Bar88, BN89, Cha84b, CS87, CM88a, Chu87, EL96, EL89, FN80, FS91, FNO87, Gat88, GR82, GT85, GLT87b, JP92, Jer88, JS89, Ker88, LB96a, Lin88, Mey81a, Rom90, Rua89, SK87, SWC89, Sun83, Ver82, Vil89, Wan93b]. Analytic [BW68, BW69, CM65, EJL98, GGM99, GS96, Hen66, Let74a, Li83a, LM67, MS91, Pri76, Sle68, Ste75a, Wil78, BGO89, GV83, GHS96]. Analytic-Numerical [GGM99]. Analytical [Bec92]. analyze [DS89a]. Angle [ADZ96, Akr99, BA76, Kří92]. Angular [PS83]. Anharmonic [IIMPL82]. Anisotropic [AM94, Bec92, CDK95, Dzi99, Neu98]. Anisotropically [Nel73, AVG89]. Antidiffusion [LQ84]. Antilimits [CJ76]. antisymmetries [DS89a]. any [Jac84]. AOR [GHY83, PSP89]. Appl [But65, Ost65]. applicable [BR80].

Application [AB87, BH70, Buz77, CMM97a, CSS97, CD98, FU69, GM78a, GHS97, Ise81b, KM93, MR77, MF97, Mor72, Mor95, Neu98, Olv65, OR66a, OR67, PT97, Pet65, VW95, VIA96, WF97, Wyn64, Asc89, BM89, BN89, EAVD79, GS88, Gol89, LT88, LT89b, Mil79b, Ste80, Whi86a]. Applications [CDK95, Cor96, Dar91, DF91, DRV94, Drm98a, Gar78, Hel74, Hic96, Lee93, LTW91, Lor64, LT68, Ma95, MP77, MS97b, Pet92b, Pol74, Qn99, Van72, Var73, Var67, BMP89, DKW82, Del88, Del89, Dou84, GZ88, Hic97, LL83, Nic87, Whi86b]. Applied [AR95, AC98, BP90, Eic93, EC93, GGS94, HV96, KG91, LTW96, Pet86, PV72, Sch96, STW97, Tan98, Tra96, Yam75, Yav95, de 86, CT86, FSU85a, Gar81, SB84]. Approach [Alb87, Bak71, Bel69, Buc76, DE72, Dut90, Faz96, GANT98, Gol98a, Gol98b, IPS91, JS68a, JT97, Kru98, Kub70, McL74, MC93, Olv78, RT92, SS99, Smi68, Smi70, VMK97, DS86, GL88, JS83]. Approaches [Hil76, Hil78]. Approximability [KS97b]. Approximants [BT99, HP80]. Approximate [AT86, AH76, AE97, Atk75, Bor76, BO86, Bry68, Cal74, CW76, CR76, Che73, Che75, CP98, Cos71, DR65, DKR68, ET71, Fra71, GANT98, Gra71, GR96, Hen69, Hen70b, HR93, Kin72, Lin74, Lin77b, MJ77a, MT75, NT92, NT94, PCW95, Ros70a, RW95, San94, Sig76, Smi73, SH66, Ste75a, Ste70, Ste71c, Wah77, WW93b, BB88, BE79, Cha85a, Hac79, Jer88, KPJ82, MK80, RS89, Tav78]. Approximated [Ben92, BC94b]. Approximates [CW76]. Approximating [CD77, Dob92, GW92, Mon91, Pru73, TW93, LM88]. Approximation [AS80, AN66, Ain96, AH74, AC96a, Ans67, AH77, Aub68, AL78, BR70, Bao95, BG95, BL94, BK97a, BK97b, BPR72, Bar73, BR73, BR78b, Bas73b, BYLP98, Ber66, BLMPT92, BDSW66, BFGP99, BBG95, BJS73, Bou98, BH75, BS77, But64, But65, CK69, CFQ83, Cap99, Cha78a, Cha92, Cha73, CL64, CH97, CKL91, CHM88, CH94, Cur66, DDS80, DPZ87, DJ99, Der71, DRS95, DNW98, Dun71, Dun75, DC80, Dur83, EW72, EW76, EH89, Est95, Fal78, Far66, FGH74, Fun90, Fun93, GM75, Gor71, GLBN97, Gri74, Ham64, Han65b, HZ90, HR86, HR88, HR90, HY96, Hud68, HN75, IT91, JK91, Jim96, Joh75, Joh76, JK68, Jup78, KHK93, Kam76, Kem75, KL76, Kir93, KPH95, LS99, Lam76, Lan64c, LV90, LK74, LPG91, LT89b, LP91, LB96b]. Approximation [Loe66, LMST69, LM92, Man71, Man78, MG95, Mas73, MS66, MS99a, Mot66, MS65c, Mou65, Mou66, Mou68, MT68b, Mur71, NdF74, NV88, NV97, Osb67, Osb76, Ost64, Ost65, Pad97, Pel67, Por73, Red76, RHH99, Rei74, RD71a, RD79, Ros73, Ros70b, Row67, Row69, Rub64, San94, Sch70, Sch74, Sch90a, Sch82, Sch72, Sch69b, Sch69a, Sch69d, Sch69c, Sch68b, ST69, Sch69f, Sew66a, Sha74, Sho75, Sim68, Sta64, Sti64, Tal91, Tay68, TXZ89, Tse96, Uso67a, Uso67b, Vel92, Wat75, Wei69, Wei72b, Wei74a, Wei74b, Wil72, WO96, Zed66, vL92, AS85, AG87, ADK82, BO87, Bea82a, Bea82b, Ber82, BCM88, Bey87, Bra85, CFT84, DS84, Des81, Dob80, Don81, Dor84, Dor86, Dun80b, Dun80c, Dur88b, DN89, EAVD79, Ell83, EYL89, Gab85]. approximation [Ged78, GT82, Hei93b, HR82, IKP91, Jac88, Kam79a, Kam79b, Kam81, Kel81, KRD78, KR88, Lus79a, Mil79a, Mil79b, Nic82, Pet89, Rus85, Sam83, Sas78, SS96a, Sma88, Swa79, Swa81, Tre83, TG83, Var81, Whi85]. Approximations [BD76, BCM91, BD99, BSTW77, BT79, BT80, CGS92, CG98, CC92, CDMW98, CDMCW99, CD90, Cle64, Cos79, CDW96, Dav94, DO92, DMP75, Doe78, Don94, DG90, Dun97, ELW99, FL71, FT74, FHM67, Fra78, FR78, Fre90, Gek75, GCB74, Gus69, Hag76, Hag85, HL97, Ise81a, JY98, JP98, Joh69, Kee70, Kut70, Lor97, LT67, MQ82a, Mat72, Mir71, MTL96, Nor78, OR76, PR95, Pat76, PV72, PS83, Prü82, QZ92, Rac73, RW89a, RW89b, Ric67, RC66, Rul96, Saa92, Sch79, Sem92, Son98, Ste76, Sto68, SP94, Tav77, TW70, Tho68b, Var71a, Var77, Wah75, Wat74, Wei71, Whe73c, Whe73a, Wig66, de 78a, Asc78, Bal86, BM89, BX89, CQ82, Cha86, Dav80, DS80, FN80, Gus81b, GK83, Hay81a, Hay81b, HHS78, Hop82, Ise79, Jam78, KK81a]. approximations [Kar78, Kro81, LN82, Lar89, Las84, MP78b, Nel86, Osh84, Rua89, Sam82, Sam86, SB88, Smi81, Str84, Wil79]. April [Hou73]. Arbitrarily [Nie99, Sec65b]. Arbitrary [CM74, CH97, Gov93, Hua75, MRW96, OS65, Swe77]. arclength [Mit86]. Arguments [AH76]. arise [SL83b]. Arising [AWZ96, CH94, HR95, Kee70, KK91, Osb76, PS80, Wim77, MR83]. Arithmetic [CO87, Cor83, Gar78, Han65a, HS67, Jac75, LO90, Olv78, RSK76, AR81, Ale84]. Aronszajn [Bea87]. Arrow [Que89]. Arrow-Hurwicz [Que89]. articles [Ano73]. Artificial [DW99, Kar96, RT95]. Aspects [HP75, Hel76, Mou65, Mou68, Wuy74, Bar88, DOR88a, DOR88b]. Associated [BIC66, Gai74, Mae82, Pla98, SLW75, dK86, DPZ87, KN85, Nie79]. Assumptions [Heg92, Wan92]. asymmetric [Rua89]. Asymptotic [AFM90, BH71, BP67, COP98, DP92, DE72, DS89b, FHK96, FR78, Gus80, Hag86, Kla98, Kla99, Küt84, Luc82, MTL96, Olv65, OS65, PFX95, SP94, Zla65, de 86, AS84, Gab85, Lee85]. Asymptotic-Induced [Kla98]. Asymptotically [DMR92, GLV91, Mah82b]. Asymptotics [HMM98, Gün88]. Asymptotiques [DS89b]. Asynchronous [El 90]. Atmosphere [SW99]. Atoms [FD98]. attainable [Cha85b]. Atteia [JS68a]. attention [BO87]. attracting [KL86]. Attractors [Lor97, Yan96]. Augmented [CGT91]. Automatic [And83a, BSW95, EJ87, GL78b, Geo80]. Average [Ale87, Joe99, Wil69]. Averaged [Bre84]. Averages [ADH99]. Averaging [Vas99]. Avoidable [BS97a]. Avoiding [BPTZ92, PT91]. Axially [Par65]. Axisymmetric [BFO96, Hei96].

B [But65, Ost65, Dah80, PR98, de 77, de 88]. B-Spline [PR98, de 88]. B-Splines [Dah80, de 77]. B.G.K. [Iss96]. Back [BF75]. Backtracking [LT98]. Backward [AE97, AL91, Car72, Cas81b, Fre98, HT82, JNT90, Mas90, Rei99, Sei96, VK90, Hay81b]. Backward-Forward [AL91]. Backwards [CSH78, KG91]. Baines [Mil92]. Bairstow [Boy77]. Balanced [GL96a, MPS98]. Ball [Nic81, Rit90, Xu89]. Banach [Pal93, Yam98]. Band [Dem77, Kee70, Kee71, de 80]. Bandwidth [GPS76]. Bank [Wal99]. Barriers [CR90, SI83, JS87]. Based [Ain96, AE97, ADH99, AMSS97, BS93b, Bar99, BS87b, Bog77, Bör90, Bro69, BC72, Cap99, CNP93, CK98, DD94, DE98, DR82, Hed79, HRR94, IT91, Lin92, Liu93, Mao95, MX95, MF97, May87, NV93, Per97, Riv84, Wan92, Jer88, Jet82, LB96a, Sas78, Sch83c, SSB85, SS80, Tru86]. Bases [BS93b, Bey92, BJS73, Can94, SLW75, Sha69, GH83, LS73]. Basic [Hes98b, IPS91, Nea92, RS89]. Basis [FD98, Lee91, Liu94, PL87, de 88, APS87, APR83, Hak82]. Basset [MS83]. Bazley [Yam75]. BDF [CLM87, Fre98, HW83]. Be [Rip92, DGB82, Fer73]. Beam [Car72]. Beams [BK73]. Behavior [CT83, Cos79, DJ99, Dun80a, LSS94, MTL96, Kau86, Lar89, MS89]. BEM [Car96, CG97, MS99c]. Bending [CS99, Kir93, MTF99, PRS90]. Bending-Dominated [Kir93]. Bergman [Eis74, BW93, Hud68]. Bergman-Vekua [Eis74]. Bernstein [Ros67]. Bessel [Der71]. Best [Aub68, Bar73, BP74, BB71, Bor76, BH75, CW76, Cle64, DMP75, DS97, FGH74, GCB74, GV74, Han65b, Hen69, Hen70b, Jim96, KL76, Lam76, Lan77, Lip73, MS73a, Mat72, MS65c, Mur71, Rei74, RD79, Sch70, Sch74, Sch65, Sch66, Ser99, TW70, Wat74, Kro81]. better [BCL89a, BCL89b]. Between [Heb71, Rit90, IT80, KL82, LOR85, Naz79, PR86b, SW74a, Wol83]. beyond [DFK87]. Bezout [MSW95, SS96b]. BFGS [Naz79]. Bi [Jia86]. Bi-Infinite [Jia86]. Bicubic [BFB92, Bia93, CF85, Hal73, Sun96, BC94a, CF89]. bicubics [DHLR84]. Bidiagonal [DDLT91]. Bidiagonalization [Pai74]. Bidimensional [Ben92]. Bifurcation [AM98, Atk77, Dav97, JP96, JD86, Rhe78, Sch93, Sim75, Web85, Wer96, WV79, JP92, WS84]. Bifurcations [Gov93, GGK97, Hon93, Kuz99, CFQ83, GMS97]. Biharmonic [Bjø83, BD74, CC66, CC67, Ehr71, EG75, Fal78, Gup75, Han93a, Jeo94, McL74, Osw92, Smi68, Smi70, Bre89, May84, Mon87, Ros80, Zha96]. Bilinear [KT97, Pan97, Tal91]. Bilinear-Constant [Pan97]. Bilinears [CK98]. binary [Whi85]. Bioreactor [Ter98]. biorthogonalization [Saa82]. Biot [MTL96]. Birkhoff [AS69, CJMT74, DLR82, Fer74, Gla69a, Jet82, Jet87, LZ71, SP68]. Birthday [Bir83, MSW94]. Bisection [Riv84]. Bivariate [BW69, CJ75, CF90, Die90, Gor71, Man71, APS87, Mey87b]. Bivariational [RY86, RB79]. Black [Yav95]. Blast [Cho92]. Blending [Gor71, Wat77]. Blending-Function [Gor71]. Blob [Hou90]. Block [BS79, BB87, Cas82, CR86, CNP93, DD94, Ged81, Geo74, GHS97, HP75, Hel76, PSP89, PS80, PS82, PS85, SP88, She96, SW94b, Sun96, Swe77, BB88, Cha94, CM88a, Dem83, Fer86b, Saa80, Sha85b, WW88]. Block-Centered [DD94, WW88]. Block-Elimination [CR86]. Blocks [Cap99]. Blowup [Le 94, NE77]. Bocher [Mar66]. bodies [DPZ87]. Bogdanov [Gov93, Jan94]. Boltzmann [ABDH95, BI89, ELR95, Iss96, LC98, PSB97, Per90, Per92]. Bordered [Wer96].

Bound [BK98, But75, Heb71, Kam69, MSW95, PW96, STW97, Hat82]. Boundaries [HW93, Kar96, Rat95, Wah92, BG88, Len86, Rhe82]. Boundary [AP95a, AP73, AMT99, ABGM92, AD78, AL96b, BS96, BM96, BJR90, BR75, BR77, Ban77, BDH98, BHL96, BCV95, BCR75, BCM91, BL92, Bia98, BBG95, BNP91, BDLN92, BK77, BBS73, BH68, CW96, Cal74, Cal76, Can68, Can86, CS96, Cas88, Che73, Che75, CS90, CH94, CDW96, DM77, DW76a, Dia77, Doo66, Dor70, Dou93, DW99, DW76b, Dup76, Dut90, Eis74, EM88, Faz96, FHK96, FL71, FR78, FU69, Gai74, GGS94, GGS98, GK97, GW97, Gek75, Gek76, Gla69b, GG76, Gre91, GH92, Gup75, Her68, Hig84, Hig90, Hig94, DW78, HW97b, HW92, HZ94, JP67, Jen92, Jeo94, Jes78, JP98, Kac99, Kel74, KW75, KS97b, KS99, KT86, Kin72, Kin75, Kos73, KMR94, Kre91, Kus68, Lan77, LB98, LP77, LK80, LM90b, LP91, LMS87, Li91]. Boundary [Lin90, LR72, Man78, MP90, MW92, MG95, MJ77a, Mat82, MT75, McL74, Med98, MS99a, Met94, Mil69, Mir71, Mit77, Oli74, Ols77, Par99, Pol74, Red76, Reu92, Reu94, RSE69, Ros70a, RS75, Rus77, RC78, RT95, Sac71, SS95, SV96, Sch90a, Sch72, Sco75, SW77, Sha68, Sha69, Sho75, Sig76, Sin77, Smi73, Sta99, SWB90, SHR96, Tap69, TKJ94, Var71a, Var71b, Var77, VF93, Wah75, Wat79, Wei86, Whe73c, Whe74, Whe77, Wig66, WO96, WH97, Yam98, Yan94, Yse86, ZB96, ZWW94, Zla67, ddd83, dd86, dC90, dW79, Asc86, BM87, BvA94, BC94a, BR88, Bog85, Cha86, CK85b, Cle81, DO79, Doe79, DR84, Dor86, Dou84, DS89a, Dut88, FKD84, Fer86a, Gar88, Gol89, GMP93, GGT82, GLT87a, GLT87b, GGL79, GZ88, Gui87]. boundary [GHS96, Gup85a, Gup85b, Gus81b, Hag86, JB88, JS89, Ked81, KP79, KS87, KNB86, Lay83a, LOR85, Lin88, LP78, Lus79a, Lus79b, MW86, MN84, Mey81b, Nic87, OR86, Pas79, PR82, Rin84, RS87, Sch80, SS96a, Smi81, WL95, Wei89, Whi79, Wom89, Zho83]. Boundary-Layer [AP73]. Boundary-Value [DM77, MW92, MJ77a, Whi82c, Dor86, Whi79]. Bounded [FP91, Li83a, Rub64, YZ94, Yin95, Bas85]. Bounding [CK69, Han92, Jac75]. Bounds [Alf96, BR70, BI66, BJ93, BT69, CM87, Che73, Che75, CH72, CGT89, CGT91, CR79, CW78, DM98, Don73, Don70, DD96, Est95, Fie78, FHM67, Gar78, Gay81, Gay83, GT74b, Güs81a, HS68, Han65b, Hen70a, Hic96, HK91, HMR73, Hol78, Hug98, Joh69, JS69, JT71, Kut82, Küt84, LW81, LK74, Les91, Let71a, Let71b, Let74b, Lin91, Loc73, Luc74, MT75, Mil92, MP68, OS65, Pet92b, Pri76, Rei68a, Rie71, Ros70a, San94, Sec65a, Ste71c, Ste77, TT95, Wat77, WGS90, BC94a, CGT88, DKW82, DS84, Gar84, GV83, Hic97, JTW83, KPJ82, Ked81, Luc85a, RB79, Sch71b]. Boussinesq [Win82]. Box [BR87, Ker96, Str72, AD92, KL82, MN84]. Branch [VMK97]. Breaking [Jan94, CFQ83, WS84]. Brenier [LeV88]. Brent [Mar79a]. Brouwer [KLY76]. Brown [Mar79a]. Broyden [DK85, Gay79, GL81, Gri87]. Broyden-like [Gri87]. bubble [Pie95]. Bubbles [BHM^{+}99]. Buchanan [Jan83]. Burgers [E92, KK98b]. BVP [dM87]. BVPs [DOR88a, DOR88b, dK86].

C. [LE72]. Cahn [EF89]. Calculate [Pet92a]. Calculating [Bac80, GK65, Gre99, de 77]. Calculation [Dun74a, GCB74, GH92, JS85a, Lyn67, Ma95, Mid66, MT68a, NE77, OT75, TW70, Wat74, MS80]. Calculations [CS94, Kam69, JMP83]. Calculus [GW90, GL93, LVQ95, MW95, MW92, Sim69]. Can [Jam77, Rip92]. cannot [DGB82]. Canonical [IS90, SSA91, OS92]. Capacitance [Kam69, Dry83]. Capturing [LO96, Lev82]. Carathéodory [GT82, TG83]. Cardinal [CJ75, ST76, Lee85]. Carleman [Ahl66]. Carlo [CM92, LC98, MC90, MC93, Owe97, Ros67, Sas78, Spa71, YKS78]. Cartesian [CGY98, MDH^{+}99]. Case [Alo97, AL96a, BR77, Ban77, Bur90, HMM98, HMMR95, JV96, Rac68b, BO87, DS89a, ST95]. Cauchy [Atk72, BJ93, Bia89, CD67, CD77, CM88b, ED77, FR85b, FP95, Ger86, Ger89, IT80, Ioa81, Mie86, Mon86, RHH99, Ros81, Rul96, VMSB97]. Cavities [Laz82, Laz81]. Cavity [BK99]. cebysev [MS65c]. Celis [EA91]. Cell [AWY97, BS97b, KSS80, VMSB97, VA91]. Cell-Centered [AWY97]. Cell-Vertex [BS97b]. Center [Ma95, Shi66]. Centered [AWY97, DD94, Mad75, RR80, Rat80, KN85, WW88]. Centers [MS91]. Central [JLL^{+}98]. centre [JP92]. Certain [Aal72, Bak68, Buc77, CH72, Don94, Eis67a, GM69b, Gup75, HLY80, Hen69, Hil76, Hil78, Ise81b, Kee71, KMR94, MG72, MT68a, Sta64, Ste70, ST67, Tho76, Var76, Eva80, Han93b, Hat82, Her85, Sle82, Smi81]. CF [Tre83]. cg [PRS90]. Chain [BD99, Hav87]. Change [BGT97, BW90b, DW81, DW82, DW93, Ip87, WW90, DW85]. Changes [Fer74]. Changing [Epp82, Gut93, Kro73]. Characteristic [Abl72, Ber95, CM90b, DVW94, Hug98, LMS97, Ste75a]. Characteristic-Galerkin [DVW94]. Characteristics [AW95, Bro77, CK69, DL95, DR82, O’S90, SW91, DRW89, Dur88b, EYL89, Jer88, KHR89, Rus85]. Characteristics-Mixed [AW95]. Characterization [AS69, GV74, GR84, Lin77a, RR86, Ric67, Sch74, SW94a, Ste75b, Wil79, Her85]. Characterizations [Bor76]. cheater [LSY89]. Chebyshev [Asc77, BT99, BR70, Bas73b, Bas73a, BCM88, Bor76, Bra85, BQ86, BH75, Can86, Cha78a, Cha69, Chu72, Cle64, Coh71, DG94, Dun71, Dun80a, Dun80b, Dun80c, DS68, DT73, ES76, EW76, Fun87, Fun90, GTV90, Ged81, GGK98, GL83, Gün88, GL96b, GT82, Heb71, HZ90, JK68, Kam76, Kar78, KP96, KP97, Kim76, Ma98a, Ma98b, Mil66, Mou66, Mou68, Mur71, Ric67, RJ69, Rie71, Row67, Row69, Sch68b, ST95, Tad86, Tre83, Tse96, Wat75, Wil72, Wil79]. Chebyshevian [GKR70, KZ66]. Chemical [Fin71]. Chernoff [NPV93]. Choice [Pet87, SEK93, HL81, Neu88, Vog86]. Cholesky [Mei83]. Choose [Ser99]. Chorin [Wet97b]. Circle [BT99, BP80, DC80, GM75, Kah67, Let71c, ML92, Sch98b, Tse96]. Circulant [CJY91, CY92, CY93, Che87, Cha89]. Circular [BK99, Gar78, Lin92, Ban83, HB86]. Circularity [Tre83]. Class [BCO94, BGM96, BP96, Buc66, BS94, But81, Cho78, CGT89, DM92, Ebm98, Fai94, FF98, FF91, Gus70, HL84b, Hin72, DW74, IT91, JT95, KM96, Küt84, LVQ95, Lov96, Ma95, Nic72, Nic73b, Osb67, PP73b, Pri76, PFX95, Rau73, RS95, Rhe68, Rhe78, Sar90, Sim68, Tav77, Tre72, Voi71b, Wat79, Wet97a, Wid78, Xu92, Yam75, Yav95, dR81, AE84, Bun81, BNY87, CK85b, CM88a, Cle81, CGT88, DK85, Epp84, Fer86b, FST87, Gab85, GGL79, Jet82, Joh88, LSY87, MR83, Vil89]. Classes [Sch74]. Classical [NS93, Not91, Ell82, Ell84, IT80]. Classification [GM69a]. Clenshaw [LE72]. Cline [BCC78]. close [LOR85]. Closed [Atk72, For73, Gir95, SS93, SS95, SV96, Ste71c]. Closest [CR90]. Closure [LO90]. Clough [Per76]. Clusters [Ost70]. coalescent [Far85]. Codes [SZ90, Ste79a]. Codim [Kuz99]. Codimension [Gov93]. Coefficient [BR75, BR77, Ban77, EW96, SLW75, TW74, CS87, Dai92, GHY83]. Coefficients [AWY97, BCO94, BS97b, BGT97, BTW93, BDLN92, Bra94, Dej67, Día79, Eis66, EL97, For66a, Gai66, Hos95, JP67, KM96, LS72, LL94, Lyn67, MDH^{+}99, OPW65, Oet65, PK90, SSA84, She96, Vav96, WB98, AW83, Bal86, LL95, Sch89b, Sun79, Vin86]. coercive [Des81]. Coexistence [LGMMV92]. Cohn [Gar71]. Collapsing [Laz82, Laz81]. Collection [Ano73, Pre73]. collinear [Dav80, Sor80]. Collocation [Arc77, Arc78, AD78, AW83, Asc86, AB86, Bar99, BFB92, Bia93, Bia98, BF99, Bru90, CW96, CP91a, CW78, Dia77, Dou93, DW76b, Egg83, Fai94, Fun93, GGS98, GYF86, HKL77, DW78, HVR88, KP96, KP97, Koc84, LB98, LFB98, LSX90, LR72, PR95, Phi72, PR76, QZ92, Rin92, Rus77, RS97b, SV96, Scu98, Sin77, Sun96, SHR96, Sun99, SW74a, Swa88, Wei86, Whe77, Whe78, YF92, Zan99, dS73, AS84, AGS83, BC94a, BQ86, Bru83, Bru84, Doe79, DHLR84, Ell82, Ell84, Gar81, GL83, Hu96, Joe85, Kum88, Mie86, PR86a, PW80, Rin84, Zen85, Día79]. Collocation- [Día79]. Collocation-Finite [Whe78]. Collocation-Galerkin [Dia77, DW76b]. collocation-type [Kum88]. Color [KL89]. coloring [CM83]. Column [Var76, GN83a, GN83b]. Combination [Pfl97, War75]. Combinations [Li91]. Combinatorial [Prü82, GM78c]. Combined [Bro87, FFLMW99, GM93, HM85, PS83, Dur88b, KHR89, Li83b]. Combining [DR82, Szy88]. Comment [Næv91, de 78a]. Commentaries [CGO07]. Compact [AC83, Atk75, CS94, DF84, Fer74, HKT92, PR82, Wei74b, Gar88]. Compactly [Bey92, Can94]. Companion [Kam78]. Comparative [LS76, FR85b]. Comparing [HEFS72, HEFS74, JEH78]. Comparison [Bar70, Cle64, Fav64, LTW95a, May87, Mil76, NK97, NS93, RC66, Rus77, Str68, Wal73, Ban81, IT80, Ral80]. Comparisons [de 78a, HHS78]. Complement [Man90]. Complementarity [Les91]. Complete [Cas69, Gra71, SW94a, Sid97]. Completely [Atk73, Kam76, Pal96, SV94, Kam79a, Kam79b, Kam81]. Complex [BT99, Bar65, Bar69, Bor76, CY93, EW76, EW91, Gau70, IT95, Joh78, Kin74, Lor97, Nav65, Tse96, Wei94, Wil72, Bru80, CM88a, Mie81, Pet89, Rei88, Saa87, Tre83, Wei95, Wil79]. Complexity [Boj84, HMR73, Kun75, SS96b]. Complicated [DJ99]. components [Hig84]. componentwise [Pan85]. Composite [BH75, CM90a, DH71, ELV94, Ise84, LW98, Tur77, Cas81a, SDC88]. Composition [Rei96]. Compound [DL75, Köh95]. Compressible [EV96, KL96, KL97, Per92, San98, ZH91, ZH94, EYL89]. Computable [CW78, Gar84, MT75, SW85]. Computation [BK73, BIC66, BK97c, Bra66b, Can94, Car74, CGO07, CG93, DR90, DL95, DRV97, Dol99, DAB99, Drm98a, Fra78, FD91, GGM99, Gau70, Gov93, Gov97, Hon93, JPRT96, LGMMV92, LS73, Mal77, MR77, MS99b, Moo96, Ned78, Pat76, PA73, RSK76, Rhe82, Sch93, SP88, Spo94, Uso67a, Uso67b, War77, Wei94, Wei95, Wer96, Boj84, Bun81, Cha78b, Cuy87, GR84, Gut83, Hac79, Kun88, RR86, Ste79b, Vin86, WS84, Wil79]. Computational [Bau74, BFGP99, DOR88b, FM67, Gra87a, Gus70, JRB95, KLY76, Kun75, Mou65, Mou68, SW77, Moo78, Xu89]. Computations [BW76, Han65a, HS67, Mer66, Mül84, BR78a, Pai79]. Compute [Cha73]. Computed [DMW83, ST91]. Computer [HY96, JP92, SB77a, Sch83c]. Computer-aided [JP92]. computer-based [Sch83c]. Computers [PO87, Sch89b]. Computing [AGM92, Alo97, BT99, BG95, BD90, BS79, BB71, Col98, DMP75, Drm98b, Erd65, GBIY71, Gay83, GMS97, Gus73, HP68b, HP68a, Hav87, Hos95, Kar78, Kut74, Nob66, SS74, Tuc69, Wal96, vS87, BSW85, DK81, Hof85, Mar78]. Concept [FSU81]. Concerning [BP80, Chu72, Dan70]. Condition [BA76, BBG95, CMSW79, Dun71, Kre91, Kří92, Lin91, MTTW99, OC84, Ric66, dK86, de 88, BO88, Dem83, Dix83, GMP93, Mil79b, Osh84, Ste80, Wat81b]. Conditioned [Bog77, FH72, Pel67, Var73]. Conditioning [AMT99, LM90b, LM90a, Mat82, MS73b, PR86a, Swa88, dM87, BS89a, BB88, IT88]. Conditions [BM96, BJR90, BDH98, Bec92, Ben96, BCV95, BCM91, CDW96, Dou93, Fer73, GH92, HW97a, Hig90, Hig94, HW92, KT86, Kos73, Lin90, Luc74, MP90, MG95, Met94, Mur97, Red76, RT95, SSA91, Sco75, Sew66a, Var71b, Var77, Wil70, WO96, dC90, BCM88, Can86, DW85, Dut88, FM84, Far88, FP86, GMP93, Hag86, Hig84, Lus79a, Lus79b, SO97, Tad88]. Conduction [Mas90]. Cones [BP74, DMP75, JD86]. Configurations [Alo97, Mey87b]. Conformal [DeL94, Dri99, Nit89, Weg86, Rei88, Tru86]. Conforming [Cas97, Osw92, Sem92, BR85, Zha96]. Conic [Dav80]. Conical [BDLN92]. Conjecture [BP80, LE72]. Conjugacy [Buc76]. Conjugate [AMS90, Coh72, Dan67, Dan70, Eis83, GHK94, Gut93, KN72, Lue70, Man90, Pla98, Rei72, Rit75, FM84, For79, Gut83, JMP83, Naz79, Nic87, SSY88, Sha78, Ste83, Vat88, Win80]. conjugate-gradient-type [SSY88]. Conjugation [DRS95]. Connected [BS77]. Connecting [FD91]. connections [VSM80]. Conquer [ST91]. Conservation [BCV95, Ber87, BG98b, CN97, CM90b, Coc89, Coc90a, Coc90b, CCL95, CG96, CGY98, CF93, Cor96, GV91, GLBN97, Har90, JY98, JLL^{+}98, JX98, KR94b, LTW96, Lav91, LTW95a, LMS97, LW98, Liu93, Ma98a, Ma98b, MKT93, Mao95, NT92, NT94, Pet87, Şab97, Sch90b, STW97, SB93, TT99, TT95, Tan98, Ten92, Ten94, TW93, Vas99, Ber86, Bre84, BO88, CT86, HL81, Lav89, LeV88, Lev82, Lev85, Luc85b, MT89a, San85, Swe84, Tad89, Vil86]. Conservative [BC99, DD94, Dk99, LeV96, Mah82b, Ber86, SO86]. Considerations [Can83]. Considering [BS97a]. Consistency [BN76, PSB97, SJ77, WW93a, WW93b, You72b]. Consistent [Asc77, Cad73, Cad74, Eis67a, LPG91, You72a]. Consolidation [MTL96]. Constant [BR75, BR77, Bra94, Pan97, SLW75, hTZ97, Vér90b, AW83, Mar83]. Constants [GGM99, Gri86, Gün80, Hof76, Law67a, Let74a, SWW97]. constitutive [LB96a]. Constrained [BV92, Buc76, Byr90, CR90, Con73, CP77, EA91, Fon90, Hag75, Han65b, Hei98a, Jay96, JK68, Lue70, MS91, PT91, Pie69, Pie71, Pot93, Rei96, Rit75, Sto71, Wei92b, Yen93, Bar88, BT84, BPTZ92, BSS87, Dem87, Don81, FST87, Fon88, Gou89, KR88, NO85, Van85, Var85]. Constraint [AL96a, CKS95]. Constraints [BR78b, BM88, Bur90, CD90, CGT91, Eic93, Eld80, KP78, Kel69, Tap74a, Tap74b, BN83, Bra85, BRD81, Cor81, MP78b]. Constructed [CY92]. constructing [Mey87b, VH93]. Construction [BC89, CJ96, Doe78, EN83, Eis67a, HP77, JK68, Mil69, MP78a, Seg72, Str68, Wil81, BE88, Esp87, MP85]. Constructional [Pet65]. Constructive [DM88, GW83, KLY76, Sch72]. Consumption [BCR75]. Contact [CG97]. Contained [Fri71]. Containing [Rei68b]. Contains [Kah67]. Contaminant [BK98, DVW94, Daw98]. contamination [EYL89]. Context [NE77, GB89]. Continuation [Avi74, CM65, Dav97, FD91, Hen66, KX94, KK98a, Mit87, Say67, dG81, dR81, Car94, CHM88, Kar78, Mit86, Rhe80b]. Continued [Fie78, GW83, JS69, JT71, Mer66, ST67, Van81a, Wyn64, JTW83]. Continuity [DLR82, Kee90b, MW94, Str66, WW83, Car94, KW85]. Continuous [Atk73, BP96, CDMW98, Dia77, DRV94, DRV97, Enr93, EH98, JV96, JK68, KM99, Mey87a, PT97, San94, Sch70, Uso67a, VF93, AGS83, Bak80, Enr89, FR87a, GB89, LP78]. Continuous-Time [CDMW98, JV96]. Continuously [JW91]. Continuum [DN91]. Contour [GTV90, Bia89]. Contours [Bar65, ED77]. contractive [Lin83]. Contractivity [BJZ94, BT97, in 96, CH95]. Contrast [Ver90a]. Control [AGH97, Ben96, BJMT73, BD99, Enr93, Est95, Hag75, Hag76, Hag90, Hig89a, JNT90, JRB95, KX94, KT97, MTTW99, San98, SH95, UTK96, Don81, Enr89, EJ87, Joh88, SC87]. Controlling [WGS90]. Convection [CS98, DR82, EC93, FR92, FFLMW99, Jer88, LMV96, MG95, Ric91, CFQ83, EL89, Fis82, KHR89, SB84]. Convection-Diffusion [CS98, EC93, FR92, LMV96, MG95, Ric91, EL89, KHR89]. Convection-Dominated [DR82, Ric91, Jer88]. Convective [ELR95]. Convective-Diffusive [ELR95]. Convenient [Tad88]. Convergence [AH74, And83b, AV99, Atk68, Atk75, AC98, Bab71, BS87a, BP96, BG95, Bar65, Bar69, BHL96, BCV95, Bia98, Bol72, BSTW77, BKP94, BC95b, Bre75, BDGB83, Bro87, Bub78, Bub85, Byr90, CL89, Cal74, Cal76, Car91, Cha78a, CHL91, CM99a, CfXmZ99, CS72, CS74a, CS75, Cha73, Che88, CE97b, Che98, CCL95, Coh72, Col98, CGT89, Coo78, Cop90, CF93, CGH91, Dan70, Dav94, De 73, DK82, DKK83, Dem77, Den74a, DW81, DW82, DV97, DD93, DNW98, Dun74b, Dur88a, E92, EH75, Eic93, Eis66, Eis74, EA91, ELR95, EL93, EL97, ES98, EH98, EMB71, FF98, FX95, Fie78, Fin71, FL71, FSU81, Fra87, Fre90, Gar93, Gay79, GW74a, Gek75, Gir95, GS91a, GJL99, GC96, GLT87a, Gri86, Hag76, Hal79, Hal87a, Hed68, Hoh99, HM76b, Hou90, HLK91]. Convergence [HW92, HW93, HVR88, Iss96, IR99, JK96, Jam73, JR75, JD86, Jer73b, Jer80, Joh76, JP83, KN72, KR72, KK98a, Ker71, KKP94, KK91, KR94b, LTW96, Las84, LQ84, Les91, LB90, LB97, LTW95a, Löt92, Mad71, MN87, Mar79b, Mat98, McC94, MS98, Mey76, Mil66, MS94, Mor71, Mor97, Mor95, Nel73, Nel86, NT92, NT94, NTT94, Neu98, Nic73a, NR72, Nic92a, NV97, O’S90, OR72, Ort73, Osa90, Osh85, OR93, Ott96a, Per67a, Pet91, Pfl97, PV72, PP73b, Por71, PH73, Ral74, Rau73, RS95, Reu92, Reu94, Rhe68, RV74, Rit90, RW96, Şab97, Sam82, Sch98a, Sch73, SZ99, Sch90b, Sch96, Sha66, SZ90, SFS86, Sid86, Sid97, SF79, Ste76, Ste75b, Ste74, Sül91, ST67, Tad89, TE98, Tan98, Ten92, TYZ95, hTZ97, TW74, Tho76]. Convergence [Urb91, Van72, Vas99, Vat88, Vic80, VTG91, VA91, Voi71a, Voi71b, Wal67, Wan92, Wan93a, Wan93b, Wei93, WL95, Wig66, Wim77, Wyn66, YB96, YZ94, Yin95, Ypm84, ZZ98, de 78b, Ale84, AVG89, AS84, BI89, BD85, BBT82, Bea86, BT89, BH83, Bub86, BNY87, CC84, CGT88, CM88b, CL84, DH79, Dob80, Dun80b, Ell84, Far88, Fer86b, FST87, Fon88, For79, GV89, Gou89, Gri87, Gus81b, HLN83, Hof78, HP78b, IT88, Jac88, Jam80, Ker86, LM88, Li83b, LB96a, MM84, Mil79b, ME81, Moo78, Mur82, Neu88, Pan85, Que89, Rau85, Saa80, Sha78, SW85, SSB85, SL83b, Sor80, Toi79, Var85, WW88, Wig87, Win80]. Convergent [AP83, Bro69, CGT91, DM95, Dör96, FMOS96, KC74, Mad75, MP77, Nic81, Osa90, Rey72, Wat79, ZH94, CL92, DGB82, EHP87, Kun88, Sch86, Sty89]. Converging [CP77, Laz82, Pol74, Laz81]. Converse [Neu97]. Convex [Afr70, BC94b, DMP75, Dol99, GG90, Gir95, GW74b, HLLnWJM98, Kel66, MS91, PR77, Sch68a, Van67, Wil74, Bea82a, BO88, BNY87, Grz85, KRD78, Pet92b]. Convexity [Bar73, Cle90]. Convexity-Preserving [Cle90]. Coordinate [Eic93]. Coordinates [ZW94, Rom90]. Copositive [HY96]. Copying [DS92]. core [NW89]. Corner [BJR90, Bre98, Fra78, Wig87]. Corners [DRS95, Don69, Lee92]. Corrected [KR97]. correcting [GM89]. Correction [AC83, BV92, BPR72, Cas88, CGT89, Dan70, DL73, SA73, Bar88, Dun80c, EL89, Kau86, Pas79]. Corrections [BF75, Cas88, DM77, KP79, Ske82, Ske86]. Corrector [DH71, Hal74, Jac86, Lam67, Lam71, Zen85]. corresponding [Gün88]. Corrigenda [Gup85b, MR85]. Corrigendum [Jel79]. cosine [Hop82]. Cost [BD99, Gra87a]. Cotes [AF68, RM77]. Counterexample [Wah92]. Counting [Sch83a]. Coupled [Ben96, Ehr71, IIMPL82, LL96a, McL74, Pao99, Smi68, Smi70, Cha85a]. Coupling [BM99, Car96, CG97, CS90, GW97, Med98, MS99c, Gui87]. Covariance [Wil69]. Covolume [CK98, CKV98, DNW98, NW97]. Crank [BF99, Jam80, KW76, MH98]. Criteria [BQ76, JR75, RM77, Szy88, BB79]. Criterion [Gar71, JEH78, Jan83, NR72]. Critical [Gol77, DFK87, Rhe82]. Criticality [Mil81a, DL82]. Cross [SV70]. Cross-Positive [SV70]. Crystal [Alo97, LL89]. Crystalline [Gir95]. Crystals [DG98a]. CS [Pai84]. Cubature [Cha68, CJ75, HP77, Let70, Let71a, Let71b, Let74a, LS97b, Pri76, Str73, Tay95, Wil74, WB86, MP78a, MP85]. Cubatures [Bar69, Fra73]. cube [Duf82]. Cubic [Arc77, Atk68, BDSW66, Cle90, CM99b, EJL85, FC80, Hal73, HVR88, Huy93, Joh69, KR72, Ker71, LL98, Luc74, MST88, Per76, Pow69, Sun99, Bea86]. Cubics [BW89, HS97, Hed79, PS92]. Curl [Nic92b, NW97]. Currents [Ned78, NW89]. Curvature [BG95, CDK95, DP98, Gir95, NV97, Wal96]. Curvature-Driven [NV97]. Curve [Dzi99, Gir95, Row67, Bea82a, Tre83]. Curved [Gue92, Zla73, Zla74, BG88, Ber89, EHP87, Len86]. Curves [Atk72, DRS95, For73, Moo96, SS93, SS95, SV96, Tra96, GU88, Hof85, Wix78]. curvilinear [Rom90]. Cusps [Lee92].

Cuthill [LS76]. Cuts [LB98]. Cutting [DM95]. Cycle [Bra94, BH83, McC85, Neu98, BKP94]. Cyclic [DF76b, DH71, Hel76, McK79, Sch89b, Swe74, Swe77]. Cylindrical [Asa98, PCW95].

D [de 78a, Gui87]. D. [Wuy70]. DAEs [AL96a, BM96, AC98]. Darboux [Gen84]. Data [Bre77, BS77, BT68, Can68, CK69, HS97, Har96, HT82, Jup78, LS97a, Lan64a, LT89a, SS89, Sec65b, Wah77, BGO89, CR79, Die82, GHS96, Jet82, MM79, Rag85, Sam83, Vog86]. Davis [Cha69]. Deblurring [Car99]. Decomposition [Ain96, BGT97, Ben95, Ben96, Bia93, Cha87, CHL91, DD94, DDLT91, Dri99, Drm98a, Drm98b, ELPV93, GK97, GOQ97, Hei93a, HS99, Van76, Man90, MSW95, OPF97, PS81, TE98, TMV98, Cha84a, FQZ88, Scr91, Wan93b]. Decompositions [AMSS97, ASW93, Pai84]. Deconvolution [Lin92]. Decoupling [MH98, Ker88]. dedicated [Ano73]. Dedication [Isa86, MSW94]. Defect [AC83, Enr93, Hig89a, EL89]. defect-correction [EL89]. Deferred [Bak71, BV92, Cas88, DM77, Bar88, KP79, Ske82, Ske86]. Deficiency [wECGS94, RR86]. Deficient [DS97, LW92, Wei92b, LSY87]. Defined [OPW65, dC90, AS85, AG87, GU88]. Defining [GGK97, GR96]. definite [Chr91, GHY83]. Definiteness [Van72]. Deflated [Cha84a, CR86]. Deflation [For78, Nic87]. Deflations [LT68]. Deformation [Kir93]. Deforming [JW91]. Degeneracies [Jan94]. Degenerate [AWZ96, CH68, Ebm98, FP95, Hil68, Jam77, Kus68, LB96b, NV88, RW96, Fre87, Hof85]. Degenerating [CT83]. Degradations [CSH78]. Degree [Die90, DS97, GM78a, Gun74, HP77, HY96, Kin74, LS97a, OT75, Str67b, Str69a, SCWM71, Str72, ASW93, BE88, GL80b, Man82]. Degrees [WB86]. Delay [AP95b, BP96, EH98, ITM91, TK96, de 86, FN84, WR85]. Delay-Differential [ITM91]. Delay-Differential-Algebraic [AP95b]. Delays [TK96]. d’Éléments [Cia75]. Delves [Li83a]. Delves/Lyness [Li83a]. Dennis [EA91]. Denoising [DV97]. dense [Hor83]. Densities [Ale87]. Density [BLMPT92]. Dependence [HV70, Kre91, Pow69, Uso67a]. Dependent [ARW95, CS98, DR90, Den75, Den77b, Día79, GP98, GW90, GL93, Lin97, van84, AF86, Bal86, BR85, FN84, Hag86, Tav78]. Depending [Chu90]. Derivation [BC73]. Derivative [Bro77, Cas81b, Enr74, MS65c, Mou65, SD77, Wim70]. Derivative-Free [Wim70]. Derivatives [Don94, GM97, Hau71, JW91, KL76, MS66, MS88, Rei68b, Sch69f, Woz74b, van79, CM88b, Hu96, SB88, Swa81]. Derived [GGM99, Wyn68, SS83]. Descent [Cli76, Nas70, WM67, BT84, CG84]. Description [Kar78]. Detection [ALM92, CLMC92]. Determinant [Hel74]. Determination [Can68, Joh78, Mel87, Sim75, Sin73]. Determined [Cli76]. Determining [DT73, Sle68, Wat81a]. Deterministic [BD99, HV98]. Deux [Cia75]. Development [Bro95, Cas88]. Developments [Wyn64]. Développements [DS89b]. Deviating [AH76, Mee80]. Deviation [Row69, Sch65, Sch66, Kro81]. Deviator [Mot66]. Device [CC94, Asc89, Ker86, MS89, RS89]. Devices [ZWW94]. Diagonal [AMT99, Bea76, Bun71, Jam73, KSC99, Mor72, WS93, Bun81, Sha85b]. Diagonalization [BS79, GHS97]. diagonalize [Dem83]. Diagonally [Ale77, BD90, BT97, BJ97, in 96, vS89a, BC89]. diagrams [JP92]. Diamond [AS80]. Dichotomic [EM88]. Dichotomies [de 80]. Dichotomy [LM90a, dM87]. Difference [Abl72, AS80, ADZ96, Akr99, BH65, Bra66a, Bre75, Bre77, Bro87, BC73, Cal76, CH68, Cho78, Cim72, CJ96, CF93, Cry67, DWW98, DW76a, Dej67, Doe78, DR82, DL91, Dun97, DKR68, Dup68, Ehr71, EG75, Eis66, Eis67a, ES98, ELV94, FMOS96, FT74, Gai74, Gar66, Gek75, Gir74, Gol98a, Gol98b, Got72, GM66, GM69a, GM69b, Gun65, Gus69, GK83, HL84b, HL98, Heb71, Hed68, Her68, HMR73, HW92, Hud68, Hur67, HS99, IR99, Jam71, JY98, Kee70, KG91, Kel69, Kel74, KW75, KP79, Kel66, KP97, Kos73, Kre78a, Kut70, Kut74, KC93, LP77, Li87, Li88, LVQ95, Mar90a, Lor97, Mad75, MW92, McA66, MM73, Mil69, Mir71, MG82, Moo99, MS65b, MDH^{+}99, OR66b, Ort73, PS80, Rei68b, Rey72, Ros73, Şab97, Say74, Sch98a, Sch96]. Difference [STW97, She96, Shr73, Sim71, Spi71, Str65a, Str68, SWB90, SHR96, SW74b, Tav77, Tho67, Tho68a, Tho68b, TW93, Var71a, Var71b, Var77, Var80, Wal73, Wen68, Wet97a, Woo71, Wu96b, ZH91, Zem95, ZH94, dd86, AO82, AM89, Ber86, BS83, Cas84, Cha84b, CS87, CK85b, Dai92, Doe79, DR84, Dou83, Dut88, EHP87, Far88, FR85b, GGT82, Gus81b, HL81, HK79, Hof78, JS87, LN82, Lor84, Nel86, NV83, Osh81, Osh84, PS82, Par82, PR82, PR86b, Sle82, Str84, Tad88, VG86, WR85]. differenced [AVG89]. Differences [AWY97, DD94, Der71, For90, Mae82, Nea92, Rus77, Smi68, Smi70, Zla65, Wei89, WW88]. Differencing [BJ91, Pit78, Fun87, Tad86]. different [Boj84, GG86, Gus80, GK83, Kre79, Nie79]. Differentiability [LM88]. Differentiable [BCS78, Rob76, Ver93]. Differential [AC90, Ale94, AH76, AP91, ARW95, AP95b, BGM96, BP96, BPR72, Ben96, BB74, BP89, BC73, BP90, But81, BJ97, CLMM94, CMM97c, Cal71, CS74a, CS75, Cho78, Coo78, Coo81, CP91a, CT72, Cry74, Dav97, DO92, Den74b, Die92, Doe78, DL73, EH89, Enr74, EH76, EH98, Est95, Ewi75, Fai94, FS74, For65, For66a, FT74, FR68, Gai74, Gea65, Gea78, GP84, Gea90, Gla69b, Gus69, HKL77, HMM98, Hen69, HVR88, HS92, HRR94, HR97, HEFS72, HEFS74, IS90, ITM91, IT91, Jac81, Jac84, Jac86, JT95, JK96, JP67, Jam71, Kam81, Kat69, Kee90b, Kee70, KG91, Kel69, KW75, Kem75, KP96, Kno92, KS92a, KM96, Lam67, Lam81, Lee91, LPG91, LM90b, LM90a, Ler75, Lin72, Lin77a, LW70, Lin74, Liu99, LT67, LNS95, McA66]. Differential [Mid66, MT97, Mor95, Nan85, Olv65, OS65, OR93, Ott96b, Pan98, Pet98a, Pet98b, Pet65, Pet86, Pfl97, PR76, Pru73, Rac73, Rei81, SD77, SM96, SS88, SS99, SLW75, Sha68, Sha69, SBBL91, Sto68, Sto73, Tal91, Tav71, Tho68b, TK96, UTK96, VP93, Vel92, Ver70, VIA96, VWI97, Wat81a, Wei72a, Wer75, Whe73a, WB98, Yam75, Zed66, dC90, de 86, van79, van84, AF86, And79, AS84, Asc78, APR83, Asc89, BMS79, BR80, BH93, Cha94, CD88, DNT83, Dun80c, ED82, Ewi78, FN84, GS88, GM80, Joh88, KK81a, Kau86, KP85, Kre78a, Kre79, LM88, LM78, LP80, Maj84, Maj85, McK79, Mey87a, MG82, Pet81, RK88, Roc89, Rüm82, Sch86, ST86, SWC89, Sun79, Tav78, WR85, van81b]. Differential-Algebraic [AC90, AP91, BGM96, HMM98, JK96, KG91, KM96, LPG91, LM90b, Pet98b, BH93]. Differential/Algebraic [BP89, BP90, GP84, Kno92, Pet86]. Differentiated [Tra65]. Differentiation [Cas81b, Cul71, DGR96, Fre98, GP73, HC86, LM67, OR66a, Riv75, RS97b, Sec65a, Sne96, Wel97, WT88]. difficult [Kau86]. Diffraction [BGJ88]. Diffusion [AB99, ALM92, AM94, ABDH95, Bec92, BCR75, Ber95, BHM^{+}99, Bro95, CM90a, CMM91, CSH78, CLMC92, CDK95, CC94, Che95, CW97, CS98, Daw93, DR82, EC93, FR92, FFLMW99, Fun93, JK91, KHK93, LMV96, MMD94, MG95, Mat98, MS98, Ric91, TV96, VS97, BMP89, CEv87, CL90, Cha84b, EL89, Fis82, Hof78, Jer80, Jer84, Jer88, KHR89, SB84]. Diffusion-Absorption [MMD94]. diffusion-convection [Fis82]. Diffusions [KPH95]. Diffusive [ELR95, GJL99, JPT98, Kla98, MR98]. Diffusivity [CM99a]. Digamma [Spo94]. Digital [CSH78, Jen70]. Digits [Jen70]. dilations [DKW82]. Dimension [Alf96, CM90b, Daw91, Die90, DD96, Gar93, Mao95, Moo94, Sch98a, Swe77, Whi82c, ASW93, Gar84, GO81]. Dimensional [AR77, AV99, Asa89, BR81, Bec92, BM99, Bel97, BYLP98, BL92, Bof97, BC95b, Bro95, BT68, Cal76, CD98, Che95, CR90, CGH91, DDS80, Dar91, DW99, Dub90, EC93, GGS94, GGS98, GM69b, GM91, GM93, GL96b, HP78a, HMMR95, IK93, JB88, Jen92, Lav91, LB97, Ma95, Mad71, MMD94, MF90, NW97, PS80, Per97, Pet92a, RR96, Rhe78, San98, SB93, Ste71a, Urb91, VS97, Wah75, Whe73b, Whe77, WO96, YA95, Yin95, ZH91, ZH94, van95, AF86, AG87, Asa86, Bak84, BE88, BR88, BMP89, CR84, EF89, JP83, Lav89, LT88, LB96a, MS89, Mos83, NV79, Rom90, ST95, Vic80, Zho83]. Dimensions [Abl72, AMW99, BH71, BR75, BEK96, CF93, DSW94, GOQ97, GM69a, Gun74, GC98, KT86, KR94b, LMS97, MR77, Moo99, MS97b, PW96, Per92, RS92, RW96, Ten94, Wah74, Atk82, BCMP91, KD83, Lin85]. diminishing [Tad88]. Dirac [RS87]. Direct [BCC78, BFB92, BJS73, BP71, BGN70, BDGG71, BD74, CG73, Kut74, Mar90b, Mil76, Nic92b, SS73, Swa74, Tew67, ES86, IT80, JA83, KW84, Wol83]. Direction [BF99, CJZ83, CP91a, CR97, DF75, DF76a, Den77b, Den77a, DGP66, Dry78, EZB83, EW91, FF91, GM69b, Hub65, Rac68a, Sta91, BEL89, Hay81a, KHR89]. Directions [PP73b, Rit75, CG84]. Directly [CC78, CP77]. Dirichlet [Ban83, Bjø83, BNP91, BHZ68, BN73, BS77, Cos71, Cur66, Han93a, Jam69, Kut79, Li87, Lin85, Ros73, Swa72]. Disaggregation [CM84, Hav87]. Disc [Bru91]. Discontinuities [CS74a, Don94]. Discontinuous [Ans67, BGT97, Ben92, CGS92, CS98, LL94, MB97, Pet91, She96, Tad91, WB98, Arn82, Bub85, Jam78, LL95]. Discrepancy [Joe99, Qn99, WF97]. Discrete [AH77, Asa89, Asc78, BR73, BR78b, BYLP98, BHZ68, BT69, Bra85, BGS98, BDGG71, CE97a, CDMCW99, Dub90, Dur77, Dzi99, ES93, EC93, Esp94, GJL99, GW90, GL93, GM97, Hag76, Hei94, Hop82, HY95, ITM91, JY98, JPT98, Joe85, KT92, LFB98, Mad71, Mad75, Mah82a, Mah82b, MS73a, Moo94, NP81, Nel73, NPV93, NV97, PS77, PSB97, Pal96, Pit78, Rac73, RHH99, SS95, SS73, Swa74, Tar96, Tay95, Tho68a, Tre72, Uso67b, Var66, Vel92, Weg66, Wet97b, Wig66, Yan99, YB96, YZ94, AVG89, Asa86, Bal86, BO88, Don81, Gar88, Her85, Jer84, JP83, Kro81, Kum88, LN82, LeV88, LP78, NV79, NV83, PR86a, Rag85, Sam83, Swa79, Vic80, VG86]. Discrete-Ordinate [GJL99, LN82]. Discrete-Ordinates [Nel73, AVG89, NV83, Vic80, VG86]. Discrete-Time [CDMCW99, LFB98, Rac73]. Discrete-Velocity [JPT98, PSB97]. Discretisation [IPS91]. Discretisations [RW89a]. Discretization [AK90, AL96b, BF85, Bel97, BK99, BH68, Cas97, Cha78a, Gar93, Gup75, HR88, HR90, HMMR95, Hud68, JNT90, KM96, Nic92b, NS93, OR66a, Per67a, Ste65a, Ste71b, Uso67b, Xu96, Zla67, Dun80b, FR83, Hay81b, HR82, ST86, Wei89]. Discretizations [AR95, BG74, HW97b, Kee90a, Nie99, Ott96a, Yan96, ABPR86, Bru84, Gar88, HP80, KL86, Lus86, Sch86]. Discretized [ES76, Sch98b, VTG91, dK86, CD88, IKP91, Zha96]. disjoint [Saa83]. Disk [ST92, SS73, Sch83a]. Dispersive [Hig94]. Displacement [Bre93, Dou85, Dur88b, EW80, ER82, Rus85, Sam86]. Displacements [Den67a]. Dissection [DER76, Geo73, Geo77, GPV78, LRT79, GL78b, Geo80]. Dissipative [BJZ94, Bre75, HS94, IK93]. Dissipativity [Hil97]. Distance [Gov86, Rit90]. Distributed [GLV91]. Distribution [Hav87, Ura68, Bra80, GMP93, MP85]. Distributional [ZT67]. distributions [CLNP89, RS87]. Div [Cha92, Nic92b, NW97]. Div-Curl [Nic92b, NW97]. Divergence [Dub90, GH83, van95, BN83]. Divergence-Free [Dub90, GH83]. Divide [ST91]. Divided [Mae82, Nea92]. Dixon [Naz86]. Documented [Car99]. Does [AGH97, Heg92, Les91]. Domain [Ain96, Asa98, BGT97, Ben95, Ben96, Bia93, BW90a, Bör90, Cha87, CHL91, DD94, Dri99, ELPV93, GK97, Hei93a, Hub65, Li83a, Man90, OPF97, TMV98, ADK82, FQZ88, Jam78, LT89b, Scr91]. Domains [BS96, BNP91, BDLN92, Dub90, Gol93, GW74b, Gue92, Hei96, YZ94, Yin95, BG88, Ban83, Ber89, Cha86, EHP87, Len86, Wil81]. Dominance [Bea76, Jam73, Mor72]. Dominant [BD90, Bia89]. Dominate [CV99]. Dominated [AW95, CS99, DR82, Kir93, Ric91, Jer88]. Double [Bas73b, CKL91, Kru98, Rat95]. Double-Layer [Rat95]. Doubly [BS77, Hof85]. doubly-degenerate [Hof85]. Douglas [Jer88]. Douglas-Russell [Jer88]. Downstream [EK99]. Drift [CC94, Che95, JK91, BMP89]. Drift-Diffusion [CC94, Che95, BMP89]. Driven [BK99, NV97]. Droite [DS89b]. dual [BCL89a, BCL89b]. Dufort [Wu96a]. d’un [Cia75]. Dynamic [DM79, PS77, VP93]. Dynamical [wECGS94, DJ99, DRV97, FX95, HS94, SH95, KL86]. Dynamics [BS89b, BD99, CP98, Ein88, ES93, JRB95, KC93, MF97, Per90, Pet87, Yen93, AM89, Mun94].

Eddy [Ned78, NW89]. Edge [ALM92, BFGP99, CV99, CLMC92, MTTW99]. Edges [AH94, CH97, Hei96]. Effect [BS97a, BD76, GT74a, PT91, Pat76, Ske81, BPTZ92, GS88]. Effects [BH68, EC93, Uso67b, Woo71]. Efficiency [BDR92, Dus95, Enr93, KT76, RR87, Sin77, SW74b]. Efficient [AGP94, BCMP91, BR78b, Bic77, Bre73, BC90, Cad74, CC78, CG73, DFK87, Dra94, ER82, Gau70, IS90, LW70, LM86, RR96, ST78, SS89, Wat81b, Ell83, LSY89, MW86, Pet81, Ste80, Tru86]. Efficiently [Bur82]. Eigenelements [AC83, AT86, Cha73, Cha78b]. Eigenfunctions [NE77, PV72, Sle68, Hac79]. Eigenpair [GHS97, Zha95]. Eigenproblem [PW70]. Eigenproblems [Jen72, McC94]. Eigensystems [BD90, Erd65, KPJ82]. Eigenvalue [Asa89, Ban82, Bel69, BBG95, BH68, CMM97b, CfXmZ99, CM84, CC96b, wECGS94, Cra76, Cra78, Der71, Esp94, FH72, HP68b, HP68a, Joh69, Kau74, KS94, Kut72, LZC92, MS73d, Pel67, PV72, Ruh73, Sim71, Slo76, Ste72, Van75, BO87, BGO89, Chu87, DNT83, DL82, FNO87, Mit86, Mül84, Neu85, PR86b, SW82, Sco81, TW88]. Eigenvalues [Atk75, Bak71, BK73, Chu90, CH72, DMW83, Don69, Don70, FHM67, Kut70, Kut74, Kut79, Kut82, MP68, Nic67, Osb67, Osb76, Pru73, Sle68, WV79, BO87, DK81, Dix83, Hac79, Par82, WT88]. Eigenvector [HP68b, HP68a, Min70, BO87]. Eigenvectors [DK70, DMW83, MS88, MP68, ST91]. Elastic [DP98, FX95, Hig90]. Elasticity [Bre93, CMM97a, CMMP98, FM90a, FF93, GGS94, Lee98, Zha97, FS91]. Elasto [LB97]. Elasto-Plasticity [LB97]. elastodynamics [SDC88]. Elastoplastic [CS90]. Elastoplasticity [HR95, HDS97, LB96a]. Electromagnetic [BFGP99, CH94, JPRT96, Lee93, Urb91]. Electrostatics [Hes98a]. Element [AP95a, AMW99, Ain96, AMT99, ABGM92, AW95, AWZ96, AE94, Asa98, AK98, AL78, AW80, Bab71, BZ73, BK75, BA76, BR81, BSK81, BS87a, BS92, BCO94, Bai91, Bak76, BD76, BS93a, Bao95, BL94, BK97a, BK97b, Bel97, Ben92, BDM^{+}95, BLMPT92, BK99, BFO96, Boc97, Bof97, BFGP99, Boi97, BW90a, BNP91, BDLN92, BV90, BV96, Bre92, CM90a, CMM91, CDS99, CG98, CS96, CV99, Cas97, Cha92, CN97, CS99, CC94, Che95, CE97a, CE97b, CDMW98, CDMCW99, Cho92, Cia75, CG96, CS90, CM99b, CDW96, DG98a, Daw98, Des72, DM98, Dob92, Dör99b, DR82, DG90, DGW91, EJ91, EJ95c, EJL98, ELW91, ELW99, Fal78, FM90a, FR92, FR99, FF93, FHK96, FFLMW99, FL71, FR78, Fre90, Geo73, GM78a, Gir78, Gol94, GH92, GM91, GM93]. Element [GC98, HM76a, HR95, HW98, Hei96, Hei94, HR86, HR88, HR90, HMR73, HW97b, IB97, JV96, JK91, JN98, JNV99, JS98, JW91, Jim96, Joh77, JPRT96, KJ98, KM99, KL96, KL97, Kir93, LSS94, Li91, LB97, LL98, LTW91, LB96b, Lov96, LL99, LM92, Man78, Man80, MR98, MW94, Mit77, Mon92, MF90, Moo94, MS97b, MTL96, NP81, Ned78, Nie99, Nit89, OR76, OPF97, Osw92, Pal96, PP96, Pan98, Par95, PW78, Pet99, QZ92, Rei81, Ric91, San98, San94, SSW96, Sco75, Sem92, Sim68, Sta99, SW74b, Tha78, TXZ89, TV96, TH98, Tra77, VW95, Ver88, Wah92, Wat86, Whe77, Whe78, Woh99, WH97, Yse86, Zla73, Zla74, AF86, Arn82, AF89, BR78a, BO83, BG88, BCMP91, Ban81, BS89a, BR85, Ber89, BX89, CL90]. element [DO79, DL82, Dob80, Dor84, Dor86, Dry83, EF89, Epp84, Eri86, EJ95a, EJ95b, EYL89, FR87a, FN80, FR85b, Fre87, Gat88, Gen84, GL78a, GL78b, Geo80, GGT82, Gui87, GHS96, GH83, HR82, JTW83, KW85, KHR89, LL83, Lar89, LT88, LNOP82, LB96a, LP80, Mel87, Mil79a, Mil79b, Mon87, PS85, PW80, Pie95, Rau85, Rua89, RS87, SDC88, Sch80, SE81, SS96a, SO86, Sun79, SB84, Wah80, Win82, Zho83]. Elements [AWY97, BZ73, BDR92, BTW93, BM99, BG98a, Dur88a, DMR92, Eri80, GW97, HP78a, Jam77, Kří92, LB97, MST88, MX95, Med98, MM81, Mil81b, Mil92, Mil97, Nic72, Nic73b, PW96, PCL94, Per76, Sch89a, SS97, Vav96, Zha97, Zla73, Zla74, Arn82, BN83, BN85, Len86, MN87, Zha96]. Elimination [Bro69, Bun74a, CR86, Cli73, Geo74, Geo76, Var76, AH82, Zla80]. Ellipti [BB74]. Elliptic [AC96a, AH94, AWY97, BCO94, BS96, BR77, Ban77, BM99, Ben96, Bia98, Bör90, BEK96, BDLN92, BBS73, CDS99, CH68, CT83, CD98, CG90, CKV98, CG73, CP91a, Dör99b, Dör99a, DM79, DSW94, DKR68, Dup68, Eis74, EL97, Esp94, Fie78, FL71, FHM67, FR68, GK97, GOQ97, Gol91, Gol93, Gun65, Hil68, HW97b, HVR88, Jaf92, JP67, Jam69, JS98, Kel66, KP96, KT92, Kus68, Kut70, LL94, Li91, Li98, LB96b, Mal98, MJ77a, McA66, McC94, MDH^{+}99, NV93, Nic67, Nie99, Par95, PS80, PR95, Par99, PCL94, PW78, Pfl97, PR76, QZ92, Reu94, Ros70a, SSA84, SS93, Say74, Say67, Sch72, She96, Shr73, Sto73, SWB90, Swa74, Tar96, TYZ95, Tho68a, Tho76, Tra96, Tra77, Wan93a, WES80, Whe78, Woo71, Yav95, Yse86, Zla65]. Elliptic [Zla67, de 68, Ban81, BC94a, BW86, Bog85, BE79, BS83, Dor86, Dou84, DS89a, Dry83, Dyk87, ES86, Fre87, Hac79, HL87, KW84, KS87, KM88, LL95, LMS87, Mey81b, MG82, PS82, Par82, PS85, Pas79, PW80, RK88, SE81, Sch86, Ser83, Wan93b, WW88]. elliptic-hyperbolic [Ser83]. Elliptic-Parabolic [CH68, FR68, Hil68, Tho76]. elliptical [Ged78]. Ellipticity [Nie99]. ELR [DK81]. Embedded [Nak91, BE88]. employing [JA83]. Enclosing [RR98, FM90b]. enclosure [Kra87]. End [Luc74, Ske86].

Endpoint [Köh95]. Endpoints [Coh71]. Energy [CP98]. Enhanced [RS95]. Enhancement [OR90]. ENO [Son98]. Enthalpy [Whi82a, Whi82b]. Entropies [HLLnWJM98]. Entropy [BC99, Dk99, EL93, Lor64, OC84, Per90, Yan99, Osh84, WW93a, WW93b]. Envelope [BC94b]. Envelopes [Dol99]. Epsilon [Rie94, Wyn66]. Epsilon-Pseudospectra [Rie94]. Epstein [Lee92]. Equal [Kah69, KU71]. Equality [BV92, Eld80, KP78, MG72, Tap74a, Tap74b, Wei92b, Bar88, Cor81, MP78b, Van85, Var85]. equality-constrained [Van85]. Equally [HY95]. Equation [Abl72, ADZ96, Akr99, AWZ96, AV99, Asa98, ABDH95, Atk74, AW80, AL91, BS97a, BJR90, BGT97, Bec92, BK99, BH65, BC99, BC73, BDGG71, BD74, CKS95, CD67, CK69, CD77, Car72, CSH78, CG93, CJ96, DM93, DO92, Den77a, Die92, Dör96, Dou93, Dri99, DL91, Dun97, E92, Ehr71, EG75, Eis67b, EC93, Fal78, FMOS96, FD98, FR68, GSY92, Haa96, Hei96, Her76, Hub65, IB97, ILT69, Iss96, Jeo94, Jok96, KAD93, KM99, Kat69, Kee70, KT86, KK98b, LB98, LC98, Lee91, LVQ95, Lin72, Mar90a, Lor97, LL99, MB97, MR98, McL74, ML69, MS97a, ML92, Mor95, MC93, NdF74, Osh70, Osw92, PSB97, Pet91, Pet87, Pit78, PS83, Pru73, Rak76, Rat95, Rei81, RHH99, Sac71, SS88, Scu98, Sei96, Smi68]. Equation [Smi70, Spi65, Spi67, SH66, SW68, Sül91, SS73, VK90, Yan94, vL92, Atk82, ADK82, BI89, BvA94, Bre89, CEv87, Cha84b, CK85a, Cha86, Dai92, DL82, EF89, EL89, FR86, Fis82, Fre87, Fri81, Gen84, GL83, Hof85, Jam80, LT89b, Lin85, MS83, Mey81a, Mie86, Mon86, Mon87, Nel86, Rau85, Ros81, Sam82, ST86, Smo83, WH86, Win82]. Equations [AP95a, AT97, AMT99, Ale94, AH76, AC96a, Ans67, Arc77, Arc78, AR95, AE94, Asc77, AP91, ARW95, AP95b, Atk67, Atk69, AFM90, Avi74, AC79, Bak71, Bak76, BD76, BK78, BP96, BTW93, BM99, Ben96, Ber95, BCM91, BJ93, BFB92, Boc97, BCMM98, BMM99, Bog71, BW76, BB74, BSTW77, BP89, BC95b, Bre73, BP67, Bro95, Bru90, Bry68, BP71, BM97, But81, BJ97, BGN70, Cad73, Cad74, CM90a, CMM91, CLMM94, CMM97c, CMM97a, CMM97b, Cal71, CH68, Can83, Car74, CS96, Cas82, Cas66, Cas69, CS74a, CS75, CC94, CX98, CDMW98, CDMCW99, Cho78, Cli76, CG73, Coo78, Coo81, CP91a, CP91b, Cor96, Cos71, CGH91, CDW96, CT72, Cry74, Dan67, Dav97, Dav94, Daw93, Daw98, DWW98, Dej67, DF91, Den74b, Den67b, DF76b, DRS95, Doe78]. Equations [Don94, Dör99b, Dör99a, DM79, DD70, Dra94, Dry78, DKR68, Dup68, Dup73b, DFJ74, DS68, Ebm98, Ehr71, EES83, ES93, EH89, EL97, Enr74, EH76, EH98, Est95, EV96, Ewi75, Fai94, FK76, FR92, FX95, FR99, FHK96, FS74, FF91, For65, For66a, FT74, For75, Fun93, GPT97, Gai74, Gar75, Gay81, Gay82, Gay83, Gea65, Gea78, GOQ97, Gir78, Gla69b, GHK94, GM69a, GS76, GL96a, GR96, GM97, Gun65, GM91, GM93, GL96b, Gus69, GS91b, Hal87b, HW98, HKL77, HMM98, HP78a, Han92, Han71, HLLnWJM98, HV96, Hei98b, Hen69, Her75, Hie80, Hil68, Hip99, Hof76, HMT75, HM76b, dHW73, DW74, Hou90, HW92, HW93, HVR88, Hu97, HT82, HS92, HRR94, HR97, HEFS72, HEFS74, Hur67, Huy95, IS90, ITM91, IT91, IK93, IR99, Jac81, Jac84, Jac86]. Equations [JT95, JK94, Jam68, JP67, Jam69, Jam71, Jen76, JPT98, JX98, KN72, KX96, KJ98, Kat69, Kee90a, KG91, KW75, Kel95, Kel66, KL96, Kem75, KP96, Kin73, Kla98, Kla99, Kno92, Kre91, KS92a, KK91, KM96, KT92, Lam67, Lam81, LE97, LL96a, LL96b, LT98, LQ84, LT89a, Le 94, LL94, Lee93, LPG91, LM90b, LM90a, Ler75, LP91, Li88, LFB98, LSX90, Lin90, LTW91, LNC97, Lin97, Lin77a, LW70, Lin69, Lin74, Lin77b, Lin91, Liu99, LT67, Löt92, LNS95, LT68, MQ82a, Mad75, MA81, Mar79a, Mar90b, Mas90, Mat98, McA66, MS98, Met94, Mey68, Mid66, MW94, MT97, Mon91, Mon92, Moo94, Moo99, MDH^{+}99, MH98, MS99c, Nan85, Nas70, Neu87, Nic73a, NW97, NK97, NS93, O’S90, Olv65, OS65, OR66b, OR67, Ort73, Osb76, Osb65, OS91]. Equations [OR93, Ost64, Ost65, Ost70, Ost71, Ott96b, Pai74, PS75, PP96, Pan98, Pao87, Pao99, PS80, PCW95, PW78, Per92, Pet98a, Pet98b, Pet65, Pfl97, Phi72, Pla98, Pot93, Pre73, PR76, QZ92, RR95, Rac73, Rei72, Rey72, Rhe93, Ric78, Ric91, Rob66, RR98, RW95, SD77, SM96, Sar90, SS93, SS95, SV96, Say74, SS99, SZ99, SF84, SM81, SA73, Sha68, Sha69, Sha95, She92, She95, SW99, She78, Sim75, SBBL91, Sta91, Ste65b, Sto68, Sto73, Str74, SWB90, SB77b, Sun96, SHR96, Swa74, Tad91, TE86, Tav71, TYZ95, Tho68b, Tho76, TK96, TV96, Tra96, Tur77, UTK96, VP93, VMSB97, Var71b, Var77, Var80, Ver70, VIA96, VWI97, Wah77, Wat86, Wat81a, Wee98, Wei93, Wei72a, Wei69, Wei74a, WES80, Wer75, Whe73b, Whe73a]. Equations [Wid78, WB98, Woo71, Woz74a, Woz75, Wu96b, Yam75, YF92, Yan96, Yen93, Yin95, ZH91, Zed66, ZW94, ZH94, ZWW94, Zla65, de 86, de 68, van79, van84, vSS97, AF86, ABPR86, And79, AS84, Asc78, APR83, AGS83, AP87, BMS79, Bak80, Bak82, Bal86, BR80, Ban81, BS89a, BR85, Bia89, BQ86, Bru83, Bru84, BS83, BRD81, CL90, CQ82, Cas84, CLS86, CS87, Cha94, CD88, CFT84, CR84, DNT83, Del88, DS80, DE79, Dut88, DL83, Egg83, Egg88, Ell82, Ell84, ES86, Epp84, ED82, Ewi78, FR87a, FN84, FR83, FR85a, FR85b, FS91, FM90b, GS88, GL80a, Gea90, Ger89, GL81, GR82, GM80, GC88, GB89, HLN83, Hal79, Hal87a, HL87, Hay81b, Hei93b, HL82, Hof78, Hu96, HS87, IPZ79, IT80, Ioa81, Jam78, JS87, JS85b]. equations [JSC91, Joe85, Joh88, KK81a, Kar82, KW84, KN88, KP85, KR94a, Kre78a, Kre79, Kum88, KM88, LN82, Las84, LM88, LM78, LL95, LSY89, LP80, Lus79b, LR82, Maj84, Maj85, MR83, Mar78, Mar79b, May84, McK79, Mey87a, MG82, MS80, Mur82, NV83, PS82, PS85, PW80, Pet81, Qua87, RK88, Rhe80b, RS89, RB79, RY86, Roc89, Ros81, Rüm82, SSY88, SE81, Sch85, Sch86, Sch89b, Scr91, Ser83, Sle82, SWC89, Str84, Sun79, Tru84, Vat88, VG86, Vog86, WR85, Whi86b, Wol83, Wu96a, Xu89, YS89, Zir82, van81b]. Equidistant [Wil70]. Equidistributed [ST92]. equidistributing [Whi79]. Equidistribution [HRR94]. equilibration [Ske81]. Equilibria [Kuz99]. Equilibrium [BK97b, KC74, Rhe82]. Equivalence [FM90a, GM69b, HLY80, Hil76, Hil78, Dem83, KL82]. Equivalent [MO93, Ste70]. Erdos [BP80]. Errata [But65, Cra78, DF76a, HEFS74, Ost65]. Erratum [BCL89a, DW82, GN83a, Hic97, Hil78, Laz82, LL95, Pie71, Sob79a, Sym81, Wei95]. Error [AT97, AB99, ADZ96, Asa89, AFM90, AGH97, AK90, BK75, BR78a, BR81, BDR92, BS96, BR70, Bak76, BR87, BW91, BS93b, BS97c, Bar88, BW68, BW69, BK97b, BK98, BDH98, Bau74, BI66, BJ93, BEK96, BV96, Bra66a, BH68, BHM^{+}99, BP67, CSS97, CQ82, CV99, CC94, Che73, Che75, CDMW98, CDMCW99, Cho92, CH72, CG96, CGY98, Coo81, CW78, DS84, DM98, Don94, Don81, Dör99b, DS97, DD96, DMR92, Ebm98, Epp82, EJ87, Est95, FM87, FR92, FHK96, FFLMW99, Fie78, FT74, FP91, FS91, FR78, Gar78, GN91, Gau70, GV83, GT74b, Gri74, Gup75, Han65b, HR86, HR88, HR90, Hic96, HK91, Hol78, Hos95, Hud68, Hug98, Jac75, Jam77, Jef69, Joh69, Joh88, JNT90, JRB95, JS69, JT71, Kam69, KAD93, Kel66]. Error [Ker96, KSC99, Kre78b, Kro81, Küt84, LL83, LV90, LK74, Lee91, Let70, Let71a, Let71b, Let74a, Let74b, LSZ97, Luc74, Luc85a, Lyn78b, Maj85, MX95, MT75, Mei83, Mon86, MS99b, MV99, Moo94, Moo99, Nie99, Obi90, OS65, Olv78, Ost73, Pad97, Pet92b, PS83, Pri76, Pry84, Rei99, Rei81, Rie71, Ros70a, Row67, Rul96, San98, San94, STW97, Sec65a, SWW97, She92, She95, Sho75, Smo83, Spi71, SS97, Ste71b, Ste71c, Str65b, SC72, SH95, SP94, Tad91, TT99, TT95, Tau96, Tho76, Tra96, TW93, Ura68, Ver78, Wah75, Wal98, Wat77, Wei94, Wet97a, Wet97b, Whe73c, Whe73a, WGS90, Woh99, Woz74b, Yan96, Zla65, Zla67, AF86, Bea82a, BC94a, BE79, CL90, DKW82, DRW89, Enr89, Eri86, EJ95a, FR83, FR85a, Gar84, Gat88, Hac81, Hat82, HR82]. error [Hic97, JTW83, Ked81, KM88, Lee85, Len86, LP78, Mil79a, Pan85, Rom90, Sch71b, SK87, Ste79a, SB84, Tre83, Wei95, Zen85, BK97a]. Error-Bounding [Jac75]. Errors [Bak68, BD76, Bas73a, Cha68, Cha69, Chu72, Don73, EK99, Mer66, Pat76, Rac68a, Rac68b, RJ69, RJ74, Sec65b, Str66, vS87, BO87, JS89]. Essential [Des81, GH92, SH95, dW79]. Essentially [OS91]. Estimate [BB71, CMSW79, EL73, Epp82, EW96, Joh76, JNT90, Kel66, Sni72, TW93, War77, Whe73c, LL83, Par87b, Zen85]. Estimates [ADZ96, Arc78, Asa89, AK90, BK75, BO87, Bak76, BR87, BW91, BS93b, BS97c, BK97a, BK97b, BEK96, Bra66a, BSTW77, Bre75, Bre77, BP67, Cal76, CV99, CC94, CDMW98, CDMCW99, Cho92, CG96, CGY98, Coo81, CDW96, DE72, Don73, Dör99b, Dör99a, DS97, Dup73b, Ebm98, FR92, FHK96, FFLMW99, FT74, FP91, FR78, GN91, GC96, Gup75, HR86, HR88, Jef69, KAD93, Ker96, LV90, Lin91, MX95, Mit77, MV99, Moo99, Ost73, PV72, PS83, Rei81, SWW97, She92, She95, Sho75, Shr73, SWB90, Str65b, SC72, ST67, Tad91, TT99, Tau96, Tho68a, TW74, Tho76, Tra96, Ver78, Wah74, Wah75, Whe73b, Whe73a, Yan96, YB96, YZ94, Zla65, Zla67, BR78a, BD85, Bea82a, BS83, CL90, CQ82, DRW89, Don81, Eri86, EJ87, EJ95a]. estimates [FR85a, GMP93, Hac81, HL87, HR82, Joh88, KM88, Las84, Len86, LP78, Maj85, Mil79a, Mon86, Pan85, Sam82, Smo83, Wig87]. Estimating [BRD81, Dix83, Pru73, RJ69, Sch83c, Str66, Whe74, Hil80]. Estimation [AB99, BS96, BTW93, Bar70, BH70, Cha68, Cha69, CM83, DD96, Her76, Jam77, Lee91, MS76, MS99b, Moo94, PS77, PT79, Pri77, Sec65b, Ste71b, SW68, AF86, BE79, MM79, Ste79a, SB84]. estimations [Kro81]. Estimator [BDR92, MS76]. Estimators [AT97, BV96, DMR92, O’S90, Pad97, Woh99, Ste80]. Euclid [Dun74a]. Euler [AFM90, BC95b, CGH91, CT72, Dra94, ES76, EV96, Gen84, Hal79, Hal87a, HLLnWJM98, Hou90, HT82, Huy95, JNT90, Kre91, LQ84, Mas90, Nav65, Per92, RR95, TYZ95, Yin95]. Eulerian [EW96]. Evaluating [DM93]. Evaluation [AG68, Atk72, Atk73, AV93, For73, GA67, Gün80, Kam78, KSC99, Lan64a, Mar83, Mie81, Van81a, WF97, Woz74b, ZT67, AR81, CM88b, Gra87b, Mar85]. Evaluations [KT76]. Even [DS97, Sch69b, VS97, WB86, Bak82]. even-order [Bak82]. Evolution [Tav77, Yan99, Epp84, Lus86]. Evolutionary [GM93, Mor98]. Exact [DMR92, Hol78, Pie69, Pie71, RSK76, Yan99, Gar88]. Exactly [BFO96]. Exactness [WB86]. Example [Dup73a, HR99]. Examples [DeL94, SK87]. Exchange [Dun74b, GCB74, Tse96, Wat75]. Exclusion [Chu87]. Existence [BS76, BQ76, BB88, Jam71, Jer73a, LP91, Moo77, Nic82, Rua89, San94, SS83, Tho68b, Wei92a, FM84, Ral80, Sch83c, SW85, Ste79b, Whi85, Wil81, Xu89]. Exit [RM77]. Expansion [Str74]. Expansions [AD68, AFM90, BH71, Cle64, KR74, PK90, SP94, Zan99, de 86, DS89b, Gus80, Hag86, Lee85, Luc82]. Experiments [CC67, Geo77]. Explicit [ARW95, BP96, BS94, CV72, DD94, DM98, FR99, Gun65, IR99, KL67, PTP96, Ric91, TS93, Ver78, vS87, APS87, BH93, CLS86, Cha85b, Dai92, OS92]. Explicit/Implicit [DD94]. Exploiting [ABGM92, VVC94]. Exponential [GS96, HL97, Ise81a, Van77, MS98, Nor78, Saa92, Sch79, Sch90a, War77, BMP89, Cas81a, CO80, Ise79, Tad86]. Exponentially [Dör99b, JK74, Vos88, SS83]. Exponentials [Kam76, Kam79a, Kam79b]. Exponents [DRV97, Tal91]. Expressed [IS90]. Extended [Cas81b, Law66, Law67b, CK85a, Gün88, Var78]. Extension [AV99, EZB83, MS68, Nic67, Ros70b, SW88, Bea87]. Extensions [Dag86, Gla69a, Rei74, Ste68, SG68, DH79, SS96b]. Exterior [HZ94, ADK82, Gui87, Lin85]. External [RT95]. Extra [Haa96]. extraction [BDGB83]. Extrapolated [AK84]. Extrapolation [Bru91, CJ76, FK76, Gra65, KPH95, LSX90, Sid97, FS87, GM80, Kin79, LM78, Sid86]. Extrema [Bac80]. Extremal [GW90, GL93, Sin73, Dix83]. Extreme [Zha95, Wil81].

Faber [Cur66, Ell83]. factor [Que89, Wol83]. Factored [Ip87]. Factorization [BB87, Ben95, BBS73, BB74, Car92, Dun74a, DKR68, Dup68, Mei83, Say74, Ste77, Wil69, BM87, Sco81]. Factorizations [ABP93, BQ76, CJZ83, CN88, Mar90b, BB88, JA83, VSM80]. factorized [Sas78]. Factors [Gra71, Yav95, BvA94, Wig87]. Fading [SWW97]. Families [DAB99, Gor75, LG80, Son98, Ver79]. Family [AD78, Kin73, Mar90b, PTP96, BT84, Cha89, GM83, SSB85]. Far [Kar96, MS99b]. Far-Field [Kar96, MS99b]. Fast [AP95a, BFB92, Bia93, Bjø83, Bör90, Buz77, CG73, Cor96, DGR96, GHK94, HYS94, HKT92, Jaf92, Kel95, Li98, Pai79, SW99, SW94b, WS93, Yan94, YR99, Die82, ES86, KW84, LNOP82, May84]. fast-direct [KW84]. Faster [de 78b]. Feasibility [Avi74]. Feasible [PP73b, BPTZ92].

Feature [OR90]. Feature-Oriented [OR90]. Feedback [KK91]. Fejér [GT82, TG83]. FEM [BS97a, Car96, CG97, IB97, MS98, MS99c, Ter98]. FEM-BEM [MS99c]. Fenchel [Cor96]. Fer [Zan99]. Ferromagnetic [MV99]. Few [BG74, Hof76]. Few-Point [BG74]. Fewer [BW76]. FFT [CNP93]. FFT-Based [CNP93]. Field [Dub90, Hei94, Joh78, Kar96, MS99b, Per97, Fri81, Lin88]. Field-Based [Per97]. Fields [BJK90, GS67, YR99, Rhe80b]. Fifth [Cas66, Cas69, KL67, Sch71a, Str69a, Str72, Vos88, Hor83]. Fifth-Order [Cas66, Vos88, Hor83]. Fill [GL75, LNOP82]. fill-in [LNOP82]. Filter [Wal99]. Filtering [Jen70, Kar96, Maj84, Maj85, Sei96]. Filters [AM94, CGS92, OR90]. Finding [CJ76, Gar78, Rit90]. Finis [Cia75]. Finite [Abl72, AP95a, Ain96, ADZ96, Akr99, AW95, AWZ96, AWY97, AV99, AE94, Asa98, AK98, AL78, AW80, Bab71, BZ73, BK75, BA76, BR81, BSK81, BS87a, BDR92, BS92, BCO94, Bai91, Bak76, BD76, BS97b, BS93a, BTW93, Bao95, BL94, BK97a, BK97b, BM99, Ben92, BCV95, BDM^{+}95, BLMPT92, BG98a, BFO96, Boc97, Bof97, Boi97, BW90a, BDLN92, BV90, BH65, Bre92, Cad73, CM90a, CMM91, Cal76, CH68, CG98, CV99, Cha92, CN97, CS99, CC94, Che95, CE97a, CE97b, CDMW98, CDMCW99, Cho78, Cho92, Cia75, CCL95, CG96, CJ96, CF93, CS90, CDW96, DG98a, DD94, Daw98, DWW98, Der71, Des72, DL90, DM98, Dob92, Doe78, Doe79, DR84, Dör99b, DR82, Dou83, DG90, Dun97, Dun75, Dur88a, DGW91, DMR92, DT73, Ehr71, Eis66]. Finite [Eis67a, ES98, Epp84, EJ91, EJ95c, EJL98, ELW91, ELV94, ELW99, Fal78, FM90a, FR92, FR99, FF93, FMOS96, FFLMW99, FL71, For90, FR78, Fre90, GV91, GGS94, Gar66, GW97, Gek75, Geo73, GM78a, GG90, Gir74, Gir78, Gol94, GH92, GM91, GM93, GC98, HM76a, HR95, HW98, HL84b, HL98, Hei96, Hei94, HR82, HR86, HR88, HR90, HMR73, HW97b, HW92, Hud68, HS99, IB97, Jam71, Jam77, JV96, JK91, JY98, JN98, JNV99, JS98, JW91, Jim96, Joh77, JPRT96, KJ98, KM99, Kee70, KL96, KL97, KP97, Kir93, Kos73, KR94b, Kut70, Kut74, LL96a, LN82, LSS94, LMV96, LP77, Li88, Li91, LVQ95, LB97, LL98, LTW91, LB96b, Lor97, Lov96, LL99, LM92, Man78, Man80, MR98, MX95, Med98, MM81, Mil81b, Mil92, MW94]. Finite [Mil97, Mit77, Mon92, MF90, Moo94, Moo99, MS65b, Mor98, MDH^{+}99, MS97b, MTL96, NP81, Ned78, Nic72, Nic73b, Nie99, Nit89, OR76, OPF97, Osw92, Pal96, PP96, Pan98, Par95, PCL94, Per76, PW78, Pet99, QZ92, Rei68b, Rei81, Rey72, Rhe78, Ric91, RC66, Ros73, Rus77, Şab97, San98, San94, SDC88, SSW96, Sch80, SE81, Sch96, Sch89a, Sco75, Sem92, She96, Sim71, Smi68, Smi70, Son98, Sta99, Ste71a, Sül91, SHR96, SW74b, Tav77, Tha78, TXZ89, TV96, TH98, Tra77, TW93, VMSB97, Var80, VW95, Vav96, Wah92, Wat86, Wet97a, Whe77, Whe78, Woh99, WH97, Yse86, ZH91, ZH94, Zla65, Zla73, Zla74, dd86, AF86, AVG89, AM89, AD92, Arn82, AF89, BR78a, BO83, BG88, BCMP91, Ban81, BS89a, BR85, Ber89]. finite [BN83, BN85, BX89, CL90, Cas84, Cha84b, CK85b, DO79, DL82, Dob80, Dor84, Dor86, Dry83, EF89, EHP87, Eri86, EJ95a, EJ95b, EYL89, FR87a, FN80, FR85b, Fre87, Gat88, Gen84, GL78a, GL78b, Geo80, GGT82, Gui87, GH83, HL81, Hof78, KW85, KHR89, LL83, Lar89, Len86, LT88, LNOP82, LB96a, LP80, MN87, Mel87, Mil79a, Mil79b, Mon87, Nel86, NV83, PS85, PW80, PR82, PR86b, Rau85, Rua89, Str84, SO86, Sun79, SB84, VG86, Wah80, Wei89, WW88, Whi86b, Win82, Zho83]. Finite-Difference [Bro87, CH68, HL84b, KP97, Kut74, Lor97, ZH91, ZH94, LN82, Nel86, NV83, PR86b, VG86]. finite-differenced [AVG89]. Finite-Dimensional [Ste71a].

Finite-Element [Ain96, Bai91, Cia75, DG90, Fre90, GM91, HR90, JW91, KL97, Kir93, LB97, LTW91, MR98, MF90, MTL96, NP81, Nit89, TXZ89, WH97, Ber89, BX89, CL90, EF89, EYL89, Lar89, Rua89]. Finite-Rank [LL96a]. First [Arc78, AW83, AL78, BFO96, BCMM98, CLMM94, CMM97c, CMM97a, CMMP98, CS96, CG90, DW76a, Den74a, Dry78, Dup73a, Gup75, Han71, Her76, Hil76, Hil78, HMT75, HM76b, dHW73, LE97, Löt92, Ree90, Ric78, She92, Sig76, Smi73, SW68, Str74, ddd83, van79, BRD81, Egg83, GM80, Gup85a, Gup85b, HL82, LM78, Lay83a, LP80, Mar78, Vog86, YS89]. First-Kind [LE97]. First-Order [BFO96, BCMM98, CLMM94, CMM97c, CMM97a, CMMP98, CG90, Löt92, She92, ddd83]. Fitted [Dör99b, JK74, Vos88]. Fitting [BS87b, BT68, HS97, Bea82a, BMP89, Cas81a]. FitzHugh [Jer80]. Five [Law66]. Fixed [Atk73, BH68, CM99a, FD91, HKT92, KLY76, Mar92, Ske76, Ste75b, Ura68, Wei74b, Yam98, ST78]. Fixed-Point [HKT92, KLY76, Mar92, Ura68, Yam98]. Fixed-Stepsize [Ske76]. Flap [MS97a, Scu98]. Floquet [FJ91]. Flow [AOW93, AWZ96, BYLP98, BFO96, BJ91, CE97a, Cho92, Daw91, DG90, Dzi99, GC95, GS91b, HM76a, JS96, Kee95, LeV96, MS99a, Ott96a, Pet92a, RT95, San98, San94, Wet97a, ZH91, ZH94, Arb89, BSW85, Dou83, EYL89]. Flows [AP73, BC95a, BLMPT92, CN88, DG98b, LP91, Pad97, BN85, MN87, Wal88]. Fluid [BYLP98, BDM^{+}95, BFO96, COP98, CP98, CP91b, GC95, JRB95, San94]. Fluid-Solid [BDM^{+}95, COP98]. fluids [BSW85]. Flux [BCMM98, BMM99, CGY98, Tar96, Whe74, BO86, Swe84]. Flux-Splitting [CGY98]. Fokker [HV96, HV98, Sch98a, BTW93, BC99]. Foldpoints [DR90]. Folds [FR86, JS85a]. following [SC87]. Fonctions [DS89b]. Forced [Bra66a]. Form [AC96b, BYLP98, Ip87, IS90, JN95, KG91, RR80, van95, Ber86, CLM89, Des81, Mee80, SO86]. formation [JP92]. Formed [DS92]. Forms [KSC99, Rat80, Rei68b, KN85]. Formula [Ahl66, KG91, Lau89, MS68, NPV93, Str69a, Ver90a, Ver91, Vos88, Wil70, CL84]. Formulae [Cas88, Fre98, MS73a, Not91, Sch65, Sch66, SP94, MP85]. Formulas [Bak68, BH71, BS76, Bol72, Bro77, BC72, Cas81b, CM74, Cha68, CJ75, CR76, Enr93, EH72, EMB71, Fri71, Gre66, Gun74, HP77, HW83, Hue73, JEH78, JN95, Jet87, Kah69, Köh95, Lip73, MR77, NtB90, Sch98b, Spi71, Sta64, SS65, Str67a, Str67b, SG68, Str69b, Str70, SCWM71, SC72, Str72, Str73, Str75b, SB77b, Tre72, WB86, AE84, CM88b, Fei82, GM78c, Jac88, Jet82, KD83, Lin83, Ros80, Ske86, WR85]. Formulation [CD98, Fun90, HW98, ILT69, Lee98, Per97, Whi82a, FNO87, Whi82b]. Formulations [CS99, Kar92]. Fornberg [Weg86]. Forsythe [Hou73, Ano73]. Fort [GG76]. Fort-Frankel [GG76]. Fortran [Sin73]. Forward [AL91, VK90, YB96]. Forward-Backward [VK90]. FOSLS [CMMP98]. Foundation [Ols77]. Four [Gun74, KL89, Kin79]. Four-Color [KL89]. Fourier [BH70, CGS92, CK85a, For75, GL96b, Hei96, KO79, Lyn67, PH73, Qua87, Tad86, VIA96, VWI97, Wei93]. Fourier-Finite-Element [Hei96]. Fourth [Bel97, CJ96, Eis67b, Hei93a, Hor83, IT95, JK74, Kin73, Man78, Mor95, Sem92, SBBL91, Zla65, Zla67, NV83]. Fourth- [Hor83]. Fourth-Order [Bel97, CJ96, Hei93a, Mor95, Sem92, SBBL91, NV83]. Fox [Nic67]. Fractal [DD96]. Fraction [Mer66]. Fractional [OR93, TT95, Ten94, HP80, KR94a]. Fractions [Fie78, GW83, Jef69, JS69, JT71, ST67, Van81a, Wyn64, JTW83]. fractured [Arb89]. Framework [Har96, FR87b, MM84, Ske82]. Frankel [GG76, Wu96a]. Frankel-type [Wu96a].

Frechet [For66b]. Fredholm

[AE94, Atk67, Bru90, Han71, Her76, Joe85, Mar78, NdF74, Rak76, RB79, SW68]. Free [BCR75, BHM^{+}99, BH86, CGH91, CH94, Dub90, Jup78, Kac99, Sac71, Sch68b, Wim70, ZB96, GH83, Han93b, LNOP82, Mey81b, Wom89]. French [Cia75, DS89b]. Friedrichs [Wah74]. Front [LW81, CT86, Wom89]. front-tracking [CT86, Wom89]. frontal [AH82]. Full [Che95, RW89a, BI89]. Fully [BYLP98, CE97a, Gen86, Hei94, Hue73, ITM91, Jer84, KD83, KT92, Let71b, Let74b, MR77, Moo94, NP81, NPV93, NV97, Rüd93, Sam83, VTG91, Wet97b, Yan99, YZ94, Asa86, Bal86, Esp87, GM83, IKP91, JP83, KL79]. Fully-Discrete [CE97a, ITM91, Moo94]. Function [ADH99, AM98, Bas73a, BS87b, BYLP98, BC94b, Bru78, Coh71, Con73, Con76, CP77, Gau70, Gor71, GS96, Heb71, Hug98, Kem75, Kur67, Lan64c, Li83a, LMST69, LT67, Lyn78a, MS66, MS65c, Mou65, Mou66, Mou68, MT68b, Nor78, Pel67, Prü82, Sch69f, Sha70, Shu72, Ste75a, Str74, Wat77, Wei94, Wil69, CO80, Gou89, Gut83, Ise79, Kra87, LB96a, Pie95, Ros80, Sma88, Wei95]. Functional [CS75, CT72, Den67b, Jac81, Jac84, Jac86, Kem75, Liu99, Lyn78b, ML92, Nav65, Tav71, Tho68b, JS83, KK81a, LM88, Rag85]. functional-differential [KK81a, LM88]. Functionals [Aal72, AN66, BH70, CM74, Man71, MS99b, Sim68, Sta99, GC88, KR88]. Functions [AD68, Ahl66, BW68, BW69, BN74, BCS78, BT69, CGS92, CD85, CT83, Cox84, Doo66, DAB99, DL75, Dur77, EW72, El 90, FP91, GGK97, Gra87a, Gre64, Hei66, JS68a, Jer73a, Jer73b, JK68, Kam76, KR97, KZ66, KL76, Köh95, KR74, Let74a, Lip73, LM67, MRW96, Mal77, Man71, MF97, Mil66, MT68a, Mun73, Nav65, Not91, Olv65, Pet92b, Por71, Pri76, RJ74, Rou71, Sch70, Sch69a, Sch68b, Sec65a, Seg72, Sin73, Spo94, Str66, SP94, Uso67a, Uso67b, Wal67, Wei71, Wei72b, Wil78, YA95, AR81, Bas85, BT84, BO86, Die82, DS89b, GV83, Hak82, Hop82, Kam79a, Kam79b, Kam81, KRD78, KN85, Mie81, PR86a, Pri79, RS87, Sas78, SB88, Swa79, Swa81]. Fundamental [BL80, Bog85, MJ77a, Zem95]. Furnace [Cho92]. Further [DL73, Gar78, McC84].

G [Wuy70, PA73]. G-Transformation [PA73]. Galerkin [DVW94, DF76a, FKD84, BO87, Bak80, Bak82, Bak84, BK98, Ber95, BJMT73, BW93, BL87, BSTW77, CEv87, CG90, CX98, CM90b, CS98, CFT84, Den74b, Den74a, DF75, Den75, Den77c, Des81, DP92, DM92, Dia77, Día79, DS80, DD70, DDE79, Dry78, DW76b, Dup73b, Dup73a, DFJ74, Dup76, DN89, DHLR84, Egg88, EZB83, Ewi78, EW80, Fai72, FF91, Fin71, FS91, GANT98, Gek76, GC96, GGOS96, Güs81a, Haa96, Hac81, HL87, Hay81a, Hed79, HR93, Ioa81, Jam78, Jes78, KX96, Kar82, KM99, Kee90a, KW76, Ker96, Lay83a, Lay83b, Lin90, LMS97, Lus79b, LR82, MB97, MT89b, MX95, Mie86, Mir71, Mos83, Pan98, Pet91, PV72, Pla98, Rac73, Rus85, SSA84, Sch69e, Sch71b, Ser83, ST95, Slo76, SBBL91, SW74a, TW74, TW75, Tra96]. Galerkin [Urb91, Wah75, Whe73b, Whe73c, Whe73a, Whe74, Whe75]. Galerkin-Characteristic [Ber95]. Galerkin-type [Jam78]. Galerkin/Least [CG90]. Galerkin/Runge [Kee90a]. Gamma [Lan64c, Spo94]. Gap [Gai66]. Gas [Cho92, Ein88, Per90, Mun94]. Gases [Bou98]. gauge [LB96a]. Gauss [Mor71, BS76, Cha69, Cry67, De 73, Ehr98, ELW91, GTV90, Jet87, Kea90, Let71a, Mor72, OR66b, OR67, Por69, Por71, RJ69, Rie71, SC72, Yav95]. Gaussian [AB86, Bro69, Cha68, Chu72, DS84, FP91, GV83, Geo76, Jet82, Koc84, MRW96, Pet92b, Rab67, SS65, Zla80, dS73].

Gegenbauer [VIA96, VWI97]. Gene [CGO07]. General [But81, CMM91, CS72, CGT91, Coo78, Coo81, CF93, CP98, EZB83, Gel81, HW97a, Har96, HLLnWJM98, HZ90, Hub65, JT95, JS98, Joh78, Lin77b, McL74, OR72, PT91, PTP96, Par87a, Pet98a, SW94b, Swa72, Van75, VAF98, ZW94, BD85, DK81, Gus81b, McC85]. Generalization [Fre98, Gar66, Jan83, Mar66, Mar79a, SA73, FS87, Lev85]. Generalizations [Fos81, FMMR99, HV70, Mor72, You72a]. Generalized [BO83, BI66, BIC66, BCM88, BN73, CR86, CL64, CK98, Cli65, Cra76, Cra78, DAB99, Drm98b, Eis83, EL72, FH72, GBIY71, GG76, GR96, Gus70, Her67, Kau74, KG91, KS94, Kub70, Law67a, Let71c, Li87, LRT79, Loe66, LMST69, MRW96, MS73d, MS97a, Mou66, Mou68, MT68b, Mun73, Nob66, Osh85, PS81, Par65, PW70, Rie94, Riv84, Sch74, Scu98, Sid97, SS74, Stu79, Swe74, Tew67, Van81a, Woz75, Chu87, Dyk87, GL81, GR84, Ise79, JS83, Jam80, Kra87, ME81, Nic82, Pai79, SW82, Sco81, Sun83]. Generalizing [Van76, PR98]. Generally [Tho67]. Generated [Nic72, Nic73b]. Generating [AD68, Wil69, Sma88]. Generation [Dar91, Wel97, Ste80]. generic [ASW93]. Gennes [DG98a]. Geodesics [Gov86]. Geometric [PP73a, ZZ98, FR87b, GL88]. Geometrical [Mil97]. Geometrical-Mechanical [Mil97]. Geometrically [DM95, Kel81]. Geometries [EZB83, Ker86]. Geometry [CHL91, PS83, Asa86, LN82]. George [Ano73, DER76, Hou73]. Gerschgorin [HV70, Loe70, Var70]. Gerschgorin-Type [Loe70]. Gibbs [GS96]. Ginzburg [DNW98, DW99, Lor97, MH98]. given [Bia89]. Glimm [LTW95a, Luc85a, PS88, WW93a]. Global [Arc78, BP96, BNY87, Car91, Che98, CGT88, CGT89, CR97, EH75, EA91, ES93, Est95, Hil97, KM93, Mor71, Mur82, Pan97, RR87, Shu72, Ste71b, Bas82, Bas85, BT89, CL92, HP78b, SSB85]. Globally [CGT91, Nic81, Pol74, Wat79, Sch86]. GMRES [Mor97]. Godunov [Daw91, Daw93, BGS98, EL96, Ein88, Lev85, LTW95a, LTW95b, Luc85a, Mun94, NTT94, Vil86, Vil89, WW93a]. Godunov-Mixed [Daw91, Daw93]. Godunov-Type [Ein88, Mun94, Vil89]. Golub [CGO07]. Good [Rip92, Esp87]. Gordon [Dun97, LVQ95]. Governed [Ben96, KK91]. Grad [Cha92]. Grad-Div [Cha92]. Gradient [AMS90, Bar70, Car91, CD90, Coh72, Dan67, Dan70, Eis83, FMMR99, GHK94, Gut93, HS94, KN72, KS94, Man90, Mil97, TXZ89, dT93, FM84, For79, JMP83, Naz79, SSY88, Sha78, Ste83, Win80]. Gradient-Type [KS94]. Gradient-Weighted [Mil97]. Gradients [Rei72, Nic87]. Grading [CD85, YS89]. graph [CM83]. Graphs [Bau74]. Grau [Car92]. Gravity [Shi66]. Greatest [Bra66b]. Green [BT69, PR86a]. Grid [AD92, BGJ88, BGT97, Ber87, CDS99, CGH91, Dar91, DWW98, Geo77, GPV78, Löt92, MX95, MDH^{+}99, Ott96a, Par99, Xu96, Die82, Dou84, Hac79, Par87b, San85]. Grid-Aligned [Ott96a]. Grid-Free [CGH91]. Grids [AV99, CM90a, CKV98, CGY98, ELPV93, ELV94, For90, HMR73, JLL^{+}98, JW91, MS94, Moo99, NV93, Wu96b]. Ground [CE97b]. Groundwater [CE97a]. group [ED82]. Groups [BT80]. Growth [CSS97, Gai66]. GSV [Pai84]. Guang [Pai84]. Guide [Hol78]. Guided [JPRT96]. Gummel [Gar93, Ker86].

H [CGO07]. Haar [Dun71, Liu94]. Half [Atk69, BH93, CK69]. Half-explicit [BH93]. Half-Line [Atk69]. Half-Strip [CK69]. Halfplane [Ahl66]. Halley [Pet89]. Halley-like [Pet89]. Hamilton [Cor96, JX98, OS91]. Hamiltonian [CSS97, DDLT91, Jay96, Rei96]. Hammerstein [KX96, Kum88, Sar90]. Hankel [KK93]. Hard [Hoh99]. Hardening [Car96, Joh77]. Harmonic [Alo97, Bao95, BS76, BS77, Doo66, GW74b, GM97, Hud68, SCWM71, Str72, Tay95, HB86]. Having [Aal72, ST69, Tay68, Cas82, DK85, JS85b, JSC91, LM77, Olv65, OS65]. Heat [AL91, CKS95, CD67, Can68, CK69, Hub65, KT86, MB97, MC93, Osh70, Sac71, Sei96, Tar96, VK90, GL83, Jam80, LT89b, Mon86]. Heatless [Bou98]. Hellerman [EGLP85]. Helmholtz [AW80, ADK82, BS97a, CD98, IB97, Lin85, WH97, Zem95]. Helmholtz-Type [WH97]. Hencky [Ben92]. Henrici [Bir83, Nic67]. Henry [de 78a]. Hermite [SP68, AS69, Aus71, BFB92, Bia93, BC94a, CJMT74, DHLR84, Fer74, Hed79, Sun96, Sun99]. Hessenberg [KG91]. Hessian [Byr90, NO85, PT79]. Heterogeneous [GK97]. Hexagonal [LM77]. Hierarchic [BS96]. Hierarchical [BS93b, Osw92, Woh99]. High [AR77, BS97a, Bur87, CV72, Enr93, EV96, FF98, FN84, For90, GC88, HW81, HL84b, HK79, dHW73, DW74, IB97, JLL^{+}98, KR97, KW84, LeV96, LB97, LNC97, LO96, OC84, OS91, Swe84, Var71b, Ver93, Yan99, van80, van84, GL88, HO87, Sch80]. High-accuracy [HK79]. High-Order [Enr93, EV96, FF98, For90, KR97, LB97, LNC97, OS91, Ver93, GC88, KW84, HO87, Sch80]. High-Resolution [JLL^{+}98, LeV96]. Higher [AL96a, Bal86, BP89, BF91, GAM71, GM97, HR88, PV72, PA73, Rei68b, SC87, Sin77, BE88, Fos81, GZ88]. Higher-Order [BF91, GM97]. Highly [AP73, AC79, Cas88, Cry74, Den77c, EL97, Güs81a, PT97, BS89a, Pet81]. Hilbert [Aub68, De 73, For79, Gri87, Hil76, Hil78, KT97, MM79, Neu88, RD71a, Sch98b, Tau96]. Hill [Seg72]. Hilliard [EF89]. History [MR85, MR84]. HIV [ILM^{+}96]. HIV/AIDS [ILM^{+}96]. Hodgkin [Mas90].

Hold [Les91]. Holder [Heb71, Car94]. holomorphic [Mül84]. Homoclinic [AM98, BK97c, Sch93]. Homogeneous [KT86]. Homogenization [DE98, NW89]. Homotopies [LW92, LZC92, VVC94, VH93]. homotopy [LSY89]. Hood [BF91]. Hood-Taylor [BF91]. Hopf [GHK94, GGK97, GMS97, JP96, LNC97, Wer96]. Hopscotch [VS97, ED82]. Horner [Rei68a]. Householder [Tsa75]. Hurwicz [Que89]. Huxley [Mas90]. Hybrid [ELV89, Gea65, KMR94]. Hydrodynamic [JRB95, Mey81a]. Hydrodynamical [HK91]. hydrodynamics [GH83]. Hydrogen [FD98]. Hydrogen-like [FD98]. Hydrology [CE97a]. Hyperbolic [Arc78, AV99, Bak76, BD76, BS97b, BL98, Bre75, Bre77, CJR97, Che75, Che77, CK92, CDW96, DF75, DF76a, Don94, Dry78, Dup73b, EH89, Fai94, FR92, FR99, FF91, For75, FP95, GGOS96, GM66, GM69a, GL96a, GYF86, Gus69, Har90, Jam71, JLL^{+}98, KS92a, KKP94, Lay83b, LTW95b, Liu93, MR85, Ott96b, Pet91, Şab97, SW91, SW74b, Tad91, TE86, Tur77, TW93, Wal99, Whe73b, Bal86, CQ82, CT86, DS80, FR87a, GGT82, GLT87a, GLT87b, GG86, Gus80, GK83, HL81, Hig84, JN83, JS87, Luc85b, Lus79a, Lus86, MR84, San85, Ser83, Swe84, van81b]. Hyperbolic-Parabolic [KKP94]. Hyperbolicity [HLLnWJM98]. Hyperbolics [Den74a, Dup73a]. Hypercube [Sob79a, Sob79b]. Hyperelastic [GW97]. Hyperpower [GBIY71]. Hypersingular [FHK96]. Hypothesis [Den69]. Hysteretic [CW97].

IBVP [GGOS96]. ICCG [PO87]. Icosahedral [BF85]. ideal [MP78a]. Identical [Lam71]. Identification [BM88, Bur90, KS92b, Ler75, dT93]. if [OS92]. Ignoring [DR65, ET71, Mye91]. II [ALM92, AG68, AV90, Ban77, Bar67b, BK97b, BMM99, BPTZ92, Bur90, CMM97c, CDMCW99, Coc90a, DK80b, DOR88b, EL96, EJ95a, For66a, Fri81, Gol98b, GM80, GL93, Hal79, HS67, HR86, IB97, IIMPL82, Jer73b, JNV99, LP80, Maj85, MG72, MS89, Mil81b, Nic73b, RD71b, Rob76, Sch66, Sch68b, ST69, Sha69, SW94b, Smi70, Spi67, Str70, Uso67b, Ver88, Wan93b, WW93b, Yam75, Yan99, Zla74]. II. [FP95]. IIA [AC98]. III [BDLN92, Coc90b, CGY98, DK70, GS96, Hal87a, HR88, JSC91]. Ill [Bog77, FH72, Liu94, Neu97, Pla98, Qn99, SEK93, Var73, Car94, Mon86, Tru84]. Ill-Conditioned [Bog77, FH72, Var73]. Ill-Posed [Liu94, Neu97, Pla98, Qn99, Var73, SEK93, Car94, Mon86, Tru84]. Image [ALM92, AM94, AV97, CSH78, Car99, CLMC92, CM99a, KLM94, OR90]. Imbedded [JS92, SV94, Ver79, GM83]. Imbedding [BW90a, Bör90, Gol77, Mey68, KL82, Mey81b]. Immersed [BL92, LL94, WB98, LL95]. impedance [Bub85]. imperfect [JP92]. Implementable [Bur82, HW81]. Implementation [BJ97, Dra94, Pot93, Sto73, van95, Bar88, GL80b, Mos83, Tru86, Var85]. Implementing [DG94, Tsa75]. implications [Naz79]. Implicit [Ale77, AP91, ARW95, AFM90, BF99, Bic77, BP89, Bro77, BCM94, BJ97, Coo72, DD94, DL91, EW91, Hof76, Liu99, MPS98, Pet98a, Pet86, Roc89, Rul96, Sac71, Say74, Sch96, STW97, Sta91, Sto68, Ver70, vS89a, vS89b, BB79, BC89, CH95, FSU85a, FSU85b, GG86, HS87, SWC89, Zir82]. Implicit-Explicit [ARW95]. implicitly [AS85, AG87]. Improper [AG68, GA67, GW92]. Improved [BK98, BR73, CH72, NV93, Sch71a, Sni72, YR99, Bra80, DRW89, Kar78, Wig87]. Improvement [Gra71, Lin72]. Improving [AC98, DMW83, Kov70, BX89]. Inaccurate [Oet65, CR79, DW85, MM79]. including [BH83]. Inclusion [GK86]. Inclusions [Vel92]. Incomplete [CJZ83, DDE79, GPV78, VSM80, Rus85]. Incompressible [BYLP98, BFO96, Dob92, Dra94, DG90, GS91b, HW93, Kre91, LeV96, Lin97, Wet97a, Yan96, ZW94, Zha97, Dou83, Rua89]. Increasing [MF98]. Increment [Gar75]. Indefinite [BKP94, BP71, PS75, Wan93a, Xu92, Yse86, Axe79, Ban81, CW93, Saa83, Wan93b]. Independence [DP92, LL96b, ABPR86, AB87, DM82]. Independent [Löt92, Wat86]. Index [AL96a, AC98, BM96, BP89, CO87, HMM98, LO90, Pet98b, BH93]. Index- [AL96a, HMM98]. indices [Gea90]. Indirect [Jeo94]. Induced [Drm98a, Kla98, RD79]. Inequalities [BW94, Dan73, Dan75, HR95, Hie80, Hop87, MG72, Mit87, Rob75, Rob76, Sig76, Bea82b]. Inequality [Boh77, BPTZ92, HN75, Joh76, Mea69, Yan99, Mar83, MP78b]. Inertial [GANT98, HR93, Lor97]. Inexact [BPV97, Bro87, Car91, Che98, DES82, EG94, Fon90, Ypm84, Naz86]. Inexact-Newton [Bro87]. Inexact-Newton/Finite-Difference [Bro87]. inexpensive [Mar85]. inf [BCM88]. inf-sup [BCM88]. Infinite [Asa98, Dun74b, Faz96, Jia86, KU71, Ma95, PCW95, CG84, Gab85, LK80]. Infinite-Dimensional [Ma95]. Influence [Eps76]. Information [Byr90, Car91, Woz75]. Infrapolynomials [Shi66]. Inhomogeneous [GH92]. Initial [Bel69, CW96, Cal74, Cal76, Che75, Che77, Dut90, EH76, Gea65, Gek75, Gek76, GG76, Gra65, Her68, HT82, Kin72, MM73, Mil69, NT94, PL87, Rus70, TW74, Var71a, Wat81a, Wer75, Whi82c, ZWW94, dC90, van80, GLT87a, GLT87b, Gus81b, HK79, Pas82, PR82, Sam83, TK84, Ver82, Wat81b]. Initial-Boundary [Cal74, Cal76, Gek75, GG76, GLT87a, GLT87b, Gus81b, PR82]. Initial/Boundary [Whi82c]. Initial/Boundary-Value [Whi82c]. Initialization [LPG91, Wat81b]. Initially [GLV91]. Inner [GGK98, RD79, Gra87b, Nel86]. Instabilities [PS88]. Instability [Die90, HW83, TT87]. Integer [Eri80, MR77]. Integral [AC83, AT86, Ans67, AE94, Atk67, Atk69, Atk74, Atk75, Bak71, BK78, BHL96, BJ93, BW76, Bru90, Bry68, CS96, CX98, Cos71, Dav94, DF91, DRS95, Far66, FHK96, GGS94, GGS98, GTV90, GHK94, HP78a, Han71, Her76, Her75, HMT75, HM76b, dHW73, DW74, Hu97, ILT69, Jam68, Jeo94, Kel95, LB98, LSX90, Lin69, Lin91, ML69, Mot66, NdF74, PCW95, Pre73, Rak76, Ric78, SS95, SV96, SW68, Str74, Tra96, Yan94, vSS97, AGS83, AP87, Bia89, Bru84, BRD81, Cha86, Cha78b, Cle81, Del88, Del89, DE79, DL83, Egg83, Egg88, Ell82, Ell84, Fri81, Gea90, Ger89, GC88, Gui87, HLN83, IT80, Ioa81, Joe85, JS89, KN88, LP80, Lus79a, Mar78, Mie86, RB79, RY86, Sle82, Vog86, Whi86b, Wol83, YS89]. Integrals [AG68, AV93, DM93, ED77, For73, GA67, GW92, CM88b, Gab85, Gen86, Ger86]. Integrand [Lyn78a, Jac88]. integrand-approximation [Jac88]. integrands [Duf82]. Integrated [PK90, Sew66a]. Integration [And83a, BR70, BH71, BF75, Bog71, Bro77, BJ97, CM92, CSS97, CW76, Cas81b, Cas88, Col98, CP76, DR65, DM92, Die92, DS92, DGR96, Ehr98, ET71, EJ95c, For75, Fra71, Fri71, Gre91, Gun74, Hab77, Hes98b, JS92, JD93, Kea73, LLNS96, LW70, MR77, MC90, Nav65, NE77, NV88, NtB90, Rab67, RS84, Rei96, RD71b, Ric70, Ros67, SV96, SS99, Sec65b, SP92, Str67a, Str67b, SG68, Str69a, Str69b, Str70, SB77b, VAF98, Wer75, YKS78, BE88, Dag86, DLN83, Esp87, GM83, GM78c, KD83, KP85, MS83, Sas78, SS82, SK87, SS80]. Integration-to-Blowup [NE77]. Integrators [DRV94, MQT98, Rei99]. Integro [Fai94, FS74, For65, For66a, HKL77, IT91, Lin74, SS88, BMS79, McK79, ST86]. Integro-Differential [Fai94, FS74, For65, For66a, HKL77, IT91, Lin74, SS88, BMS79, McK79, ST86]. Integrodifferential [Jok96, LT89a, Lin90, LTW91, Mar90a, PP96, YF92, Hu96]. intended [AE84]. intensity [BvA94, Wig87]. Interdiffusion [FL92]. Interface [CG98, Car96, CS90, GW97, ILT69, LL94, Lee93, Li98, WB98, FQZ88, Hof85, LL95]. Interfaces [Ber87, MDH^{+}99, NV97, LMS87]. Interference [DM88]. Interior [BS83, Shr73, Tho68a, Whe78, Arn82, Bak84]. Intermediate [JD93, Lip73, Bea87]. Interpolants [Bub78, PT97, Ver93, BZ88, Cuy87, Del93, Her85]. Interpolated [Sco75]. Interpolates [KR72]. Interpolating [Dur77, Luc74, Nie74, RJ74, Sew66a, FP86, Luc82, LS73]. Interpolation [Alf85, Atk68, AS69, Aus71, BW68, BW69, BS76, BZ85, BH70, Bru78, BP80, Bus85, CF85, CJMT74, CM74, CJ75, Cle90, CF90, Dem76, DGR96, DLR82, EJL85, Eps76, FC80, Gor71, GW74b, Gre64, Hal73, Hau71, Her67, Hes98a, Huy93, Ise81b, Jam77, Jia86, LS97a, Lin72, LMST69, LZ71, MS73c, Nic72, Nic73b, Nie80, PR77, Pra70, PV79, RD71a, Rip92, Rob79, SSA84, Sch83b, Sec65a, Sha85a, SP68, Sni72, Ste79a, SCWM71, Str72, Str75b, ST92, SV73, Vér90b, Wal98, Wei74a, Wer75, Wix78, Wuy74, AD92, BBT82, Bea86, Ber89, CF89, DL83, GU88, Gün88, Lee85, PS92, SWC89]. Interpolations [CY77, Hak82]. Interpolative [Rob66]. Interpolatory [HJ99, Ker71, Küt84, RS97a, SP92, CLM89, Dag86, Gen86, SS82, Var81]. Interpretation [McC82, Mil97]. Intersection [Ale87]. Interval [Coh71, Cor83, CR97, Gay82, GS96, Han65a, HS67, Han92, Jac75, Kea90, KX94, Kee86, KN85, MQ82b, Neu87, NK97, Oli74, RC66, Roh93, Sha95, AR81, Ale84, Sch86, Sch89b]. Interval-Arithmetic [Cor83, Ale84]. Intervals [Faz96, Gus69, KU71, CK85a, Cha85b, LK80, Saa83]. Introduced [IPS91]. Introducing [Hof76]. Introduction [Lan64b]. Invariant [Cli68, DL95, Eis67a, FD91, Gol77, GM78c, KL82, LVQ95, Moo96, Rou71, Ste71c, TKJ94, DH79, Jer84, Jer88, Mey81b]. Inverse [Ban82, CfXmZ99, CC96b, Chu92, Cli65, DR83, Eri80, GBIY71, GK65, Hoh99, Nob66, Roh93, SS74, Sym80, Sym81, YB96, Zha95, BBT82, Bub86, FNO87, Neu85, Szy88]. Inverses [BI66, BIC66, Cli68, Dem77, GP73, MT68a, RSK76, Mee80]. Inversion [Che93, Cop90, Hug98, Lau89, MG66, ML69, Obi90, Sch98b, Spi66, Tew67, Bub85]. Inversions [EL72]. Involving [Fin71, GBIY71, HK91, Len86, Pas82]. Irreducible [HP68a, Bun81]. Irregular [BDGG71, BD74, Gir74, MW92, Moo99, Olv65, OS65, GL78b, Geo80, GO83, May84]. Isentropic [ZH91]. isola [JP92]. Isolated [BHZ68, Kel81]. isoparametric [Len86]. Isospectral [CN88]. Issues [Die92]. Itô [Pet98a]. Iterated [Bru84, Cas88, CX98, DF91, KX96, LSX90, Slo76, van84, AP87, Egg83, SB88]. Iterated-Collocation [LSX90]. iterates [BPTZ92]. Iteration [Ban82, BS89b, Cor83, DAB99, DL91, EW91, Gel81, JT70, Kel65, Str74, VP93, Web85, Wim70, Woz75, ALY88, DDE79, Gar81, Mur82, Neu85, Nie79, Rei88, Rus85, Szy88]. Iterations [BJZ94, BKP94, FL71, GGK98, Jam73, KT76, Mor72, OR67, PP73a, Sha66, Ste75b, Ste74, Cha78b, Nel86]. Iterative [AMW99, AK84, Bas82, BQ76, BB87, BIC66, BP74, BB71, BW86, Chr91, DV97, DSW94, DL83, ELV89, EES83, Erd65, ELPV93, GBIY71, Gun65, HLY80, JK68, Kin74, KS94, Kov70, KMR94, KK93, Li98, Man81, May87, Med98, MP77, MJ77b, Nic73a, OR66b, Ort73, Ost73, PS80, PW96, PRS90, Per67b, Rat95, Rhe68, Saa83, Sch73, Ser99, SA73, She78, SW94b, SS74, Sto68, Sun96, VK90, Van72, Voi71a, Voi71b, WS93, Wei69, Woo71, Woz74a, Xu92, Zha95, de 78b, de 68, Ban81, BEL89, FQZ88, Ger89, Gre84, Han93b, Jer80, Man82, ME81, Moo78, PSP89, Par87a, PS82, PS85, Str84, Wan93b, You72b, Zla82]. Ito [KPH95]. IV [EJ95b, HR90, SS96b]. IVPs [Jac88].

J [BCL89a, But65, CGT89, Cra78, DF76a, DW82, GN83a, Gup85b, Hic97, HL82, Hil78, HEFS74, Jel79, Laz82, LL95, MR85, Ost65, Pie71, Sob79a, Sym81, Wei95, de 78a]. J. [Sew66b]. Jacobi [Cor96, DL91, Fer86b, Hua75, JX98, MG66, OS91, Por69, Por71, PH73, SS80]. Jacobi-Type [Hua75]. Jacobian [CM83, DK85]. January [Hou73]. Ji [Pai84]. Johnson [Med98]. Joseph [Wid66]. Julia [DAB99].

Kantorovich [Den69, GT74b, Ral80]. Karmarkar [Fra87, GM89]. keeping [LT88]. Keller [MN84]. Kernel [AE97, BN74, Fer73, Hil76, Hil78, Hud68, Yan94, Egg83, Tru86, YS89]. Kernel-Based [AE97]. Kernels [Ans67, CY92, DW74, Lin69, Ric78, Sle68, Bru83, RY86, Sle82]. Kind [AE94, Atk67, Bas73a, CS96, CX98, DW76a, Gar75, Han71, Her76, Hil76, Hil78, HMT75, HM76b, dHW73, LE97, LSX90, Lin77b, NdF74, PCW95, Ric78, SW68, Str74, Wil73, Yan94, vSS97, BRD81, Egg83, GC88, HLN83, HL82, Joe85, JS89, LP80, Mar78, Vog86, Wol83, YS89]. Kinetic [CP91b, EV96, GPT97, HV96, HL98, JPT98, Kla99, MA81, SZ99, Vas99]. Kinetics [BK98, DVW94]. Klein [Dun97, LVQ95]. Knot [Jia86]. Knots [HY95, Jup78, Sch68b, Far85, Mey87b]. Korovkin [Cap99]. Korovkin-Based [Cap99]. Korteweg [CK85a]. Krawczyk [AP83, Sch85, Wol80]. Krawczyk-Like [AP83, Sch85]. Kreiss [SW88]. Kronecker [Ste68]. Kronrod [Ehr98]. Krylov [BSW95, HL97, JK94, Saa92, de 99]. Kublanovskaya [JS83]. Kutta [OS92, vS89a, Alb87, Alb96, Ale77, AC90, Ale94, AP91, BP96, Bar99, BZ88, BJZ94, BT97, Bic77, BH93, BB79, Bur82, Bur87, BP90, BCM94, But75, BC90, Cas66, Cas69, CS74a, CK92, CV72, Enr89, Enr93, FSU85a, FSU85b, Gup85a, Gup85b, HW81, HL84a, HW97a, Hig89a, Hor83, Hos95, HS94, HS87, JRF91, JT95, JEH78, JN95, Jay96, Kee90a, KL67, Law66, Law67a, Law67b, LNS95, Nak91, PTP96, PT97, Pet86, Roc89, SSA91, SW94a, Sha85a, SV94, SO97, Ver78, Ver79, Ver90a, Ver91, Ver93, in 96, van79, vS87, vS89b].

L [Sew66b]. Lacunary [Dem76, MS73c, PV79, SV73]. Ladyzhenskaya [DG90]. Lag [BP96, vS89b]. Lagged [CM99a]. Lagrange [BK98, CY77, Gün88, Kar82, LQ84, LK74, Lee93, MTF99, Nic72, Nic73b, RR95, Vér90b, Woh99]. Lagrangian [CGT91, EW96, FF98, Mun94, Riv75]. lags [Tav78]. Laguerre [Fos81, Kah67, Mas73, SC72]. Laminar [AP95a, AP73]. Lanczos [Saa80, Saa82, Wid78, Nas84]. Landau [DG98a, BC99, DNW98, DW99, Lor97, MH98]. Landau-de [DG98a]. Landweber [Str74]. Laplace [Dou93, Atk82, BvA94, BK99, CD77, Che93, Cop90, MG66, Obi90, RHH99, Spi66, Wim77, WH97]. Laplace- [WH97]. Laplacian [Kut74, ALY88, BL94]. Large [Dav97, Dax85, DM92, EL73, Gov97, Hof76, JK94, Jen72, Law67a, Lev82, Rak76, BSW95, CFT84, Lev85, PSP89, Saa82, Ste83, WW93a, WW93b]. Large-Scale [Dav97, PSP89]. Latent [Kub70, JS83]. Lattice [DS92, ELR95, Hic96, JS92, JD93, SK87, Hic97]. Lattices [CY77]. Laurent [JS68a]. Laws [BCV95, BG98b, CM90b, Coc89, Coc90a, Coc90b, CCL95, CG96, CGY98, CF93, CP98, Cor96, GV91, GLBN97, Har90, JY98, JLL^{+}98, JX98, KR94b, LTW96, Lav91, LTW95a, LMS97, LW98, Liu93, Ma98a, Ma98b, MKT93, NT92, NT94, Şab97, Sch90b, STW97, SB93, TT99, TT95, Tan98, Ten92, Ten94, TW93, Vas99, Bre84, BO88, CT86, HL81, Lav89, LeV88, Lev82, Lev85, LB96a, Luc85b, MT89a, San85, Swe84, Tad89, Vil86]. Lawson [WGS90]. Lax [MR85, Pet65, Gar66, MR84]. Lax-Milgram [Pet65]. Layer [AP73, Gue92, LP91, Rat95, Smi81]. Layers [JT97, LP77, WL95]. leading [Neu88]. Least [AL91, BV92, BP74, Boc97, BCMM98, BMM99, BW90b, BN73, BT68, CLMM94, CMM97c, CMM97a, CMMP98, CG98, Car74, CNP93, CG90, CN97, CC96b, CD90, Cli73, DW81, DW82, DW85, DW93, DS68, Eld80, Erd65, GP73, GV80, Ip87, Kah67, Kam79b, KP78, Kin72, Lan77, Liu94, MR98, Mor68, Osb65, Osb75, PS77, PCL94, Per67b, Pow69, Saa87, Sch65, Sch66, SM81, Sta99, Sto71, WW90, Wei92b, Asc78, Bar88, Cor81, Dem87, DS86, FS91, GN83a, GN83b, GM78b, LP78, LP80, Mar83, Pai79, PSP89, Van85]. Least-Change [BW90b, DW81, DW82, DW93, Ip87, WW90, DW85]. Least-Square [Erd65]. Least-Squares [Boc97, BCMM98, BMM99, BT68, CG98, CN97, Kin72, Liu94, MR98, PCL94, SM81, Sta99, Sto71, Cor81, DS86, GM78b, PSP89, Van85]. Lebesgue [Bru78, Gün80, Vér90b]. Leg [Dah83]. Legendre [BYLP98, Can86, Cor96, DG94, Let71a, Ma98a, Ma98b, MKT93, PR95, QZ92, ST95]. Lemma [Gre99, Pet65]. Lemmas [Vas99]. Length [Gol77]. Leonard [Wid66]. Level [AL96b, Bra94, CDK95, CO87, DFJ74, GM66, KL89, LT98, LO90, Mal98, McC94, MS99c, Rac73, Var80, Yse86, Dai92, JS87, Vog86, Zha96]. Level-Index [CO87, LO90]. LeVeque [Luc85a]. Levinson [KW93]. Liapunov [GGM99]. Liftings [MS97b]. Like [AP83, Bro69, CM74, DAB99, Gar78, BR80, FD98, FM90b, Gri87, Grz85, Pet89, Sch85, Vat88]. Limit [Bak71, BJZ94, FGH74, Gus73, Kla98, Kla99]. limiters [Swe84]. Limits [CJ76, Doo66]. Line [Atk69, AV93, Cop90, PT91, PS80, VS97, YR99, BBT82, BPTZ92, DS89b, GLL86, Naz86]. Linear [AN66, AK84, ADH99, AV99, Asc77, AL96a, ABDH95, Aub68, Bab72, BDR92, BR73, BR78b, BCC78, BCS78, Bec92, BP74, BF99, BG74, BJS73, BH70, Bre93, Bro95, Buc77, BP71, BC72, Cad73, Cad74, CMM97a, CMM97b, CMMP98, Car99, Cas82, CM74, CS75, Cha73, Che87, Che73, Che75, Chr91, Cli73, Cli76, Con76, Coo72, Coo78, Coo81, Dan67, Dan73, Dan75, Dur88a, DS68, EW72, EES83, Eld80, EJ91, FGH74, For66b, FR78, Gar83, Gay82, Gek76, Geo74, Gre66, Gri74, HMM98, HP78a, Han92, Hel76, Hes98b, Hie80, HMT75, HM76b, Jer73a, Jer73b, JP98, Jok96, KHK93, KN72, KS97a, Kel65, Kel69, Ker96, Kří92, KK91, KM96, Kus68, LQ84, LC98, Lee98, LB90, Lin74, Mah82a, Man81, Man71, MG95, Mat82, MP77, MM73, Mid66]. Linear [MS99b, Nas70, NdF74, Neu97, Neu87, Nic73a, Nic92a, NK97, OPW65, Oet65, Osb65, Pai74, PS75, PS77, Phi72, Pot93, Rei72, Rei81, Rip92, Rob75, RS92, RR98, Ros70a, Rou71, SZ99, Sch72, SW77, Sha95, SWW97, SF79, Sta64, Sta91, Ste68, Ste71c, TE98, TXZ89, TK96, Van72, Var73, Var76, Vel92, Wah77, Wal98, Wid78, Xu96, Zed66, dK86, AS85, AG87, BM87, BCL89a, BCL89b, Ber82, Boj84, Bra85, BDGB83, Cas81a, Cha78b, CL92, CR79, DM82, DOR88a, DOR88b, Dob80, Eva80, FKD84, Fer86a, Fra87, GL81, GC88, Gre84, Kin79, Man82, MM79, Mey87a, MK80, Moo94, Nel86, Pai79, Par87a, Rei88, RY86, Saa83, Saa87, SSY88, Tru84, Var78, Vat88, Vog86, Wu96a, You72b, Zho83, Zla82]. Linearly [Rit75, CH95, SWC89, Zir82]. Lines [ELW91, KS92a, Tho76, Mey81a, Mey81b, MP85, RK88]. Liouville [GB89, GM98, Joh69, Pru73, PFX95]. Lipschitz [BO88, Kee90b, Law67a, Sch74, Sew66a]. Lipschitzian [BK98, Gri87]. Liquid [Alo97, DG98a, LL89]. Local [BG98a, Bro87, CS98, Dem77, Dun71, FR92, FR85a, Fon88, JP96, KR72, Mil92, MF90, PW78, Pow69, RV74, Rom90, Ste75b, Ste71b, SH95, Tad91, Tra96, Ver78, Yam98, Ypm84, BX89, CC84, Gri87, LS73, Zen85]. localization [Bas82]. Localized [EW96, LT88]. Locally [AMT99, Cal76, ELPV93, GM69b, SSW96]. Locating [Li83a]. Location [Row69, CG84]. Locational [KC74]. Locking [BS92, Boi97]. Logarithmic [GS88, SF79, Str75a, Yan94, YS89]. Logarithmically [Osa90, DGB82]. logarithms [Mie81]. Long [EJ95c, Gus69, RR96, Rip92, Lar89]. Long-Time [EJ95c, Lar89]. Lotka [Wal83]. Low [Cal71, CV99, GC95, Kla99, Str66, BN85]. Lower [Alf96]. lowers [Dem87]. Lowest [Bre92, Mar85]. Lowest-Order [Bre92]. LP [HM85]. lubrication [Mey81a]. Lumped [DP98]. Lyapunov [DRV97, JK94, RW95, Tal91]. Lyness [Li83a].

M [de 78a, GL93]. m-Dependent [GL93]. MAC [HW98, Nic92a]. Mach [GC95, Kla99]. Maclaurin [Nav65]. Magnetic [Hei94, Fri81]. Magnetohydrodynamics [van95]. Magnetostatic [Per97, Rog93]. Magnus [Zan99]. Majorization [Ros80]. Making [LM86]. Mandelbrot [DAB99].

Mandelbrot-Like [DAB99]. manifold [AS85, FR86]. Manifolds [DR90, FD91, GANT98, HR93, Lor97, Ma95, TKJ94, Yen93, Rhe82]. Many [TK96]. Mapping [Alo97, DeL94, Dri99, Weg86, Rei88, Tru86]. Mappings [Ben95, Che98, Nit89]. Maps [BK97c, Kee86]. Maratos [BPTZ92, PT91]. March [Osb79]. Marching [BR77, Ban77, Buz77]. Markov [BD99, Hav87]. Marti [HL82, HL82]. Mass [CN97, Mil92, CEv87]. Match [Coh71]. Matching [Cim72]. Materials [Dob92]. Math [But65, Ost65]. Mathematical [CC66, CC67, For65, For66a, For66b, Fri81, GH77, JPRT96, Lav91, Rog93, Sew66b, Lav89]. Matrices [ABP93, BD90, BW90b, Buc66, Bun74b, Cap99, CY93, Chu90, wECGS94, Cli65, Dem77, Des72, EH75, Eri80, FU69, GL75, Gor75, Kee70, Kee71, Kub70, LZC92, LS76, Loe70, MS73b, MS65b, Pai74, PT79, Rak76, RSK76, RS97b, SV70, Sne96, Var79, Var70, Wel97, Wer96, Wil73, Zha95, de 80, CM88a, CM83, Dix83, EGLP85, GZ88, JS83, KPJ82, Mee80, Mel87, PR86a, Sha85b, Ste80, VSM80, WT88]. Matrix [AC96b, BQ76, BB71, Bra66b, BH86, Cap99, CGO07, CN88, CD90, Cli68, CMSW79, Dej67, DO92, DTW76, DTW78, GBIY71, Geo76, GPS76, GK65, GV74, GS67, Hal87b, HP68b, HP68a, Han65a, HS67, HYS94, HL97, Hua75, Jam73, Joh78, Kam78, LS72, Van77, MG72, Mar90b, Mea69, MP77, Mil92, Min70, MS73d, Mor72, Nob66, Pel67, Roh93, Saa92, Sch90a, SA73, Ste77, SW88, Tew67, Van75, War77, Bun81, DK81, Dem83, Dry83, Fun87, GHY83, GL78a, JA83]. Matrix-Free [BH86]. Matrix-Vector [HYS94]. Max [GG90]. Maxima [Pie69, Pie71]. Maximal [DS92, HP68b, HP68a, Min70, Woz74a, Woz75]. Maximum [BJZ94, BT97, Bre75, EL93, Gre99, Hud68, Kří92, Liu93, LO96, Pal96, Reu92, Reu94, Shu72, Var66, Wah74, Wah75, in 96, Bas82, Bas85, RR86]. Maximum-Norm [Bre75]. Maxwell [GS91a, GM97, Hip99, Mon91, Mon92]. McKee [LS76]. Mean [BG95, BS87b, CDK95, Dun75, Sew66a, Wal96]. Mean-Square [Dun75]. Mean-Value [BS87b]. Means [Cos71, ML69, PT91, BPTZ92, Hac79]. Measure [Pry84, Bra80]. Measured [Far66]. Measurements [BS87b]. Mechanical [Eic93, Mil97]. Mechanics [DDLT91]. Media [AWZ96, BK97a, BK97b, CSH78, CJ96, AVG89, Dou83, Dur88b, EW80, ER82, EYL89, Rus85, Sam86, Vic80]. medium [BSW85]. Membrane [BH68]. Membranes [Don69]. Memory [SWW97, Ano73, CL90]. Merit [LS97b]. Mesh [AH94, BS97c, BL98, CD85, Cho78, DP92, GT74a, Geo73, HRR94, HR97, JN98, Köh95, LL96b, Osh70, Rhe80a, Riv84, RC78, TH98, Wat86, YS89, ABPR86, AB87, CK85b, Jam80]. mesh-independence [ABPR86]. Mesh-Independent [Wat86]. Mesh-Refinement [Rhe80a]. Meshes [Abg98, Ain96, AMT99, Güs81a, JV96, JNV99, MW92, SSW96, Sül91, BS89a, Whi79]. Method [AD68, Abl72, AP73, Afr70, AGM92, AE97, AW95, AWZ96, Arc77, AR96, Asa89, Asa98, AC93, Atk74, AV97, AL96b, AW80, AL91, Bab71, BZ73, BK75, BA76, BSK81, BS87a, BS92, Bad91, BS97b, BGT97, BR75, BR87, BK98, BCC78, BDH98, BS87b, BHL96, Ben95, Ben96, Ben92, BJ93, BL92, BF99, Boi97, BJS73, BNP91, Boy77, BK77, BN73, Bra66b, Bre92, Bre93, Bro69, Bun71, CJ76, CL89, CM90a, CMM91, CH68, Car92, CM99a, CfXmZ99, CG90, CN97, CC78, Cha69, Che93, CC94, CE97b, CK92, CW97, Cho78, CK98, CD90, Cia75, Cli73, Cli76, CCL95, CS98, Col98, CP77, Cop90, Cor83, CGH91, CT72, Dan67, Dan70, DM77, DP98, Daw98, Dk99, DK80a, DK80b, Den77a, Den77c, Den67a, Den67b, Den69, DR83, DM98, DF76b]. Method [Dia77, Día79, DL95, Dob92, DV97, DG94, Dör99b, Dou93, DR82, Dri99, DER76, Dun74b, DGW91, Eic93, EH92, Eis74, Eis83, EZB83, EH89, EL97, FK76, Fal78, FR92, FFLMW99, FF91, Fin71, For75, For90, FMMR99, GGS94, GGS98, GANT98, GBIY71, GGM99, Gay79, GOQ97, GG90, Gir74, Gir78, GS91a, GJL99, GC96, GS76, GW92, GM91, GM93, GC95, GL96b, GC98, Gus73, Hag75, Hag85, HR95, HW98, Han71, Har90, Hed79, Heg92, Hei98a, Hei96, Hei94, HR93, Hip99, Hof76, Hoh99, Hou90, HLK91, Hua75, HT82, HS92, Iss96, ITM91, IK93, IR99, JK74, Jeo94, JN98, JNV99, JS98, JW91, Jim96, JS68b, Joh77, JPRT96, Kah67, KN72, KJ98, KM99, KP78, Kea90, KL96, KL97, Ker96, KP96, KLM94, KO79, KS92a, KL89, LT98]. Method [LL94, Lee91, Lee93, Li83a, LB97, LSX90, LTW95a, LTW95b, Lin97, Lin74, Löt92, Lue70, Ma98a, Ma98b, MKT93, Mad71, MB97, MTF99, MX95, MF98, Mas90, Mat98, Med98, MS91, MS65a, Mey71, MW94, MTTW99, Mon91, Mon92, MF90, Mor95, MH98, MS99c, MS97b, Nel73, Nic81, Nob66, Ols77, Omo77, OR66a, Ort69, Osa90, Pan98, Par99, PW96, Pet91, Pet92a, Pet87, Phi72, PV72, Pie69, Pie71, Pla98, Pol74, PL87, Pre73, PFX95, Ral74, RK88, RR96, Rat95, Red78, Ric91, Rin90, Rit75, SD77, SS88, Sco75, Scu98, She95, SK93, Shu72, Sim75, Sle68, Slo76, SBBL91, SS74, SW91, Swa74, Tap69, Tap74a, Tap74b, Ten92, TYZ95, hTZ97, Tew67, Tha78, Tho76, TV96, Tra96, Tra77, Urb91, Van67, VMSB97, VS97, VIA96, VWI97, Was82]. Method [Weg66, Weg86, Wet97b, Whe75, Whe77, Whe78, WM67, Wid78, WB98, WH97, Yam98, ZB96, ZWW94, Zla65, Zla73, Zla74, dT93, dG81, AF86, Ale84, And83b, AM89, Arn82, AF89, ADK82, BI89, BM87, BG88, BCMP91, Bak80, BSW85, Ber86, BR85, Bia89, Bog85, BL87, BH83, Bre89, Cha85a, CK85b, CC84, CL92, DRW89, DK81, DO79, DK82, DKK83, DK85, Des81, DL82, DH79, DOR88a, DOR88b, Dor84, Dor86, DDE79, Dry83, Dur88b, EL96, Ell82, Ell84, EF89, EHP87, Eri86, EL89, ED82, FM84, FKD84, FR87a, For79, Fos81, Fre87, GHY83, Gar84, Gar81, Gat88, GL81, Gou89, GGL79, Gre84, GO81, GO83, GLL86, GG86, GHS96, GT82, Hac79, Hac81, HL82, IPZ79, Ioa81, JA83, KD79, Kar82, Kar78, KW85, KS83, KS87]. method [KN88, Kin79, KHR89, Kum88, LL83, LeV88, LL95, LT88, Lev85, Li83b, LB96a, Lus79b, LR82, MT89a, Mar85, Mar79b, Mey81a, Mey81b, Mit86, Mon86, Mon87, Mos83, Nas84, NV83, OS92, Par87b, Pas79, PW80, Pet81, Rau85, Rei88, RS89, RBC79, San85, Sch71b, SS96a, Ser83, ST95, Sma88, Sor82, Ste83, Str84, SO86, Sty89, Sun79, SB84, TW88, Tre83, TG83, Tru84, Tru86, Van85, Wah80, WL95, Win80, Win82, Wom89, Zen85, Zho83]. Mèthode [Cia75]. Methods [AP95a, AMW99, AT86, ADZ96, Akr99, AK84, Alb87, Ale77, AC90, Ale94, ABGM92, AG85, AMSS97, AOW93, AD78, AR95, AE94, AP91, ARW95, AL96a, AMS90, AC98, Avi74, AC79, AV90, BCO94, BR70, Bak76, BP96, BS93a, Bar70, Bar99, BF75, BQ76, BB87, Bel97, BJZ94, BT97, BCV95, BP74, Bic77, Boc97, Bof97, BW90a, BW93, BDLN92, BV90, Bra66a, BP89, BC95b, Bre98, BF91, BH86, Bro87, Bru90, BP71, BM97, Bur82, BP90, BS94, BCM94, But75, But81, BC90, BJ97, BGN70, Byr90, CW96, CMM97b, CDS99, Cal71, CF88, Car74, CS96, CV99, Cas97, CS72, CS74a, CS75, Cha73, Che98, CX98, CM90b, CKV98, Chr91, Chu92, CG96, CGY98, CG73, Coo72, CV72, Coo78, Coo81, CS90, CP76, CW78, Cry74, CR97]. Methods [Dah83, Daw91, Daw93, DVW94, DW76a, DeL94, DF91, DES82, Den74b, Den74a, DF75, DF76a, Den77b, DW81, DW82, DW93, DL90, DP92, DM92, DRS95, DG94, DH71, DM79, DGP66, DD70, DR82, Dra94, Dry78, DW76b, Dup73b, Dup73a, DFJ74, Dup76, Dut90, E92, ES76, EN83, ELV89, Ein88, EH92, EES83, EC93, ELR95, Enr74, EH76, EH98, EJ91, EJ95c, EJL98, ELW91, EW96, FM87, FM67, Fai72, Fai94, FM90a, FR99, FS74, Fon90, GLV91, Gar75, Gea65, GT74a, GW74a, Gea78, GP84, Gek76, Geo77, Gol93, Gol94, GP98, Gor71, GG76, GGOS96, GM69a, GM69b, GYF86, GW90, GL93, GH92, Haa96, HLY80, Hag90, HW81, HL84a, HW97a, HKL77, Hav87, HV96, HV98, Her75, Hil97, HV97, HMT75, HM76b, dHW73, DW74, DW78]. Methods [HK94, HVR88, HZ94, Hud68, HEFS72, HEFS74, HS94, Huy95, HS99, IT91, Jac84, Jac86, JRF91, JT95, JK96, JN95, Jaf92, JK94, JP67, Jan79, JV97, Jay96, Jel77, Jel79, Jen76, Jes78, Joh69, KP87, KX96, KX94, KG91, Kel69, Kel74, KW75, KT86, Kin74, Kin72, Kin73, Kin75, KPH95, KS94, Kov70, KMR94, KM93, KK81b, KK93, KS92b, KT97, KM96, KT92, Kut74, KC93, LS72, Lam81, LL96a, Lay83b, LMV96, Lee98, LPG91, Ler75, Li87, Li91, LFB98, LL89, Lin90, Lin77a, LW70, Lin69, Lov96, LR72, MT89b, MR85, MF97, Mar79a, Mar90b, Mar92, MMD94, Mar94, Mas73, May87, McC82, MR82, McC94, Mey73b, MP77, MT97, MPS98, MJ77b, Moo94, Moo99, Mor71, Mor98, Mur97, Nak91, NP81, Ned78, NK97, Nit89, O’S90]. Methods [OPF97, OR66b, OR67, OR72, OR93, Osw92, Pal96, PP96, Pao87, PT97, Par95, Par65, PRS90, Per67b, Pet86, PP73b, PO87, Por69, Por71, Pot93, Rac68a, Red76, RS95, Rei96, Reu92, Reu94, RV74, Rhe78, Rüd93, RS75, Şab97, San98, Sar90, SS93, SS95, SV96, Say74, SSW96, Sch68a, SW94a, SZ99, SF84, Sch90b, Sch96, Sch69e, SS89, SM81, SA73, Sha69, Sha85a, Sha70, She92, She78, Sim71, Sin77, Ske76, SJ77, Sta99, Sti64, Str74, Str65a, de 99, SW74a, SW74b, TE98, TE86, TE89, TS93, TT95, Tan98, Tau96, Tav71, Ten94, TW74, TW75, TK96, TT87, Tur77, VK90, Van72, VP93, Var71b, VW95, Ver78, Ver79, Ver93, VTG91, VA91, Wah92, Wal99, WW90, Wal73, Wat77, Wei93, Wet97a, Whe73b, Wim77, Woh99]. Methods [Woz74a, Xu92, Yam75, YF92, YZ94, Yin95, Ypm84, ZZ98, dR81, in 96, van79, van80, van84, vS87, vS89a, vS89b, AVG89, AS84, AP87, BO83, BO87, Bak82, Bak84, BR80, Ban81, BCL89a, BCL89b, Bas82, BZ88, BW86, BEL89, BH93, BQ86, Bru84, Bub85, Bub86, Bun81, BB79, Bur87, BC89, BSW95, BNY87, BN89, CLM87, CLM89, CL90, Can86, Cas81a, CK85a, CT86, CH95, Cha94, Cha85b, CHM88, CR84, CL84, DLN83, DW85, DH79, Doe79, DR84, Dou83, DHLR84, Dyk87, Egg88, ES86, Enr89, Epp84, EJ95a, EJ95b, Ewi78, EW80, ER82, FN84, FR85b, FST87, Fon88, Fos81, FS91, FSU85a, FSU85b, FNO87, FM90b, FQZ88, GV89, Gen84, Ger89, GT85, GGT82, GLT87a, GLT87b, GM80, GC88, GZ88, Gri87, GM78c, Grz85]. methods [Gui87, HLN83, Hal79, HM85, Hal87a, HP80, Han93b, HL87, Hay81a, Hof78, HS87, IT80, Ise84, Joe85, Joh88, JS89, KW84, KP79, Kre78a, KNB86, Kun88, KM88, LM78, Lay83a, LOR85, LMS87, LP78, Luc85a, MM84, MMR88, Man82, MR84, McC84, McC85, McK79, MS83, Mie86, ME81, MG82, Moo78, Nie79, Pan85, PS82, PS85, Pas82, PR86b, Qua87, Rhe80b, RY86, Roc89, Rom90, Rou82, RS87, Saa80, Saa82, Saa83, Sam83, SDC88, SSY88, Sch80, SE81, Sch86, SJ83, SK87, SWC89, Tad86, Tad89, TK84, Var78, Vat88, Ver82, VG86, Wan93b, WH86, Whi86b, Wol83, Wu96a, You72b, Zha96, van81b]. MGR [Par87b]. MHD [MS99a]. Micromagnetics [LM92]. Microscopic [CLNP89]. Microstructure [LL98, LM92]. Microstructures [Col98]. Midpoint [de 86]. Mild [LP77]. Mildly [Ros73, Sim71]. Milestones [CGO07]. Milgram [Pet65]. Mimetic [HS99]. Min [GG90]. Min-Max [GG90]. Mindlin [AF89, DGW91, GN91, Lov96]. Mini [Ver88]. Mini-Element [Ver88]. Minimal [BP80, Fer74, Fra73, RD71a, RD71b, Ric70, Var70, GL80b, Sid86, TD82]. Minimax [CC78, DC80, FGH74, Mas73, Ged78]. Minimization [BG71, BCS78, Ber66, Bog77, Buc76, DG98a, Jer73a, Jer73b, Lue70, OR72, Sha70, BN89, EAVD79, Gou89, KD79, Nas84, Naz86, SW82, SSB85, Toi79, Var85]. Minimizing [Sim68]. Minimum [AK98, Cad73, Cad74, CR76, DS72, FP86, GM78a, Her76, KS97a, Nie80, SW68, Tre72, GL80b, Mar78]. Minimum-RMS [Her76, SW68]. miscible [BSW85, Dou85, Dur88b, EW80, ER82, Rus85, Sam86]. MITC [SS97]. Mixed [AW95, AWZ96, AWY97, Bre92, Bre93, CKV98, CDW96, Daw91, Daw93, Daw98, Dob92, ELW91, ELW99, Fal78, FF93, GV91, GW97, Gir78, HR95, HW98, Her68, HW97b, Joh77, Lee98, Lov96, LL99, MW94, Mon91, OR76, Pan98, Par95, PCL94, Per97, Sch96, SM81, Sta99, Str65a, Tar96, Var71a, Ver84, Ver88, Wig66, ZWW94, BO83, BvA94, BX89, Cle81, DO79, Dur88b, EYL89, FN80, Gus81b, KM88, Mar85, Mon87, PS88, PR82]. Mixed-Type [SM81]. MMPDEs [HRR94]. Mode [GGS94]. Model [BL92, Che95, DG90, DN91, DNW98, Dub90, DGW91, EJ91, ILM^{+}96, JK91, JN98, JNV99, KC93, LW81, Neu98, PSB97, Wal83, AM89, EL89, Lin88, SDC88, Sor82]. Modeling [Bou98, GH77, HR93]. Modelled [CG93]. Models [BS96, BC95a, BFGP99, COP98, Dax85, DW99, Boj84, BMP89, IKP91]. Moderate [WES80]. Modes [BDM^{+}95]. Modification [Liu93, Wol80, EGLP85, Sor82]. Modifications [Ree90, Ale84]. Modified [AD78, Coh71, Gar75, Hal87b, Qn99, Sha70, van81b, DRW89, Hay81b, KD79, KHR89, Mur82]. Modulus [Gre99, HY95]. Moler [Nic67]. Moment [Gus70, GKR70, SZ99]. Moments [Pri77]. Mono [BCM94]. Mono-Implicit [BCM94]. Monosplines [Sch65, Sch66, BL80]. Monotone [AO82, Bad91, Bea82a, CF85, CGY98, EJL85, FC80, HS97, Huy93, Jen92, Jok96, LB96b, OR67, PR77, RD79, Rou71, Şab97, Sch73, Sha66, Sha69, hTZ97, Wil73, ZZ98, CF89, KRD78]. Monotonic [Kam76, Kam79a, Kam79b, Kam81]. Monotonicity [AK90, BZ85, BW89, Van72]. Monte [CM92, LC98, MC90, MC93, Owe97, Ros67, Sas78, Spa71, YKS78]. Moore [Qi82, Ral80, SS74]. Morphological [CDK95]. Mortar [AP95a, AMW99, Bel97, CDS99, Woh99]. Motion [BP67, CDK95, Gir95, Jen76, Pot93, Wal96, Yen93]. Motions [BG95]. Motivate [DAB99]. motivated [Scr91]. Mountain [IZ81]. Moursund [Wuy70]. Movement [CC96a]. Moving [Bai91, BS93a, BTW93, HRR94, HR97, Jim96, MM81, Mil81b, Mil92, Mil97, TH98, Wat86, Wil69, AF86, Hig84, Lin88, MN84, San85]. MR [CGT89, Hic97, Jel79, Laz82, LL95, Sob79a, Sym81]. Multi [Bro77, DDS80, Dou84, Yse86, Hac79]. Multi-Derivative [Bro77]. Multi-Dimensional [DDS80]. Multi-grid [Dou84, Hac79]. Multi-Level [Yse86]. Multibody [Yen93]. Multicell [Dax85]. Multicolor [PO87]. Multiderivative [Jel77, Jel79, TK84, Cas81a]. Multidimensional [BB74, CCL95, CGY98, Got72, HJ99, JS92, JD93, Kea73, Mey73a, RS97a, Sas78, Sch69e, Sto68, VA91, GV89, KL79]. Multidimensions [Daw93, Dör99a]. Multifluid [BC95a]. Multigrid [AC93, BV90, BKP94, Bra94, Bre92, Bre93, Bre98, CMM97b, DD93, EL97, Gol93, Han93a, Hip99, Hop87, HZ94, JV96, KP87, Lee98, Löt92, MM84, McC82, MR82, McC84, McC85, NV93, Ott96a, PRS90, Reu92, Reu94, Rüd93, SB93, Wan92, Wan93a, Web85, Yav95, ZW94, BD85, BH83, Bre89, DS89a, Gar81, Gol89, Gre84, MMR88, Par87b]. multigroup [AVG89]. Multilevel [AV90, AK98, Ban82, BM97, Cap99, DD93, HK94, HW97b, Kel95, LL96b, McC94, RW95, Sta99, VW95, Ver84, Ver88, Ban81, Wan93b]. multiparameter [Mül84]. Multiphase [BJ91, CE97a, Wom89]. Multiple [Chu90, CP76, Fra71, GCB74, GGK97, Hon93, LSZ97, LT68, NV93, Ste74, SS65, Tse96, VF93, Wat75, BO87, Gen86, SK87]. Multiple-Exchange [Tse96]. Multiplication [HYS94, Ura68]. multiplications [CLNP89]. Multiplicative [CW93, KP85]. multiplicative-noise [KP85]. Multiplier [Hag85, Hag90, Lee93, Kar82]. Multipliers [FJ91, MTF99, Woh99]. Multiplying [Nel73]. Multipoint [Ste74, Mey87a]. Multipole [YR99]. Multiresolution [Abg98, AC96b, ADH99, CW96, HYS94, Har96]. Multiscale [JPT98, KLM94]. Multistage [BJ97, DGP66, Spa71]. Multistep [BF75, Bro77, BC72, CS75, Coo72, Cry74, Dah83, DH71, ES76, Enr74, GT74a, GW74a, Gek76, HMT75, HM76b, Jac86, Jel77, Jel79, LS72, MF97, Pot93, SZ90, TK96, VAF98, van80, AE84, BEL89, Bur87, Cas81a, CH95, CL84, Lin83, McK79, Rou82, Var78]. Multivalue [BS94, BC89]. Multivalued [Ise81a, Bre84]. Multivariate [Alf85, Alf96, BJS73, Cox84, Dah80, Far85, Gor71, Hak82, Höl82, Jan79, LLNS96, LK74, Nea92, Nie74, Sch69b, Sch69a, Sch69d, Sch69c, Wat75, Zwa73, de 78a, Cuy87, DM82, Gra87b, HHS78]. Multiwavelets [vSS97]. MUSCL [Osh85, Yan99]. Mysovskiı [DP92].

N [de 78a]. Nagumo [Jer80]. Nanbu

[BI89]. Natural

[Atk68, FF70, Hal73, JS68a, SS99, LS73]. naturally [Arb89]. Nature [Bak68]. Navier [AP95a, BR85, Boc97, BCMM98, BMM99, BC95b, BM97, CFT84, Dut88, EHP87, GR82, GM93, GL96b, GS91b, HP80, HR82, HR86, HR88, HR90, Hou90, HW92, HW93, IK93, Kar82, KJ98, KR94a, LL96b, LT98, Lin97, MQ82a, Osb76, She92, She95, TV96, Wee98, Wei93, Yan96, ZH91, ZW94, ZH94]. Near [AH94, CW76, Ged78, JD86, LSS94, LNS95, Mas73, Nav65, Bey87]. Near-Minimax [Mas73, Ged78]. Nearly [CMM97b, WGS90, Ban83, Cha84a]. necessarily [Dag86]. Necessary [FM84, Wil70]. Nedelec [Med98]. Negative [Geo76, SS93]. Nekrasov [CG93]. Nerve [Mas90]. Nessyahu [AV99]. Nested [DER76, Geo73, GPV78, HS68, LRT79, GL78b]. net [Owe97]. Nets [LLNS96, LC98]. Network [Por69]. Networks [Gir74, Wix78]. Neumann [BH65, Fun90, ILT69, MG95, Str75b, TMV98]. Neutral [AP95b, EH98, Jac81, Jac84, Jac86, Liu99, KK81a]. Neutron [Asa89, Asa98, Mad75, MR98, Pit78, Asa86, JP83, Vic80]. Newton [Nas84, Sch73, She78, AT86, Ale84, BR80, Bro69, BNY87, BN89, CM87, Cha85a, CfXmZ99, CC84, DK80a, DK80b, DK82, DKK83, DES82, Den67a, Den67b, Den69, DH79, DP92, FST87, FM90b, GT74b, Gra71, GO81, GO83, GLL86, Grz85, HM85, Hoh99, IT88, JS96, JA83, KD79, KS83, KS87, KN88, Mar90b, Mar92, Mor71, MT68a, Mur82, Nic81, OR66a, Pan85, Ral74, Red78, RS89, RM77, Sha70, Sor82, Tap69, Tap74a, Tap74b, TW88, Van67, VVC94, Weg66, Ypm84]. Newton-Iterative [She78]. Newton-Like [Bro69, BR80, FM90b, Grz85]. Newton-type [Nas84]. Newton/Finite [Bro87]. Newtonian [Pad97]. Nicolson [KW76, BF99, Jam80, MH98]. Nicolson- [KW76]. Nine [HP77, MST88]. Nine-Node [MST88]. no [BCL89a, CGT89, Cra78, DF76a, DW82, GN83a, Gup85b, Hic97, HL82, Hil78, HEFS74, Jel79, Laz82, LL95, MR85, Pie71, Sob79a, Sym81, Wei95, de 78a]. Nodal [Hes98a, Bak84]. Node [BTW93, CR90, MST88, Sch93]. Nodes [Ehr98, Moo99, SP92, SS65, ST92, Gün88, MP85]. Noise [MT97, KP85]. Noisy [Lan64a, Wah77, Rag85, Vog86]. Non [BK98, Chr91, CGY98, EH76, GKR70, LTW95b, Moo94, Pet65]. Non- [Pet65]. Non-Cartesian [CGY98]. Non-Chebyshevian [GKR70]. Non-linear [Moo94]. Non-Lipschitzian [BK98]. Non-Stiff [EH76]. Non-Strictly [LTW95b]. Nonclassical [HL98, Pri79]. noncompact [Mil79a, Mil79b]. Nonconforming [BZ73, Bel97, BV90, Bre93, DM98, Gol94, JPRT96, KJ98, Li91, MDH^{+}99, Pad97, VW95, Zha97, Bre89, EF89, Li83b]. Nonconvergence [Boy77]. Nonconvex [Bur90, Gre94, MW95]. Nondegenerate [Zwa73]. Nondifferentiable [Che98, Con76]. Nondifferentiale [Con73]. Nonequilibrium [BK97a, DVW94]. Nonhomogeneous [CJ96]. Nonideal [MS99a]. nonimpulsive [Bub86]. Noninterpolatory [EH72]. Nonisentropic [ZH94]. nonisolated [Kel81]. Noniterative [DO92]. Nonlinear [AGP94, ALM92, AWZ96, AG68, Avi74, AC79, AL96b, AK98, Bad91, Bar70, Bog71, Boh77, BJMT73, BW76, Bre73, Bru90, Bry68, CW96, Car99, CG97, Cas88, CLMC92, Che77, Dan67, Daw98, DWW98, DF91, Den75, Den77b, Den77c, Den67a, Den67b, DR83, DM92, Don94, Dou93, Dun97, Dun75, EJ95b, Esp94, FX95, FFLMW99, FS74, Fin71, Fon90, FT74, GPT97, Gai74, GGS94, Gar75, Gay81, Gay83, GP73, Gov97, GS76, GA67, GW90, GL93, GSY92, HP75, Hag90, Hei94, Hen69, HR93, IS90, IR99, Jam71, Jan94, Jer73a, Jer73b, KHK93, KAD93, KM99, KP78, Kee90b, Kel74, Kin73, KKP94, KT92, Kus68, Lam67, Le 94, LP77, Li87, Li88, LW92, LVQ95, Lin90, LTW95a, Mar90a, LR72, Ma98a, Ma98b, MKT93, Mal77, Man80, MT89b, MX95, Mar79a, Mar90b]. Nonlinear [MN84, Mey68, Mey73b, Mit77, Moo77, MQ82b, Mor72, MS99c, NT92, NPV93, O’S90, OR66b, OR67, Osb75, Osh81, Pao99, Par95, Per67b, Por69, Qn99, Qi82, Rac73, Rau73, Red76, Rey72, RD79, Rit75, RSE69, Rob66, Rob76, Ros73, Ruh73, SSA84, SEK93, SF84, Sch96, ST95, She78, Sim71, Sim75, Tad91, TE98, Tho76, Tuc69, Wat79, Wei69, Wei74a, Wei74b, WB98, Xu96, dG81, AP87, Ban83, BR80, CL90, Cha85a, CT86, CC84, CG84, Cor81, DPZ87, Dob80, Doe79, EAVD79, Epp84, EJ87, Ewi78, FR83, FR85a, Fon88, Fri81, FM90b, GM78b, GGL79, HM85, Hay81b, Hof78, IPZ79, JB88, JS85b, JSC91, Jer88, LeV88, LM79, Lus79a, Lus79b, Mar79b, MS83, Mit86, Moo78, MS80, MK80, Neu85, Pas82, Rhe80b, Rua89, San85, Sch85]. nonlinear [Sch86, SW85, Sma88, Tad88, Tad89, Ver82, WH86, Whi86a, Wu96a, Xu89, Zir82]. Nonlinearly [CS94, BSS87, NO85]. Nonlocal [CKS95]. Nonmatching [CDS99]. Nonmonotone [PT91, BPTZ92, GLL86]. Nonnegative [Bra66b, HP68a, Rou71, Wil74, Bun81, BRD81]. nonnormal [KPJ82]. Nonoscillatory [JLL^{+}98, LO96, OS91, HO87]. Nonoverlapping [Dri99]. Nonperiodic [VWI97]. Nonpolynomial [BBT82, Bru83]. nonrectangular [Hay81a]. nonreversible [GK96]. Nonself [Osb67]. Nonself-Adjoint [Osb67]. Nonselfadjoint [Osb76, Wan93a, Wan93b, Xu92, ES86, Las84]. Nonseparable [CG73, ES86]. Nonsmooth [Bre77, HKT92, LT89a, Ric78, FN84, Sam83, Sch80]. Nonsquare [BW90b]. Nonstaggered [JLL^{+}98]. nonstandard [FP86]. Nonstationary [HR86, HR88, HR90, Kla98, LV90, HR82, KR94a]. Nonsymmetric [BKP94, EES83, KS97a, LZC92, Sta91, Wid78, Ban81, CW93, Lus79a, RY86, Saa87]. Nontriviality [Joh75]. Nonuniform [BK75, Güs81a, MS94, Sül91]. nonuniformity [Bra80]. nonunique [MG82]. Nonuniqueness [AC96a].

Nordsieck

[BS94, CLM87, CLM89, SJ77, SJ83]. Norm [AH77, BJZ94, BT97, Bra94, Bre75, BP80, Cad74, DKW82, Gut93, Heb71, Hei98a, Loc73, Nie80, Pal96, Reu92, Reu94, Wah74, Wah75, in 96, BC94a, EAVD79, FP86, Mar78]. Norm-preserving [DKW82]. Normal [CH72, EH75]. Normalization [Kuz99]. Norms [Buc66, FF70, Heb71, Kre78b, RD79, Str75a, Gün88, KRD78]. Note [BV92, Coh71, Dur88a, Eis83, GTV90, GL75, Her67, JS68a, Ker71, Lam76, Mor97, Pet91, Qi82, Ral74, Rie71, SV73, Tsa75, Wil73, in 96, Cor81, Gat88, Hil80, Pai84, PR86b]. Notions [DP98]. Novel [Aus71, Faz96, GANT98]. nuclear [EYL89]. null [GO81]. Number [CMSW79, CP76, GC95, IB97, Joe99, KSS80, Kla99, Dem83, GL80a, Ste79b]. Number-Theoretic [Joe99]. Numbers [AF68, BS97a, KR74, Kun75, Lin91, MSW95, Dix83, GMP93]. Numer [BCL89a, But65, CGT89, Cra78, DF76a, DW82, GN83a, Gup85b, Hic97, HL82, Hil78, HEFS74, Jel79, Laz82, LL95, MR85, Ost65, Pie71, Sob79a, Sym81, Wei95, de 78a]. Numerical [AP73, AGP94, And79, AOW93, AP95b, AM98, Atk67, Atk69, Atk72, Atk73, Atk74, Atk77, AV93, AW80, Bab72, BvA94, BK78, BGJ88, BF75, Bec92, BW94, Ben92, BCV95, BCR75, BC95a, Ber82, BK97c, Bjø83, BL87, BK77, BP89, BP67, Bru90, Bub86, BC99, BC73, CM65, CC66, CC67, CSS97, Can83, Cas81b, Cas88, CS72, Cha78b, Che93, CC92, Chu92, Cia75, CGY98, CKL91, CG73, Cop90, CP91b, CH94, CT72, Cul71, Dav97, Dav94, DeL94, Dej67, DM92, Die92, Dol99, DE98, Dor70, Dou93, DR82, DN91, DW99, Dub90, Dus95, EK99, Eis67b, Eis74, EH98, Ewi75, Faz96, FS74, FD98, For73, FR68, FP95, FL92, FD91, Fri71, Gar71, Gar83, GGM99, Gau67, Gea78, Geo77, Gla69b, Gol98a, Gol98b, GJL99, GL96a]. Numerical [GM98, Gre94, Gre64, GSY92, Gue92, Hab77, HDS97, HP78a, Han71, Har90, HV96, HHS78, HC86, Her76, Her75, HK91, Hil68, HR99, Hu97, Hug98, HEFS72, HEFS74, ILM^{+}96, ILT69, IIMPL82, IT91, Jac81, Jam68, JP67, Jam69, Jan94, JP96, Jen92, JX98, Joh78, Kam78, Kat69, KC74, Kin75, Kla99, KP85, KSC99, KK81b, KNB86, Kru98, Kus68, Kuz99, Lam67, LE97, LLNS96, LT89a, Ler75, LB90, LSY87, Lin72, Lin69, Liu99, LGMMV92, LL99, LM67, LT68, Ma95, MF97, Mas90, MQT98, MS99a, Met94, Mey71, Mey73b, MG66, MT97, MS97a, MV99, ML92, NP81, NR72, NV88, NtB90, Omo77, PP96, Pao87, Pao99, Par65, Pet98a, Pet92a, Pre73, Pri77, PFX95, RR95, Rak76, Rei99, Rhe78, Ric78, Rin90, Riv75, Rog93, Rüm82, Rus70, RS75, RT95]. Numerical [SM96, Sam86, SS88, Say67, Sch93, SS99, Sec65b, Seg72, SW99, Sim75, Sle68, Smi68, Smi70, SS74, Spi66, Sto71, Sto73, SW68, Swa72, Sym80, Sym81, Tar96, Tav71, TK96, UTK96, Var73, Var67, VF93, War77, Weg86, Wei92a, Wei72a, Wei74a, WV79, Whi82c, Wil72, WO96, Yam98, Yan94, ZT67, ZB96, de 78a, dW79, van95, Asc89, Atk82, BM87, BGO89, BMS79, Bak80, Bey87, Bia89, BR88, BO86, Bub85, Cle81, CM88b, Des81, DFK87, DE79, Ell83, ED82, FR87b, Fis82, FNO87, GM83, GL78a, GGT82, GGL79, IT80, Ioa81, Ise84, Jac88, JS85b, JSC91, LK80, Lin85, Lin88, Luc85b, MVV89, MW86, MR83, Mon86, Pet81, RBC79, Ros81, SL83b, Smi81, Sty89, WL95, Whi82b, Whi85, Wol83]. Numerically [BW76, HB86, Pai79]. Numerics [JRB95]. Numérique [Cia75]. Numerov [DM77]. Nutshell [Alb96]. Nyström [Bia89, Cha85b, Ger89, OS92, Rat95, van81b, vS87, vS89a].

O. [FJ91]. O.D.E [Ale77, Zir82]. oblique [Saa82]. Obreshkov [MS68]. Observability [CD88]. Obstacle [HK94, Pet92a, ZZ98, CHM88]. Obtain [Tar96]. Obtained [Man90]. Obtaining [Cad74, JS68a, Mat72]. occasion [MSW94]. Odd [DS97, VS97]. Odd-Even-Line [VS97]. ODE [And83a, BH86, GP84, Hof76, Kuz99, PT97, PR86b, TS93]. ODEs [AE84, BM87, BP90, Ise84, Jac88, Yen93, dK86]. Oil [GH77]. Old [Sti64]. One [BR81, BG71, BH71, Bak84, BC95a, BL92, CGS92, Cal76, Che95, CM90b, Dah83, Dar91, Daw91, Dol99, EC93, Fer73, Gar93, GM69b, HMMR95, Jac84, Kah67, KC93, Mao95, MMD94, Mey68, MF90, Moo94, Olv65, PT97, Per92, Rac68b, San98, Sch98a, Tav71, Wah75, Wat79, Whe73b, Whe77, Whi82c, WO96, ZH91, Zen85, ZH94, ddd83, AF86, Bak82, BO88, Cha94, Cho78, CR84, Dun80a, Gar84, Geo80, HLN83, IT88, Joh88, KL86, Lav89, LB96a, MS89, NV79, Osh81, ST95, SWC89, Wat81b]. one-block [Cha94]. One-Dimensional [BR81, BL92, Cal76, Che95, Dar91, EC93, GM69b, HMMR95, MMD94, MF90, San98, Wah75, Whe73b, WO96, ZH91, ZH94, Bak84, AF86, CR84, Lav89, LB96a, MS89, NV79, ST95]. One-Leg [Dah83]. One-Parameter [Mey68]. One-Pressure [BC95a]. One-Sex [KC93]. One-Sided [CGS92, BO88, Dun80a, Osh81]. One-Space [Rac68b]. One-Step [Jac84, PT97, Tav71, ddd83, Zen85, HLN83, Joh88, KL86, SWC89]. one-way [Geo80]. Only [Hof76, OS92]. Open [Tra96]. Operations [Cha73, CO87]. Operator [AR95, AR96, BG98a, Dan67, HL97, KN72, LTW96, LL96a, Lin77b, Mal98, Mey68, Nas70, O’S90, Osb76, Pet65, Phi72, RW95, Saa92, Ste71c, Tan98, Wah77, WW83, Woz74a, Woz75, Zem95, ABPR86, Des81, Dun80a, Fun87, GK96, GL83, Hop82, Tru84, Xu89]. Operator-Splitting [Tan98]. Operators [AC83, AT86, Atk73, Atk75, Aub68, Bey92, Boh77, BW93, BT80, Can94, CH72, De 73, DF84, Don70, EW72, For66b, FHM67, Hes98b, IIMPL82, Kar96, Kee90b, KP97, KK91, Kut70, Mil69, MP68, Mun73, Nic67, OR66a, Osb67, Pal93, Rou71, SLW75, Shr73, Ste68, Tho67, Tho68a, Van67, Wei74b, Wen68, Yav95, dC90, Cha78b, Del89, Hac79, LM79, LP80, MM79, Mil79a, Mil79b, Par82, Pas82]. Optimal [Arc78, BS87a, Bar67a, Bar67b, Bas85, Ben96, Ber89, BG74, Cas97, CJY91, CK92, CH97, EJ95a, EW96, Fra78, GK86, GN91, GT74b, Hac81, Hag75, Hag90, HJ99, Heg92, Hes98a, HZ94, IT88, KAD93, Kea73, KS97b, Kra87, KT76, LLNS96, Len86, Lin77a, LP78, Man71, Man82, MO93, MS73b, Med98, MM79, MM73, Mil79a, MT68a, Pet99, Pie95, RW96, Şab97, SEK93, Sei96, Sha85b, Sha95, SP92, Son98, Sta91, de 99, hTZ97, UTK96, Vér90b, Vog86, Wei72a, Wei72b, Whe73b, Whe73c, Wil78, WGS90, Bre89, BRD81, DKW82, Hil80, Neu88, SS96a]. Optimal-Order [EW96, Bre89]. Optimality [Hin72]. Optimally [Riv75]. Optimization [Byr90, Con73, CGT89, CGT91, CR97, EA91, Fon90, Les91, MW95, Mir71, Mor68, RR87, Tap74a, Wal73, BT84, BSS87, CGT88, FST87, Fon88, HM85, MP78b, NO85, Sor80, Ste83]. optimizers [Dav80]. Optimum [CP77, Den74a, GHY83, RS92]. Orbital [BP67]. Orbits [AM98, BK97c, CSS97, Kee86, Sch83c]. Order [And83a, Arc78, AD78, Atk68, AL78, BCO94, Bak76, BD76, BJR90, BM99, Bel97, BFO96, BCMM98, Bre92, BF91, BC73, BP90, BCM94, But75, CLMM94, CMM97c, CMM97a, CMMP98, Cal71, CV99, Cas66, Cas69, CG90, CS74a, CS75, CH97, CJ96, CV72, Coo78, CDW96, Den74a, Die90, Dry78, Dup73b, Dup73a, Eis67b, EW72, EJL85, Enr93, EV96, EW96, FF98, FX95, FF91, For90, FSU85a, Gai74, GW74a, Gea78, Gor75, GAM71, GM98, GM97, HW81, HLN83, HW97a, Heg92, Hei93a, HR88, HR90, dHW73, DW74, HW93, HZ94, HS92, Jac84, JK74, Jen76, KR97, KAD93, Kee70, Kin73, Köh95, KL67, KT76, KC93, Law66, Law67b, LB90, LB97, LNC97, LO96, Löt92, Man78, MF98, Mor95, Mur97, Olv65, OS65, OS91, Par95, PCL94, Per92, Pet86]. Order [Pfl97, PV72, PA73, Rac73, Rei68b, SSA91, SS93, Say74, Sch71a, Sch69b, Sem92, SZ90, SB88, She92, She96, Sin77, SBBL91, Sne96, Var71b, Var80, Vel92, Ver93, Vos88, Wah75, Wei86, Whe73b, Woz75, Yam75, Zla65, Zla67, ddd83, van79, van80, van84, AE84, AW83, Bak82, Bal86, BN85, BE79, Bre89, Bur87, Cha85b, DS80, FKD84, FN84, Fos81, FM90b, GT85, GK96, GM80, GC88, GZ88, Gup85a, Gup85b, HO87, HR82, Hor83, KW84, Kin79, KM88, LM78, Lay83a, LeV88, Lin83, MW86, Mar85, NV83, Par82, Sch80, SC87, Ske86, TW88, Var78, Vil89, WT88, Zir82, SO97]. Order-Preserving [Köh95, SB88]. Ordered [Esp94, Van67]. Ordering [EC93, LS76, You72a]. Orderings [Van72, GN83a, GN83b]. Orders [OR93, PTP96, Voi71a, Whe73b]. Ordinary [Ale94, But81, BJ97, Cal71, Coo78, Coo81, DNT83, Doe78, ED77, Enr74, EH76, Est95, Gai74, Gea65, Gea78, Gra65, HEFS72, HEFS74, Hur67, IS90, JT95, Kel69, KW75, Kno92, Lam67, Lam81, LW70, LT67, Mor95, Rei81, SD77, Sha68, Sha69, UTK96, Wat81a, Wer75, Yam75, Zed66, And79, Asc78, Cha94, GS88, Joh88, KP79, Kre78a, Kre79, Maj84, Maj85, Mey87a, Pet81]. Ordinate [GJL99, Mad71, Mad75, Pit78, LN82]. Ordinates [Asa89, Nel73, AVG89, NV79, NV83, Vic80, VG86]. Oriented [OR90]. Original [Wet97b]. originating [DL83]. Orthogonal [Bar99, BFB92, Bia93, Bia98, BF99, BB71, BJN81, FM87, Fra71, Grz85, HP77, HS99, LFB98, ML69, PS77, PR76, Pri79, Str67a, Str69b, Str70, SHR96, YF92, BC94a, Saa83, Ste80, Vin86]. Orthogonality [ST91]. Orthonormal [Can94]. Orthonormality [Kov70].

Orthonormalization [DRV94, SW77, Mey87a]. Oscillating [vS87, Mos83]. Oscillation [GM98, Fis82]. Oscillations [LTW95b]. Oscillator [IIMPL82]. Oscillators [IIMPL82, Mah82a, Mah82b]. Oscillatory [EH89, EL97, VAF98, vS89a, Gab85, Pet81]. Osculatory [Bus85]. Oseen [AT97]. OSLO [LM86]. other [Saa82]. Outer [GGK98]. Outflow [Kre91]. output [Hor83]. Over-Determined [Cli76]. Overcoming [Bre98, Car94]. Overdetermined [BCC78, BCS78, BCL89a, BCL89b]. Overestimation [Neu87]. Overlapping [CDS99, Par99, Wu96b]. Overrelaxation [HP75, JR75, CM88a, GHY83, Hat82]. Overspecified [Can68].

P.D [Pet65]. Package [de 77]. Padé [Ged81, HP80, Ise79, Nor78, WW83, Wyn68]. Pairs [Eri80, PTP96, SV94, Ver90a, Ver91, Wil73]. panel [Rom90]. Paper [CGT89, Gla69a, Nic67, Wuy70]. Parabolic [ADZ96, Akr99, AE97, AWZ96, Arc77, BL94, BF99, BJS73, BSTW77, Buz77, Cal74, Cal76, CH68, Car72, Car74, Che73, Cho78, DD94, DWW98, DF75, DF76a, Día79, DD70, DFJ74, Ebm98, ES93, EJ91, EJ95c, EJL98, ELPV93, ELV94, FT74, FR68, Gek75, Gol98a, Gol98b, Gol91, GP98, GG76, GYF86, Hil68, HT82, Jam69, Joh76, JNT90, Kee90a, KW76, Kin72, Kin75, KKP94, KK91, LSS94, LT89a, Le 94, Lin90, Mey73b, Mil69, MF90, Moo94, Moo99, NV88, Pal96, PP96, Pan98, Pao87, Pao99, Rac68a, Rac68b, Rac73, Rey72, RW96, Sho75, TE89, TW75, TXZ89, Tho76, Var71a, Var71b, Var77, Var80, Wah75, Wei92a, Whe73b, Whe73a, Whe75, Whe77, Bak80, Bak82, Ber82, BEL89, BQ86, Cas84, Dob80, DDE79, EJ87, EJ95a, EJ95b, GM80]. parabolic [Hac81, Hay81b, Hof85, Jam78, JTW83, Lar89, Las84, LM78, Lus79b, LR82, MR83, Mey81b, PS82, Rou82, Sam82, Sam83, ST86, Ste79b, Wah80]. Parallel [ABP93, BGM96, BP96, BW94, DM88, DD93, El 90, Hel74, KM93, KT97, OPF97, SB77a, TS93, Zha95, Cha94, DS89a, Whi86a]. Parallelism [JN95]. Parameter [BR80, Bar70, DR90, JS85a, KS92b, Mey68, SEK93, Sha70, Hil80, Neu88, RK88]. Parameter-Dependent [DR90]. Parameterized [Rhe93]. Parameters [Chu90, Hon93, Kea73, LW92, Sta91, JS85b, JSC91, Man82]. Parametric [CM99b, Eps76]. parametrically [GU88]. Parametrisation [Moo96]. Parametrization [Eps76]. Parametrizations [Lee92]. parametrized [FR83, FR85a, FR86]. Parrott [Næv91]. Part [Cas88, JN95, Lyn78a, Lyn78b, de 78b, Par87a, BK97a, BK97b, BK99, BCMM98, BMM99, CLMM94, CMM97c, CDMCW99, Coc89, Coc90a, Coc90b, CGY98, FP95, Han65a, HS67, HR86, HR88, HR90, IB97, IPS91, JS85a, JSC91, Rob75, Rob76, SW94b, WS93]. Partage [DS89b]. Parter [MSW94]. Partial [ARW95, AS69, BQ76, Ben96, BB74, Bun74b, CLMM94, CMM97c, CP91a, Dav97, Den74b, Erd65, Ewi75, Fai94, FT74, FR68, GS96, HVR88, HRR94, HR97, Kee70, KP96, KS94, Mar90a, McA66, OR93, Ott96b, Pan98, PR76, Rac73, SS88, Sto68, Sto73, Van72, Wal73, Wei72a, Whe73a, YF92, dC90, van84, AF86, AS84, BR80, CD88, ED82, Ewi78, GM80, LM78, MG82, Sch86, Ske81]. Partially [Hof76, Van67, WB86]. Particle [BC95b, CR84, Dra94, EH89, GS91a, HV96, HV98, Iss96, MG95, MC93, Ten92, VTG91, VA91, GV89]. Particle-in-Cell [VA91]. Particles [BG98b, GLV91, Nel73]. particular [BO87]. partition [DS89b]. Partitioned [Jay96, Loe70, Mur97, Var70, BW86]. Partitioning [Ste65b, Wei72b, BSW95, GL78a]. Partitions [Zwa73]. Passage [CP91b, JX98]. Past [NV97]. Patch [Sha69, Stu79, Hay81a, Hay81b]. path [SC87]. pathfollowing [DFK87]. Pathological [HR99]. Patterns [HV70, MS99b]. PC [Ste79a]. PC-codes [Ste79a]. PDE [Ter98]. PDEs [CW96, CD98, Xu96, Wal99]. Peano [Fer73, SC72]. PECE [Sch71a]. Penalized [KL97]. Penalties [Whe78]. Penalty [BZ73, Con73, Con76, CP77, Den74b, Dus95, FK76, NP81, She95, Arn82, Gou89, Pas79]. pencils [Dem83]. Penrose [SS74]. Perfect [ST76]. performance [BO83, DHLR84]. Period [GGM99, Wal83]. Periodic [Bao95, BC73, CSS97, Car74, Kee86, Mah82b, Sch90b, VP93, Wil78, WO96, Ber82, FN80, Luc82, MT89a, Sch83c, SB88, Ste79b, TK84]. Periodicity [Woo71, Cha85b]. Perpendicular [Alf85]. Perturbation [Bog77, DK70, EH92, EI95, Eld80, Gay81, Gay83, KS97a, KK81b, Ott96a, RSE69, Sem92, Ste77, Sun83, Wei92b, AO82, AW83, Ber86, Chu87, Dem87, Osh81, Smo83, SO86, Sty89]. Perturbations [Dan73, Dan75, Dor70, Har90, HMMR95, MS88]. Perturbed [AB99, Dör99b, Dör99a, FMOS96, GK97, Gol91, KS97b, KS99, Kin75, Mat98, Rei81, Sch90a, TKJ94, Far88, Fer86a, Gol89, Rin84, Scr91]. Peter [Bir83]. Petrov [BW93, CX98, DN89, Ker96]. Phase [DN91, Gre94, Hol78, JN98, JNV99, Mey71, Tar96, vS87, vS89b, Arb89, Bey87, Dou83, Lin88, MVV89, RBC79]. Phase-Lag [vS89b]. Phases [Cia75]. Phenomena [BDH98, Rin90]. Phenomenon [GS96, TT87, Fis82]. Physical [Var67]. physically [Scr91]. Piecewise [AGS83, BJK90, BCS78, BW89, BFB92, BDSW66, CF85, Cle90, Dia77, Die90, EJL85, FC80, Ger86, GS96, HS97, Joh69, Ker96, Pat76, TT99, hTZ97, TXZ89, AS85, AG87, BGO89, Bia89, BC94a, CF89, GHS96, Smo83]. piecewise-linear [AS85]. Piecewise-polynomial [Ger86]. Piessen [Cop90]. Pipeline [Kee95]. pivotal [Zla80].

Pivoting [Ale87, Bun71, Bun74b, Ske81]. Planar [Bre93, CMMP98, Fra73, HP77, Hue73, Nic92b, vL92]. Planck [BC99, BTW93, HV96, HV98, Sch98a]. Plane [CG90, Cha92, Cur66, DM95, EW76, IT95, Ste71a, Wil72, Bru80, Saa87, Wil79]. Plastic [HM76a]. Plasticity [Car96, Joh77, LB97]. Plate [DGW91, GN91, MTF99, PRS90, SS97, TMV98, AF89]. Plates [DP98, Lov96]. Plus [KK93]. Point [AD78, BG74, BK77, BPV97, Cas88, CM99a, Che98, CGH91, DM77, Dia77, DW76b, Dup76, EG94, Fra73, Gla69b, Gre91, HKT92, HLK91, JP98, Kel74, KLY76, KS97b, Lam81, Lan77, LP77, Mar92, MT75, Oli74, Ols77, Red76, Reu92, RSE69, Rus77, SSW96, SW77, Sin77, Ste75b, Ura68, VF93, Wat79, Whe73c, Whe74, Whe77, Yam98, YA95, ddd83, dd86, ALY88, AD92, Asc86, BM87, CK85b, Doe79, DR84, FKD84, Far88, Fer86a, JP92, Ked81, KNB86, Lay83a, LOR85, Nic82, ST78, Sch80, Whi79]. Points [AB86, Atk73, BDLN92, DK80a, DK80b, DG94, FD91, GR96, Gun74, Jan94, Koc84, LSS94, LNS95, MS65c, Row69, Sch93, Wei74b, WF97, dS73, Bak84, Bey87, DK82, DKK83, DFK87, FR87b, GR84, KS83, MS80, RR86, WS84]. Pointwise [BT69, DN89, GN91, Mad71, Mit77, RB79, Son98, TT99, Gar84, KN88]. Poised [AS69]. Poisson [Gen84, HV96, HV98, Sch98a, AR96, Bia93, BL87, BH65, BGN70, BDGG71, BD74, Buz77, CR84, DF76b, Dör96, GV89, GLV91, Hei96, May84, Rin92, Spi65, Spi67, SB77b, Sül91, SS73, VTG91, VA91, WO96]. Polar [Ben95]. Pollution [BS97a]. Polygon [Dou93, Kel66]. Polygonal [BNP91, BDLN92, Cha86]. Polygons [DF75, DF76a, LB98, NtB90]. polyhedra [SS96a]. Polyhedral [Rat95]. Polylogarithmic [PW96]. polymer [MR83]. Polynomial [AH74, AGM92, BT99, Bru78, Bru91, CJ76, Cle64, CW78, Dem76, Dia77, Dun74a, Dur83, DGR96, For78, GM75, GK90, Gra71, Gre99, Hen70a, Hes98a, Joh75, JMP83, Kah67, LW92, LMST69, MS97b, Ost70, Ost71, Pat76, Rub64, ST92, VMK97, VVC94, Wat77, Woz74b, de 78a, Bru80, DS84, GL80a, Ged78, Ger86, GT82, Hak82, HHS78, Hu96, LSY87, LSY89, MP78a, Pet89, Sid86, VH93]. Polynomials [Bac80, Bar99, Bas73a, BDSW66, BJN81, Car92, Cur66, DTW76, DTW78, Die90, Fra71, Gar78, HP77, JT70, Kim76, LLNS96, Mas73, Mat72, MG66, ML69, Mur71, Not91, PK90, Por73, Rei88, Riv64, Ros67, Row67, Row69, ST69, Sew66a, Str65a, Str67a, Str69b, Str70, Tay68, Tse96, de 88, Gün88, Pri79, Saa83, Saa87, Sch83a, ST95, SL83b, SS80, TD82, Vin86]. Polytopes [VVC94]. Population [KC93, AM89, IKP91]. Porous [AWZ96, BK97a, BK97b, BSW85, Dou83, Dur88b, EW80, ER82, EYL89, Rus85, Sam86]. Portion [JP67]. portraits [Bey87]. Posed [Liu94, Neu97, Pla98, Qn99, Var73, Wen68, Car94, HK79, Mon86, SEK93, Tru84]. Positive [EW72, Esp94, Gun74, HP68b, Kat69, Mat72, Min70, Pla98, RS92, SV70, Van72, GHY83]. Positivity [AF68, EV96]. Positivity-Preserving [EV96]. Possessing [Mil66, Rei72]. Post [BW89, BX89]. Post-Processing [BW89, BX89]. Posteriori [AB99, BDR92, BS97c, CV99, FHK96, JNT90, MT75, Moo99, SEK93, Woh99, BE79, Gat88, Han93b, Ked81, Neu88, AT97, BR81, BS96, BW91, BS93b, BEK96, BV96, Est95, JT71, MS99b, Moo94, Ost73, Pad97, Rei81]. Postprocessing [GANT98]. Potential [Can83, CC92, CKL91, Dav94, Dub90, GGS98, JN95, LT68, Pie69, Pie71, Rat95, YR99, MN87]. Potentials [Gue92, Par65]. Powell [Buc76, Naz86]. Power [AD68, DR83, EL72, Gai66, LS99, PP73a]. Powers [Buc66, Cli68, MP77, Far85].

Practical [DM95, Wah77]. Precision [Fra73, Lan64c, LO90]. Preconditioned [CJY91, EG94, GHK94, KS94, KK93, PRS90, Cha89, ME81, Que89]. Preconditioner [Ain96, Mal98]. Preconditioners [AMW99, Cha87, CY92, CY93, CNP93, Kea90, LNC97, MO93, Ott96b, SW94b, WS93, Gol89, JMP83]. Preconditioning [AMT99, Axe79, AV90, ES86, Gol91, GMP93, Gol94, Han93a, Hes98b, KP96, KP97, MP90, PR95, QZ92, SHR96, BCMP91, Fun87]. Predator [Wal83]. Predator-Prey [Wal83]. Prediction [SL83a]. Predictor [DH71, Hal74, Jac86, Lam67, Lam71, Zen85]. Predictor-Corrector [DH71, Hal74, Jac86, Lam67, Lam71, Zen85]. predictors [SC87]. Prescribed [LS97b, Shi66]. Presence [CR90, ES98, Gau67, Nie99, Rab67, RS84]. Preserve [MQT98]. Preserving [BZ85, Cle90, CF90, EV96, Kla99, Köh95, Sch83b, DKW82, GU88, SB88]. Pressure [BC95a, BCM91, CP98, Dou85]. Prey [Wal83]. Primal [BCL89a, BCL89b]. Primitive [Kar92, SW99]. Principal [ED77, CM88b]. Principle [HRR94, Hud68, LL96b, Liu93, LO96, Qn99, Tra65, Var66, ABPR86, AB87]. Principles [BCMM98, BMM99, MA81]. Priori [BHM^{+}99, Dör99a, CL90, CDMW98, CDMCW99, CGY98, CDW96, Sho75, Sig76, Whe73a, YB96]. Priver [Gla69a]. Probability [Hug98, SS96b, Wat79]. Problem [AH94, AH77, AR96, Atk77, BG71, BJK90, BJR90, BW91, Bel69, Ben92, BCR75, Bjø83, Bof97, BBG95, BJS73, BNP91, BH65, BHZ68, BH68, Bre93, CMMP98, CD67, CD77, CG97, CD98, CN97, CC78, CK98, Cho92, Cia75, CKL91, CS90, Cos71, CH94, Cra76, Cra78, Cur66, DG98a, Der71, Dia77, Dub90, Eld80, Epp82, EJ91, Esp94, FH72, FP95, Fre90, FL92, GGS94, Gar83, GW97, GG90, GV80, GW90, GL93, Gre94, GM97, Han93a, Hei94, HR86, HR88, HR90, ILT69, Jam69, JNT90, KC74, Kau74, KK91, Kub70, Kut72, LV90, Lee98, Li87, MF98, MG95, McL74, MS99a, MTL96, NP81, Neu98, Nic92a, Pan97, PS80, Pel67, Per97, RHH99, Rin92, Rog93, Ros73, Ruh73, Sac71, Sem92, SWW97, Sig76, Sim69, Smi73]. Problem [Ste72, Sti64, Str75b, Swa72, Sym80, Sym81, Tar96, TXZ89, Tra77, VF93, Wei92a, Wei92b, Whi82a, Wuy74, Zha95, Ban83, Ber86, Ber82, BCM88, CFQ83, Chu87, CC84, DO79, DNT83, DL82, DDE79, DS89a, Dry83, DN89, Far88, Fer86a, GM78b, Gui87, GK83, Gut83, HR82, JS83, LT88, LT89b, LB96a, Lin85, Lor84, MVV89, Mon86, NW89, Neu85, Pas82, Pie95, RBC79, SW82, SO86, Sty89, Sun83, TW88, Toi79, Whi82b, Whi85, Whi86a, Zha96, Zho83]. Problème [Cia75]. Problems [AB99, AE97, AW95, AWY97, Arc78, AD78, AS69, AL96b, AK98, Bab72, BR81, BCO94, BS96, Bad91, BS97b, BR75, BR77, Ban77, Ban82, Bar70, BV92, Bel69, Bel97, Bia98, BF99, BG74, Bog77, BEK96, BJMT73, BD99, BDLN92, BK77, BBS73, Bra66a, BN73, BKP94, BPV97, BHM^{+}99, Buz77, CW96, CMM97b, CDS99, Cal74, Cal76, CM65, CC66, CC67, CG98, Car72, Car96, Cas88, CNP93, CfXmZ99, CM84, CC96b, Che98, Che73, Che75, Che77, CC92, CKV98, Chu92, wECGS94, Cli73, CR90, DR90, DM77, Daw91, DD94, DW76a, DF75, Den75, DF76a, Den77b, Den77c, Día79, Dor70, DR82, DSW94, DW76b, Dup76, Dut90, EH92, Eis74, EI95, EG94, EM88, EJ91, EJ95c, EJL98, ELPV93, ELV94, EW96, Fai72, FM90a, FF93, Faz96, FFLMW99, Fin71, FL71]. Problems [FR78, FU69, Fun90, Gai74, GGS98, GK97, Gea65, Gek75, Gek76, Geo77, GPV78, Gla69b, Gol77, Gol91, Gol93, GP73, GP98, Got72, GG76, Gra65, GM98, Gre91, GYF86, Gup75, Gus70, GKR70, Hag75, Hag76, HDS97, HKT92, Hei98a, Hei93a, Her68, HK91, Hil76, Hil78, DW78, HK94, HW97b, HZ94, IT95, Jaf92, Jen92, Jer73a, Jer73b, Jes78, JS98, JP98, Joh77, KHK93, Kac99, KP78, Kel74, KW75, KS97b, KS99, KW76, Kin72, Kin75, KS94, KMR94, Kre79, KK81b, Kru98, KS92b, KT97, Kus68, Lan77, LSS94, LMV96, LP77, LM90b, Li91, LZC92, LB97, Li98, LL89, Liu94, LR72, Lue70, MTF99, Man78, MW92, MR85, MMD94, MJ77a, Mat82, MT75, MR82, Mey71, Mey73a, Mey73b, MM73, Mil69, Mir71, Mit77, MS73d]. Problems [MV99, Mor98, Neu97, Nic92b, Nie99, NV88, Oli74, Ols77, Osb75, Ott96a, Pal96, PT91, Par95, PR95, Par99, PCL94, PRS90, Per67b, PV72, Pla98, Pol74, Por69, PL87, Pri77, Pru73, PFX95, Qn99, Rac68a, Rac68b, Rau73, Red76, Red78, Reu92, Reu94, Rhe78, Rit75, RSE69, Ros70a, Rul96, RW96, Rus70, RS75, Rus77, RC78, RT95, Sch68a, SEK93, Sch90a, Sch82, Sch72, Sch69e, SW77, Sha68, Sha69, She96, Sho75, Sim71, Sin77, Slo76, Ste71a, Sto71, Str65a, SW91, SHR96, SW74b, TE98, TE89, TMV98, Tap69, Tap74a, Tap74b, Tav77, TW74, TW75, TKJ94, TH98, Urb91, Var71a, Var73, Var67, Vav96, Ver84, Ver88, VS97, VAF98, Wah75, Wan93a, Wat79, Wei86, Wen68, Wer75, Whe73c, Whe74, Whe75, Whe77, Whi82c, Wig66, WH97]. Problems [Xu92, Yam98, YB96, Yse86, ZZ98, ZB96, ZWW94, Zla67, ddd83, dd86, dW79, van80, vS89a, AO82, AW83, Asc86, Asc89, Axe79, BM87, BGO89, BvA94, Bar88, Bea87, BC94a, BW86, BR88, Bog85, BPTZ92, BL87, BE79, BEL89, Bub86, BNY87, CW93, Car94, CK85b, Cle81, CM83, CL92, CHM88, Cor81, DK85, DPZ87, Dem87, DS86, Dob80, Doe79, DR84, Don81, Dor86, Dou84, Dyk87, EHP87, EJ87, EJ95a, EJ95b, EW80, ER82, FKD84, FNO87, Gar88, GL78b, Geo80, GN83a, GN83b, Gol89, GGL79, GZ88, GO81, Gri87, GG86, GHS96, Gup85a, Gup85b, Gus80, Gus81b, GK83, Hag86, HK79, JN83, JB88, KW85, Ked81, KP79, KS87, KNB86, KHR89, Lar89, Lay83a, LK80, LOR85, LNOP82, LMS87, Lin88, LP78, MMR88, MW86]. problems [MR84, McC84, McC85, MN84, Mey81b, Mey87a, Mit86, MP78b, Mül84, Nic82, Nic87, OR86, Osh81, Pai79, PSP89, Pas79, PR82, PR86b, Rin84, Rou82, Sam83, Sch80, Sco81, Smi81, Ste79b, SB84, TK84, Van85, Ver82, Wan93b, Wei89, WW88, Whi79, Wom89]. Procedure [Aus71, AK98, Bai91, Ban82, BJMT73, BBS73, BB74, Cad74, DKR68, Dup68, FM67, KW76, Obi90, PW78, Spa71, Sto73, VAF98, Whe74, CEv87, Cle81, FQZ88, LL83, LSY89, Lus79a]. Procedures [Buz77, DD94, DR82, Gek76, Gre64, Gup75, Ost73, SSA84, SK93, Voi71a, Voi71b, Wei69]. Process [Bic77, Law66, Law67b, Med98, Sid97, Sim69, Wil69, FS87, Sam86]. Processes [BK97a, BK97b, FF70, Fav64, Law67a, Loc73, Nic73a, Ort73, Rhe68, Rhe80a, Riv84, Wim70]. Processing [BW89, GL96a, BX89]. Product [Atk74, Bol72, Drm98a, DL75, Ehr98, Let70, Let71a, Let71c, MS83, MSW95, RS84, SV96, SS80, SP92, Ste68, Wei71, CM88b, Dag86, Dyk87, SS82]. Product-Decomposition [MSW95]. Product-Induced [Drm98a]. Product-integration [SS80]. Product-Type [Bol72]. Products [GBIY71, RD79, Gra87b]. Profile [GPS76, AH82]. Profiles [BGS98, Kar92, Bub85, CLNP89]. Program [Sin73]. Programming [Afr70, Asc77, CC66, CC67, Che73, Con76, For65, For66a, For66b, Lav91, MR77, Rau73, Rit75, BT89, CC84, CG84, Fra87, Lav89]. Programs [Man81, MS91]. Progressive [Afr70]. Projected [AP91, CD90, NO85, dT93, TW88]. Projection [AP87, Bro87, DF91, EL96, Eic93, GOQ97, HW93, Phi72, Sar90, She92, SK93, WL95, Wet97b, Saa82]. Projections [BIC66, Dem77, GM75, Güs81a, LTW91, Bru80, DS86, Grz85]. Projective [Ste76]. Proof [BG95, KLY76, Pet65, BI89, BH83, HP78b, Ker86]. Propagation [BL98, Don94, GSY92, Hig90, FN80, Wal88]. Properties [AE84, BM89, But81, CM99b, DS72, Fer74, FU69, Gay79, IIMPL82, Nea92, Ost71, RD71b, Ric70, Sch71a, SL83a, SS82, SS74, And83b, BB88, BZ88, For79, FSU85b, NV79, SSB85, Sid86, Var85, Vic80]. Property [HR88, Jim96, Per90, Rei72, Riv64, dT93, LR82, You72a]. Providing [Prü82]. proving [Ske82]. Prufer [BK73]. Prune [VMK97]. Pryce [Gov86]. Pseudo [FT74, GK65, GP73, GW74b, KK98a, MQ82a, Nak91, Mit86]. pseudo-arclength [Mit86]. Pseudo-Harmonic [GW74b]. Pseudo-Inverse [GK65]. Pseudo-Inverses [GP73]. Pseudo-Parabolic [FT74]. Pseudo-Runge [Nak91]. Pseudo-Spectral [MQ82a]. Pseudo-Transient [KK98a]. pseudodifferential [Pas82]. Pseudoinverse [Her67]. pseudoparabolic [Qua87]. Pseudoperiods [Hei66]. Pseudospectra [Rie94]. Pseudospectral [For90, Fun90, GGS94, GM91, GM93, Hes98b, HS92, IK93, MKT93, Wel97, CQ82, Hei93b]. Pure [Bre93, CMMP98, Lee98, Wil69]. pyramid [Duf82].

QL [HP78b]. Quadrangulations [LS99]. Quadratic [BD99, JT70, KK91, MS91, Rob79, Sch83b, Sha74, BT89]. Quadratically [AP83, Bro69, Kun88, MS91, CL92]. Quadrature [Bak68, BD76, BJ93, Bol72, Chu72, DE72, Don73, Duf82, DL75, ED77, EH72, EMB71, FP91, Gau67, GTV90, Hic96, Hic97, Hue73, Jet87, Kah69, KU71, KR97, Köh95, Let74b, LNC97, Lip73, Loc73, LM77, Lyn78a, Lyn78b, MRW96, MS65a, OT75, PP96, Pat76, Pet92b, RJ69, Rie71, RM77, SS93, Sch65, Sch66, Str65b, SS65, Str66, SC72, SP94, Wil78, Wil70, YA95, DS84, GV83, Jet82, JS89, KL79, Owe97, SL83b, Wol83]. Quadratures [Bar65, Bar67a, Bar67b, Cha69, Gue92, Küt84, Ver70, FP86, Ger86]. Quadrilateral [BG98a, Jam77, Zha97]. Quadrilaterals [ELW99]. Quadtrees [Gar83]. Qualitative [HDS97, JS96, GT85]. Quantitative [Bra94, San98]. Quantum [AR95, Rin90]. Quartic [Per76, APS87]. Quasi [Ain96, Bar73, CM87, Cas97, CfXmZ99, Che75, GN91, HDS97, JN98, LQ84, LC98, Mar90b, Mar92, MC93, Pad97, Ros70a, ST92, BNY87, BN89, CC84, FST87, HM85, IT88, JA83, KS87, KN88]. Quasi-Convexity [Bar73]. Quasi-Equidistributed [ST92]. Quasi-Linear [Che75, LQ84, Ros70a]. Quasi-Monte [LC98, MC93]. Quasi-Newton [CM87, CfXmZ99, Mar90b, Mar92, BNY87, BN89, CC84, FST87, HM85, IT88, JA83, KS87, KN88]. Quasi-Newtonian [Pad97]. Quasi-Optimal [Cas97, GN91]. Quasi-Static [HDS97]. Quasi-Uniform [Ain96, JN98]. Quasilinear [AC96a, Arc77, Arc78, Che73, FR78, Jac86, Jen92, LB96b, McA66, Wei92a, DDE79, MR83, Rin84]. quasilinearization [And83b]. Quasimonotone [Coc89, Coc90a, Coc90b]. Quench [LW81]. Question [HR93]. Quintic [PV79]. Quotient [BS89b, Gel81, Szy88]. Quotients [Tra65].

R [Pet92b]. R-Convex [Pet92b]. Radau [AC98]. radial [Smo83]. radiating [DPZ87]. Radiation [Hig90, Hig94]. Radiative [NS93]. Radius [Gol98a, Gol98b, HS68, MTTW99, Bun81]. Radius-Edge [MTTW99]. Radon [Lau89]. Random [BC95b, CSH78, HV96, Jen70, KC74, Pet87, HL81, Ste80, Sun79]. Randomization [CP76]. Range [Joh75, Mil66, Nav65, Bea82b, Des81, KRD78]. Ranges [ST69, Tay68]. Rank [CJMT74, wECGS94, Cli68, DL90, DS92, Dol99, JD93, LL96a, Olv65, OS65, Wei92b, Dem87, IT88, RR86, Sid86, Whi86b]. Rank-Deficient [Wei92b]. Rank-One [Dol99]. Raphson [MT68a, Weg66]. rapidly [Hig84]. Rarefactions [NT94]. Rarick [EGLP85]. Rate [Bab71, BS87a, Bre75, Cha78a, CfXmZ99, Coh72, Dan70, Eis74, ES98, Hed68, NT92, NT94, NTT94, PP73b, Şab97, Sch90b, SZ90, TE98, hTZ97, Dun80b, Fer86b, Gus81b, Li83b]. Rates [Atk75, Cal74, Den74a, Ell84, EL93, Hag76, Hoh99, KK91, LTW95a, Por71, RW96, Voi71b, BD85, BBT82, DKK83, LM88, Neu88, Saa80]. Ratio [Riv64]. Rational [BPR72, Bar73, BT79, BT80, CL64, Cle90, EW76, HZ90, Ise81b, Lam76, MT68b, Pel67, RW89a, RW89b, Sch82, Wal67, Wuy74, de 78a, AR81, Bru80, Cuy87, Del93, Dun80c, Ell83, Gut83, HHS78, KN85, TG83, Var81, Wil79]. Rationals [Loe66]. Raviart [BV96, Bre92, Mar85]. Rayleigh [BS89b, BDSW66, CFQ83, Gel81, Joh69, Omo77, PV72, Sch69e, Sch71b, Sim69, Szy88, Yam75]. Reaction [AB99, CW97, EW96, Fin71, MS98, Hof78, Jer80, Jer84]. Reaction-Diffusion [AB99, CW97, MS98, Hof78, Jer84]. Reactive [BK97a, BK97b, CE97b, DG98b, LP91]. Real [AGM92, BT99, Cop90, Dun74a, Gun74, GT82, JT70, KR74, LZC92, DS89b, Sch83a, TG83]. realistic [Ker86]. Reconstruction [ADH99, HB86]. Recovering [GS96]. Recovery [AV97]. Rectangle [Bia93]. Rectangles [BF99, Bjø83]. Rectangular [BD74, Cap99, DF75, DF76a, DMR92, Die82, Dun80b, Jer84]. Recurrence [Hos95, LB90]. Recursive [SK93, Cuy87]. Red [Yav95]. Red-Black [Yav95]. Redistribution [CD85]. Reduced [Col98, KS92b, Lee91, Ost71, PL87, Was82, vS87, GK83, Li83b, Sid86]. Reducing [BS79, GPS76]. Reduction [BG71, BP90, DF76b, Hel76, Jen92, Spa71, Spi71, Swe74, Swe77, GS88, JB88, Sch89b]. Réelle [DS89b]. Reentrant [Don69]. Refine [AC83]. Refined [AMT99, DP92, ELPV93, JNV99, Mei83, BS89a]. Refinement [AT86, AH94, BL98, DM93, Dax85, ELV94, MF90, Rhe80a, Riv84, AF86, AB87, Del88, Del89, Zla82]. Refinements [Osh70]. Reflection [YB96, Bub86]. Reflections [Kar96]. Reflectors [SP88]. Regime [AP95a]. Regimes [GJL99, MR98]. Region [Car91, CGT89, EA91, Hei98a, Law66, Law67b, Les91, Swa72, Tra77, YA95, BSS87, CGT88, Ged78, SSB85, Sor82, Var85]. Regions [BK78, BMS79, BS77, BDGG71, BD74, Fie78, Fra73, Fri71, HP77, Hue73, Lam71, LM77, MJ77b, Str67b, Tha78, AD92, BW86, Hay81a, Jer84, Jer88, JTW83, May84, Sch83c, Ste83]. Regression [For78, dG81]. Regular [Geo73, HMR73, LM77, NtB90, Wer75]. Regularity [BGO89, BH68, HR82, SWB90, Wan92, BS83]. Regularization [AL96a, BG98a, Che93, Cul71, EL93, GG90, Heg92, Hil76, Hil78, LE97, Lin97, LP80, Neu97, O’S90, Qn99, RHH99, SEK93, Tau96, Axe79, Han93b, Hil80, Neu88]. Regularizations [Wet97b]. Regularized [Hoh99]. Regulator [BJS73, KK91]. Reissner [AF89, DGW91, GN91, Lov96]. Related [AG68, BJZ94, CHL91, Die92, GA67, Gre94, LTW91, Pet92b, Sch82, Str74, WM67, BL80, DH79, Gut83, Mie81]. Relating [COP98, PV72]. Relation [Jel76, SW74a, BO83, Wol83]. Relations [HL98, LB90]. relationship [Naz79, PR86b]. relationships [LOR85]. Relative [EI95, Enr93, Gov86, Gri74, KR74, Not91, Pry84, Sin77, SW74b]. Relativistic [GS91a]. Relaxation [BJZ94, CJR97, CP98, EC93, GPT97, JK96, JV96, JV97, JN98, JNV99, JPT98, JX98, Kac99, Lei98, LL89, Sch68a, Yav95, de 68, FQZ88]. Relaxed [DP98, Fre90, Zan99]. Reliable [AB99]. Remainder [CM74, GTV90, Sta64]. Remainders [DE72, Lee85]. Remark [EK99, Ura68, Wah80]. Remarks [Ahl66, CJMT74, Dan75, Doo66, DL73, HL82, Kut72, Sew66b, Wuy70]. Remes [Gri74]. Remes-Type [Gri74]. Remez [BT99, Ree90]. Removal [CSH78]. reordering [EGLP85]. repetition [Wol83]. Representation [Aal72, Abg98, Bas73a, Bey92, Dub90, GTV90, Har96, Let70, Lyn78a, Lyn78b, Mil66]. Representations [Cli65]. Reproducing [BN74, Hil76, Hil78]. Requirements [Haa96, Ste70, Was82, AH82]. requiring [KW85]. Researches [Sew66b]. Reservoir [BJ91, Pet87, Arb89]. Residual [AK98, BHM^{+}99, CN97, KPJ82, Lue70, Neu85, PW78, Ste76, Ske81, Vat88]. Residual-Free [BHM^{+}99]. residual-like [Vat88]. Residuals [CV99, Pla98]. Residue [RSK76]. Resolution [AS80, HL84b, Jaf92, JLL^{+}98, LeV96, OC84, Yan99, GL88, Swe84]. resolvent [Mil79b]. Resonant [LTW95a, LTW95b]. Resonators [BFGP99]. Respect [SSW96, Des81]. Restart [Mey76]. Restoration [AM94, CM99a]. Restricted [AH77, Bea82b, GCB74, HZ90, Joh75, KL76, KRD78, Nor78, ST69, Tay68, TW70, Wat74]. Restrictions [Var80]. Result [Geo76, OR72, PSB97, Pal93, Pai84]. Resultants [AGM92]. Results [BCM94, CHL91, DeL94, EH75, Eic93, EH76, GW83, GS67, KLY76, LP91, Neu97, Pet86, PV72, Ros70b, Ter98, BL80, BE79, Dag86, FSU85a, MVV89, Mar79b, McC84, Ske82, Win80]. Retarded [AP95b, Cry74, Dav94, EH98, LM88, SWC89]. Retards [FMMR99]. Retorting [GH77]. Reverse [LS76]. Reversible [CSS97, Mae82]. Reversing [MQT98]. Revised [Li83a]. Revisited [IZ81, Asc86]. Reynolds [KSS80, Mey81a]. Rezoning [BC95b]. rheology [MR83]. Riccati [DO92, DOR88a, DOR88b, Die92, OR86, RW95, SS99, SH66]. Rice [Ros70b]. Richardson [FS87, Rei88, Sid97]. Richtmyer [Gar66]. Riemann [BO86, CP98, FR85b, Kur67, Osh84, Ros81, VMSB97, WW93b]. Rigorous [Bra94]. Risky [Rob79]. Ritz [BN73, LTW91, Sch69e, Sch71b, BDSW66, BJMT73, Gar84, Hag75, Jes78, Joh69, Omo77, PV72, Sim69, Yam75]. Ritz-Least-Squares [BN73]. Ritz-Volterra [LTW91]. RK [JN95]. RLW [LL99]. RMS [Her76, SW68]. Robust [AB99, Hig89a, MS98]. Robustness [BS92, NV93]. Rods [FX95]. Romberg [MS65a, Str65b]. Root [AOW93, Bra66b, DK85]. Root-Soil [AOW93]. Rootfinding [Hin72]. Roots [AH74, Ost70]. Rosenberg [Mil81a]. Rosenbrock [OR93, Ver82]. Rotated [CK98]. Rotation [DK70]. Rotationally [FP95]. Rough [BCO94, HT82, MDH^{+}99]. Rounding [Bau74, Rac68a, Rac68b, Ura68, Woz74b]. Roundoff [Mil76, Spi71, Ura68]. Row [Var76, GN83a, GN83b]. Rule [DS72, GTV90, Hab77, Kre78b, Let71c, LG80, Ura68, de 86, Bea87, Han93b]. Rules [DS92, DE72, DL75, ED77, Hic96, JS92, JD93, Joe99, KR97, Let70, Let71a, Let71b, Let74b, LNC97, LM77, LS97b, MRW96, RD71b, Ric70, RM77, BE88, Dag86, Esp87, GM83, Gen86, Hic97, KL79, MP78a, SS82, SS80]. Runge [HL84a, SW94a, Alb87, Alb96, Ale77, AC90, Ale94, AP91, BP96, Bar99, BZ88, BJZ94, BT97, Bic77, BH93, BB79, Bur82, Bur87, BP90, BCM94, But75, BC90, Cas66, Cas69, CS74a, CK92, CV72, Enr89, Enr93, FSU85a, FSU85b, Gup85a, Gup85b, HW81, HW97a, Hig89a, Hor83, Hos95, HS94, HS87, JRF91, JT95, JEH78, JN95, Jay96, Kee90a, KL67, Law66, Law67a, Law67b, LNS95, Nak91, OS92, PTP96, PT97, Pet86, Roc89, SSA91, Sha85a, SV94, SO97, Ver78, Ver79, Ver90a, Ver91, Ver93, in 96, van79, vS87, vS89a, vS89b]. Russell [Jer88]. Ryabenki [EL96].

S [de 78a, Chr91, IS90]. s-Step [Chr91].

S-System [IS90]. Saddle

[BPV97, Che98, EG94, Sch93, Nic82].

Saddle-Node [Sch93]. Safe

[Lam81, MJ77b]. Salutation [Bir83]. Sard [Rei74, BN74, Fra78]. Sard-Schoenberg [Rei74]. Sarkovskiı [Kee86]. Satellite [MF97]. Satisfaction [Dk99]. satisfy [FP86]. Satisfying [Liu93, LO96]. Saturation [But64, But65, Neu97]. Scalable [MTF99]. Scalar [BG98b, CM90b, Coc89, Coc90a, Coc90b, CG96, CGY98, JS85b, JY98, KSC99, KR94b, LMS97, Liu93, NT92, Pet91, Sch90b, TT99, Ten92, Vas99, Bre84, BO88, Lev82, Swa79]. Scale [Dav97, PSP89, Ste83]. Scaled [BD90, Hor83]. Scales [Tau96, GG86, Gus80, GK83, Kre79, Neu88]. Scaling [AMT99, GK90, GV74, Sha85b, Sor80]. scalings [Dav80]. Scattered [HS97, LS97a]. Scattering [Hoh99, Nel73, Sym80, Sym81, AVG89]. Scharfetter [Gar93]. Scharfetter-Gummel [Gar93]. Schauder [Tho68a]. Scheifele [MF97]. Scheme [AV99, AFM90, BG95, Cal76, CDK95, DWW98, ELV94, FX95, FMOS96, Fun93, GL96a, HW92, HW93, Jam71, KC74, Kla98, Kla99, Kos73, LQ84, LC98, Mar90a, MS94, MV99, MDH^{+}99, Nic92a, Pan97, Sch98a, STW97, She96, SW99, hTZ97, Ter98, TW93, Wah74, Wee98, ZH91, ZH94, Asa86, Cuy87, Dai92, Far88, HL81, Hof85, IKP91, Jam80, JP83, KL82, KR94a, Luc85b, MN84, PS88, SS83]. Schemes [Bre75, Bre77, BC99, CJR97, Car72, Cim72, Coc89, Coc90a, Coc90b, CS94, CGY98, CJ96, CF93, Den75, DL91, EG75, Eis66, Eis67a, ES98, Enr93, EV96, FF98, GPT97, GV91, Gar66, Gol98a, Gol98b, GJL99, Got72, GM66, HJ99, HL84b, HL98, Hed68, Her68, Hig89a, Hub65, JLL^{+}98, JPT98, JX98, KHK93, Kac99, KR94b, LMS97, LW98, LO96, Mad75, MW92, MM73, Mir71, NTT94, OC84, Osh85, OS91, PS80, Per90, Per92, Rey72, RS97a, RS92, SM96, SSA91, She92, Str68, Sül91, Tuc69, Var80, Yan99, ddd83, dd86, AO82, Bea86, BC94a, Cas84, Cha84b, CLS86, CS87, Dut88, Fer86b, GR82, GK96, GL88, GH83, HO87, Hat82, HK79, JS87, Lor84, Mun94, Osh81, PSP89, PR82, Rin84, SO97, Swe84, Tad88, Vil86, Vil89, WW93a]. schemes [WW93b]. Schmidt [de 78a]. Schoenberg [Rei74]. Schröder [DAB99]. Schrödinger [CLS86, CS87, Dai92, Den77a, FD98, IIMPL82, IR99, KAD93, KM99, LFB98, Mal98, Smo83, WH86, Wu96a]. Schrödinger-Type [Mal98]. Schroeder [Gar78]. Schroeder-Like [Gar78]. Schubert [Mar79b]. Schumaker [Ros70b]. Schur [Gar71, Man90]. Schwarz [Bad91, CW93, CF88, Cas97, DSW94, GC98, HV97, Mat98, Sto73, Wan93b, ZZ98, Zha96]. scrambled [Owe97]. sea [LT88]. sea-keeping [LT88]. Search [AR77, PT91, Sob79a, Sob79b, BBT82, Bas85, BPTZ92, GLL86]. searches [Naz86]. Secant [BW90b, DW81, DW82, DW93, Fon90, GS76, Ip87, Jan79, Li88, Mar90b, Pol74, WW90, DW85, Fon88, JA83, Mar94]. Secant/Finite [Li88]. Second [And83a, AD78, AE94, Atk67, BCO94, Bak76, BD76, BJR90, Bas73a, BM99, BC73, CLMM94, CMM97c, Cas81b, CX98, CDW96, Dup73b, Enr74, FX95, FF91, Gai74, Gar75, Gea78, HR90, HW93, Jen76, Kin74, KC93, LeV88, LB90, LSX90, Lin77b, NdF74, Olv65, OS65, Par95, PCL94, PCW95, Per92, Pfl97, Rac73, SD77, Say74, She96, Sne96, SCWM71, Var80, Vel92, Wah75, Wei86, Whe73b, Yam75, Yan94, van79, vSS97, AE84, Bal86, BE79, DS80, FKD84, GT85, GC88, HLN83, HR82, Joe85, JS89, KM88, Lin83, Man82, MW86, Par82, Vil89, WT88, Wol83, Zir82]. Second-Degree [Kin74, Man82]. Second-Kind [CX98, JS89]. Second-Order [And83a, BJR90, CLMM94, CMM97c, CDW96, FX95, FF91, HR90, HW93, KC93, LB90, Olv65, OS65, Par95, PCL94, Per92, Pfl97, She96, Sne96, Vel92, DS80, FKD84, HR82, KM88, Lin83, Vil89, WT88]. Sectioning [Jen72]. Sectorial [Pal93]. Seeking [Shu72]. Segmentation [KLM94]. Seidel [De 73, Kea90, Mor71, Mor72, OR66b, OR67, Por69, Por71, Yav95]. Seismology [YB96]. Selected [CGO07, MP85]. Selection [CR97, RC78, Sha70, APR83, BR80, Pie95, Whi79]. Selective [ALM92, CLMC92]. Self [Don70, DKR68, Fri71, JS98, Laz81, Laz82, LO96, GM89]. Self-Adjoint [Don70, DKR68, JS98]. Self-Contained [Fri71]. self-correcting [GM89]. Self-Similar [Laz82, LO96, Laz81]. Semi [DC80, ELV89, FF98, Gun65, Moo94, NP81, STW97, GG86, LK80]. Semi- [Moo94]. Semi-Circle [DC80]. Semi-Discrete [NP81]. Semi-Explicit [Gun65]. Semi-Implicit [STW97, GG86]. semi-infinite [LK80]. Semi-Iterative [ELV89]. Semi-Lagrangian [FF98]. Semicoarsened [NV93]. Semiconductor [CC94, Che95, JK91, MW94, ZWW94, Asc89, Ker88, MS89, RS89]. Semiconductors [HN75]. Semidefinite [Kel65, Pla98]. Semidiscrete [BSTW77, Kin75, Kur67, MTL96, Las84, Sam82, van81b]. Semidiscretisations [RW89b]. Semidiscretization [Le 94, Vas99]. semidiscretizations [JN83, Jer80]. Semigroup [IT91]. Semigroups [BT79, EJL98, LM88]. Semilinear [ES93, FMOS96, Kee90a, LSS94, LT89a, Pao87, Tav77, TW75, HL87, KM88, Lar89, Sty89]. seminorm [Bas85]. semistable [BN85]. Semistrict [Bea76].

Semiterative [EN83]. Sense [BCC78, MT97, Mot66, BCL89a, BCL89b]. Sensitivity [Liu94, Van77, Rhe93, Ste72, Pai84]. Separable [CP91a, KP78, Swa74, Var77, Cor81, Dyk87, KW84]. Separably [Lam81]. Separate [GP73]. Separated [And83a, And79]. Separation [Var79, BR88]. September [Bir83]. Sequence [Hen70b, Jia86, Kee86, KR74, Bra80]. Sequences [CJ76, Lin77a, Osa90, SFS86, Wal67, Wim77, DGB82, GB89, Kin79]. Sequential [AL96a, KT97, LE97, Lin97, Shu72, BT89, Gou89]. Ser [But65, Ost65]. Serial [DD93]. Series [AD68, Bas73b, Cle64, Coh71, EL72, Gai66, LLNS96, Mil66, Ols77, PH73, Wyn68, Smo83]. Set [Dun74b, GK86, Gra71, Oet65, WF97, DGB82, Kra87]. Sets [Aal72, AR77, CDK95, DAB99, Dun75, Hes98a, Mat72, Var70, Wil74, Dun80b, Grz85, KL86, TD82]. Seven [Fra73, HP77]. Seventh [Str67b]. Several [Chu90, Coh72, CF93, GM69a, Joh69, Wei71, JS85b, JSC91]. Sex [KC93, AM89]. Seymour [MSW94]. Shading [AC93, RT92]. Shale [GH77]. Shallow [CDMW98, CDMCW99]. Shape [AC93, CF90, GU88, RT92, Sch83b, Swa72]. Shape-From-Shading [RT92, AC93]. Shape-Preserving [CF90, GU88]. Shaping [CH94]. Sharp [BD85, SWW97]. Sharply [OPW65]. sharpness [Tre83]. Sheets [CL89, HLK91, Pet99]. Shell [TMV98]. Shells [CS99, Kir93]. Shield [Rog93]. Shift [Can94, DR83, War75]. Shock [AM94, BGS98, CS94, Dk99, Kar92, LO96, Mao95, OR90, CLNP89, Lev82, Lor84]. shock-capturing [Lev82]. Shock-Tracking [Dk99]. Shocks [EK99, ES98, HL98, HK91, JY98, Laz82, hTZ97, Laz81]. Shooting [KM93, Lan77, MMD94, GGL79, PR86b]. Shortening [Dzi99]. SIAM [BCL89a, CGT89, Cra78, DF76a, DW82, GN83a, Gup85b, Hic97, HL82, Hil78, HEFS74, Jel79, Laz82, LL95, MR85, Pie71, Sob79a, Sym81, Wei95, de 78a]. Sided [CGS92, Sha66, ZZ98, BO88, Dun80a, Osh81]. Sign [Fer73, Fer74, RD79]. Sign-Monotone [RD79]. Signal [AM94]. Signorini [CG97]. Similar [Laz82, LO96, Laz81]. Simple [Atk72, BG95, CGT89, CGT91, For73, Gir95, JP92, KT86, Lin72, WS93, WV79, CGT88, MK80]. Simplest [Sim69]. Simplex [Gra87a, Hes98a, Lyn78a, Lyn78b, LG80, Str69a, Wal98, Gra87b, GM78c]. Simplices [Riv84]. Simplified [Vil86]. simploids [DM82]. Simulation [AOW93, BJ91, DG98b, GLV91, Kee95, MS89, MC93, Rin90, Arb89, Asc89, BI89, Dou85, Ker88]. Simulations [Pet98a]. Simultaneous [AH74, Car92, Den67a, Gra71, Mar79a, MS66, GL81, IPZ79, Pet89]. Sinc [Mor95, Bia89, SBBL91, BL87]. sinc-Galerkin [BL87]. Sinc-Nyström [Bia89]. Single [Cal71, ILT69, KW76, Whe75, Arb89, Bak80, Bal86]. Single-Space [KW76]. Singular [BJ93, Bog77, BK77, CMM97b, Chu92, DK80a, DK80b, DDLT91, DL90, DRS95, Dor70, Drm98a, Drm98b, Eis66, EH92, EI95, GK65, Han71, DW74, DW78, HMMR95, IT91, Jam68, JP67, Jes78, KN72, KR97, Kel65, KL67, KK81b, LL94, Lin69, Van76, LT68, Mal98, MP77, Nas70, Ott96a, PS81, PFX95, Red78, Rus70, RS75, Sem92, Str74, Wei86, WB98, AO82, AW83, Ber86, Bia89, BL87, Bru83, Cha84a, CK85b, DK82, DKK83, DK85, DR84, DE79, Ell82, Ell84, FR87b, Fri81, Ger86, Ger89, GMP93, GO81, IT80, Ioa81, KS83, LL95, Mie86, Osh81, RR86, Sch80, SO86, Sty89, Sun83, Wei89, Zho83]. Singular-Function [Str74]. Singularities [AS80, AL96a, BDLN92, BHZ68, Bre98, Gov97, Hu97, Köh95, NV97, WH97, BvA94, GO83, Hu96, Kun88, LMS87, SS96a]. Singularity [DR65, DW76a, ET71, Gau67, Gol91, Mil66, Mye91, Nav65, Olv65, OS65, Rab67, RS84, YA95, dW79, Duf82]. Singularly [AB99, Dör99b, Dör99a, FMOS96, GK97, KS97b, KS99, Mat98, Rei81, Sch90a, TKJ94, Far88, Fer86a, Gol89, Rin84, Scr91]. singularly-perturbed [Fer86a]. SIP [Che88]. Situ [GH77]. situations [Kau86]. Six [LS97a, Law67b]. Sixth [GM98, Sch71a]. Size [GT74a, Kro73, IKP91, Rou82, Ske81]. size-structured [IKP91]. Sizing [DW93]. Skew [Buc77]. Slab [Nel73, PS83, AVG89, LN82, NV83]. slicing [Bea87]. Slightly [EK99, Mee80]. slopes [KN85]. Small [Hof76, MT97, Nie99, Yan99, RK88]. Smallest [Osb65, Dem87]. SMF [MF98]. Smooth [KR72, SS93, SS95, SW74a, SP94, TT99, Bia89]. Smoothed [KR88]. Smoothers [Neu98]. Smoothing [ALM92, BS97c, CLMC92, Cox84, Dur77, Gre66, HR88, HR97, Mun73, Nie74, SS89, Tre72, Wan92, Yav95, BD85, BRD81, Die82, LR82, LS73, Rag85, Wah80]. Smoothness [Die90, Haa96, Heg92, HY95]. Snakes [CC96a]. Sobolev [BM89, BH70, Dur83, Ewi75, Ewi78, Mar83, Sch69a]. Soc [But65, Ost65]. Soil [AOW93]. Solenoidal [BJK90]. Solid [BDM^{+}95, COP98, Gre94, SWW97]. Solid-Solid [Gre94]. Solitary [LL99]. Solutes [BK97a, BK97b]. Solution [AE97, Asc77, AP95b, Atk67, Atk69, Atk74, Atk77, AK98, AW80, BJR90, Bar70, BCS78, Bel69, BW94, BCR75, BJ93, Bic77, Bjø83, Bog71, Bou98, BBS73, BB74, BP89, BP67, Bru90, Buc77, BC73, BDGG71, BD74, Cad73, Cad74, CC66, CC67, CK69, CD77, Cas82, Cas69, Che87, CC96b, Cli73, Cli76, CR90, CKL91, CG73, Cos71, CT72, DO92, Den75, DF76b, Dor70, Dou93, DW99, Dup68, DT73, Eis74, EH98, Ewi75, Faz96, FD98, FL92, GK86, GP84, Gla69b, GM98, GR96, GM97, Gun65, Gus70, Hal87b, HP78a, Han92, Hen69, Her76, Hil68, IB97, ILT69, IS90, IR99, Jac81, Jam68, Jam69, Jen72, Jen92, KHK93, Kac99, Kat69, Kel65, Kin72, Kin75, Kre91, Kub70, KM96, Kus68, LE97, Lan77, Lav89, Lav91, LT89a]. Solution [Le 94, Lee91, LB90, Lin72, Lin74, Lin77b, Liu93, Liu99, LL99, LT68, Man81, Man78, Man80, Man90, MW92, MR98, Mas90, MJ77a, MT75, Mey73b, Mit77, MS97a, ML92, NP81, NdF74, Oet65, Pai74, PS75, PRS90, PCW95, Pre73, Pri77, RR95, Rak76, Rat95, Rei81, Rhe80b, Ric78, Rin92, Rob66, Ros70a, RW95, Ros73, Rus70, RT95, SD77, Say74, SW77, SA73, Sha95, She78, SB93, SW94b, Sle82, Smi68, Smi70, Smi73, SH66, Sto71, Sto68, SW68, Swa72, SS73, Swa74, Tav71, TH98, UTK96, Var73, VWI97, WS93, Wei72a, Wei69, Wei74a, Whi82c, Woo71, Woz74a, Yan94, ZW94, Zha95, dW79, de 68, And79, Atk82, BM87, BGO89, Bak80, Bia89, BW86, BR88, BL87, Cha86, Cle81, Cor81, DE79, DL83, Eva80]. solution [ED82, FR86, Fis82, GL78a, Ger89, GM78b, GGL79, IT80, Ioa81, Ise84, Jac88, JS83, JS85b, JSC91, LK80, LSY87, Lin85, MW86, Mar85, MR83, May84, OR86, RS89, RBC79, Saa83, Sch85, Smi81, Tav78, Vog86, WH86, Whi82b, Zir82, Zla82]. Solutions [AGP94, AGM92, AH76, BR81, Ban83, BC95a, BP74, Bor76, Bry68, Cal74, Cal71, CG98, Car74, Cas66, Che73, Che75, CK92, Cur66, DS68, EH89, Esp94, FHK96, HM76a, Hen70b, HR86, HK91, Hig89b, Hu97, IPS91, JS85a, Jok96, KS97b, Kem75, KL67, KKP94, KK98b, Lam67, LSS94, Laz82, LT67, LNS95, Mah82b, MJ77a, Mid66, Moo77, NT92, NT94, NW97, OPW65, Olv65, OS65, Osb76, Osb65, Rhe93, RR98, RT92, Rul96, San94, Sch72, Sig76, Ste70, Tad91, TT99, hTZ97, Tho68b, Wah77, Wei92a, Whe73c, Wig66, Wu96b, Yam98, Zed66, Zem95, de 78b, vS87, vL92, Bog85, BE79, Bre84, BRD81, Cha84a, CR79, Dor86, FN84, FM90b, GL80a, GC88, HR82, Hof85, Kel81, Lar89, Laz81, LP80, Mar78, MG82, MK80, Par87a, RB79]. solutions [Sch80, Smo83, Ste79b, Vic80, Wil81, Xu89]. Solve [CC78, Geo77, Kau74, DL82, Vat88]. Solvents [DTW78]. Solver [AP95a, Bia93, LP77, OPF97, Gup85a, Gup85b, LNOP82]. Solvers [BFB92, Bör90, Buz77, CP98, NV93, PT97, Rat95, BO86, Osh84, WW93b]. Solving [BR75, BCC78, BW76, Bre73, BP71, BGN70, CSH78, Cho78, DKR68, Ehr71, Gar75, Gay82, GPV78, GG90, GS76, GM97, GM91, Gup75, Han71, HV98, Hie80, Hil76, Hil78, HMT75, HZ94, Iss96, JK94, KMR94, KK93, Li88, Li91, LW92, LZC92, Mar79a, Mat98, McL74, Mey68, Moo99, Nic73a, NK97, Per67b, Phi72, Pla98, Pot93, RC78, SB77a, SEK93, Sco81, Sin77, Sti64, Sun96, SHR96, Swe77, TMV98, VMK97, VP93, Var76, VVC94, VTG91, Was82, WES80, Zla65, APR83, BCL89a, BCL89b, Boj84, DOR88a, DOR88b, Ell84, GL81, Gre84, HL82, IPZ79, LOR85, LMS87, LSY89, Mar79b, McK79, NV83, Saa82, Saa87, Str84, Toi79, Tru84, VG86]. Some [AH74, AH94, AF68, BR70, BR87, BS93a, BE79, BSTW77, Bre73, Bro77, CM65, CC66, CC67, CS96, CV72, DRW89, Den75, Den77c, DS72, DMP75, Die92, DW76b, EG75, Eri86, EL72, For79, FU69, Gay79, Gla69a, GS67, Hed68, Hel76, Hur67, Kac99, KL67, Kov70, LB96b, LR72, Mas73, Mou65, NK97, Not91, Osa90, Osb75, Ost71, Pad97, RS87, Saa92, SS95, SW85, Str67b, SG68, SCWM71, Str73, Str75b, TE98, Tho68b, Var67, Ver90a, Ver91, Win80, Woz74b, Wuy70, Wuy74, Wyn64, YF92, Zha97, Ale84, Bea86, BM89, CW93, Dag86, EH75, FS91, GZ88, Mee80, ME81, SL83b, Zla80]. Sor [Sch73, ALY88, FL71, JV97, Mar94, Wan92]. Sor-Newton [Sch73]. SOR-Secant [Mar94]. Sound [Hoh99]. Sound-Hard [Hoh99]. Source [BG74, Can68, GL96a, GLBN97, GSY92, LTW96, TT95, Tan98]. Sources [LL94, WB98, Bub86, LL95].

Space [BS93a, BEK96, BW93, BM97, Cho78, CF93, Daw91, De 73, ELV94, Fai72, FF91, GM69a, GYF86, Hil76, Hil78, KM99, KW76, KT86, LFB98, LMS97, Mao95, Moo94, Moo99, Per92, Rac68b, TE98, Wet97b, Whe75, Whe77, Whi82c, ASW93, Bak80, Bak82, DPZ87, For79, GO81]. Space-Time [BS93a, KM99]. Spaced [HY95, Sec65b]. Spaces [Alf96, Arc78, Aub68, BN74, BH70, Cap99, Dia77, Die90, Dur83, For66b, Fra78, HJ99, Kim76, KT97, LTW91, Man71, Pal93, RD71a, Sch69a, Van67, Yse86, BM89, Gri87, MM79]. Spacing [Köh95]. Sparse [BB87, Bun74a, Geo74, GL75, Geo76, GPS76, Li88, LS76, PS75, PT79, VVC94, BB88, CM83, DW85, EGLP85, GN83a, GN83b, Mar79b, Sco81, Toi79, Zla80, Zla82]. Spatial [HR88, HR97, JV96, Moo99, Pit78, PS83, Wet97a, HR82]. Special [BCO94, Buc66, BT68, Cas82, MF98, Osb75]. specification [CEv87]. Specified [CK69]. Spectra [CJY91, EW91, Wat86, CM88a, Mil79b]. Spectral [AR95, BM99, Bel97, BYLP98, BCM91, BK99, BBG95, CGS92, CF88, Cas97, CD90, COP98, DF84, DG98b, Dut90, E92, EH92, GGS98, GOQ97, GS96, Gre91, GC95, GL96b, HS68, IIMPL82, ITM91, KP96, KP97, LV90, Ma98a, Ma98b, MQ82a, PR95, Pas82, PW96, QZ92, Rin90, Rin92, Sch90b, Sne96, SW91, Sun99, TE86, TE89, TT87, BCM88, BM89, Bun81, CQ82, Del89, GLT87a, GLT87b, Ker88, Lus86, MT89a, Mil79a, Qua87, Swa79, Tad89, WT88]. Spectral-Viscosity [Sch90b]. Spectrally [VWI97]. Spectrum [Mil92, Bea87, Cha89, Des81, GL83]. Speed [AH74, ST67]. Sphere [BF85, Spi65, Spi67, Str73, Str75b, Tay95, KD83, Swa79, Swa81]. Spherical [GC95, Kir93, KSC99, Tay95]. spherically [SE81]. Spline [Alf96, Arc77, Atk68, BFB92, Bia93, Bia98, BF99, BJS73, BH70, Cox84, DDS80, Dem76, Dem77, Dur77, Fai94, Gra87a, GYF86, GZ88, Gre64, Hal73, HVR88, JS68a, Jer73a, Jer73b, Jia86, KR72, KK81a, KZ66, Kem75, LB98, Lau89, LFB98, Lip73, LT67, Mal77, Mun73, NdF74, PR77, PR98, Ros70b, RS97b, Sch69a, Sch68b, Sch83b, Sec65a, Sun99, YF92, de 88, AS84, APR83, Bea86, BC94a, Bru83, Die82, DS89b, Hak82, Hu96, Lee85, Rag85]. Splines [BC72, Cal71, CS74b, Dah80, DS89b, DS97, Fer74, Güs81a, Höl82, HY95, Jup78, Ker71, Koc84, LS99, Luc74, MS73a, MS73c, Nie74, Pow69, PV79, Ric67, Sch70, Sch74, SLW75, Sch69b, Sch89a, ST76, SV73, SW74a, SW74b, Wer75, Zwa73, de 77, dC90, APS87, ASW93, AD92, Bea82a, Bea82b, DM82, Far85, Gra87b, Luc82, LS73, Mey87b, Pet95, SB88]. Split [BT97, GM69a, KW93, WH86, in 96]. Split-step [WH86]. Splitting [AR95, AR96, CGY98, Geo76, Gup75, Hei93b, Hei98b, KT86, KT97, LTW96, LM79, Pet98a, Spi71, Tan98, Yse86, van80, van84, GK96, Gut83]. Splittings [BN76, May87, de 78b, Nie79]. Spread [ILM^{+}96]. Spuriosity [AGH97]. Spurious [BDM^{+}95, IPS91]. SQP [KS92b]. Square [Dun75, Erd65, Sch66]. Squares [AL91, BV92, BP74, Boc97, BCMM98, BMM99, BN73, BT68, CLMM94, CMM97c, CMM97a, CMMP98, CG98, Car74, CNP93, CG90, CN97, CC96b, CD90, Cli73, DS68, Eld80, GP73, GV80, KP78, Kin72, Lan77, Liu94, MR98, Mor68, Osb65, Osb75, PS77, PCL94, Per67b, Pow69, SM81, Sta99, Sto71, Wei92b, Asc78, Bar88, Cor81, Dem87, DS86, FS91, GN83a, GN83b, GM78b, Kam79b, LP78, LP80, Pai79, PSP89, Saa87, Van85]. Stabilised [SW94b, WS93]. Stability [Ale94, AB86, Bab72, BK78, Bar99, Bec92, BZ88, Boh77, BN83, BF91, BB79, But81, CKS95, Cha84b, CS87, CE97b, DLN83, Dav94, Dej67, DL90, Dun80c, Dus95, Eis66, EM88, FSU85b, Gar71, GT74a, GW74a, Gea78, GHS97, GP98, Gor75, GLT87b, Gre66, Gri86, HL84a, Hal74, Hen70a, HR86, Hof78, HM76b, Hur67, IR99, Jam80, Jan83, Jel76, Jel77, Jel79, JN83, JRB95, KR94a, Kos73, KO79, KK98b, Lam67, Lam71, LL96a, Law66, Law67b, Lin92, MB97, MS65b, Omo77, Ort73, Osb76, Pal93, PS92, Pet98b, Pet87, RS95, RHH99, Rob75, Rob76, SM96, Sch71a, Sid97, Ste65a, SI83, SWC89, SH95, Sym80, Sym81, TS93, Ter98, TK96, Tre72, Tuc69, Var71a, Var71b, Var77, Var80, VS97, Wei89, Wyn66, ddd83, dd86, van81c, BMS79, CLM87]. stability [CLM89, CK85a, Gar88, Jer84, Rua89, Sid86, SJ83, WR85, Whi86b, Wol83, Mil79b]. Stabilization [BDH98, EH92, SK93, BEL89, Pie95]. Stabilized [CS99, Osb79, van79]. Stabilizing [Pet98a]. Stable [Alo97, Bog71, BN85, Bur82, Buz77, Car72, Cas88, CLS86, Cim72, CS94, Coo72, Cra76, Cra78, Cry74, Dut88, Gol98a, Gol98b, GS76, HW81, HL84b, Hen70a, Hil97, Jen76, Kee86, KT86, KL86, KW93, Law67a, Lay83b, Met94, Osb79, Riv75, Tho67, Van81a, Vav96, Wen68, Bur87, BC89, Cha94, Dai92, DN89, EHP87, EGLP85, Fei82, GT85, HB86, HK79, Luc85b, Pai79, Var78, Zir82, SW94a]. Stage [JT70, Nic73a]. Staggered [For90]. Standard [AC96b, JN95]. start [VH93]. Started [SZ90]. Starting [MJ77b, MT68a]. State [CS99, CK92, Hag75, HM76a, KKP94, Lav91, Ric91, SB93, Tar96, Wu96b, ZB96, FN84, Lav89, Tav78]. state-dependent [FN84, Tav78]. States [LGMMV92, Sim75]. Static [HDS97]. Stationary [AK98, Den67b, DG90, FK76, Gir78, Hav87, KJ98, LSS94, LNS95, MS99a, MW94, Osb76, Woz74a, Bey87, Kar82, Lor84, Man82, Par87a, You72b]. Steady [BGS98, CK92, HM76a, Kre91, KKP94, KK98b, Lav91, Ric91, SB93, Tar96, Wu96b, ZB96, Lav89]. Steady-State [CK92, Lav91, SB93, Tar96, Wu96b, Lav89]. Steckin [Gre99]. Steepest [Nas70, WM67]. Stefan [Cia75, Epp82, LT89b, MVV89, Mey71, Mey73a, RBC79, Tar96, Whi82a, Whi82b, Whi86a]. Steffensen [JS68b]. Stein [Mil81a]. Stekloff [Kut72, Kut82]. Stencils [Liu93]. Step [BS94, Cal71, Chr91, DT73, HW97a, Jac84, Jac86, JRF91, JT95, KX94, Kro73, PT97, TT95, Tav71, Ten94, UTK96, ddd83, Bal86, DLN83, EJ87, GT85, HLN83, HP80, Joh88, KL86, KR94a, Lev82, Lev85, SWC89, WW93a, WW93b, WH86, Zen85, Rou82]. Steplength [dR81, SC87]. Stepping [GP98, LSZ97, DDE79, Ewi78, ER82, Rus85]. steps [DLN83, Lin83]. Stepsize [Lin77a, SZ90, Ske76, CLM87, CLM89, CL84, SJ83]. Stepwise [CKS95]. Sticky [BG98b]. Stiefel [Dun74b]. Stieltjes [Hu96, Not91, Wyn68]. Stiff [Ale77, Ale94, AFM90, BH86, BP90, BC90, Cas81b, Enr74, EH76, Hof76, Ise81b, Jel76, Jel77, Jel79, Kno92, Lam81, Lin77a, LW70, LNS95, MPS98, SD77, STW97, Tan98, VIA96, Wat81a, BSW95, FSU85a, Ise84, Joh88, Kre78a, KNB86, Maj84, Maj85, Smi81, SWC89, Ver82, Wat81b]. Stiffly [Coo72, Jen76, Var78]. Stochastic [MT97, MPS98, Pet98a, SM96, Tal91, KP85, KR88, Rüm82]. Stokes [AP95a, AT97, BJK90, BW91, BR85, BCM88, BCM91, Boc97, BCMM98, BMM99, Bof97, BC95b, BM97, CMM97a, CN97, CK98, CFT84, DN89, Dut88, EHP87, FK76, FF93, Gir78, GR82, Gui87, GM93, GL96b, GS91b, HP80, HW98, Hei93b, Hei98b, HR82, HR86, HR88, HR90, Hou90, HW92, HW93, IK93, Kar82, KJ98, KL96, KL97, KR94a, LV90, LL96b, LT98, Lin97, MQ82a, Met94, Osb76, Pan97, Pie95, She92, She95, SW94b, Str84, TV96, WS93, Wee98, Wei93, Yan96, ZH91, ZW94, ZH94]. stopping [Han93b]. Storage [AH82, Was82, GL80b]. Strain [DP98, RS95]. Strand [Her76]. Strang [Got72]. Strang-type [Got72]. Strategies [Bun74b, RC78, de 99, AB87, Enr89, Zla80]. Strategy [Ser99, BT89]. Stratified [CM92]. Stream [BYLP98]. Streamline [TV96].

Stress [San94, Wig87]. Stresses [GH92]. Strict [Bar73, DT73, Les91]. Strictly [Bre75, Bre77, LTW95b]. Strip [CK69, Kra87]. stripe [Mel87]. Strong [Gar88, May87, Ste65a, SSB85]. Strongly [SS93, Say74, Tra96]. Structural [Pet87]. structurally [EGLP85]. Structure [BT68, Cas82, GGM99, Ged81, LVQ95, Row67, KL79]. Structured [AC79, IKP91]. Structures [Bao95, Mel87]. Studies [Don94]. Study [Car99, FP95, HR99, Osb76, Rog93, Ste65a, Van72, FR87b, FR85b, Fri81]. Sturm [GB89, GM98, Joh69, Pru73, PFX95]. Sub [GS96]. Sub-Interval [GS96]. Subcritical [Nel73]. Subdivision [CR97, HJ99, PR98, RS97a]. Subdomain [CG90]. Subdomains [Bad91]. Subdominant [Mid66]. Subject [Jam73, Kel69, CEv87]. submatrix [Dem87]. Suboptimal [Sch79]. subsequences [BDGB83]. Subset [RC66]. Subsets [Bac80]. Subspace [Byr90, HL97, JK94, Jan94, KS94, Rit90, Saa92, de 99]. Subspace-Breaking [Jan94]. Subspaces [Aal72, BS79, Eis67a, Rit90, Ste71c]. substructures [BW86]. Substructuring [AMW99, DSW94, MTF99, PW96]. Subsystems [And83a, And79]. success [SS96b]. Successive [Ber66, Cos79, Dax85, HP75, JR75, MQ82b, CM88a, DS86, Hat82, Jer80, SB88]. Sufficient [Far88, Fer73, Wil70, FM84]. Suitable [Erd65]. Sum [GS96, Rit90, LM79]. Summation [Fav64, MS73a]. Sums [Cli65, Kam76, Rei68a, Kam79a, Kam79b]. Sun [Pai84]. sup [BCM88]. Super [CJY91, Ma98b, Sch89a]. Super-Optimal [CJY91]. Superconductivity [DNW98]. Superconvergence [Dou85, Dup76, ELW91, ELW99, GGOS96, Hu97, KX96, Pan97, SSW96, TXZ89, Wah92, Bak84, DS80, LN82, Zen85].

Superconvergent [Cha86, JP98]. Superlinear [Che98, Mor97, Sor80, Toi79, Win80]. Supersonic [DG98b]. Supersplines [LS97a]. Support [Aal72, Afr70, Fer74]. Supported [Bey92, Can94]. Suppress [AGH97]. suppressing [Hig84]. Suppression [Kar96, LTW95b]. Surface [BZ85, Far66, GC95, Ned78, Str73, DL83, KD83, Pet95, Rom90]. Surfaces [HS97, PCW95, AG87]. Survey [DeL94]. SVD [Vog86]. Swapping [CH97]. switching [SS83]. Symm [Dri99]. Symmetric [BBS73, Buc77, BP71, Bun74b, FR99, Fra73, Fri71, HP77, Hue73, Jen72, Kat69, KS94, Let71b, Let74b, MR77, MP68, NtB90, Par65, Pla98, Say74, SSW96, Ser99, Str67b, WB86, Zha95, Chr91, DNT83, Esp87, GM83, Gen86, KL79, KD83, Mee80, Saa83, SE81, Sco81, TW88]. Symmetries [MQT98, DS89a]. Symmetrize [HV97]. Symmetry [ABGM92, Hon93, LM77, CFQ83, JSC91, WS84]. symmetry-breaking [WS84]. Symplectic [AMSS97, Dun97, Jay96, LSZ97, Mur97, Rei96, SO97]. synthesis [BBT82]. Synthetic [ABDH95, Bro95]. System [Asc77, BCS78, BCMM98, Bou98, Cad73, Cad74, CLMM94, CMM97c, CMM97a, CMMP98, CDMW98, CDMCW99, FP95, GS91a, HR99, Jam71, KL97, LTW95b, Man90, Oet65, Sch98a, STW97, Ter98, TW93, WO96, Bia89, IS90]. Systematic [AR77, Sob79a, Sob79b, Spa71]. Systems [AK84, AGP94, AC90, AGM92, And83a, AOW93, AC79, AL78, BGM96, BCC78, BJZ94, Ben96, BL98, BP74, BDM^{+}95, Bog71, Bre75, Bre77, Bre73, BH86, Buc77, BP71, BP90, BC90, CJR97, CSS97, Cas81b, Cas82, CJY91, CG90, Cha92, Che87, CK92, CW97, Chr91, wECGS94, Cli76, CS98, Dan73, DRV97, Dry78, Eic93, EES83, EW91, GLV91, GP84, Geo74, GM78a, Gol98a, Gol98b, GM66, Gov97, HV98, Hel76, Hie80, Hig89b, Hof76, HS94, Ise81b, JK96, Jay96, JS85a, JX98, KS97a, Kel65, Kel69, KKP94, KK93, Lam81, LL96a, Law67a, Lay83b, Lee91, Li88, LW92, LTW95a, LW70, LB96b, LT68, Ma95, MRW96, Mae82, Man80, MP77, MPS98, Moo77, MQ82b, MF90, OPW65, PS75, PS77, Pao99, Pet86, Qi82, QZ92, Rei96, Rei72, Rob66, Rob75, Rob76, SB77a]. Systems [Sch96, Ser99, She78, SW94b, Sim75, Sta91, Ste65b, SWB90, SH95, Sun99, Swe77, Tal91, VMK97, Van72, Var71a, Var73, Var76, VVC94, VTG91, VA91, WW90, WS93, Wat81a, Wid78, dK86, de 78b, And79, AW83, BCL89a, BCL89b, Boj84, BH93, BS83, BSW95, Cha84a, Cha85a, Cha89, CT86, CR79, DL83, Eva80, FSU85a, GHY83, GV89, GL80a, GL78a, GT85, GK96, GGT82, GLT87a, GLT87b, Gre84, Hig84, Hof78, Ise84, Jer80, Jer84, Jer88, KL86, Lay83a, LeV88, Lev85, LSY87, LSY89, Lus79a, Lus86, Maj84, Maj85, MW86, Moo78, MK80, Par87a, PS88, Rei88, Saa82, Saa83, Saa87, Sch85, Sch89b, SW85, Tav78, VH93, Vil86, Wat81b, Zir82, Zla82]. Szego [Tru86].

T [HL82]. Table [Ged81, Wyn68]. Tadmor [AV99]. Takens [Gov93, Jan94]. Tangent [Drm98b]. Tanh [Hab77]. Tapia [EA91]. Tau [Ort69]. Taxonomy [AMS90]. Taylor [BF91, CM74]. Taylor-Like [CM74]. Tchebycheff [Sti64]. Tearing [Ste65b]. Technique [Mao95, Pfl97, Rin92, RSE69, RW95, Sch73, BX89, Gre84, GLL86, Zla80]. Techniques [ADH99, BCS78, Bog71, CC66, CC67, DRV94, EI95, ELPV93, Gun65, Hei98b, HW97b, Kuz99, ST91, Wei69, Xu96, BSW95, Hei93b, Lev82, Maj84, Maj85, ST78, Sch83c, Zir82]. Temporal [JW91]. Tension [CM99b, Del93]. Tensor [AWY97, Dyk87, SF84]. Term [GTV90, GSY92]. Termination [NR72]. Terms [GL96a, GLBN97, LTW96, TT95, Tan98, Rom90]. Test [EH76, Moo77, MQ82b, Qi82, Stu79, Moo78, MK80, Pan85, Xu89]. tests [SW85]. Tetragonal [LL98]. Tetrahedral [Kří92, MW94, ASW93]. Tetrahedron [MS97b]. th [GK96]. th-order [GK96]. Their [Bak68, DE72, FU69, GGM99, JS85a, LL96a, Shi66, ABPR86, BO83, Bru84, DKW82, GZ88, Kel81, KL86, LK80, PT97, Saa87, Swa81, Wei89]. them [KL82]. Theorem [BP96, Cry67, Gar66, GT74b, Hel74, HV70, Jam71, KLY76, KSC99, Lee92, Mar66, Mil81a, Næv91, Ost70, SW88, Yam98, BL80, DP92, SS96b]. Theorems [DW81, DW82, Hur67, HS99, Loe70, May87, Chu87, DH79, Naz86, Ral80]. Theoretic [CP76, Joe99]. Theoretical [Alb87, Can83, DOR88a, JEH78, JT97, MVV89, NV79, Ols77, Ske82]. Theory [Alb96, Ans67, Bog77, Bro87, CS72, CF93, Dar91, DTW76, DD93, Dup76, EA91, Eld80, Gir74, GM98, Her68, HN75, IPS91, Jan79, JK91, Kat69, Lin77b, LT68, McC94, PP73a, Rau73, Rei74, Rhe68, Rhe80a, Ric66, Rob75, Rob76, SP88, Sch69b, Sch69a, Sch69d, Sch69c, Sid97, Str74, Ter98, VW95, VA91, Wei92b, Wim77, Wyn64, Dor84, FST87, MM84, MMR88, McC85, MP78a, SK87]. Thereof [RC66]. Thermal [Kee95, LP91]. Thermoelasticity [Boi97]. Thickness [Pet99]. Thin [BS96, Kir93, Rip92, Rog93]. Third [Gun74, HS92, IT95, Kee70, SCWM71]. Third-Order [HS92]. Thomas [BV96, Bre92, Mar85]. Three [Bec92, BM99, Bof97, BEK96, Bro95, Che77, CGH91, DSW94, DW99, Dub90, DFJ74, GGS94, GGS98, GC98, JT70, LMS97, Mad71, MR77, Moo99, MS97b, NW97, PS80, PW96, RS92, Var80, VS97, Yin95, de 78b, Atk82, BE88, BR88, Dai92, GT85, Lin85, Par87b, Rom90, Vic80]. Three-Dimensional [Bec92, BM99, Bof97, Bro95, CGH91, DW99, Dub90, GGS94, GGS98, Mad71, NW97, PS80, VS97, Yin95, BE88, BR88, Rom90, Vic80]. three-grid [Par87b]. Three-Level [DFJ74, Dai92]. Three-Part [de 78b]. Three-Stage [JT70]. three-step [GT85]. Threshold [WF97]. Threshold-Accepting [WF97]. Tikhonov [LE97, Neu88, Neu97, Qn99, SEK93]. Time [ARW95, BS93a, Bao95, BM97, Car74, CSH78, CDMW98, CDMCW99, CS98, Den75, Den77b, DM92, Día79, EJ95c, Ewi78, ELV94, GP98, GM97, Gus69, HR86, HR90, Hof76, IPS91, JV96, KM99, Le 94, Lin97, LSZ97, NS93, RR96, Rus85, ST86, TE86, TE89, UTK96, VP93, Wet97b, Yan99, van84, AF86, Bak80, Bal86, BR85, CFT84, DDE79, EJ87, ER82, GG86, Gus80, GK83, Hag86, Hay81b, Jam78, JS87, Joh88, Kre79, Lar89, LeV88, Lev82, Lev85, LFB98, Lus86, Rac73, Tav78, WW93a, WW93b]. Time-Constants [Hof76]. Time-Dependent [ARW95, CS98, Den75, Den77b, GP98, Lin97, van84, BR85, Hag86]. time-discrete [LeV88]. Time-Harmonic [GM97]. Time-Periodic [Car74, VP93]. time-scales [GG86, Gus80]. time-step [Joh88]. Time-Stepping [GP98, Ewi78, DDE79, ER82]. Timestep [JNT90, Lei98]. Timestepping [EYL89]. Tocher [Per76]. Toeplitz [BW93, Cap99, Cha89, CJY91, CY93, CNP93, Eva80, KK93, Mee80, Ser99]. Toeplitz-Block [CNP93]. Toeplitz-Plus-Hankel [KK93]. tool [BN89]. Topological [OT75]. Topologies [Ham64]. Topping [Bac80]. Tori [DL95, Moo96]. Total [CM99a, DV97, Gla69b, GV80, HL84b, Dem87, Tad88]. Total-Variation-Stable [HL84b]. trace [SW82]. Tracking [Dk99, Mao95, CT86, Wom89]. Traction [CMMP98, Lee98]. Transfer [BM96, GJL99, NS93, Pet98b]. Transferable [LM90b]. Transform [Atk72, Che93, Cop90, Cor96, Lau89, Lin92, MG66, Obi90, Sch98b, Spi66, Tay95, Wim77, Ell83]. Transform-Based [Lin92]. Transformation [BK73, LL98, PA73, SL83a, Tsa75, Bun81, DOR88a, DOR88b, OR86]. Transformations [AG68, GA67, GAM71, PS77, Dem83]. Transformed [Koc84]. transformer [NW89]. Transforms [BH70, ZT67]. Transient [KK98a, MS89]. Transition [DN91]. Transitions [Gre94]. Transmission [Ben96, CG97, GHS96]. transonic [MN87]. Transport [AW95, AR95, Asa89, Asa98, ABDH95, BK97a, BK97b, BK98, Ber95, Bro95, CE97b, DVW94, Daw98, Kla98, Mad75, MR98, Pit78, PS83, Pri77, Asa86, DL82, Jer88, JP83, Nel86, NV83, Vic80, VG86]. Transport-Diffusion [Ber95, Jer88]. Trapezoidal [Atk74, CM92, DS72, KR97, Kre78b, LG80, MC90]. Travelling [Wee98]. Treating [GH92, Hor83]. Treatment [BGM96, BK78, Can83, Her75, Jan94, Ver88, BvA94, BMS79, GGT82, Ros81, Rüm82]. Trefftz [Hag75]. trial [RS87]. Triangle [Liu93]. Triangles [Jam77, Nie80, Rip92]. Triangular [BDR92, BG98a, Bre92, CKV98, Hig89b, Kee71, MST88, SB77a, YA95, Her85]. Triangularizing [Hua75]. Triangulated [LS99]. Triangulation [Bör90, CR90, WGS90, PS92]. Triangulations [CMM91, CH97, DM82]. Tricomi [Kat69, Tra77]. Tridiagonal [ABP93, Cas82, FU69, Hel76, Swe77, Zha95, Eva80, HP78b, Mee80, Sch89b]. Trigamma [Spo94]. Trigonometric [Bub78, Cap99, Riv64, SV96, Sni72, Gut83, Var81]. Truncated [Cle64, Coh71, Far85, Vog86]. Truncation [Hos95, Jef69, JS69, JT71, JTW83, Mer66, de 99, Ver78, RBC79, Vog86]. Trust [Car91, CGT89, EA91, Hei98a, Les91, BSS87, CGT88, SSB85, Sor82, Ste83, Var85].

Trust-Region [EA91]. trust-region-based [SSB85]. Tunneling [Rin90]. Turbulence [HR93]. Turning [GR96, VF93, Far88, GR84, MS80]. TVD [GL88]. Two [Abl72, AMW99, AD78, AV99, Asa89, AL96b, Bad91, BR75, BF85, Bel97, BYLP98, BEK96, BK77, Bra94, BC95b, BMP89, BT68, Car72, Cas84, Cas88, CD98, wECGS94, Cia75, CR90, DM77, DWW98, Den74a, Dia77, DW76b, Dup76, DL75, Fai72, FF91, GOQ97, Gla69b, GM66, GW83, Gre91, GYF86, Gun74, GM91, GM93, GL96b, HW97a, HP78a, IIMPL82, IK93, JRF91, JT95, JS85a, JP98, Kel74, KS97b, KT86, KR94b, KT76, KL89, Lan77, Lav91, LT98, LP77, LB97, LFB98, LMS97, Mal98, MR77, MX95, MT75, McC94, Mey71, Mil76, MS99c, Nic73a, Oli74, Ols77, Per92, Per97, Pet92a, Rac73, RR96, Red76, Rei74, Reu92, RSE69, RS92, RW96, Rus77, SSY88, SW77, Sha66, SB93, Sin77, Sta64, Str69b, Str72, Tar96, Ten94, Urb91, Var79]. Two [Wah74, Wat79, Whe73c, Whe74, Whe77, WM67, Xu96, YA95, ZZ98, Zha96, ddd83, dd86, van95, AG87, AM89, Asa86, Asc86, BM87, BCMP91, Ban81, CK85b, DLN83, Doe79, DR84, Dou83, EF89, FKD84, Fer86a, GT85, JS87, JP83, Ked81, KNB86, Lay83a, LT88, LOR85, LM79, MVV89, Mos83, PS88, Ros80, Saa83, Sch80, TD82, Whi79, Zho83]. two- [GT85]. Two-Coupled [IIMPL82]. Two-Dimensional [AV99, Asa89, Bel97, BYLP98, BC95b, BT68, CD98, CR90, GM91, GM93, GL96b, HP78a, IK93, Lav91, LB97, Per97, Pet92a, RR96, SB93, Urb91, YA95, van95, BMP89, AG87, Asa86, EF89, JP83, LT88, Mos83, Zho83]. Two-Grid [DWW98, MX95, Xu96]. Two-Level [AL96b, Bra94, GM66, KL89, LT98, Mal98, McC94, MS99c, Rac73, Zha96]. Two-Phase [Mey71, Tar96, Dou83, MVV89]. Two-Point [Cas88, DM77, DW76b, Dup76, Gla69b, Gre91, JP98, Kel74, KS97b, Lan77, LP77, MT75, Oli74, Ols77, Red76, Reu92, RSE69, Rus77, SW77, Sin77, Wat79, Whe73c, Whe74, Whe77, ddd83, dd86, Asc86, BM87, CK85b, Doe79, DR84, FKD84, Fer86a, Ked81, KNB86, Lay83a, LOR85, Sch80, Whi79]. two-sex [AM89]. Two-Sided [Sha66, ZZ98]. Two-Sphere [BF85]. Two-Stage [Nic73a]. Two-Step [HW97a, JRF91, JT95, DLN83]. Type [AT86, AP95b, Aus71, BN74, Bol72, BSTW77, Can83, Cha92, Che77, Den74a, Ein88, ELPV93, GV91, Gov93, Gri74, Gut93, Hal74, Hua75, Jac81, KS94, LT89a, Loe70, Mal98, MF98, NTT94, OR66b, Per90, SM81, WH97, CLS86, CS87, Dag86, DO79, Egg88, Fei82, FR85b, Got72, IT80, Ioa81, Jam78, Jet82, Kum88, Mun94, Nas84, Pan85, PS88, Que89, SSY88, Ser83, ST86, Vil89, Wu96a].

Ultra [CD98]. Ultraspherical [AF68, PK90]. Unbalanced [Mil69, Str65a]. Unbounded [EW72, Gol93, Gor75, Osb67, Tha78]. uncentered [Ske86]. Unconditional [BT97, in 96]. Unconditionally [Tho67, Dai92, EHP87]. Unconstrained [Hag76, OR72, BN89, KD79, Naz86, SSB85, Sor80]. Undamped [Jen76]. Underdetermined [WW90]. underlying [PS92]. Underwater [Pet92a]. unfolded [Kun88]. Unified [DE72, DD93, Dup76, IPS91, Rhe68, Ros81]. Uniform [Ain96, Ans67, BGJ88, BKP94, Cli76, Dör99b, Dör99a, FT74, Gar93, HR86, JN98, Lam76, LMST69, Mat98, Mou65, MT68b, PH73, RC66, Sch68b, Sch69f, Wei72b, CK85b, DS89b, Far88, Gün88, Kro81, Zen85]. Uniforme [DS89b]. Uniformly [CJR97, FMOS96, HO87, JP98, Kut70, McA66, AF89, Ber86, SO86, Sty89]. Unique [CY77]. Uniqueness [AC96a, Jet87, KSS80, Sch69f, Wei92a, Wu96b, Yam98, Nic82, SW85, SS83, Whi85].

Unit [BT99, BP80, Bru91, GM75, Sch98b, ST92, Tse96, Sch83a]. Unitary [DRV94, EH75]. Unknown [BS87b, Can68, Wat81b]. unrestricted [CO80]. Unstable [BGS98, SK93, Sch83c]. Unsteady [BYLP98, Hei98b, She95, Hei93b]. Unstructured [Abg98, AV99, Cap99]. Unsymmetric [Neu98, EGLP85, Saa82, SSY88]. Updata [AC79]. Update [DW81, DW82, RV74, WW90, DW85]. Updates [BW90b, CM87, Ip87, Mar90b, IT88, JA83]. updating [NO85]. Upon [Bro69, Wyn68]. Upper [Alf96, Don73, Hen70a, Tal91]. Upstream [BJ91]. Upwind [Daw98, Huy95, KR94b, GR82]. Upwind-Mixed [Daw98]. Use [AC83, BK73, BF75, CG73, Gar83, GM97, Heg92, LT68, MG66, Phi72, Rei72, Sec65a, Wei69, Zla82, MP78b, Saa87]. Using [AP95a, AM94, BS97c, BTW93, BL98, Bre98, Buc76, CW76, Car91, Cas88, Con73, DM93, Dia77, DS89a, DW99, Dun74a, DS68, For78, Geo77, Gre99, HP77, HS97, HP78a, Han71, HW92, JT70, Kat69, LS97a, LC98, Liu94, Mas73, MS76, MDH^{+}99, Mou66, Mou68, MT68b, NV88, OR90, PS77, SW94b, SW74a, VMK97, Vas99, WS93, AD92, Die82, Hay81a, Hay81b, MP78a, RS87, Saa83, Sch83c, ST95, Ske86, Swe84]. Uzawa [BPV97, Che98, EG94].

V [MSW94, BKP94, EJ95c]. V-Cycle [BKP94]. Validation [AGP94]. Value [AD78, AL96b, BS96, BR75, BR77, Ban77, BS87b, Bel69, Bia98, BDLN92, BK77, BBS73, CW96, Cal74, Cal76, Cas88, Che73, Che75, Che77, Chu92, DM77, DW76a, DDLT91, Dia77, Dor70, Drm98a, Drm98b, DW76b, Dut90, Eis74, EI95, ED77, EM88, EH76, Faz96, FL71, FR78, FU69, Gai74, GK97, Gea65, Gek75, Gek76, Gla69b, GG76, Gra65, Gre91, Gup75, Her68, DW78, HW97b, HZ94, Jen92, Jes78, JP98, Kel74, KW75, KS97b, KS99, Kin72, Kin75, KMR94, Kub70, Kus68, Lan77, LM90b, Li91, Van76, LR72, MW92, MR85, MJ77a, Mat82, MT75, McL74, MM73, Mil69, Mir71, Mit77, Oli74, Ols77, PS81, Par99, Pol74, PL87, Red76, Reu92, Reu94, RSE69, Ros70a, Rus70, RS75, Rus77, RC78, Sch90a, Sch72, SW77, Sha68]. Value [Sha69, Sho75, Sig76, Sin77, Smi73, SHR96, Tap69, TW74, TKJ94, Var71a, Wah75, Wat79, Wei86, Wer75, Whe73c, Whe74, Whe77, Whi82c, Wig66, WH97, Yam98, Yse86, ZWW94, Zla67, ddd83, dd86, dC90, dW79, van80, Asc86, BM87, BvA94, BC94a, BR88, Bog85, CK85b, Cle81, CM88b, DO79, Doe79, DR84, Dor86, Dou84, DS89a, FKD84, Fer86a, Gar88, Gol89, GLT87a, GLT87b, GGL79, GZ88, Gup85a, Gup85b, Gus81b, HK79, JS83, JB88, Ked81, KP79, KS87, KNB86, Lay83a, LK80, LOR85, LP78, MW86, MR84, Nic87, OR86, Pas79, Pas82, PR82, Rin84, Sch80, Smi81, Sun83, TK84, Ver82, Wei89, Whi79, Zho83]. Valued [Ben95]. Values [Coh71, GK65, GS67, Gus73, Han71, Joh78, MT68a, Wat81a, GMP93]. Vanishing [BP96, MT89a]. Variable [BG71, BS97b, Ban77, BTW93, BS94, Cho78, EW96, For66a, GT74a, GW74a, GW90, GL93, Hou90, Jac86, KW76, KM96, LS72, Pet99, Rac68b, Rou82, TYZ95, TW74, Whe75, Bak80, Bak82, CLM87, CLM89, CS87, CL84, DLN83, Dai92, Jam78, Jam80, LT89b, Lin83, SJ83]. Variable-Coefficient [EW96]. Variable-Elliptic-Vortex [TYZ95]. variable-formula [CL84]. Variable-Step [BS94]. variable-stepsize [CLM87, CLM89, CL84, SJ83]. Variables [BG71, Fai72, FF91, GP73, GYF86, LFB98, Sta64, Str69b, Wei71, BR88, Ros80, TD82]. Variance [CR76, Spa71, Tre72, Owe97]. Variant [GS76, Ger89]. Variation [CM99a, DV97, FP91, HL84b, SZ90, Tad88]. Variational [AV97, BW94, CD98, CC92, CKL91, EES83, Fre90, Fun90, Gre94, HR95, Her75, Hop87, HN75, Joh76, Kin75, KLM94, MA81, MR82, Mey73b, Mit87, Ste71a, TH98, Bea82b, MM84, MMR88, McC84, McC85]. Variations [GW90, GL93, Lee92, MW95, Sim69]. Various [Li91, Luc74]. Varying [BLMPT92]. Vector [BJK90, Ben95, BJN81, CJ76, Dub90, HYS94, PO87, SFS86, Sch89b, Swa79, Swa81]. Vector-Valued [Ben95]. Vectors [Pry84]. Vekua [Eis74]. Velocity [BCMM98, BMM99, ELW91, Hol78, JPT98, PSB97]. Velocity-Flux [BCMM98, BMM99]. Verification [Yam98]. Version [BSK81, BS87a, GC98, IB97, LB97, MS97b, Weg66, BG88, BCMP91, Dor84, Dor86, Eri86, GM89, GHS96, KW85, LB96a, SS96a, Win82]. Vertex [BS97b, VMSB97, BvA94, Duf82]. Vertices [Wal98, PS92]. VI [EJL98]. Via [Con76, SW77, BL87, Cap99, DP92, Dor86, Gar88, LS73, Nas84, de 78b]. Vibrating [BK73]. Vibration [BDM^{+}95, Bra66a, Fai72]. Vibrations [COP98]. Visco [HM76a]. Visco-Plastic [HM76a]. Viscoelastic [San94]. Viscoelasticity [Jok96, SWW97]. Viscoplasticity [Car96]. Viscosity [BCV95, Ma98a, Ma98b, MKT93, RT92, Sch90b, hTZ97, MT89a]. Viscous [BLMPT92, DG90, EK99, Har90, Kar92, RT95, Wet97a, BN85]. Vlasov [CR84, GV89, GLV91, GS91a, HV96, HV98, Sch98a, VTG91, VA91, WO96]. Volterra [LTW91, BK78, BMS79, BW76, Bru83, Bru84, Bru90, CT72, Egg83, FS74, Gar75, HLN83, HL84a, HMT75, HM76b, dHW73, DW74, Hu96, Hu97, Jac81, Jam68, LE97, Lin69, MR83, McK79, SWW97, Tav71, Tav77, Tav78, Wal83, Wol83]. Volume [AV99, BS97b, BCV95, CM90a, CMM91, CCL95, FFLMW99, GV91, JS98, KR94b, LMV96, MTTW99, Mor98, Son98, Sül91, VMSB97]. Vortex [AG85, CL89, CGH91, Hou90, HLK91, RR96, TYZ95, YZ94, Yin95, Hal79, Hal87a]. Vorticity [GM91, HW92, Wal88]. Vries [CK85a].

W [LE72]. Walsh [LLNS96, Sew66b, Wid66]. waste [EYL89]. Water [BHL96, CG93, CE97b, CDMW98, CDMCW99]. Waterhammer [NP81]. Wathen [Mil92]. Wave [Abl72, BS97a, BJR90, BGT97, BL98, CJ96, DL91, Eis67b, GSY92, Haa96, Hig90, Hol78, IB97, Wee98, vL92, FN80, Rau85, Ros81, Wal88]. Wave-Propagation [BL98]. Waveform [BJZ94, JK96, JV96, JV97, Lei98]. Waveguide [GM97]. Wavelet [Can94, DE98, FD98, GC96, Jaf92, SP94]. Wavelet-Based [DE98]. Wavelets [Bey92, DM93]. Waves [Bao95, BHL96, BGS98, CG93, Hig94, JPRT96, CLNP89]. Wavewise [Yan99]. way [Geo80]. Weak [CD98, KPH95, LP80, MT97, Ros73, Ste65a, Tap69, TD82]. Weakly [DW74, KW93, Bru83]. Weierstrass [Hei66]. Weight [Kah69, KU71, LMST69, Mou66, Mou68, MT68b, Not91, Pri79]. Weighted [ADH99, AL91, Gir95, Iss96, Kin72, Mil97, Pra70, YKS78, BM89, DS84, Ioa81]. weighted- [DS84]. weighting [Van85]. Weights [Gun74]. Weinstein [Yam75]. Well [CKL91, GL96a, Kru98, Pel67, Wen68, dM87, HK79]. Well-Balanced [GL96a]. Well-Conditioned [Pel67]. Well-Posed [Wen68, HK79]. Wells [CC92]. Wendroff [MR85, MR84]. Westwater [Her76]. Which [Heg92, Kah67, Dem87, FP86, Jam77, Jam78, Ste75a, Wei69]. while [Die82]. Whittaker [Olv65]. Whose [GP73, JP67]. Wide [ADZ96, Akr99]. Wide-Angle [ADZ96, Akr99]. Widths [KS99]. Wiener [GHK94, LNC97]. Wigner [AR96, Rin92]. Wild [Vav96]. Windowing [AH94]. Winds [IZ81]. within [Li83a, Sch83a]. Without [BDM^{+}95, Wan92, GMP93, Sco81, van79]. Works [CGO07]. wrapping [GS88]. Wrong [Kos73].

Yee [MS94].

Zero [CN97, Kah67, Prü82, Kra87]. Zeros [Dun74a, Gar78, Li83a, RJ74, Shi66, Ste74, Pet89, Sch83a, SL83b, SS80]. Zeta [Hei66]. Zolotarev [IT95].

References

Aalto:1972:RFS

[Aal72] S. K. Aalto. The representation of functionals on subspaces by functionals having certain support sets. SIAM Journal on Numerical Analysis, 9(1):35–39, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1986:SCG

[AB86] U. Ascher and G. Bader. Stability of collocation at Gaussian points. SIAM Journal on Numerical Analysis, 23(2): 412–422, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Allgower:1987:AMI

[AB87] E. L. Allgower and K. Böhmer. Application of the mesh independence principle to mesh refinement strategies. SIAM Journal on Numerical Analysis, 24(6):1335–1351, December 1987. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:1999:RRP

[AB99] Mark Ainsworth and Ivo Babuška. Reliable and robust A posteriori error estimation for singularly perturbed reaction-diffusion problems. SIAM Journal on Numerical Analysis, 36(2):331–353, April 1999. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32187.

Ashby:1995:LAA

[ABDH95] S. F. Ashby, P. N. Brown, M. R. Dorr, and A. C. Hindmarsh. A linear algebraic analysis of diffusion synthetic acceleration for the Boltzmann transport equation. SIAM Journal on Numerical Analysis, 32(1):128– 178, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Abgrall:1998:MRU

[Abg98] Rémi Abgrall. Multiresolution representation in unstructured meshes. SIAM Journal on Numerical Analysis, 35(6):2128–2146, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/31505.

Allgower:1992:ESB

[ABGM92] Eugene L. Allgower, Klaus Böhmer, Kurt Georg, and Rick Miranda. Exploiting symmetry in boundary element methods. SIAM Journal on Numerical Analysis, 29 (2):534–552, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ablow:1972:CFD

[Abl72] C. M. Ablow. A characteristic finite difference method for the wave equation in two dimensions. SIAM Journal on Numerical Analysis, 9(1): 152–164, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Amodio:1993:PFT

[ABP93] P. Amodio, L. Brugnano, and T. Politi. Parallel factorizations for tridiagonal matrices. SIAM Journal on Numerical Analysis, 30(3):813–823, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Allgower:1986:MIP

[ABPR86] E. L. Allgower, K. Böhmer, F.-A. Potra, and W. C. Rheinboldt. A meshindependence principle for operator equations and their discretizations. SIAM Journal on Numerical Analysis, 23(1): 160–169, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Avila:1979:UMH

[AC79] J. H. Avila and P. Concus.

Updata methods for highly structured systems of nonlinear equations. SIAM Journal on Numerical Analysis, 16 (2):260–269, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ahues:1983:UDC

[AC83] Mario Ahués and Françoise Chatelin. The use of defect correction to refine the eigenelements of compact integral operators. SIAM Journal on Numerical Analysis, 20(6):1087–1093, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alexander:1990:RKM

[AC90] Roger K. Alexander and James J. Coyle. Runge–Kutta methods and differentialalgebraic systems. SIAM Journal on Numerical Analysis, 27(3):736–752, June 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1993:MMS

[AC93] Uri M. Ascher and Paul M.

Carter. A multigrid method for shape-from-shading. SIAM Journal on Numerical Analysis, 30(1):102–115, February 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Andre:1996:UNA

[AC96a] N. André and M. Chipot. Uniqueness and nonuniqueness for the approximation of quasilinear elliptic equations. SIAM Journal on Numerical Analysis, 33(5): 1981–1994, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26740.

Arandiga:1996:MSF

[AC96b] Francesc Aràndiga and Vicente F. Candela. Multiresolution standard form of a matrix. SIAM Journal on Numerical Analysis, 33(2): 417–434, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aubry:1998:ICR

[AC98] A. Aubry and P. Chartier. On improving the convergence of Radau IIA methods applied to index 2 DAEs. SIAM Journal on Numerical Analysis, 35 (4):1347–1367, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29653.

Abate:1968:NMG

[AD68] J. Abate and H. Dubner.

A new method for generating power series expansions of functions. SIAM Journal on Numerical Analysis, 5(1): 102–112, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Archer:1978:FMC

[AD78] David Archer and Julio Cesar Diaz. A family of modified collocation methods for second order two point boundary value problems. SIAM Journal on Numerical Analysis, 15 (2):242–254, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arge:1992:GPI

[AD92] Erlend Arge and Morten Dæhlen. Grid point interpolation on finite regions using C^{1} box splines. SIAM Journal on Numerical Analysis, 29(4): 1136–1153, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arandiga:1999:MBW

[ADH99] Francesc Aràndiga, Rosa Donat, and Ami Harten. Multiresolution based on weighted averages of the hat function I: Linear reconstruction techniques. SIAM Journal on Numerical Analysis, 36 (1):160–203, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30877.

Aziz:1982:NAM

[ADK82] A. K. Aziz, M. R. Dorr, and R. B. Kellogg. A new approximation method for the Helmholtz equation in an exterior domain. SIAM Journal on Numerical Analysis, 19(5): 899–908, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Akrivis:1996:EEF

[ADZ96] G. D. Akrivis, V. A. Dougalis, and G. E. Zouraris. Error estimates for finite difference methods for a wide-angle “parabolic” equation. SIAM Journal on Numerical Analysis, 33(6): 2488–2509, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26635.

Addison:1984:PMF

[AE84] C. A. Addison and W. H. Enright. Properties of multistep formulas intended for a class of second order ODEs. SIAM Journal on Numerical Analysis, 21(2):327–339, April 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Asadzadeh:1994:AFE

[AE94] Mohammad Asadzadeh and Kenneth Eriksson. On adaptive finite element methods for Fredholm integral equations of the second kind. SIAM Journal on Numerical Analysis, 31(3):831–855, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ames:1997:KBM

[AE97] Karen A. Ames and James F.

Epperson. A kernel-based method for the approximate solution of backward parabolic problems. SIAM Journal on Numerical Analysis, 34(4):1357–1390, August 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27678.

Askey:1968:PCN

[AF68] Richard Askey and James Fitch. Positivity of the Cotes numbers for some ultraspherical abscissas. SIAM Journal on Numerical Analysis, 5 (2):199–201, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Adjerid:1986:MFE

[AF86] Slimane Adjerid and Joseph E. Flaherty. A moving finite element method with error estimation and refinement for one-dimensional time dependent partial differential equations. SIAM Journal on Numerical Analysis, 23(4): 778–796, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arnold:1989:UAF

[AF89] Douglas N. Arnold and Richard S. Falk. A uniformly accurate finite element method for the Reissner– Mindlin plate. SIAM Journal on Numerical Analysis, 26(6): 1276–1290, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Auzinger:1990:AEE

[AFM90] W. Auzinger, R. Frank, and F. Macsek. Asymptotic error expansions for stiff equations: The implicit Euler scheme. SIAM Journal on Numerical Analysis, 27(1): 67–104, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Afriat:1970:PSM

[Afr70] S. N. Afriat. The progressive support method for convex programming. SIAM Journal on Numerical Analysis, 7 (3):447–457, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atchison:1968:NTR

[AG68] T. A. Atchison and H. L.

Gray. Nonlinear transformations related to the evaluation of improper integrals. II. SIAM Journal on Numerical Analysis, 5(2):451–459, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Anderson:1985:VM

[AG85] Christopher Anderson and Claude Greengard. On vortex methods. SIAM Journal on Numerical Analysis, 22 (3):413–440, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Allgower:1987:APL

[AG87] Eugene L. Allgower and Stefan Gnutzmann. An algorithm for piecewise linear approximation of implicitly defined two-dimensional surfaces. SIAM Journal on Numerical Analysis, 24(2): 452–469, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aves:1997:DEC

[AGH97] Mark A. Aves, David F.

Griffiths, and Desmond J.

Higham. Does error control suppress spuriosity? SIAM Journal on Numerical Analysis, 34(2):756–778, April 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27698.

Allgower:1992:MRC

[AGM92] Eugene L. Allgower, Kurt Georg, and Rick Miranda. The method of resultants for computing real solutions of polynomial systems. SIAM Journal on Numerical Analysis, 29(3):831–844, June 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alefeld:1994:ENV

[AGP94] G. Alefeld, A. Gienger, and F. Potra. Efficient numerical validation of solutions of nonlinear systems. SIAM Journal on Numerical Analysis, 31(1): 252–260, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1983:PCC

[AGS83] K. Atkinson, I. Graham, and I. Sloan. Piecewise continuous collocation for integral equations. SIAM Journal on Numerical Analysis, 20(1):172– 186, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alefeld:1974:CSS

[AH74] G. Alefeld and J. Herzberger.

On the convergence speed of some algorithms for the simultaneous approximation of polynomial roots. SIAM Journal on Numerical Analysis, 11(2):237–243, April 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Allinger:1976:ASD

[AH76] Glenn Allinger and Myron Henry. Approximate solutions of differential equations with deviating arguments. SIAM Journal on Numerical Analysis, 13(3):412–426, June 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Armstrong:1977:ARD

[AH77] Ronald D. Armstrong and John W. Hultz. An algorithm for a restricted discrete approximation problem in the L_{1} norm. SIAM Journal on Numerical Analysis, 14 (3):555–565, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Amit:1982:SRP

[AH82] R. Amit and C. Hall. Storage requirements for profile and frontal elimination. SIAM Journal on Numerical Analysis, 19(1):205–218, February 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Apel:1994:MRW

[AH94] Th. Apel and B. Heinrich. Mesh refinement and windowing near edges for some elliptic problem. SIAM Journal on Numerical Analysis, 31 (3):695–708, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ahlfors:1966:RCF

[Ahl66] Lars V. Ahlfors. Remarks on Carleman’s formula for functions in a halfplane. SIAM Journal on Numerical Analysis, 3(2):183–187, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:1996:PBD

[Ain96] Mark Ainsworth. A preconditioner based on domain decomposition for hp finite-element approximation on quasi-uniform meshes. SIAM Journal on Numerical Analysis, 33(4):1358– 1376, August 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25822.

Albrecht:1984:EIM

[AK84] P. Albrecht and M. P. Klein.

Extrapolated iterative methods for linear systems. SIAM Journal on Numerical Analysis, 21(1):192–201, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Axelsson:1990:MDE

[AK90] O. Axelsson and L. Kolotilina.

Monotonicity and discretization error estimates. SIAM Journal on Numerical Analysis, 27(6):1591–1611, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Axelsson:1998:MRA

[AK98] O. Axelsson and I. E. Kaporin. Minimum residual adaptive multilevel finite element procedure for the solution of nonlinear stationary problems. SIAM Journal on Numerical Analysis, 35(3):1213–1229, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28642.

Akrivis:1999:FDM

[Akr99]  Georgios Akrivis.  Finite

difference methods for the wide-angle “parabolic” equation. SIAM Journal on Numerical Analysis, 36(1):317– 329, February 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30851.

Aziz:1978:FEA

[AL78] A. K. Aziz and S. Leventhal. Finite element approximation for first order systems. SIAM Journal on Numerical Analysis, 15(6):1103–1111, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aziz:1991:WLS

[AL91] A. K. Aziz and J.-L. Liu. A weighted least squares method for the backwardforward heat equation. SIAM Journal on Numerical Analysis, 28(1):156–167, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1996:SRM

[AL96a] Uri M. Ascher and Ping Lin. Sequential regularization methods for higher index DAEs with constraint singularities: The linear index-2 case. SIAM Journal on Numerical Analysis, 33(5): 1921–1940, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25325.

Axelsson:1996:TLM

[AL96b] O. Axelsson and W. Layton. A two-level method for the discretization of nonlinear boundary value problems. SIAM Journal on Numerical Analysis, 33(6): 2359–2374, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24710.

Albrecht:1987:NTA

[Alb87] Peter Albrecht. A new theoretical approach to Runge– Kutta methods. SIAM Journal on Numerical Analysis, 24 (2):391–406, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Albrecht:1996:RKT

[Alb96] Peter Albrecht. The Runge– Kutta theory in a nutshell. SIAM Journal on Numerical Analysis, 33(5): 1712–1735, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26087.

Alexander:1977:DIR

[Ale77] Roger Alexander. Diagonally implicit Runge–Kutta methods for stiff O.D.E.’s. SIAM Journal on Numerical Analysis, 14(6):1006–1021, December 1977. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Alefeld:1984:CSI

[Ale84] G. Alefeld. On the convergence of some intervalarithmetic modifications of Newton’s method. SIAM Journal on Numerical Analysis, 21(2):363–372, April 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alexander:1987:AIP

[Ale87] J. C. Alexander. Average intersection and pivoting densities. SIAM Journal on Numerical Analysis, 24(1):129– 146, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alexander:1994:SRK

[Ale94] Roger K. Alexander. Stability of Runge–Kutta methods for stiff ordinary differential equations. SIAM Journal on Numerical Analysis, 31(4): 1147–1168, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alfeld:1985:MPI

[Alf85]  Peter Alfeld.  Multivari-

ate perpendicular interpolation. SIAM Journal on Numerical Analysis, 22(1):95– 106, February 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alfeld:1996:ULB

[Alf96] Peter Alfeld. Upper and lower bounds on the dimension of multivariate spline spaces. SIAM Journal on Numerical Analysis, 33(2): 571–588, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alvarez:1992:ISS

[ALM92] Luis Alvarez, Pierre-Louis Lions, and Jean-Michel Morel. Image selective smoothing and edge detection by nonlinear diffusion. II. SIAM Journal on Numerical Analysis, 29 (3):845–866, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alouges:1997:NAC

[Alo97] François Alouges. A new algorithm for computing liquid crystal stable configurations: The harmonic mapping case. SIAM Journal on Numerical Analysis, 34 (5):1708–1726, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26424.

Adams:1988:ASI

[ALY88] Loyce M. Adams, Randall J.

LeVeque, and David M.

Young. Analysis of the SOR iteration for the 9-point Laplacian. SIAM Journal on Numerical Analysis, 25 (5):1156–1180, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arbogast:1989:FDM

[AM89] Todd Arbogast and Fabio A. Milner. A finite difference method for a two-sex model of population dynamics. SIAM Journal on Numerical Analysis, 26(6):1474–1486, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alvarez:1994:SIR

[AM94] Luis Alvarez and Luis Mazorra. Signal and image restoration using shock filters and anisotropic diffusion. SIAM Journal on Numerical Analysis, 31(2): 590–605, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ashwin:1998:NBF

[AM98] Peter Ashwin and Zhen Mei. A numerical bifurcation function for homoclinic orbits. SIAM Journal on Numerical Analysis, 35 (5):2055–2069, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29816.

Ashby:1990:TCG

[AMS90] Steven F. Ashby, Thomas A.

Manteuffel, and Paul E. Saylor. A taxonomy for conjugate gradient methods. SIAM Journal on Numerical Analysis, 27(6):1542–1568, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Araujo:1997:SMB

[AMSS97] A. L. Araújo, A. Murua, and J. M. Sanz-Serna. Symplectic methods based on decompositions. SIAM Journal on Numerical Analysis, 34 (5):1926–1947, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29212.

Ainsworth:1999:CBE

[AMT99] Mark Ainsworth, William McLean, and Thanh Tran. The conditioning of boundary element equations on locally refined meshes and preconditioning by diagonal scaling. SIAM Journal on Numerical Analysis, 36(6): 1901–1932, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33080.

Achdou:1999:ISP

[AMW99] Yves Achdou, Yvon Maday, and Olof Widlund. Iterative substructuring preconditioners for mortar element methods in two dimensions. SIAM Journal on Numerical Analysis, 36(2):551–580, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32100.

Ahlberg:1966:ALF

[AN66] J. H. Ahlberg and E. N. Nilson. The approximation of linear functionals. SIAM Journal on Numerical Analysis, 3(2):173–182, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Andrus:1979:NSS

[And79] J. F. Andrus. Numerical solution of systems of ordinary differential equations separated into subsystems. SIAM Journal on Numerical Analysis, 16(4):605–611, August 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Andrus:1983:AIS

[And83a] J. F. Andrus. Automatic integration of systems of secondorder ODE’s separated into subsystems. SIAM Journal on Numerical Analysis, 20(4):

815–827, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Andrus:1983:CAP

[And83b] J. F. Andrus. Convergence and accuracy properties of the method of quasilinearization. SIAM Journal on Numerical Analysis, 20(4):828– 839, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Anonymous:1973:CAD

[Ano73] Anonymous. Collection of articles dedicated to the memory of George E. Forsythe. SIAM Journal on Numerical Analysis, 10(2):i–xi + 241–432, 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Anselone:1967:UAT

[Ans67] P. M. Anselone. Uniform approximation theory for integral equations with discontinuous kernels. SIAM Journal on Numerical Analysis, 4 (2):245–253, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Abrahamsson:1982:MDS

[AO82] Leif Abrahamsson and Stanley Osher. Monotone difference schemes for singular perturbation problems. SIAM Journal on Numerical Analysis, 19(5):979–992, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arbogast:1993:NMS

[AOW93] Todd Arbogast, Mandri Obeyesekere, and Mary F. Wheeler. Numerical methods for the simulation of flow in root-soil systems. SIAM Journal on Numerical Analysis, 30(6):1677–1702, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ackerberg:1973:NMH

[AP73] R. C. Ackerberg and J. H. Phillips. A numerical method for highly accelerated laminar boundary-layer flows. SIAM Journal on Numerical Analysis, 10(1):147–160, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alefeld:1983:QCK

[AP83] G. Alefeld and L. Platzöder. A quadratically convergent Krawczyk-like algorithm. SIAM Journal on Numerical Analysis, 20(1):210–219, February 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1987:PIP

[AP87] Kendall E. Atkinson and Florian A. Potra. Projection and iterated projection methods for nonlinear integral equations. SIAM Journal on Numerical Analysis, 24(6):1352– 1373, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1991:PIR

[AP91] Uri M. Ascher and Linda R.

Petzold. Projected implicit Runge–Kutta methods for differential-algebraic equations. SIAM Journal on Numerical Analysis, 28(4):1097– 1120, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Achdou:1995:FSN

[AP95a] Y. Achdou and O. Pironneau. A fast solver for Navier– Stokes equations in the laminar regime using mortar finite element and boundary element methods. SIAM Journal on Numerical Analysis, 32(4): 985–1016, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1995:NSD

[AP95b] Uri M. Ascher and Linda R.

Petzold. The numerical solution of delay-differential-algebraic equations of retarded and neutral type. SIAM Journal on Numerical Analysis, 32(5):1635– 1657, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1983:SBS

[APR83] U. Ascher, S. Pruess, and R. D. Russell. On spline basis selection for solving differential equations. SIAM Journal on Numerical Analysis, 20(1): 121–142, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alfeld:1987:EBQ

[APS87] Peter Alfeld, Bruce Piper, and L. L. Schumaker. An explicit basis for C^{1} quartic bivariate splines. SIAM Journal on Numerical Analysis, 24(4): 891–911, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aird:1977:SSH

[AR77] Thomas J. Aird and John R. Rice. Systematic search in high dimensional sets. SIAM Journal on Numerical Analysis, 14(2):296–312, April 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alefeld:1981:ERF

[AR81] G. Alefeld and J. G. Rokne. On the evaluation of rational functions in interval arithmetic. SIAM Journal on Numerical Analysis, 18(5):862– 870, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arnold:1995:OSM

[AR95] Anton Arnold and Christian Ringhofer. Operator splitting methods applied to spectral discretizations of quantum transport equations. SIAM Journal on Numerical Analysis, 32(6):1876–1894, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arnold:1996:OSM

[AR96] Anton Arnold and Christian Ringhofer. An operator splitting method for the Wigner– Poisson problem. SIAM Journal on Numerical Analysis, 33 (4):1622–1643, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23882.

Arbogast:1989:ASS

[Arb89] Todd Arbogast. Analysis of the simulation of single phase flow through a naturally fractured reservoir. SIAM Journal on Numerical Analysis, 26(1):12–29, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Archer:1977:CSC

[Arc77] David Archer. An O(h^{4}) cubic spline collocation method for quasilinear parabolic equations. SIAM Journal on Numerical Analysis, 14(4):620– 637, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Archer:1978:CSF

[Arc78] David Archer. Collocation in C^{0} spaces for first order hyperbolic equations. I: Optimal order global estimates for quasilinear problems. SIAM Journal on Numerical Analysis, 15(2): 271–281, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arnold:1982:IPF

[Arn82] Douglas N. Arnold. An interior penalty finite element method with discontinuous elements. SIAM Journal on Numerical Analysis, 19(4): 742–760, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1995:IEM

[ARW95] Uri M. Ascher, Steven J. Ruuth, and Brian T. R. Wetton. Implicit-explicit methods for time-dependent partial differential equations. SIAM Journal on Numerical Analysis, 32 (3):797–823, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1969:PCP

[AS69] K. Atkinson and A. Sharma.

A partial characterization of poised Hermite–Birkhoff interpolation problems. SIAM Journal on Numerical Analysis, 6(2):230–235, June 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Abu-Shumays:1980:RSD

[AS80] I. K. Abu-Shumays. Resolution of singularities of the diamond difference approximation. SIAM Journal on Numerical Analysis, 17(6):794– 805, December 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arnold:1984:ACS

[AS84] Douglas N. Arnold and Jukka Saranen. On the asymptotic convergence of spline collocation methods for partial differential equations. SIAM Journal on Numerical Analysis, 21(3):459–472, June 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Allgower:1985:APL

[AS85] Eugene L. Allgower and Phillip H. Schmidt. An algorithm for piecewise-linear approximation of an implicitly defined manifold. SIAM Journal on Numerical Analysis, 22 (2):322–346, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Asadzadeh:1986:AFD

[Asa86] Mohammad Asadzadeh. Analysis of a fully discrete scheme for neutron transport in twodimensional geometry. SIAM Journal on Numerical Analysis, 23(3):543–561, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Asadzadeh:1989:EEE

[Asa89] Mohammad Asadzadeh. L_{p} and eigenvalue error estimates for the discrete ordinates method for two-dimensional neutron transport. SIAM Journal on Numerical Analysis, 26(1):66–87, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Asadzadeh:1998:FEM

[Asa98] Mohammad Asadzadeh. A finite element method for the neutron transport equation in an infinite cylindrical domain. SIAM Journal on Numerical Analysis, 35 (4):1299–1314, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23811.

Ascher:1977:LPA

[Asc77] Uri Ascher. Linear programming algorithms for the Chebyshev solution to a system of consistent linear equations. SIAM Journal on Numerical Analysis, 14(3): 519–526, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1978:DLS

[Asc78] Uri Ascher. Discrete least squares approximations for ordinary differential equations. SIAM Journal on Numerical Analysis, 15(3): 478–496, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1986:CTP

[Asc86]  Uri Ascher.  Collocation

for two-point boundary value problems revisited. SIAM Journal on Numerical Analysis, 23(3):596–609, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1989:NDA

[Asc89] Uri Ascher. On numerical differential algebraic problems with application to semiconductor device simulation. SIAM Journal on Numerical Analysis, 26(3):517–538, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Alfeld:1993:GDS

[ASW93] Peter Alfeld, Larry L. Schumaker, and Walter Whiteley. The generic dimension of the space of C^{1} splines of degree d ≥ 8 on tetrahedral decompositions. SIAM Journal on Numerical Analysis, 30 (3):889–920, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ahues:1986:RMN

[AT86] Mario Ahués and Mauricio Telias. Refinement methods of Newton type for approximate eigenelements of integral operators. SIAM Journal on Numerical Analysis, 23(1): 144–159, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:1997:PEE

[AT97] Mark Ainsworth and J. Tinsley Oden. A Posteriori error estimators for the Stokes and Oseen equations. SIAM Journal on Numerical Analysis, 34 (1):228–245, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26409.

Atkinson:1967:NSF

[Atk67] Kendall E. Atkinson. The numerical solution of Fredholm integral equations of the second kind. SIAM Journal on Numerical Analysis, 4 (3):337–348, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1968:OCN

[Atk68] Kendall E. Atkinson. On the order of convergence of natural cubic spline interpolation. SIAM Journal on Numerical Analysis, 5(1):89–101, March 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1969:NSI

[Atk69]  Kendall Atkinson.  The

numerical solution of integral equations on the halfline. SIAM Journal on Numerical Analysis, 6(3):375– 397, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1972:NEC

[Atk72] Kendall Atkinson. The numerical evaluation of the Cauchy transform on simple closed curves. SIAM Journal on Numerical Analysis, 9 (2):284–299, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1973:NEF

[Atk73] Kendall E. Atkinson. The numerical evaluation of fixed points for completely continuous operators. SIAM Journal on Numerical Analysis, 10(5): 799–807, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1974:NSA

[Atk74] Kendall E. Atkinson. The numerical solution of an Abel integral equation by a product trapezoidal method. SIAM Journal on Numerical Analysis, 11(1):97–101, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1975:CRA

[Atk75] Kendall Atkinson. Convergence rates for approximate eigenvalues of compact integral operators. SIAM Journal on Numerical Analysis, 12 (2):213–222, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1977:NSB

[Atk77] Kendall E. Atkinson. The numerical solution of a bifurcation problem. SIAM Journal on Numerical Analysis, 14 (4):584–599, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1982:NSL

[Atk82] K. E. Atkinson. The numerical solution of Laplace’s equation in three dimensions. SIAM Journal on Numerical Analysis, 19(2): 263–274, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aubin:1968:BAL

[Aub68] Jean-Pierre Aubin. Best approximation of linear operators in Hilbert spaces. SIAM Journal on Numerical Analysis, 5(3):518–521, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Auslaender:1971:NHT

[Aus71] G. J. Auslaender. A novel, Hermite type, interpolation procedure. SIAM Journal on Numerical Analysis, 8(2): 465–472, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Axelsson:1990:AMP

[AV90] O. Axelsson and P. S. Vassilevski. Algebraic multilevel preconditioning methods, II. SIAM Journal on Numerical Analysis, 27(6):1569– 1590, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Atkinson:1993:NEL

[AV93] K. Atkinson and E. Venturino. Numerical evaluation of line integrals. SIAM Journal on Numerical Analysis, 30 (3):882–888, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aubert:1997:VMI

[AV97] Gilles Aubert and Luminita Vese. A variational method in image recovery. SIAM Journal on Numerical Analysis, 34 (5):1948–1979, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29230.

Arminjon:1999:CFV

[AV99] Paul Arminjon and Marie-Claude Viallon. Convergence of a finite volume extension of the Nessyahu–Tadmor scheme on unstructured grids for a two-dimensional linear hyperbolic equation. SIAM Journal on Numerical Analysis, 36(3):738–771, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27889.

Allen:1989:CFD

[AVG89] E. J. Allen, H. D. Victory, Jr., and K. Ganguly. On the convergence of finite-differenced multigroup, discrete-ordinates methods for anisotropically scattering slab media. SIAM Journal on Numerical Analysis, 26(1): 88–106, February 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Avila:1974:FCM

[Avi74] J. H. Avila, Jr. The feasibility of continuation methods for nonlinear equations. SIAM Journal on Numerical Analysis, 11(1):102–122, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aziz:1980:NSH

[AW80] A. K. Aziz and A. Werschulz.

On the numerical solution of Helmholtz’s equation by the finite element method. SIAM Journal on Numerical Analysis, 17(5):681–686, October 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ascher:1983:CSP

[AW83] U. Ascher and R. Weiss. Collocation for singular perturbation problems. I. first order systems with constant coefficients. SIAM Journal on Numerical Analysis, 20(3): 537–557, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arbogast:1995:CMF

[AW95] Todd Arbogast and Mary F.

Wheeler. A characteristicsmixed finite element method for advection-dominated transport problems. SIAM Journal on Numerical Analysis, 32 (2):404–424, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Arbogast:1997:MFE

[AWY97] Todd Arbogast, Mary F. Wheeler, and Ivan Yotov. Mixed finite elements for elliptic problems with tensor coefficients as cell-centered finite differences. SIAM Journal on Numerical Analysis, 34(2):828–852, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26258.

Arbogast:1996:NMF

[AWZ96] Todd Arbogast, Mary F. Wheeler, and Nai-Ying Zhang. A nonlinear mixed finite element method for a degenerate parabolic equation arising in flow in porous media. SIAM Journal on Numerical Analysis, 33(4):1669– 1687, August 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26672.

Axelsson:1979:PIP

[Axe79] O. Axelsson. Preconditioning of indefinite problems by regularization. SIAM Journal on Numerical Analysis, 16(1): 58–69, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1976:ACF

[BA76] I. Babuška and A. K. Aziz.

On the angle condition in the finite element method. SIAM Journal on Numerical Analysis, 13(2):214–226, April 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1971:RCF

[Bab71] Ivo Babuška. The rate of convergence for the finite element method. SIAM Journal on Numerical Analysis, 8 (2):304–315, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1972:NSP

[Bab72] Ivo Babuška. Numerical stability in problems of linear algebra. SIAM Journal on Numerical Analysis, 9(1): 53–77, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Backus:1980:TAC

[Bac80] George E. Backus. The topping algorithm for calculating extrema of polynomials on subsets of R^{n}. SIAM Journal on Numerical Analysis, 17(4):

479–511, August 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Badea:1991:SAM

[Bad91] Lori Badea. On the Schwarz alternating method with more than two subdomains for nonlinear monotone problems. SIAM Journal on Numerical Analysis, 28(1):179– 204, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baines:1991:AMF

[Bai91] M. J. Baines. An analysis of the moving finite-element procedure. SIAM Journal on Numerical Analysis, 28(5):1323– 1349, October 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1968:NCQ

[Bak68] Christopher T. H. Baker. On the nature of certain quadrature formulas and their errors. SIAM Journal on Numerical Analysis, 5(4):783– 804, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1971:DAL

[Bak71] Christopher T. H. Baker. The deferred approach to the limit for eigenvalues of integral equations. SIAM Journal on Numerical Analysis, 8 (1):1–10, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1976:EEF

[Bak76] Garth A. Baker. Error estimates for finite element methods for second order hyperbolic equations. SIAM Journal on Numerical Analysis, 13 (4):564–576, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bakker:1980:NSP

[Bak80] Miente Bakker. On the numerical solution of parabolic equations in a single space variable by the continuous time Galerkin method. SIAM Journal on Numerical Analysis, 17(1):162–177, February 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bakker:1982:GME

[Bak82] Miente Bakker. Galerkin methods for even-order parabolic equations in one space variable. SIAM Journal on Numerical Analysis, 19(3): 571–587, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bakker:1984:ODG

[Bak84]  Miente Bakker.  One-

dimensional Galerkin methods and superconvergence at interior nodal points. SIAM Journal on Numerical Analysis, 21(1):101–110, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bales:1986:HOS

[Bal86] Laurence A. Bales. Higher order single step fully discrete approximations for second order hyperbolic equations with time dependent coefficients. SIAM Journal on Numerical Analysis, 23(1):27–43, February 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1977:MAEb

[Ban77] Randolph E. Bank. Marching algorithms for elliptic boundary value problems. II: The variable coefficient case. SIAM Journal on Numerical Analysis, 14(5):950– 970, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1981:CTM

[Ban81] Randolph E. Bank. A comparison of two multilevel iterative methods for nonsymmetric and indefinite elliptic finite element equations. SIAM Journal on Numerical Analysis, 18(4): 724–743, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1982:AMI

[Ban82] Randolph E. Bank. Analysis of a multilevel inverse iteration procedure for eigenvalue problems. SIAM Journal on Numerical Analysis, 19(5):886–898, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bandle:1983:SND

[Ban83] Catherine Bandle. Solutions of a nonlinear Dirichlet problem for nearly circular domains. SIAM Journal on Numerical Analysis, 20(6):1094– 1098, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bao:1995:FEA

[Bao95] Gang Bao. Finite element approximation of time harmonic waves in periodic structures. SIAM Journal on Numerical Analysis, 32(4):1155– 1169, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barnhill:1965:CQC

[Bar65] Robert E. Barnhill. The convergence of quadratures on complex contours. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):

321–336, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Barnhill:1967:OQ

[Bar67a] Robert E. Barnhill. Optimal quadratures in L^{2}(E_{ρ}). I. SIAM Journal on Numerical Analysis, 4(3):390– 397, September 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barnhill:1967:OQI

[Bar67b] Robert E. Barnhill. Optimal quadratures in L^{2}(E_{ρ}). II. SIAM Journal on Numerical Analysis, 4(4):534– 541, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barnhill:1969:CCC

[Bar69] Robert E. Barnhill. The convergence of complex cubatures. SIAM Journal on Numerical Analysis, 6(1): 82–89, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bard:1970:CGM

[Bar70] Yonathan Bard. Comparison of gradient methods for the solution of nonlinear parameter estimation problems. SIAM Journal on Numerical Analysis, 7(1):157– 186, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrodale:1973:BRA

[Bar73] Ian Barrodale. Best rational approximation and strict quasi-convexity. SIAM Journal on Numerical Analysis, 10 (1):8–12, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barlow:1988:EAI

[Bar88] Jesse L. Barlow. Error analysis and implementation aspects of deferred correction for equality constrained least squares problems. SIAM Journal on Numerical Analysis, 25(6):1340–1358, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrio:1999:SRK

[Bar99] Roberto Barrio. On the A-stability of Runge–Kutta collocation methods based on orthogonal polynomials. SIAM Journal on Numerical Analysis, 36(4):1291– 1303, August 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32098.

Basu:1973:ERF

[Bas73a] N. K. Basu. Errors in the representation of a function by Chebyshev polynomials of the second kind. SIAM Journal on Numerical Analysis, 10 (3):485–488, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Basu:1973:DCS

[Bas73b] N. K. Basu. On double Chebyshev series approximation. SIAM Journal on Numerical Analysis, 10(3): 496–505, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Basso:1982:IML

[Bas82]  Patricio Basso.  Iterative

methods for the localization of the global maximum. SIAM Journal on Numerical Analysis, 19(4): 781–792, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Basso:1985:OSG

[Bas85]  Patricio Basso.  Optimal

search for the global maximum of functions with bounded seminorm. SIAM Journal on Numerical Analysis, 22(5):888–903, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bauer:1974:CGR

[Bau74] F. L. Bauer. Computational graphs and rounding error. SIAM Journal on Numerical Analysis, 11(1):87–96, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In memoriam George Forsythe.

Bjorck:1971:IAC

[BB71] Å. Björck and C. Bowie. An iterative algorithm for computing the best estimate of an orthogonal matrix. SIAM Journal on Numerical Analysis, 8(2):358–364, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bracha-Barak:1974:FPS

[BB74] Amnon Bracha-Barak. A factorization procedure for the solution of multidimensional ellipti partial differential equations. SIAM Journal on Numerical Analysis, 11(5): 887–893, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1979:SCI

[BB79] Kevin Burrage and J. C.

Butcher. Stability criteria for implicit Runge–Kutta methods. SIAM Journal on Numerical Analysis, 16(1):46– 57, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beauwens:1987:SBF

[BB87] R. Beauwens and M. Ben Bouzid. On sparse block factorization iterative methods. SIAM Journal on Numerical Analysis, 24(5):1066– 1076, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beauwens:1988:ECP

[BB88] Robert Beauwens and Mustapha Ben Bouzid. Existence and conditioning properties of sparse approximate block factorizations. SIAM Journal on Numerical Analysis, 25(4): 941–956, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bonnet-Bendhia:1995:SAB

[BBG95] Anne-Sophie Bonnet-Bendhia and Nabil Gmati. Spectral approximation of a boundary condition for an eigenvalue problem. SIAM Journal on Numerical Analysis, 32(4): 1263–1279, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bracha-Barak:1973:SFP

[BBS73] Amnon Bracha-Barak and Paul E. Saylor. A symmetric factorization procedure for the solution of elliptic boundary value problems. SIAM Journal on Numerical Analysis, 10(1):190–206, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barzilai:1982:NII

[BBT82] J. Barzilai and A. Ben-Tal.

Nonpolynomial and inverse interpolation for line search: synthesis and convergence rates. SIAM Journal on Numerical Analysis, 19(6):1263– 1277, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Byrne:1972:LMF

[BC72] George D. Byrne and Donald N. H. Chi. Linear multistep formulas based on g-splines. SIAM Journal on Numerical Analysis, 9(2):316–324, June 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burkowski:1973:NDP

[BC73] F. J. Burkowski and D. D.

Cowan. The numerical derivation of a periodic solution of a second order differential difference equation. SIAM Journal on Numerical Analysis, 10 (3):489–495, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1989:CSD

[BC89] Kevin Burrage and Fred Chipman. Construction of A-stable diagonally implicit multivalue methods. SIAM Journal on Numerical Analysis, 26(2):397–413, April 1989. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Butcher:1990:TER

[BC90] J. C. Butcher and J. R. Cash. Towards efficient Runge–Kutta methods for stiff systems. SIAM Journal on Numerical Analysis, 27 (3):753–761, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bialecki:1994:NEB

[BC94a] Bernard Bialecki and Xiao-Chuan Cai. H^{1}-norm error bounds for piecewise Hermite bicubic orthogonal spline collocation schemes for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 31(4):1128– 1146, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brighi:1994:ACE

[BC94b] Bernard Brighi and Michel Chipot. Approximated convex envelope of a function. SIAM Journal on Numerical Analysis, 31(1):128– 148, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berger:1995:NSO

[BC95a] Fabienne Berger and Jean-François Colombeau. Numerical solutions of one-pressure models in multifluid flows. SIAM Journal on Numerical Analysis, 32(4):1139– 1154, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenier:1995:CPM

[BC95b] Y. Brenier and G.-H. Cottet.

Convergence of particle methods with random rezoning for the two-dimensional Euler and Navier–Stokes equations. SIAM Journal on Numerical Analysis, 32(4):1080– 1097, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buet:1999:NAC

[BC99] C. Buet and S. Cordier. Numerical analysis of conservative and entropy schemes for the Fokker–Planck–Landau equation. SIAM Journal on Numerical Analysis, 36(3): 953–973, June 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32210.

Bartels:1978:CDM

[BCC78] Richard H. Bartels, Andrew R. Conn, and Christakis Charalambous. On Cline’s direct method for solving overdetermined linear systems in the l_{∞} sense. SIAM Journal on Numerical Analysis, 15(2):255–270, April 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bartels:1989:EPM

[BCL89a] Richard H. Bartels, Andrew R. Conn, and Yuying Li. Erratum: “Primal methods are better than dual methods for solving overdetermined linear systems in the L_{∞} sense?” [SIAM J. Numer. Anal. 26 (1989), no. 3, 693–726]. SIAM Journal on Numerical Analysis, 26(5): 1266, October 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [BCL89b].

Bartels:1989:PMB

[BCL89b] Richard H. Bartels, Andrew R. Conn, and Yuying Li. Primal methods are better than dual methods for solving overdetermined linear systems in the L_{∞} sense. SIAM Journal on Numerical Analysis, 26(3):693–726, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [BCL89a].

Bernardi:1988:GIS

[BCM88] Christine Bernardi, Claudio Canuto, and Yvon Maday. Generalized inf-sup conditions for Chebyshev spectral approximation of the Stokes problem. SIAM Journal on Numerical Analysis, 25(6):

1237–1271, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bernardi:1991:SAS

[BCM91] Christine Bernardi, Claudio Canuto, and Yvon Maday. Spectral approximations of the Stokes equations with boundary conditions on the pressure. SIAM Journal on Numerical Analysis, 28(2): 333–362, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1994:ORM

[BCM94] K. Burrage, F. H. Chipman, and P. H. Muir. Order results for mono-implicit Runge– Kutta methods. SIAM Journal on Numerical Analysis, 31 (3):876–891, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bochev:1998:AVF

[BCMM98] P. Bochev, Z. Cai, T. A.

Manteuffel, and S. F. Mc-Cormick. Analysis of velocity-flux first-order system least-squares principles for the Navier–Stokes equations: Part I. SIAM Journal on Numerical Analysis, 35(3):990–1009, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31359.

Babuska:1991:EPV

[BCMP91] I. Babuška, A. Craig, J. Mandel, and J. Pitkäranta. Efficient preconditioning for the p-version finite element method in two dimensions. SIAM Journal on Numerical Analysis, 28(3):624–661, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1994:SFE

[BCO94] Ivo Babuška, Gabriel Caloz, and John E. Osborn. Special finite element methods for a class of second order elliptic problems with rough coefficients. SIAM Journal on Numerical Analysis, 31(4): 945–981, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berger:1975:NSD

[BCR75] Alan E. Berger, Melvyn Ciment, and Joel C. W. Rogers. Numerical solution of a diffusion consumption problem with a free boundary. SIAM Journal on Numerical Analysis, 12(4):646–672, September 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bartels:1978:MTP

[BCS78] Richard H. Bartels, Andrew R. Conn, and James W.

Sinclair. Minimization techniques for piecewise differentiable functions: The l_{1} solution to an overdetermined linear system. SIAM Journal on Numerical Analysis, 15 (2):224–241, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Benharbit:1995:NVC

[BCV95] S. Benharbit, A. Chalabi, and J. P. Vila. Numerical viscosity and convergence of finite volume methods for conservation laws with boundary conditions. SIAM Journal on Numerical Analysis, 32 (3):775–796, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buzbee:1974:DSB

[BD74] B. L. Buzbee and Fred W.

Dorr. The direct solution of the biharmonic equation on rectangular regions and the Poisson equation on irregular regions. SIAM Journal on Numerical Analysis, 11 (4):753–763, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1976:EQE

[BD76] Garth A. Baker and Vassilios A. Dougalis. The effect of quadrature errors on finite element approximations for second order hyperbolic equations. SIAM Journal on Numerical Analysis, 13(4):577– 598, September 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1985:SEM

[BD85] Randolph E. Bank and Craig C. Douglas. Sharp estimates for multigrid rates of convergence with general smoothing and acceleration. SIAM Journal on Numerical Analysis, 22(4): 617–633, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barlow:1990:CAE

[BD90] Jesse Barlow and James Demmel. Computing accurate eigensystems of scaled diagonally dominant matrices. SIAM Journal on Numerical Analysis, 27(3):762–791, June 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boue:1999:MCA

[BD99] Michelle Boué and Paul Dupuis. Markov chain approximations for deterministic control problems with affine dynamics and quadratic cost in the control. SIAM Journal on Numerical Analysis, 36(3):667–695, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32352.

Brezinski:1983:CAE

[BDGB83] C. Brezinski, J. P. Delahaye, and B. Germain-Bonne. Convergence acceleration by extraction of linear subsequences. SIAM Journal on Numerical Analysis, 20(6): 1099–1105, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buzbee:1971:DSD

[BDGG71] B. L. Buzbee, F. W. Dorr, J. A. George, and G. H. Golub. The direct solution of the discrete Poisson equation on irregular regions. SIAM Journal on Numerical Analysis, 8(4):722–736, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barucq:1998:MAB

[BDH98] H. Barucq, F. Delaurens, and B. Hanouzet. Method of absorbing boundary conditions: Phenomena of error stabilization. SIAM Journal on Numerical Analysis, 35(3):1113–1129, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29780.

Bourlard:1992:CSE

[BDLN92] Maryse Bourlard, Monique Dauge, Mbaro-Saman Lubuma, and Serge Nicaise. Coefficients of the singularities for elliptic boundary value problems on domains with conical points III. finite element methods on polygonal domains. SIAM Journal on Numerical Analysis, 29(1):136– 155, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bermudez:1995:FEV

[BDM^{+}95] A. Bermúdez, R. Durán, M. A. Muschietti, R. Rodríguez, and J. Solomin. Finite element vibration analysis of fluid-solid systems without spurious modes. SIAM Journal on Numerical Analysis, 32 (4):1280–1295, August 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1992:AEP

[BDR92] Ivo Babuška, Ricardo Durán, and Rodolfo Rodríguez. Analysis of the efficiency of an A posteriori error estimator for linear triangular finite elements. SIAM Journal on Numerical Analysis, 29(4):947– 964, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Birkhoff:1966:RRA

[BDSW66] G. Birkhoff, C. De Boor, B. Swartz, and B. Wendroff. Rayleigh–Ritz approximation by piecewise cubic polynomials. SIAM Journal on Numerical Analysis, 3(2): 188–203, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brady:1979:SRP

[BE79] Stephen W. Brady and Alan R. Elcrat. Some results on a posteriori error estimation for approximate solutions of second order elliptic problems. SIAM Journal on Numerical Analysis, 16(6):877– 889, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berntsen:1988:CHD

[BE88] Jarle Berntsen and Terje O.

Espelid. On the construction of higher degree threedimensional embedded integration rules. SIAM Journal on Numerical Analysis, 25(1): 222–234, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beauwens:1976:SDD

[Bea76] Robert Beauwens. Semistrict diagonal dominance. SIAM Journal on Numerical Analysis, 13(1):109–112, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beatson:1982:MCA

[Bea82a] R. K. Beatson. Monotone and convex approximation by splines: error estimates and a curve fitting algorithm. SIAM Journal on Numerical Analysis, 19(6):1278–1285, December 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beatson:1982:RRA

[Bea82b] R. K. Beatson. Restricted range approximation by splines and variational inequalities. SIAM Journal on Numerical Analysis, 19(2):372–380, April 1982. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beatson:1986:CSC

[Bea86] R. K. Beatson. On the convergence of some cubic spline interpolation schemes. SIAM Journal on Numerical Analysis, 23(4):903–912, August 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beattie:1987:EAR

[Bea87] Christopher Beattie. An extension of Aronszajn’s rule: slicing the spectrum for intermediate problems. SIAM Journal on Numerical Analysis, 24(4):828–843, August 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beckers:1992:ALN

[Bec92]  J. M. Beckers.  Analyti-

cal linear numerical stability conditions for an anisotropic three-dimensional advectiondiffusion equation. SIAM Journal on Numerical Analysis, 29(3):701–713, June 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bornemann:1996:PEE

[BEK96] Folkmar A. Bornemann, Bodo Erdmann, and Ralf Kornhuber. A Posteriori error estimates for elliptic problems in two and three space dimensions. SIAM Journal on Numerical Analysis, 33(3): 1188–1204, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Belford:1969:IVP

[Bel69] Geneva G. Belford. An initial value problem approach to the solution of eigenvalue problems. SIAM Journal on Numerical Analysis, 6(1): 99–103, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1989:ADM

[BEL89] James H. Bramble, Richard E.

Ewing, and Gang Li. Alternating direction multistep methods for parabolic problems — iterative stabilization. SIAM Journal on Numerical Analysis, 26(4):904– 919, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Belhachmi:1997:NME

[Bel97] Zakaria Belhachmi. Nonconforming mortar element methods for the spectral discretization of two-dimensional fourth-order problems. SIAM Journal on Numerical Analysis, 34(4):1545–1573, August 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28436.

Bendhia:1992:NAB

[Ben92] Hachmi Bendhia. Numerical analysis of a bidimensional hencky problem approximated by a discontinuous finite element method. SIAM Journal on Numerical Analysis, 29(4):1059– 1073, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Benamou:1995:DDM

[Ben95] J.-D. Benamou. A domain decomposition method for the polar factorization of vectorvalued mappings. SIAM Journal on Numerical Analysis, 32(6):1808–1838, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Benamou:1996:DDM

[Ben96] Jean-David Benamou. A domain decomposition method with coupled transmission conditions for the optimal control of systems governed by elliptic partial differential equations. SIAM Journal on Numerical Analysis, 33(6): 2401–2416, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26710.

Berman:1966:MSA

[Ber66] Gerald Berman. Minimization by successive approximation. SIAM Journal on Numerical Analysis, 3(1):123– 133, March 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bernardi:1982:NAP

[Ber82] Christine Bernardi. Numerical approximation of a periodic linear parabolic problem. SIAM Journal on Numerical Analysis, 19(6):1196– 1207, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berger:1986:CUA

[Ber86] Alan E. Berger. A conservative uniformly accurate difference method for a singular perturbation problem in conservation form. SIAM Journal on Numerical Analysis, 23(6):1241–1253, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berger:1987:CGI

[Ber87]  Marsha J. Berger.  On

conservation at grid interfaces. SIAM Journal on Numerical Analysis, 24(5):967– 984, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). ICASE Report No. 84-43, September, 1984.

Bernardi:1989:OFE

[Ber89] Christine Bernardi. Optimal finite-element interpolation on curved domains. SIAM Journal on Numerical Analysis, 26(5):1212–1240, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bermejo:1995:GCA

[Ber95] Rodolfo Bermejo. A Galerkincharacteristic algorithm for transport-diffusion equations. SIAM Journal on Numerical Analysis, 32(2):425–454, April 1995. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beyn:1987:NAP

[Bey87] W.-J. Beyn. On the numerical approximation of phase portraits near stationary points. SIAM Journal on Numerical Analysis, 24(5):1095– 1113, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beylkin:1992:ROB

[Bey92] G. Beylkin. On the representation of operators in bases of compactly supported wavelets. SIAM Journal on Numerical Analysis, 29(6): 1716–1740, December 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beaudet:1975:UBC

[BF75] P. Beaudet and T. Feagin.

The use of back corrections in multistep methods of numerical integration. SIAM Journal on Numerical Analysis, 12(6):895–918, December 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baumgardner:1985:IDT

[BF85] John R. Baumgardner and Paul O. Frederickson. Icosahedral discretization of the two-sphere. SIAM Journal on Numerical Analysis, 22(6): 1107–1115, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brezzi:1991:SHO

[BF91] Franco Brezzi and Richard S. Falk. Stability of higher-order hood-taylor methods. SIAM Journal on Numerical Analysis, 28(3):581–590, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bialecki:1999:OSC

[BF99] Bernard Bialecki and Ryan I. Fernandes. An orthogonal spline collocation alternating direction implicit Crank– Nicolson method for linear parabolic problems on rectangles. SIAM Journal on Numerical Analysis, 36(5): 1414–1434, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31038.

Bialecki:1992:FDS

[BFB92] Bernard Bialecki, Graeme Fairweather, and Karin R. Bennett. Fast direct solvers for piecewise Hermite bicubic orthogonal spline collocation equations. SIAM Journal on Numerical Analysis, 29(1): 156–173, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boffi:1999:CME

[BFGP99] D. Boffi, P. Fernandes, L. Gastaldi, and I. Perugia. Computational models of electromagnetic resonators: Analysis of edge element approximation. SIAM Journal on Numerical Analysis, 36 (4):1264–1290, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31853.

Bernstein:1996:FOE

[BFO96] Barry Bernstein, Kathleen A.

Feigl, and Elwood T. Olsen. A first-order exactly incompressible finite element for axisymmetric fluid flow. SIAM Journal on Numerical Analysis, 33(5):1736– 1758, October 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24309.

Bacopoulos:1971:RPM

[BG71] A. Bacopoulos and B. Gaff.

On the reduction of a problem of minimization of n + 1 variables to a problem of one variable. SIAM Journal on Numerical Analysis, 8(1): 97–103, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Birkhoff:1974:OFP

[BG74] Garrett Birkhoff and Surender Gulati. Optimal fewpoint discretizations of linear source problems. SIAM Journal on Numerical Analysis, 11(4):700–728, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1988:VFE

[BG88] I. Babuška and B. Q. Guo. The h-p version of the finite element method for domains with curved boundaries. SIAM Journal on Numerical Analysis, 25(4):837– 861, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barles:1995:SPC

[BG95] Guy Barles and Christine Georgelin. A simple proof of convergence for an approximation scheme for computing motions by mean curvature. SIAM Journal on Numerical Analysis, 32(2): 484–500, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bernardi:1998:LRO

[BG98a] C. Bernardi and V. Girault. A local regularization operator for triangular and quadrilateral finite elements. SIAM Journal on Numerical Analysis, 35(5): 1893–1916, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29376.

Brenier:1998:SPS

[BG98b] Yann Brenier and Emmanuel Grenier. Sticky particles and scalar conservation laws. SIAM Journal on Numerical Analysis, 35(6):2317– 2328, December 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31735.

Bamberger:1988:NDU

[BGJ88] A. Bamberger, J. C. Guillot, and P. Joly. Numerical diffraction by a uniform grid. SIAM Journal on Numerical Analysis, 25(4):753– 783, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bahi:1996:PTC

[BGM96] J. Bahi, E. Griepentrog, and J. C. Miellou. Parallel treatment of a class of differential-algebraic systems. SIAM Journal on Numerical Analysis, 33(5): 1969–1980, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25810.

Buzbee:1970:DMS

[BGN70] B. L. Buzbee, G. H. Golub, and C. W. Nielson. On direct methods for solving Poisson’s equations. SIAM Journal on Numerical Analysis, 7 (4):627–656, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1989:RNS

[BGO89] I. Babuška, B. Q. Guo, and J. E. Osborn. Regularity and numerical solution of eigenvalue problems with piecewise analytic data. SIAM Journal on Numerical Analysis, 26(6):1534–1560, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bultelle:1998:UGD

[BGS98] Matthieu Bultelle, Magali Grassin, and Denis Serre. Unstable Godunov discrete profiles for steady shock waves. SIAM Journal on Numerical Analysis, 35(6): 2272–2297, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31228.

Bamberger:1997:DDM

[BGT97] Alain Bamberger, Roland Glowinski, and Quang Huy Tran. A domain decomposition method for the acoustic wave equation with discontinuous coefficients and grid change. SIAM Journal on Numerical Analysis, 34(2):603–639, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26151.

Bramble:1965:FDA

[BH65] J. H. Bramble and B. E.

Hubbard. A finite difference analog of the Neumann problem for Poisson’s equation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):1– 14, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Bramble:1968:EBR

[BH68] J. H. Bramble and B. E. Hubbard. Effects of boundary regularity on the discretization error in the fixed membrane eigenvalue problem. SIAM Journal on Numerical Analysis, 5(4):835–863, December 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1970:ELF

[BH70] J. H. Bramble and S. R.

Hilbert. Estimation of linear functionals on Sobolev spaces with application to Fourier transforms and spline interpolation. SIAM Journal on Numerical Analysis, 7(1): 112–124, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1971:AEI

[BH71] Christopher T. H. Baker and Graham S. Hodgson. Asymptotic expansions for integration formulas in one or more dimensions. SIAM Journal on Numerical Analysis, 8(2): 473–480, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brown:1975:BCC

[BH75] J. A. Brown and M. S. Henry. Best Chebyshev composite approximation. SIAM Journal on Numerical Analysis, 12(3):336–344, June 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Braess:1983:NCP

[BH83] D. Braess and W. Hackbusch. A new convergence proof for the multigrid method including the V -cycle. SIAM Journal on Numerical Analysis, 20(5):967–975, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brown:1986:MFM

[BH86] Peter N. Brown and Alan C. Hindmarsh. Matrix-free methods for stiff systems of ODE’s. SIAM Journal on Numerical Analysis, 23(3): 610–638, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brasey:1993:HER

[BH93] V. Brasey and E. Hairer. Halfexplicit Runge–Kutta methods for differential-algebraic systems of index 2. SIAM Journal on Numerical Analysis, 30(2):538–552, April 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beale:1996:CBI

[BHL96] J. Thomas Beale, Thomas Y.

Hou, and John Lowengrub. Convergence of a boundary integral method for water waves. SIAM Journal on Numerical Analysis, 33 (5):1797–1843, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24575.

Brezzi:1999:PEA

[BHM^{+}99] F. Brezzi, T. J. R. Hughes, L. D. Marini, A. Russo, and E. Süli. A priori error analysis of residual-free bubbles for advection-diffusion problems. SIAM Journal on Numerical Analysis, 36(6): 1933–1948, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34236.

Bramble:1968:DAD

[BHZ68] J. H. Bramble, B. E. Hubbard, and M. Zlamal. Discrete analogues of the Dirichlet problem with isolated singularities. SIAM Journal on Numerical Analysis, 5(1): 1–25, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ben-Israel:1966:EBG

[BI66] Adi Ben-Israel. On error bounds for generalized inverses. SIAM Journal on Numerical Analysis, 3(4):585– 592, December 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babovsky:1989:CPN

[BI89] Hans Babovsky and Reinhard Illner. A convergence proof for Nanbu’s simulation method for the full Boltzmann equation. SIAM Journal on Numerical Analysis, 26(1):45–65, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bialecki:1989:SNM

[Bia89]  Bernard Bialecki.  Sinc-

Nyström method for numerical solution of a dominant system of Cauchy singular integral equations given on a piecewise smooth contour. SIAM Journal on Numerical Analysis, 26(5):1194– 1211, October 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bialecki:1993:FDD

[Bia93] Bernard Bialecki. A fast domain decomposition Poisson solver on a rectangle for Hermite bicubic orthogonal spline collocation. SIAM Journal on Numerical Analysis, 30(2): 425–434, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bialecki:1998:CAO

[Bia98] Bernard Bialecki. Convergence analysis of orthogonal spline collocation for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 35(2): 617–631, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30540.

Ben-Israel:1966:ICG

[BIC66] Adi Ben-Israel and Dan Cohen. On iterative computation of generalized inverses and associated projections. SIAM Journal on Numerical Analysis, 3(3):410– 419, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bickart:1977:ESP

[Bic77] Theodore A. Bickart. An efficient solution process for implicit Runge–Kutta methods. SIAM Journal on Numerical Analysis, 14(6):1022– 1027, December 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Birkhoff:1983:SPH

[Bir83] Garrett Birkhoff. Salutation to Peter Henrici on his 60th birthday: September 13, 1983. SIAM Journal on Numerical Analysis, 20(6):??, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //www.jstor.org/stable/
                                        pdf/2157141.pdf. Volume

20(6) is dedicated to Peter Henrici.

Brenier:1991:UDM

[BJ91] Yann Brenier and Jérôme Jaffré. Upstream differencing for multiphase flow in reservoir simulation. SIAM Journal on Numerical Analysis, 28 (3):685–696, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berthold:1993:NEB

[BJ93] Dietmar Berthold and Peter Junghanns. New error bounds for the quadrature method for the solution of Cauchy singular integral equations. SIAM Journal on Numerical Analysis, 30(5):1351–1372, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Butcher:1997:IDI

[BJ97] J. C. Butcher and Z. Jackiewicz. Implementation of diagonally implicit multistage integration methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 34(6): 2119–2141, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28250.

Baker:1990:PSV

[BJK90] Garth A. Baker, Wadi N.

Jureidini, and Ohannes A. Karakashian. Piecewise solenoidal vector fields and the Stokes problem. SIAM Journal on Numerical Analysis, 27(6):1466–1485, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bosarge:1973:RGP

[BJMT73] W. E. Bosarge, Jr., O. G.

Johnson, R. S. McKnight, and W. P. Timlake. The Ritz–Galerkin procedure for nonlinear control problems. SIAM Journal on Numerical Analysis, 10(1):94–111, March 1973. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burley:1981:VOP

[BJN81] S. K. Burley, S. O. John, and J. Nuttall. Vector orthogonal polynomials. SIAM Journal on Numerical Analysis, 18(5): 919–924, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bjorstad:1983:FNS

[Bjø83] Petter Bjørstad. Fast numerical solution of the biharmonic Dirichlet problem on rectangles. SIAM Journal on Numerical Analysis, 20(1): 59–71, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bamberger:1990:SOA

[BJR90] Alain Bamberger, Patrick Joly, and Jean E. Roberts. Second-order absorbing boundary conditions for the wave equation: A solution for the corner problem. SIAM Journal on Numerical Analysis, 27 (2):323–352, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bosarge:1973:DMA

[BJS73] W. E. Bosarge, Jr., O. G. Johnson, and C. L. Smith. A direct method approximation to the linear parabolic regulator problem over multivariate spline bases. SIAM Journal on Numerical Analysis, 10 (1):35–49, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bellen:1994:CWR

[BJZ94] A. Bellen, Z. Jackiewicz, and M. Zennaro. Contractivity of waveform relaxation Runge–Kutta iterations and related limit methods for dissipative systems in the maximum norm. SIAM Journal on Numerical Analysis, 31(2): 499–523, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Banks:1973:CEV

[BK73] D. O. Banks and G. J.

Kurowski. Computation of eigenvalues for vibrating beams by use of a prufer transformation. SIAM Journal on Numerical Analysis, 10(5):918–932, October 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1975:NEE

[BK75] I. Babuška and R. B. Kellogg. Nonuniform error estimates for the finite element method. SIAM Journal on Numerical Analysis, 12 (6):868–875, December 1975.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brabston:1977:NMS

[BK77] D. C. Brabston and H. B. Keller. A numerical method for singular two point boundary value problems. SIAM Journal on Numerical Analysis, 14(5):779–791, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1978:SRN

[BK78] Christopher T. H. Baker and Malcolm S. Keech. Stability regions in the numerical treatment of Volterra integral equations. SIAM Journal on Numerical Analysis, 15 (2):394–417, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrett:1997:FEAa

[BK97a] John W. Barrett and Peter Knabner. Finite element approximation of the transport of reactive solutes in porous media. Part 1: Error estimates for nonequilibrium adsorption processes. SIAM Journal on Numerical Analysis, 34(1):201–227, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/24902.

Barrett:1997:FEAb

[BK97b] John W. Barrett and Peter Knabner. Finite element approximation of the transport of reactive solutes in porous media. Part II: Error estimates for equilibrium adsorption processes. SIAM Journal on Numerical Analysis, 34(2):455–479, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25819.

Beyn:1997:NCH

[BK97c] W.-J. Beyn and J.-M. Kleinkauf.

The numerical computation of homoclinic orbits for maps. SIAM Journal on Numerical Analysis, 34 (3):1207–1236, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28169.

Barrett:1998:IEB

[BK98] John W. Barrett and Peter Knabner. An improved error bound for a Lagrange– Galerkin method for contaminant transport with nonlipschitzian adsorption kinetics. SIAM Journal on Numerical Analysis, 35(5): 1862–1882, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30151.

Bernardi:1999:SED

[BK99] C. Bernardi and A. Karageorghis. Spectral element discretization of the circular driven cavity, Part I: The Laplace equation. SIAM Journal on Numerical Analysis, 36(5):1435– 1465, October 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32379.

Bramble:1994:UCM

[BKP94] James H. Bramble, Do Y. Kwak, and Joseph E. Pasciak. Uniform convergence of multigrid V-Cycle iterations for indefinite and nonsymmetric problems. SIAM Journal on Numerical Analysis, 31(6):1746–1763, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrar:1980:FTA

[BL80] R. B. Barrar and H. L. Loeb. Fundamental theorem of algebra for monosplines and related results. SIAM Journal on Numerical Analysis, 17 (6):874–882, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bowers:1987:NSS

[BL87] Kenneth L. Bowers and John Lund. Numerical solution of singular Poisson problems via the sinc-Galerkin method. SIAM Journal on Numerical Analysis, 24(1):36–51, February 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beyer:1992:AOD

[BL92] R. P. Beyer and R. J. Leveque. Analysis of a onedimensional model for the immersed boundary method. SIAM Journal on Numerical Analysis, 29(2):332–364, April 1992. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrett:1994:FEA

[BL94] John W. Barrett and W. B.

Liu. Finite element approximation of the parabolic p-Laplacian. SIAM Journal on Numerical Analysis, 31(2): 413–428, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berger:1998:AMR

[BL98] Marsha J. Berger and Randall J. LeVeque. Adaptive mesh refinement using wave-propagation algorithms for hyperbolic systems. SIAM Journal on Numerical Analysis, 35(6):

2298–2316, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31597.

Bernardi:1992:FEA

[BLMPT92] Christine Bernardi, Frédéric Laval, Brigitte Métivet, and Bernadette Pernaud-Thomas. Finite element approximation of viscous flows with varying density. SIAM Journal on Numerical Analysis, 29 (5):1203–1243, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1987:FMN

[BM87] I. Babuška and V. Majer. The factorization method for the numerical solution of twopoint boundary value problems for linear ODEs. SIAM Journal on Numerical Analysis, 24(6):1301–1334, December 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burke:1988:IAC

[BM88] James V. Burke and Jorge J. Moré. On the identification of active constraints. SIAM Journal on Numerical Analysis, 25(5):1197–1211, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bernardi:1989:PSW

[BM89] Christine Bernardi and Yvon Maday. Properties of some weighted Sobolev spaces and application to spectral approximations. SIAM Journal on Numerical Analysis, 26(4): 769–829, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Balla:1996:TBC

[BM96] Katalin Balla and Roswitha März. Transfer of boundary conditions for DAEs of index 1. SIAM Journal on Numerical Analysis, 33(6): 2318–2332, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24234.

Burie:1997:MMS

[BM97] J. B. Burie and M. Marion.

Multilevel methods in space and time for the Navier– Stokes equations. SIAM Journal on Numerical Analysis, 34 (4):1574–1599, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26798.

Belgacem:1999:CSF

[BM99] Faker Ben Belgacem and Yvon Maday. Coupling spectral and finite elements for second order elliptic three-dimensional equations. SIAM Journal on Numerical Analysis, 36(4):1234– 1263, August 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31919.

Bochev:1999:AVF

[BMM99] Pavel Bochev, Thomas A. Manteuffel, and Stephen F. McCormick. Analysis of velocity-flux least-squares principles for the Navier– Stokes equations: Part II. SIAM Journal on Numerical Analysis, 36(4):1125– 1144, August 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32497.

Brezzi:1989:TDE

[BMP89] Franco Brezzi, Luisa Donatella Marini, and Paola Pietra. Two-dimensional exponential fitting and applications to drift-diffusion models. SIAM Journal on Numerical Analysis, 26(6):1342– 1355, December 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1979:RSN

[BMS79] Christopher T. H. Baker, Athena Makroglou, and Edward Short. Regions of stability in the numerical treatment of Volterra integro-differential equations. SIAM Journal on Numerical Analysis, 16 (6):890–910, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1973:GRL

[BN73] James H. Bramble and Joachim A. Nitsche. A generalized ritz-least-squares method for Dirichlet problems. SIAM Journal on Numerical Analysis, 10(1): 81–93, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barnhill:1974:RKF

[BN74] Robert E. Barnhill and Gregory M. Nielson. Reproducing kernel functions for Sard spaces of type B. SIAM Journal on Numerical Analysis, 11 (1):37–44, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berman:1976:CS

[BN76] Abraham Berman and Michael Neumann. Consistency and splittings. SIAM Journal on Numerical Analysis, 13 (6):877–888, December 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boland:1983:SFE

[BN83] J. M. Boland and R. A. Nicolaides. Stability of finite elements under divergence constraints. SIAM Journal on Numerical Analysis, 20(4): 722–731, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boland:1985:SSL

[BN85] J. M. Boland and R. A. Nicolaides. Stable and semistable low order finite elements for viscous flows. SIAM Journal on Numerical Analysis, 22 (3):474–492, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Byrd:1989:TAQ

[BN89] Richard H. Byrd and Jorge Nocedal. A tool for the analysis of quasi-Newton methods with application to unconstrained minimization. SIAM Journal on Numerical Analysis, 26(3):727–739, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bourlard:1991:ABE

[BNP91] M. Bourlard, S. Nicaise, and L. Paquet. An adapted boundary element method for the Dirichlet problem in polygonal domains. SIAM Journal on Numerical Analysis, 28(3):728–743, June 1991.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Byrd:1987:GCC

[BNY87] Richard H. Byrd, Jorge Nocedal, and Ya Xiang Yuan. Global convergence of a class of quasi-Newton methods on convex problems. SIAM Journal on Numerical Analysis, 24 (5):1171–1190, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1983:GFE

[BO83] I. Babuška and J. E. Osborn. Generalized finite element methods: their performance and their relation to mixed methods. SIAM Journal on Numerical Analysis, 20 (3):510–536, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenier:1986:ARS

[BO86] Yann Brenier and Stanley Osher. Approximate Riemann solvers and numerical flux functions. SIAM Journal on Numerical Analysis, 23 (2):259–273, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1987:EEE

[BO87] I. Babuška and J. E. Osborn. Estimates for the errors in eigenvalue and eigenvector approximation by Galerkin methods, with particular attention to the case of multiple eigenvalues. SIAM Journal on Numerical Analysis, 24(6): 1249–1276, December 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenier:1988:DOS

[BO88] Yann Brenier and Stanley Osher. The discrete one-sided Lipschitz condition for convex scalar conservation laws. SIAM Journal on Numerical Analysis, 25(1):8–23, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bochev:1997:ALS

[Boc97] Pavel B. Bochev. Analysis of least-squares finite element methods for the Navier– Stokes equations. SIAM Journal on Numerical Analysis, 34 (5):1817–1844, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27600.

Boffi:1997:TDF

[Bof97]  Daniele Boffi.  Three-

dimensional finite element methods for the Stokes problem. SIAM Journal on Numerical Analysis, 34(2): 664–670, April 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27019.

Boggs:1971:SNS

[Bog71] Paul T. Boggs. The solution of nonlinear systems of equations by A-stable integration techniques. SIAM Journal on Numerical Analysis, 8 (4):767–785, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boggs:1977:ABS

[Bog77]  Paul T. Boggs.  An al-

gorithm, based on singular perturbation theory, for ill-conditioned minimization problems. SIAM Journal on Numerical Analysis, 14 (5):830–843, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bogomolny:1985:FSM

[Bog85] Alexander Bogomolny. Fundamental solutions method for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 22(4): 644–669, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bohl:1977:SIN

[Boh77] Erich Bohl. On a stability inequality for nonlinear operators. SIAM Journal on Numerical Analysis, 14(2):

242–253, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boillat:1997:LFE

[Boi97] Eric Boillat. On the locking of the finite element method in thermoelasticity. SIAM Journal on Numerical Analysis, 34(6): 2319–2334, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27101.

Bojanczyk:1984:CSL

[Boj84] A. Bojańczyk. Complexity of solving linear systems in different models of computation. SIAM Journal on Numerical Analysis, 21(3):591–603, June 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boland:1972:CPT

[Bol72] W. Robert Boland. The convergence of product-type quadrature formulas. SIAM Journal on Numerical Analysis, 9(1):6–13, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Borowsky:1976:CBC

[Bor76] Michael S. Borowsky. Characterizations of best complex Chebyshev approximate solutions of av = b. SIAM Journal on Numerical Analysis, 13 (3):324–336, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Borgers:1990:TAF

[Bör90] Christoph Börgers. A triangulation algorithm for fast elliptic solvers based on domain imbedding. SIAM Journal on Numerical Analysis, 27 (5):1187–1196, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bourdarias:1998:ASS

[Bou98] C. Bourdarias. Approximation of the solution to a system modeling heatless adsorption of gases. SIAM Journal on Numerical Analysis, 35(1):13–30, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24824.

Boyd:1977:NBM

[Boy77] David W. Boyd. Nonconvergence in Bairstow’s method. SIAM Journal on Numerical Analysis, 14(3):571–574, June 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brooks:1967:AEE

[BP67] John D. Brooks and David A.

Pope. Asymptotic error estimates and the numerical solution of the equations of orbital motion. SIAM Journal on Numerical Analysis, 4 (3):446–456, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bunch:1971:DMS

[BP71] J. R. Bunch and B. N. Parlett. Direct methods for solving symmetric indefinite systems of linear equations. SIAM Journal on Numerical Analysis, 8(4):639–655, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Berman:1974:CIM

[BP74] Abraham Berman and Robert J. Plemmons. Cones and iterative methods for best least squares solutions of linear systems. SIAM Journal on Numerical Analysis, 11(1): 145–154, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brutman:1980:ECC

[BP80] L. Brutman and A. Pinkus. On the Erdős conjecture concerning minimal norm interpolation on the unit circle. SIAM Journal on Numerical Analysis, 17(3):373–375, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenan:1989:NSH

[BP89] Kathryn E. Brenan and Linda R. Petzold. The numerical solution of higher index differential/algebraic equations by implicit methods. SIAM Journal on Numerical Analysis, 26(4):976– 996, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1990:ORR

[BP90] Kevin Burrage and Linda Petzold. On order reduction for Runge–Kutta methods applied to differential/algebraic systems and to stiff systems of ODEs. SIAM Journal on Numerical Analysis, 27(2): 447–456, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1996:GCT

[BP96] Christopher T. H. Baker and Christopher A. H. Paul. A global convergence theorem for a class of parallel continuous explicit Runge– Kutta methods and vanishing lag delay differential equations. SIAM Journal on Numerical Analysis, 33 (4):1559–1576, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25141.

Barrodale:1972:DCA

[BPR72] I. Barrodale, M. J. D. Powell, and F. D. K. Roberts. The differential correction algorithm for rational ↕_{∞}- approximation. SIAM Journal on Numerical Analysis, 9 (3):493–504, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bonnans:1992:AME

[BPTZ92] J. Frédéric Bonnans, Eliane R. Panier, André L. Tits, and Jian L. Zhou. Avoiding the maratos effect by means of a nonmonotone line search. II. inequality constrained problems — feasible iterates. SIAM Journal on Numerical Analysis, 29(4):1187– 1202, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1997:AIU

[BPV97] James H. Bramble, Joseph E. Pasciak, and Apostol T. Vassilev. Analysis of the inexact Uzawa algorithm for saddle point problems. SIAM Journal on Numerical Analysis, 34(3):1072–1092, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27334.

Beauwens:1976:ECP

[BQ76] Robert Beauwens and Lena Quenon. Existence criteria for partial matrix factorizations in iterative methods. SIAM Journal on Numerical Analysis, 13(4):615–643, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bressan:1986:ACC

[BQ86] N. Bressan and A. Quarteroni. Analysis of Chebyshev collocation methods for parabolic equations. SIAM Journal on Numerical Analysis, 23(6):1138–1154, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baker:1970:EBS

[BR70] Christopher T. H. Baker and Pauline A. Radcliffe. Error bounds for some Chebyshev methods of approximation and integration. SIAM Journal on Numerical Analysis, 7(2):317–327, June 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrodale:1973:IAD

[BR73] I. Barrodale and F. D. K.

Roberts. An improved algorithm for discrete l_{1} linear approximation. SIAM Journal on Numerical Analysis, 10(5): 839–848, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1975:MSC

[BR75] Randolph E. Bank and Donald J. Rose. An O(n^{2}) method for solving constant coefficient boundary value problems in two dimensions. SIAM Journal on Numerical Analysis, 12(4):529–540, September 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1977:MAEa

[BR77] Randolph E. Bank and Donald J. Rose. Marching algorithms for elliptic boundary value problems. I: The constant coefficient case. SIAM Journal on Numerical Analysis, 14(5):792–829, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1978:EEA

[BR78a] I. Babuška and W. C. Rheinboldt. Error estimates for adaptive finite element computations. SIAM Journal on Numerical Analysis, 15(4): 736–754, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrodale:1978:EAD

[BR78b] I. Barrodale and F. D. K. Roberts. An efficient algorithm for discrete l_{1} linear approximation with linear constraints. SIAM Journal on Numerical Analysis, 15 (3):603–611, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1980:PSN

[BR80] R. E. Bank and D. J.

Rose. Parameter selection for Newton-like methods applicable to nonlinear partial differential equations. SIAM Journal on Numerical Analysis, 17(6):806–822, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1981:PEA

[BR81] Ivo Babuška and Werner C.

Rheinboldt. A Posteriori error analysis of finite element solutions for one-dimensional problems. SIAM Journal on Numerical Analysis, 18(3): 565–589, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bernardi:1985:CFE

[BR85] Christine Bernardi and Geneviève Raugel. A conforming finite element method for the time-dependent Navier– Stokes equations. SIAM Journal on Numerical Analysis, 22 (3):455–473, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1987:SEE

[BR87] Randolph E. Bank and Donald J. Rose. Some error estimates for the box method. SIAM Journal on Numerical Analysis, 24(4): 777–787, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Blum:1988:NST

[BR88] E. K. Blum and G. J. Reid. On the numerical solution of three-dimensional boundary value problems by separation of variables. SIAM Journal on Numerical Analysis, 25(1):75–90, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1966:EED

[Bra66a] J. H. Bramble. Error estimates for difference methods in forced vibration problems. SIAM Journal on Numerical Analysis, 3(1):1–12, March 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brauer:1966:MCG

[Bra66b] Alfred Brauer. A method for the computation of the greatest root of a nonnegative matrix. SIAM Journal on Numerical Analysis, 3 (4):564–569, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Braaten:1980:IMN

[Bra80] Eric Braaten. An improved measure of the nonuniformity of the distribution of a sequence. SIAM Journal on Numerical Analysis, 17(1):31– 32, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brannigan:1985:DCA

[Bra85]  Michael Brannigan.  Dis-

crete Chebyshev approximation with linear constraints. SIAM Journal on Numerical Analysis, 22(1):1–15, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brandt:1994:RQA

[Bra94] Achi Brandt. Rigorous quantitative analysis of multigrid, I: Constant coefficients twolevel cycle with L_{2}-norm. SIAM Journal on Numerical Analysis, 31(6):1695– 1730, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Butler:1981:ESF

[BRD81] J. P. Butler, J. A. Reeds, and S. V. Dawson. Estimating solutions of first kind integral equations with nonnegative constraints and optimal smoothing. SIAM Journal on Numerical Analysis, 18 (3):381–397, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brent:1973:SEA

[Bre73] Richard P. Brent. Some efficient algorithms for solving systems of nonlinear equations. SIAM Journal on Numerical Analysis, 10(2): 327–344, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1975:MNE

[Bre75] Philip Brenner. Maximumnorm estimates of the rate of convergence of dissipative difference schemes for strictly hyperbolic systems. SIAM Journal on Numerical Analysis, 12(5):726–740, October 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1977:EDS

[Bre77] Philip Brenner. L_{p}-estimates of difference schemes for strictly hyperbolic systems with nonsmooth data. SIAM Journal on Numerical Analysis, 14(6):1126–1144, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenier:1984:AMS

[Bre84] Yann Brenier. Averaged multivalued solutions for scalar conservation laws. SIAM Journal on Numerical Analysis, 21(6):1013–1037, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1989:OON

[Bre89] Susanne C. Brenner. An optimal-order nonconforming multigrid method for the biharmonic equation. SIAM Journal on Numerical Analysis, 26(5):1124–1138, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1992:MAL

[Bre92]  Susanne C. Brenner.  A

multigrid algorithm for the lowest-order Raviart–Thomas mixed triangular finite element method. SIAM Journal on Numerical Analysis, 29 (3):647–678, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1993:NMM

[Bre93] Susanne C. Brenner. A nonconforming mixed multigrid method for the pure displacement problem in planar linear elasticity. SIAM Journal on Numerical Analysis, 30(1): 116–135, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1998:OCS

[Bre98] Susanne C. Brenner. Overcoming corner singularities using multigrid methods. SIAM Journal on Numerical Analysis, 35(5):1883– 1892, October 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30802.

Brown:1969:QCN

[Bro69]  Kenneth M. Brown.  A

quadratically convergent Newtonlike method based upon Gaussian elimination. SIAM Journal on Numerical Analysis, 6 (4):560–569, December 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brown:1977:SCI

[Bro77] R. Leonard Brown. Some characteristics of implicit multistep multi-derivative integration formulas. SIAM Journal on Numerical Analysis, 14(6):982–993, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brown:1987:LCT

[Bro87] Peter N. Brown. A local convergence theory for combined inexact-Newton/finite-difference projection methods. SIAM Journal on Numerical Analysis, 24(2):

407–434, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brown:1995:LAD

[Bro95] Peter N. Brown. A linear algebraic development of diffusion synthetic acceleration for three-dimensional transport equations. SIAM Journal on Numerical Analysis, 32 (1):179–214, February 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brutman:1978:LFP

[Bru78] L. Brutman. On the Lebesgue function for polynomial interpolation. SIAM Journal on Numerical Analysis, 15(4): 694–704, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brutman:1980:PRP

[Bru80] L. Brutman. On the polynomial and rational projections in the complex plane. SIAM Journal on Numerical Analysis, 17(3):366–372, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brunner:1983:NSC

[Bru83]  Hermann Brunner.  Non-

polynomial spline collocation for Volterra equations with weakly singular kernels. SIAM Journal on Numerical Analysis, 20(6):1106– 1119, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brunner:1984:ICM

[Bru84] Hermann Brunner. Iterated collocation methods and their discretizations for Volterra integral equations. SIAM Journal on Numerical Analysis, 21(6):1132–1145, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brunner:1990:NSN

[Bru90] Hermann Brunner. On the numerical solution of nonlinear Volterra–Fredholm integral equations by collocation methods. SIAM Journal on Numerical Analysis, 27(4): 987–1000, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brutman:1991:PEU

[Bru91] L. Brutman. Polynomial extrapolation from [−1, 1] to the unit disc. SIAM Journal on Numerical Analysis, 28(2): 573–579, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bryan:1968:ASN

[Bry68] Charles A. Bryan. Approximate solutions to nonlinear integral equations. SIAM Journal on Numerical Analysis, 5(1):151–155, March 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barrow:1976:EGH

[BS76] David L. Barrow and A. H.

Stroud. Existence of Gauss harmonic interpolation formulas. SIAM Journal on Numerical Analysis, 13(1): 18–26, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bruch:1977:HAD

[BS77] John C. Bruch, Jr. and James M. Sloss. Harmonic approximation with Dirichlet data on doubly connected regions. SIAM Journal on Numerical Analysis, 14(6):994– 1005, December 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bavely:1979:ACR

[BS79] Connice A. Bavely and G. W.

Stewart. An algorithm for computing reducing subspaces by block diagonalization. SIAM Journal on Numerical Analysis, 16(2): 359–367, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bube:1983:IRE

[BS83] K. P. Bube and J. C. Strikwerda. Interior regularity estimates for elliptic systems of difference equations. SIAM Journal on Numerical Analysis, 20(4):653–670, August 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1987:OCR

[BS87a] I. Babuška and Manil Suri. The optimal convergence rate of the p-version of the finite element method. SIAM Journal on Numerical Analysis, 24(4): 750–776, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Baszenski:1987:MFU

[BS87b] Günter Baszenski and Larry L. Schumaker. On a method for fitting an unknown function based on mean-value measurements. SIAM Journal on Numerical Analysis, 24(3): 725–736, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1989:CFE

[BS89a] Randolph E. Bank and L. Ridgway Scott. On the conditioning of finite element equations with highly refined meshes. SIAM Journal on Numerical Analysis, 26(6): 1383–1394, December 1989.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Batterson:1989:DRQ

[BS89b] Steve Batterson and John Smillie. The dynamics of Rayleigh quotient iteration. SIAM Journal on Numerical Analysis, 26(3):624–636, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1992:LRF

[BS92] Ivo Babuška and Manil Suri.

On locking and robustness in the finite element method. SIAM Journal on Numerical Analysis, 29(5):1261– 1293, October 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1993:ASM

[BS93a] Randolph E. Bank and Rafael F. Santos. Analysis of some moving spacetime finite element methods. SIAM Journal on Numerical Analysis, 30(1):1–18, February 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1993:PEE

[BS93b] Randolph E. Bank and R. Kent Smith. A Posteriori error estimates based on hierarchical bases. SIAM Journal on Numerical Analysis, 30(4):921–935, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1994:CVS

[BS94] K. Burrage and P. W. Sharp. A class of variable-step explicit Nordsieck multivalue methods. SIAM Journal on Numerical Analysis, 31 (5):1434–1451, October 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1996:PEE

[BS96] I. Babuška and C. Schwab. A Posteriori error estimation for hierarchic models of elliptic boundary value problems on thin domains. SIAM Journal on Numerical Analysis, 33(1): 221–246, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1997:PEF

[BS97a] Ivo M. Babuška and Stefan A. Sauter. Is the pollution effect of the FEM avoidable for the Helmholtz equation considering high wave numbers. SIAM Journal on Numerical Analysis, 34(6): 2392–2423, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26918.

Balland:1997:ACV

[BS97b] Philippe Balland and Endre Süli. Analysis of the cell-vertex finite volume method for hyperbolic problems with variable coefficients. SIAM Journal on Numerical Analysis, 34 (3):1127–1151, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26488.

Bank:1997:MSU

[BS97c] Randolph E. Bank and R. Kent Smith. Mesh smoothing using A posteriori error estimates. SIAM Journal on Numerical Analysis, 34(3):979–997, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26529.

Babuska:1981:VFE

[BSK81] I. Babuška, B. A. Szabø, and I. N. Katz. The p-version of the finite element method. SIAM Journal on Numerical Analysis, 18(3):515–545, June 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Byrd:1987:TRA

[BSS87] Richard H. Byrd, Robert B.

Schnabel, and Gerald A. Shultz. A trust region algorithm for nonlinearly constrained optimization. SIAM Journal on Numerical Analysis, 24(5):1152–1170, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1977:SCE

[BSTW77] J. H. Bramble, A. H. Schatz, V. Thomee, and L. B. Wahlbin. Some convergence estimates for semidiscrete Galerkin type approximations for parabolic equations. SIAM Journal on Numerical Analysis, 14(2): 218–241, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bell:1985:ANM

[BSW85] John B. Bell, Gregory R. Shubin, and Mary Fanett Wheeler. Analysis of a new method for computing the flow of miscible fluids in a porous medium. SIAM Journal on Numerical Analysis, 22(6):1041–1050, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buttner:1995:MAP

[BSW95] Mario Büttner, Bernhard A. Schmitt, and Rüdiger Weiner. W -methods with automatic partitioning by Krylov techniques for large stiff systems. SIAM Journal on Numerical Analysis, 32(1):260– 284, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buchanan:1968:LSF

[BT68] John E. Buchanan and Donald H. Thomas. On least-squares fitting of twodimensional data with a special structure. SIAM Journal on Numerical Analysis, 5 (2):252–257, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1969:PBD

[BT69] J. H. Bramble and V. Thomee.

Pointwise bounds for discrete Green’s functions. SIAM Journal on Numerical Analysis, 6(4):583–590, December 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1979:RAS

[BT79] Philip Brenner and Vidar Thomée. On rational approximations of semigroups. SIAM Journal on Numerical Analysis, 16(4):683–694, August 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brenner:1980:RAG

[BT80] Philip Brenner and Vidar Thomée. On rational approximations of groups of operators. SIAM Journal on Numerical Analysis, 17(1):119– 125, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boggs:1984:FDF

[BT84] Paul T. Boggs and Jon W. Tolle. A family of descent functions for constrained optimization. SIAM Journal on Numerical Analysis, 21(6): 1146–1161, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boggs:1989:SGC

[BT89] Paul T. Boggs and Jon W. Tolle. A strategy for global convergence in a sequential quadratic programming algorithm. SIAM Journal on Numerical Analysis, 26(3): 600–623, June 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bellen:1997:UCM

[BT97] A. Bellen and L. Torelli. Unconditional contractivity in the maximum norm of diagonally split Runge–Kutta methods. SIAM Journal on Numerical Analysis, 34(2): 528–543, April 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26757.

Bailly:1999:CCP

[BT99] B. Le Bailly and J. P. Thiran. Computing complex polynomial Chebyshev approximants on the unit circle by the real Remez algorithm. SIAM Journal on Numerical Analysis, 36(6): 1858–1877, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33951.

Banks:1993:EVC

[BTW93] H. T. Banks, H. T. Tran, and D. E. Woodward. Estimation of variable coefficients in the Fokker–Planck equations using moving node finite elements. SIAM Journal on Numerical Analysis, 30(6):1574– 1602, December 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bube:1978:CTI

[Bub78] Kenneth P. Bube. C^{m} convergence of trigonometric interpolants. SIAM Journal on Numerical Analysis, 15(6): 1258–1268, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bube:1985:CNI

[Bub85] Kenneth P. Bube. Convergence of numerical inversion methods for discontinuous impedance profiles. SIAM Journal on Numerical Analysis, 22(5):924–946, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bube:1986:NMR

[Bub86]  Kenneth P. Bube.  Nu-

merical methods for reflection inverse problems: convergence and nonimpulsive sources. SIAM Journal on Numerical Analysis, 23(2): 227–258, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buchanan:1966:NPM

[Buc66] Mary Louise Buchanan. Norms of powers of matrices in a special class. SIAM Journal on Numerical Analysis, 3 (4):616–623, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buckley:1976:CMU

[Buc76] A. G. Buckley. Constrained minimization using Powell’s conjugacy approach. SIAM Journal on Numerical Analysis, 13(4):520–535, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Buckley:1977:SCS

[Buc77] A. Buckley. On the solution of certain skew symmetric linear systems. SIAM Journal on Numerical Analysis, 14(3):

566–570, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bunch:1971:ADP

[Bun71] J. R. Bunch. Analysis of the diagonal pivoting method. SIAM Journal on Numerical Analysis, 8(4):656–680, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bunch:1974:ASE

[Bun74a] James R. Bunch. Analysis of sparse elimination. SIAM Journal on Numerical Analysis, 11(5):847–873, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bunch:1974:PPS

[Bun74b] James R. Bunch. Partial pivoting strategies for symmetric matrices. SIAM Journal on Numerical Analysis, 11(3): 521–528, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bunse:1981:CDT

[Bun81] Wolfgang Bunse. A class of diagonal transformation methods for the computation of the spectral radius of a nonnegative irreducible matrix. SIAM Journal on Numerical Analysis, 18(4):693–704, August 1981. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1982:EIA

[Bur82] Kevin Burrage. Efficiently implementable algebraically stable Runge–Kutta methods. SIAM Journal on Numerical Analysis, 19(2): 245–258, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burrage:1987:HOA

[Bur87]  K. Burrage.  High or-

der algebraically stable multistep Runge–Kutta methods. SIAM Journal on Numerical Analysis, 24(1):106– 115, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burke:1990:IAC

[Bur90] Jim Burke. On the identification of active constraints II: The nonconvex case. SIAM Journal on Numerical Analysis, 27(4):1081–1102, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Busch:1985:OI

[Bus85] J. R. Busch. Osculatory interpolation in R^{n}. SIAM Journal on Numerical Analysis, 22 (1):107–113, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Butzer:1964:SA

[But64] P. L. Butzer. Saturation and approximation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):2– 10, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic). See errata [But65].

Butzer:1965:ESA

[But65] P. L. Butzer. Errata: “Saturation and Approximation” [J. Soc. Ind. Appl. Math., Ser. B Numer. anal. 1] (1964), 2-10. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):203, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic). See [But64].

Butcher:1975:OBR

[But75]  J. C. Butcher.  An or-

der bound for Runge–Kutta methods. SIAM Journal on Numerical Analysis, 12(3): 304–315, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Butcher:1981:SPG

[But81] J. C. Butcher. Stability properties for a general class of methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 18(1): 37–44, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

Buzbee:1977:AFP

[Buz77]  B. L. Buzbee.  Applica-

tion of fast Poisson solvers to A-stable marching procedures for parabolic problems. SIAM Journal on Numerical Analysis, 14(2): 205–217, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Braess:1990:MMN

[BV90] D. Braess and R. Verfürth. Multigrid methods for nonconforming finite element methods. SIAM Journal on Numerical Analysis, 27(4): 979–986, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barlow:1992:NDC

[BV92] Jesse L. Barlow and Udaya B. Vemulapati. A note on deferred correction for equality constrained least squares problems. SIAM Journal on Numerical Analysis, 29(1): 249–256, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Braess:1996:PEE

[BV96] D. Braess and R. Verfürth. A Posteriori error estimators for the Raviart–Thomas element. SIAM Journal on Numerical Analysis, 33(6): 2431–2444, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26407.

Babuska:1994:NTV

[BvA94] I. Babuška, T. von Petersdorff, and B. Andersson. Numerical treatment of vertex singularities and intensity factors for mixed boundary value problems for the Laplace equation in r^{3}. SIAM Journal on Numerical Analysis, 31(5):1265–1288, October 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barnhill:1968:EAI

[BW68] Robert E. Barnhill and James A. Wixom. An error analysis for interpolation of analytic functions. SIAM Journal on Numerical Analysis, 5(3):522–529, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Barnhill:1969:EAB

[BW69] Robert E. Barnhill and James A. Wixom. An error analysis for the bivariate interpolation of analytic functions. SIAM Journal on Numerical Analysis, 6(3):450– 457, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bownds:1976:NSN

[BW76] John M. Bownds and Bruce Wood. On numerically solving nonlinear Volterra integral equations with fewer computations. SIAM Journal on Numerical Analysis, 13(5): 705–719, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bjorstad:1986:IMS

[BW86] Petter E. Bjørstad and Olof B. Widlund. Iterative methods for the solution of elliptic problems on regions partitioned into substructures. SIAM Journal on Numerical Analysis, 23(6): 1097–1120, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beatson:1989:PPP

[BW89] R. K. Beatson and H. Wolkowicz. Post-processing piecewise cubics for monotonicity. SIAM Journal on Numerical Analysis, 26(2): 480–502, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Borgers:1990:FED

[BW90a] Christoph Börgers and Olof B. Widlund. On finite element domain imbedding methods. SIAM Journal on Numerical Analysis, 27(4):963– 978, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bourji:1990:LCS

[BW90b] Samih K. Bourji and Homer F.

Walker. Least-change secant updates of nonsquare matrices. SIAM Journal on Numerical Analysis, 27(5):1263– 1294, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bank:1991:PEE

[BW91] Randolph E. Bank and Bruno D. Welfert. A Posteriori error estimates for the Stokes problem. SIAM Journal on Numerical Analysis, 28 (3):591–623, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bottcher:1993:GPM

[BW93] Albrecht Böttcher and Hartmut Wolf. Galerkin–Petrov methods for Bergman space Toeplitz operators. SIAM Journal on Numerical Analysis, 30(3):846–863, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Benassi:1994:PNS

[BW94] M. Benassi and R. E. White. Parallel numerical solution of variational inequalities. SIAM Journal on Numerical Analysis, 31(3):813–830, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bramble:1989:LPP

[BX89] James H. Bramble and Jin Chao Xu. A local post-processing technique for improving the accuracy in mixed finite-element approximations. SIAM Journal on Numerical Analysis, 26(6): 1267–1275, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ben-Yu:1998:FDL

[BYLP98] Guo Ben-Yu and He Li-Ping. The fully discrete Legendre spectral approximation of two-dimensional unsteady incompressible fluid flow in stream function form. SIAM Journal on Numerical Analysis, 35(1):146–176, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/29708.

Byrd:1990:CCO

[Byr90] Richard H. Byrd. On the convergence of constrained optimization methods with accurate Hessian information on a subspace. SIAM Journal on Numerical Analysis, 27(1): 141–153, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babuska:1973:NEF

[BZ73] Ivo Babuška and Milos Zlamal. Nonconforming elements in the finite element method with penalty. SIAM Journal on Numerical Analysis, 10(5): 863–875, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beatson:1985:MPS

[BZ85] R. K. Beatson and Z. Ziegler.

Monotonicity preserving surface interpolation. SIAM Journal on Numerical Analysis, 22(2):401–411, April 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bellen:1988:SPI

[BZ88] Alfredo Bellen and Marino Zennaro. Stability properties of interpolants for Runge– Kutta methods. SIAM Journal on Numerical Analysis, 25 (2):411–432, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cadzow:1973:FAM

[Cad73] James A. Cadzow. A finite algorithm for the minimum l_{∞} solution to a system of consistent linear equations. SIAM Journal on Numerical Analysis, 10(4):607–617, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cadzow:1974:EAP

[Cad74]  James A. Cadzow.  An

efficient algorithmic procedure for obtaining a minimum l_{∞}-norm solution to a system of consistent linear equations. SIAM Journal on Numerical Analysis, 11(6):1151– 1165, December 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Callender:1971:SSM

[Cal71] E. David Callender. Single step methods and low order splines for solutions of ordinary differential equations. SIAM Journal on Numerical Analysis, 8(1):61–66, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caldwell:1974:RCA

[Cal74] C. S. Caldwell. Rates of convergence of approximate solutions of parabolic initialboundary value problems. SIAM Journal on Numerical Analysis, 11(6):1121– 1135, December 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caldwell:1976:CEL

[Cal76] C. S. Caldwell. Convergence estimates for a locally one-dimensional finite difference scheme for parabolic initial-boundary value problems. SIAM Journal on Numerical Analysis, 13(4):514– 519, September 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1968:DUH

[Can68] J. R. Cannon. Determination of an unknown heat source from overspecified boundary data. SIAM Journal on Numerical Analysis, 5(2): 275–286, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cantor:1983:NTP

[Can83] Murray Cantor. Numerical treatment of potential type equations on R^{n}: Theoretical considerations. SIAM Journal on Numerical Analysis, 20(1): 72–85, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Canuto:1986:BCC

[Can86] Claudio Canuto. Boundary conditions in Chebyshev and Legendre methods. SIAM Journal on Numerical Analysis, 23(4):815–831, August 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Candela:1994:CSO

[Can94] Vicente F. Candela. Computation of shift operators in orthonormal compactly supported wavelet bases. SIAM Journal on Numerical Analysis, 31(3):768–787, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Capizzano:1999:KBA

[Cap99] Stefano Serra Capizzano. A Korovkin-based approximation of multilevel Toeplitz matrices (with rectangular unstructured blocks) via multilevel trigonometric matrix spaces. SIAM Journal on Numerical Analysis, 36(6): 1831–1857, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32249.

Carasso:1972:BBE

[Car72] Alfred Carasso. The backward beam equation: Two Astable schemes for parabolic problems. SIAM Journal on Numerical Analysis, 9 (3):406–434, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carasso:1974:LSM

[Car74] Alfred Carasso. On least squares methods for parabolic equations and the computation of time-periodic solutions. SIAM Journal on Numerical Analysis, 11(6):1181– 1192, December 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carter:1991:GCT

[Car91] Richard G. Carter. On the global convergence of trust region algorithms using inexact gradient information. SIAM Journal on Numerical Analysis, 28(1):251–265, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:1992:GMS

[Car92]  Carsten Carstensen.  On

Grau’s method for simultaneous factorization of polynomials. SIAM Journal on Numerical Analysis, 29(2): 601–613, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carasso:1994:OHC

[Car94] Alfred S. Carasso. Overcoming Hölder continuity in ill-posed continuation problems. SIAM Journal on Numerical Analysis, 31(6):1535– 1557, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:1996:CFB

[Car96] Carsten Carstensen. Coupling of FEM and BEM for interface problems in viscoplasticity and plasticity with hardening. SIAM Journal on Numerical Analysis, 33(1):171– 207, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carasso:1999:LNI

[Car99]  Alfred S. Carasso.  Lin-

ear and nonlinear image deblurring: A documented study. SIAM Journal on Numerical Analysis, 36(6): 1659–1689, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32041.

Cassity:1966:SFO

[Cas66] C. R. Cassity. Solutions of the fifth-order Runge–Kutta equations. SIAM Journal on Numerical Analysis, 3 (4):598–606, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cassity:1969:CSF

[Cas69] C. R. Cassity. The complete solution of the fifth order Runge–Kutta equations. SIAM Journal on Numerical Analysis, 6(3):432– 436, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cash:1981:EFC

[Cas81a] J. R. Cash. On the exponential fitting of composite, multiderivative linear multistep methods. SIAM Journal on Numerical Analysis, 18(5): 808–821, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cash:1981:SDE

[Cas81b] J. R. Cash. Second derivative extended backward differentiation formulas for the numerical integration of stiff systems. SIAM Journal on Numerical Analysis, 18(1):21– 36, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

Cash:1982:SBT

[Cas82] J. R. Cash. On the solution of block tridiagonal systems of linear algebraic equations having a special structure. SIAM Journal on Numerical Analysis, 19(6):1220– 1232, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cash:1984:TNF

[Cas84] J. R. Cash. Two new finite difference schemes for parabolic equations. SIAM Journal on Numerical Analysis, 21(3):433–446, June 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cash:1988:NIN

[Cas88] J. R. Cash. On the numerical integration of nonlinear twopoint boundary value problems using iterated deferred corrections. part 2: The development and analysis of highly stable deferred correction formulae. SIAM Journal on Numerical Analysis, 25(4): 862–882, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Casarin:1997:QOS

[Cas97] Mario A. Casarin. Quasioptimal Schwarz methods for the conforming spectral element discretization. SIAM Journal on Numerical Analysis, 34(6):2482– 2502, December 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29228.

Cannon:1966:NSS

[CC66] J. R. Cannon and Maria M.

Cecchi. The numerical solution of some biharmonic problems by mathematical programming techniques. SIAM Journal on Numerical Analysis, 3(3):451–466, September 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1967:NES

[CC67] J. R. Cannon and Maria M.

Cecchi. Numerical experiments on the solution of some biharmonic problems by mathematical programming techniques. SIAM Journal on Numerical Analysis, 4 (2):147–154, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Charalambous:1978:EMS

[CC78] C. Charalambous and A. R.

Conn. An efficient method to solve the minimax problem directly. SIAM Journal on Numerical Analysis, 15(1): 162–187, February 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coleman:1984:LCQ

[CC84] Thomas F. Coleman and Andrew R. Conn. On the local convergence of a quasi-Newton method for the nonlinear programming problem. SIAM Journal on Numerical Analysis, 21(4):755– 769, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chipot:1992:NAV

[CC92] Michel Chipot and Charles Collins. Numerical approximations in variational problems with potential wells. SIAM Journal on Numerical Analysis, 29(4):1002– 1019, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1994:EEF

[CC94] Zhangxin Chen and Bernardo Cockburn. Error estimates for a finite element method for the drift-diffusion semiconductor device equations. SIAM Journal on Numerical Analysis, 31(4):1062– 1089, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caselles:1996:SM

[CC96a] V. Caselles and B. Coll. Snakes in movement. SIAM Journal on Numerical Analysis, 33(6):2445–2456, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27504.

Chen:1996:LSS

[CC96b] Xuzhou Chen and Moody T.

Chu. On the least squares solution of inverse eigenvalue problems. SIAM Journal on Numerical Analysis, 33(6): 2417–2430, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26474.

Cockburn:1995:CFV

[CCL95] B. Cockburn, F. Coquel, and P. G. LeFloch. Convergence of the finite volume method for multidimensional conservation laws. SIAM Journal on Numerical Analysis, 32(3): 687–705, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1967:CPH

[CD67] J. R. Cannon and Jim Douglas, Jr. The Cauchy problem for the heat equation. SIAM Journal on Numerical Analysis, 4(3):317–336, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1977:ASC

[CD77] J. R. Cannon and P. DuChateau.

Approximating the solution to the Cauchy problem for Laplace’s equation. SIAM Journal on Numerical Analysis, 14(3):473–483, June 1977.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carey:1985:GFM

[CD85] Graham F. Carey and Hung T. Dinh. Grading functions and mesh redistribution. SIAM Journal on Numerical Analysis, 22(5):1028–1040, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cohn:1988:ODP

[CD88] Stephen E. Cohn and Dick P. Dee. Observability of discretized partial differential equations. SIAM Journal on Numerical Analysis, 25(3): 586–617, June 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chu:1990:PGM

[CD90] Moody T. Chu and Kenneth R. Driessel. The projected gradient method for least squares matrix approximations with spectral constraints. SIAM Journal on Numerical Analysis, 27(4): 1050–1060, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cessenat:1998:AUW

[CD98] Olivier Cessenat and Bruno Despres. Application of an ultra weak variational formulation of elliptic PDEs to the two-dimensional Helmholtz problem. SIAM Journal on Numerical Analysis, 35 (1):255–299, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28587.

Catte:1995:MSM

[CDK95] Francine Catté, Françoise Dibos, and Georges Koepfler. A morphological scheme for mean curvature motion and applications to anisotropic diffusion and motion of level sets. SIAM Journal on Numerical Analysis, 32(6):1895– 1909, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chippada:1999:FEA

[CDMCW99] S. Chippada, C. N. Dawson, M. L. Martínez-Canales, and M. F. Wheeler. Finite element approximations to the system of shallow water equations, Part II: Discrete-time a priori error estimates. SIAM Journal on Numerical Analysis, 36 (1):226–250, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31415.

Chippada:1998:FEA

[CDMW98] S. Chippada, C. N. Dawson, M. L. Martinez, and M. F.

Wheeler. Finite element approximations to the system of shallow water equations I: Continuous-time a priori error estimates. SIAM Journal on Numerical Analysis, 35(2):692–711, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29651.

Cai:1999:ONG

[CDS99] Xiao-Chuan Cai, Maksymilian Dryja, and Marcus Sarkis. Overlapping nonmatching grid mortar element methods for elliptic problems. SIAM Journal on Numerical Analysis, 36(2):581–606, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32358.

Cowsar:1996:PEM

[CDW96] Lawrence C. Cowsar, Todd F. Dupont, and Mary F. Wheeler. A Priori estimates for mixed finite element approximations of second-order hyperbolic equations with absorbing boundary conditions. SIAM Journal on Numerical Analysis, 33(2):492–504, April 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1997:FDF

[CE97a] Zhangxin Chen and Richard E.

Ewing. Fully-discrete finite element analysis of multiphase flow in groundwater hydrology. SIAM Journal on Numerical Analysis, 34(6): 2228–2253, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29006.

Chen:1997:SCF

[CE97b] Zhangxin Chen and Richard E.

Ewing. Stability and convergence of a finite element method for reactive transport in ground water. SIAM Journal on Numerical Analysis, 34(3):881–904, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27286.

Cannon:1987:GPD

[CEv87] John R. Cannon, Salvador Pérez Esteva, and John van der Hoek. A Galerkin procedure for the diffusion equation subject to the specification of mass. SIAM Journal on Numerical Analysis, 24 (3):499–515, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carlson:1985:MPB

[CF85] R. E. Carlson and F. N. Fritsch. Monotone piecewise bicubic interpolation. SIAM Journal on Numerical Analysis, 22(2):386–400, April 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Canuto:1988:SAS

[CF88] C. Canuto and D. Funaro. The Schwarz algorithm for spectral methods. SIAM Journal on Numerical Analysis, 25(1):24–40, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carlson:1989:AMP

[CF89] R. E. Carlson and F. N. Fritsch. An algorithm for monotone piecewise bicubic interpolation. SIAM Journal on Numerical Analysis, 26(1): 230–238, February 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Costantini:1990:SPB

[CF90] Paolo Costantini and Ferruccio Fontanella. Shapepreserving bivariate interpolation. SIAM Journal on Numerical Analysis, 27(2): 488–506, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coquel:1993:CFD

[CF93] Frédéric Coquel and Philippe Le Floch. Convergence of finite difference schemes for conservation laws in several space dimensions: A general theory. SIAM Journal on Numerical Analysis, 30(3):675–700, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Canuto:1983:ASB

[CFQ83] Claudio Canuto, Hiroshi Fujii, and Alfio Quarteroni. Approximation of symmetry breaking bifurcations for the Rayleigh convection problem. SIAM Journal on Numerical Analysis, 20(5):873– 884, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Constantin:1984:LTG

[CFT84] P. Constantin, C. Foiaş, and R. Temam. On the large time Galerkin approximation of the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 21(4):615– 634, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1999:CRQ

[CfXmZ99] Raymond H. Chan, Shu fang Xu, and Hao min Zhou. On the convergence rate of a quasi-Newton method for inverse eigenvalue problems. SIAM Journal on Numerical Analysis, 36(2): 436–441, April 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32705.

Concus:1973:UFD

[CG73] Paul Concus and Gene H. Golub. Use of fast direct methods for the efficient numerical solution of nonseparable elliptic equations. SIAM Journal on Numerical Analysis, 10(6):1103–1120, December 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conn:1984:LDI

[CG84] Andrew R. Conn and Nicholas I. M. Gould. On the location of directions of infinite descent for nonlinear programming algorithms. SIAM Journal on Numerical Analysis, 21(6):1162–1179, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chang:1990:SGL

[CG90] Ching-Lung Chang and Max D. Gunzburger. A subdomain Galerkin/least squares method for first-order elliptic systems in the plane. SIAM Journal on Numerical Analysis, 27(5):1197–1211, October 1990. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Chandler:1993:CWW

[CG93] G. A. Chandler and I. G.

Graham. The computation of water waves modelled by Nekrasov’s equation. SIAM Journal on Numerical Analysis, 30(4):1041–1065, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cockburn:1996:EEF

[CG96] Bernardo Cockburn and Pierre-Alain Gremaud. Error estimates for finite element methods for scalar conservation laws. SIAM Journal on Numerical Analysis, 33 (2):522–554, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:1997:FBC

[CG97] Carsten Carstensen and Joachim Gwinner. FEM and BEM coupling for a nonlinear transmission problem with Signorini contact. SIAM Journal on Numerical Analysis, 34(5):1845– 1864, October 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28113.

Cao:1998:LSF

[CG98] Yanzhao Cao and Max D. Gunzburger. Least-squares finite element approximations to solutions of interface problems. SIAM Journal on Numerical Analysis, 35 (1):393–405, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30324.

Cottet:1991:CGF

[CGH91] Georges-Henri Cottet, Jonathan Goodman, and Thomas Y. Hou. Convergence of the grid-free point vortex method for the three-dimensional Euler equations. SIAM Journal on Numerical Analysis, 28 (2):291–307, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:2007:MMC

[CGO07] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3 pp. LCCN QA188 .G67

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0737/2007276086-d.html.

Cai:1992:OSF

[CGS92] Wei Cai, David Gottlieb, and Chi-Wang Shu. On one-sided filters for spectral Fourier approximations of discontinuous functions. SIAM Journal on Numerical Analysis, 29(4): 905–916, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conn:1988:GCC

[CGT88] A. R. Conn, N. I. M.

Gould, and Ph. L. Toint. Global convergence of a class of trust region algorithms for optimization with simple bounds. SIAM Journal on Numerical Analysis, 25(2):433–460, April 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See correction [CGT89].

Conn:1989:CPG

[CGT89] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Correction to the paper on “Global Convergence of a Class of Trust Region Algorithms for Optimization with Simple Bounds” [SIAM J. Numer. Anal. 25 (1988), no. 2, 433– 460, MR 89h:90192]. SIAM Journal on Numerical Analysis, 26(3):764–767, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [CGT88].

Conn:1991:GCA

[CGT91] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. A globally convergent augmented Lagrangian algorithm for optimization with general constraints and simple bounds. SIAM Journal on Numerical Analysis, 28(2): 545–572, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cockburn:1998:PEE

[CGY98] Bernardo Cockburn, Pierre-Alain Gremaud, and Jimmy Xiangrong Yang. A Priori error estimates for numerical methods for scalar conservation laws Part III: Multidimensional flux-splitting monotone schemes on non-Cartesian grids. SIAM Journal on Numerical Analysis, 35 (5):1775–1803, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31616.

Cannon:1968:FDM

[CH68] J. R. Cannon and C. Denson Hill. A finite-difference method for degenerate ellipticparabolic equations. SIAM Journal on Numerical Analysis, 5(2):211–218, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cochran:1972:IEB

[CH72] James Alan Cochran and Erold W. Hinds. Improved error bounds for the eigenvalues of certain normal operators. SIAM Journal on Numerical Analysis, 9(3):446– 453, September 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coulaud:1994:NAF

[CH94] O. Coulaud and A. Henrot.

Numerical approximation of a free boundary problem arising in electromagnetic shaping. SIAM Journal on Numerical Analysis, 31(4):1109– 1127, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Charron:1995:CLI

[CH95] Richard J. Charron and Min Hu. A-contractivity of linearly implicit multistep methods. SIAM Journal on Numerical Analysis, 32(1):285– 295, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chui:1997:SEA

[CH97] Charles K. Chui and Dong Hong. Swapping edges of arbitrary triangulations to achieve the optimal order of approximation. SIAM Journal on Numerical Analysis, 34 (4):1472–1482, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27353.

Chawla:1968:EEG

[Cha68] M. M. Chawla. On the estimation of errors of Gaussian cubature formulas. SIAM Journal on Numerical Analysis, 5(1):172–181, March 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chawla:1969:DME

[Cha69] M. M. Chawla. On Davis’s method for the estimation of errors of Gauss–Chebyshev quadratures. SIAM Journal on Numerical Analysis, 6(1): 108–117, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chatelin:1973:CAM

[Cha73] Françoise Chatelin. Convergence of approximation methods to compute eigenelements of linear operations. SIAM Journal on Numerical Analysis, 10(5):939–948, October 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chalmers:1978:RCD

[Cha78a] Bruce A. Chalmers. On the rate of convergence of discretization in Chebyshev approximation. SIAM Journal on Numerical Analysis, 15(3): 612–617, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chatelin:1978:NCE

[Cha78b] Françoise E. Chatelin. Numerical computation of the eigenelements of linear integral operators by iterations. SIAM Journal on Numerical Analysis, 15(6):1112– 1124, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1984:DDS

[Cha84a] Tony F. Chan. Deflated decomposition of solutions of nearly singular systems. SIAM Journal on Numerical Analysis, 21(4):738– 754, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1984:SAF

[Cha84b] Tony F. Chan. Stability analysis of finite difference schemes for the advectiondiffusion equation. SIAM Journal on Numerical Analysis, 21(2):272–284, April 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1985:ANM

[Cha85a] Tony F. Chan. An approximate Newton method for coupled nonlinear systems. SIAM Journal on Numerical Analysis, 22(5):904–913, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chawla:1985:OAI

[Cha85b] M. M. Chawla. On the order and attainable intervals of periodicity of explicit Nyström methods for y = f(t, y). SIAM Journal on Numerical Analysis, 22(1):127– 131, February 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chandler:1986:SAS

[Cha86] G. A. Chandler. Superconvergent approximations to the solution of a boundary integral equation on polygonal domains. SIAM Journal on Numerical Analysis, 23(6): 1214–1229, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1987:APD

[Cha87] Tony F. Chan. Analysis of preconditioners for domain decomposition. SIAM Journal on Numerical Analysis, 24 (2):382–390, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1989:SFC

[Cha89] Raymond H. Chan. The spectrum of a family of circulant preconditioned Toeplitz systems. SIAM Journal on Numerical Analysis, 26(2): 503–506, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chang:1992:FEA

[Cha92] Ching Lung Chang. Finite element approximation for grad-div type systems in the plane. SIAM Journal on Numerical Analysis, 29(2): 452–461, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chartier:1994:SPO

[Cha94] P. Chartier. L-stable parallel one-block methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 31(2): 552–571, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cheung:1973:QPB

[Che73] To-Yat Cheung. Quasilinear parabolic boundary value problems. approximate solutions and error bounds by linear programming. SIAM Journal on Numerical Analysis, 10(6):1061–1079, December 1973. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Cheung:1975:ASE

[Che75] To-Yat Cheung. Approximate solutions and error bounds for quasi-linear hyperbolic initial boundary value problems. SIAM Journal on Numerical Analysis, 12(1):37–45, March 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cheung:1977:TNI

[Che77] To-Yat Cheung. Three nonlinear initial value problems of the hyperbolic type. SIAM Journal on Numerical Analysis, 14(3):484–491, June 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1987:SCL

[Che87] Ming Kui Chen. On the solution of circulant linear systems. SIAM Journal on Numerical Analysis, 24(3): 668–683, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1988:CS

[Che88] Pei Xian Chen. Convergence of SIP. SIAM Journal on Numerical Analysis, 25 (2):342–350, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1993:RMN

[Che93] Wei Dong Chen. A regularization method for the numerical inversion of the Laplace transform. SIAM Journal on Numerical Analysis, 30(3): 759–773, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1995:FEA

[Che95]  Zhangxin Chen.  Finite

element analysis of the one-dimensional full driftdiffusion semiconductor model. SIAM Journal on Numerical Analysis, 32(2):455–483, April 1995. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:1998:GSC

[Che98] Xiaojun Chen. Global and superlinear convergence of inexact Uzawa methods for saddle point problems with nondifferentiable mappings. SIAM Journal on Numerical Analysis, 35(3):1130– 1148, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29578.

Chan:1991:GRC

[CHL91] Tony F. Chan, Thomas Y.

Hou, and P. L. Lions. Geometry related convergence results for domain decomposition algorithms. SIAM Journal on Numerical Analysis, 28 (2):378–391, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conrad:1988:AOP

[CHM88] Francis Conrad, Raphaèle Herbin, and Hans D. Mittelmann. Approximation of obstacle problems by continuation methods. SIAM Journal on Numerical Analysis, 25(6): 1409–1431, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chong:1978:VMF

[Cho78] T. H. Chong. A variable mesh finite difference method for solving a class of parabolic differential equations in one space variable. SIAM Journal on Numerical Analysis, 15(4): 835–857, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chow:1992:FEE

[Cho92] S. Chow. Finite element error estimates for a blast furnace gas flow problem. SIAM Journal on Numerical Analysis, 29 (3):769–780, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chronopoulos:1991:SIM

[Chr91]  A. T. Chronopoulos.  s-

step iterative methods for (non)symmetric (in)definite linear systems. SIAM Journal on Numerical Analysis, 28(6): 1776–1789, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chui:1972:CGC

[Chu72] Charles K. Chui. Concerning Gaussian–Chebyshev quadrature errors. SIAM Journal on Numerical Analysis, 9(2): 237–240, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chu:1987:ETP

[Chu87] King-Wah Eric Chu. Exclusion theorems and the perturbation analysis of the generalized eigenvalue problem. SIAM Journal on Numerical Analysis, 24(5):1114– 1125, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chu:1990:MEM

[Chu90] King-Wah Eric Chu. On multiple eigenvalues of matrices depending on several parameters. SIAM Journal on Numerical Analysis, 27(5):1368– 1385, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chu:1992:NMI

[Chu92] Moody T. Chu. Numerical methods for inverse singular value problems. SIAM Journal on Numerical Analysis, 29 (3):885–903, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ciavaldini:1975:ANP

[Cia75] J. F. Ciavaldini. Analyse numérique d’un problème de Stefan à deux phases par une Mèthode d’éléments finis [French: Numerical analysis of a problem of Stefan in two phases by a finiteelement method]. SIAM Journal on Numerical Analysis, 12 (3):464–487, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ciment:1972:SMD

[Cim72]  Melvyn Ciment.  Sta-

ble matching of difference schemes. SIAM Journal on Numerical Analysis, 9 (4):695–701, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chawla:1975:BCI

[CJ75] M. M. Chawla and N. Jayarajan. Bivariate cardinal interpolation and cubature formulas. SIAM Journal on Numerical Analysis, 12(4):605– 616, September 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cabay:1976:PEM

[CJ76] S. Cabay and L. W. Jackson. A polynomial extrapolation method for finding limits and antilimits of vector sequences. SIAM Journal on Numerical Analysis, 13(5): 734–752, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cohen:1996:CAF

[CJ96] Gary Cohen and Patrick Joly.

Construction and analysis of fourth-order finite difference schemes for the acoustic wave equation in nonhomogeneous media. SIAM Journal on Numerical Analysis, 33 (4):1266–1302, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24644.

Chalmers:1974:RRH

[CJMT74] B. L. Chalmers, D. J. Johnson, F. T. Metcalf, and G. D. Taylor. Remarks on the rank of Hermite–Birkhoff interpolation. SIAM Journal on Numerical Analysis, 11(2): 254–259, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caflisch:1997:UAS

[CJR97] Russel E. Caflisch, Shi Jin, and Giovanni Russo. Uniformly accurate schemes for hyperbolic systems with relaxation. SIAM Journal on Numerical Analysis, 34 (1):246–281, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26809.

Chan:1991:SSO

[CJY91] Raymond H. Chan, Xiao-Qing Jin, and Man-Chung Yeung. The spectra of super-optimal circulant preconditioned Toeplitz systems. SIAM Journal on Numerical Analysis, 28(3):871–879, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1983:ADI

[CJZ83] Tony F. Chan, Kenneth R. Jackson, and Ben Ren Zhu. Alternating-direction incomplete factorizations. SIAM Journal on Numerical Analysis, 20(2):239–257, April 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1969:ASH

[CK69] J. R. Cannon and George H. Knightly. The approximation of the solution to the heat equation in a half-strip from data specified on the bounding characteristics. SIAM Journal on Numerical Analysis, 6(1):149–159, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1985:FME

[CK85a] Tony F. Chan and Tom Kerkhoven. Fourier methods with extended stability intervals for the Korteweg–de Vries equation. SIAM Journal on Numerical Analysis, 22 (3):441–454, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chawla:1985:UMF

[CK85b] M. M. Chawla and C. P.

Katti. A uniform mesh finite difference method for a class of singular two-point boundary value problems. SIAM Journal on Numerical Analysis, 22(3):561–565, June 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chiu:1992:ORK

[CK92] Chichia Chiu and David A.

Kopriva. An optimal Runge– Kutta method for steadystate solutions of hyperbolic systems. SIAM Journal on Numerical Analysis, 29(2): 425–438, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chou:1998:CMB

[CK98] S. H. Chou and D. Y. Kwak. A covolume method based on rotated bilinears for the generalized Stokes problem. SIAM Journal on Numerical Analysis, 35(2): 494–507, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29996.

Collins:1991:NAS

[CKL91] Charles Collins, David Kinderlehrer, and Mitchell Luskin. Numerical approximation of the solution of a variational problem with a double well potential. SIAM Journal on Numerical Analysis, 28(2): 321–332, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cahlon:1995:SSH

[CKS95] Baruch Cahlon, Devadatta M. Kulkarni, and Peter Shi. Stepwise stability for the heat equation with a nonlocal constraint. SIAM Journal on Numerical Analysis, 32(2): 571–593, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chou:1998:MCM

[CKV98] So-Hsiang Chou, Do Y. Kwak, and Panayot S. Vassilevski. Mixed covolume methods for elliptic problems on triangular grids. SIAM Journal on Numerical Analysis, 35(5):1850– 1861, October 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32128.

Cheney:1964:GRA

[CL64] E. W. Cheney and H. L. Loeb.

Generalized rational approximation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):11– 25, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Crouzeix:1984:CVS

[CL84] M. Crouzeix and F. J. Lisbona. The convergence of variable-stepsize, variableformula, multistep methods. SIAM Journal on Numerical Analysis, 21(3):512–534, June 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caflisch:1989:CVM

[CL89] Russel E. Caflisch and John S.

Lowengrub. Convergence of the vortex method for vortex sheets. SIAM Journal on Numerical Analysis, 26(5):1060– 1080, October 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1990:PEE

[CL90] J. R. Cannon and Yanping Lin. A priori L^{2} error estimates for finite-element methods for nonlinear diffusion equations with memory. SIAM Journal on Numerical Analysis, 27(3):595–607, June 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coleman:1992:GQC

[CL92] Thomas F. Coleman and Yuying Li. A global and quadratically convergent method for linear L_{∞} problems. SIAM Journal on Numerical Analysis, 29(4):1166–1186, August 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clenshaw:1964:CBP

[Cle64] C. W. Clenshaw. A comparison of “best” polynomial approximations with truncated Chebyshev series expansions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):26– 37, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Clements:1981:BIP

[Cle81] D. L. Clements. A boundary integral procedure for the numerical solution of a class of mixed boundary value problems. SIAM Journal on Numerical Analysis, 18(4):664– 680, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clements:1990:CPP

[Cle90] John C. Clements. Convexitypreserving piecewise rational cubic interpolation. SIAM Journal on Numerical Analysis, 27(4):1016–1023, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cline:1965:RGI

[Cli65] Randall E. Cline. Representations for the generalized inverse of sums of matrices. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):99– 114, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Cline:1968:IRI

[Cli68] Randall E. Cline. Inverses of rank invariant powers of a matrix. SIAM Journal on Numerical Analysis, 5(1): 182–197, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cline:1973:EMS

[Cli73] A. K. Cline. An elimination method for the solution of linear least squares problems. SIAM Journal on Numerical Analysis, 10(2): 283–289, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cline:1976:DMU

[Cli76] A. K. Cline. A descent method for the uniform solution to over-determined systems of linear equations. SIAM Journal on Numerical Analysis, 13(3):293–309, June 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Calvo:1987:SVS

[CLM87] M. Calvo, F. Lisbona, and J. Montijano. On the stability of variable-stepsize Nordsieck BDF methods. SIAM Journal on Numerical Analysis, 24(4): 844–854, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Calvo:1989:SIV

[CLM89] M. Calvo, F. Lisbona, and J. Montijano. On the stability of interpolatory variablestepsize Adams methods in Nordsieck form. SIAM Journal on Numerical Analysis, 26(4):946–962, August 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Catte:1992:ISS

[CLMC92] Francine Catté, Pierre-Louis Lions, Jean-Michel Morel, and Tomeu Coll. Image selective smoothing and edge detection by nonlinear diffusion. SIAM Journal on Numerical Analysis, 29(1):182– 193, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:1994:FOS

[CLMM94] Z. Cai, R. Lazarov, T. A.

Manteuffel, and S. F. Mc-Cormick. First-order system least squares for second-order partial differential equations: Part I. SIAM Journal on Numerical Analysis, 31(6):1785– 1799, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Colombeau:1989:MPS

[CLNP89] J. F. Colombeau, A. Y. LeRoux, A. Noussair, and B. Perrot. Microscopic profiles of shock waves and ambiguities in multiplications of distributions. SIAM Journal on Numerical Analysis, 26(4):871– 883, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1986:SES

[CLS86] Tony F. Chan, Ding Lee, and Long Jun Shen. Stable explicit schemes for equations of the Schrödinger type. SIAM Journal on Numerical Analysis, 23(2):274–281, April 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cannon:1965:SPN

[CM65] J. R. Cannon and Keith Miller. Some problems in numerical analytic continuation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):87– 98, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Chalmers:1974:TLR

[CM74] Bruce L. Chalmers and Frederic T. Metcalf. Taylor-like remainder formulas for interpolation by arbitrary linear functionals. SIAM Journal on Numerical Analysis, 11(5): 950–964, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coleman:1983:ESJ

[CM83] Thomas F. Coleman and Jorge J. Moré. Estimation of sparse Jacobian matrices and graph coloring problems. SIAM Journal on Numerical Analysis, 20(1):187– 209, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chatelin:1984:ADE

[CM84] Françoise Chatelin and Willard L. Miranker. Aggregation/ disaggregation for eigenvalue problems. SIAM Journal on Numerical Analysis, 21(3): 567–582, June 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Calamai:1987:QNU

[CM87] Paul H. Calamai and Jorge J.

Moré. Quasi-Newton updates with bounds. SIAM Journal on Numerical Analysis, 24(6): 1434–1441, December 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chin:1988:ABS

[CM88a] R. C. Y. Chin and Thomas A.

Manteuffel. An analysis of block successive overrelaxation for a class of matrices with complex spectra. SIAM Journal on Numerical Analysis, 25(3):564–585, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Criscuolo:1988:CPF

[CM88b] Giuliana Criscuolo and Giuseppe Mastroianni. On the convergence of product formulas for the numerical evaluation of derivatives of Cauchy principal value integrals. SIAM Journal on Numerical Analysis, 25(3):713–727, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:1990:AFV

[CM90a] Zhi Qiang Cai and Steve McCormick. On the accuracy of the finite volume element method for diffusion equations on composite grids. SIAM Journal on Numerical Analysis, 27(3):636–655, June 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Childs:1990:CGM

[CM90b] P. N. Childs and K. W. Morton. Characteristic Galerkin methods for scalar conservation laws in one dimension. SIAM Journal on Numerical Analysis, 27(3):553–594, June 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cambanis:1992:TSM

[CM92] Stamatis Cambanis and Elias Masry. Trapezoidal stratified Monte Carlo integration. SIAM Journal on Numerical Analysis, 29(1):284– 301, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1999:CLD

[CM99a] Tony F. Chan and Pep Mulet. On the convergence of the lagged diffusivity fixed point method in total variation image restoration. SIAM Journal on Numerical Analysis, 36(2):354–367, April 1999.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32707.

Costantini:1999:PCE

[CM99b] Paolo Costantini and Carla Manni. A parametric cubic element with tension properties. SIAM Journal on Numerical Analysis, 36(2):607–628, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32072.

Cai:1991:FVE

[CMM91] Zhiqiang Cai, Jan Mandel, and Steve McCormick. The finite volume element method for diffusion equations on general triangulations. SIAM Journal on Numerical Analysis, 28(2):392–402, April 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:1997:FOSb

[CMM97a] Z. Cai, T. A. Manteuffel, and S. F. McCormick. Firstorder system least squares for the Stokes equations, with application to linear elasticity. SIAM Journal on Numerical Analysis, 34(5): 1727–1741, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27299.

Cai:1997:MMN

[CMM97b] Zhiqiang Cai, Jan Mandel, and Steve McCormick. Multigrid methods for nearly singular linear equations and eigenvalue problems. SIAM Journal on Numerical Analysis, 34 (1):178–200, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26113.

Cai:1997:FOSa

[CMM97c] Zhiqiang Cai, Thomas A. Manteuffel, and Stephen F. McCormick. First-order system least squares for secondorder partial differential equations: Part II. SIAM Journal on Numerical Analysis, 34(2):425–454, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26606.

Cai:1998:FOS

[CMMP98] Zhiqiang Cai, Thomas A. Manteuffel, Stephen F. Mc-Cormick, and Seymour V. Parter. First-order system least squares (FOSLS) for planar linear elasticity: Pure traction problem. SIAM Journal on Numerical Analysis, 35(1):320– 335, February 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29493.

Cline:1979:ECN

[CMSW79] A. K. Cline, C. B. Moler, G. W. Stewart, and J. H. Wilkinson. An estimate for the condition number of a matrix. SIAM Journal on Numerical Analysis, 16(2): 368–375, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chu:1988:IFA

[CN88] Moody T. Chu and Larry K.

Norris. Isospectral flows and abstract matrix factorizations. SIAM Journal on Numerical Analysis, 25(6):1383– 1391, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chang:1997:LSF

[CN97] C. L. Chang and John J.

Nelson. Least-squares finite element method for the Stokes problem with zero residual of mass conservation. SIAM Journal on Numerical Analysis, 34(2): 480–489, April 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27336.

Chan:1993:FBP

[CNP93] Raymond H. Chan, James G. Nagy, and Robert J. Plemmons. FFT-based preconditioners for Toeplitz-block least squares problems. SIAM Journal on Numerical Analysis, 30(6):1740–1768, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clenshaw:1980:UAE

[CO80] C. W. Clenshaw and F. W. J. Olver. An unrestricted algorithm for the exponential function. SIAM Journal on Numerical Analysis, 17(2): 310–331, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clenshaw:1987:LIA

[CO87] C. W. Clenshaw and F. W. J. Olver. Level-index arithmetic operations. SIAM Journal on Numerical Analysis, 24(2): 470–485, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cockburn:1989:QSS

[Coc89] Bernardo Cockburn. Quasimonotone schemes for scalar conservation laws. Part I. SIAM Journal on Numerical Analysis, 26(6):1325– 1341, December 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cockburn:1990:QSSa

[Coc90a] Bernardo Cockburn. Quasimonotone schemes for scalar conservation laws. Part II. SIAM Journal on Numerical Analysis, 27(1):247–258, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cockburn:1990:QSSb

[Coc90b] Bernardo Cockburn. Quasimonotone schemes for scalar conservation laws. Part III. SIAM Journal on Numerical Analysis, 27(1):259–276, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cohen:1971:NTC

[Coh71] Edgar A. Cohen, Jr. Note on a truncated Chebyshev series modified to match function values at interval endpoints. SIAM Journal on Numerical Analysis, 8(4):754– 756, December 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cohen:1972:RCS

[Coh72] Arthur I. Cohen. Rate of convergence of several conjugate gradient algorithms. SIAM Journal on Numerical Analysis, 9(2):248–259, June 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Collins:1998:CRI

[Col98] Charles R. Collins. Convergence of a reduced integration method for computing microstructures. SIAM Journal on Numerical Analysis, 35(3):1271–1298, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30554.

Conn:1973:COU

[Con73]  Andrew R. Conn.  Con-

strained optimization using a nondifferentiale penalty function. SIAM Journal on Numerical Analysis, 10(4):760– 784, September 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conn:1976:LPN

[Con76] A. R. Conn. Linear programming via a nondifferentiable penalty function. SIAM Journal on Numerical Analysis, 13(1):145–154, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cooke:1972:SSI

[Coo72] Charlie H. Cooke. On stiffly stable implicit linear multistep methods. SIAM Journal on Numerical Analysis, 9 (1):29–34, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cooper:1978:OCG

[Coo78] G. J. Cooper. The order of convergence of general linear methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 15(4): 643–661, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cooper:1981:EEG

[Coo81] G. J. Cooper. Error estimates for general linear methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 18(1): 65–82, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

Cope:1990:CPM

[Cop90] Davis K. Cope. Convergence of piessens’ method for numerical inversion of the Laplace transform on the real line. SIAM Journal on Numerical Analysis, 27(5):1345– 1354, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conca:1998:AAR

[COP98] Carlos Conca, Axel Osses, and Jacques Planchard. Asymptotic analysis relating spectral models in fluid-solid vibrations. SIAM Journal on Numerical Analysis, 35 (3):1020–1048, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30480.

Corradi:1981:NSS

[Cor81] Corrado Corradi. A note on the solution of separable nonlinear least-squares problems with separable nonlinear equality constraints. SIAM Journal on Numerical Analysis, 18(6):1134–1138, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cornelius:1983:AIA

[Cor83] Herbert Cornelius. On the acceleration of an intervalarithmetic iteration method. SIAM Journal on Numerical Analysis, 20(5):1010– 1022, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Corrias:1996:FLF

[Cor96] Lucilla Corrias. Fast Legendre– Fenchel transform and applications to Hamilton–Jacobi equations and conservation laws. SIAM Journal on Numerical Analysis, 33(4): 1534–1558, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26020.

Costello:1971:ASD

[Cos71] George A. Costello. The approximate solution of the Dirichlet problem by means of integral equations. SIAM Journal on Numerical Analysis, 8(1):135–136, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cosnard:1979:BSA

[Cos79] M. Y. Cosnard. On the behavior of successive approximations. SIAM Journal on Numerical Analysis, 16(2): 300–310, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cox:1984:MSS

[Cox84] Dennis D. Cox. Multivariate smoothing spline functions. SIAM Journal on Numerical Analysis, 21(4):789– 813, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cranley:1976:RNT

[CP76] R. Cranley and T. N. L. Patterson. Randomization of number theoretic methods for multiple integration. SIAM Journal on Numerical Analysis, 13(6):904–914, December 1976. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Conn:1977:PFM

[CP77] Andrew R. Conn and Tomasz Pietrzykowski. A penalty function method converging directly to a constrained optimum. SIAM Journal on Numerical Analysis, 14(2): 348–375, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cooper:1991:ADC

[CP91a] K. D. Cooper and P. M. Prenter. Alternating direction collocation for separable elliptic partial differential equations. SIAM Journal on Numerical Analysis, 28 (3):711–727, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coron:1991:NPK

[CP91b] F. Coron and B. Perthame. Numerical passage from kinetic to fluid equations. SIAM Journal on Numerical Analysis, 28(1):26–42, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coquel:1998:REA

[CP98] Frédéric Coquel and Benoît Perthame. Relaxation of energy and approximate Riemann solvers for general pressure laws in fluid dynamics. SIAM Journal on Numerical Analysis, 35(6): 2223–2249, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31852.

Canuto:1982:EES

[CQ82] C. Canuto and A. Quarteroni. Error estimates for spectral and pseudospectral approximations of hyperbolic equations. SIAM Journal on Numerical Analysis, 19(3): 629–642, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chawla:1976:MVA

[CR76] M. M. Chawla and T. R.

Ramakrishnan. Minimum variance approximate formulas. SIAM Journal on Numerical Analysis, 13(1):113– 128, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cope:1979:BSL

[CR79] J. E. Cope and B. W. Rust.

Bounds on solutions of linear systems with inaccurate data. SIAM Journal on Numerical Analysis, 16(6):950– 963, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cottet:1984:PMO

[CR84] G.-H. Cottet and P.-A. Raviart. Particle methods for the one-dimensional Vlasov– Poisson equations. SIAM Journal on Numerical Analysis, 21(1):52–76, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1986:GDB

[CR86] Tony F. Chan and Diana C. Resasco. Generalized deflated block-elimination. SIAM Journal on Numerical Analysis, 23(5):913–924, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cline:1990:CTD

[CR90] A. K. Cline and R. J. Renka. A constrained twodimensional triangulation and the solution of closest node problems in the presence of barriers. SIAM Journal on Numerical Analysis, 27(5): 1305–1321, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Csendes:1997:SDS

[CR97] T. Csendes and D. Ratz. Subdivision direction selection in interval methods for global optimization. SIAM Journal on Numerical Analysis, 34(3):922–938, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28152.

Crawford:1976:SGE

[Cra76] C. R. Crawford. A stable generalized eigenvalue problem. SIAM Journal on Numerical Analysis, 13(6):854– 860, December 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See errata [Cra78].

Crawford:1978:ESG

[Cra78] C. R. Crawford. Errata:

“A Stable Generalized Eigenvalue Problem” [SIAM J. Numer. Anal. 13 (1976), no. 6, 854–860]. SIAM Journal on Numerical Analysis, 15(5): 1070, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Cra76].

Cryer:1967:DAG

[Cry67] C. W. Cryer. The difference analogue of Gauss’ theorem. SIAM Journal on Numerical Analysis, 4(2):155–162, June 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cryer:1974:HSM

[Cry74] Colin W. Cryer. Highly stable multistep methods for retarded differential equations. SIAM Journal on Numerical Analysis, 11(4):788– 797, September 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chartres:1972:GTC

[CS72] Bruce Chartres and Robert Stepleman. A general theory of convergence for numerical methods. SIAM Journal on Numerical Analysis, 9 (3):476–492, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chartres:1974:AOC

[CS74a] Bruce A. Chartres and Robert S. Stepleman. Actual order of convergence of Runge–Kutta methods on differential equations with discontinuities. SIAM Journal on Numerical Analysis, 11(6): 1193–1206, December 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chui:1974:S

[CS74b] Charles K. Chui and Philip W. Smith. On H^{m,∞}-splines. SIAM Journal on Numerical Analysis, 11(3):554–558, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chartres:1975:OCL

[CS75] Bruce A. Chartres and Robert S. Stepleman. Order of convergence of linear multistep methods for functional differential equations. SIAM Journal on Numerical Analysis, 12(6):876–886, December 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1987:SAD

[CS87] Tony F. Chan and Long Jun Shen. Stability analysis of difference schemes for variable coefficient Schrödinger type equations. SIAM Journal on Numerical Analysis, 24(2): 336–349, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Costabel:1990:CFB

[CS90] Martin Costabel and Ernst P.

Stephan. Coupling of finite and boundary element methods for an elastoplastic interface problem. SIAM Journal on Numerical Analysis, 27 (5):1212–1226, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cockburn:1994:NSC

[CS94] Bernardo Cockburn and Chi-Wang Shu. Nonlinearly stable compact schemes for shock calculations. SIAM Journal on Numerical Analysis, 31(3): 607–627, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:1996:ABE

[CS96] Carsten Carstensen and Ernst P. Stephan. Adaptive boundary element methods for some first kind integral equations. SIAM Journal on Numerical Analysis, 33(6): 2166–2183, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25350.

Cockburn:1998:LDG

[CS98] Bernardo Cockburn and Chi-Wang Shu. The local discontinuous Galerkin method for time-dependent convectiondiffusion systems. SIAM Journal on Numerical Analysis, 35(6):2440–2463, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31671.

Chapelle:1999:SFE

[CS99] Dominique Chapelle and Rolf Stenberg. Stabilized finite element formulations for shells in a bending dominated state. SIAM Journal on Numerical Analysis, 36(1):32–73, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30291.

Carasso:1978:DRR

[CSH78] Alfred S. Carasso, James G.

Sanderson, and James M. Hyman. Digital removal of random media image degradations by solving the diffusion equation backwards in time. SIAM Journal on Numerical Analysis, 15(2): 344–367, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cano:1997:EGN

[CSS97] B. Cano and J. M. Sanz-Serna. Error growth in the numerical integration of periodic orbits, with application to Hamiltonian and reversible systems. SIAM Journal on Numerical Analysis, 34 (4):1391–1417, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28115.

Cryer:1972:NSV

[CT72] Colin W. Cryer and Lucio Tavernini. The numerical solution of Volterra functional differential equations by Euler’s method. SIAM Journal on Numerical Analysis, 9(1): 105–129, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carlson:1983:DBE

[CT83] B. C. Carlson and John Todd. The degenerating behavior of elliptic functions. SIAM Journal on Numerical Analysis, 20(6):1120–1129, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Charrier:1986:FTM

[CT86] P. Charrier and B. Tessieras. On front-tracking methods applied to hyperbolic systems of nonlinear conservation laws. SIAM Journal on Numerical Analysis, 23(3): 461–472, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cullum:1971:NDR

[Cul71]  Jane Cullum.  Numerical

differentiation and regularization. SIAM Journal on Numerical Analysis, 8(2): 254–265, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Curtiss:1966:SDP

[Cur66] J. H. Curtiss. Solutions of the Dirichlet problem in the plane by approximation with Faber polynomials. SIAM Journal on Numerical Analysis, 3(2): 204–228, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cuyt:1987:RCS

[Cuy87] A. Cuyt. A recursive computation scheme for multivariate rational interpolants. SIAM Journal on Numerical Analysis, 24(1):228–239, February 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cooper:1972:SER

[CV72] G. J. Cooper and J. H.

Verner. Some explicit Runge– Kutta methods of high order. SIAM Journal on Numerical Analysis, 9(3):389– 405, September 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:1999:ERD

[CV99] Carsten Carstensen and Rüdiger Verfürth. Edge residuals dominate A posteriori error estimates for low order finite element methods. SIAM Journal on Numerical Analysis, 36(5): 1571–1587, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32334.

Carroll:1976:AAI

[CW76] M. P. Carroll and W. R.

Woodward. An algorithm for approximate integration using near best approximates. SIAM Journal on Numerical Analysis, 13(1): 104–108, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cruickshank:1978:CEB

[CW78] D. M. Cruickshank and K. Wright. Computable error bounds for polynomial collocation methods. SIAM Journal on Numerical Analysis, 15 (1):134–151, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:1993:MSA

[CW93] Xiao-Chuan Cai and Olof B. Widlund. Multiplicative Schwarz algorithms for some nonsymmetric and indefinite problems. SIAM Journal on Numerical Analysis, 30 (4):936–952, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:1996:AMC

[CW96] Wei Cai and Jianzhong Wang. Adaptive multiresolution collocation methods for initial boundary value problems of nonlinear PDEs. SIAM Journal on Numerical Analysis, 33 (3):937–970, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chiu:1997:AMH

[CW97] Chichia Chiu and Noel Walkington. An ADI method for hysteretic reaction-diffusion systems. SIAM Journal on Numerical Analysis, 34 (3):1185–1206, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27018.

Chen:1998:PGI

[CX98] Zhongying Chen and Yuesheng Xu. The Petrov–Galerkin and iterated Petrov–Galerkin methods for second-kind integral equations. SIAM Journal on Numerical Analysis, 35 (1):406–434, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29721.

Chung:1977:LAU

[CY77] K. C. Chung and T. H.

Yao. On lattices admitting unique Lagrange interpolations. SIAM Journal on Numerical Analysis, 14(4):735– 743, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1992:CPC

[CY92] Raymond H. Chan and Man-Chung Yeung. Circulant preconditioners constructed from kernels. SIAM Journal on Numerical Analysis, 29(4): 1093–1103, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chan:1993:CPC

[CY93] Raymond H. Chan and Man-Chung Yeung. Circulant preconditioners for complex Toeplitz matrices. SIAM Journal on Numerical Analysis, 30(4):1193–1207, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Drakopoulos:1999:GCS

[DAB99] V. Drakopoulos, N. Argyropoulos, and A. Böhm. Generalized computation of Schröder iteration functions to motivate families of Julia and Mandelbrot-like sets. SIAM Journal on Numerical Analysis, 36(2):417–435, April 1999. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31736.

Dagnino:1986:ESR

[Dag86] Catterina Dagnino. Extensions of some results for interpolatory product integration rules to rules not necessarily of interpolatory type. SIAM Journal on Numerical Analysis, 23(6):1284–1289, December 1986. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahmen:1980:MBS

[Dah80] Wolfgang Dahmen. On multivariate B-splines. SIAM Journal on Numerical Analysis, 17 (2):179–191, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahlquist:1983:OLM

[Dah83] Germund Dahlquist. On oneleg multistep methods. SIAM Journal on Numerical Analysis, 20(6):1130–1138, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dai:1992:UST

[Dai92] Wei Zhong Dai. An unconditionally stable three-level explicit difference scheme for the Schrödinger equation with a variable coefficient. SIAM Journal on Numerical Analysis, 29(1):174–181, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Daniel:1967:CGM

[Dan67] James W. Daniel. The conjugate gradient method for linear and nonlinear operator equations. SIAM Journal on Numerical Analysis, 4 (1):10–26, March 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Daniel:1970:CCC

[Dan70] James W. Daniel. A correction concerning the convergence rate for the conjugate gradient method. SIAM Journal on Numerical Analysis, 7 (2):277–280, June 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Daniel:1973:PSL

[Dan73] James W. Daniel. On perturbations in systems of linear inequalities. SIAM Journal on Numerical Analysis, 10 (2):299–307, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Daniel:1975:RPL

[Dan75] James W. Daniel. Remarks on perturbations in linear inequalities. SIAM Journal on Numerical Analysis, 12(5): 770–772, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Daripa:1991:NTO

[Dar91] Prabir Daripa. A new theory for one-dimensional adaptive grid generation and its applications. SIAM Journal on Numerical Analysis, 28(6): 1635–1660, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davidon:1980:CAC

[Dav80] William C. Davidon. Conic approximations and collinear scalings for optimizers. SIAM Journal on Numerical Analysis, 17(2):268–281, April 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davies:1994:NSC

[Dav94] Penny J. Davies. Numerical stability and convergence of approximations of retarded potential integral equations. SIAM Journal on Numerical Analysis, 31(3):856–875, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davidson:1997:LSC

[Dav97] Bryan D. Davidson. Largescale continuation and numerical bifurcation for partial differential equations. SIAM Journal on Numerical Analysis, 34(5):2008– 2027, October 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27328.

Dawson:1991:GMM

[Daw91] Clint N. Dawson. Godunovmixed methods for advective flow problems in one space dimension. SIAM Journal on Numerical Analysis, 28 (5):1282–1309, October 1991.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dawson:1993:GMM

[Daw93] Clint Dawson. Godunovmixed methods for advectiondiffusion equations in multidimensions. SIAM Journal on Numerical Analysis, 30 (5):1315–1332, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dawson:1998:AUM

[Daw98] Clint Dawson. Analysis of an upwind-mixed finite element method for nonlinear contaminant transport equations. SIAM Journal on Numerical Analysis, 35 (5):1709–1724, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25942.

Dax:1985:SRL

[Dax85] Achiya Dax. Successive refinement of large multicell models. SIAM Journal on Numerical Analysis, 22(5): 865–887, October 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1980:MAS

[DC80] Charles B. Dunham and Charles R. Crawford. Minimax approximation by a semicircle. SIAM Journal on Numerical Analysis, 17(1):63– 65, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deFigueiredo:1990:SDP

[dC90] Rui J. P. de Figueiredo and Guanrong Chen. PDL_{g} splines defined by partial differential operators with initial and boundary value conditions. SIAM Journal on Numerical Analysis, 27(2): 519–528, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1970:GMP

[DD70] Jim Douglas, Jr. and Todd Dupont. Galerkin methods for parabolic equations. SIAM Journal on Numerical Analysis, 7(4):575–626, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1986:SFD

[dd86] C. de Boor and F. de Hoog.

Stability of finite difference schemes for two-point boundary value problems. SIAM Journal on Numerical Analysis, 23(5):925–935, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1993:UCT

[DD93] Craig C. Douglas and Jim Douglas, Jr. A unified convergence theory for abstract multigrid or multilevel algorithms, serial and parallel. SIAM Journal on Numerical Analysis, 30(1):136– 158, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dawson:1994:EIC

[DD94] Clint N. Dawson and Todd F. Dupont. Explicit/implicit, conservative domain decomposition procedures for parabolic problems based on block-centered finite differences. SIAM Journal on Numerical Analysis, 31(4):1045– 1061, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dubuc:1996:EBE

[DD96] B. Dubuc and S. Dubuc. Error bounds on the estimation of fractal dimension. SIAM Journal on Numerical Analysis, 33(2): 602–626, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1983:SOS

[ddd83] C. de Boor, F. de Hoog, and H. B. de Keller. The stability of one-step schemes for first-order two-point boundary value problems. SIAM Journal on Numerical Analysis, 20(6):1139–1146, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1979:IIT

[DDE79] Jim Douglas, Jr., Todd Dupont, and Richard E. Ewing. Incomplete iteration for time-stepping a Galerkin method for a quasilinear parabolic problem. SIAM Journal on Numerical Analysis, 16(3):503–522, June 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deift:1991:BSV

[DDLT91] Percy Deift, James Demmel, Luen Chau Li, and Carlos Tomei. The bidiagonal singular value decomposition and Hamiltonian mechanics. SIAM Journal on Numerical Analysis, 28(5):1463– 1516, October 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahmen:1980:MDS

[DDS80] W. Dahmen, R. De Vore, and K. Scherer. Multidimensional spline approximation. SIAM Journal on Numerical Analysis, 17(3): 380–402, June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

delaValleePoussin:1968:ARA

[de 68] F. de la Vallée Poussin. An accelerated relaxation algorithm for iterative solution of elliptic equations. SIAM Journal on Numerical Analysis, 5 (2):340–351, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

DePillis:1973:GSC

[De 73] John De Pillis. Gauss–Seidel convergence for operators on Hilbert space. SIAM Journal on Numerical Analysis, 10(1): 112–122, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1977:PCB

[de 77]  Carl de Boor.  Package

for calculating with B-splines. SIAM Journal on Numerical Analysis, 14(3):441–472, June 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1978:CNC

[de 78a] Carl de Boor. A comment on “Numerical Comparisons of Algorithms for Polynomial and Rational Multivariate Approximations” [SIAM J. Numer. Anal. 15 (1978), no. 6, 1197–1207, by J. N. Henry, M. S. Henry and D. Schmidt]. SIAM Journal on Numerical Analysis, 15(6): 1208–1211, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [HHS78].

dePillis:1978:FCI

[de 78b] John de Pillis. Faster convergence for iterative solutions to systems via threepart splittings. SIAM Journal on Numerical Analysis, 15(5): 888–911, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1980:DBM

[de 80] Carl de Boor. Dichotomies for band matrices. SIAM Journal on Numerical Analysis, 17(6):894–907, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deGee:1986:AEM

[de 86] Maarten de Gee. Asymptotic expansions for the midpoint rule applied to delay differential equations. SIAM Journal on Numerical Analysis, 23(6):1254–1272, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1988:CBS

[de 88] Carl de Boor. The condition of the B-spline basis for polynomials. SIAM Journal on Numerical Analysis, 25(1):

148–152, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sturler:1999:TSO

[de 99] Eric de Sturler. Truncation strategies for optimal Krylov subspace methods. SIAM Journal on Numerical Analysis, 36(3):864–889, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31595.

Donaldson:1972:UAQ

[DE72] J. D. Donaldson and David Elliott. A unified approach to quadrature rules with asymptotic estimates of their remainders. SIAM Journal on Numerical Analysis, 9 (4):573–602, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dow:1979:NSS

[DE79] M. L. Dow and David Elliott. The numerical solution of singular integral equations over (−1, 1). SIAM Journal on Numerical Analysis, 16(1): 115–134, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorobantu:1998:WBN

[DE98] Mihai Dorobantu and Björn Engquist. Wavelet-based numerical homogenization. SIAM Journal on Numerical Analysis, 35(2):540–559, April 1998. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29888.

Dejon:1967:NSD

[Dej67] B. Dejon. Numerical stability of difference equations with matrix coefficients. SIAM Journal on Numerical Analysis, 4(1):119–128, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dellwo:1988:ARA

[Del88] David R. Dellwo. Accelerated refinement with applications to integral equations. SIAM Journal on Numerical Analysis, 25(6):1327–1339, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dellwo:1989:ASR

[Del89] David R. Dellwo. Accelerated spectral refinement with applications to integral operators. SIAM Journal on Numerical Analysis, 26(5):1184– 1193, October 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Delbourgo:1993:ARI

[Del93]  Roger Delbourgo.  Accu-

rate C^{2} rational interpolants in tension. SIAM Journal on Numerical Analysis, 30(2): 595–607, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

DeLillo:1994:ANC

[DeL94] Thomas K. DeLillo. The accuracy of numerical conformal mapping methods: A survey of examples and results. SIAM Journal on Numerical Analysis, 31(3):788–812, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Demko:1976:LPS

[Dem76] Stephen Demko. Lacunary polynomial spline interpolation. SIAM Journal on Numerical Analysis, 13(3): 369–381, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Demko:1977:IBM

[Dem77] Stephen Demko. Inverses of band matrices and local convergence of spline projections. SIAM Journal on Numerical Analysis, 14(4):616– 619, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Demmel:1983:CNE

[Dem83] James Demmel. The condition number of equivalence transformations that block diagonalize matrix pencils. SIAM Journal on Numerical Analysis, 20(3):599–610, June 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Demmel:1987:SPS

[Dem87] James Weldon Demmel. The smallest perturbation of a submatrix which lowers the rank and constrained total least squares problems. SIAM Journal on Numerical Analysis, 24(1):199–206, February 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1967:NMN

[Den67a]  J. E. Dennis, Jr.  On

Newton’s method and nonlinear simultaneous displacements. SIAM Journal on Numerical Analysis, 4(1):103– 108, March 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1967:SNM

[Den67b] J. E. Dennis, Jr. A stationary Newton method for nonlinear functional equations. SIAM Journal on Numerical Analysis, 4(2):222–232, June 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1969:KHN

[Den69] J. E. Dennis, Jr. On the Kantorovich hypothesis for Newton’s method. SIAM Journal on Numerical Analysis, 6 (3):493–507, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1974:TMG

[Den74a] J. E. Dendy. Two methods of Galerkin type achieving optimum L^{2} rates of convergence for first order hyperbolics. SIAM Journal on Numerical Analysis, 11(3):637–653, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1974:PGM

[Den74b] J. E. Dendy, Jr. Penalty Galerkin methods for partial differential equations. SIAM Journal on Numerical Analysis, 11(3):604–636, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1975:ASG

[Den75] J. E. Dendy, Jr. An analysis of some Galerkin schemes for the solution of nonlinear timedependent problems. SIAM Journal on Numerical Analysis, 12(4):541–565, September 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1977:ADMb

[Den77a] J. E. Dendy, Jr. An alternating direction method for Schrödinger’s equation. SIAM Journal on Numerical Analysis, 14(6):1028–1032, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1977:ADMa

[Den77b] J. E. Dendy, Jr. Alternating direction methods for nonlinear time-dependent problems. SIAM Journal on Numerical Analysis, 14(2): 313–326, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1977:GMS

[Den77c] J. E. Dendy, Jr. Galerkin’s method for some highly nonlinear problems. SIAM Journal on Numerical Analysis, 14 (2):327–347, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dershem:1971:ABE

[Der71] Herbert L. Dershem. Approximation of the Bessel eigenvalue problem by finite differences. SIAM Journal on Numerical Analysis, 8(4):706– 716, December 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duff:1976:GND

[DER76] Iain S. Duff, A. M. Erisman, and J. K. Reid. On George’s nested dissection method. SIAM Journal on Numerical Analysis, 13(5): 686–695, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Descloux:1972:FEM

[Des72] Jean Descloux. On finite element matrices. SIAM Journal on Numerical Analysis, 9 (2):260–265, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Descloux:1981:ENR

[Des81] Jean Descloux. Essential numerical range of an operator with respect to a coercive form and the approximation of its spectrum by the Galerkin method. SIAM Journal on Numerical Analysis, 18(6):1128–1133, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dembo:1982:INM

[DES82] Ron S. Dembo, Stanley C. Eisenstat, and Trond Steihaug. Inexact Newton methods. SIAM Journal on Numerical Analysis, 19(2): 400–408, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dendy:1975:ADG

[DF75] J. E. Dendy, Jr. and G. Fairweather. Alternating-direction Galerkin methods for parabolic and hyperbolic problems on rectangular polygons. SIAM Journal on Numerical Analysis, 12(2):144–163, April 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See errata [DF76a].

Dendy:1976:EAD

[DF76a] J. E. Dendy, Jr. and G. Fairweather. Errata: “Alternating-Direction Galerkin Methods for Parabolic and Hyperbolic Problems on Rectangular Polygons” [SIAM J. Numer. Anal. 12 (1975), no. 2, 144–163]. SIAM Journal on Numerical Analysis, 13(4): 644, September 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [DF75].

Diamond:1976:CRM

[DF76b] Martin A. Diamond and Dulcinea L. V. Ferreira. On a cyclic reduction method for the solution of Poisson’s equations. SIAM Journal on Numerical Analysis, 13(1): 54–70, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dellwo:1984:ASA

[DF84] David R. Dellwo and Morton B. Friedman. Accelerated spectral analysis of compact operators. SIAM Journal on Numerical Analysis, 21(6): 1115–1131, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dellwo:1991:API

[DF91] David R. Dellwo and Morton B. Friedman. Accelerated projection and iterated projection methods with applications to nonlinear integral equations. SIAM Journal on Numerical Analysis, 28(1): 236–250, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dupont:1974:TLG

[DFJ74] Todd Dupont, Graeme Fairweather, and J. Peter Johnson. Three-level Galerkin methods for parabolic equations. SIAM Journal on Numerical Analysis, 11(2): 392–410, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deuflhard:1987:ENP

[DFK87] P. Deuflhard, B. Fiedler, and P. Kunkel. Efficient numerical pathfollowing beyond critical points. SIAM Journal on Numerical Analysis, 24(4): 912–927, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deVilliers:1981:CMN

[dG81] Noël de Villiers and David Glasser. A continuation method for nonlinear regression. SIAM Journal on Numerical Analysis, 18(6):1139– 1154, December 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:1990:FEA

[DG90] Qiang Du and Max D. Gunzburger. Finite-element approximations of a ladyzhenskaya model for stationary incompressible viscous flow. SIAM Journal on Numerical Analysis, 27(1):1–19, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Don:1994:CLM

[DG94] Wai-Sun Don and David Gottlieb. The Chebyshev– Legendre method: Implementing Legendre methods on Chebyshev points. SIAM Journal on Numerical Analysis, 31(6):1519–1534, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davis:1998:FEA

[DG98a] Timothy A. Davis and Eugene C. Gartland, Jr. Finite element analysis of the landau-de gennes minimization problem for liquid crystals. SIAM Journal on Numerical Analysis, 35(1):336– 362, February 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29744.

Don:1998:SSS

[DG98b] Wai Sun Don and David Gottlieb. Spectral simulation of supersonic reactive flows. SIAM Journal on Numerical Analysis, 35(6): 2370–2384, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31896.

Delahaye:1982:SLC

[DGB82] J. P. Delahaye and B. Germain-Bonne. The set of logarithmically convergent sequences cannot be accelerated. SIAM Journal on Numerical Analysis, 19(4):840–844, August 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1966:MAD

[DGP66] Jim Douglas, Jr., A. O. Garder, and Carl Pearcy. Multistage alternating direction methods. SIAM Journal on Numerical Analysis, 3 (4):570–581, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dutt:1996:FAP

[DGR96] A. Dutt, M. Gu, and V. Rokhlin. Fast algorithms for polynomial interpolation, integration, and differentiation. SIAM Journal on Numerical Analysis, 33(5):1689– 1711, October 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duran:1991:FEM

[DGW91] Ricardo Durán, Adriana Ghioldi, and Noemí Wolanski. A finite element method for the Mindlin–Reissner plate model. SIAM Journal on Numerical Analysis, 28(4):1004– 1014, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Donelson:1971:CCM

[DH71] John Donelson, III and Eldon Hansen. Cyclic composite multistep predictor-corrector methods. SIAM Journal on Numerical Analysis, 8(1): 137–157, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deuflhard:1979:AIC

[DH79] P. Deuflhard and G. Heindl.

Affine invariant convergence theorems for Newton’s method and extensions to related methods. SIAM Journal on Numerical Analysis, 16(1):

1–10, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dyksen:1984:PCG

[DHLR84] W. R. Dyksen, E. N. Houstis, R. E. Lynch, and J. R. Rice. The performance of the collocation and Galerkin methods with Hermite bicubics. SIAM Journal on Numerical Analysis, 21(4):695–715, August 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoog:1973:HOM

[dHW73] Frank de Hoog and Richard Weiss. High order methods for Volterra integral equations of the first kind. SIAM Journal on Numerical Analysis, 10 (4):647–664, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Diaz:1977:CGM

[Dia77]  Julio Cesar Diaz.  A

collocation-Galerkin method for the two point boundary value problem using continuous piecewise polynomial spaces. SIAM Journal on Numerical Analysis, 14(5):844– 858, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Diaz:1979:CGM

[Día79] Julio César Díaz. Collocation-H^{−1}-Galerkin method for parabolic problems with time dependent coefficients. SIAM Journal on Numerical Analysis, 16(6):911–922, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dierckx:1982:FAS

[Die82] Paul Dierckx. A fast algorithm for smoothing data on a rectangular grid while using spline functions. SIAM Journal on Numerical Analysis, 19(6):1286–1304, December 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Diener:1990:IDS

[Die90] Dwight Diener. Instability in the dimension of spaces of bivariate piecewise polynomials of degree 2r and smoothness order r. SIAM Journal on Numerical Analysis, 27 (2):543–551, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dieci:1992:NID

[Die92] Luca Dieci. Numerical integration of the differential Riccati equation and some related issues. SIAM Journal on Numerical Analysis, 29(3): 781–815, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dixon:1983:EEE

[Dix83] John D. Dixon. Estimating extremal eigenvalues and condition numbers of matrices. SIAM Journal on Numerical Analysis, 20(4):812– 814, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dellnitz:1999:ACD

[DJ99] Michael Dellnitz and Oliver Junge. On the approximation of complicated dynamical behavior. SIAM Journal on Numerical Analysis, 36(2):491–515, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31300.

Davis:1970:REP

[DK70] Chandler Davis and W. M. Kahan. The rotation of eigenvectors by a perturbation. III. SIAM Journal on Numerical Analysis, 7(1):1–46, March 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Decker:1980:NMSa

[DK80a] D. W. Decker and C. T. Kelley. Newton’s method at singular points. I. SIAM Journal on Numerical Analysis, 17(1):

66–70, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Decker:1980:NMSb

[DK80b] D. W. Decker and C. T. Kelley. Newton’s method at singular points. II. SIAM Journal on Numerical Analysis, 17 (3):465–471, June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dax:1981:EMC

[DK81] A. Dax and S. Kaniel. The ELR method for computing the eigenvalues of a general matrix. SIAM Journal on Numerical Analysis, 18(4):597– 605, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Decker:1982:CAN

[DK82] D. W. Decker and C. T. Kelley. Convergence acceleration for Newton’s method at singular points. SIAM Journal on Numerical Analysis, 19(1): 219–229, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Decker:1985:BMC

[DK85] D. W. Decker and C. T. Kelley. Broyden’s method for a class of problems having singular Jacobian at the root. SIAM Journal on Numerical Analysis, 22(3):566–574, June 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1986:CLS

[dK86] C. de Boor and H.-O. Kreiss. On the condition of the linear systems associated with discretized BVPs of ODEs. SIAM Journal on Numerical Analysis, 23(5):936–939, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

De-kang:1999:ESC

[Dk99]  Mao De-kang.  Entropy

satisfaction of a conservative shock-tracking method. SIAM Journal on Numerical Analysis, 36(2):529–550, April 1999. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31975.

Decker:1983:CRN

[DKK83] D. W. Decker, H. B. Keller, and C. T. Kelley. Convergence rates for Newton’s method at singular points. SIAM Journal on Numerical Analysis, 20(2): 296–314, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dupont:1968:AFP

[DKR68] Todd Dupont, Richard P.

Kendall, and H. H. Rachford, Jr. An approximate factorization procedure for solving self-adjoint elliptic difference equations. SIAM Journal on Numerical Analysis, 5 (3):559–573, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davis:1982:NPD

[DKW82] Chandler Davis, W. M.

Kahan, and H. F. Weinberger. Norm-preserving dilations and their applications to optimal error bounds. SIAM Journal on Numerical Analysis, 19(3):445–469, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dua:1973:FRD

[DL73] S. N. Dua and H. L. Loeb.

Further remarks on the differential correction algorithm. SIAM Journal on Numerical Analysis, 10(1): 123–126, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duris:1975:CQR

[DL75] C. S. Duris and J. N. Lyness.

Compound quadrature rules for the product of two functions. SIAM Journal on Numerical Analysis, 12(5):681– 697, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Descloux:1982:FEM

[DL82] Jean Descloux and Mitchell Luskin. On a finite element method to solve the criticality eigenvalue problem for the transport equation. SIAM Journal on Numerical Analysis, 19(6):1208–1219, December 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dyn:1983:ISS

[DL83] Nira Dyn and David Levin. Iterative solution of systems originating from integral equations and surface interpolation. SIAM Journal on Numerical Analysis, 20(2): 377–390, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deshpande:1990:SSF

[DL90] Lalita N. Deshpande and Balmohan V. Limaye. On the stability of singular finite rank methods. SIAM Journal on Numerical Analysis, 27(3): 792–803, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duncan:1991:JII

[DL91] D. B. Duncan and M. A. M. Lynch. Jacobi iteration in implicit difference schemes for the wave equation. SIAM Journal on Numerical Analysis, 28(6):1661–1679, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dieci:1995:CIT

[DL95] Luca Dieci and Jens Lorenz.

Computation of invariant tori by the method of characteristics. SIAM Journal on Numerical Analysis, 32(5):1436– 1474, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahlquist:1983:STS

[DLN83] Germund G. Dahlquist, Werner Liniger, and Olavi Nevanlinna. Stability of two-step methods for variable integration steps. SIAM Journal on Numerical Analysis, 20 (5):1071–1085, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dyn:1982:CBI

[DLR82] N. Dyn, G. G. Lorentz, and S. D. Riemenschneider. Continuity of the Birkhoff interpolation. SIAM Journal on Numerical Analysis, 19(3): 507–509, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Daniel:1977:NMD

[DM77] James W. Daniel and Andrew J. Martin. Numerov’s method with deferred corrections for two-point boundaryvalue problems. SIAM Journal on Numerical Analysis, 14(6):1033–1050, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Doss:1979:DAM

[DM79] Said Doss and Keith Miller. Dynamic ADI methods for elliptic equations. SIAM Journal on Numerical Analysis, 16(5):837–856, October 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahmen:1982:LIM

[DM82] Wolfgang A. Dahmen and Charles A. Micchelli. On the linear independence of multivariate B-splines. I. triangulations of simploids. SIAM Journal on Numerical Analysis, 19(5):993–1012, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deHoog:1987:DWC

[dM87] F. R. de Hoog and R. M. M. Mattheij. On dichotomy and well conditioning in BVP. SIAM Journal on Numerical Analysis, 24(1):89– 105, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1988:CIP

[DM88] Craig C. Douglas and Willard L.

Miranker. Constructive interference in parallel algorithms. SIAM Journal on Numerical Analysis, 25(2): 376–398, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Devulder:1992:CNA

[DM92] Christophe Devulder and Martine Marion. A class of numerical algorithms for large time integration: The nonlinear Galerkin methods. SIAM Journal on Numerical Analysis, 29(2):462–483, April 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahmen:1993:URE

[DM93] Wolfgang Dahmen and Charles A.

Micchelli. Using the refinement equation for evaluating integrals of wavelets. SIAM Journal on Numerical Analysis, 30(2):507–537, April 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dempster:1995:PGC

[DM95] M. A. H. Dempster and R. R.

Merkovsky. A practical geometrically convergent cutting plane algorithm. SIAM Journal on Numerical Analysis, 32 (2):631–644, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Destuynder:1998:EEB

[DM98] Philippe Destuynder and Brigitte Métivet. Explicit error bounds for a nonconforming finite element method. SIAM Journal on Numerical Analysis, 35(5): 2099–2115, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30019.

Deutsch:1975:SAC

[DMP75] Frank Deutsch, John H. Mc-Cabe, and George M. Phillips. Some algorithms for computing best approximations from convex cones. SIAM Journal on Numerical Analysis, 12 (3):390–403, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duran:1992:AEE

[DMR92] Ricardo Durán, María Amelia Muschietti, and Rodolfo Rodríguez. Asymptotically exact error estimators for rectangular finite elements. SIAM Journal on Numerical Analysis, 29(1):78–88, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dongarra:1983:IAC

[DMW83] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. SIAM Journal on Numerical Analysis, 20(1): 23–45, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic). URL http://www.
netlib.org/utk/people/JackDongarra/
PAPERS/Improving-the-Accuracy-
of-Computed-Eigenvalues-
and-Eigenvectors.pdf. Cited

in [?].

Duran:1989:PAS

[DN89] Ricardo G. Durán and Ricardo H. Nochetto. Pointwise accuracy of a stable Petrov– Galerkin approximation to the Stokes problem. SIAM Journal on Numerical Analysis, 26(6):1395–1406, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:1991:NAC

[DN91] Qiang Du and R. A. Nicolaides. Numerical analysis of a continuum model of phase transition. SIAM Journal on Numerical Analysis, 28 (5):1310–1322, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deift:1983:ODE

[DNT83] P. Deift, T. Nanda, and C. Tomei. Ordinary differential equations and the symmetric eigenvalue problem. SIAM Journal on Numerical Analysis, 20(1):1–22, February 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:1998:ACC

[DNW98] Qiang Du, R. A. Nicolaides, and Xiaonan Wu. Analysis and convergence of a covolume approximation of the Ginzburg–Landau model of superconductivity. SIAM Journal on Numerical Analysis, 35(3):1049– 1072, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30285.

Deacon:1979:FEM

[DO79] Arthur G. Deacon and Stanley Osher. A finite element method for a boundary value problem of mixed type. SIAM Journal on Numerical Analysis, 16(5):756–778, October 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Delfour:1992:NAS

[DO92] M. C. Delfour and A. Ouansafi. Noniterative approximations to the solution of the matrix Riccati differential equation. SIAM Journal on Numerical Analysis, 29(6):1648– 1693, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dobrowolski:1980:CLF

[Dob80] Manfred Dobrowolski. L^{∞}- convergence of linear finite element approximation to nonlinear parabolic problems. SIAM Journal on Numerical Analysis, 17(5):663– 674, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dobrowolski:1992:MFE

[Dob92]  Manfred Dobrowolski.  A

mixed finite element method for approximating incompressible materials. SIAM Journal on Numerical Analysis, 29(2):365–389, April 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Doedel:1978:CFD

[Doe78] Eusebius J. Doedel. The construction of finite difference approximations to ordinary differential equations. SIAM Journal on Numerical Analysis, 15(3):450–465, June 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Doedel:1979:FDC

[Doe79]  Eusebius J. Doedel.  Fi-

nite difference collocation methods for nonlinear twopoint boundary value problems. SIAM Journal on Numerical Analysis, 16(2): 173–185, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dolzmann:1999:NCR

[Dol99] Georg Dolzmann. Numerical computation of rank-one convex envelopes. SIAM Journal on Numerical Analysis, 36 (5):1621–1635, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32558.

Donnelly:1969:EMR

[Don69] J. D. P. Donnelly. Eigenvalues of membranes with reentrant corners. SIAM Journal on Numerical Analysis, 6 (1):47–61, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Donnelly:1970:BES

[Don70] J. D. P. Donnelly. Bounds for the eigenvalues of selfadjoint operators. SIAM Journal on Numerical Analysis, 7(3):458–478, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Donaldson:1973:EUB

[Don73] J. D. Donaldson. Estimates of upper bounds for quadrature errors. SIAM Journal on Numerical Analysis, 10(1): 13–22, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dontchev:1981:EED

[Don81] Asen L. Dontchev. Error estimates for a discrete approximation to constrained control problems. SIAM Journal on Numerical Analysis, 18 (3):500–514, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Donat:1994:SEP

[Don94] Rosa Donat. Studies on error propagation for certain nonlinear approximations to hyperbolic equations: Discontinuities in derivatives. SIAM Journal on Numerical Analysis, 31(3):655–679, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Doob:1966:RBL

[Doo66] J. L. Doob. Remarks on the boundary limits of harmonic functions. SIAM Journal on Numerical Analysis, 3 (2):229–235, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorr:1970:NSS

[Dor70] Fred W. Dorr. The numerical solution of singular perturbations of boundary value problems. SIAM Journal on Numerical Analysis, 7(2): 281–313, June 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorr:1984:ATV

[Dor84] Milo R. Dorr. The approximation theory for the p-version of the finite element method. SIAM Journal on Numerical Analysis, 21(6): 1180–1207, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorr:1986:ASE

[Dor86] Milo R. Dorr. The approximation of solutions of elliptic boundary-value problems via the p-version of the finite element method. SIAM Journal on Numerical Analysis, 23(1): 58–77, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dieci:1988:RTMa

[DOR88a] Luca Dieci, Michael R. Osborne, and Robert D. Russell. A Riccati transformation method for solving linear BVPs. I. theoretical aspects. SIAM Journal on Numerical Analysis, 25(5):1055– 1073, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dieci:1988:RTMb

[DOR88b] Luca Dieci, Michael R. Osborne, and Robert D. Russell. A Riccati transformation method for solving linear BVPs. II. computational aspects. SIAM Journal on Numerical Analysis, 25(5):1074– 1092, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorfler:1996:CAA

[Dör96] Willy Dörfler. A convergent adaptive algorithm for Poisson’s equation. SIAM Journal on Numerical Analysis, 33(3): 1106–1124, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorfler:1999:UPE

[Dör99a]  W. Dörfler.  Uniform A

priori estimates for singularly perturbed elliptic equations in multidimensions. SIAM Journal on Numerical Analysis, 36(6):1878– 1900, December 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34132.

Dorfler:1999:UEE

[Dör99b] W. Dörfler. Uniform error estimates for an exponentially fitted finite element method for singularly perturbed elliptic equations. SIAM Journal on Numerical Analysis, 36(6): 1709–1738, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34131.

Douglas:1983:FDM

[Dou83] Jim Douglas, Jr. Finite difference methods for twophase incompressible flow in porous media. SIAM Journal on Numerical Analysis, 20(4): 681–696, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1984:MGA

[Dou84] Craig C. Douglas. Multigrid algorithms with applications to elliptic boundary value problems. SIAM Journal on Numerical Analysis, 21 (2):236–254, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1985:SPS

[Dou85] Jim Douglas, Jr. Superconvergence in the pressure in the simulation of miscible displacement. SIAM Journal on Numerical Analysis, 22(5): 962–969, October 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Doucette:1993:CMN

[Dou93] Robert L. Doucette. A collocation method for the numerical solution of laplace’s equation with nonlinear boundary conditions on a polygon. SIAM Journal on Numerical Analysis, 30(3):717–732, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deuflhard:1992:AMI

[DP92] Peter Deuflhard and Florian A. Potra. Asymptotic mesh independence of Newton–Galerkin methods via a refined Mysovskiĭ theorem. SIAM Journal on Numerical Analysis, 29(5):1395– 1412, October 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davini:1998:RNC

[DP98] Cesare Davini and Igino Pitacco. Relaxed notions of curvature and a lumped strain method for elastic plates. SIAM Journal on Numerical Analysis, 35(2): 677–691, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29610.

Delfour:1987:ANP

[DPZ87] M. C. Delfour, G. Payre, and J.-P. Zolésio. Approximation of nonlinear problems associated with radiating bodies in space. SIAM Journal on Numerical Analysis, 24(5):1077– 1094, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Davis:1965:ISA

[DR65] Philip J. Davis and Philip Rabinowitz. Ignoring the singularity in approximate integration. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):367– 383, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

denHeijer:1981:SAC

[dR81] C. den Heijer and W. C. Rheinboldt. On steplength algorithms for a class of continuation methods. SIAM Journal on Numerical Analysis, 18(5):925–948, October 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1982:NMC

[DR82] Jim Douglas, Jr. and Thomas F. Russell. Numerical methods for convection-dominated diffusion problems based on combining the method of characteristics with finite element or finite difference procedures. SIAM Journal on Numerical Analysis, 19(5): 871–885, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Descloux:1983:NIP

[DR83] Jean Descloux and Jacques Rappaz. A nonlinear inverse power method with shift. SIAM Journal on Numerical Analysis, 20(6):1147– 1152, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Doedel:1984:FDM

[DR84] E. J. Doedel and G. W.

Reddien. Finite difference methods for singular twopoint boundary value problems. SIAM Journal on Numerical Analysis, 21(2): 300–313, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dai:1990:CMF

[DR90] Rui-Xiu Dai and Werner C.

Rheinboldt. On the computation of manifolds of foldpoints for parameterdependent problems. SIAM Journal on Numerical Analysis, 27(2):437–446, April 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Draghicescu:1994:EIP

[Dra94] C. I. Draghicescu. An efficient implementation of particle methods for the incompressible Euler equations. SIAM Journal on Numerical Analysis, 31(4):1090–1108, August 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Driscoll:1999:NDD

[Dri99] Tobin A. Driscoll. A nonoverlapping domain decomposition method for Symm’s equation for conformal mapping. SIAM Journal on Numerical Analysis, 36(3): 922–934, June 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32416.

Drmac:1998:ACP

[Drm98a]  Zlatko Drmac.  Accurate

computation of the productinduced singular value decomposition with applications. SIAM Journal on Numerical Analysis, 35(5): 1969–1994, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29263.

Drmac:1998:TAC

[Drm98b] Zlatko Drmac. A tangent algorithm for computing the generalized singular value decomposition. SIAM Journal on Numerical Analysis, 35 (5):1804–1832, October 1998.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28988.

Didenko:1995:AMS

[DRS95] V. D. Didenko, S. Roch, and B. Silbermann. Approximation methods for singular integral equations with conjugation on curves with corners. SIAM Journal on Numerical Analysis, 32(6):1910– 1939, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dieci:1994:UIA

[DRV94] Luca Dieci, Robert D. Russell, and Erik S. Van Vleck. Unitary integrators and applications to continuous orthonormalization techniques. SIAM Journal on Numerical Analysis, 31(1):261–281, February 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dieci:1997:CLE

[DRV97] Luca Dieci, Robert D. Russell, and Erik S. Van Vleck. On the computation of Lyapunov exponents for continuous dynamical systems. SIAM Journal on Numerical Analysis, 34(1):402–423, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/24731.

Dawson:1989:SIE

[DRW89] C. N. Dawson, T. F. Russell, and M. F. Wheeler. Some improved error estimates for the modified method of characteristics. SIAM Journal on Numerical Analysis, 26(6):1487– 1512, December 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dryja:1978:ADG

[Dry78] M. Dryja. Alternating direction Galerkin methods for the systems of hyperbolic equations of first order. SIAM Journal on Numerical Analysis, 15(1):81–89, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dryja:1983:FEC

[Dry83] M. Dryja. A finite element - capacitance matrix method for the elliptic problem. SIAM Journal on Numerical Analysis, 20(4):671–680, August 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duris:1968:CSL

[DS68] C. S. Duris and V. P. Sreedharan. Chebyshev and l^{1}- solutions of linear equations using least squares solutions. SIAM Journal on Numerical Analysis, 5(3):491– 505, September 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1972:SMP

[DS72] J. E. Dennis, Jr. and Roland A. Sweet. Some minimum properties of the trapezoidal rule. SIAM Journal on Numerical Analysis, 9(2): 230–236, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deBoor:1973:CGP

[dS73] Carl de Boor and Blair Swartz. Collocation at Gaussian points. SIAM Journal on Numerical Analysis, 10 (4):582–606, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dougalis:1980:SGA

[DS80] Vassilios A. Dougalis and Steven M. Serbin. On the superconvergence of Galerkin approximations to secondorder hyperbolic equations. SIAM Journal on Numerical Analysis, 17(3):431–446, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

DeVore:1984:EBG

[DS84] Ronald A. DeVore and L. Ridgway Scott. Error bounds for Gaussian quadrature and weighted-L^{1} polynomial approximation. SIAM Journal on Numerical Analysis, 21(2):400–412, April 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1986:SPA

[DS86] J. E. Dennis, Jr. and Trond Steihaug. On the successive projections approach to least-squares problems. SIAM Journal on Numerical Analysis, 23(4):717–733, August 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Douglas:1989:USA

[DS89a] Craig C. Douglas and Barry F. Smith. Using symmetries and antisymmetries to analyze a parallel multigrid algorithm: the elliptic boundary value problem case. SIAM Journal on Numerical Analysis, 26(6): 1439–1461, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dubeau:1989:DAF

[DS89b] François Dubeau and Jean Savoie. Développements asymptotiques de fonctions splines avec partage uniforme de la droite Réelle. (French) [asymptotic expansions of spline functions with uniform partition of the real line]. SIAM Journal on Numerical Analysis, 26(2): 468–479, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Disney:1992:LIR

[DS92] Shaun Disney and Ian H.

Sloan. Lattice integration rules of maximal rank formed by copying rank 1 rules. SIAM Journal on Numerical Analysis, 29(2): 566–577, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dubeau:1997:BEE

[DS97] François Dubeau and Jean Savoie. Best error estimates for odd and even degree deficient splines. SIAM Journal on Numerical Analysis, 34(3):1167–1184, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25928.

Dryja:1994:SAI

[DSW94] Maksymilian Dryja, Barry F.

Smith, and Olof B. Widlund. Schwarz analysis of iterative substructuring algorithms for elliptic problems in three dimensions. SIAM Journal on Numerical Analysis, 31(6):1662–1694, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duris:1973:FSA

[DT73] C. S. Duris and M. G. Temple. A finite step algorithm for determining the “strict” Chebyshev solution to ax = b. SIAM Journal on Numerical Analysis, 10(4):690–699, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deAngelis:1993:IPP

[dT93] P. L. de Angelis and G. Toraldo. On the identification property of a projected gradient method. SIAM Journal on Numerical Analysis, 30 (5):1483–1497, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1976:ATM

[DTW76] J. E. Dennis, Jr., J. F. Traub, and R. P. Weber. The algebraic theory of matrix polynomials. SIAM Journal on Numerical Analysis, 13(6):831– 845, December 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1978:ASM

[DTW78] J. E. Dennis, Jr., J. F. Traub, and R. P. Weber. Algorithms for solvents of matrix polynomials. SIAM Journal on Numerical Analysis, 15(3): 523–533, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dubois:1990:DVP

[Dub90] François Dubois. Discrete vector potential representation of a divergence-free vector field in three-dimensional domains: Numerical analysis of a model problem. SIAM Journal on Numerical Analysis, 27(5):1103–1141, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duffy:1982:QPC

[Duf82] Michael G. Duffy. Quadrature over a pyramid or cube of integrands with a singularity at a vertex. SIAM Journal on Numerical Analysis, 19(6):1260–1262, December 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1971:CAL

[Dun71] Charles B. Dunham. Chebyshev approximation with the local Haar condition. SIAM Journal on Numerical Analysis, 8(4):749–753, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunaway:1974:CZR

[Dun74a] Donna K. Dunaway. Calculation of zeros of a real polynomial through factorization using Euclid’s algorithm. SIAM Journal on Numerical Analysis, 11(6):1087–1104, December 1974. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1974:CSE

[Dun74b] Charles B. Dunham. Convergence of Stiefel’s exchange method on an infinite set. SIAM Journal on Numerical Analysis, 11(4):729–731, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1975:NMS

[Dun75] C. B. Dunham. Nonlinear mean-square approximation on finite sets. SIAM Journal on Numerical Analysis, 12(1):105–110, March 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1980:BOS

[Dun80a] Charles B. Dunham. Behavior of the one-sided alternating Chebyshev operator. SIAM Journal on Numerical Analysis, 17(3):428–430, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1980:RCD

[Dun80b] Charles B. Dunham. The rate of convergence of discretization in Chebyshev approximation on rectangular sets. SIAM Journal on Numerical Analysis, 17(5):635– 638, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunham:1980:SDC

[Dun80c] Charles B. Dunham. Stability of differential correction for rational Chebyshev approximation. SIAM Journal on Numerical Analysis, 17(5): 639–640, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duncan:1997:SFD

[Dun97]  D. B. Duncan.  Sym-

plectic finite difference approximations of the nonlinear Klein–Gordon equation. SIAM Journal on Numerical Analysis, 34(5): 1742–1760, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24310.

Dupont:1968:FPS

[Dup68] Todd Dupont. A factorization procedure for the solution of elliptic difference equations. SIAM Journal on Numerical Analysis, 5(4):753– 782, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dupont:1973:GMF

[Dup73a] Todd Dupont. Galerkin methods for first order hyperbolics: An example. SIAM Journal on Numerical Analysis, 10(5): 890–899, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dupont:1973:EGM

[Dup73b] Todd Dupont. L^{2}-estimates for Galerkin methods for second order hyperbolic equations. SIAM Journal on Numerical Analysis, 10(5):880– 889, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dupont:1976:UTS

[Dup76] Todd Dupont. A unified theory of superconvergence for Galerkin methods for twopoint boundary problems. SIAM Journal on Numerical Analysis, 13(3):362–368, June 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duris:1977:DIS

[Dur77] C. S. Duris. Discrete interpolating and smoothing spline functions. SIAM Journal on Numerical Analysis, 14 (4):686–698, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duran:1983:PAS

[Dur83]  Ricardo G. Durán.  On

polynomial approximation in Sobolev spaces. SIAM Journal on Numerical Analysis, 20(5):985–988, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duran:1988:NCL

[Dur88a] Ricardo G. Durán. A note on the convergence of linear finite elements. SIAM Journal on Numerical Analysis, 25 (5):1032–1036, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duran:1988:AMD

[Dur88b] Ricardo G. Durán. On the approximation of miscible displacement in porous media by a method of characteristics combined with a mixed method. SIAM Journal on Numerical Analysis, 25(5): 989–1001, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dussault:1995:NSE

[Dus95] Jean-Pierre Dussault. Numerical stability and efficiency of penalty algorithms. SIAM Journal on Numerical Analysis, 32(1):296–317, February 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dutt:1988:SBC

[Dut88] Pravir Dutt. Stable boundary conditions and difference schemes for Navier–Stokes equations. SIAM Journal on Numerical Analysis, 25(2): 245–267, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dutt:1990:SMI

[Dut90] P. Dutt. Spectral methods for initial boundary value problems — an alternative approach. SIAM Journal on Numerical Analysis, 27(4):885– 903, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dobson:1997:CIM

[DV97] David C. Dobson and Curtis R. Vogel. Convergence of an iterative method for total variation denoising. SIAM Journal on Numerical Analysis, 34(5): 1779–1791, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28701.

Dawson:1994:CGM

[DVW94] C. N. Dawson, C. J. Van Duijn, and M. F. Wheeler. Characteristic-galerkin methods for contaminant transport with nonequilibrium adsorption kinetics. SIAM Journal on Numerical Analysis, 31(4): 982–999, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoog:1974:HOM

[DW74] Frank De Hoog and Richard Weiss. High order methods for a class of Volterra integral equations with weakly singular kernels. SIAM Journal on Numerical Analysis, 11(6): 1166–1180, December 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

DeHoog:1976:DMB

[DW76a] Frank R. De Hoog and Richard Weiss. Difference methods for boundary value problems with a singularity of the first kind. SIAM Journal on Numerical Analysis, 13(5): 775–813, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dunn:1976:SCG

[DW76b] Roderick J. Dunn, Jr. and Mary Fanett Wheeler. Some collocation-Galerkin methods for two-point boundary value problems. SIAM Journal on Numerical Analysis, 13(5): 720–733, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoog:1978:CMS

[DW78] Frank R. De Hoog and Richard Weiss. Collocation methods for singular boundary value problems. SIAM Journal on Numerical Analysis, 15(1):198–217, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

deHoog:1979:NSB

[dW79] Frank R. de Hoog and Richard Weiss. The numerical solution of boundary value problems with an essential singularity. SIAM Journal on Numerical Analysis, 16(4): 637–669, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1981:CTL

[DW81] J. E. Dennis, Jr. and Homer F. Walker. Convergence theorems for leastchange secant update methods. SIAM Journal on Numerical Analysis, 18(6):949– 987, December 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [DW82].

Dennis:1982:ECT

[DW82] J. E. Dennis, Jr. and Homer F. Walker. Erratum: “Convergence Theorems for Least-Change Secant Update Methods” [SIAM J. Numer. Anal. 18 (1981), no. 6, 949–987]. SIAM Journal on Numerical Analysis, 19(2):443, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [DW81].

Dennis:1985:LCS

[DW85] J. E. Dennis, Jr. and Homer F. Walker. Leastchange sparse secant update methods with inaccurate secant conditions. SIAM Journal on Numerical Analysis, 22(4):760–778, August 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1993:SLC

[DW93] John E. Dennis, Jr. and H. Wolkowicz. Sizing and least-change secant methods. SIAM Journal on Numerical Analysis, 30(5):1291– 1314, October 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:1999:NST

[DW99] Qiang Du and Xiaonan Wu.

Numerical solution of the three-dimensional Ginzburg– Landau models using artificial boundary. SIAM Journal on Numerical Analysis, 36 (5):1482–1506, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33031.

Dawson:1998:TGF

[DWW98] Clint N. Dawson, Mary F.

Wheeler, and Carol S. Woodward. A two-grid finite difference scheme for nonlinear parabolic equations.

SIAM Journal on Numerical Analysis, 35(2):435–452, April 1998. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/29349.

Dyksen:1987:TPG

[Dyk87] Wayne R. Dyksen. Tensor product generalized ADI methods for separable elliptic problems. SIAM Journal on Numerical Analysis, 24(1): 59–76, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dziuk:1999:DAC

[Dzi99]  Gerhard Dziuk.  Discrete

anisotropic curve shortening flow. SIAM Journal on Numerical Analysis, 36(6): 1808–1830, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33753.

E:1992:CSM

[E92] Weinan E. Convergence of spectral methods for Burgers’ equation. SIAM Journal on Numerical Analysis, 29(6): 1520–1541, December 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

El-Alem:1991:GCT

[EA91]  Mahmoud El-Alem.  A

global convergence theory for the Celis–Dennis–Tapia trustregion algorithm for constrained optimization. SIAM Journal on Numerical Analysis, 28(1):266–290, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

El-Attar:1979:ANM

[EAVD79] R. A. El-Attar, M. Vidyasagar, and S. R. K. Dutta. An algorithm for L_{1}-norm minimization with application to nonlinear L_{1}-approximation. SIAM Journal on Numerical Analysis, 16(1):70–86, February 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ebmeyer:1998:EEC

[Ebm98] Carsten Ebmeyer. Error estimates for a class of degenerate parabolic equations. SIAM Journal on Numerical Analysis, 35(3):1095– 1112, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30520.

Elman:1993:OER

[EC93] Howard C. Elman and Michael P. Chernesky. Ordering effects on relaxation methods applied to the discrete one-dimensional convectiondiffusion equation. SIAM Journal on Numerical Analysis, 30(5):1268–1290, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elliott:1977:QRO

[ED77] David Elliott and J. D. Donaldson. On quadrature rules for ordinary and Cauchy principal value integrals over contours. SIAM Journal on Numerical Analysis, 14(6):1078– 1087, December 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Evans:1982:NGH

[ED82] D. J. Evans and A. Danaee. A new group hopscotch method for the numerical solution of partial differential equations. SIAM Journal on Numerical Analysis, 19(3):588–598, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisenstat:1983:VIM

[EES83] Stanley C. Eisenstat, Howard C. Elman, and Martin H. Schultz. Variational iterative methods for nonsymmetric systems of linear equations. SIAM Journal on Numerical Analysis, 20(2): 345–357, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elliott:1989:NFE

[EF89] Charles M. Elliott and Donald A. French. A nonconforming finite-element method for the two-dimensional Cahn– Hilliard equation. SIAM Journal on Numerical Analysis, 26(4):884–903, August 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ehrlich:1975:SDS

[EG75] Louis W. Ehrlich and Murli M.

Gupta. Some difference schemes for the biharmonic equation. SIAM Journal on Numerical Analysis, 12(5): 773–790, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elman:1994:IPU

[EG94] Howard C. Elman and Gene H. Golub. Inexact and preconditioned Uzawa algorithms for saddle point problems. SIAM Journal on Numerical Analysis, 31(6):1645– 1661, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eggermont:1983:CVI

[Egg83] P. P. B. Eggermont. Collocation for Volterra integral equations of the first kind with iterated kernel. SIAM Journal on Numerical Analysis, 20(5):1032–1048, October 1983. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Eggermont:1988:GMA

[Egg88] Paul P. B. Eggermont. On Galerkin methods for Abel-type integral equations. SIAM Journal on Numerical Analysis, 25(5):1093– 1117, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Erisman:1985:SSM

[EGLP85] A. M. Erisman, R. G. Grimes, J. G. Lewis, and W. G. Poole, Jr. A structurally stable modification of Hellerman– Rarick’s P^{4} algorithm for reordering unsymmetric sparse matrices. SIAM Journal on Numerical Analysis, 22(2): 369–385, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Epstein:1972:NQF

[EH72] M. P. Epstein and R. W. Hamming. Noninterpolatory quadrature formulas. SIAM Journal on Numerical Analysis, 9(3):464–475, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eberlein:1975:GCA

[EH75] P. J. Eberlein and C. P. Huang. Global convergence of the QR algorithm for unitary matrices with some results for normal matrices. SIAM Journal on Numerical Analysis, 12 (1):97–104, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Enright:1976:TRI

[EH76] W. H. Enright and T. E. Hull.

Test results on initial value methods for non-stiff ordinary differential equations. SIAM Journal on Numerical Analysis, 13(6):944–961, December 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Engquist:1989:PMA

[EH89] Björn Engquist and Thomas Y.

Hou. Particle method approximation of oscillatory solutions to hyperbolic differential equations. SIAM Journal on Numerical Analysis, 26 (2):289–319, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisen:1992:NMS

[EH92] Henner Eisen and Wilhelm Heinrichs. A new method of stabilization for singular perturbation problems with spectral methods. SIAM Journal on Numerical Analysis, 29(1): 107–122, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Enright:1998:CAS

[EH98] W. H. Enright and H. Hayashi. Convergence analysis of the solution of retarded and neutral delay differential equations by continuous numerical methods. SIAM Journal on Numerical Analysis, 35(2):572–585, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30204.

Ellison:1987:USC

[EHP87] J. H. Ellison, C. A. Hall, and T. A. Porsching. An unconditionally stable convergent finite difference method for Navier–Stokes problems on curved domains. SIAM Journal on Numerical Analysis, 24(6):1233–1248, December 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ehrlich:1971:SBE

[Ehr71] Louis W. Ehrlich. Solving the biharmonic equation as coupled finite difference equations. SIAM Journal on Numerical Analysis, 8(2): 278–287, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ehrich:1998:PIG

[Ehr98]  Sven Ehrich.  On prod-

uct integration with Gauss– Kronrod nodes. SIAM Journal on Numerical Analysis, 35(1):78–92, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29579.

Eisenstat:1995:RPT

[EI95] Stanley C. Eisenstat and Ilse C. F. Ipsen. Relative perturbation techniques for singular value problems. SIAM Journal on Numerical Analysis, 32(6):1972–1988, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eich:1993:CRC

[Eic93] Edda Eich. Convergence results for a coordinate projection method applied to mechanical systems with algebraic constraints. SIAM Journal on Numerical Analysis, 30 (5):1467–1482, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Einfeldt:1988:GTM

[Ein88] Bernd Einfeldt. On Godunovtype methods for gas dynamics. SIAM Journal on Numerical Analysis, 25(2): 294–318, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisen:1966:SCF

[Eis66] Dennis Eisen. Stability and convergence of finite difference schemes with singular coefficients. SIAM Journal on Numerical Analysis, 3 (4):545–552, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisen:1967:CCF

[Eis67a] Dennis Eisen. On the construction of consistent finite difference schemes with certain invariant subspaces. SIAM Journal on Numerical Analysis, 4(3):349–356, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisen:1967:NAF

[Eis67b] Dennis Eisen. On the numerical analysis of a fourth order wave equation. SIAM Journal on Numerical Analysis, 4(3):457–464, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisenstat:1974:RCB

[Eis74] Stanley C. Eisenstat. On the rate of convergence of the bergman-vekua method for the numerical solution of elliptic boundary value problems. SIAM Journal on Numerical Analysis, 11(3): 654–680, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisenstat:1983:NGC

[Eis83] Stanley C. Eisenstat. A note on the generalized conjugate gradient method. SIAM Journal on Numerical Analysis, 20 (2):358–361, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1987:EEA

[EJ87] Kenneth Eriksson and Claes Johnson. Error estimates and automatic time step control for nonlinear parabolic problems. I. SIAM Journal on Numerical Analysis, 24(1): 12–23, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1991:AFE

[EJ91] Kenneth Eriksson and Claes Johnson. Adaptive finite element methods for parabolic problems. I: A linear model problem. SIAM Journal on Numerical Analysis, 28(1): 43–77, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1995:AFEa

[EJ95a] Kenneth Eriksson and Claes Johnson. Adaptive finite element methods for parabolic problems. II. optimal error estimates in L_{∞}L_{2} and L_{∞}L_{∞}. SIAM Journal on Numerical Analysis, 32(3):706–740, June 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1995:AFEb

[EJ95b] Kenneth Eriksson and Claes Johnson. Adaptive finite element methods for parabolic problems. IV. nonlinear problems. SIAM Journal on Numerical Analysis, 32(6):1729– 1749, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1995:AFEc

[EJ95c] Kenneth Eriksson and Claes Johnson. Adaptive finite element methods for parabolic problems V: Long-time integration. SIAM Journal on Numerical Analysis, 32(6): 1750–1763, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisenstat:1985:OMP

[EJL85] S. C. Eisenstat, K. R. Jackson, and J. W. Lewis. The order of monotone piecewise cubic interpolation. SIAM Journal on Numerical Analysis, 22(6):1220–1237, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1998:AFE

[EJL98] Kenneth Eriksson, Claes Johnson, and Stig Larsson. Adaptive finite element methods for parabolic problems VI: Analytic semigroups. SIAM Journal on Numerical Analysis, 35(4): 1315–1325, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31021.

Efraimsson:1999:RNE

[EK99] Gunilla Efraimsson and Gunilla Kreiss. A remark on numerical errors downstream of slightly viscous shocks. SIAM Journal on Numerical Analysis, 36(3):853–863, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31328.

ElBaz:1990:FPA

[El 90] Didier El Baz. M-functions and parallel asynchronous algorithms. SIAM Journal on Numerical Analysis, 27(1): 136–140, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Estes:1972:SGP

[EL72] R. H. Estes and E. R. Lancaster. Some generalized power series inversions. SIAM Journal on Numerical Analysis, 9(2):241–247, June 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elliott:1973:EL

[EL73] David Elliott and Binh Lam. An estimate of E_{n}(f) for large n. SIAM Journal on Numerical Analysis, 10(6):1091– 1102, December 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ervin:1989:ADC

[EL89] V. Ervin and W. Layton. An analysis of a defectcorrection method for a model convection-diffusion equation. SIAM Journal on Numerical Analysis, 26(1):169–179, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Engl:1993:CRM

[EL93] Heinz W. Engl and Gerhard Landl. Convergence rates for maximum entropy regularization. SIAM Journal on Numerical Analysis, 30(5):1509– 1536, October 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

E:1996:PMI

[EL96] Weinan E and Jian-Guo Liu. Projection method.

II. Godunov–Ryabenki analysis. SIAM Journal on Numerical Analysis, 33(4): 1597–1621, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26450.

Engquist:1997:CMM

[EL97] Bjorn Engquist and Erding Luo. Convergence of a multigrid method for elliptic equations with highly oscillatory coefficients. SIAM Journal on Numerical Analysis, 34(6): 2254–2273, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28940.

Elden:1980:PTL

[Eld80] Lars Eldén. Perturbation theory for the least squares problem with linear equality constraints. SIAM Journal on Numerical Analysis, 17(3): 338–350, June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elliott:1982:CCM

[Ell82] David Elliott. The classical collocation method for singular integral equations. SIAM Journal on Numerical Analysis, 19(4):816–832, August 1982. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Ellacott:1983:FTE

[Ell83] S. W. Ellacott. On the Faber transform and efficient numerical rational approximation. SIAM Journal on Numerical Analysis, 20(5):989– 1000, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elliott:1984:RCM

[Ell84] David Elliott. Rates of convergence for the method of classical collocation for solving singular integral equations. SIAM Journal on Numerical Analysis, 21(1):136– 148, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1993:DDT

[ELPV93] Richard E. Ewing, Raytcho D. Lazarov, Joseph E. Pasciak, and Panayot S. Vassilevski. Domain decomposition type iterative techniques for parabolic problems on locally refined grids. SIAM Journal on Numerical Analysis, 30(6):1537–1557, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elton:1995:CCD

[ELR95] Bracy H. Elton, C. David Levermore, and Garry H.

Rodrigue. Convergence of convective-diffusive lattice Boltzmann methods. SIAM Journal on Numerical Analysis, 32(5):1327–1354, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eiermann:1989:HSI

[ELV89] M. Eiermann, X. Li, and R. S. Varga. On hybrid semiiterative methods. SIAM Journal on Numerical Analysis, 26(1):152–168, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1994:FDS

[ELV94] Richard E. Ewing, Raytcho D.

Lazarov, and Apostol T. Vassilev. Finite difference scheme for parabolic problems on composite grids with refinement in time and space. SIAM Journal on Numerical Analysis, 31(6):1605– 1622, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1991:SVA

[ELW91] R. E. Ewing, R. D. Lazarov, and J. Wang. Superconvergence of the velocity along the Gauss lines in mixed finite element methods. SIAM Journal on Numerical Analysis, 28(4): 1015–1029, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1999:SMF

[ELW99] Richard E. Ewing, Michael M. Liu, and Junping Wang. Superconvergence of mixed finite element approximations over quadrilaterals. SIAM Journal on Numerical Analysis, 36(3):772–787, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32280.

England:1988:BVP

[EM88] R. England and R. M. M. Mattheij. Boundary value problems and dichotomic stability. SIAM Journal on Numerical Analysis, 25(5):1037– 1054, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Espinosa-Maldonado:1971:CQF

[EMB71] Ruben J. Espinosa-Maldonado and George D. Byrne. On the convergence of quadrature formulas. SIAM Journal on Numerical Analysis, 8(1): 110–114, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eiermann:1983:CSM

[EN83] Michael Eiermann and Wilhelm Niethammer. On the construction of semiterative methods. SIAM Journal on Numerical Analysis, 20(6): 1153–1160, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Enright:1974:SDM

[Enr74]  W. H. Enright.  Second

derivative multistep methods for stiff ordinary differential equations. SIAM Journal on Numerical Analysis, 11(2): 321–331, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Enright:1989:AEC

[Enr89]  W. H. Enright.  Analy-

sis of error control strategies for continuous Runge– Kutta methods. SIAM Journal on Numerical Analysis, 26 (3):588–599, June 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Enright:1993:REA

[Enr93] W. H. Enright. The relative efficiency of alternative defect control schemes for high-order continuous Runge–Kutta formulas. SIAM Journal on Numerical Analysis, 30(5):1419– 1445, October 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Epperson:1982:EEC

[Epp82] James F. Epperson. An error estimate for changing the Stefan problem. SIAM Journal on Numerical Analysis, 19(1): 114–120, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Epperson:1984:FEM

[Epp84] James F. Epperson. Finite element methods for a class of nonlinear evolution equations. SIAM Journal on Numerical Analysis, 21(6):1066– 1079, December 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Epstein:1976:IPP

[Eps76] M. P. Epstein. On the influence of parametrization in parametric interpolation. SIAM Journal on Numerical Analysis, 13(2): 261–268, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1982:ETS

[ER82] Richard E. Ewing and Thomas F. Russell. Efficient timestepping methods for miscible displacement problems in porous media. SIAM Journal on Numerical Analysis, 19 (1):1–67, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Erdelyi:1965:ILS

[Erd65] Ivan Erdélyi. An iterative least-square algorithm suitable for computing partial eigensystems. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3): 421–436, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Ericksen:1980:IPM

[Eri80]  W. S. Ericksen.  Inverse

pairs of matrices with integer elements. SIAM Journal on Numerical Analysis, 17 (3):474–477, June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:1986:SEE

[Eri86]  Kenneth Eriksson.  Some

error estimates for the pversion of the finite element method. SIAM Journal on Numerical Analysis, 23(2): 403–411, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ehile:1976:DEC

[ES76] Gerd-Peter Ehile and Hubert Schwetlick. Discretized Euler–Chebyshev multistep methods. SIAM Journal on Numerical Analysis, 13(3): 432–447, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elman:1986:PFD

[ES86] Howard C. Elman and Martin H. Schultz. Preconditioning by fast direct methods for nonselfadjoint nonseparable elliptic equations. SIAM Journal on Numerical Analysis, 23(1):44–57, February 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Elliott:1993:GDD

[ES93] C. M. Elliott and A. M. Stuart. The global dynamics of discrete semilinear parabolic equations. SIAM Journal on Numerical Analysis, 30(6): 1622–1663, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Engquist:1998:CRF

[ES98] Bjorn Engquist and Björn Sjögreen. The convergence rate of finite difference schemes in the presence of shocks. SIAM Journal on Numerical Analysis, 35(6): 2464–2485, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31758.

Espelid:1987:CGF

[Esp87]  Terje O. Espelid.  On

the construction of good fully symmetric integration rules. SIAM Journal on Numerical Analysis, 24(4): 855–881, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Espinoza:1994:POS

[Esp94] Pedro C. Espinoza. Positive ordered solutions of a discrete analogue of a nonlinear elliptic eigenvalue problem. SIAM Journal on Numerical Analysis, 31(3):760–767, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Estep:1995:PEB

[Est95] Donald Estep. A Posteriori error bounds and global error control for approximation of ordinary differential equations. SIAM Journal on Numerical Analysis, 32(1): 1–48, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

El-Tom:1971:ISA

[ET71] M. E. A. El-Tom. On ignoring the singularity in approximate integration. SIAM Journal on Numerical Analysis, 8 (2):412–424, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Estivalezes:1996:HOP

[EV96] J. L. Estivalezes and P. Villedieu.

High-order positivity-preserving kinetic schemes for the compressible Euler equations. SIAM Journal on Numerical Analysis, 33(5):2050– 2067, October 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27100.

Evans:1980:SCT

[Eva80] D. J. Evans. On the solution of certain Toeplitz tridiagonal linear systems. SIAM Journal on Numerical Analysis, 17(5): 675–680, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisenberg:1972:OAU

[EW72] S. Eisenberg and B. Wood. On the order of approximation of unbounded functions by positive linear operators. SIAM Journal on Numerical Analysis, 9(2):266–276, June 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ellacott:1976:RCA

[EW76] S. Ellacott and Jack Williams. Rational Chebyshev approximation in the complex plane. SIAM Journal on Numerical Analysis, 13(3):310–323, June 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1980:GMM

[EW80] Richard E. Ewing and Mary Fanett Wheeler. Galerkin methods for miscible displacement problems in porous media. SIAM Journal on Numerical Analysis, 17(3):351–365, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ellner:1991:ADI

[EW91] Nancy S. Ellner and Eugene L. Wachspress. Alternating direction implicit iteration for systems with complex spectra. SIAM Journal on Numerical Analysis, 28 (3):859–870, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1996:OOE

[EW96] Richard E. Ewing and Hong Wang. An optimal-order estimate for Eulerian–Lagrangian localized adjoint methods for variable-coefficient advectionreaction problems. SIAM Journal on Numerical Analysis, 33(1):318–348, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1975:NSS

[Ewi75] Richard E. Ewing. Numerical solution of Sobolev partial differential equations. SIAM Journal on Numerical Analysis, 12(3):345–363, June 1975.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1978:TSG

[Ewi78] Richard E. Ewing. Timestepping Galerkin methods for nonlinear Sobolev partial differential equations. SIAM Journal on Numerical Analysis, 15(6):1125–1150, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ewing:1989:TAC

[EYL89] Richard E. Ewing, Yi Rang Yuan, and Gang Li. Timestepping along characteristics for a mixed finite-element approximation for compressible flow of contamination from nuclear waste in porous media. SIAM Journal on Numerical Analysis, 26(6):1513– 1524, December 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

El-Zorkany:1983:EAD

[EZB83] H. I. El-Zorkany and R. Balasubramanian. An extension of the alternating direction Galerkin method to more general geometries. SIAM Journal on Numerical Analysis, 20 (2):258–278, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fairweather:1972:GMV

[Fai72] Graeme Fairweather. Galerkin methods for vibration problems in two space variables. SIAM Journal on Numerical Analysis, 9(4):702– 714, December 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fairweather:1994:SCM

[Fai94] Graeme Fairweather. Spline collocation methods for a class of hyperbolic partial integro-differential equations. SIAM Journal on Numerical Analysis, 31(2): 444–460, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1978:ABE

[Fal78] Richard S. Falk. Approximation of the biharmonic equation by a mixed finite element method. SIAM Journal on Numerical Analysis, 15 (3):556–567, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Farrell:1966:AMS

[Far66] O. J. Farrell. On approximation measured by a surface integral. SIAM Journal on Numerical Analysis, 3 (2):236–247, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Farwig:1985:MTP

[Far85] R. Farwig. Multivariate truncated powers and B-splines with coalescent knots. SIAM Journal on Numerical Analysis, 22(3):592–603, June 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Farrell:1988:SCU

[Far88] Paul A. Farrell. Sufficient conditions for the uniform convergence of a difference scheme for a singularly perturbed turning point problem. SIAM Journal on Numerical Analysis, 25(3):618–643, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Favard:1964:CPS

[Fav64] J. Favard. On the comparison of the processes of summation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):38– 52, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Fazio:1996:NAN

[Faz96] Riccardo Fazio. A novel approach to the numerical solution of boundary value problems on infinite intervals. SIAM Journal on Numerical Analysis, 33(4): 1473–1483, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25204.

Fritsch:1980:MPC

[FC80] F. N. Fritsch and R. E. Carlson. Monotone piecewise cubic interpolation. SIAM Journal on Numerical Analysis, 17 (2):238–246, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Friedman:1991:NCC

[FD91] Mark J. Friedman and Eusebius J. Doedel. Numerical computation and continuation of invariant manifolds connecting fixed points. SIAM Journal on Numerical Analysis, 28(3):789–808, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fischer:1998:NSS

[FD98] Patrick Fischer and Mireille Defranceschi. Numerical solution of the Schrödinger equation in a wavelet basis for hydrogen-like atoms. SIAM Journal on Numerical Analysis, 35(1):1–12, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28455.

Feinberg:1982:SFA

[Fei82] Robert B. Feinberg. A_{0}- stable formulas of Adams type. SIAM Journal on Numerical Analysis, 19(2): 259–262, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ferguson:1973:SCP

[Fer73] David Ferguson. Sufficient conditions for Peano’s kernel to be of one sign. SIAM Journal on Numerical Analysis, 10(6):1047–1054, December 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ferguson:1974:SCM

[Fer74] David R. Ferguson. Sign changes and minimal support properties of Hermite– Birkhoff splines with compact support. SIAM Journal on Numerical Analysis, 11 (4):769–779, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ferguson:1986:ASP

[Fer86a] Warren E. Ferguson, Jr. Analysis of a singularly-perturbed linear two-point boundary value problem. SIAM Journal on Numerical Analysis, 23(5): 940–947, October 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ferguson:1986:RCC

[Fer86b] Warren E. Ferguson, Jr.

The rate of convergence of a class of block Jacobi schemes. SIAM Journal on Numerical Analysis, 23(2): 297–303, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Faddeev:1970:NNA

[FF70] D. K. Faddeev and V. N. Faddeeva. Natural norms in algebraic processes. SIAM Journal on Numerical Analysis, 7 (4):520–531, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fernandes:1991:ADG

[FF91] Ryan I. Fernandes and Graeme Fairweather. An alternating direction Galerkin method for a class of secondorder hyperbolic equations in two space variables. SIAM Journal on Numerical Analysis, 28(5):1265–1281, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Farhloul:1993:NMF

[FF93] M. Farhloul and M. Fortin. A new mixed finite element for the Stokes and elasticity problems. SIAM Journal on Numerical Analysis, 30(4):971– 990, September 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falcone:1998:CAC

[FF98] Maurizio Falcone and Roberto Ferretti. Convergence analysis for a class of highorder semi-Lagrangian advection schemes. SIAM Journal on Numerical Analysis, 35(3):909–940, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27351.

Feistauer:1999:EEC

[FFLMW99] Miloslav Feistauer, Jirí Felcman, Mária Lukácová-Medvid’ová, and Gerald Warnecke. Error estimates for a combined finite volume–finite element method for nonlinear convection–diffusion problems. SIAM Journal on Numerical Analysis, 36(5): 1528–1548, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31469.

Fletcher:1974:LMA

[FGH74] R. Fletcher, J. A. Grant, and M. D. Hebden. Linear minimax approximation as the limit of best L_{p}- approximation. SIAM Journal on Numerical Analysis, 11(1):123–136, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fix:1972:AIC

[FH72] G. Fix and R. Heiberger.

An algorithm for the illconditioned generalized eigenvalue problem. SIAM Journal on Numerical Analysis, 9 (1):78–88, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Feistauer:1996:APE

[FHK96] Miloslav Feistauer, George C.

Hsiao, and Ralph E. Kleinman. Asymptotic and A posteriori error estimates for boundary element solutions of hypersingular integral equations. SIAM Journal on Numerical Analysis, 33(2): 666–685, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fox:1967:ABE

[FHM67] L. Fox, P. Henrici, and C. Moler. Approximations and bounds for eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4(1):89–102, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Field:1978:EBE

[Fie78] David A. Field. Error bounds for elliptic convergence regions for continued fractions.

SIAM Journal on Numerical Analysis, 15(3):444–449, June 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Finlayson:1971:CGM

[Fin71] Bruce A. Finlayson. Convergence of the Galerkin method for nonlinear problems involving chemical reaction. SIAM Journal on Numerical Analysis, 8(2):316–324, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fisk:1982:OPN

[Fis82] Robert S. Fisk. On an oscillation phenomenon in the numerical solution of the diffusion-convection equation. SIAM Journal on Numerical Analysis, 19(4):721– 724, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fairgrieve:1991:KFM

[FJ91] Thomas F. Fairgrieve and Allan D. Jepson. O. K. Floquet multipliers. SIAM Journal on Numerical Analysis, 28 (5):1446–1462, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1976:PEM

[FK76] Richard S. Falk and J. Thomas King. A penalty and extrapolation method for the stationary Stokes equations. SIAM Journal on Numerical Analysis, 13(5):814–829, October 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fairweather:1984:GMS

[FKD84] Graeme Fairweather, Patrick Keast, and Julio César Díaz. On the H^{−1}-galerkin method for second-order linear twopoint boundary value problems. SIAM Journal on Numerical Analysis, 21(2): 314–326, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fix:1971:CSI

[FL71] George J. Fix and Kate Larsen. On the convergence of SOR iterations for finite element approximations to elliptic boundary value problems. SIAM Journal on Numerical Analysis, 8(3):536– 547, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fresnel:1992:NSI

[FL92] Christophe Fresnel and Marie-Noëlle Le Roux. Numerical solution of an interdiffusion problem. SIAM Journal on Numerical Analysis, 29(1): 194–208, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fairweather:1967:NCP

[FM67] G. Fairweather and A. R. Mitchell. A new computational procedure for A.D.I. methods. SIAM Journal on Numerical Analysis, 4(2): 163–170, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Faber:1984:NSC

[FM84] Vance Faber and Thomas Manteuffel. Necessary and sufficient conditions for the existence of a conjugate gradient method. SIAM Journal on Numerical Analysis, 21 (2):352–362, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Faber:1987:OEM

[FM87] Vance Faber and Thomas A. Manteuffel. Orthogonal error methods. SIAM Journal on Numerical Analysis, 24(1): 170–187, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1990:EFE

[FM90a] Richard S. Falk and Mary E. Morley. Equivalence of finite element methods for problems in elasticity. SIAM Journal on Numerical Analysis, 27(6): 1486–1505, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Frommer:1990:ONL

[FM90b] Andreas Frommer and Günter Mayer. On the R-order of Newton-like methods for enclosing solutions of nonlinear equations. SIAM Journal on Numerical Analysis, 27(1): 105–116, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Friedlander:1999:GMR

[FMMR99] A. Friedlander, J. M. Martínez, B. Molina, and M. Raydan. Gradient method with retards and generalizations. SIAM Journal on Numerical Analysis, 36 (1):275–289, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27315.

Farrell:1996:UCF

[FMOS96] Paul A. Farrell, John J. H.

Miller, Eugene O’Riordan, and Grigori I. Shishkin. A uniformly convergent finite difference scheme for a singularly perturbed semilinear equation. SIAM Journal on Numerical Analysis, 33(3): 1135–1149, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fix:1980:AMF

[FN80] G. J. Fix and R. A. Nicolaides. An analysis of mixed finite element approximations for periodic acoustic wave propagation. SIAM Journal on Numerical Analysis, 17 (6):779–786, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Feldstein:1984:HOM

[FN84] Alan Feldstein and Kenneth W. Neves. High order methods for state-dependent delay differential equations with nonsmooth solutions. SIAM Journal on Numerical Analysis, 21(5):844–863, October 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Friedland:1987:FAN

[FNO87] S. Friedland, J. Nocedal, and M. L. Overton. The formulation and analysis of numerical methods for inverse eigenvalue problems. SIAM Journal on Numerical Analysis, 24 (3):634–667, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fontecilla:1988:LCS

[Fon88] Rodrigo Fontecilla. Local convergence of secant methods for nonlinear constrained optimization. SIAM Journal on Numerical Analysis, 25(3): 692–712, June 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fontecilla:1990:ISM

[Fon90] Rodrigo Fontecilla. Inexact secant methods for nonlinear constrained optimization. SIAM Journal on Numerical Analysis, 27(1):154– 165, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ford:1965:MPI

[For65] Wayne T. Ford. Mathematical programming and integrodifferential equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):171–202, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Ford:1966:MPI

[For66a] Wayne T. Ford. Mathematical programming and integrodifferential equations: II. variable coefficients. SIAM Journal on Numerical Analysis, 3 (3):383–389, September 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ford:1966:MPL

[For66b]  Wayne T. Ford.  Math-

ematical programming and linear operators in frechet spaces. SIAM Journal on Numerical Analysis, 3(3):367– 371, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fornaro:1973:NEI

[For73] Robert J. Fornaro. Numerical evaluation of integrals around simple closed curves. SIAM Journal on Numerical Analysis, 10(4):623–634, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fornberg:1975:FMI

[For75] Bengt Fornberg. On a Fourier method for the integration of hyperbolic equations. SIAM Journal on Numerical Analysis, 12(4):509–528, September 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ford:1978:PDU

[For78] Wayne T. Ford. Polynomial deflation using regression analysis. SIAM Journal on Numerical Analysis, 15 (2):219–223, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fortuna:1979:SCP

[For79] Zenon Fortuna. Some convergence properties of the conjugate gradient method in Hilbert space. SIAM Journal on Numerical Analysis, 16 (3):380–384, June 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fornberg:1990:HOF

[For90] Bengt Fornberg. High-order finite differences and the pseudospectral method on staggered grids. SIAM Journal on Numerical Analysis, 27(4):904–918, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Foster:1981:GLM

[Fos81] Leslie V. Foster. Generalizations of Laguerre’s method: higher order methods. SIAM Journal on Numerical Analysis, 18(6):1004–1018, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Finney:1986:MNQ

[FP86] Patricia Hillhouse Finney and Thomas E. Price, Jr. Minimum norm quadratures which satisfy nonstandard interpolating conditions. SIAM Journal on Numerical Analysis, 23 (1):210–216, February 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Forster:1991:EEG

[FP91] Klaus-Jürgen Förster and Knut Petras. Error estimates in Gaussian quadrature for functions of bounded variation. SIAM Journal on Numerical Analysis, 28(3):

880–889, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Freistuhler:1995:NSR

[FP95] Heinrich Freistühler and E. Bruce Pitman. A numerical study of a rotationally degenerate hyperbolic system. Part II. the Cauchy problem. SIAM Journal on Numerical Analysis, 32(3):741–753, June 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Funaro:1988:IPI

[FQZ88] Daniele Funaro, Alfio Quarteroni, and Paola Zanolli. An iterative procedure with interface relaxation for domain decomposition methods. SIAM Journal on Numerical Analysis, 25(6):1213–1236, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Franklin:1968:NAE

[FR68] Joel N. Franklin and Eugene R. Rodemich. Numerical analysis of an ellipticparabolic partial differential equation. SIAM Journal on Numerical Analysis, 5 (4):680–716, December 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Frehse:1978:AEE

[FR78] Jens Frehse and Rolf Rannacher. Asymptotic L^{∞}- error estimates for linear finite element approximations of quasilinear boundary value problems. SIAM Journal on Numerical Analysis, 15(2): 418–431, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fink:1983:DEP

[FR83] James P. Fink and Werner C.

Rheinboldt. On the discretization error of parametrized nonlinear equations. SIAM Journal on Numerical Analysis, 20(4):732–746, August 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fink:1985:LEE

[FR85a] James P. Fink and Werner C.

Rheinboldt. Local error estimates for parametrized nonlinear equations. SIAM Journal on Numerical Analysis, 22(4):729–735, August 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fix:1985:CSF

[FR85b] George J. Fix and Milton E.

Rose. A comparative study of finite element and finite difference methods for Cauchy–Riemann type equations. SIAM Journal on Numerical Analysis, 22(2): 250–261, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fink:1986:FSM

[FR86] James P. Fink and Werner C. Rheinboldt. Folds on the solution manifold of a parametrized equation. SIAM Journal on Numerical Analysis, 23(4):693–706, August 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1987:ACF

[FR87a] Richard S. Falk and Gerard R. Richter. Analysis of a continuous finite element method for hyperbolic equations. SIAM Journal on Numerical Analysis, 24(2): 257–278, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fink:1987:GFN

[FR87b] James P. Fink and Werner C. Rheinboldt. A geometric framework for the numerical study of singular points. SIAM Journal on Numerical Analysis, 24(3):618–633, June 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1992:LEE

[FR92] Richard S. Falk and Gerard R. Richter. Local error estimates for a finite element method for hyperbolic and convection-diffusion equations. SIAM Journal on Numerical Analysis, 29(3): 730–754, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1999:EFE

[FR99] Richard S. Falk and Gerard R. Richter. Explicit finite element methods for symmetric hyperbolic equations. SIAM Journal on Numerical Analysis, 36(3): 935–952, June 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32946.

Franke:1971:OPA

[Fra71] Richard Franke. Orthogonal polynomials and approximate multiple integration. SIAM Journal on Numerical Analysis, 8(4):757–766, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Franke:1973:MPC

[Fra73] Richard Franke. Minimal point cubatures of precision seven for symmetric planar regions. SIAM Journal on Numerical Analysis, 10(5):849– 862, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Franke:1978:COA

[Fra78] Richard Franke. On the computation of optimal approximations in Sard corner spaces. SIAM Journal on Numerical Analysis, 15(4):791– 800, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Franklin:1987:CKA

[Fra87] Joel Franklin. Convergence in Karmarkar’s algorithm for linear programming. SIAM Journal on Numerical Analysis, 24(4):928–945, August 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

French:1987:FEM

[Fre87] Donald A. French. The finite element method for a degenerate elliptic equation. SIAM Journal on Numerical Analysis, 24(4):788–815, August 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

French:1990:CFE

[Fre90] Donald A. French. On the convergence of finite-element approximations of a relaxed variational problem. SIAM Journal on Numerical Analysis, 27(2):419–436, April 1990. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Fredebeul:1998:BGB

[Fre98]  Christoph Fredebeul.  A-

BDF: A generalization of the backward differentiation formulae. SIAM Journal on Numerical Analysis, 35(5): 1917–1938, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30621.

Fritsch:1971:SCN

[Fri71]  F. N. Fritsch.  On self-

contained numerical integration formulas for symmetric regions. SIAM Journal on Numerical Analysis, 8(2): 213–221, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Friedman:1981:MSN

[Fri81] Mark J. Friedman. Mathematical study of the nonlinear singular integral magnetic field equation. II. SIAM Journal on Numerical Analysis, 18(4):644–653, August 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Feldstein:1974:NMN

[FS74] Alan Feldstein and John R.

Sopka. Numerical methods for nonlinear Volterra integrodifferential equations. SIAM Journal on Numerical Analysis, 11(4):826–846, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ford:1987:AGR

[FS87] William F. Ford and Avram Sidi. An algorithm for a generalization of the Richardson extrapolation process. SIAM Journal on Numerical Analysis, 24(5):1212–1232, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Franca:1991:EAS

[FS91] Leopoldo P. Franca and Rolf Stenberg. Error analysis of some Galerkin least squares methods for the elasticity equations. SIAM Journal on Numerical Analysis, 28(6): 1680–1697, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fontecilla:1987:CTC

[FST87] Rodrigo Fontecilla, Trond Steihaug, and Richard A. Tapia. A convergence theory for a class of quasi-Newton methods for constrained optimization. SIAM Journal on Numerical Analysis, 24 (5):1133–1151, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Frank:1981:CC

[FSU81] Reinhard Frank, Josef Schneid, and Christoph W. Ueberhuber. The concept of Bconvergence. SIAM Journal on Numerical Analysis, 18(5): 753–780, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Frank:1985:ORI

[FSU85a] Reinhard Frank, Josef Schneid, and Christoph W. Ueberhuber. Order results for implicit Runge–Kutta methods applied to stiff systems. SIAM Journal on Numerical Analysis, 22(3):515–534, June 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Frank:1985:SPI

[FSU85b] Reinhard Frank, Josef Schneid, and Christoph W. Ueberhuber. Stability properties of implicit Runge–Kutta methods. SIAM Journal on Numerical Analysis, 22(3): 497–514, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ford:1974:UEE

[FT74] William H. Ford and T. W.

Ting. Uniform error estimates for difference approximations to nonlinear pseudoparabolic partial differential equations. SIAM Journal on Numerical Analysis, 11(1): 155–169, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

FroeseFischer:1969:PST

[FU69] Charlotte Froese Fischer and Riaz A. Usmani. Properties of some tridiagonal matrices and their application to boundary value problems. SIAM Journal on Numerical Analysis, 6(1):127–142, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Funaro:1987:PMC

[Fun87]  Daniele Funaro.  A pre-

conditioning matrix for the Chebyshev differencing operator. SIAM Journal on Numerical Analysis, 24(5):1024– 1031, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Funaro:1990:VFC

[Fun90] Daniele Funaro. A variational formulation for the Chebyshev pseudospectral approximation of Neumann problems. SIAM Journal on Numerical Analysis, 27(3): 695–703, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Funaro:1993:NSA

[Fun93]  Daniele Funaro.  A new

scheme for the approximation of advection-diffusion equations by collocation. SIAM Journal on Numerical Analysis, 30(6):1664–1676, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Falk:1995:CSO

[FX95] Richard S. Falk and Jian-Ming Xu. Convergence of a second-order scheme for the nonlinear dynamical equations of elastic rods. SIAM Journal on Numerical Analysis, 32(4):1185–1209, August 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gray:1967:NTR

[GA67] H. L. Gray and T. A. Atchison. Nonlinear transformations related to the evaluation of improper integrals. I. SIAM Journal on Numerical Analysis, 4(3):363–371, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gabutti:1985:AAC

[Gab85] Bruno Gabutti. An asymptotic approximation for a class of oscillatory infinite integrals. SIAM Journal on Numerical Analysis, 22(6):1191– 1199, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gaier:1966:CGG

[Gai66] Dieter Gaier. On the coefficients and the growth of gap power series. SIAM Journal on Numerical Analysis, 3 (2):248–265, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gaines:1974:DEA

[Gai74] Robert Gaines. Difference equations associated with boundary value problems for second order nonlinear ordinary differential equations. SIAM Journal on Numerical Analysis, 11(2): 411–434, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gray:1971:HOT

[GAM71] H. L. Gray, T. A. Atchison, and G. V. McWilliams. Higher order G-transformations. SIAM Journal on Numerical Analysis, 8(2):365–381, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garcia-Archilla:1998:PGM

[GANT98] Bosco García-Archilla, Julia Novo, and Edriss S. Titi. Postprocessing the Galerkin method: a novel approach to approximate inertial manifolds. SIAM Journal on Numerical Analysis, 35(3):941–972, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29609.

Gary:1966:GLR

[Gar66] John Gary. A generalization of the Lax–Richtmyer theorem on finite difference schemes. SIAM Journal on Numerical Analysis, 3 (3):467–473, September 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gargantini:1971:NSS

[Gar71] Irene Gargantini. The numerical stability of the Schur– Cohn criterion. SIAM Journal on Numerical Analysis, 8 (1):24–29, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garey:1975:SNS

[Gar75] L. Garey. Solving nonlinear second kind Volterra equations by modified increment methods. SIAM Journal on Numerical Analysis, 12(3): 501–508, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gargantini:1978:FAC

[Gar78] Irene Gargantini. Further applications of circular arithmetic: Schroeder-like algorithms with error bounds for finding zeros of polynomials. SIAM Journal on Numerical Analysis, 15(3):497–510, June 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gary:1981:MIA

[Gar81] John Gary. The multigrid iteration applied to the collocation method. SIAM Journal on Numerical Analysis, 18 (2):211–224, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gargantini:1983:ULQ

[Gar83] Irene Gargantini. The use of linear quadtrees in a numerical problem. SIAM Journal on Numerical Analysis, 20(6): 1161–1169, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gartland:1984:CPE

[Gar84] E. C. Gartland, Jr. Computable pointwise error bounds and the Ritz method in one dimension. SIAM Journal on Numerical Analysis, 21(1): 84–100, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gartland:1988:SSC

[Gar88] Eugene C. Gartland, Jr. Strong stability of compact discrete boundary value problems via exact discretizations. SIAM Journal on Numerical Analysis, 25(1):111– 123, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gartland:1993:UCS

[Gar93] Eugene C. Gartland, Jr. On the uniform convergence of the scharfetter-gummel discretization in one dimension. SIAM Journal on Numerical Analysis, 30(3):749–758, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gatica:1988:NPE

[Gat88] Gabriel N. Gatica. A note on the a posteriori error analysis for the finite element method. SIAM Journal on Numerical Analysis, 25(1):41–45, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gautschi:1967:NQP

[Gau67] Walter Gautschi. Numerical quadrature in the presence of a singularity. SIAM Journal on Numerical Analysis, 4(3):357–362, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gautschi:1970:ECC

[Gau70] Walter Gautschi. Efficient computation of the complex error function. SIAM Journal on Numerical Analysis, 7(1): 187–198, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Wei95].

Gay:1979:SCP

[Gay79] David M. Gay. Some convergence properties of Broyden’s method. SIAM Journal on Numerical Analysis, 16(4): 623–630, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gay:1981:PBN

[Gay81] David M. Gay. Perturbation bounds for nonlinear equations. SIAM Journal on Numerical Analysis, 18(4):654– 663, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gay:1982:SIL

[Gay82] David M. Gay. Solving interval linear equations. SIAM Journal on Numerical Analysis, 19(4):858–870, August 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gay:1983:CPB

[Gay83] David M. Gay. Computing perturbation bounds for nonlinear algebraic equations. SIAM Journal on Numerical Analysis, 20(3):638–651, June 1983. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Greenberg:1989:CAS

[GB89] L. Greenberg and I. Babuška.

A continuous analogue of Sturm sequences in the context of Sturm–Liouville equations. SIAM Journal on Numerical Analysis, 26(4):920– 945, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garnett:1971:HIM

[GBIY71] James M. Garnett, III, Adi Ben-Israel, and Stephen S. Yau. A hyperpower iterative method for computing matrix products involving the generalized inverse. SIAM Journal on Numerical Analysis, 8(1): 104–109, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Graham:1988:HOM

[GC88] I. G. Graham and G. A.

Chandler. High-order methods for linear functionals of solutions of second kind integral equations. SIAM Journal on Numerical Analysis, 25 (5):1118–1137, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:1995:SMF

[GC95] Ben Yu Guo and Wei Ming Cao. A spectral method for the fluid flow with low Mach number on the spherical surface. SIAM Journal on Numerical Analysis, 32(6):1764– 1777, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gomes:1996:CEW

[GC96] Sônia M. Gomes and Elsa Cortina. Convergence estimates for the wavelet Galerkin method. SIAM Journal on Numerical Analysis, 33(1):149–161, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:1998:ASM

[GC98] Benqi Guo and Weiming Cao. An additive Schwarz method for the h-b version of the finite element method in three dimensions. SIAM Journal on Numerical Analysis, 35(2):632–654, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29943.

Gimlin:1974:MEA

[GCB74] D. R. Gimlin, R. K. Cavin, III, and M. C. Budge, Jr. A multiple exchange algorithm for calculation of best restricted approximations. SIAM Journal on Numerical Analysis, 11(2): 219–231, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gear:1965:HMI

[Gea65] C. W. Gear. Hybrid methods for initial value problems in ordinary differential equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):69– 86, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Gear:1978:SNM

[Gea78] C. William Gear. The stability of numerical methods for second order ordinary differential equations. SIAM Journal on Numerical Analysis, 15 (1):188–197, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gear:1990:DAE

[Gea90] C. W. Gear. Differential algebraic equations, indices, and integral algebraic equations. SIAM Journal on Numerical Analysis, 27(6):1527– 1534, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Geddes:1978:NMP

[Ged78] K. O. Geddes. Near-minimax polynomial approximation in an elliptical region. SIAM Journal on Numerical Analysis, 15(6):1225–1233, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Geddes:1981:BSC

[Ged81] K. O. Geddes. Block structure in the Chebyshev–Padé table. SIAM Journal on Numerical Analysis, 18(5):844– 861, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gekeler:1975:CFD

[Gek75] E. Gekeler. A-convergence of finite difference approximations of parabolic initialboundary value problems. SIAM Journal on Numerical Analysis, 12(1):1–12, March 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gekeler:1976:LMM

[Gek76] E. Gekeler. Linear multistep methods and Galerkin procedures for initial boundary value problems. SIAM Journal on Numerical Analysis, 13(4):536–548, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Geltner:1981:GRQ

[Gel81]  Peter B. Geltner.  Gen-

eral Rayleigh quotient iteration. SIAM Journal on Numerical Analysis, 18(5):839– 843, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Genis:1984:FEM

[Gen84] Anatoly M. Genis. On finite element methods for the Euler–Poisson–Darboux equation. SIAM Journal on Numerical Analysis, 21(6): 1080–1106, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Genz:1986:FSI

[Gen86] Alan Genz. Fully symmetric interpolatory rules for multiple integrals. SIAM Journal on Numerical Analysis, 23(6): 1273–1283, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1973:NDR

[Geo73] Alan George. Nested dissection of a regular finite element mesh. SIAM Journal on Numerical Analysis, 10(2): 345–363, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1974:BES

[Geo74] Alan George. On block elimination for sparse linear systems. SIAM Journal on Numerical Analysis, 11(3): 585–603, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1976:NRS

[Geo76] Alan George. A negative result on sparse matrix splitting and Gaussian elimination. SIAM Journal on Numerical Analysis, 13(6):846– 853, December 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1977:NEU

[Geo77]  Alan George.  Numerical

experiments using dissection methods to solve n by n grid problems. SIAM Journal on Numerical Analysis, 14 (2):161–179, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1980:AOW

[Geo80]  Alan George.  An auto-

matic one-way dissection algorithm for irregular finite element problems. SIAM Journal on Numerical Analysis, 17 (6):740–751, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gerasoulis:1986:PPQ

[Ger86] Apostolos Gerasoulis. Piecewisepolynomial quadratures for Cauchy singular integrals. SIAM Journal on Numerical Analysis, 23(4):891– 902, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gerasoulis:1989:NIV

[Ger89] Apostolos Gerasoulis. Nyström’s iterative variant methods for the solution of Cauchy singular integral equations. SIAM Journal on Numerical Analysis, 26(2):430–441, April 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:1976:GDF

[GG76] David Gottlieb and Bertil Gustafsson. Generalized du fort-frankel methods for parabolic initial-boundary value problems. SIAM Journal on Numerical Analysis, 13(1):129–144, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guerra:1986:SIM

[GG86] Jaime Guerra and Bertil Gustafsson. A semi-implicit method for hyperbolic problems with different timescales. SIAM Journal on Numerical Analysis, 23(4): 734–749, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gigola:1990:RMS

[GG90] Cristina Gígola and Susana Gómez. A regularization method for solving the finite convex min-max problem. SIAM Journal on Numerical Analysis, 27(6):1621– 1634, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Govaerts:1997:DFM

[GGK97] W. Govaerts, J. Guckenheimer, and A. Khibnik. Defining functions for multiple Hopf bifurcations. SIAM Journal on Numerical Analysis, 34(3): 1269–1288, June 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27216.

Giladi:1998:IOI

[GGK98] Eldar Giladi, Gene H. Golub, and Joseph B. Keller. Inner and outer iterations for the Chebyshev algorithm. SIAM Journal on Numerical Analysis, 35(1):300–319, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29374.

Granas:1979:SMN

[GGL79] A. Granas, R. B. Guenther, and J. W. Lee. The shooting method for the numerical solution of a class of nonlinear boundary value problems. SIAM Journal on Numerical Analysis, 16(5):828– 836, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gasull:1999:ANM

[GGM99] Armengol Gasull, Antoni Guillamon, and Víctor Mañosa. An analytic-numerical method for computation of the Liapunov and period constants derived from their algebraic structure. SIAM Journal on Numerical Analysis, 36 (4):1030–1043, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29747.

Gottlieb:1996:SGM

[GGOS96] David Gottlieb, Bertil Gustafsson, Pelle Olsson, and Bo Strand. On the superconvergence of Galerkin methods for hyperbolic IBVP. SIAM Journal on Numerical Analysis, 33(5):1778– 1796, October 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25768.

Ganesh:1994:PTD

[GGS94] M. Ganesh, I. G. Graham, and J. Sivaloganathan. A pseudospectral three-dimensional boundary integral method applied to a nonlinear mode problem from finite elasticity. SIAM Journal on Numerical Analysis, 31(5):1378– 1414, October 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ganesh:1998:NSB

[GGS98] M. Ganesh, I. G. Graham, and J. Sivaloganathan. A new spectral boundary integral collocation method for three-dimensional potential problems. SIAM Journal on Numerical Analysis, 35(2):778–805, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30123.

Gottlieb:1982:NBT

[GGT82] David Gottlieb, Max Gunzburger, and Eli Turkel. On numerical boundary treatment of hyperbolic systems for finite difference and finite element methods. SIAM Journal on Numerical Analysis, 19(4):671–682, August 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1977:MMS

[GH77] J. H. George and H. G. Harris. Mathematical modeling of in situ oil shale retorting. SIAM Journal on Numerical Analysis, 14(1):137–151, March 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gustafson:1983:DFB

[GH83] Karl Gustafson and Robert Hartman. Divergence-free bases for finite element schemes in hydrodynamics. SIAM Journal on Numerical Analysis, 20(4):697– 721, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gunzburger:1992:TIE

[GH92] Max D. Gunzburger and Steven L. Hou. Treating inhomogeneous essential boundary conditions in finite element methods and the calculation of boundary stresses. SIAM Journal on Numerical Analysis, 29(2): 390–424, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gohberg:1994:FPC

[GHK94] I. Gohberg, M. Hanke, and I. Koltracht. Fast preconditioned conjugate gradient algorithms for Wiener–Hopf integral equations. SIAM Journal on Numerical Analysis, 31 (2):429–443, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:1996:VBE

[GHS96] Benqi Guo, Norbert Heuer, and Ernst P. Stephan. The h-p version of the boundary element method for transmission problems with piecewise analytic data. SIAM Journal on Numerical Analysis, 33 (2):789–808, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ghosh:1997:AES

[GHS97] Nilotpal Ghosh, William W. Hager, and Purandar Sarmah. The application of eigenpair stability to block diagonalization. SIAM Journal on Numerical Analysis, 34(3):1255–1268, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26125.

Gaitanos:1983:OAO

[GHY83] N. Gaïtanos, A. Hadjidimos, and A. Yeyios. Optimum accelerated overrelaxation (AOR) method for systems with positive definite coefficient matrix. SIAM Journal on Numerical Analysis, 20(4):774–783, August 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Girault:1974:TFD

[Gir74] V. Girault. Theory of a finite difference method on irregular networks. SIAM Journal on Numerical Analysis, 11 (2):260–282, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Girault:1978:MFE

[Gir78] V. Girault. A mixed finite element method for the stationary Stokes equations. SIAM Journal on Numerical Analysis, 15(3):534–555, June 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Girao:1995:CCA

[Gir95] Pedro Martins Girão. Convergence of a crystalline algorithm for the motion of a simple closed convex curve by weighted curvature. SIAM Journal on Numerical Analysis, 32(3):886–899, June 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Golse:1999:CNT

[GJL99] François Golse, Shi Jin, and C. David Levermore. The convergence of numerical transfer schemes in diffusive regimes I: Discreteordinate method. SIAM Journal on Numerical Analysis, 36 (5):1333–1369, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31598.

Golub:1965:CSV

[GK65] G. H. Golub and W. Kahan. Calculating the singular values and pseudo-inverse of a matrix. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):205– 224, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Gustafsson:1983:DAH

[GK83] Bertil Gustafsson and Heinz-Otto Kreiss. Difference approximations of hyperbolic problems with different time scales. I. the reduced problem. SIAM Journal on Numerical Analysis, 20(1):46–58, February 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garloff:1986:OIS

[GK86] J. Garloff and R. Krawczyk. Optimal inclusion of a solution set. SIAM Journal on Numerical Analysis, 23(1): 217–226, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ghaderpanah:1990:PS

[GK90] S. Ghaderpanah and S. Klasa. Polynomial scaling. SIAM Journal on Numerical Analysis, 27(1):117–135, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldman:1996:OOS

[GK96] Daniel Goldman and Tasso J.

Kaper. Nth-order operator splitting schemes and nonreversible systems. SIAM Journal on Numerical Analysis, 33 (1):349–367, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garbey:1997:HDD

[GK97] Marc Garbey and Hans G.

Kaper. Heterogeneous domain decomposition for singularly perturbed elliptic boundary value problems. SIAM Journal on Numerical Analysis, 34(4):1513– 1544, August 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28558.

Gustafson:1970:NCM

[GKR70] S.-A. Gustafson, K. O. Kortanek, and W. Rom. Non-Chebyshevian moment problems. SIAM Journal on Numerical Analysis, 7(3):335– 342, September 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1975:NFS

[GL75] Alan George and Wai-Hung Liu. A note on fill for sparse matrices. SIAM Journal on Numerical Analysis, 12 (3):452–455, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1978:AMP

[GL78a] Alan George and Joseph W. H. Liu. Algorithms for matrix partitioning and the numerical solution of finite element systems. SIAM Journal on Numerical Analysis, 15 (2):297–327, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1978:AND

[GL78b] Alan George and Joseph W. H. Liu. An automatic nested dissection algorithm for irregular finite element problems. SIAM Journal on Numerical Analysis, 15 (5):1053–1069, October 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garcia:1980:NSP

[GL80a] C. B. García and T. Y. Li. On the number of solutions to polynomial systems of equations. SIAM Journal on Numerical Analysis, 17(4):540– 546, August 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1980:MSI

[GL80b] Alan George and Joseph W. H. Liu. A minimal storage implementation of the minimum degree algorithm. SIAM Journal on Numerical Analysis, 17(2):282–299, April 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gerber:1981:GBM

[GL81] Richard R. Gerber and Franklin T. Luk. A generalized Broyden’s method for solving simultaneous linear equations. SIAM Journal on Numerical Analysis, 18(5): 882–890, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:1983:SCC

[GL83] David Gottlieb and Liviu Lustman. The spectrum of the Chebyshev collocation operator for the heat equation. SIAM Journal on Numerical Analysis, 20(5):909– 921, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goodman:1988:GAH

[GL88] Jonathan B. Goodman and Randall J. LeVeque. A geometric approach to high resolution TVD schemes. SIAM Journal on Numerical Analysis, 25(2):268–284, April 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gregory:1993:DVM

[GL93] John Gregory and Cantian Lin. Discrete variable methods for the m-dependent variable nonlinear, extremal problem in the calculus of variations II. SIAM Journal on Numerical Analysis, 30 (3):871–881, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Greenberg:1996:WBS

[GL96a] J. M. Greenberg and A. Y. Leroux. A well-balanced scheme for the numerical processing of source terms in hyperbolic equations. SIAM Journal on Numerical Analysis, 33(1):1–16, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:1996:FCS

[GL96b] Ben-Yu Guo and Jian Li. Fourier–Chebyshev spectral method for the two-dimensional Navier–Stokes equations. SIAM Journal on Numerical Analysis, 33(3):1169–1187, June 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gladwell:1969:SEP

[Gla69a] Ian Gladwell. Some extensions of a paper by Birkhoff and Priver. SIAM Journal on Numerical Analysis, 6(2): 294–298, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Glasser:1969:NST

[Gla69b] David Glasser. Numerical solution of two-point boundary value problems on total differential equations. SIAM Journal on Numerical Analysis, 6 (4):591–597, December 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Greenberg:1997:AAC

[GLBN97] J. M. Greenberg, A. Y. LeRoux, R. Baraille, and A. Noussair. Analysis and approximation of conservation laws with source terms. SIAM Journal on Numerical Analysis, 34 (5):1980–2007, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28675.

Grippo:1986:NLS

[GLL86] L. Grippo, F. Lampariello, and S. Lucidi. A nonmonotone line search technique for Newton’s method. SIAM Journal on Numerical Analysis, 23(4):707–716, August 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:1987:CSM

[GLT87a] David Gottlieb, Liviu Lustman, and Eitan Tadmor.

Convergence of spectral methods for hyperbolic initialboundary value systems. SIAM Journal on Numerical Analysis, 24(3):532–537, June 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:1987:SAS

[GLT87b] David Gottlieb, Liviu Lustman, and Eitan Tadmor. Stability analysis of spectral methods for hyperbolic initial-boundary value systems. SIAM Journal on Numerical Analysis, 24(2): 241–256, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ganguly:1991:SMV

[GLV91] Keshab Ganguly, J. Todd Lee, and H. D. Victory, Jr. On simulation methods for Vlasov– Poisson systems with particles initially asymptotically distributed. SIAM Journal on Numerical Analysis, 28(6): 1574–1609, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gourlay:1966:TLD

[GM66] A. R. Gourlay and A. R. Mitchell. Two-level difference schemes for hyperbolic systems. SIAM Journal on Numerical Analysis, 3(3):474– 485, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gourlay:1969:CSD

[GM69a] A. R. Gourlay and A. R.

Mitchell. A classification of split difference methods for hyperbolic equations in several space dimensions. SIAM Journal on Numerical Analysis, 6(1):62–71, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gourlay:1969:ECA

[GM69b] A. R. Gourlay and A. R.

Mitchell. The equivalence of certain alternating direction and locally one-dimensional difference methods. SIAM Journal on Numerical Analysis, 6(1):37–46, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Geddes:1975:PAP

[GM75] K. O. Geddes and J. C. Mason. Polynomial approximation by projections on the unit circle. SIAM Journal on Numerical Analysis, 12(1): 111–120, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

George:1978:AMD

[GM78a] Alan George and David R.

McIntyre. On the application of the minimum degree algorithm to finite element systems. SIAM Journal on Numerical Analysis, 15(1): 90–112, February 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gill:1978:ASN

[GM78b] Philip E. Gill and Walter Murray. Algorithms for the solution of the nonlinear least-squares problem. SIAM Journal on Numerical Analysis, 15(5):977–992, October 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Grundmann:1978:IIF

[GM78c] Axel Grundmann and H. M. Möller. Invariant integration formulas for the nsimplex by combinatorial methods. SIAM Journal on Numerical Analysis, 15(2): 282–290, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gourlay:1980:EFO

[GM80] A. R. Gourlay and J. Ll. Morris. The extrapolation of first order methods for parabolic partial differential equations. II. SIAM Journal on Numerical Analysis, 17(5):641– 655, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Genz:1983:IFF

[GM83] A. C. Genz and A. A. Malik. An imbedded family of fully symmetric numerical integration rules. SIAM Journal on Numerical Analysis, 20 (3):580–588, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldfarb:1989:SCV

[GM89] Donald Goldfarb and Sanjay Mehrotra. A self-correcting version of Karmarkar’s algorithm. SIAM Journal on Numerical Analysis, 26(4):1006– 1015, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:1991:PFE

[GM91] Ben-Yu Guo and He-Ping Ma. A pseudospectral finiteelement method for solving two-dimensional vorticity equations. SIAM Journal on Numerical Analysis, 28(1): 113–132, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:1993:CFE

[GM93] Ben Yu Guo and He Ping Ma. Combined finite element and pseudospectral method for the two-dimensional evolutionary Navier–Stokes equations. SIAM Journal on Numerical Analysis, 30(4):1066– 1083, September 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guillaume:1997:STH

[GM97] Philippe Guillaume and Mohamed Masmoudi. Solution to the time-harmonic Maxwell’s equations in a waveguide; use of higher-order derivatives for solving the discrete problem. SIAM Journal on Numerical Analysis, 34 (4):1306–1330, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27207.

Greenberg:1998:OTN

[GM98] Leon Greenberg and Marco Marletta. Oscillation theory and numerical solution of sixth order Sturm–Liouville problems. SIAM Journal on Numerical Analysis, 35 (5):2070–2098, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31645.

Goldstein:1993:PBC

[GMP93] C. I. Goldstein, Thomas A. Manteuffel, and Seymour V. Parter. Preconditioning and boundary conditions without H_{2} estimates: L_{2} condition numbers and the distribution of the singular values. SIAM Journal on Numerical Analysis, 30(2): 343–376, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guckenheimer:1997:CHB

[GMS97] John Guckenheimer, Mark Myers, and Bernd Sturmfels. Computing Hopf bifurcations. I. SIAM Journal on Numerical Analysis, 34(1):1–21, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25346.

George:1983:ERC

[GN83a] Alan George and Esmond Ng.

Erratum: “On row and column orderings for sparse least squares problems” [SIAM J. Numer. Anal. 20 (1983), no. 2, 326–344]. SIAM Journal on Numerical Analysis, 20 (4):872, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [GN83b].

George:1983:RCO

[GN83b] Alan George and Esmond Ng.

On row and column orderings for sparse least squares problems. SIAM Journal on Numerical Analysis, 20(2): 326–344, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [GN83a].

Gastaldi:1991:QOP

[GN91] Lucia Gastaldi and Ricardo H. Nochetto. Quasioptimal pointwise error estimates for the Reissner– Mindlin plate. SIAM Journal on Numerical Analysis, 28 (2):363–377, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griewank:1981:NMS

[GO81] Andreas Griewank and M. R. Osborne. Newton’s method for singular problems when the dimension of the null space is > 1. SIAM Journal on Numerical Analysis, 18(1): 145–149, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griewank:1983:ANM

[GO83] A. Griewank and M. R. Osborne. Analysis of Newton’s method at irregular singularities. SIAM Journal on Numerical Analysis, 20(4):747– 773, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldberg:1977:IIC

[Gol77] Michael A. Goldberg. Invariant imbedding and critical length problems. SIAM Journal on Numerical Analysis, 14(1):152–160, March 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldstein:1989:AAM

[Gol89]  C. I. Goldstein.  Analy-

sis and application of multigrid preconditioners for singularly perturbed boundary value problems. SIAM Journal on Numerical Analysis, 26 (5):1090–1123, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldstein:1991:PSP

[Gol91] C. I. Goldstein. Preconditioning singularity perturbed elliptic and parabolic problems. SIAM Journal on Numerical Analysis, 28(5):1386– 1418, October 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldstein:1993:MME

[Gol93] Charles I. Goldstein. Multigrid methods for elliptic problems in unbounded domains. SIAM Journal on Numerical Analysis, 30(1):159– 183, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldstein:1994:PNF

[Gol94] Charles I. Goldstein. Preconditioning nonconforming finite element methods. SIAM Journal on Numerical Analysis, 31(6):1623–1644, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goldberg:1998:SDSa

[Gol98a] Moshe Goldberg. Stable difference schemes for parabolic systems — A numerical radius approach. SIAM Journal on Numerical Analysis, 35(2):478–493, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31113.

Goldberg:1998:SDSb

[Gol98b] Moshe Goldberg. Stable difference schemes for parabolic systems — A numerical radius approach II. SIAM Journal on Numerical Analysis, 35 (5):1995–2003, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32825.

Gervasio:1997:SPD

[GOQ97] P. Gervasio, E. Ovtchinnikov, and A. Quarteroni. The spectral projection decomposition method for elliptic equations in two dimensions. SIAM Journal on Numerical Analysis, 34 (4):1616–1639, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26533.

Gordon:1971:BFM

[Gor71] William J. Gordon. Blendingfunction methods of bivariate and multivariate interpolation and approximation. SIAM Journal on Numerical Analysis, 8(1):158– 177, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gorelick:1975:SFM

[Gor75] Michael R. Gorelick. Stability of families of matrices of unbounded order. SIAM Journal on Numerical Analysis, 12 (2):188–202, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:1972:STD

[Got72] David Gottlieb. Strang-type difference schemes for multidimensional problems. SIAM Journal on Numerical Analysis, 9(4):650–661, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gould:1989:CSP

[Gou89] Nicholas Ian Mark Gould.

On the convergence of a sequential penalty function method for constrained minimization. SIAM Journal on Numerical Analysis, 26(1):

107–128, February 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Govaerts:1986:GPR

[Gov86] W. Govaerts. The geodesics of Pryce’s relative distance in R_{∞}^{2}. SIAM Journal on Numerical Analysis, 23(6):1295– 1302, December 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Govaerts:1993:CTB

[Gov93] W. Govaerts. Computation of Takens–Bogdanov type bifurcations with arbitrary codimension. SIAM Journal on Numerical Analysis, 30(4): 1121–1133, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Govaerts:1997:CSL

[Gov97] W. Govaerts. Computation of singularities in large nonlinear systems. SIAM Journal on Numerical Analysis, 34(3):867–880, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28218.

Golub:1973:DPI

[GP73] G. H. Golub and V. Pereyra. The differentiation of pseudoinverses and nonlinear least squares problems whose variables separate. SIAM Journal on Numerical Analysis, 10 (2):413–432, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gear:1984:OMS

[GP84] C. W. Gear and L. R. Petzold.

ODE methods for the solution of differential/algebraic systems. SIAM Journal on Numerical Analysis, 21(4):716– 728, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gonzalez:1998:STS

[GP98] C. González and C. Palencia. Stability of time-stepping methods for abstract timedependent parabolic problems. SIAM Journal on Numerical Analysis, 35(3): 973–989, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28341.

Gibbs:1976:ARB

[GPS76] Norman E. Gibbs, William G.

Poole, Jr., and Paul K. Stockmeyer. An algorithm for reducing the bandwidth and profile of a sparse matrix. SIAM Journal on Numerical Analysis, 13(2): 236–250, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gabetta:1997:RSN

[GPT97] E. Gabetta, L. Pareschi, and G. Toscani. Relaxation schemes for nonlinear kinetic equations. SIAM Journal on Numerical Analysis, 34(6): 2168–2194, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28776.

George:1978:IND

[GPV78] Alan George, William G. Poole, Jr., and Robert G. Voigt. Incomplete nested dissection for solving n by n grid problems. SIAM Journal on Numerical Analysis, 15(4): 662–673, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Girault:1982:AUS

[GR82] V. Girault and P.-A. Raviart. An analysis of upwind schemes for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 19(2): 312–333, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griewank:1984:CCG

[GR84] A. Griewank and G. W. Reddien. Characterization and computation of generalized turning points. SIAM Journal on Numerical Analysis, 21(1): 176–185, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griewank:1996:ASD

[GR96] A. Griewank and G. W.

Reddien. The approximate solution of defining equations for generalized turning points. SIAM Journal on Numerical Analysis, 33 (5):1912–1920, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25866.

Gragg:1965:EAO

[Gra65] William B. Gragg. On extrapolation algorithms for ordinary initial value problems. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):384–403, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Grau:1971:SNI

[Gra71] A. A. Grau. The simultaneous Newton improvement of a complete set of approximate factors of a polynomial. SIAM Journal on Numerical Analysis, 8(2):425–438, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Grandine:1987:CCS

[Gra87a] Thomas A. Grandine. The computational cost of simplex spline functions. SIAM Journal on Numerical Analysis, 24(4):887–890, August 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Grandine:1987:EIP

[Gra87b] Thomas A. Grandine. The evaluation of inner products of multivariate simplex splines. SIAM Journal on Numerical Analysis, 24(4): 882–886, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Greville:1964:NPI

[Gre64] T. N. E. Greville. Numerical procedures for interpolation by spline functions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??): 53–68, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Greville:1966:SLS

[Gre66]  T. N. E. Greville.  On

stability of linear smoothing formulas. SIAM Journal on Numerical Analysis, 3(1): 157–170, March 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Greenbaum:1984:AMM

[Gre84] Anne Greenbaum. Analysis of a multigrid method as an iterative technique for solving linear systems. SIAM Journal on Numerical Analysis, 21 (3):473–485, June 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Greengard:1991:SIT

[Gre91] L. Greengard. Spectral integration and two-point boundary value problems. SIAM Journal on Numerical Analysis, 28(4):1071–1080, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gremaud:1994:NAN

[Gre94] Pierre-Alain Gremaud. Numerical analysis of a nonconvex variational problem related to solid-solid phase transitions. SIAM Journal on Numerical Analysis, 31(1):111– 127, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Green:1999:CMM

[Gre99]  J. J. Green.  Calculat-

ing the maximum modulus of a polynomial using Steckin’s lemma. SIAM Journal on Numerical Analysis, 36 (4):1022–1029, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33133.

Griesel:1974:LRT

[Gri74]  Martha Ann Griesel.  A

linear remes-type algorithm for relative error approximation. SIAM Journal on Numerical Analysis, 11(1):170– 173, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Grigorieff:1986:CSC

[Gri86] Rolf Dieter Grigorieff. On the convergence of stability constants. SIAM Journal on Numerical Analysis, 23(4): 832–836, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griewank:1987:LCB

[Gri87] Andreas Griewank. The local convergence of Broydenlike methods on Lipschitzian problems in Hilbert spaces. SIAM Journal on Numerical Analysis, 24(3):684–705, June 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Grzegorski:1985:OPC

[Grz85] Stanisław M. Grzegórski. Orthogonal projections on convex sets for Newton-like methods. SIAM Journal on Numerical Analysis, 22(6): 1208–1219, December 1985.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gries:1967:SRF

[GS67] D. Gries and J. Stoer. Some results on fields of values of a matrix. SIAM Journal on Numerical Analysis, 4(2): 283–300, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gragg:1976:SVS

[GS76] W. B. Gragg and G. W. Stewart. A stable variant of the secant method for solving nonlinear equations. SIAM Journal on Numerical Analysis, 13 (6):889–903, December 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gambill:1988:LRW

[GS88] Thomas N. Gambill and Robert D. Skeel. Logarithmic reduction of the wrapping effect with application to ordinary differential equations. SIAM Journal on Numerical Analysis, 25(1):153– 162, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Glassey:1991:CPM

[GS91a] Robert Glassey and Jack Schaeffer. Convergence of a particle method for the relativistic Vlasov–Maxwell system. SIAM Journal on Numerical Analysis, 28(1): 1–25, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gustafsson:1991:NSE

[GS91b] Bertil Gustafsson and Hans Stoor. Navier–Stokes equations for almost incompressible flow. SIAM Journal on Numerical Analysis, 28(6): 1523–1547, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:1996:GPI

[GS96] David Gottlieb and Chi-Wang Shu. On the Gibbs phenomenon III: Recovering exponential accuracy in a subinterval from a spectral partial sum of a piecewise analytic function. SIAM Journal on Numerical Analysis, 33(1): 280–290, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griffiths:1992:NWP

[GSY92] D. F. Griffiths, A. M. Stuart, and H. C. Yee. Numerical wave propagation in an advection equation with a nonlinear source term. SIAM Journal on Numerical Analysis, 29(5):1244–1260, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gear:1974:EVM

[GT74a] C. W. Gear and K. W. Tu.

The effect of variable mesh size on the stability of multistep methods. SIAM Journal on Numerical Analysis, 11 (5):1025–1043, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gragg:1974:OEB

[GT74b] W. B. Gragg and R. A.

Tapia. Optimal error bounds for the Newton–Kantorovich theorem. SIAM Journal on Numerical Analysis, 11(1): 10–13, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gutknecht:1982:RPC

[GT82] Martin H. Gutknecht and Lloyd N. Trefethen. Real polynomial Chebyshev approximation by the Carathéodory– Fejér method. SIAM Journal on Numerical Analysis, 19 (2):358–371, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gladwell:1985:QAT

[GT85] I. Gladwell and R. M.

Thomas. A qualitative analysis of two- and three-step methods for stable second order systems. SIAM Journal on Numerical Analysis, 22(3): 535–560, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gautschi:1990:NCI

[GTV90] Walter Gautschi, E. Tychopoulos, and R. S. Varga. A note on the contour integral representation of the remainder term for a Gauss– Chebyshev quadrature rule. SIAM Journal on Numerical Analysis, 27(1):219–224, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Goodman:1988:SPI

[GU88] T. N. T. Goodman and K. Unsworth. Shapepreserving interpolation by parametrically defined curves. SIAM Journal on Numerical Analysis, 25(6):1453– 1465, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guermond:1992:NQL

[Gue92] Jean-Luc Guermond. Numerical quadratures for layer potentials over curved domains in R^{3}. SIAM Journal on Numerical Analysis, 29(5):1347– 1369, October 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guirguis:1987:CBI

[Gui87] Georges H. Guirguis. On the coupling boundary integral and finite element methods for the exterior Stokes problem in 3-D. SIAM Journal on Numerical Analysis, 24 (2):310–322, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gunn:1965:SED

[Gun65] James E. Gunn. The solution of elliptic difference equations by semi-explicit iterative techniques. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):24– 45, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Gunther:1974:TDI

[Gun74]  C. Gunther.  Third de-

gree integration formulas with four real points and positive weights in two dimensions. SIAM Journal on Numerical Analysis, 11(3):480–493, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gunttner:1980:ELC

[Gün80] R. Günttner. Evaluation of Lebesgue constants. SIAM Journal on Numerical Analysis, 17(4):512–520, August 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gunttner:1988:AUN

[Gün88] R. Günttner. On asymptotics for the uniform norms of the Lagrange interpolation polynomials corresponding to extended Chebyshev nodes. SIAM Journal on Numerical Analysis, 25(2): 461–469, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gupta:1975:DEE

[Gup75]  Murli M. Gupta.  Dis-

cretization error estimates for certain splitting procedures for solving first biharmonic boundary value problems. SIAM Journal on Numerical Analysis, 12(3):364–377, June 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gupta:1985:ABV

[Gup85a] Suchitra Gupta. An adaptive boundary value Runge–Kutta solver for first order boundary value problems. SIAM Journal on Numerical Analysis, 22(1):114–126, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See corrigenda [Gup85b].

Gupta:1985:CAB

[Gup85b]  Suchitra Gupta.  Corri-

genda: “An adaptive boundary value Runge–Kutta solver for first order boundary value problems” [SIAM J. Numer. Anal. 22 (1985), no. 1, 114– 126]. SIAM Journal on Numerical Analysis, 22(6): 1255, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Gup85a].

Gustafsson:1969:DAH

[Gus69] Bertil Gustafsson. On difference approximations to hyperbolic differential equations over long time intervals. SIAM Journal on Numerical Analysis, 6(3):508– 522, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gustafson:1970:CSC

[Gus70] S.-A. Gustafson. On the computational solution of a class of generalized moment problems. SIAM Journal on Numerical Analysis, 7(3):343– 357, September 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gustafson:1973:MCL

[Gus73]  Sven-Åke Gustafson.  A

method of computing limit values. SIAM Journal on Numerical Analysis, 10(6):1080– 1090, December 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gustafsson:1980:AEH

[Gus80] Bertil Gustafsson. Asymptotic expansions for hyperbolic problems with different time-scales. SIAM Journal on Numerical Analysis, 17(5): 623–634, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gusmann:1981:BGP

[Güs81a] Bernd Güsmann. Bounds of Galerkin projections on splines with highly nonuniform meshes. SIAM Journal on Numerical Analysis, 18(6): 1109–1119, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gustafsson:1981:CRD

[Gus81b] Bertil Gustafsson. The convergence rate for difference approximations to general mixed initial-boundary value problems. SIAM Journal on Numerical Analysis, 18(2): 179–190, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gutknecht:1983:CCT

[Gut83] Martin H. Gutknecht. On the computation of the conjugate trigonometric rational function and on a related splitting problem. SIAM Journal on Numerical Analysis, 20(6): 1198–1205, December 1983.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gutknecht:1993:CNC

[Gut93] Martin H. Gutknecht. Changing the norm in conjugate gradient type algorithms. SIAM Journal on Numerical Analysis, 30(1):40–56, February 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Golub:1974:CBS

[GV74] G. H. Golub and J. M. Varah.

On a characterization of the best l_{2}-scaling of a matrix. SIAM Journal on Numerical Analysis, 11(3):472–479, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Golub:1980:ATLb

[GV80] Gene H. Golub and Charles F.

Van Loan. An analysis of the total least squares problem. SIAM Journal on Numerical Analysis, 17(6):883– 893, December 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic). URL http://www.
jstor.org/stable/2156807.

Reprinted in [CGO07].

Gautschi:1983:EBG

[GV83] Walter Gautschi and Richard S.

Varga. Error bounds for Gaussian quadrature of analytic functions. SIAM Journal on Numerical Analysis, 20(6):

1170–1186, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ganguly:1989:CPM

[GV89] Keshab Ganguly and H. D. Victory, Jr. On the convergence of particle methods for multidimensional Vlasov– Poisson systems. SIAM Journal on Numerical Analysis, 26 (2):249–288, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gallouet:1991:FVS

[GV91] T. Gallouët and J.-P. Vila. Finite volume schemes for conservation laws of mixed type. SIAM Journal on Numerical Analysis, 28(6):1548– 1573, December 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gear:1974:SCV

[GW74a] C. W. Gear and D. S. Watanabe. Stability and convergence of variable order multistep methods. SIAM Journal on Numerical Analysis, 11 (5):1044–1058, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gordon:1974:PHI

[GW74b] William J. Gordon and James A. Wixom. Pseudoharmonic interpolation on convex domains. SIAM Journal on Numerical Analysis, 11(5):909–933, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gragg:1983:TCR

[GW83] W. B. Gragg and D. D.

Warner. Two constructive results in continued fractions. SIAM Journal on Numerical Analysis, 20(6):1187– 1197, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gregory:1990:DVM

[GW90] John Gregory and R. S.

Wang. Discrete variable methods for the m-dependent variable, nonlinear extremal problem in the calculus of variations. SIAM Journal on Numerical Analysis, 27(2): 470–487, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gray:1992:NMA

[GW92] H. L. Gray and Suojin Wang. A new method for approximating improper integrals. SIAM Journal on Numerical Analysis, 29(1):271– 283, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gatica:1997:CMF

[GW97] Gabriel N. Gatica and Wolfgang L. Wendland. Coupling of mixed finite elements and boundary elements for a hyperelastic interface problem. SIAM Journal on Numerical Analysis, 34(6): 2335–2356, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29131.

Greenwell-Yanik:1986:ASC

[GYF86] C. E. Greenwell-Yanik and G. Fairweather. Analyses of spline collocation methods for parabolic and hyperbolic problems in two space variables. SIAM Journal on Numerical Analysis, 23(2): 282–296, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gregory:1988:SMT

[GZ88] John Gregory and Marvin Zeman. Spline matrices and their applications to some higher order methods for boundary value problems. SIAM Journal on Numerical Analysis, 25(2): 399–410, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Haase:1996:ESR

[Haa96]  Mark C. Haase.  Extra

smoothness requirements for Galerkin methods for the wave equation. SIAM Journal on Numerical Analysis, 33 (5):1962–1968, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26781.

Haber:1977:TRN

[Hab77] Seymour Haber. The tanh rule for numerical integration. SIAM Journal on Numerical Analysis, 14(4):668– 685, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hackbusch:1979:CAE

[Hac79]  W. Hackbusch.  On the

computation of approximate eigenvalues and eigenfunctions of elliptic operators by means of a multi-grid method. SIAM Journal on Numerical Analysis, 16(2): 201–215, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hackbusch:1981:OEE

[Hac81]  W. Hackbusch.  Optimal

H^{p, p/2} error estimates for a parabolic Galerkin method. SIAM Journal on Numerical Analysis, 18(4):681– 692, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hager:1975:RTM

[Hag75] William W. Hager. The Ritz–Trefftz method for state and control constrained optimal control problems. SIAM Journal on Numerical Analysis, 12(6):854–867, December 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hager:1976:RCD

[Hag76] William W. Hager. Rates of convergence for discrete approximations to unconstrained control problems. SIAM Journal on Numerical Analysis, 13(4):449–472, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hager:1985:AMM

[Hag85] William W. Hager. Approximations to the multiplier method. SIAM Journal on Numerical Analysis, 22(1): 16–46, February 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hagstrom:1986:AEB

[Hag86] Thomas M. Hagstrom. Asymptotic expansions and boundary conditions for timedependent problems. SIAM Journal on Numerical Analysis, 23(5):948–958, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hager:1990:MMN

[Hag90] William W. Hager. Multiplier methods for nonlinear optimal control. SIAM Journal on Numerical Analysis, 27(4): 1061–1080, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hakopian:1982:MSF

[Hak82] Hakop Hakopian. Multivariate spline functions, B-spline basis and polynomial interpolations. SIAM Journal on Numerical Analysis, 19(3): 510–517, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hall:1973:NCB

[Hal73] C. A. Hall. Natural cubic and bicubic spline interpolation. SIAM Journal on Numerical Analysis, 10(6):1055– 1060, December 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hall:1974:SAP

[Hal74] George Hall. Stability analysis of predictor-corrector algorithms of Adams type. SIAM Journal on Numerical Analysis, 11(3):494–505, June 1974.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hald:1979:CVM

[Hal79] Ole H. Hald. Convergence of vortex methods for Euler’s equations. II. SIAM Journal on Numerical Analysis, 16(5): 726–755, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hald:1987:CVM

[Hal87a] Ole H. Hald. Convergence of vortex methods for Euler’s equations. III. SIAM Journal on Numerical Analysis, 24 (3):538–582, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Haley:1987:SMM

[Hal87b] Stephen B. Haley. Solution of modified matrix equations. SIAM Journal on Numerical Analysis, 24(4):946– 951, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hammer:1964:TA

[Ham64] Preston C. Hammer. Topologies of approximation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):69–75, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Hansen:1965:IAM

[Han65a] Eldon Hansen. Interval arithmetic in matrix computations, Part I. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 308–320, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Hanson:1965:EBC

[Han65b] M. A. Hanson. Error bounds in constrained best approximation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):473– 479, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Hanson:1971:NMS

[Han71] Richard J. Hanson. A numerical method for solving Fredholm integral equations of the first kind using singular values. SIAM Journal on Numerical Analysis, 8(3):616– 622, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hansen:1992:BSI

[Han92] E. R. Hansen. Bounding the solution of interval linear equations. SIAM Journal on Numerical Analysis, 29 (5):1493–1503, October 1992.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hanisch:1993:MPB

[Han93a]  M. R. Hanisch.  Multi-

grid preconditioning for the biharmonic Dirichlet problem. SIAM Journal on Numerical Analysis, 30(1):184– 214, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hanke:1993:FPS

[Han93b] Martin Hanke. An ϵ-free a posteriori stopping rule for certain iterative regularization methods. SIAM Journal on Numerical Analysis, 30(4): 1208–1228, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harabetian:1990:NMV

[Har90] Eduard Harabetian. A numerical method for viscous perturbations of hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 27(4):870–884, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harten:1996:MRD

[Har96] Ami Harten. Multiresolution representation of data: A general framework. SIAM Journal on Numerical Analysis, 33(3):1205–1256, June 1996.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hatcher:1982:EBC

[Hat82] Theodore R. Hatcher. An error bound for certain successive overrelaxation schemes. SIAM Journal on Numerical Analysis, 19(5):930–941, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Haussmann:1971:ID

[Hau71]  Werner Haussmann.  On

interpolation with derivatives. SIAM Journal on Numerical Analysis, 8(3):483– 485, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Haviv:1987:ADM

[Hav87]  Moshe Haviv.  Aggrega-

tion/disaggregation methods for computing the stationary distribution of a Markov chain. SIAM Journal on Numerical Analysis, 24(4):952– 966, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hayes:1981:GAD

[Hay81a] Linda J. Hayes. Galerkin alternating-direction methods for nonrectangular regions using patch approximations. SIAM Journal on Numerical Analysis, 18(4): 627–643, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hayes:1981:MBT

[Hay81b] Linda J. Hayes. A modified backward time discretization for nonlinear parabolic equations using patch approximations. SIAM Journal on Numerical Analysis, 18(5):781– 793, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hawkins:1986:NSC

[HB86] W. G. Hawkins and H. H. Barrett. A numerically stable circular harmonic reconstruction algorithm. SIAM Journal on Numerical Analysis, 23(4): 873–890, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Herceg:1986:ND

[HC86] Dragoslav Herceg and Ljiljana Cvetković. On a numerical differentiation. SIAM Journal on Numerical Analysis, 23 (3):686–691, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Han:1997:QNA

[HDS97] Weimin Han, B. Daya Reddy, and Gregory C. Schroeder. Qualitative and numerical analysis of quasi-static problems in elastoplasticity. SIAM Journal on Numerical Analysis, 34(1):143–177, February 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/26538.

Hebden:1971:BDB

[Heb71] M. D. Hebden. A bound on the difference between the Chebyshev norm and the holder norms of a function. SIAM Journal on Numerical Analysis, 8(2):270–277, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hedstrom:1968:RCS

[Hed68] G. W. Hedstrom. The rate of convergence of some difference schemes. SIAM Journal on Numerical Analysis, 5 (2):363–406, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hedstrom:1979:GMB

[Hed79]  G. W. Hedstrom.  The

Galerkin method based on Hermite cubics. SIAM Journal on Numerical Analysis, 16 (3):385–393, June 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hull:1972:CNM

[HEFS72] T. E. Hull, W. H. Enright, B. M. Fellen, and A. E. Sedgwick. Comparing numerical methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 9(4):603–637, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See errata [HEFS74].

Hull:1974:ECN

[HEFS74] T. E. Hull, W. H. Enright, B. M. Fellen, and A. E. Sedgwick. Errata: “Comparing Numerical Methods for Ordinary Differential Equations” [SIAM J. Numer. Anal. 9 (1972), no. 4, 603–637]. SIAM Journal on Numerical Analysis, 11(3):681, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [HEFS72].

Hegland:1992:OOR

[Heg92] Markus Hegland. An optimal order regularization method which does not use additional smoothness assumptions. SIAM Journal on Numerical Analysis, 29(5):1446– 1461, October 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heins:1966:PWZ

[Hei66] Maurice Heins. On the pseudoperiods of the Weierstrass zeta functions. SIAM Journal on Numerical Analysis, 3 (2):266–268, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heinrichs:1993:DDF

[Hei93a] Wilhelm Heinrichs. Domain decomposition for fourthorder problems. SIAM Journal on Numerical Analysis, 30 (2):435–453, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heinrichs:1993:STP

[Hei93b] Wilhelm Heinrichs. Splitting techniques for the pseudospectral approximation of the unsteady Stokes equations. SIAM Journal on Numerical Analysis, 30(1):19– 39, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heise:1994:AFD

[Hei94] Bodo Heise. Analysis of a fully discrete finite element method for a nonlinear magnetic field problem. SIAM Journal on Numerical Analysis, 31(3):745–759, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heinrich:1996:FFE

[Hei96] Bernd Heinrich. The Fourierfinite-element method for Poisson’s equation in axisymmetric domains with edges. SIAM Journal on Numerical Analysis, 33(5): 1885–1911, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26610.

Heinkenschloss:1998:TRM

[Hei98a] Matthias Heinkenschloss. A trust region method for norm constrained problems. SIAM Journal on Numerical Analysis, 35(4):1594– 1620, August 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27398.

Heinrichs:1998:STU

[Hei98b] Wilhelm Heinrichs. Splitting techniques for the unsteady Stokes equations. SIAM Journal on Numerical Analysis, 35 (4):1646–1662, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30878.

Heller:1974:DTA

[Hel74] Don Heller. A determinant theorem with applications to parallel algorithms. SIAM Journal on Numerical Analysis, 11(3):559–568, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heller:1976:SAC

[Hel76] Don Heller. Some aspects of the cyclic reduction algorithm for block tridiagonal linear systems. SIAM Journal on Numerical Analysis, 13 (4):484–496, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Henrici:1966:AAC

[Hen66] Peter Henrici. An algorithm for analytic continuation. SIAM Journal on Numerical Analysis, 3(1):67–78, March 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Henry:1969:BAS

[Hen69] Myron S. Henry. A best approximate solution of certain nonlinear differential equations. SIAM Journal on Numerical Analysis, 6(1):143– 148, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Henrici:1970:UBA

[Hen70a] Peter Henrici. Upper bounds for the abscissa of stability of a stable polynomial. SIAM Journal on Numerical Analysis, 7(4):538–544, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Henry:1970:SBA

[Hen70b] Myron S. Henry. A sequence of best approximate solutions of y′′ = F (x, y, y′). SIAM Journal on Numerical Analysis, 7(1):129–133, March 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Herring:1967:NGI

[Her67] Gary P. Herring. A note on generalized interpolation and the pseudoinverse. SIAM Journal on Numerical Analysis, 4(4):548–556, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hersh:1968:TDS

[Her68]  R. Hersh.  On the the-

ory of difference schemes for mixed initial boundary value problems. SIAM Journal on Numerical Analysis, 5(2): 436–450, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hertling:1975:NTA

[Her75] Jorg Hertling. Numerical treatment of algebraic integral equations by variational methods. SIAM Journal on Numerical Analysis, 12(2): 203–212, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Herr:1976:SWM

[Her76] David G. Herr. On Strand and Westwater’s minimum-RMS estimation of the numerical solution of a Fredholm integral equation of the first kind. SIAM Journal on Numerical Analysis, 13(3): 427–431, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Herron:1985:CCD

[Her85] Gary Herron. A characterization of certain C^{1} discrete triangular interpolants. SIAM Journal on Numerical Analysis, 22(4):811–819, August 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hesthaven:1998:EAO

[Hes98a] J. S. Hesthaven. From electrostatics to almost optimal nodal sets for polynomial interpolation in a simplex. SIAM Journal on Numerical Analysis, 35(2): 655–676, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30587.

Hesthaven:1998:IPP

[Hes98b] J. S. Hesthaven. Integration preconditioning of pseudospectral operators. I. basic linear operators. SIAM Journal on Numerical Analysis, 35 (4):1571–1593, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31918.

Henry:1978:NCA

[HHS78] J. N. Henry, M. S. Henry, and D. Schmidt. Numerical comparisons of algorithms for polynomial and rational multivariate approximations. SIAM Journal on Numerical Analysis, 15(6):1197– 1207, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See errata [de 78a].

Hickernell:1996:QEB

[Hic96] Fred J. Hickernell. Quadrature error bounds with applications to lattice rules. SIAM Journal on Numerical Analysis, 33(5):1995– 2016, October 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26143. See er-

ratum [Hic97].

Hickernell:1997:EQE

[Hic97] Fred J. Hickernell. Erratum: “Quadrature error bounds with applications to lattice rules” [SIAM J. Numer. Anal. 33 (1996), no. 5, 1995–2016; MR 1 411 860]. SIAM Journal on Numerical Analysis, 34(2):853–866, April 1997.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97400.  See

[Hic96].

Hiebert:1980:SSL

[Hie80] Kathie L. Hiebert. Solving systems of linear equations and inequalities. SIAM Journal on Numerical Analysis, 17 (3):447–464, June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Higdon:1984:BCS

[Hig84] Robert L. Higdon. Boundary conditions for suppressing rapidly moving components in hyperbolic systems. SIAM Journal on Numerical Analysis, 21(3):413–432, June 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Higham:1989:RDC

[Hig89a] Desmond J. Higham. Robust defect control with Runge– Kutta schemes. SIAM Journal on Numerical Analysis, 26 (5):1175–1183, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Higham:1989:AST

[Hig89b] Nicholas J. Higham. The accuracy of solutions to triangular systems. SIAM Journal on Numerical Analysis, 26 (5):1252–1265, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Higdon:1990:RBC

[Hig90] Robert L. Higdon. Radiation boundary conditions for elastic wave propagation. SIAM Journal on Numerical Analysis, 27(4):831–869, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Higdon:1994:RBC

[Hig94] Robert L. Higdon. Radiation boundary conditions for dispersive waves. SIAM Journal on Numerical Analysis, 31(1): 64–100, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hill:1968:NSD

[Hil68] C. Denson Hill. On the numerical solution of degenerate elliptic-parabolic equations. SIAM Journal on Numerical Analysis, 5(4):717– 724, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hilgers:1976:ERC

[Hil76] John W. Hilgers. On the equivalence of regularization and certain reproducing kernel Hilbert space approaches for solving first kind problems. SIAM Journal on Numerical Analysis, 13(2): 172–184, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Hil78].

Hilgers:1978:EER

[Hil78] J. W. Hilgers. Erratum: “On the Equivalence of Regularization and Certain Reproducing Kernel Hilbert Space Approaches for Solving First Kind Problems” [SIAM J. Numer. Anal. 13 (1976), no. 2, 172–184]. SIAM Journal on Numerical Analysis, 15(6): 1301, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Hil76].

Hilgers:1980:NEO

[Hil80] John W. Hilgers. A note on estimating the optimal regularization parameter. SIAM Journal on Numerical Analysis, 17(3):472–473, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hill:1997:GDS

[Hil97] A. T. Hill. Global dissipativity for A-stable methods. SIAM Journal on Numerical Analysis, 34(1):119– 142, February 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27097.

Hindmarsh:1972:OCR

[Hin72] A. C. Hindmarsh. Optimality in a class of rootfinding algorithms. SIAM Journal on Numerical Analysis, 9(2): 205–214, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hiptmair:1999:MMM

[Hip99]  R. Hiptmair.  Multigrid

method for Maxwell’s equations. SIAM Journal on Numerical Analysis, 36(1):204– 225, February 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32620.

Han:1999:OIS

[HJ99] Bin Han and Rong-Qing Jia. Optimal interpolatory subdivision schemes in multidimensional spaces. SIAM Journal on Numerical Analysis, 36 (1):105–124, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32561.

Hersh:1979:HAS

[HK79] Reuben Hersh and Tosio Kato. High-accuracy stable difference schemes for well-posed initial value problems. SIAM Journal on Numerical Analysis, 16(4):670– 682, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hicks:1991:EBN

[HK91] D. L. Hicks and K. L. Kuttler. Error bounds for numerical solutions to hydrodynamical problems involving shocks. SIAM Journal on Numerical Analysis, 28(3):662–684, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoppe:1994:AMM

[HK94] R. H. W. Hoppe and R. Kornhuber. Adaptive multilevel methods for obstacle problems. SIAM Journal on Numerical Analysis, 31(2): 301–323, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hangelbroek:1977:CMI

[HKL77] Rutger J. Hangelbroek, Hans G.

Kaper, and Gary K. Leaf. Collocation methods for integro-differential equations. SIAM Journal on Numerical Analysis, 14(3):377–390, June 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heinkenschloss:1992:FAN

[HKT92] M. Heinkenschloß, C. T. Kelley, and H. T. Tran. Fast algorithms for nonsmooth compact fixed-point problems. SIAM Journal on Numerical Analysis, 29(6):1769– 1792, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harten:1981:RCF

[HL81] Amiram Harten and Peter D. Lax. A random choice finite difference scheme for hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 18(2): 289–315, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hickey:1982:RMM

[HL82] Kevin R. Hickey and Glenn R. Luecke. Remarks on Marti’s method for solving first kind equations [SIAM J. Numer. Anal. 15 (1978), no. 6, 1071– 1076, by J. T. Marti]. SIAM Journal on Numerical Analysis, 19(3):623–628, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Mar78].

Hairer:1984:SVR

[HL84a] E. Hairer and Ch. Lubich. On the stability of Volterra–Runge–Kutta methods. SIAM Journal on Numerical Analysis, 21(1):123– 135, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harten:1984:CHR

[HL84b] Ami Harten and Peter D. Lax.

On a class of high resolution total-variation-stable finitedifference schemes. SIAM Journal on Numerical Analysis, 21(1):1–23, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). With an appendix by Peter D. Lax.

Harrell:1987:EGM

[HL87] E. M. Harrell and W. J. Layton. L^{2} estimates for Galerkin methods for semilinear elliptic equations. SIAM Journal on Numerical Analysis, 24(1): 52–58, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hochbruck:1997:KSA

[HL97] Marlis Hochbruck and Christian Lubich. On Krylov subspace approximations to the matrix exponential operator. SIAM Journal on Numerical Analysis, 34(5): 1911–1925, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28057.

Hayes:1998:NSK

[HL98] Brian T. Hayes and Philippe G.

LeFloch. Nonclassical shocks and kinetic relations: Finite difference schemes. SIAM Journal on Numerical Analysis, 35(6):2169–2194, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31599.

Hou:1991:CPV

[HLK91] Thomas Y. Hou, John Lowengrub, and Robert Krasny. Convergence of a point vortex method for vortex sheets. SIAM Journal on Numerical Analysis, 28(2):308–320, April 1991. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harten:1998:CEH

[HLLnWJM98] Ami Harten, Peter D. Lax, C. David Levermore, and newline William J. Morokoff. Convex entropies and hyperbolicity for general Euler equations. SIAM Journal on Numerical Analysis, 35(6): 2117–2127, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31670.

Hairer:1983:OCO

[HLN83] E. Hairer, Ch. Lubich, and S. P. Nørsett. Order of convergence of one-step methods for Volterra integral equations of the second kind. SIAM Journal on Numerical Analysis, 20 (3):569–579, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hageman:1980:ECI

[HLY80] Louis A. Hageman, Franklin T.

Luk, and David M. Young. On the equivalence of certain iterative acceleration methods. SIAM Journal on Numerical Analysis, 17(6):852– 873, December 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hamilton:1976:FES

[HM76a] J. D. Hamilton and H. D.

Meyer. Finite element solutions for steady state viscoplastic flow. SIAM Journal on Numerical Analysis, 13 (6):940–943, December 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Holyhead:1976:SCM

[HM76b] P. A. W. Holyhead and S. Mc-Kee. Stability and convergence of multistep methods for linear Volterra integral equations of the first kind. SIAM Journal on Numerical Analysis, 13(2): 269–292, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hald:1985:CLQ

[HM85] Jørgen Hald and Kaj Madsen. Combined LP and quasi-Newton methods for nonlinear L_{1} optimization. SIAM Journal on Numerical Analysis, 22(1):68–80, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hanke:1998:ACL

[HMM98] Michael Hanke, Ebroul Izquierdo Macana, and Roswitha März. On asymptotics in case of linear index-2 differentialalgebraic equations. SIAM Journal on Numerical Analysis, 35(4):1326–1346, August 1998. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/26887.

Hu:1995:ADS

[HMMR95] X. C. Hu, T. A. Manteuffel, S. McCormick, and T. F. Russell. Accurate discretization for singular perturbations: The one-dimensional case. SIAM Journal on Numerical Analysis, 32(1):83– 109, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoffman:1973:CBR

[HMR73] Alan J. Hoffman, Michael S. Martin, and Donald J. Rose.

Complexity bounds for regular finite difference and finite element grids. SIAM Journal on Numerical Analysis, 10 (2):364–369, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Holyhead:1975:MMS

[HMT75] P. A. W. Holyhead, S. Mc-Kee, and P. J. Taylor. Multistep methods for solving linear Volterra integral equations of the first kind. SIAM Journal on Numerical Analysis, 12(5):698–711, October 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hunt:1975:VIA

[HN75] Charles Hunt and Nabil R.

Nassif. On a variational inequality and its approximation, in the theory of semiconductors. SIAM Journal on Numerical Analysis, 12 (6):938–950, December 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harten:1987:UHO

[HO87] Ami Harten and Stanley Osher. Uniformly highorder accurate nonoscillatory schemes. I. SIAM Journal on Numerical Analysis, 24(2): 279–309, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hofer:1976:PIM

[Hof76] E. Hofer. A partially implicit method for large stiff systems of ODE’s with only few equations introducing small timeconstants. SIAM Journal on Numerical Analysis, 13(5): 645–663, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoff:1978:SCF

[Hof78] David Hoff. Stability and convergence of finite difference methods for systems of nonlinear reaction-diffusion equations. SIAM Journal on Numerical Analysis, 15(6):1161– 1177, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoff:1985:SCS

[Hof85] David Hoff. A scheme for computing solutions and interface curves for a doublydegenerate parabolic equation. SIAM Journal on Numerical Analysis, 22(4):687– 712, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hohage:1999:CRR

[Hoh99]  Thorsten Hohage.  Con-

vergence rates of a regularized Newton method in sound-hard inverse scattering. SIAM Journal on Numerical Analysis, 36(1):125– 142, February 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32775.

Holley:1978:EEB

[Hol78] Richard Holley. Exact error bounds for the phase velocity in an acoustic wave guide. SIAM Journal on Numerical Analysis, 15(4):715– 735, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hollig:1982:MS

[Höl82] Klaus Höllig. Multivariate splines. SIAM Journal on Numerical Analysis, 19(5):1013– 1031, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hong:1993:CMB

[Hon93] Bin Hong. On the computation of multiple bifurcations with multiple parameters and symmetry. SIAM Journal on Numerical Analysis, 30(4):1134–1154, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoppe:1982:DAC

[Hop82] Ronald H. W. Hoppe. Discrete approximations of cosine operator functions. I.

SIAM Journal on Numerical Analysis, 19(6):1110– 1128, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoppe:1987:MAV

[Hop87] Ronald H. W. Hoppe. Multigrid algorithms for variational inequalities. SIAM Journal on Numerical Analysis, 24 (5):1046–1065, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Horn:1983:FFO

[Hor83] M. K. Horn. Fourth- and fifthorder, scaled Runge–Kutta algorithms for treating dense output. SIAM Journal on Numerical Analysis, 20(3): 558–568, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hosea:1995:NRC

[Hos95] M. E. Hosea. A new recurrence for computing Runge– Kutta truncation error coefficients. SIAM Journal on Numerical Analysis, 32(6):1989– 2001, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Householder:1973:GFJ

[Hou73] A. S. Householder. George E. Forsythe (January 8, 1917– April 9, 1972). SIAM Journal on Numerical Analysis, 10: viii–xi, 1973. CODEN SJ-NAAM. ISSN 1095-7170. Collection of articles dedicated to the memory of George E. Forsythe.

Hou:1990:CVB

[Hou90]  Thomas Y. Hou.  Con-

vergence of a variable blob vortex method for the Euler and Navier–Stokes equations. SIAM Journal on Numerical Analysis, 27(6):1387– 1404, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hall:1968:CMEb

[HP68a] C. A. Hall and T. A.

Porsching. Computing the maximal eigenvalue and eigenvector of a nonnegative irreducible matrix. SIAM Journal on Numerical Analysis, 5(3):470–474, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hall:1968:CMEa

[HP68b] C. A. Hall and T. A.

Porsching. Computing the maximal eigenvalue and eigenvector of a positive matrix. SIAM Journal on Numerical Analysis, 5(2): 269–274, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hageman:1975:ANB

[HP75] L. A. Hageman and T. A. Porsching. Aspects of nonlinear block successive overrelaxation. SIAM Journal on Numerical Analysis, 12(3): 316–335, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Haegemans:1977:CCF

[HP77] Ann Haegemans and Robert Piessens. Construction of cubature formulas of degree seven and nine symmetric planar regions, using orthogonal polynomials. SIAM Journal on Numerical Analysis, 14(3): 492–508, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hanosn:1978:NST

[HP78a] Richard L. Hanson and James L. Phillips. Numerical solution of two-dimensional integral equations using linear elements. SIAM Journal on Numerical Analysis, 15(1): 113–121, February 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoffmann:1978:NPG

[HP78b] W. Hoffmann and B. N. Parlett. A new proof of global convergence for the tridiagonal QL algorithm. SIAM Journal on Numerical Analysis, 15(5):929–937, October 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hall:1980:PAF

[HP80] C. A. Hall and T. A.

Porsching. Padé approximants, fractional step methods and Navier–Stokes discretizations. SIAM Journal on Numerical Analysis, 17 (6):840–851, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heywood:1982:FEA

[HR82] John G. Heywood and Rolf Rannacher. Finite element approximation of the nonstationary Navier–Stokes problem. I. regularity of solutions and second-order error estimates for spatial discretization. SIAM Journal on Numerical Analysis, 19(2): 275–311, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heywood:1986:FEA

[HR86] John G. Heywood and Rolf Rannacher. Finite element approximation of the nonstationary Navier–Stokes problem, Part II: Stability of solutions and error estimates uniform in time. SIAM Journal on Numerical Analysis, 23(4): 750–777, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heywood:1988:FEA

[HR88] John G. Heywood and Rolf Rannacher. Finite element approximation of the nonstationary Navier–Stokes problem, Part III. smoothing property and higher order error estimates for spatial discretization. SIAM Journal on Numerical Analysis, 25(3): 489–512, June 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heywood:1990:FEA

[HR90] John G. Heywood and Rolf Rannacher. Finite-element approximation of the nonstationary Navier–Stokes problem. Part IV: Error analysis for second-order time discretization. SIAM Journal on Numerical Analysis, 27(2): 353–384, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heywood:1993:QTM

[HR93] John G. Heywood and Rolf Rannacher. On the question of turbulence modeling by approximate inertial manifolds and the nonlinear Galerkin method. SIAM Journal on Numerical Analysis, 30(6): 1603–1621, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Han:1995:FEM

[HR95] Weimin Han and B. Daya Reddy. On the finite element method for mixed variational inequalities arising in elastoplasticity. SIAM Journal on Numerical Analysis, 32(6): 1778–1807, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huang:1997:AMM

[HR97] Weizhang Huang and Robert D.

Russell. Analysis of moving mesh partial differential equations with spatial smoothing. SIAM Journal on Numerical Analysis, 34 (3):1106–1126, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25644.

Hoch:1999:NSP

[HR99] Philippe Hoch and Michel Rascle. A numerical study of a pathological example of p-system. SIAM Journal on Numerical Analysis, 36 (5):1588–1603, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32318.

Huang:1994:MMP

[HRR94] Weizhang Huang, Yuhe Ren, and Robert D. Russell. Moving mesh partial differential equations (MMPDEs) based on the equidistribution principle. SIAM Journal on Numerical Analysis, 31(3): 709–730, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hansen:1967:IAM

[HS67] Eldon Hansen and Roberta Smith. Interval arithmetic in matrix computations, Part II. SIAM Journal on Numerical Analysis, 4(1):1–9, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hall:1968:NBS

[HS68] C. A. Hall and J. Spanier. Nested bounds for the spectral radius. SIAM Journal on Numerical Analysis, 5(1): 113–125, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hundsdorfer:1987:AEI

[HS87] W. H. Hundsdorfer and M. N. Spijker. On the algebraic equations in implicit Runge– Kutta methods. SIAM Journal on Numerical Analysis, 24 (3):583–594, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huang:1992:PMT

[HS92] Wei Zhang Huang and David M. Sloan. The pseudospectral method for thirdorder differential equations. SIAM Journal on Numerical Analysis, 29(6):1626– 1647, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Humphries:1994:RKM

[HS94] A. R. Humphries and A. M.

Stuart. Runge–Kutta methods for dissipative and gradient dynamical systems. SIAM Journal on Numerical Analysis, 31(5):1452–1485, October 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Han:1997:FMS

[HS97] Lu Han and Larry L. Schumaker. Fitting monotone surfaces to scattered data using C^{1} piecewise cubics. SIAM Journal on Numerical Analysis, 34(2): 569–585, April 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26858.

Hyman:1999:ODT

[HS99] James M. Hyman and Mikhail Shashkov. The orthogonal decomposition theorems for mimetic finite difference methods. SIAM Journal on Numerical Analysis, 36(3):788–818, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31404.

Huang:1982:BEM

[HT82] Mingyou Huang and Vidar Thomee. On the backward Euler method for parabolic equations with rough initial data. SIAM Journal on Numerical Analysis, 19(3): 599–603, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Teng:1997:ORC

[hTZ97] Zhen huan Teng and Pingwen Zhang. Optimal L^{1}-rate of convergence for the viscosity method and monotone scheme to piecewise constant solutions with shocks. SIAM Journal on Numerical Analysis, 34(3):959–978, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26886.

Hu:1996:SDP

[Hu96] Qiya Hu. Stieltjes derivatives and β-polynomial spline collocation for Volterra integrodifferential equations with singularities. SIAM Journal on Numerical Analysis, 33(1):

208–220, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hu:1997:SNS

[Hu97]  Qi Ya Hu.  Supercon-

vergence of numerical solutions to Volterra integral equations with singularities. SIAM Journal on Numerical Analysis, 34(5): 1698–1707, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26683.

Huang:1975:JTM

[Hua75] C. P. Huang. A Jacobi-type method for triangularizing an arbitrary matrix. SIAM Journal on Numerical Analysis, 12 (4):566–570, September 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hubbard:1965:ADS

[Hub65] Bert E. Hubbard. Alternating direction schemes for the heat equation in a general domain. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):448– 463, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Huddleston:1968:MPM

[Hud68] Robert E. Huddleston. Maximum principle methods for the discretization error in the approximation of the Bergman harmonic kernel by finite difference methods. SIAM Journal on Numerical Analysis, 5(3):613– 625, September 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huelsman:1973:QFF

[Hue73] Charles B. Huelsman, III. Quadrature formulas over fully symmetric planar regions. SIAM Journal on Numerical Analysis, 10(3): 539–552, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hughett:1998:EBN

[Hug98] Paul Hughett. Error bounds for numerical inversion of a probability characteristic function. SIAM Journal on Numerical Analysis, 35 (4):1368–1392, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31085.

Hurt:1967:SST

[Hur67] James Hurt. Some stability theorems for ordinary difference equations. SIAM Journal on Numerical Analysis, 4 (4):582–596, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huynh:1993:AMC

[Huy93]  Hung T. Huynh.  Accu-

rate monotone cubic interpolation. SIAM Journal on Numerical Analysis, 30(1):57– 100, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huynh:1995:AUM

[Huy95] H. T. Huynh. Accurate upwind methods for the Euler equations. SIAM Journal on Numerical Analysis, 32 (5):1565–1619, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoffman:1970:PDG

[HV70] Alan J. Hoffman and Richard S.

Varga. Patterns of dependence in generalizations of Gerschgorin’s theorem. SIAM Journal on Numerical Analysis, 7(4):571–574, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Havlak:1996:NAR

[HV96] Karl J. Havlak and Harold Dean Victory, Jr. The numerical analysis of random particle methods applied to Vlasov– Poisson–Fokker–Planck kinetic equations. SIAM Journal on Numerical Analysis, 33 (1):291–317, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Holst:1997:SMS

[HV97] Michael Holst and Stefan Vandewalle. Schwarz methods: To symmetrize or not to symmetrize. SIAM Journal on Numerical Analysis, 34(2):699–722, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27574.

Havlak:1998:DPM

[HV98] Karl J. Havlak and Harold Dean Victory, Jr. On deterministic particle methods for solving Vlasov–Poisson–Fokker– Planck systems. SIAM Journal on Numerical Analysis, 35 (4):1473–1519, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30252.

Houstis:1988:CCS

[HVR88] E. N. Houstis, E. A. Vavalis, and J. R. Rice. Convergence of O(h^{4}) cubic spline collocation methods for elliptic partial differential equations. SIAM Journal on Numerical Analysis, 25(1):54–74, February 1988. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Hairer:1981:ASI

[HW81] E. Hairer and G. Wanner. Algebraically stable and implementable Runge–Kutta methods of high order. SIAM Journal on Numerical Analysis, 18(6):1098–1108, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hairer:1983:IBF

[HW83] E. Hairer and G. Wanner. On the instability of the BDF formulas. SIAM Journal on Numerical Analysis, 20(6):1206– 1209, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hou:1992:CFD

[HW92] Thomas Y. Hou and Brian T. R. Wetton. Convergence of a finite difference scheme for the Navier–Stokes equations using vorticity boundary conditions. SIAM Journal on Numerical Analysis, 29 (3):615–639, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hou:1993:SOC

[HW93] Thomas Y. Hou and Brian T. R. Wetton. Second-order convergence of a projection scheme for the incompressible Navier–Stokes equations with boundaries. SIAM Journal on Numerical Analysis, 30(3): 609–629, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hairer:1997:OCG

[HW97a] Ernst Hairer and Gerhard Wanner. Order conditions for general two-step Runge– Kutta methods. SIAM Journal on Numerical Analysis, 34(6):2087–2089, December 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29814.

Hoppe:1997:AMT

[HW97b] Ronald H. W. Hoppe and Barbara Wohlmuth. Adaptive multilevel techniques for mixed finite element discretizations of elliptic boundary value problems. SIAM Journal on Numerical Analysis, 34(4):1658–1681, August 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27699.

Han:1998:NMF

[HW98] Houde Han and Xiaonan Wu. A new mixed finite element formulation and the MAC method for the Stokes equations. SIAM Journal on Numerical Analysis, 35(2):560–571, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30038.

Hu:1995:DMS

[HY95] Yingkang Hu and Xiang Ming Yu. Discrete modulus of smoothness of splines with equally spaced knots. SIAM Journal on Numerical Analysis, 32(5):1428–1435, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hu:1996:DCA

[HY96] Yingkang Hu and Xiang Ming Yu. The degree of copositive approximation and a computer algorithm. SIAM Journal on Numerical Analysis, 33 (1):388–398, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Harten:1994:FMA

[HYS94] Ami Harten and Itai Yad-Shalom. Fast multiresolution algorithms for matrix-vector multiplication. SIAM Journal on Numerical Analysis, 31(4): 1191–1218, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hettich:1990:AGR

[HZ90] R. Hettich and P. Zencke. An algorithm for general restricted rational Chebyshev approximation. SIAM Journal on Numerical Analysis, 27 (4):1024–1033, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hsiao:1994:OOM

[HZ94] George C. Hsiao and Shangyou Zhang. Optimal order multigrid methods for solving exterior boundary value problems. SIAM Journal on Numerical Analysis, 31(3):680–694, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ihlenburg:1997:FES

[IB97] Frank Ihlenburg and Ivo Babuška. Finite element solution of the Helmholtz equation with high wave number Part II: The h-p version of the FEM. SIAM Journal on Numerical Analysis, 34 (1):315–358, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27233.

Isaacson:1982:NAS

[IIMPL82] D. Isaacson, E. L. Isaacson, D. Marchesin, and P. J. Paes-Leme. Numerical analysis of the spectral properties of coupled oscillator Schrödinger operators. II. two-coupled anharmonic oscillators. SIAM Journal on Numerical Analysis, 19(1):126–141, February 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ito:1993:DPM

[IK93] Kazufumi Ito and Sungkwon Kang. A dissipative pseudospectral method for the two-dimensional Navier– Stokes equations. SIAM Journal on Numerical Analysis, 30 (5):1333–1350, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ito:1991:FDA

[IKP91] K. Ito, F. Kappel, and G. Peichl. A fully discretized approximation scheme for sizestructured population models. SIAM Journal on Numerical Analysis, 28(4):923– 954, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Iannelli:1996:NAM

[ILM^{+}96] Mimmo Iannelli, Roberto Loro, Fabio A. Milner, Andrea Pugliese, and Guglielmo Rabbiolo. Numerical analysis of a model for the spread of HIV/AIDS. SIAM Journal on Numerical Analysis, 33 (3):864–882, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ikebe:1969:NSI

[ILT69] Yasuhiko Ikebe, M. Stuart Lynn, and William P. Timlake. The numerical solution of the integral equation formulation of the single interface Neumann problem. SIAM Journal on Numerical Analysis, 6(3):334– 346, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

intHout:1996:NUM

[in 96] K. J. in ’t Hout. A note on unconditional maximum norm contractivity of diagonally split Runge–Kutta methods. SIAM Journal on Numerical Analysis, 33(3): 1125–1134, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ioakimidis:1981:WGM

[Ioa81] N. I. Ioakimidis. On the weighted Galerkin method of numerical solution of Cauchy type singular integral equations. SIAM Journal on Numerical Analysis, 18(6):1120– 1127, December 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ip:1987:LCS

[Ip87] Chi Ming Ip. On least-change secant updates in factored form. SIAM Journal on Numerical Analysis, 24(5):1126– 1132, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Iserles:1991:UAS

[IPS91] A. Iserles, A. T. Peplow, and A. M. Stuart. A unified approach to spurious solutions introduced by time discretisation. Part I: Basic theory. SIAM Journal on Numerical Analysis, 28(6):1723– 1751, December 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Incerti:1979:NMS

[IPZ79] Sandro Incerti, Valerio Parisi, and Francesco Zirilli. A new method for solving nonlinear simultaneous equations. SIAM Journal on Numerical Analysis, 16(5):779– 789, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ivanauskas:1999:CSE

[IR99] F. Ivanauskas and M. Radziunas. On convergence and stability of the explicit difference method for solution of nonlinear Schrödinger equations. SIAM Journal on Numerical Analysis, 36(5): 1466–1481, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31972.

Irvine:1990:ESN

[IS90] Douglas H. Irvine and Michael A. Savageau. Efficient solution of nonlinear ordinary differential equations expressed in S-System canonical form. SIAM Journal on Numerical Analysis, 27(3):704–735, June 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Isaacson:1986:D

[Isa86] Eugene Isaacson. Dedication. SIAM Journal on Numerical Analysis, 23(5):i, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Iserles:1979:GPA

[Ise79] Arieh Iserles. On the generalized Padé approximations to the exponential function. SIAM Journal on Numerical Analysis, 16(4):631– 636, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Iserles:1981:MEA

[Ise81a] Arieh Iserles. On multivalued exponential approximations. SIAM Journal on Numerical Analysis, 18(3):480–499, June 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Iserles:1981:RIA

[Ise81b] Arieh Iserles. Rational interpolation to exp(−x) with application to certain stiff systems. SIAM Journal on Numerical Analysis, 18(1):1– 12, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

Iserles:1984:CMN

[Ise84]  Arieh Iserles.  Composite

methods for numerical solution of stiff systems of ODEs. SIAM Journal on Numerical Analysis, 21(2): 340–351, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Issautier:1996:CWP

[Iss96]  D. Issautier.  Conver-

gence of a weighted particle method for solving the Boltzmann (B.G.K.) equation. SIAM Journal on Numerical Analysis, 33(6): 2099–2119, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26685.

Ioakimidis:1980:CBD

[IT80] N. I. Ioakimidis and P. S. Theocaris. A comparison between the direct and the classical numerical methods for the solution of Cauchy type singular integral equations. SIAM Journal on Numerical Analysis, 17(1):115– 118, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

IP:1988:OCC

[IT88] Chi Ming Ip and Michael J. Todd. Optimal conditioning and convergence in rank one quasi-Newton updates. SIAM Journal on Numerical Analysis, 25(1):206–221, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ito:1991:NMC

[IT91] Kazufumi Ito and Janos Turi. Numerical methods for a class of singular integro-differential equations based on semigroup approximation. SIAM Journal on Numerical Analysis, 28(6):1698–1722, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Istace:1995:TFZ

[IT95] M.-P. Istace and J.-P. Thiran. On the third and fourth Zolotarev problems in the complex plane. SIAM Journal on Numerical Analysis, 32(1): 249–259, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ito:1991:FDS

[ITM91] K. Ito, H. T. Tran, and A. Manitius. A fully-discrete spectral method for delaydifferential equations. SIAM Journal on Numerical Analysis, 28(4):1121–1140, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Isaacson:1981:MWR

[IZ81] Eugene Isaacson and Gideon Zwas. Mountain winds revisited. SIAM Journal on Numerical Analysis, 18(2): 277–288, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1983:QNM

[JA83] George W. Johnson and Nieves H. Austria. A quasi-Newton method employing direct secant updates of matrix factorizations. SIAM Journal on Numerical Analysis, 20(2):315–325, April 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackson:1975:IAE

[Jac75] L. W. Jackson. Interval arithmetic error-bounding algorithms. SIAM Journal on Numerical Analysis, 12(2): 223–238, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackiewicz:1981:NSV

[Jac81] Zdzisław Jackiewicz. The numerical solution of Volterra functional differential equations of neutral type. SIAM Journal on Numerical Analysis, 18(4):615–626, August 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackiewicz:1984:OSM

[Jac84] Zdzisław Jackiewicz. Onestep methods of any order for neutral functional differential equations. SIAM Journal on Numerical Analysis, 21 (3):486–511, June 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackiewicz:1986:QMM

[Jac86] Zdzisław Jackiewicz. Quasilinear multistep methods and variable step predictorcorrector methods for neutral functional differential equations. SIAM Journal on Numerical Analysis, 23(2): 423–452, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackson:1988:CIA

[Jac88] Kenneth R. Jackson. The convergence of integrandapproximation formulas for the numerical solution of IVPs for ODEs. SIAM Journal on Numerical Analysis, 25 (1):163–188, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jaffard:1992:WMF

[Jaf92] Stéphane Jaffard. Wavelet methods for fast resolution of elliptic problems. SIAM Journal on Numerical Analysis, 29(4):965–986, August 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

James:1968:NSS

[Jam68] Ralph L. James. The numerical solution of singular Volterra integral equations. SIAM Journal on Numerical Analysis, 5(2):352–362, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jamet:1969:NSD

[Jam69] Pierre Jamet. Numerical solution of the Dirichlet problem for elliptic parabolic equations. SIAM Journal on Numerical Analysis, 6(3):458– 466, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jamet:1971:FDS

[Jam71] Pierre Jamet. A finite difference scheme and an existence theorem for a nonlinear hyperbolic system of differential equations. SIAM Journal on Numerical Analysis, 8 (3):524–535, September 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

James:1973:CMI

[Jam73] K. R. James. Convergence of matrix iterations subject to diagonal dominance. SIAM Journal on Numerical Analysis, 10(3):478–484, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jamet:1977:EIE

[Jam77] Pierre Jamet. Estimation of the interpolation error for quadrilateral finite elements which can degenerate into triangles. SIAM Journal on Numerical Analysis, 14(5):925– 930, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jamet:1978:GTA

[Jam78]  Pierre Jamet.  Galerkin-

type approximations which are discontinuous in time for parabolic equations in a variable domain. SIAM Journal on Numerical Analysis, 15(5): 912–928, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jamet:1980:SCG

[Jam80] Pierre Jamet. Stability and convergence of a generalized Crank–Nicolson scheme on a variable mesh for the heat equation. SIAM Journal on Numerical Analysis, 17(4): 530–539, August 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jankowska:1979:TMS

[Jan79] Janina Jankowska. Theory of multivariate secant methods. SIAM Journal on Numerical Analysis, 16(4):547– 562, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janssen:1983:GBS

[Jan83] Rainer Janssen. A generalization of Buchanan’s stability criterion. SIAM Journal on Numerical Analysis, 20(2): 234–238, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janovska:1994:NTS

[Jan94]  Dáša Janovská.  Numeri-

cal treatment of subspacebreaking Takens–Bogdanov points with nonlinear degeneracies. SIAM Journal on Numerical Analysis, 31(5):1415– 1433, October 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jay:1996:SPR

[Jay96] Laurent Jay. Symplectic partitioned Runge–Kutta methods for constrained Hamiltonian systems. SIAM Journal on Numerical Analysis, 33(1): 368–387, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jensen:1988:DRN

[JB88] Søren Jensen and Ivo Babuška. Dimensional reduction for nonlinear boundary value problems. SIAM Journal on Numerical Analysis, 25(3): 644–669, June 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jepson:1986:CCN

[JD86] A. D. Jepson and D. W. Decker. Convergence cones near bifurcation. SIAM Journal on Numerical Analysis, 23(5):959–975, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Joe:1993:IRL

[JD93] S. Joe and S. A. R. Disney. Intermediate rank lattice rules for multidimensional integration. SIAM Journal on Numerical Analysis, 30(2):

569–582, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jefferson:1969:TEE

[Jef69] Thomas H. Jefferson. Truncation error estimates for T -fractions. SIAM Journal on Numerical Analysis, 6 (3):359–364, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackson:1978:TCC

[JEH78] K. R. Jackson, W. H. Enright, and T. E. Hull. A theoretical criterion for comparing Runge–Kutta formulas. SIAM Journal on Numerical Analysis, 15(3):618–641, June 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jeltsch:1976:SSR

[Jel76] Rolf Jeltsch. Stiff stability and its relation to A_{0}-and A(0)-stability. SIAM Journal on Numerical Analysis, 13 (1):8–17, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jeltsch:1977:SSM

[Jel77]  R. Jeltsch.  Stiff stabil-

ity of multistep multiderivative methods. SIAM Journal on Numerical Analysis, 14 (4):760–772, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See corrigendum [Jel79].

Jeltsch:1979:CSS

[Jel79] Rolf Jeltsch. Corrigendum: “Stiff Stability of Multistep Multiderivative Methods” [SIAM J. Numer. Anal. 14 (1977), no. 4, 760–772, MR 57 #11058]. SIAM Journal on Numerical Analysis, 16 (2):339–345, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Jel77].

Jennings:1970:DFR

[Jen70] Gray Jennings. Digital filtering of random digits. SIAM Journal on Numerical Analysis, 7(2):238–247, June 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jensen:1972:SLS

[Jen72] Paul S. Jensen. The solution of large symmetric eigenproblems by sectioning. SIAM Journal on Numerical Analysis, 9(4):534–545, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jensen:1976:SSM

[Jen76] Paul S. Jensen. Stiffly stable methods for undamped second order equations of motion. SIAM Journal on Numerical Analysis, 13(4):549– 563, September 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jensen:1992:ADR

[Jen92] Søren Jensen. Adaptive dimensional reduction numerical solution of monotone quasilinear boundary value problems. SIAM Journal on Numerical Analysis, 29 (5):1294–1320, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jeon:1994:IBI

[Jeo94] Youngmok Jeon. An indirect boundary integral equation method for the biharmonic equation. SIAM Journal on Numerical Analysis, 31 (2):461–476, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1973:MPLa

[Jer73a] Joseph W. Jerome. Minimization problems and linear and nonlinear spline functions. I: Existence. SIAM Journal on Numerical Analysis, 10(5): 808–819, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1973:MPLb

[Jer73b] Joseph W. Jerome. Minimization problems and linear and nonlinear spline functions. II: Convergence. SIAM Journal on Numerical Analysis, 10(5): 820–830, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1980:CSI

[Jer80] Joseph W. Jerome. Convergence of successive iterative semidiscretizations for FitzHugh–Nagumo reaction diffusion systems. SIAM Journal on Numerical Analysis, 17(2):192–206, April 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1984:FDS

[Jer84] Joseph W. Jerome. Fully discrete stability and invariant rectangular regions for reaction-diffusion systems. SIAM Journal on Numerical Analysis, 21(6):1054– 1065, December 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1988:CDN

[Jer88] Joseph W. Jerome. Convectiondominated nonlinear systems: analysis of the douglas-russell transport-diffusion algorithm based on approximate characteristics and invariant regions. SIAM Journal on Numerical Analysis, 25(4): 815–836, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jespersen:1978:RGM

[Jes78]  Dennis Jespersen.  Ritz–

Galerkin methods for singular boundary value problems. SIAM Journal on Numerical Analysis, 15(4):813– 834, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jetter:1982:NCG

[Jet82] Kurt Jetter. A new class of Gaussian quadrature formulas based on Birkhoff type data. SIAM Journal on Numerical Analysis, 19(5):1081– 1089, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jetter:1987:UGB

[Jet87]  Kurt Jetter.  Uniqueness

of Gauss–Birkhoff quadrature formulas. SIAM Journal on Numerical Analysis, 24(1): 147–154, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jia:1986:SIB

[Jia86] Rong Qing Jia. Spline interpolation at a bi-infinite knot sequence. SIAM Journal on Numerical Analysis, 23 (3):653–662, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jimack:1996:BAP

[Jim96]  P. K. Jimack.  A best

approximation property of the moving finite element method. SIAM Journal on Numerical Analysis, 33(6): 2286–2302, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25543.

Jones:1968:ICC

[JK68] R. C. Jones and L. A. Karlovitz. Iterative construction of constrained Chebyshev approximation of continuous functions. SIAM Journal on Numerical Analysis, 5 (3):574–585, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackson:1974:FOE

[JK74] L. W. Jackson and Surender Kumar Kenue. A fourth order exponentially fitted method. SIAM Journal on Numerical Analysis, 11(5): 965–978, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1991:FEA

[JK91] Joseph W. Jerome and Thomas Kerkhoven. A finite element approximation theory for the drift diffusion semiconductor model. SIAM Journal on Numerical Analysis, 28(2):403–422, April 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jaimoukha:1994:KSM

[JK94] Imad M. Jaimoukha and Ebrahim M. Kasenally. Krylov subspace methods for solving large Lyapunov equations. SIAM Journal on Numerical Analysis, 31(1):227– 251, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackiewicz:1996:CWR

[JK96] Z. Jackiewicz and M. Kwapisz.

Convergence of waveform relaxation methods for differentialalgebraic systems. SIAM Journal on Numerical Analysis, 33(6):2303–2317, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/23309.

Jiang:1998:HRN

[JLL^{+}98] G.-S. Jiang, D. Levy, C.-T.

Lin, S. Osher, and E. Tadmor. High-resolution nonoscillatory central schemes with nonstaggered grids for hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 35(6):2147– 2168, December 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31756.

Johnson:1983:PPC

[JMP83] Olin G. Johnson, Charles A. Micchelli, and George Paul. Polynomial preconditioners for conjugate gradient calculations. SIAM Journal on Numerical Analysis, 20(2): 362–376, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jeltsch:1983:SSH

[JN83] Rolf Jeltsch and Olavi Nevanlinna. Stability of semidiscretizations of hyperbolic problems. SIAM Journal on Numerical Analysis, 20(6): 1210–1218, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackson:1995:PPR

[JN95] K. R. Jackson and S. P. Nørsett. The potential for parallelism in Runge–Kutta methods. part 1: RK formulas in standard form. SIAM Journal on Numerical Analysis, 32(1):49–82, February 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jiang:1998:FEM

[JN98] Xun Jiang and Ricardo H. Nochetto. A P^{1}-P^{1} finite element method for a phase relaxation model I: Quasiuniform mesh. SIAM Journal on Numerical Analysis, 35(3):1176–1190, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29778.

Johnson:1990:PEE

[JNT90] Claes Johnson, Yi-Yong Nie, and Vidar Thomée. An A posteriori error estimate and adaptive timestep control for a backward Euler discretization of a parabolic problem. SIAM Journal on Numerical Analysis, 27(2): 277–291, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jiang:1999:FEM

[JNV99] Xun Jiang, Ricardo H. Nochetto, and Claudio Verdi. A P^{1}–P^{1} finite element method for a phase relaxation model II: Adaptively refined meshes. SIAM Journal on Numerical Analysis, 36(3):974–999, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31759.

Joe:1985:DCM

[Joe85] S. Joe. Discrete collocation methods for second kind Fredholm integral equations. SIAM Journal on Numerical Analysis, 22(6):1167– 1177, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Joe:1999:ADN

[Joe99] Stephen Joe. An average L_{2} discrepancy for numbertheoretic rules. SIAM Journal on Numerical Analysis, 36(6): 1949–1961, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33760.

Johnson:1969:EBS

[Joh69]  Olin G. Johnson.  Error

bounds for Sturm–Liouville eigenvalue approximations by several piecewise cubic Rayleigh–Ritz methods. SIAM Journal on Numerical Analysis, 6(3):317–333, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1975:NRR

[Joh75]  Darell J. Johnson.  On

the nontriviality of restricted range polynomial approximation. SIAM Journal on Numerical Analysis, 12(2): 183–187, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1976:CEA

[Joh76] Claes Johnson. A convergence estimate for an approximation of a parabolic variational inequality. SIAM Journal on Numerical Analysis, 13 (4):599–606, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1977:MFE

[Joh77] C. Johnson. A mixed finite element method for plasticity problems with hardening. SIAM Journal on Numerical Analysis, 14(4):575– 583, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1978:NDF

[Joh78] Charles R. Johnson. Numerical determination of the field of values of a general complex matrix. SIAM Journal on Numerical Analysis, 15(3): 595–602, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1988:EEA

[Joh88] Claes Johnson. Error estimates and adaptive time-step control for a class of onestep methods for stiff ordinary differential equations. SIAM Journal on Numerical Analysis, 25(4):908–926, August 1988. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Jokinen:1996:MSI

[Jok96] Olli Jokinen. On monotone solutions of an integrodifferential equation in linear viscoelasticity. SIAM Journal on Numerical Analysis, 33 (4):1410–1424, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25126.

Jamet:1967:NME

[JP67] Pierre Jamet and Seymour V. Parter. Numerical methods for elliptic differential equations whose coefficients are singular on a portion of the boundary. SIAM Journal on Numerical Analysis, 4(2): 131–146, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1983:CFD

[JP83] Claes Johnson and Juhani Pitkäranta. Convergence of a fully discrete scheme for twodimensional neutron transport. SIAM Journal on Numerical Analysis, 20(5):951– 966, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janovsky:1992:CAA

[JP92] Vladimír Janovský and Petr Plecháč. Computer-aided analysis of imperfect bifurcation diagrams. I. simple bifurcation point and isola formation centre. SIAM Journal on Numerical Analysis, 29 (2):498–512, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janovsky:1996:LNA

[JP96] Vladimír Janovský and Petr Plecháč. Local numerical analysis of Hopf bifurcation. SIAM Journal on Numerical Analysis, 33(3): 1150–1168, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jin:1998:USA

[JP98] HongSung Jin and Steven Pruess. Uniformly superconvergent approximations for linear two-point boundary value problems. SIAM Journal on Numerical Analysis, 35 (1):363–375, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29720.

Joly:1996:NNF

[JPRT96] P. Joly, C. Poirier, J. E.

Roberts, and P. Trouve. A new nonconforming finite element method for the computation of electromagnetic guided waves I: Mathematical analysis. SIAM Journal on Numerical Analysis, 33 (4):1494–1525, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25681.

Jin:1998:DRS

[JPT98] Shi Jin, Lorenzo Pareschi, and Giuseppe Toscani. Diffusive relaxation schemes for multiscale discrete-velocity kinetic equations. SIAM Journal on Numerical Analysis, 35(6):2405–2439, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31596.

James:1975:CCS

[JR75] K. R. James and W. Riha. Convergence criteria for successive overrelaxation. SIAM Journal on Numerical Analysis, 12(2):137–143, April 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1995:NHS

[JRB95] Claes Johnson, Rolf Rannacher, and Mats Boman. Numerics and hydrodynamic stability: Toward error control in computational fluid dynamics. SIAM Journal on Numerical Analysis, 32(4):1058– 1079, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackiewicz:1991:TSR

[JRF91] Z. Jackiewicz, R. Renaut, and A. Feldstein. Twostep Runge–Kutta methods. SIAM Journal on Numerical Analysis, 28(4):1165– 1182, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jerome:1968:NON

[JS68a] J. W. Jerome and L. L. Schumaker. A note on obtaining natural spline functions by the abstract approach of atteia and laurent. SIAM Journal on Numerical Analysis, 5(4):657–663, December 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1968:SM

[JS68b] L. W. Johnson and D. R.

Scholz. On Steffensen’s method. SIAM Journal on Numerical Analysis, 5(2): 296–302, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jones:1969:TEB

[JS69] William B. Jones and R. I.

Snell. Truncation error bounds for continued fractions. SIAM Journal on Numerical Analysis, 6(2): 210–221, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jain:1983:KAS

[JS83] Narendra K. Jain and Kishore Singhal. On Kublanovskaya’s approach to the solution of the generalized latent value problem for functional λmatrices. SIAM Journal on Numerical Analysis, 20 (5):1062–1070, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jepson:1985:FST

[JS85a] A. Jepson and A. Spence. Folds in solutions of two parameter systems and their calculation. Part I. SIAM Journal on Numerical Analysis, 22 (2):347–368, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jepson:1985:NSN

[JS85b] Allan D. Jepson and Alastair Spence. The numerical solution of nonlinear equations having several parameters. I. scalar equations. SIAM Journal on Numerical Analysis, 22(4):736–759, August 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jeltsch:1987:ABT

[JS87] Rolf Jeltsch and J. H. Smit.

Accuracy barriers of two time level difference schemes for hyperbolic equations. SIAM Journal on Numerical Analysis, 24(1):1–11, February 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Johnson:1989:AQE

[JS89] Claes G. L. Johnson and L. Ridgway Scott. An analysis of quadrature errors in second-kind boundary integral methods. SIAM Journal on Numerical Analysis, 26(6): 1356–1382, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Joe:1992:ILR

[JS92] Stephen Joe and Ian H.

Sloan. Imbedded lattice rules for multidimensional integration. SIAM Journal on Numerical Analysis, 29(4):1119– 1135, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janovsky:1996:QAN

[JS96] Vladimír Janovský and Viktor Seige. Qualitative analysis of Newton’s flow. SIAM Journal on Numerical Analysis, 33 (5):2068–2097, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24185.

Jianguo:1998:FVE

[JS98] Huang Jianguo and Xi Shitong. On the finite volume element method for general self-adjoint elliptic problems. SIAM Journal on Numerical Analysis, 35(5): 1762–1774, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26469.

Jepson:1991:NSN

[JSC91] A. D. Jepson, A. Spence, and K. A. Cliffe. The numerical solution of nonlinear equations having several parameters. Part III. equations with Z_{2}-symmetry. SIAM Journal on Numerical Analysis, 28 (3):809–832, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jenkins:1970:TSA

[JT70] M. A. Jenkins and J. F. Traub. A three-stage algorithm for real polynomials using quadratic iteration. SIAM Journal on Numerical Analysis, 7(4):545–566, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jones:1971:PBT

[JT71] William B. Jones and W. J.

Thron. A Posteriori bounds for the truncation error of continued fractions. SIAM Journal on Numerical Analysis, 8(4):693–705, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jackiewicz:1995:GCT

[JT95] Z. Jackiewicz and S. Tracogna.

A general class of two-step Runge–Kutta methods for ordinary differential equations. SIAM Journal on Numerical Analysis, 32(5):1390– 1427, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Joly:1997:NTA

[JT97] Patrick Joly and Jukka Tuomela. A new theoretical approach to absorbing layers. SIAM Journal on Numerical Analysis, 34(2):671–698, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26174.

Jones:1983:TEB

[JTW83] William B. Jones, W. J.

Thron, and Haakon Waadeland. Truncation error bounds for continued fractions K(a_{n}/1) with parabolic element regions. SIAM Journal on Numerical Analysis, 20(6):1219–1230, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jupp:1978:ADS

[Jup78] David L. B. Jupp. Approximation to data by splines with free knots. SIAM Journal on Numerical Analysis, 15 (2):328–343, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janssen:1996:MWR

[JV96] Jan Janssen and Stefan Vandewalle. Multigrid waveform relaxation on spatial finite element meshes: The continuous-time case. SIAM Journal on Numerical Analysis, 33(2):456–474, April 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Janssen:1997:SWR

[JV97] Jan Janssen and Stefan Vandewalle. On SOR waveform relaxation methods. SIAM Journal on Numerical Analysis, 34(6):2456– 2481, December 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29429.

Jimack:1991:TDF

[JW91] P. K. Jimack and A. J.

Wathen. Temporal derivatives in the finite-element method on continuously deforming grids. SIAM Journal on Numerical Analysis, 28(4): 990–1003, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jin:1998:NPS

[JX98] Shi Jin and Zhouping Xin.

Numerical passage from systems of conservation laws to Hamilton–Jacobi equations, and relaxation schemes. SIAM Journal on Numerical Analysis, 35(6):2385– 2404, December 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31436.

Jiang:1998:DSF

[JY98] Guang-Shan Jiang and Shih-Hsien Yu. Discrete shocks for finite difference approximations to scalar conservation laws. SIAM Journal on Numerical Analysis, 35(2):749–772, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30709.

Kacur:1999:SSF

[Kac99]  J. Kacur.  Solution of

some free boundary problems by relaxation schemes. SIAM Journal on Numerical Analysis, 36(1):290–316, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/31114.

Karakashian:1993:OOE

[KAD93] Ohannes Karakashian, Georgios D. Akrivis, and Vassilios A. Dougalis. On optimal order error estimates for the nonlinear Schrödinger equation. SIAM Journal on Numerical Analysis, 30(2): 377–400, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kahan:1967:LMC

[Kah67]  W. Kahan.  Laguerre’s

method and a circle which contains at least one zero of a polynomial. SIAM Journal on Numerical Analysis, 4 (3):474–482, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kahaner:1969:EAE

[Kah69]  David K. Kahaner.  On

equal and almost equal weight quadrature formulas. SIAM Journal on Numerical Analysis, 6(4):551–556, December 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammler:1969:EBC

[Kam69] David W. Kammler. An error bound for capacitance calculations. SIAM Journal on Numerical Analysis, 6(2): 254–257, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammler:1976:CAC

[Kam76] David W. Kammler. Chebyshev approximation of completely monotonic functions by sums of exponentials. SIAM Journal on Numerical Analysis, 13(5):761–774, October 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammler:1978:NEW

[Kam78] David W. Kammler. Numerical evaluation of exp(tA) when A is a companion matrix. SIAM Journal on Numerical Analysis, 15(6):1077– 1102, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammler:1979:ACM

[Kam79a] David W. Kammler. L_{1}- approximation of completely monotonic functions by sums of exponentials. SIAM Journal on Numerical Analysis, 16(1):30–45, February 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammler:1979:LSA

[Kam79b] David W. Kammler. Least squares approximation of completely monotonic functions by sums of exponentials. SIAM Journal on Numerical Analysis, 16(5):801– 818, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammler:1981:DAC

[Kam81] David W. Kammler. Differential approximation of completely monotonic functions. SIAM Journal on Numerical Analysis, 18(5):900– 918, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karon:1978:CIC

[Kar78] John M. Karon. Computing improved Chebyshev approximations by the continuation method. I. description of an algorithm. SIAM Journal on Numerical Analysis, 15(6): 1269–1288, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karakashian:1982:GLM

[Kar82] Ohannes A. Karakashian. On a Galerkin–Lagrange multiplier method for the stationary Navier–Stokes equations. SIAM Journal on Numerical Analysis, 19(5):909– 923, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karni:1992:VSP

[Kar92] S. Karni. Viscous shock profiles and primitive formulations. SIAM Journal on Numerical Analysis, 29(6):1592– 1609, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karni:1996:FFF

[Kar96] Smadar Karni. Far-field filtering operators for suppression of reflections from artificial boundaries. SIAM Journal on Numerical Analysis, 33(3): 1014–1047, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Katsanis:1969:NST

[Kat69] Theodore Katsanis. Numerical solution of Tricomi equation using theory of symmetric positive differential equations. SIAM Journal on Numerical Analysis, 6(2): 236–253, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaufman:1974:ASG

[Kau74] Linda Kaufman. The LZalgorithm to solve the generalized eigenvalue problem. SIAM Journal on Numerical Analysis, 11(5):997– 1024, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaufman:1986:BDC

[Kau86] Edwin H. Kaufman, Jr. The behavior of differential correction in difficult situations. SIAM Journal on Numerical Analysis, 23(2): 453–460, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Katz:1974:ACN

[KC74] I. Norman Katz and Leon Cooper. An always-convergent numerical scheme for a random locational equilibrium problem. SIAM Journal on Numerical Analysis, 11(4): 683–692, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kwon:1993:SOA

[KC93] Yonghoon Kwon and Chung-Ki Cho. Second-order accurate difference methods for a one-sex model of population dynamics. SIAM Journal on Numerical Analysis, 30 (5):1385–1399, October 1993.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaniel:1979:MNM

[KD79] Shmuel Kaniel and Achiya Dax. A modified Newton’s method for unconstrained minimization. SIAM Journal on Numerical Analysis, 16 (2):324–331, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keast:1983:FSI

[KD83] Patrick Keast and Julio C.

Díaz. Fully symmetric integration formulas for the surface of the sphere in S dimensions. SIAM Journal on Numerical Analysis, 20(2): 406–419, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keast:1973:OPM

[Kea73] P. Keast. Optimal parameters for multidimensional integration. SIAM Journal on Numerical Analysis, 10(5):831– 838, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kearfott:1990:PIG

[Kea90] R. Baker Kearfott. Preconditioners for the interval Gauss– Seidel method. SIAM Journal on Numerical Analysis, 27 (3):804–822, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kedem:1981:PEB

[Ked81] Gershon Kedem. A posteriori error bounds for twopoint boundary value problems. SIAM Journal on Numerical Analysis, 18(3): 431–448, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keeping:1970:BMA

[Kee70] A. J. Keeping. Band matrices arising from finite difference approximations to a third order partial differential equation. SIAM Journal on Numerical Analysis, 7(1): 142–156, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keeping:1971:CAT

[Kee71] A. J. Keeping. On certain almost-triangular band matrices. SIAM Journal on Numerical Analysis, 8(2): 288–303, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keener:1986:SSS

[Kee86]  James P. Keener.  The

sarkovskiĭ sequence and stable periodic orbits of maps of the interval. SIAM Journal on Numerical Analysis, 23(5): 976–985, October 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). Dedicated to Professor Herbert B. Keller on the occasion of his 60th birthday.

Keeling:1990:GRK

[Kee90a] Stephen L. Keeling. Galerkin/ Runge–Kutta discretizations for semilinear parabolic equations. SIAM Journal on Numerical Analysis, 27(2): 394–418, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keeling:1990:LCN

[Kee90b] Stephen L. Keeling. On Lipschitz continuity of nonlinear differential operators. SIAM Journal on Numerical Analysis, 27(2):385–393, April 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keenan:1995:TSP

[Kee95] Philip T. Keenan. Thermal simulation of pipeline flow. SIAM Journal on Numerical Analysis, 32(4):1225– 1262, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:1965:SSS

[Kel65] Herbert B. Keller. On the solution of singular and semidefinite linear systems by iteration. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):281– 290, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Kellogg:1966:EEE

[Kel66]  R. B. Kellogg.  An er-

ror estimate for elliptic difference equations on a convex polygon. SIAM Journal on Numerical Analysis, 3 (1):79–90, March 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:1969:ADM

[Kel69] Herbert B. Keller. Accurate difference methods for linear ordinary differential systems subject to linear constraints. SIAM Journal on Numerical Analysis, 6(1):8–30, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:1974:ADM

[Kel74] Herbert B. Keller. Accurate difference methods for nonlinear two-point boundary value problems. SIAM Journal on Numerical Analysis, 11 (2):305–320, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:1981:GIN

[Kel81] Herbert B. Keller. Geometrically isolated nonisolated solutions and their approximation. SIAM Journal on Numerical Analysis, 18(5):822– 838, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kelley:1995:FMA

[Kel95] C. T. Kelley. A fast multilevel algorithm for integral equations. SIAM Journal on Numerical Analysis, 32(2): 501–513, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kemper:1975:SFA

[Kem75] Gene A. Kemper. Spline function approximation for solutions of functional differential equations. SIAM Journal on Numerical Analysis, 12 (1):73–88, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kershaw:1971:NCI

[Ker71] D. Kershaw. A note on the convergence of interpolatory cubic splines. SIAM Journal on Numerical Analysis, 8 (1):67–74, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kerkhoven:1986:PCG

[Ker86] Thomas Kerkhoven. A proof of convergence of Gummel’s algorithm for realistic device geometries. SIAM Journal on Numerical Analysis, 23(6): 1121–1137, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kerkhoven:1988:SAD

[Ker88] Thomas Kerkhoven. A spectral analysis of the decoupling algorithm for semiconductor simulation. SIAM Journal on Numerical Analysis, 25(6): 1299–1312, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kerkhoven:1996:PLP

[Ker96] Thomas Kerkhoven. Piecewise linear Petrov–Galerkin error estimates for the box method. SIAM Journal on Numerical Analysis, 33(5): 1864–1884, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24182.

Keiper:1991:AGB

[KG91] J. B. Keiper and C. W. Gear. The analysis of generalized backwards difference formula methods applied to Hessenberg form differentialalgebraic equations. SIAM Journal on Numerical Analysis, 28(3):833–858, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kacur:1993:SND

[KHK93] J. Kačur, A. Handlovičová, and M. Kačurová. Solution of nonlinear diffusion problems by linear approximation schemes. SIAM Journal on Numerical Analysis, 30(6): 1703–1722, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Krishnamachari:1989:FEA

[KHR89] S. V. Krishnamachari, L. J.

Hayes, and T. F. Russell. A finite element alternatingdirection method combined with a modified method of characteristics for convectiondiffusion problems. SIAM Journal on Numerical Analysis, 26(6):1462–1473, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kimchi:1976:CSP

[Kim76] E. Kimchi. Chebyshev spaces of polynomials. SIAM Journal on Numerical Analysis, 13 (1):44–50, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

King:1972:ASP

[Kin72] J. Thomas King. The approximate solution of parabolic initial boundary value problems by weighted leastsquares methods. SIAM Journal on Numerical Analysis, 9 (2):215–229, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

King:1973:FFO

[Kin73] Richard F. King. A family of fourth order methods for nonlinear equations. SIAM Journal on Numerical Analysis, 10(5):876–879, October 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kincaid:1974:CSD

[Kin74] David R. Kincaid. On complex second-degree iterative methods. SIAM Journal on Numerical Analysis, 11(2): 211–218, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

King:1975:SPV

[Kin75] J. Thomas King. Semidiscrete perturbed variational methods for the numerical solution of parabolic boundary value problems. SIAM Journal on Numerical Analysis, 12 (1):59–72, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

King:1979:EMO

[Kin79] Richard F. King. An extrapolation method of order four for linear sequences. SIAM Journal on Numerical Analysis, 16(5):719–725, October 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kirmse:1993:BDD

[Kir93] Andreas Kirmse. Bendingdominated deformation of thin spherical shells: Analysis and finite-element approximation. SIAM Journal on Numerical Analysis, 30(4):1015– 1040, September 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karakashian:1998:NFE

[KJ98] Ohannes A. Karakashian and Wadi N. Jureidini. A nonconforming finite element method for the stationary Navier– Stokes equations. SIAM Journal on Numerical Analysis, 35(1):93–120, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29719.

Kappel:1981:SAN

[KK81a] F. Kappel and K. Kunisch.

Spline approximations for neutral functional-differential equations. SIAM Journal on Numerical Analysis, 18(6):

1058–1080, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1981:NMS

[KK81b] Barbro Kreiss and Heinz-Otto Kreiss. Numerical methods for singular perturbation problems. SIAM Journal on Numerical Analysis, 18(2): 262–276, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kroller:1991:CRF

[KK91] M. Kroller and K. Kunisch. Convergence rates for the feedback operators arising in the linear quadratic regulator problem governed by parabolic equations. SIAM Journal on Numerical Analysis, 28(5):1350–1385, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ku:1993:PIM

[KK93] Ta-Kang Ku and C.-C. Jay Kuo. Preconditioned iterative methods for solving Toeplitz-Plus-Hankel systems. SIAM Journal on Numerical Analysis, 30(3):824–845, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kelley:1998:CAP

[KK98a] C. T. Kelley and David E. Keyes. Convergence analysis of pseudo-transient continuation. SIAM Journal on Numerical Analysis, 35(2):508–523, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30479.

Kreiss:1998:SSS

[KK98b] Gunilla Kreiss and Heinz-Otto Kreiss. Stability of steady solutions of Burgers’s equation. SIAM Journal on Numerical Analysis, 35(6): 2329–2349, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31600.

Kreiss:1994:CSS

[KKP94] Gunilla Kreiss, Heinz-Otto Kreiss, and N. Anders Petersson. On the convergence to steady state of solutions of nonlinear hyperbolicparabolic systems. SIAM Journal on Numerical Analysis, 31(6):1577–1604, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Konen:1967:SSE

[KL67] H. P. Konen and H. A.

Luther. Some singular explicit fifth order Runge–Kutta solutions. SIAM Journal on Numerical Analysis, 4(4):607– 619, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kimchi:1976:RBA

[KL76] E. Kimchi and D. Leviatan. On restricted best approximation to functions with restricted derivatives. SIAM Journal on Numerical Analysis, 13(1):51–53, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keast:1979:SFS

[KL79] P. Keast and J. N. Lyness. On the structure of fully symmetric multidimensional quadrature rules. SIAM Journal on Numerical Analysis, 16(1): 11–29, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:1982:IIB

[KL82] Herbert B. Keller and Marianela Lentini. Invariant imbedding, the box scheme and an equivalence between them. SIAM Journal on Numerical Analysis, 19(5):942– 962, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kloeden:1986:SAS

[KL86] P. E. Kloeden and J. Lorenz. Stable attracting sets in dynamical systems and in their one-step discretizations.

SIAM Journal on Numerical Analysis, 23(5):986–995, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuo:1989:TLF

[KL89] C.-C. Jay Kuo and Bernard C.

Levy. A two-level four-color SOR method. SIAM Journal on Numerical Analysis, 26(1): 129–151, February 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kellogg:1996:FEM

[KL96] R. Bruce Kellogg and Biyue Liu. A finite element method for the compressible Stokes equations. SIAM Journal on Numerical Analysis, 33(2): 780–788, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kellogg:1997:PFE

[KL97] R. Bruce Kellogg and Biyue Liu. A penalized finiteelement method for a compressible Stokes system. SIAM Journal on Numerical Analysis, 34(3):1093–1105, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27327.

Klar:1998:AIS

[Kla98] Axel Klar. An asymptoticinduced scheme for nonstationary transport equations in the diffusive limit. SIAM Journal on Numerical Analysis, 35(3):1073– 1094, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30555.

Klar:1999:APN

[Kla99]  Axel Klar.  An asymp-

totic preserving numerical scheme for kinetic equations in the low Mach number limit. SIAM Journal on Numerical Analysis, 36(5): 1507–1527, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32176.

Koepfler:1994:MAI

[KLM94] G. Koepfler, C. Lopez, and J. M. Morel. A multiscale algorithm for image segmentation by variational method. SIAM Journal on Numerical Analysis, 31(1): 282–299, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kellogg:1976:CPB

[KLY76] R. B. Kellogg, T. Y. Li, and J. Yorke. A constructive proof of the Brouwer fixedpoint theorem and computational results. SIAM Journal on Numerical Analysis, 13 (4):473–483, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kwon:1988:EEM

[KM88] Yonghoon Kwon and Fabio A.

Milner. L^{∞}-error estimates for mixed methods for semilinear second-order elliptic equations. SIAM Journal on Numerical Analysis, 25(1): 46–53, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kramer:1993:AGM

[KM93] M. E. Kramer and R. M. M.

Mattheij. Application of global methods in parallel shooting. SIAM Journal on Numerical Analysis, 30(6): 1723–1739, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kunkel:1996:NCD

[KM96] Peter Kunkel and Volker Mehrmann. A new class of discretization methods for the solution of linear differentialalgebraic equations with variable coefficients. SIAM Journal on Numerical Analysis, 33 (5):1941–1961, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24036.

Karakashian:1999:STF

[KM99] Ohannes Karakashian and Charalambos Makridakis. A space-time finite element method for the nonlinear Schrödinger equation: The continuous Galerkin method. SIAM Journal on Numerical Analysis, 36(6):1779– 1807, December 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33011.

Kozlov:1994:CHI

[KMR94] Vladimir Kozlov, Vladimir Maz’ya, and Leonid Rozin. On certain hybrid iterative methods for solving boundary value problems. SIAM Journal on Numerical Analysis, 31(1):101–110, February 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammerer:1972:CCG

[KN72] W. J. Kammerer and M. Z. Nashed. On the convergence of the conjugate gradient method for singular linear operator equations. SIAM Journal on Numerical Analysis, 9(1):165–181, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Krawczyk:1985:ISR

[KN85] R. Krawczyk and A. Neumaier. Interval slopes for rational functions and associated centered forms. SIAM Journal on Numerical Analysis, 22(3):604–616, June 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kelley:1988:PQN

[KN88] C. T. Kelley and J. I.

Northrup. A pointwise quasi-Newton method for integral equations. SIAM Journal on Numerical Analysis, 25 (5):1138–1155, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1986:NMS

[KNB86] Heinz-Otto Kreiss, N. K.

Nichols, and David L. Brown. Numerical methods for stiff two-point boundary value problems. SIAM Journal on Numerical Analysis, 23(2): 325–368, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Knorrenschild:1992:DAE

[Kno92] Michael Knorrenschild. Differential/algebraic equations as stiff ordinary differential equations. SIAM Journal on Numerical Analysis, 29(6): 1694–1715, December 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1979:SFM

[KO79] Heinz-Otto Kreiss and Joseph Oliger. Stability of the Fourier method. SIAM Journal on Numerical Analysis, 16 (3):421–433, June 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Koch:1984:CST

[Koc84] Per Erik Koch. Collocation by L-splines at transformed Gaussian points. SIAM Journal on Numerical Analysis, 21(6):1107–1114, December 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kohler:1995:OPM

[Köh95] P. Köhler. Order-preserving mesh spacing for compound quadrature formulas and functions with endpoint singularities. SIAM Journal on Numerical Analysis, 32(2): 671–686, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Koster:1973:SFD

[Kos73]  D. E. Koster.  Stability

of a finite difference scheme with “wrong” boundary conditions. SIAM Journal on Numerical Analysis, 10(6):1039– 1046, December 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kovarik:1970:SIM

[Kov70] Zdislav Kovarik. Some iterative methods for improving orthonormality. SIAM Journal on Numerical Analysis, 7 (3):386–389, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaufman:1978:MSN

[KP78] Linda Kaufman and Victor Pereyra. A method for separable nonlinear least squares problems with separable nonlinear equality constraints. SIAM Journal on Numerical Analysis, 15(1):12–20, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:1979:DMD

[KP79] H. B. Keller and V. Pereyra.

Difference methods and deferred corrections for ordinary boundary value problems. SIAM Journal on Numerical Analysis, 16(2): 241–259, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Klauder:1985:NIM

[KP85] John R. Klauder and Wesley P. Petersen. Numerical integration of multiplicativenoise stochastic differential equations. SIAM Journal on Numerical Analysis, 22(6): 1153–1166, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kamowitz:1987:MM

[KP87] David Kamowitz and Seymour V. Parter. On MGR[ν] multigrid methods. SIAM Journal on Numerical Analysis, 24(2):366–381, April 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kim:1996:PCS

[KP96] Sang Dong Kim and Seymour V. Parter. Preconditioning Chebyshev spectral collocation method for elliptic partial differential equations. SIAM Journal on Numerical Analysis, 33(6): 2375–2400, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27599.

Kim:1997:PCS

[KP97] Sang Dong Kim and Seymour V. Parter. Preconditioning Chebyshev spectral collocation by finitedifference operators. SIAM Journal on Numerical Analysis, 34(3):939–958, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28503.

Kloeden:1995:EMW

[KPH95] P. E. Kloeden, E. Platen, and N. Hofmann. Extrapolation methods for the weak approximation of Ito diffusions. SIAM Journal on Numerical Analysis, 32(5):1519– 1534, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kahan:1982:RBA

[KPJ82] W. Kahan, B. N. Parlett, and E. Jiang. Residual bounds on approximate eigensystems of nonnormal matrices. SIAM Journal on Numerical Analysis, 19(3):470–484, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kammerer:1972:LCS

[KR72] W. J. Kammerer and G. W.

Reddien, Jr. Local convergence of smooth cubic spline interpolates. SIAM Journal on Numerical Analysis, 9 (4):687–694, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Krabill:1974:ERN

[KR74] D. Krabill and M. Reichaw.

Expansions of real numbers relative to a sequence of functions. SIAM Journal on Numerical Analysis, 11(1):

75–86, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreimer:1988:SFC

[KR88] J. Kreimer and R. Y. Rubinstein. Smoothed functionals and constrained stochastic approximation. SIAM Journal on Numerical Analysis, 25 (2):470–487, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kloucek:1994:SFS

[KR94a] Petr Klouček and Franz S. Rys. Stability of the fractional step Θ-scheme for the nonstationary Navier–Stokes equations. SIAM Journal on Numerical Analysis, 31(5):1312– 1335, October 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kroner:1994:CUF

[KR94b] Dietmar Kröner and Mirko Rokyta. Convergence of upwind finite volume schemes for scalar conservation laws in two dimensions. SIAM Journal on Numerical Analysis, 31 (2):324–343, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kapur:1997:HOC

[KR97] Sharad Kapur and Vladimir Rokhlin. High-order corrected trapezoidal quadrature rules for singular functions. SIAM Journal on Numerical Analysis, 34(4): 1331–1356, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28784.

Krawczyk:1987:OEG

[Kra87] R. Krawczyk. Optimal enclosure of a generalized zero set of a function strip. SIAM Journal on Numerical Analysis, 24(5):1202–1211, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kimchi:1978:RRA

[KRD78] E. Kimchi and N. Richter-Dyn. Restricted range approximation of k-convex functions in monotone norms. SIAM Journal on Numerical Analysis, 15(5):1030– 1038, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1978:DMS

[Kre78a] Heinz-Otto Kreiss. Difference methods for stiff ordinary differential equations. SIAM Journal on Numerical Analysis, 15(1):21–58, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kress:1978:ENT

[Kre78b] Rainer Kress. On error norms of the trapezoidal rule. SIAM Journal on Numerical Analysis, 15(3):433–443, June 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1979:PDT

[Kre79] Heinz-Otto Kreiss. Problems with different time scales for ordinary differential equations. SIAM Journal on Numerical Analysis, 16(6):980– 998, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1991:DOB

[Kre91] Gunilla Kreiss. The dependence on the outflow boundary condition of the solution of steady, incompressible Euler equations. SIAM Journal on Numerical Analysis, 28 (5):1242–1264, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Krizek:1992:MAC

[Kří92]  Michal Křížek.  On the

maximum angle condition for linear tetrahedral elements. SIAM Journal on Numerical Analysis, 29(2): 513–520, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Krogh:1973:ACS

[Kro73]  Fred T. Krogh.  Algo-

rithms for changing the step size. SIAM Journal on Numerical Analysis, 10(5):949– 965, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kroo:1981:EED

[Kro81]  A. Kroó.  Error estima-

tions for deviation of best uniform, discrete and L_{q}- approximations. SIAM Journal on Numerical Analysis, 18(5):891–896, October 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kruzik:1998:NAD

[Kru98]  Martin Kruzík.  Numeri-

cal approach to double well problems. SIAM Journal on Numerical Analysis, 35 (5):1833–1849, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28647.

Kelley:1983:NAM

[KS83] C. T. Kelley and R. Suresh. A new acceleration method for Newton’s method at singular points. SIAM Journal on Numerical Analysis, 20(5):1001– 1009, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kelley:1987:QNM

[KS87] C. T. Kelley and E. W. Sachs. A quasi-Newton method for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 24(3): 516–531, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kreiss:1992:MLH

[KS92a] Heinz-Otto Kreiss and Godela Scherer. Method of lines for hyperbolic differential equations. SIAM Journal on Numerical Analysis, 29(3): 640–646, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kunisch:1992:RSM

[KS92b] K. Kunisch and E. W. Sachs. Reduced SQP methods for parameter identification problems. SIAM Journal on Numerical Analysis, 29(6):1793– 1820, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Knyazev:1994:PGT

[KS94] Andrew V. Knyazev and Alexander L. Skorokhodov. Preconditioned gradient-type iterative methods in a subspace for partial generalized symmetric eigenvalue problems. SIAM Journal on Numerical Analysis, 31(4):1226– 1239, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kasenally:1997:AMP

[KS97a] Ebrahim M. Kasenally and Valeria Simoncini. Analysis of a minimum perturbation algorithm for nonsymmetric linear systems. SIAM Journal on Numerical Analysis, 34(1):48–66, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26684.

Kellogg:1997:OAS

[KS97b] R. Bruce Kellogg and Martin Stynes. Optimal approximability of solutions of singularly perturbed twopoint boundary value problems. SIAM Journal on Numerical Analysis, 34(5): 1808–1816, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29026.

Kellogg:1999:WSP

[KS99] R. Bruce Kellogg and Martin Stynes. n-widths and singularly perturbed boundary value problems. SIAM Journal on Numerical Analysis, 36 (5):1604–1620, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32725.

Koc:1999:EAN

[KSC99] S. Koc, Jiming Song, and W. C. Chew. Error analysis for the numerical evaluation of the diagonal forms of the scalar spherical addition theorem. SIAM Journal on Numerical Analysis, 36(3):906–921, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32811.

Kellogg:1980:UCR

[KSS80] R. B. Kellogg, G. R. Shubin, and A. B. Stephens. Uniqueness and the cell Reynolds number. SIAM Journal on Numerical Analysis, 17 (6):733–739, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kung:1976:OOE

[KT76] H. T. Kung and J. F. Traub. Optimal order and efficiency for iterations with two evaluations. SIAM Journal on Numerical Analysis, 13(1): 84–99, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Khaliq:1986:SSM

[KT86] A. Q. M. Khaliq and E. H.

Twizell. L_{0}-stable splitting methods for the simple heat equation in two space dimensions with homogeneous boundary conditions. SIAM Journal on Numerical Analysis, 23(3):473–484, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuo:1992:DMF

[KT92] Hung Ju Kuo and Neil S.

Trudinger. Discrete methods for fully nonlinear elliptic equations. SIAM Journal on Numerical Analysis, 29(1): 123–135, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kunisch:1997:SPS

[KT97] K. Kunisch and X.-C. Tai.

Sequential and parallel splitting methods for bilinear control problems in Hilbert spaces. SIAM Journal on Numerical Analysis, 34(1):91– 118, February 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25525.

Kahaner:1971:EWQ

[KU71] David K. Kahaner and J. L.

Ullman. Equal weight quadrature on infinite intervals.

SIAM Journal on Numerical Analysis, 8(1):75–79, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kublanovskaya:1970:ASG

[Kub70] V. N. Kublanovskaya. On an approach to the solution of the generalized latent value problem for λ-matrices. SIAM Journal on Numerical Analysis, 7(4):532–537, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kumar:1988:DCT

[Kum88]  Sunil Kumar.  A dis-

crete collocation-type method for Hammerstein equations. SIAM Journal on Numerical Analysis, 25(2):328–341, April 1988. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kung:1975:CCA

[Kun75] H. T. Kung. The computational complexity of algebraic numbers. SIAM Journal on Numerical Analysis, 12 (1):89–96, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kunkel:1988:QCM

[Kun88] P. Kunkel. Quadratically convergent methods for the computation of unfolded singularities. SIAM Journal on Numerical Analysis, 25(6):1392– 1408, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kurowski:1967:SRF

[Kur67]  G. J. Kurowski.  The

semidiscrete Riemann function. SIAM Journal on Numerical Analysis, 4(4):489– 498, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kushner:1968:NSD

[Kus68] Harold J. Kushner. On the numerical solution of degenerate linear and nonlinear elliptic boundary value problems. SIAM Journal on Numerical Analysis, 5(4):664– 679, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuttler:1970:FDA

[Kut70] James R. Kuttler. Finite difference approximations for eigenvalues of uniformly elliptic operators. SIAM Journal on Numerical Analysis, 7 (2):206–232, June 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuttler:1972:RSE

[Kut72]  James R. Kuttler.  Re-

marks on a Stekloff eigenvalue problem. SIAM Journal on Numerical Analysis, 9 (1):1–5, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuttler:1974:DMC

[Kut74] J. R. Kuttler. Direct methods for computing eigenvalues of the finite-difference laplacian. SIAM Journal on Numerical Analysis, 11(4):732– 740, September 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuttler:1979:DE

[Kut79]  J. R. Kuttler.  Dirichlet

eigenvalues. SIAM Journal on Numerical Analysis, 16(2): 332–338, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuttler:1982:BSE

[Kut82] J. R. Kuttler. Bounds for Stekloff eigenvalues. SIAM Journal on Numerical Analysis, 19(1):121–125, February 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kutz:1984:AEB

[Küt84] Martin Kütz. Asymptotic error bounds for a class of interpolatory quadratures. SIAM Journal on Numerical Analysis, 21(1):167–175, February 1984. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuznetsov:1999:NNT

[Kuz99]  Yu. A. Kuznetsov.  Nu-

merical normalization techniques for all codim 2 bifurcations of equilibria in ODE’s. SIAM Journal on Numerical Analysis, 36(4): 1104–1124, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33500.

Keller:1975:DMB

[KW75] H. B. Keller and A. B.

White, Jr. Difference methods for boundary value problems in ordinary differential equations. SIAM Journal on Numerical Analysis, 12(5): 791–802, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kendall:1976:CNG

[KW76] Richard P. Kendall and Mary F. Wheeler. A Crank– Nicolson-H^{−1}–Galerkin procedure for parabolic problems in a single-space variable. SIAM Journal on Numerical Analysis, 13(6):861– 876, December 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaufman:1984:HOF

[KW84] Linda Kaufman and Daniel D. Warner. High-order, fastdirect methods for separable elliptic equations. SIAM Journal on Numerical Analysis, 21(4):672–694, August 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Katz:1985:VFE

[KW85] I. Norman Katz and Douglas W. Wang. The p-version of the finite element method for problems requiring C^{1}- continuity. SIAM Journal on Numerical Analysis, 22(6): 1082–1106, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Krishna:1993:SLA

[KW93] Hari Krishna and Yunbiao Wang. The split Levinson algorithm is weakly stable. SIAM Journal on Numerical Analysis, 30(5):1498– 1508, October 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kearfott:1994:ISC

[KX94] R. Baker Kearfott and Zhaoyun Xing. An interval step control for continuation methods. SIAM Journal on Numerical Analysis, 31 (3):892–914, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaneko:1996:SIG

[KX96] Hideaki Kaneko and Yuesheng Xu. Superconvergence of the iterated Galerkin methods for Hammerstein equations. SIAM Journal on Numerical Analysis, 33(3): 1048–1064, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karlin:1966:CSF

[KZ66] Samuel Karlin and Zvi Ziegler. Chebyshevian spline functions. SIAM Journal on Numerical Analysis, 3 (3):514–543, September 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lambert:1967:ANS

[Lam67] Robert J. Lambert. An analysis of the numerical stability of predictor-corrector solutions of nonlinear ordinary differential equations. SIAM Journal on Numerical Analysis, 4(4):597–606, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lambert:1971:PCA

[Lam71] J. D. Lambert. Predictorcorrector algorithms with identical regions of stability. SIAM Journal on Numerical Analysis, 8(2):337–344, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lam:1976:NBU

[Lam76] Binh Lam. A note on best uniform rational approximation. SIAM Journal on Numerical Analysis, 13(6):962– 965, December 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lambert:1981:SPM

[Lam81] J. D. Lambert. Safe point methods for separably stiff systems of ordinary differential equations. SIAM Journal on Numerical Analysis, 18(1): 83–101, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

Lanczos:1964:END

[Lan64a] C. Lanczos. Evaluation of noisy data. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):76– 85, ???? 1964. ISSN 0887-459X (print), 1095-7170 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/2949766.

Lanczos:1964:I

[Lan64b] Cornelius Lanczos. Introduction. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):1, ???? 1964. ISSN 0887-459X (print), 1095-7170 (elec-

tronic). URL http://www.
jstor.org/stable/2949759.

Lanczos:1964:PAG

[Lan64c] Cornelius Lanczos. A precision approximation of the gamma function. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(1): 86–96, ???? 1964. ISSN 0887-459X (print), 1095-7170 (elec-

tronic). URL http://www.
jstor.org/stable/2949767.

Langford:1977:SAB

[Lan77] W. F. Langford. A shooting algorithm for the best least squares solution of twopoint boundary value problems. SIAM Journal on Numerical Analysis, 14(3): 527–542, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Larsson:1989:LTB

[Lar89] Stig Larsson. The long-time behavior of finite-element approximations of solutions to semilinear parabolic problems. SIAM Journal on Numerical Analysis, 26(2): 348–365, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lasiecka:1984:CES

[Las84]  Irena Lasiecka.  Conver-

gence estimates for semidiscrete approximations of nonselfadjoint parabolic equations. SIAM Journal on Numerical Analysis, 21(5):894– 909, October 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lautsch:1989:SIF

[Lau89] Michael Lautsch. A spline inversion formula for the Radon transform. SIAM Journal on Numerical Analysis, 26(2): 456–467, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lavery:1989:SSS

[Lav89] J. E. Lavery. Solution of steady-state one-dimensional conservation laws by mathematical programming. SIAM Journal on Numerical Analysis, 26(5):1081–1089, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lavery:1991:SSS

[Lav91] John E. Lavery. Solution of steady-state, two-dimensional conservation laws by mathematical programming. SIAM Journal on Numerical Analysis, 28(1):141–155, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lawson:1966:OFR

[Law66] J. Douglas Lawson. An order five Runge–Kutta process with extended region of stability. SIAM Journal on Numerical Analysis, 3(4):593– 597, December 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lawson:1967:GRK

[Law67a] J. Douglas Lawson. Generalized Runge–Kutta processes for stable systems with large Lipschitz constants. SIAM Journal on Numerical Analysis, 4(3):372–380, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lawson:1967:OSR

[Law67b] J. Douglas Lawson. An order six Runge–Kutta process with extended region of stability. SIAM Journal on Numerical Analysis, 4(4):620– 625, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Layton:1983:GMT

[Lay83a] William J. Layton. Galerkin methods for two-point boundary value problems for first order systems. SIAM Journal on Numerical Analysis, 20(1): 161–171, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Layton:1983:SGM

[Lay83b] William J. Layton. Stable Galerkin methods for hyperbolic systems. SIAM Journal on Numerical Analysis, 20 (2):221–233, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lazarus:1981:SSS

[Laz81]  Roger B. Lazarus.  Self-

similar solutions for converging shocks and collapsing cavities. SIAM Journal on Numerical Analysis, 18(2): 316–371, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Laz82].

Lazarus:1982:ESS

[Laz82] Roger B. Lazarus. Erratum: “Self-Similar Solutions for Converging Shocks and Collapsing Cavities” [SIAM J. Numer. Anal. 18 (1981), no. 2, 316–371, MR 82i:76054]. SIAM Journal on Numerical Analysis, 19(5):1090, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Laz81].

Levrie:1990:CAN

[LB90] Paul Levrie and Adhemar Bultheel. Convergence acceleration for the numerical solution of second-order linear recurrence relations. SIAM Journal on Numerical Analysis, 27(1):166–177, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1996:CAV

[LB96a] Y. Li and I. Babuška.

A convergence analysis of a p-version finite element method for one-dimensional elastoplasticity problem with constitutive laws based on the gauge function method. SIAM Journal on Numerical Analysis, 33(2):809–842, April 1996. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:1996:FEA

[LB96b] W. B. Liu and John W.

Barrett. Finite element approximation of some degenerate monotone quasilinear elliptic systems. SIAM Journal on Numerical Analysis, 33(1): 88–106, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1997:CAV

[LB97] Yiwei Li and Ivo Babuška.

A convergence analysis of an H-version finite-element method with high-order elements for two-dimensional elasto-plasticity problems. SIAM Journal on Numerical Analysis, 34(3):998– 1036, June 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26357.

Laubin:1998:SCB

[LB98] Pascal Laubin and Marc Baiwir. Spline collocation for a boundary integral equation on polygons with cuts. SIAM Journal on Numerical Analysis, 35 (4):1452–1472, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31835.

Lecot:1998:QMC

[LC98] Christian Lécot and Ibrahim Coulibaly. A quasi-Monte Carlo scheme using nets for a linear Boltzmann equation. SIAM Journal on Numerical Analysis, 35(1):51– 70, February 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29005.

LeRoux:1994:STN

[Le 94] Marie-Noëlle Le Roux. Semidiscretization in time of nonlinear parabolic equations with blowup of the solution. SIAM Journal on Numerical Analysis, 31(1):170–195, February 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lam:1972:CCW

[LE72] Binh Lam and David Elliott.

On a conjecture of C. W. Clenshaw. SIAM Journal on Numerical Analysis, 9(1): 44–52, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lamm:1997:NSF

[LE97] Patricia K. Lamm and Lars Eldén. Numerical solution of first-kind Volterra equations by sequential Tikhonov regularization. SIAM Journal on Numerical Analysis, 34 (4):1432–1450, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28081.

Lee:1985:RAE

[Lee85] Daniel Lee. On remainders and asymptotic error expansions in cardinal spline interpolation. SIAM Journal on Numerical Analysis, 22(6): 1238–1242, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lee:1991:EER

[Lee91] Meei-Yow Lin Lee. Estimation of the error in the reduced basis method solution of differential algebraic equation systems. SIAM Journal on Numerical Analysis, 28 (2):512–528, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lee:1992:CCP

[Lee92] E. T. Y. Lee. Corners, cusps, and parametrizations: Variations on a theorem of epstein. SIAM Journal on Numerical Analysis, 29(2): 553–565, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lee:1993:LMM

[Lee93] Ping Lee. A Lagrange multiplier method for the interface equations from electromagnetic applications. SIAM Journal on Numerical Analysis, 30(2):478–506, April 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lee:1998:MMP

[Lee98]  Chang-Ock Lee.  Multi-

grid methods for the pure traction problem of linear elasticity: Mixed formulation. SIAM Journal on Numerical Analysis, 35(1):121– 145, February 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28283.

Leimkuhler:1998:TAW

[Lei98] B. Leimkuhler. Timestep acceleration of waveform relaxation. SIAM Journal on Numerical Analysis, 35 (1):31–50, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28002.

Lenoir:1986:OIF

[Len86] M. Lenoir. Optimal isoparametric finite elements and error estimates for domains involving curved boundaries. SIAM Journal on Numerical Analysis, 23(3):562–580, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lermit:1975:NMI

[Ler75] R. Jonathan Lermit. Numerical methods for the identification of differential equations. SIAM Journal on Numerical Analysis, 12(3):488–500, June 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lescrenier:1991:CTR

[Les91] M. Lescrenier. Convergence of trust region algorithms for optimization with bounds when strict complementarity does not hold. SIAM Journal on Numerical Analysis, 28(2): 476–495, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lether:1970:ERP

[Let70] Frank Lether. An error representation for product cubature rules. SIAM Journal on Numerical Analysis, 7 (3):363–365, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lether:1971:CEB

[Let71a] Frank G. Lether. Cubature error bounds for Gauss– Legendre product rules. SIAM Journal on Numerical Analysis, 8(1):36–42, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lether:1971:EBF

[Let71b]  Frank G. Lether.  Error

bounds for fully symmetric cubature rules. SIAM Journal on Numerical Analysis, 8 (1):49–60, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lether:1971:GPR

[Let71c] Frank G. Lether. A generalized product rule for the circle. SIAM Journal on Numerical Analysis, 8(2): 249–253, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lether:1974:CEC

[Let74a] F. G. Lether. Cubature error constants for analytic functions. SIAM Journal on Numerical Analysis, 11(2): 232–236, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lether:1974:EBF

[Let74b] F. G. Lether. Error bounds for fully symmetric quadrature rules. SIAM Journal on Numerical Analysis, 11 (1):1–9, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Leveque:1982:LTS

[Lev82] Randall J. Leveque. Large time step shock-capturing techniques for scalar conservation laws. SIAM Journal on Numerical Analysis, 19(6): 1091–1109, December 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Leveque:1985:LTS

[Lev85]  Randall J. Leveque.  A

large time step generalization of Godunov’s method for systems of conservation laws. SIAM Journal on Numerical Analysis, 22(6):1051– 1073, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

LeVeque:1988:SOA

[LeV88] Randall J. LeVeque. Second order accuracy of Brenier’s time-discrete method for nonlinear systems of conservation laws. SIAM Journal on Numerical Analysis, 25(1): 1–7, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

LeVeque:1996:HRC

[LeV96] Randall J. LeVeque. Highresolution conservative algorithms for advection in incompressible flow. SIAM Journal on Numerical Analysis, 33 (2):627–665, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1998:DTO

[LFB98] Bingkun Li, Graeme Fairweather, and Bernard Bialecki. Discrete-time orthogonal spline collocation methods for Schrödinger equations in two space variables. SIAM Journal on Numerical Analysis, 35(2): 453–477, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30239.

Lyness:1980:STR

[LG80] J. N. Lyness and A. C. Genz. On simplex trapezoidal rule families. SIAM Journal on Numerical Analysis, 17(1): 126–147, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lopez-Gomez:1992:NCC

[LGMMV92] Julián López-Gómez, Marcela Molina-Meyer, and Mónica Villarreal. Numerical computation of coexistence states. SIAM Journal on Numerical Analysis, 29(4):1074– 1092, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1983:LAZ

[Li83a] Tien-Yien Li. On locating all zeros of an analytic function within a bounded domain by a revised Delves/ Lyness method. SIAM Journal on Numerical Analysis, 20(4):865–871, August 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1983:RRC

[Li83b] Zi Cai Li. On the reduced rate of convergence for a nonconforming combined method. SIAM Journal on Numerical Analysis, 20(1):86–93, February 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1987:GDM

[Li87] Ronghua Li. Generalized difference methods for a nonlinear Dirichlet problem. SIAM Journal on Numerical Analysis, 24(1):77–88, February 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1988:SFD

[Li88] Guangye Li. The secant/ finite difference algorithm for solving sparse nonlinear systems of equations. SIAM Journal on Numerical Analysis, 25(5):1181–1196, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1991:NCV

[Li91] Zi-Cai Li. On nonconforming combinations of various finite element methods for solving elliptic boundary value problems. SIAM Journal on Numerical Analysis, 28(2): 446–475, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1998:FIA

[Li98] Zhilin Li. A fast iterative algorithm for elliptic interface problems. SIAM Journal on Numerical Analysis, 35 (1):230–254, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29132.

Linz:1969:NMV

[Lin69] Peter Linz. Numerical methods for Volterra integral equations with singular kernels. SIAM Journal on Numerical Analysis, 6(3):365– 374, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lindberg:1972:SIA

[Lin72] Bengt Lindberg. A simple interpolation algorithm for improvement of the numerical solution of a differential equation. SIAM Journal on Numerical Analysis, 9(4):662– 668, December 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Linz:1974:MAS

[Lin74] Peter Linz. A method for the approximate solution of linear integro-differential equations. SIAM Journal on Numerical Analysis, 11(1):137– 144, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lindberg:1977:COS

[Lin77a] Bengt Lindberg. Characterization of optimal stepsize sequences for methods for stiff differential equations. SIAM Journal on Numerical Analysis, 14(5):859–887, September 1977. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Linz:1977:GTA

[Lin77b] Peter Linz. A general theory for the approximate solution of operator equations of the second kind. SIAM Journal on Numerical Analysis, 14 (3):543–554, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liniger:1983:CSO

[Lin83] Werner Liniger. The Acontractive second-order multistep formulas with variable steps. SIAM Journal on Numerical Analysis, 20(6):1231– 1238, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1985:NSH

[Lin85] Tzu Chu Lin. The numerical solution of Helmholtz’s equation for the exterior Dirichlet problem in three dimensions. SIAM Journal on Numerical Analysis, 22(4):670– 686, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1988:NAP

[Lin88] J. T. Lin. The numerical analysis of a phase field model in moving boundary problems. SIAM Journal on Numerical Analysis, 25(5):1015– 1031, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1990:GMN

[Lin90] Yanping Lin. Galerkin methods for nonlinear parabolic integrodifferential equations with nonlinear boundary conditions. SIAM Journal on Numerical Analysis, 27(3): 608–621, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Linz:1991:BEC

[Lin91] Peter Linz. Bounds and estimates for condition numbers of integral equations. SIAM Journal on Numerical Analysis, 28(1):227–235, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Linzer:1992:STB

[Lin92] Elliot Linzer. On the stability of transform-based circular deconvolution. SIAM Journal on Numerical Analysis, 29(5):1482–1492, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1997:SRM

[Lin97]  Ping Lin.  A sequential

regularization method for time-dependent incompressible Navier–Stokes equations.

SIAM Journal on Numerical Analysis, 34(3):1051– 1071, June 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27052.

Lipow:1973:SFI

[Lip73] Peter R. Lipow. Spline functions and intermediate best quadrature formulas. SIAM Journal on Numerical Analysis, 10(1):127–136, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:1993:MPS

[Liu93] Xu-Dong Liu. A maximum principle satisfying modification of triangle based adaptive stencils for the solution of scalar hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 30(3): 701–716, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:1994:SAL

[Liu94] Jun Liu. A sensitivity analysis for least-squares ill-posed problems using the Haar basis. SIAM Journal on Numerical Analysis, 31(5):1486– 1496, October 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:1999:NSI

[Liu99]  Yunkang Liu.  Numerical

solution of implicit neutral functional differential equations. SIAM Journal on Numerical Analysis, 36(2): 516–528, April 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31867.

Leaf:1974:EBM

[LK74] Gary K. Leaf and Hans G.

Kaper. L^{∞}-error bounds for multivariate Lagrange approximation. SIAM Journal on Numerical Analysis, 11(2): 363–381, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lentini:1980:BVP

[LK80] Marianela Lentini and Herbert B. Keller. Boundary value problems on semiinfinite intervals and their numerical solution. SIAM Journal on Numerical Analysis, 17(4):577–604, August 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ladeveze:1983:EEP

[LL83] P. Ladevèze and D. Leguillon. Error estimate procedure in the finite element method and applications. SIAM Journal on Numerical Analysis, 20 (3):485–509, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1989:RML

[LL89] San Yih Lin and Mitchell Luskin. Relaxation methods for liquid crystal problems. SIAM Journal on Numerical Analysis, 26(6):1310– 1324, December 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

LeVeque:1994:IIM

[LL94] Randall J. LeVeque and Zhi Lin Li. The immersed interface method for elliptic equations with discontinuous coefficients and singular sources. SIAM Journal on Numerical Analysis, 31(4): 1019–1044, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [LL95].

LeVeque:1995:EII

[LL95] Randall J. LeVeque and Zhi Lin Li. Erratum: “The immersed interface method for elliptic equations with discontinuous coefficients and singular sources” [SIAM J. Numer. Anal. 31 (1994), no. 4, 1019–1044, MR 95g:65139]. SIAM Journal on Numerical Analysis, 32(5):1704, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [LL94].

Largillier:1996:FRM

[LL96a] Alain Largillier and Balmohan V. Limaye. Finite-rank methods and their stability for coupled systems of operator equations. SIAM Journal on Numerical Analysis, 33 (2):707–728, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Layton:1996:MMI

[LL96b] W. Layton and H. W. J.

Lenferink. A multilevel mesh independence principle for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 33(1):17–30, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1998:FEA

[LL98] Bo Li and Mitchell Luskin.

Finite element analysis of microstructure for the cubic to tetragonal transformation. SIAM Journal on Numerical Analysis, 35 (1):376–392, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30111.

Luo:1999:MFE

[LL99] Zhendong Luo and Ruxun Liu. Mixed finite element analysis and numerical solitary solution for the RLW equation. SIAM Journal on Numerical Analysis, 36 (1):89–104, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31299.

Larcher:1996:OPN

[LLNS96] G. Larcher, A. Lauss, H. Niederreiter, and W. Ch. Schmid. Optimal polynomials for (t, m, s)-nets and numerical integration of multivariate Walsh series. SIAM Journal on Numerical Analysis, 33(6): 2239–2253, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26470.

Lyness:1967:NDA

[LM67] J. N. Lyness and C. B. Moler. Numerical differentiation of analytic functions. SIAM Journal on Numerical Analysis, 4(2):202–210, June 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyness:1977:QRR

[LM77] J. N. Lyness and G. Monegato. Quadrature rules for regions having regular hexagonal symmetry. SIAM Journal on Numerical Analysis, 14 (2):283–295, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lawson:1978:EFO

[LM78] J. D. Lawson and J. Ll. Morris. The extrapolation of first order methods for parabolic partial differential equations. I. SIAM Journal on Numerical Analysis, 15(6):1212– 1224, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lions:1979:SAS

[LM79] P. L. Lions and B. Mercier.

Splitting algorithms for the sum of two nonlinear operators. SIAM Journal on Numerical Analysis, 16(6):964– 979, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyche:1986:MOA

[LM86] T. Lyche and K. Mørken.

Making the OSLO algorithm more efficient. SIAM Journal on Numerical Analysis, 23 (3):663–675, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lasiecka:1988:DCR

[LM88] I. Lasiecka and A. Manitius. Differentiability and convergence rates of approximating semigroups for retarded functional-differential equations. SIAM Journal on Numerical Analysis, 25(4): 883–907, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lentini:1990:CDD

[LM90a] M. Lentini and R. März. Conditioning and dichotomy in differential algebraic equations. SIAM Journal on Numerical Analysis, 27(6):1519– 1526, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lentini:1990:CBV

[LM90b] M. Lentini and R. März. The conditioning of boundary value problems in transferable differential-algebraic equations. SIAM Journal on Numerical Analysis, 27(4): 1001–1015, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Luskin:1992:AFE

[LM92] Mitchell Luskin and Ling Ma. Analysis of the finite element approximation of microstructure in micromagnetics. SIAM Journal on Numerical Analysis, 29(2): 320–331, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1987:BMS

[LMS87] Zi-Cai Li, Rudolf Mathon, and Pavol Sermer. Boundary methods for solving elliptic problems with singularities and interfaces. SIAM Journal on Numerical Analysis, 24 (3):487–498, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1997:CGS

[LMS97] Peixiong Lin, K. W. Morton, and E. Süli. Characteristic Galerkin schemes for scalar conservation laws in two and three space dimensions. SIAM Journal on Numerical Analysis, 34(2):779–796, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25929.

Loeb:1969:UGW

[LMST69] H. L. Loeb, D. G. Moursund, L. L. Schumaker, and G. D. Taylor. Uniform generalized weight function polynomial approximation with interpolation. SIAM Journal on Numerical Analysis, 6(2): 284–293, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lazarov:1996:FVM

[LMV96] R. D. Lazarov, Ilya D.

Mishev, and P. S. Vassilevski. Finite volume methods for convection-diffusion problems. SIAM Journal on Numerical Analysis, 33(1): 31–55, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Larsen:1982:FDA

[LN82] Edward W. Larsen and Paul Nelson. Finite-difference approximations and superconvergence for the discreteordinate equations in slab geometry. SIAM Journal on Numerical Analysis, 19(2): 334–348, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1997:PWH

[LNC97] Fu-Rong Lin, Michael K. Ng, and Raymond H. Chan. Preconditioners for Wiener–Hopf equations with high-order quadrature rules. SIAM Journal on Numerical Analysis, 34 (4):1418–1431, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27098.

Li:1982:FSF

[LNOP82] M. R. Li, B. Nour-Omid, and B. N. Parlett. A fast solver free of fill-in for finite element problems. SIAM Journal on Numerical Analysis, 19(6):1233–1242, December 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lubich:1995:RKS

[LNS95] Ch. Lubich, K. Nipp, and D. Stoffer. Runge–Kutta solutions of stiff differential equations near stationary points. SIAM Journal on Numerical Analysis, 32(4): 1296–1307, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lozier:1990:CPL

[LO90] D. W. Lozier and F. W. J.

Olver. Closure and precision in level-index arithmetic. SIAM Journal on Numerical Analysis, 27(5):1295– 1304, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:1996:NHO

[LO96] Xu-Dong Liu and Stanley Osher. Nonoscillatory high order accurate self-similar maximum principle satisfying shock capturing schemes I. SIAM Journal on Numerical Analysis, 33(2):760– 779, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Locher:1973:NBQ

[Loc73] Franz Locher. Norm bounds of quadrature processes.

SIAM Journal on Numerical Analysis, 10(4):553–558, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Loeb:1966:AGR

[Loe66] H. L. Loeb. Approximation by generalized rationals. SIAM Journal on Numerical Analysis, 3(1):34–55, March 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Loewy:1970:GTT

[Loe70] Raphael Loewy. Gerschgorintype theorems for partitioned matrices. SIAM Journal on Numerical Analysis, 7(1): 125–128, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lorentz:1964:EA

[Lor64] G. G. Lorentz. Entropy and its applications. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??): 97–103, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Lorenz:1984:ADS

[Lor84] Jens Lorenz. Analysis of difference schemes for a stationary shock problem. SIAM Journal on Numerical Analysis, 21(6):1038–1053, December 1984. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Lentini:1985:CRB

[LOR85] Marianela Lentini, Michael R.

Osborne, and Robert D. Russell. The close relationships between methods for solving two-point boundary value problems. SIAM Journal on Numerical Analysis, 22(2): 280–309, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lord:1997:AIM

[Lor97] Gabriel James Lord. Attractors and inertial manifolds for finite-difference approximations of the complex Ginzburg–Landau equation. SIAM Journal on Numerical Analysis, 34(4): 1483–1512, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28554.

Lotstedt:1992:GIC

[Löt92] Per Lötstedt. Grid independent convergence of the multigrid method for firstorder equations. SIAM Journal on Numerical Analysis, 29 (5):1370–1394, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lovadina:1996:NCM

[Lov96] C. Lovadina. A new class of mixed finite element methods for Reissner–Mindlin plates. SIAM Journal on Numerical Analysis, 33(6): 2457–2467, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26506.

Lentini:1977:AFD

[LP77] M. Lentini and V. Pereyra. An adaptive finite difference solver for nonlinear two-point boundary problems with mild boundary layers. SIAM Journal on Numerical Analysis, 14 (1):91–111, March 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Locker:1978:OEE

[LP78] John Locker and P. M. Prenter. Optimal L^{2} and L^{∞} error estimates for continuous and discrete least squares methods for boundary value problems. SIAM Journal on Numerical Analysis, 15(6): 1151–1160, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Locker:1980:RDO

[LP80] John Locker and P. M. Prenter. Regularization with differential operators. II. weak least squares finite element solutions to first kind integral equations. SIAM Journal on Numerical Analysis, 17 (2):247–267, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lesaint:1991:EAR

[LP91] P. Lesaint and J. Pousin. Existence and approximation results for thermal boundary layer equations of reactive flows. SIAM Journal on Numerical Analysis, 28(4):1030– 1046, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Leimkuhler:1991:AMC

[LPG91] B. Leimkuhler, L. R. Petzold, and C. W. Gear. Approximation methods for the consistent initialization of differential-algebraic equations. SIAM Journal on Numerical Analysis, 28(1):205– 226, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

LeRoux:1984:CAL

[LQ84] A. Y. Le Roux and P. Quesseveur. Convergence of an antidiffusion Lagrange–Euler scheme for quasi-linear equations. SIAM Journal on Numerical Analysis, 21(5):985– 994, October 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lucas:1972:SCM

[LR72] Thomas R. Lucas and George W. Reddien, Jr. Some collocation methods for nonlinear boundary value problems. SIAM Journal on Numerical Analysis, 9(2):341–356, June 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Luskin:1982:SPG

[LR82] Mitchell Luskin and Rolf Rannacher. On the smoothing property of the Galerkin method for parabolic equations. SIAM Journal on Numerical Analysis, 19(1):93– 113, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lipton:1979:GND

[LRT79] Richard J. Lipton, Donald J. Rose, and Robert Endre Tarjan. Generalized nested dissection. SIAM Journal on Numerical Analysis, 16(2): 346–358, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lambert:1972:MMV

[LS72] J. D. Lambert and S. T. Sigurdsson. Multistep methods with variable matrix coefficients. SIAM Journal on Numerical Analysis, 9(4):715– 733, December 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyche:1973:CSI

[LS73] Tom Lyche and Larry L.

Schumaker. Computation of smoothing and interpolating natural splines via local bases. SIAM Journal on Numerical Analysis, 10(6):1027– 1038, December 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:1976:CAC

[LS76] Wai-Hung Liu and Andrew H.

Sherman. Comparative analysis of the Cuthill–McKee and the reverse Cuthill–McKee ordering algorithms for sparse matrices. SIAM Journal on Numerical Analysis, 13(2): 198–213, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lai:1997:SDI

[LS97a] Ming-Jun Lai and Larry L.

Schumaker. Scattered data interpolation using C^{2} supersplines of degree six. SIAM Journal on Numerical Analysis, 34(3):905–921, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27317.

Lyness:1997:CRP

[LS97b] J. N. Lyness and I. H. Sloan. Cubature rules of prescribed merit. SIAM Journal on Numerical Analysis, 34(2):586–602, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26748.

Lai:1999:APS

[LS99] Ming-Jun Lai and Larry L. Schumaker. On the approximation power of splines on triangulated quadrangulations. SIAM Journal on Numerical Analysis, 36 (1):143–159, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32208.

Larsson:1994:BFE

[LSS94] S. Larsson and J. M. Sanz-Serna. The behavior of finite element solutions of semilinear parabolic problems near stationary points. SIAM Journal on Numerical Analysis, 31(4):1000–1018, August 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1990:EIC

[LSX90] Q. Lin, I. H. Sloan, and R. Xie. Extrapolation of the iterated-collocation method for integral equations of the second kind. SIAM Journal on Numerical Analysis, 27(6): 1535–1541, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1987:NSC

[LSY87] Tien-Yien Li, Tim Sauer, and James A. Yorke. Numerical solution of a class of deficient polynomial systems. SIAM Journal on Numerical Analysis, 24(2): 435–451, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1989:CHE

[LSY89] T. Y. Li, Tim Sauer, and J. A. Yorke. The cheater’s homotopy: an efficient procedure for solving systems of polynomial equations. SIAM Journal on Numerical Analysis, 26(5):1241–1251, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Littell:1997:EAS

[LSZ97] Todd R. Littell, Robert D.

Skeel, and Meiqing Zhang. Error analysis of symplectic multiple time stepping. SIAM Journal on Numerical Analysis, 34(5):1792– 1807, October 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28709.

Loscalzo:1967:SFA

[LT67] Frank R. Loscalzo and Thomas D. Talbot. Spline function approximations for solutions of ordinary differential equations. SIAM Journal on Numerical Analysis, 4 (3):433–445, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lynn:1968:UMD

[LT68] M. Stuart Lynn and William P. Timlake. The use of multiple deflations in the numerical solution of singular systems of equations, with applications to potential theory. SIAM Journal on Numerical Analysis, 5(2):303–322, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lenoir:1988:LFE

[LT88] M. Lenoir and A. Tounsi. The localized finite element method and its application to the two-dimensional seakeeping problem. SIAM Journal on Numerical Analysis, 25(4):729–752, August 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

LeRoux:1989:NSS

[LT89a] Marie-Noëlle Le Roux and Vidar Thomée. Numerical solution of semilinear integrodifferential equations of parabolic type with nonsmooth data. SIAM Journal on Numerical Analysis, 26(6): 1291–1309, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lesaint:1989:AHE

[LT89b] P. Lesaint and R. Touzani.

Approximation of the heat equation in a variable domain with application to the Stefan problem. SIAM Journal on Numerical Analysis, 26(2): 366–379, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Layton:1998:TLM

[LT98] W. Layton and L. Tobiska.

A two-level method with backtracking for the Navier– Stokes equations. SIAM Journal on Numerical Analysis, 35 (5):2035–2054, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30230.

Lin:1991:RVP

[LTW91] Yan Ping Lin, Vidar Thomée, and Lars B. Wahlbin. Ritzvolterra projections to finiteelement spaces and applications to integrodifferential and related equations. SIAM Journal on Numerical Analysis, 28(4):1047–1070, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1995:CCR

[LTW95a] L. Lin, J. B. Temple, and J. Wang. A comparison of convergence rates for Godunov’s method and Glimm’s method in resonant nonlinear systems of conservation laws. SIAM Journal on Numerical Analysis, 32(3):824–840, June 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:1995:SOG

[LTW95b] Long Wei Lin, Blake Temple, and Jing Hua Wang. Suppression of oscillations in Godunov’s method for a resonant non-strictly hyperbolic system. SIAM Journal on Numerical Analysis, 32(3): 841–864, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Langseth:1996:COS

[LTW96] J. O. Langseth, A. Tveito, and R. Winther. On the convergence of operator splitting applied to conservation laws with source terms. SIAM Journal on Numerical Analysis, 33(3):843–863, June 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lucas:1974:EBI

[Luc74] Thomas R. Lucas. Error bounds for interpolating cubic splines under various end conditions. SIAM Journal on Numerical Analysis, 11(3): 569–584, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lucas:1982:AEI

[Luc82] Thomas R. Lucas. Asymptotic expansions for interpolating periodic splines. SIAM Journal on Numerical Analysis, 19(5):1051–1066, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lucier:1985:EBM

[Luc85a] Bradley J. Lucier. Error bounds for the methods of Glimm, Godunov and LeVeque. SIAM Journal on Numerical Analysis, 22(6):1074– 1081, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lucier:1985:SAN

[Luc85b] Bradley J. Lucier. A stable adaptive numerical scheme for hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 22(1):180– 203, February 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Luenberger:1970:CRM

[Lue70] David G. Luenberger. The conjugate residual method for constrained minimization problems. SIAM Journal on Numerical Analysis, 7 (3):390–398, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Luskin:1979:APN

[Lus79a] Mitchell Luskin. An approximation procedure for nonsymmetric, nonlinear hyperbolic systems with integral boundary conditions. SIAM Journal on Numerical Analysis, 16(1):145–164, February 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Luskin:1979:GMN

[Lus79b] Mitchell Luskin. A Galerkin method for nonlinear parabolic equations with nonlinear boundary conditions. SIAM Journal on Numerical Analysis, 16(2):284–299, April 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lustman:1986:TES

[Lus86] Liviu Lustman. The time evolution of spectral discretizations of hyperbolic systems. SIAM Journal on Numerical Analysis, 23(6):1193– 1198, December 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Landriani:1990:EES

[LV90] Giovanni Sacchi Landriani and Hervé Vandeven. Error estimates for the spectral approximation of the nonstationary Stokes problem. SIAM Journal on Numerical Analysis, 27(5):1160– 1186, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1995:FDC

[LVQ95] S. Li and L. Vu-Quoc. Finite difference calculus invariant structure of a class of algorithms for the nonlinear Klein–Gordon equation. SIAM Journal on Numerical Analysis, 32(6):1839– 1875, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liniger:1970:EIM

[LW70] Werner Liniger and Ralph A.

Willoughby. Efficient integration methods for stiff systems of ordinary differential equations. SIAM Journal on Numerical Analysis, 7(1): 47–66, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Laquer:1981:BMQ

[LW81] H. Turner Laquer and Burton Wendroff. Bounds for the model quench front. SIAM Journal on Numerical Analysis, 18(2):225–241, April 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1992:NHS

[LW92] T. Y. Li and Xiao Shen Wang. Nonlinear homotopies for solving deficient polynomial systems with parameters. SIAM Journal on Numerical Analysis, 29(4):1104– 1118, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liska:1998:CSC

[LW98] Richard Liska and Burton Wendroff. Composite schemes for conservation laws. SIAM Journal on Numerical Analysis, 35(6): 2250–2271, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31097.

Lyness:1967:CFC

[Lyn67] J. N. Lyness. The calculation of Fourier coefficients. SIAM Journal on Numerical Analysis, 4(2):301–315, June 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyness:1978:QSPa

[Lyn78a] J. N. Lyness. Quadrature over a simplex: Part 1. A representation for the integrand function. SIAM Journal on Numerical Analysis, 15(1):122– 133, February 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyness:1978:QSPb

[Lyn78b] J. N. Lyness. Quadrature over a simplex: Part 2. A representation for the error functional. SIAM Journal on Numerical Analysis, 15(5): 870–887, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lorentz:1971:BI

[LZ71] G. G. Lorentz and K. L.

Zeller. Birkhoff interpolation. SIAM Journal on Numerical Analysis, 8(1):43–48, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:1992:SEP

[LZC92] T. Y. Li, Zhong Gang Zeng, and Luan Cong. Solving eigenvalue problems of real nonsymmetric matrices with real homotopies. SIAM Journal on Numerical Analysis, 29 (1):229–248, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marchuk:1981:KEV

[MA81] G. I. Marchuk and V. I. Agoshkov. Kinetic equations and variational principles. SIAM Journal on Numerical Analysis, 18(2): 242–261, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ma:1995:NCC

[Ma95] Fu Ming Ma. Numerical calculation of center manifolds for a class of infinitedimensional systems with applications. SIAM Journal on Numerical Analysis, 32(3): 952–968, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ma:1998:CLSa

[Ma98a]  Heping Ma.  Chebyshev–

Legendre spectral viscosity method for nonlinear conservation laws. SIAM Journal on Numerical Analysis, 35(3):869–892, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29390.

Ma:1998:CLSb

[Ma98b]  Heping Ma.  Chebyshev–

Legendre super spectral viscosity method for nonlinear conservation laws. SIAM Journal on Numerical Analysis, 35(3):893–908, June 1998.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29391.

Madsen:1971:PCT

[Mad71] N. K. Madsen. Pointwise convergence of the threedimensional discrete ordinate method. SIAM Journal on Numerical Analysis, 8 (2):266–269, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Madsen:1975:CCD

[Mad75]  Niel K. Madsen.  Con-

vergent centered difference schemes for the discrete ordinate neutron transport equations. SIAM Journal on Numerical Analysis, 12(2): 164–176, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Maeztu:1982:DDA

[Mae82] J. I. Maeztu. Divided differences associated with reversible systems in R^{2}. SIAM Journal on Numerical Analysis, 19(5):1032–1040, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mahar:1982:DAL

[Mah82a] T. J. Mahar. Discrete almostlinear oscillators. SIAM Journal on Numerical Analysis, 19 (2):237–244, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mahar:1982:DCO

[Mah82b] T. J. Mahar. Discrete conservative oscillators: Periodic and asymptotically periodic solutions. SIAM Journal on Numerical Analysis, 19 (2):231–236, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Majda:1984:FTS

[Maj84] George Majda. Filtering techniques for systems of stiff ordinary differential equations. I. SIAM Journal on Numerical Analysis, 21(3): 535–566, June 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Majda:1985:FTS

[Maj85] George Majda. Filtering techniques for systems of stiff ordinary differential equations. II. error estimates. SIAM Journal on Numerical Analysis, 22(6):1116–1134, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Malcolm:1977:CNS

[Mal77] Michael A. Malcolm. On the computation of nonlinear spline functions. SIAM Journal on Numerical Analysis, 14 (2):254–282, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Maliassov:1998:TLP

[Mal98]  Serguei Maliassov.  A

two-level preconditioner for Schrödinger-type singular elliptic operator. SIAM Journal on Numerical Analysis, 35 (1):217–229, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27505.

Mansfield:1971:OAL

[Man71] Lois E. Mansfield. On the optimal approximation of linear functionals in spaces of bivariate functions. SIAM Journal on Numerical Analysis, 8(1): 115–126, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mansfield:1978:ABF

[Man78] Lois Mansfield. Approximation of the boundary in the finite element solution of fourth order problems. SIAM Journal on Numerical Analysis, 15 (3):568–579, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mansfield:1980:SNF

[Man80] Lois Mansfield. On the solution of nonlinear finite element systems. SIAM Journal on Numerical Analysis, 17 (6):752–765, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mangasarian:1981:ISL

[Man81] O. L. Mangasarian. Iterative solution of linear programs. SIAM Journal on Numerical Analysis, 18(4):606– 614, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manteuffel:1982:OPL

[Man82] T. A. Manteuffel. Optimal parameters for linear second-degree stationary iterative methods. SIAM Journal on Numerical Analysis, 19(4): 833–839, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mansfield:1990:CGS

[Man90] Lois Mansfield. On the conjugate gradient solution of the Schur complement system obtained from domain decomposition. SIAM Journal on Numerical Analysis, 27(6):1612– 1620, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mao:1995:STT

[Mao95] De Kang Mao. A shock tracking technique based on conservation in one space dimension. SIAM Journal on Numerical Analysis, 32(5):1677– 1703, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marden:1966:GTB

[Mar66] Morris Marden. A generalization of a theorem of bocher. SIAM Journal on Numerical Analysis, 3(2):269–275, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marti:1978:ACM

[Mar78] J. T. Marti. An algorithm for computing minimum norm solutions of Fredholm integral equations of the first kind. SIAM Journal on Numerical Analysis, 15(6):1071– 1076, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See remarks [HL82].

Martinez:1979:GMB

[Mar79a] José Mario Martínez. Generalization of the methods of Brent and Brown for solving nonlinear simultaneous equations. SIAM Journal on Numerical Analysis, 16(3): 434–448, June 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marwil:1979:CRS

[Mar79b] Earl Marwil. Convergence results for Schubert’s method for solving sparse nonlinear equations. SIAM Journal on Numerical Analysis, 16(4): 588–604, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marti:1983:ELC

[Mar83] J. T. Marti. Evaluation of the least constant in Sobolev’s inequality for H^{1}(0, s). SIAM Journal on Numerical Analysis, 20(6):1239–1242, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marini:1985:IME

[Mar85] Luisa Donatella Marini. An inexpensive method for the evaluation of the solution of the lowest order Raviart– Thomas mixed method. SIAM Journal on Numerical Analysis, 22(3):493–496, June 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lopez-Marcos:1990:DSN

[Mar90a] J. C. López Marcos. A difference scheme for a nonlinear partial integrodifferential equation. SIAM Journal on Numerical Analysis, 27(1): 20–31, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Martinez:1990:FQN

[Mar90b] José Mario Martínez. A family of quasi-Newton methods for nonlinear equations with direct secant updates of matrix factorizations. SIAM Journal on Numerical Analysis, 27(4):1034–1049, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Martinez:1992:FPQ

[Mar92] José Mario Martínez. Fixedpoint quasi-Newton methods. SIAM Journal on Numerical Analysis, 29(5):1413– 1434, October 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Martinez:1994:SSM

[Mar94] José Mario Martínez. SOR-Secant methods. SIAM Journal on Numerical Analysis, 31 (1):217–226, February 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mason:1973:SMN

[Mas73] J. C. Mason. Some methods of near-minimax approximation using Laguerre polynomials. SIAM Journal on Numerical Analysis, 10(3):470–477, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mascagni:1990:BEM

[Mas90]  Michael Mascagni.  The

backward Euler method for numerical solution of the Hodgkin–Huxley equations of nerve conduction. SIAM Journal on Numerical Analysis, 27(4):941–962, August 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Matthews:1972:AOB

[Mat72] J. W. Matthews. An algorithm for obtaining best approximations on sets of positive polynomials. SIAM Journal on Numerical Analysis, 9 (4):734–742, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mattheij:1982:CLB

[Mat82] R. M. M. Mattheij. The conditioning of linear boundary value problems. SIAM Journal on Numerical Analysis, 19(5):963–978, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mathew:1998:UCS

[Mat98] T. P. Mathew. Uniform convergence of the Schwarz alternating method for solving singularly perturbed advectiondiffusion equations. SIAM Journal on Numerical Analysis, 35(4):1663–1683, August 1998. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29648.

Mayo:1984:FSP

[May84] Anita Mayo. The fast solution of Poisson’s and the biharmonic equations on irregular regions. SIAM Journal on Numerical Analysis, 21(2): 285–299, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mayer:1987:CTI

[May87] Günter Mayer. Comparison theorems for iterative methods based on strong splittings. SIAM Journal on Numerical Analysis, 24(1):215– 227, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Makridakis:1997:SDG

[MB97] Ch. G. Makridakis and I. Babuška. On the stability of the discontinuous Galerkin method for the heat equation. SIAM Journal on Numerical Analysis, 34 (1):389–401, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26165.

Masry:1990:TMC

[MC90] Elias Masry and Stamatis Cambanis. Trapezoidal Monte Carlo integration. SIAM Journal on Numerical Analysis, 27(1):225–246, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Morokoff:1993:QMC

[MC93] William J. Morokoff and Russel E. Caflisch. A quasi-Monte Carlo approach to particle simulation of the heat equation. SIAM Journal on Numerical Analysis, 30(6):1558– 1573, December 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McAllister:1966:QUE

[McA66] G. T. McAllister. Quasilinear uniformly elliptic partial differential equations and difference equations. SIAM Journal on Numerical Analysis, 3 (1):13–33, March 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCormick:1982:AIM

[McC82] S. F. McCormick. An algebraic interpretation of multigrid methods. SIAM Journal on Numerical Analysis, 19 (3):548–560, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCormick:1984:MMV

[McC84] S. F. McCormick. Multigrid methods for variational problems: further results. SIAM Journal on Numerical Analysis, 21(2):255–263, April 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCormick:1985:MMV

[McC85] S. F. McCormick. Multigrid methods for variational problems: general theory for the V -cycle. SIAM Journal on Numerical Analysis, 22(4): 634–643, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCormick:1994:MAM

[McC94] Stephen F. McCormick. Multilevel adaptive methods for elliptic eigenproblems: A two-level convergence theory. SIAM Journal on Numerical Analysis, 31(6):1731– 1745, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McKee:1979:CMM

[McK79]  S. McKee.  Cyclic mul-

tistep methods for solving Volterra integro-differential equations. SIAM Journal on Numerical Analysis, 16(1): 106–114, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McLaurin:1974:GCE

[McL74] Johnnie William McLaurin. A general coupled equation approach for solving the biharmonic boundary value problem. SIAM Journal on Numerical Analysis, 11(1): 14–33, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moskow:1999:FDS

[MDH^{+}99] Shari Moskow, Vladimir Druskin, Tarek Habashy, Ping Lee, and Sofia Davydycheva. A finite difference scheme for elliptic equations with rough coefficients using a Cartesian grid nonconforming to interfaces. SIAM Journal on Numerical Analysis, 36(2):442–464, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31854.

Missirlis:1981:CSG

[ME81] Nikolaos M. Missirlis and David J. Evans. On the convergence of some generalized preconditioned iterative methods. SIAM Journal on Numerical Analysis, 18(4): 591–596, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meany:1969:MI

[Mea69] R. K. Meany. A matrix inequality. SIAM Journal on Numerical Analysis, 6(1): 104–107, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meddahi:1998:OIP

[Med98] Salim Meddahi. An optimal iterative process for the Johnson–Nedelec method of coupling boundary and finite elements. SIAM Journal on Numerical Analysis, 35 (4):1393–1415, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30076.

Meek:1980:ISM

[Mee80]  D. Meek.  The inverses

of some matrices deviating slightly from a symmetric, tridiagonal, Toeplitz form. SIAM Journal on Numerical Analysis, 17(1):39–43, February 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meinguet:1983:REA

[Mei83] Jean Meinguet. Refined error analyses of Cholesky factorization. SIAM Journal on Numerical Analysis, 20(6):1243– 1250, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Melhem:1987:DSS

[Mel87] Rami Melhem. Determination of stripe structures for finite element matrices. SIAM Journal on Numerical Analysis, 24(6):1419–1433, December 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Merkes:1966:TEC

[Mer66] E. P. Merkes. On truncation errors for continued fraction computations. SIAM Journal on Numerical Analysis, 3 (3):486–496, September 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meth:1994:SNB

[Met94] Kalman Z. Meth. Stable numerical boundary conditions for Stokes equations. SIAM Journal on Numerical Analysis, 31(5):1336–1351, October 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1968:SNE

[Mey68] Gunter H. Meyer. On solving nonlinear equations with a one-parameter operator imbedding. SIAM Journal on Numerical Analysis, 5 (4):739–752, December 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1971:NMT

[Mey71] Gunter H. Meyer. A numerical method for two-phase Stefan problems. SIAM Journal on Numerical Analysis, 8 (3):555–568, September 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1973:MSP

[Mey73a] Gunter H. Meyer. Multidimensional Stefan problems. SIAM Journal on Numerical Analysis, 10(3):522–538, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1973:NSN

[Mey73b]  H. D. Meyer.  The nu-

merical solution of nonlinear parabolic problems by variational methods. SIAM Journal on Numerical Analysis, 10 (4):700–722, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1976:CAR

[Mey76] R. R. Meyer. On the convergence of algorithms with restart. SIAM Journal on Numerical Analysis, 13(5):696– 704, October 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1981:AML

[Mey81a] Gunter H. Meyer. An analysis of the method of lines for the Reynolds equation in hydrodynamic lubrication. SIAM Journal on Numerical Analysis, 18(1):165–177, February 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1981:MLI

[Mey81b]  Gunter H. Meyer.  The

method of lines and invariant imbedding for elliptic and parabolic free boundary problems. SIAM Journal on Numerical Analysis, 18(1):150– 164, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1987:COM

[Mey87a] Gunter H. Meyer. Continuous orthonormalization for multipoint problems for linear ordinary differential equations. SIAM Journal on Numerical Analysis, 24(6):1288– 1300, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyling:1987:ACC

[Mey87b] R. H. J. Gmelig Meyling. An algorithm for constructing configurations of knots for bivariate B-splines. SIAM Journal on Numerical Analysis, 24(3):706–724, June 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1990:LRF

[MF90] Peter K. Moore and Joseph E.

Flaherty. A local refinement finite-element method for one-dimensional parabolic systems. SIAM Journal on Numerical Analysis, 27(6): 1422–1444, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Martin:1997:MNM

[MF97] Pablo Martín and José M.

Ferrándiz. Multistep numerical methods based on the scheifele G-functions with application to satellite dynamics. SIAM Journal on Numerical Analysis, 34 (1):359–375, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26505.

Martin:1998:IOS

[MF98] Pablo Martín and José M.

Farto. Increasing the order of the SMF method for a special type of problem. SIAM Journal on Numerical Analysis, 35(2):773–777, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30693.

Miller:1966:NIL

[MG66] Max K. Miller and W. T. Guy, Jr. Numerical inversion of the Laplace transform by use of Jacobi polynomials. SIAM Journal on Numerical Analysis, 3(4):624–635, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marcus:1972:AEC

[MG72] Marvin Marcus and William R. Gordon. An analysis of equality in certain matrix inequalities. II. SIAM Journal on Numerical Analysis, 9(1): 130–136, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Molchanov:1982:DME

[MG82] I. N. Molchanov and Ē. F. Galba. Difference methods for elliptic partial differential equations with nonunique solutions. SIAM Journal on Numerical Analysis, 19(3): 531–547, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mas-Gallic:1995:PAL

[MG95]  S. Mas-Gallic.  Particle

approximation of a linear convection-diffusion problem with Neumann boundary conditions. SIAM Journal on Numerical Analysis, 32(4):1098– 1125, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mu:1998:ACN

[MH98] Mo Mu and Yunqing Huang.

An alternating Crank–Nicolson method for decoupling the Ginzburg–Landau equations. SIAM Journal on Numerical Analysis, 35(5):1740– 1761, October 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30309.

Midgley:1966:CSS

[Mid66] J. E. Midgley. Calculation of subdominant solutions of linear differential equations. SIAM Journal on Numerical Analysis, 3(1):56–66, March 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miel:1981:ECL

[Mie81] George J. Miel. Evaluation of complex logarithms and related functions. SIAM Journal on Numerical Analysis, 18(4):744–750, August 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miel:1986:GCM

[Mie86] George Miel. On the Galerkin and collocation methods for a Cauchy singular integral equation. SIAM Journal on Numerical Analysis, 23(1): 135–143, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1966:CCS

[Mil66] G. F. Miller. On the convergence of the Chebyshev series for functions possessing a singularity in the range of representation. SIAM Journal on Numerical Analysis, 3 (3):390–409, September 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1969:CUD

[Mil69] John J. H. Miller. The construction of unbalanced difference operators for parabolic initial boundary value problems. SIAM Journal on Numerical Analysis, 6(3):476– 479, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1976:RAD

[Mil76] Webb Miller. Roundoff analysis by direct comparison of two algorithms. SIAM Journal on Numerical Analysis, 13 (3):382–392, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mills:1979:OEE

[Mil79a] Wendell H. Mills, Jr. Optimal error estimates for the finite element spectral approximation of noncompact operators. SIAM Journal on Numerical Analysis, 16(4):704– 718, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mills:1979:RSC

[Mil79b] Wendell H. Mills, Jr. The resolvent stability condition for spectra convergence with application to the finite element approximation of noncompact operators. SIAM Journal on Numerical Analysis, 16(4): 695–703, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Milaszewicz:1981:CSR

[Mil81a] J. P. Milaszewicz. On criticality and the Stein–Rosenberg theorem. SIAM Journal on Numerical Analysis, 18(3): 559–564, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1981:MFEb

[Mil81b] Keith Miller. Moving finite elements. II. SIAM Journal on Numerical Analysis, 18(6): 1033–1057, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1992:MMS

[Mil92] Keith Miller. On the mass matrix spectrum bounds of wathen and the local moving finite elements of baines.

SIAM Journal on Numerical Analysis, 29(1):89–106, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1997:GMI

[Mil97] Keith Miller. A geometricalmechanical interpretation of gradient-weighted moving finite elements. SIAM Journal on Numerical Analysis, 34(1):67–90, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26088.

Minc:1970:MEP

[Min70] Henryk Minc. On the maximal eigenvector of a positive matrix. SIAM Journal on Numerical Analysis, 7 (3):424–427, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miranker:1971:GAO

[Mir71] W. L. Miranker. Galerkin approximations and the optimization of difference schemes for boundary value problems. SIAM Journal on Numerical Analysis, 8(3):486– 496, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mittelmann:1977:PEF

[Mit77]  H. D. Mittelmann.  On

pointwise estimates for a finite element solution of nonlinear boundary value problems. SIAM Journal on Numerical Analysis, 14(4):773– 778, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mittelmann:1986:PAC

[Mit86] Hans D. Mittelmann. A pseudo-arclength continuation method for nonlinear eigenvalue problems. SIAM Journal on Numerical Analysis, 23(5):1007–1016, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mittelmann:1987:CVI

[Mit87] H. D. Mittelmann. On continuation for variational inequalities. SIAM Journal on Numerical Analysis, 24(6):1374– 1381, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mathon:1977:ASE

[MJ77a] Rudolf Mathon and R. L.

Johnston. The approximate solution of elliptic boundaryvalue problems by fundamental solutions. SIAM Journal on Numerical Analysis, 14 (4):638–650, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1977:SSR

[MJ77b] R. E. Moore and S. T. Jones. Safe starting regions for iterative methods. SIAM Journal on Numerical Analysis, 14(6): 1051–1065, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1980:STA

[MK80] R. E. Moore and J. B. Kioustelidis. A simple test for accuracy of approximate solutions to nonlinear (or linear) systems. SIAM Journal on Numerical Analysis, 17(4): 521–529, August 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Maday:1993:LPV

[MKT93] Yvon Maday, Sidi M. Ould Kaber, and Eitan Tadmor. Legendre pseudospectral viscosity method for nonlinear conservation laws. SIAM Journal on Numerical Analysis, 30(2):321–342, April 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Minerbo:1969:IAI

[ML69] Gerald N. Minerbo and Maurice E. Levy. Inversion of Abel’s integral equation by means of orthogonal polynomials. SIAM Journal on Numerical Analysis, 6(4):598– 616, December 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Morlet:1992:NSF

[ML92] Anne C. Morlet and Jens Lorenz. Numerical solution of a functional equation on a circle. SIAM Journal on Numerical Analysis, 29(6):1741– 1768, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Micchelli:1973:ODS

[MM73] C. A. Micchelli and W. L.

Miranker. Optimal difference schemes for linear initial value problems. SIAM Journal on Numerical Analysis, 10(6):983–1009, December 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Melkman:1979:OEL

[MM79] Avraham A. Melkman and Charles A. Micchelli. Optimal estimation of linear operators in Hilbert spaces from inaccurate data. SIAM Journal on Numerical Analysis, 16(1): 87–105, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1981:MFEa

[MM81] Keith Miller and Robert N.

Miller. Moving finite elements. I. SIAM Journal on Numerical Analysis, 18(6): 1019–1032, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Maitre:1984:MMC

[MM84] Jean-François Maitre and François Musy. Multigrid methods: convergence theory in a variational framework. SIAM Journal on Numerical Analysis, 21(4): 657–671, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Martinez:1994:SMO

[MMD94] V. Martínez, A. Marquina, and R. Donat. Shooting methods for one-dimensional diffusion-absorption problems. SIAM Journal on Numerical Analysis, 31(2): 572–589, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mandel:1988:ATM

[MMR88] Jan Mandel, Steve F. Mc-Cormick, and John W. Ruge. An algebraic theory for multigrid methods for variational problems. SIAM Journal on Numerical Analysis, 25(1): 91–110, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meek:1984:NMB

[MN84] Peter C. Meek and John Norbury. Nonlinear moving boundary problems and a Keller box scheme. SIAM Journal on Numerical Analysis, 21(5):883–893, October 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mandel:1987:CFE

[MN87] Jan Mandel and Jindřich Nečas. Convergence of finite elements for transonic potential flows. SIAM Journal on Numerical Analysis, 24(5): 985–996, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manteuffel:1993:OEP

[MO93] Thomas Manteuffel and James Otto. Optimal equivalent preconditioners. SIAM Journal on Numerical Analysis, 30(3): 790–812, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Monk:1986:EEN

[Mon86] Peter Monk. Error estimates for a numerical method for an ill-posed Cauchy problem for the heat equation. SIAM Journal on Numerical Analysis, 23(6):1155–1172, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Monk:1987:MFE

[Mon87] Peter Monk. A mixed finite element method for the biharmonic equation. SIAM Journal on Numerical Analysis, 24(4):737–749, August 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Monk:1991:MMA

[Mon91] Peter B. Monk. A mixed method for approximating Maxwell’s equations. SIAM Journal on Numerical Analysis, 28(6):1610–1634, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Monk:1992:AFE

[Mon92] Peter Monk. Analysis of a finite element method for Maxwell’s equations. SIAM Journal on Numerical Analysis, 29(3):714–729, June 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1977:TES

[Moo77] R. E. Moore. A test for existence of solutions to nonlinear systems. SIAM Journal on Numerical Analysis, 14 (4):611–615, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1978:CTC

[Moo78] R. E. Moore. A computational test for convergence of iterative methods for nonlinear systems. SIAM Journal on Numerical Analysis, 15(6): 1194–1196, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1994:PEE

[Moo94] Peter K. Moore. A Posteriori error estimation with finite element semi- and fullydiscrete methods for nonlinear parabolic equations in one space dimension. SIAM Journal on Numerical Analysis, 31(1):149–169, February 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1996:CPI

[Moo96] Gerald Moore. Computation and parametrisation of invariant curves and tori. SIAM Journal on Numerical Analysis, 33(6):2333– 2358, December 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26288.

Moore:1999:FDM

[Moo99] Peter K. Moore. Finite difference methods and spatial A posteriori error estimates for solving parabolic equations in three space dimensions on grids with irregular nodes. SIAM Journal on Numerical Analysis, 36 (4):1044–1064, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32207.

Morrison:1968:OLS

[Mor68] David D. Morrison. Optimization by least squares. SIAM Journal on Numerical Analysis, 5(1):83–88, March 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

More:1971:GCN

[Mor71] Jorge J. More. Global convergence of Newton–Gauss– Seidel methods. SIAM Journal on Numerical Analysis, 8 (2):325–336, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

More:1972:NGM

[Mor72] Jorge J. More. Nonlinear generalizations of matrix diagonal dominance with application to Gauss–Seidel iterations. SIAM Journal on Numerical Analysis, 9(2): 357–378, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Morlet:1995:CSM

[Mor95] Anne C. Morlet. Convergence of the sinc method for a fourth-order ordinary differential equation with an application. SIAM Journal on Numerical Analysis, 32(5):1475– 1503, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moret:1997:NSC

[Mor97] Igor Moret. A note on the superlinear convergence of GMRES. SIAM Journal on Numerical Analysis, 34(2):513–516, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25979.

Morton:1998:AFV

[Mor98] K. W. Morton. On the analysis of finite volume methods for evolutionary problems. SIAM Journal on Numerical Analysis, 35(6): 2195–2222, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31696.

Moss:1983:TDO

[Mos83] William F. Moss. The twodimensional oscillating airfoil: a new implementation of the Galerkin method. SIAM Journal on Numerical Analysis, 20(2):391–399, April 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Motzkin:1966:ASD

[Mot66] Theodore S. Motzkin. Approximation in the sense of a deviator integral. SIAM Journal on Numerical Analysis, 3 (2):276–286, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moursund:1965:SCA

[Mou65] D. G. Moursund. Some computational aspects of the uniform approximation of a function and its derivative. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):464–472, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Moursund:1966:CAU

[Mou66] David G. Moursund. Chebyshev approximation using a generalized weight function. SIAM Journal on Numerical Analysis, 3(3):435– 450, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moursund:1968:CAC

[Mou68] D. G. Moursund. Computational aspects of Chebyshev approximation using a generalized weight function. SIAM Journal on Numerical Analysis, 5(1):126–137, March 1968. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Moler:1968:BEE

[MP68] C. B. Moler and L. E. Payne.

Bounds for eigenvalues and eigenvectors of symmetric operators. SIAM Journal on Numerical Analysis, 5(1): 64–70, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1977:CPM

[MP77] Carl D. Meyer, Jr. and R. J.

Plemmons. Convergent powers of a matrix with applications to iterative methods for singular linear systems. SIAM Journal on Numerical Analysis, 14(4):699–705, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Morrow:1978:CAC

[MP78a] C. R. Morrow and T. N. L.

Patterson. Construction of algebraic cubature rules using polynomial ideal theory. SIAM Journal on Numerical Analysis, 15(5):953– 976, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mukai:1978:UAA

[MP78b] H. Mukai and E. Polak. On the use of approximations in algorithms for optimization problems with equality and inequality constraints. SIAM Journal on Numerical Analysis, 15(4):674–693, August 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Morrow:1985:CAC

[MP85] C. R. Morrow and T. N. L. Patterson. The construction of algebraic cubature formulae by the distribution of nodes along selected lines. SIAM Journal on Numerical Analysis, 22(6):1178– 1190, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manteuffel:1990:PBC

[MP90] Thomas A. Manteuffel and Seymour V. Parter. Preconditioning and boundary conditions. SIAM Journal on Numerical Analysis, 27(3): 656–694, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Milstein:1998:BIM

[MPS98] G. N. Milstein, E. Platen, and H. Schurz. Balanced implicit methods for stiff stochastic systems. SIAM Journal on Numerical Analysis, 35(3):1010–1019, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27352.

Maday:1982:SPS

[MQ82a] Y. Maday and A. Quarteroni. Spectral and pseudospectral approximations of the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 19(4):761– 780, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1982:SIT

[MQ82b] R. E. Moore and L. Qi. A successive interval test for nonlinear systems. SIAM Journal on Numerical Analysis, 19(4): 845–850, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McLachlan:1998:NIP

[MQT98] R. I. McLachlan, G. R. W.

Quispel, and G. S. Turner. Numerical integrators that preserve symmetries and reversing symmetries. SIAM Journal on Numerical Analysis, 35(2):586–599, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29580.

Mantel:1977:AIP

[MR77] Francis Mantel and Philip Rabinowitz. The application of integer programming to the computation of fully symmetric integration formulas in two and three dimensions. SIAM Journal on Numerical Analysis, 14(3):391–425, June 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCormick:1982:MMV

[MR82] S. F. McCormick and J. W. Ruge. Multigrid methods for variational problems. SIAM Journal on Numerical Analysis, 19(5):924–929, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Markowich:1983:NSC

[MR83] Peter Markowich and Michael Renardy. The numerical solution of a class of quasilinear parabolic Volterra equations arising in polymer rheology. SIAM Journal on Numerical Analysis, 20(5):890– 908, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Markowich:1984:LWM

[MR84] Peter Markowich and Michael Renardy. Lax–Wendroff methods for hyperbolic history value problems. SIAM Journal on Numerical Analysis, 21(1):24–51, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See corrigenda [MR85].

Markowich:1985:CLW

[MR85] Peter Markowich and Michael Renardy. Corrigenda: “Lax– Wendroff Methods for Hyperbolic History Value Problems” [SIAM J. Numer. Anal. 21 (1984), no. 1, 24–51]. SIAM Journal on Numerical Analysis, 22(1):204, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [MR84].

Manteuffel:1998:LSF

[MR98] Thomas A. Manteuffel and Klaus J. Ressel. Least-squares finite-element solution of the neutron transport equation in diffusive regimes. SIAM Journal on Numerical Analysis, 35(2):806–835, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29970.

Ma:1996:GGQ

[MRW96] J. Ma, V. Rokhlin, and S. Wandzura. Generalized Gaussian quadrature rules for systems of arbitrary functions. SIAM Journal on Numerical Analysis, 33(3): 971–996, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meir:1965:MRQ

[MS65a] A. Meir and A. Sharma.

On the method of Romberg quadrature. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 250–258, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Morton:1965:SFD

[MS65b] K. W. Morton and S. Schechter. On the stability of finite difference matrices. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1): 119–128, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Moursund:1965:BCA

[MS65c] D. G. Moursund and A. H. Stroud. The best çebyšev approximation to a function and its derivative on n + 2 points. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):15– 23, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Meir:1966:SAF

[MS66] A. Meir and A. Sharma. Simultaneous approximation of a function and its derivatives. SIAM Journal on Numerical Analysis, 3(4):553– 563, December 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meir:1968:EOF

[MS68] A. Meir and A. Sharma. An extension of Obreshkov’s formula. SIAM Journal on Numerical Analysis, 5(3):488– 490, September 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mangasarian:1973:BSF

[MS73a] O. L. Mangasarian and L. L.

Schumaker. Best summation formulae and discrete splines. SIAM Journal on Numerical Analysis, 10(3):448–459, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCarthy:1973:OCM

[MS73b] Charles McCarthy and Gilbert Strang. Optimal conditioning of matrices. SIAM Journal on Numerical Analysis, 10 (2):370–388, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meir:1973:LIS

[MS73c] A. Meir and A. Sharma. Lacunary interpolation by splines. SIAM Journal on Numerical Analysis, 10(3):433–442, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moler:1973:AGM

[MS73d] C. B. Moler and G. W.

Stewart. An algorithm for generalized matrix eigenvalue problems. SIAM Journal on Numerical Analysis, 10(2): 241–256, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCormick:1976:UEE

[MS76] G. F. McCormick and V. A. Sposito. Using the L_{2}- estimator in L_{1}-estimation. SIAM Journal on Numerical Analysis, 13(3):337–343, June 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moore:1980:CTP

[MS80] G. Moore and A. Spence. The calculation of turning points of nonlinear equations. SIAM Journal on Numerical Analysis, 17(4):567–576, August 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McKee:1983:PIM

[MS83] S. McKee and A. Stokes. Product integration methods for the nonlinear Basset equation. SIAM Journal on Numerical Analysis, 20(1):143– 160, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meyer:1988:DPE

[MS88] Carl D. Meyer and G. W. Stewart. Derivatives and perturbations of eigenvectors.

SIAM Journal on Numerical Analysis, 25(3):679–691, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

MartinezGamba:1989:STB

[MS89] Irene Martínez Gamba and Maria Cristina J. Squeff. Simulation of the transient behavior of a one-dimensional semiconductor device. II. SIAM Journal on Numerical Analysis, 26(3):539–552, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mehrotra:1991:MAC

[MS91] Sanjay Mehrotra and Jie Sun.

A method of analytic centers for quadratically constrained convex quadratic programs. SIAM Journal on Numerical Analysis, 28(2): 529–544, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Monk:1994:CAY

[MS94] Peter Monk and Endre Süli.

A convergence analysis of Yee’s scheme on nonuniform grids. SIAM Journal on Numerical Analysis, 31(2): 393–412, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Monegato:1997:NSG

[MS97a] G. Monegato and I. H. Sloan. Numerical solution of the generalized airfoil equation for an airfoil with a flap. SIAM Journal on Numerical Analysis, 34(6): 2288–2305, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29505.

Munoz-Sola:1997:PLT

[MS97b] Rafael Muñoz-Sola. Polynomial liftings on a tetrahedron and applications to the h-p version of the finite element method in three dimensions. SIAM Journal on Numerical Analysis, 34 (1):282–314, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26755.

Melenk:1998:FRD

[MS98] J. M. Melenk and C. Schwab. hp FEM for reaction-diffusion equations I: Robust exponential convergence. SIAM Journal on Numerical Analysis, 35 (4):1520–1557, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31760.

Meir:1999:ANA

[MS99a] A. J. Meir and Paul G.

Schmidt. Analysis and numerical approximation of a stationary MHD flow problem with nonideal boundary. SIAM Journal on Numerical Analysis, 36(4):1304– 1332, August 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32615.

Monk:1999:ACF

[MS99b] Peter Monk and Endre Süli.

The adaptive computation of far-field patterns by a posteriori error estimation of linear functionals. SIAM Journal on Numerical Analysis, 36 (1):251–274, February 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31517.

Mund:1999:ATL

[MS99c] Patrick Mund and Ernst P.

Stephan. An adaptive twolevel method for the coupling of nonlinear FEM-BEM equations. SIAM Journal on Numerical Analysis, 36 (4):1001–1021, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31649.

Manohar:1988:NNT

[MST88] R. Manohar, J. W. Stephenson, and G. M. Trojan. Ninenode triangular cubic elements. SIAM Journal on Numerical Analysis, 25(6):1432– 1441, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manteuffel:1994:DSV

[MSW94] Thomas A. Manteuffel, Michael Steuerwalt, and Burton Wendroff. Dedication to Seymour V. Parter on the occasion of his 65th birthday. SIAM Journal on Numerical Analysis, 31(6):vii–ix, December 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Morgan:1995:PDB

[MSW95] Alexander P. Morgan, Andrew J. Sommese, and Charles W. Wampler. A product-decomposition bound for Bezout numbers. SIAM Journal on Numerical Analysis, 32(4):1308–1325, August 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moursund:1968:OSV

[MT68a] D. G. Moursund and G. D. Taylor. Optimal starting values for the Newton–Raphson calculation of inverses of certain functions. SIAM Journal on Numerical Analysis, 5(1):

138–150, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Moursund:1968:URA

[MT68b] D. G. Moursund and G. D.

Taylor. Uniform rational approximation using a generalized weight function. SIAM Journal on Numerical Analysis, 5(4):882–889, December 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

McCarthy:1975:CPE

[MT75] Mary Anne McCarthy and R. A. Tapia. Computable a posteriori L_{∞}-error bounds for the approximate solution of two-point boundary value problems. SIAM Journal on Numerical Analysis, 12 (6):919–937, December 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Maday:1989:ASV

[MT89a] Yvon Maday and Eitan Tadmor. Analysis of the spectral vanishing viscosity method for periodic conservation laws. SIAM Journal on Numerical Analysis, 26(4): 854–870, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marion:1989:NGM

[MT89b] Martine Marion and Roger Temam. Nonlinear Galerkin methods. SIAM Journal on Numerical Analysis, 26 (5):1139–1157, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Milstein:1997:NMW

[MT97] G. N. Milstein and M. V. Tret′yakov. Numerical methods in the weak sense for stochastic differential equations with small noise. SIAM Journal on Numerical Analysis, 34(6):2142–2167, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27896.

Mandel:1999:SSM

[MTF99] Jan Mandel, Radek Tezaur, and Charbel Farhat. A scalable substructuring method by Lagrange multipliers for plate bending problems. SIAM Journal on Numerical Analysis, 36(5):1370– 1391, October 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28989.

Murad:1996:ABS

[MTL96] Márcio A. Murad, Vidar Thomée, and Abimael F. D.

Loula. Asymptotic behavior of semidiscrete finite-element approximations of biot’s consolidation problem. SIAM Journal on Numerical Analysis, 33(3):1065–1083, June 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1999:REC

[MTTW99] Gary L. Miller, Dafna Talmor, Shang-Hua Teng, and Noel Walkington. On the radiusedge condition in the control volume method. SIAM Journal on Numerical Analysis, 36(6):1690–1708, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31185.

Muller:1984:CHM

[Mül84] Rolf E. Müller. Computations of holomorphic multiparameter eigenvalue problems. SIAM Journal on Numerical Analysis, 21(2): 373–387, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Munteanu:1973:GSS

[Mun73] Marie-Jeanne Munteanu. Generalized smoothing spline functions for operators. SIAM Journal on Numerical Analysis, 10(1):28–34, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Munz:1994:GTS

[Mun94] C. D. Munz. On Godunovtype schemes for Lagrangian gas dynamics. SIAM Journal on Numerical Analysis, 31(1): 17–42, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Murty:1971:BAC

[Mur71] V. N. Murty. Best approximation with Chebyshev polynomials. SIAM Journal on Numerical Analysis, 8(4):717– 721, December 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Murota:1982:GCM

[Mur82] Kazuo Murota. Global convergence of a modified Newton iteration for algebraic equations. SIAM Journal on Numerical Analysis, 19(4): 793–799, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Murua:1997:OCP

[Mur97]  Ander Murua.  On or-

der conditions for partitioned symplectic methods. SIAM Journal on Numerical Analysis, 34(6):2204– 2211, December 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28516.

Monk:1999:EEN

[MV99] P. B. Monk and O. Vacus. Error estimates for a numerical scheme for ferromagnetic problems. SIAM Journal on Numerical Analysis, 36(3):696–718, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32422.

Magenes:1989:TNR

[MVV89] E. Magenes, C. Verdi, and A. Visintin. Theoretical and numerical results on the twophase Stefan problem. SIAM Journal on Numerical Analysis, 26(6):1425–1438, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manteuffel:1986:ENS

[MW86] Thomas A. Manteuffel and Andrew B. White, Jr. On the efficient numerical solution of systems of second order boundary value problems. SIAM Journal on Numerical Analysis, 23(5):996– 1006, October 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manteuffel:1992:CDS

[MW92] Thomas A. Manteuffel and Andrew B. White, Jr. A calculus of difference schemes for the solution of boundaryvalue problems on irregular meshes. SIAM Journal on Numerical Analysis, 29 (5):1321–1346, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miller:1994:TMF

[MW94] J. J. H. Miller and S. Wang. A tetrahedral mixed finite element method for the stationary semiconductor continuity equations. SIAM Journal on Numerical Analysis, 31(1): 196–216, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ma:1995:ANO

[MW95] Ling Ma and Noel J. Walkington. On algorithms for nonconvex optimization in the calculus of variations. SIAM Journal on Numerical Analysis, 32(3):900–923, June 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marion:1995:EEN

[MX95] Martine Marion and Jinchao Xu. Error estimates on a new nonlinear Galerkin method based on two-grid finite elements. SIAM Journal on Numerical Analysis, 32(4):1170– 1184, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Myerson:1991:IS

[Mye91] G. Myerson. On ignoring the singularity. SIAM Journal on Numerical Analysis, 28(6): 1803–1807, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Naevdal:1991:CPT

[Næv91] Geir Nævdal. A comment on Parrott’s theorem. SIAM Journal on Numerical Analysis, 28(5):1517–1522, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nakashima:1991:EPR

[Nak91] Masaharu Nakashima. Embedded pseudo-Runge–Kutta methods. SIAM Journal on Numerical Analysis, 28(6): 1790–1802, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nanda:1985:DEA

[Nan85] T. Nanda. Differential equations and the QR algorithm. SIAM Journal on Numerical Analysis, 22(2): 310–321, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nashed:1970:SDS

[Nas70] M. Z. Nashed. Steepest descent for singular linear operator equations. SIAM Journal on Numerical Analysis, 7 (3):358–362, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nash:1984:NTM

[Nas84] Stephen G. Nash. Newtontype minimization via the Lánczos method. SIAM Journal on Numerical Analysis, 21(4):770–788, August 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Navot:1965:EMF

[Nav65]  I. Navot.  The Euler–

Maclaurin functional for functions with a complex singularity near the range of integration. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):259– 264, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Nazareth:1979:RBB

[Naz79] Larry Nazareth. A relationship between the BFGS and conjugate gradient algorithms and its implications for new algorithms. SIAM Journal on Numerical Analysis, 16(5): 794–800, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nazareth:1986:ADP

[Naz86] J. L. Nazareth. Analogues of Dixon’s and Powell’s theorems for unconstrained minimization with inexact line searches. SIAM Journal on Numerical Analysis, 23(1): 170–177, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Netravali:1974:SAS

[NdF74] Arun N. Netravali and Rui J. P. de Figueiredo. Spline approximation to the solution of the linear Fredholm integral equation of the second kind. SIAM Journal on Numerical Analysis, 11(3):538–549, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nelson:1977:CEC

[NE77] Paul Nelson and Ira T. Elder.

Calculation of eigenfunctions in the context of integrationto-blowup. SIAM Journal on Numerical Analysis, 14(1): 124–136, March 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neamtu:1992:MDD

[Nea92] M. Neamtu. Multivariate divided differences I: Basic properties. SIAM Journal on Numerical Analysis, 29 (5):1435–1445, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nedelec:1978:CEC

[Ned78] J. C. Nedelec. Computation of eddy currents on a surface in R^{3} by finite element methods. SIAM Journal on Numerical Analysis, 15 (3):580–594, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nelson:1973:CDO

[Nel73] Paul Nelson, Jr. Convergence of the discrete-ordinates method for anisotropically scattering multiplying particles in a subcritical slab. SIAM Journal on Numerical Analysis, 10(1):175– 181, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nelson:1986:CII

[Nel86] Paul Nelson. Convergence of inner iterations for finitedifference approximations to the linear transport equation. SIAM Journal on Numerical Analysis, 23(5):1017– 1022, October 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neumaier:1985:RII

[Neu85] A. Neumaier. Residual inverse iteration for the nonlinear eigenvalue problem. SIAM Journal on Numerical Analysis, 22(5):914–923, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neumaier:1987:OLI

[Neu87] A. Neumaier. Overestimation in linear interval equations. SIAM Journal on Numerical Analysis, 24(1):207– 214, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neubauer:1988:PPC

[Neu88] Andreas Neubauer. An a posteriori parameter choice for Tikhonov regularization in Hilbert scales leading to optimal convergence rates. SIAM Journal on Numerical Analysis, 25(6):1313–1326, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neubauer:1997:CSR

[Neu97]  Andreas Neubauer.  On

converse and saturation results for Tikhonov regularization of linear ill-posed problems. SIAM Journal on Numerical Analysis, 34(2):517–527, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25392.

Neuss:1998:CCU

[Neu98] Nicolas Neuss. V -cycle convergence with unsymmetric smoothers and application to an anisotropic model problem. SIAM Journal on Numerical Analysis, 35 (3):1201–1212, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31084.

Nickel:1967:ERP

[Nic67] Karl L. E. Nickel. Extension of a recent paper by Fox, Henrici and Moler on eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4(4):483– 488, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic). URL https://www.
jstor.org/stable/2949713.

Nicolaides:1972:CFE

[Nic72] R. A. Nicolaides. On a class of finite elements generated by Lagrange interpolation. SIAM Journal on Numerical Analysis, 9(3):435– 445, September 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nichols:1973:CTS

[Nic73a] Nancy K. Nichols. On the convergence of two-stage iterative processes for solving linear equations. SIAM Journal on Numerical Analysis, 10 (3):460–469, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nicolaides:1973:CFE

[Nic73b] R. A. Nicolaides. On a class of finite elements generated by Lagrange interpolation. II. SIAM Journal on Numerical Analysis, 10(1): 182–189, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nickel:1981:GCB

[Nic81] Karl L. Nickel. A globally convergent ball Newton method. SIAM Journal on Numerical Analysis, 18 (6):988–1003, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nicolaides:1982:EUA

[Nic82] R. A. Nicolaides. Existence, uniqueness and approximation for generalized saddle point problems. SIAM Journal on Numerical Analysis, 19 (2):349–357, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nicolaides:1987:DCG

[Nic87] R. A. Nicolaides. Deflation of conjugate gradients with applications to boundary value problems. SIAM Journal on Numerical Analysis, 24(2): 355–365, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nicolaides:1992:ACM

[Nic92a] R. A. Nicolaides. Analysis and convergence of the MAC scheme I. the linear problem. SIAM Journal on Numerical Analysis, 29(6):1579– 1591, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nicolaides:1992:DDP

[Nic92b] R. A. Nicolaides. Direct discretization of planar divcurl problems. SIAM Journal on Numerical Analysis, 29(1): 32–56, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nielson:1974:MSI

[Nie74] Gregory M. Nielson. Multivariate smoothing and interpolating splines. SIAM Journal on Numerical Analysis, 11 (2):435–446, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Niethammer:1979:DSA

[Nie79] Wilhelm Niethammer. On different splittings and the associated iteration methods. SIAM Journal on Numerical Analysis, 16(2): 186–200, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nielson:1980:MNI

[Nie80] G. Nielson. Minimum norm interpolation in triangles. SIAM Journal on Numerical Analysis, 17(1):44–62, February 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nielsen:1999:FED

[Nie99] Bjørn Fredrik Nielsen. Finite element discretizations of elliptic problems in the presence of arbitrarily small ellipticity: An error analysis. SIAM Journal on Numerical Analysis, 36(2): 368–392, April 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31943.

Nitsche:1989:FEM

[Nit89] Joachim A. Nitsche. Finiteelement methods for conformal mappings. SIAM Journal on Numerical Analysis, 26(6): 1525–1533, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ning:1997:CSM

[NK97] S. Ning and R. B. Kearfott. A comparison of some methods for solving linear interval equations. SIAM Journal on Numerical Analysis, 34 (4):1289–1305, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27099.

Nocedal:1985:PHU

[NO85] Jorge Nocedal and Michael L. Overton. Projected Hessian updating algorithms for nonlinearly constrained optimization. SIAM Journal on Numerical Analysis, 22(5):821– 850, October 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Noble:1966:MCG

[Nob66] Ben Noble. A method for computing the generalized inverse of a matrix. SIAM Journal on Numerical Analysis, 3 (4):582–584, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Norsett:1978:RPA

[Nor78] Syvert P. Norsett. Restricted Padé approximations to the exponential function. SIAM Journal on Numerical Analysis, 15(5):1008–1029, October 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Notaris:1991:SNF

[Not91] Sotirios E. Notaris. Some new formulae for the Stieltjes polynomials relative to classical weight functions. SIAM Journal on Numerical Analysis, 28(4):1196–1206, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nassif:1981:SDF

[NP81] Nabil R. Nassif and Fabrice Pini. Semi-discrete and fully discrete finite-element methods with penalty for the numerical solution of the waterhammer problem. SIAM Journal on Numerical Analysis, 18(1):111–128, February 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nochetto:1993:FDA

[NPV93] R. H. Nochetto, M. Paolini, and C. Verdi. A fully discrete adaptive nonlinear Chernoff formula. SIAM Journal on Numerical Analysis, 30(4): 991–1014, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nickel:1972:TCN

[NR72] K. Nickel and K. Ritter. Termination criterion and numerical convergence. SIAM Journal on Numerical Analysis, 9 (2):277–283, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nkaoua:1993:NTD

[NS93] T. N’Kaoua and R. Sentis. A new time discretization for the radiative transfer equations: Analysis and comparison with the classical discretization. SIAM Journal on Numerical Analysis, 30(3): 733–748, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nessyahu:1992:CRA

[NT92] Haim Nessyahu and Eitan Tadmor. The convergence rate of approximate solutions for nonlinear scalar conservation laws. SIAM Journal on Numerical Analysis, 29(6): 1505–1519, December 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nessyahu:1994:CRA

[NT94] Haim Nessyahu and Tamir Tassa. Convergence rate of approximate solutions to conservation laws with initial rarefactions. SIAM Journal on Numerical Analysis, 31(3):

628–654, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nooijen:1990:SNI

[NtB90] M. Nooijen, G. te Velde, and E. J. Baerends. Symmetric numerical integration formulas for regular polygons. SIAM Journal on Numerical Analysis, 27(1):198– 218, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nessyahu:1994:CRG

[NTT94] Haim Nessyahu, Eitan Tadmor, and Tamir Tassa. The convergence rate of Godunov type schemes. SIAM Journal on Numerical Analysis, 31 (1):1–16, February 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nelson:1979:TPO

[NV79] Paul Nelson, Jr. and H. Dean Victory, Jr. Theoretical properties of one-dimensional discrete ordinates. SIAM Journal on Numerical Analysis, 16 (2):270–283, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neta:1983:NFO

[NV83] Beny Neta and H. D. Victory, Jr. A new fourth-order finitedifference method for solving discrete-ordinates slab transport equations. SIAM Journal on Numerical Analysis, 20(1):94–105, February 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nochetto:1988:ADP

[NV88] Ricardo H. Nochetto and Claudio Verdi. Approximation of degenerate parabolic problems using numerical integration. SIAM Journal on Numerical Analysis, 25(4): 784–814, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Naik:1993:IRM

[NV93] Naomi H. Naik and John Van Rosendale. The improved robustness of multigrid elliptic solvers based on multiple semicoarsened grids. SIAM Journal on Numerical Analysis, 30(1):215–229, February 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nochetto:1997:CPS

[NV97] Ricardo H. Nochetto and Claudio Verdi. Convergence past singularities for a fully discrete approximation of curvature-driven interfaces. SIAM Journal on Numerical Analysis, 34(2): 490–512, April 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26952.

Nedelec:1989:HPE

[NW89] J.-C. Nédélec and S. Wolf.

Homogenization of the problem of eddy currents in a transformer core. SIAM Journal on Numerical Analysis, 26(6):1407–1424, December 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nicolaides:1997:CST

[NW97] Roy A. Nicolaides and Xiaonan Wu. Covolume solutions of three-dimensional div-curl equations. SIAM Journal on Numerical Analysis, 34(6): 2195–2203, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27728.

Obi:1990:EAL

[Obi90] Wilson C. Obi. Error analysis of a Laplace transform inversion procedure. SIAM Journal on Numerical Analysis, 27 (2):457–469, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1984:HRS

[OC84] Stanley Osher and Sukumar Chakravarthy. High resolution schemes and the entropy condition. SIAM Journal on Numerical Analysis, 21(5): 955–984, October 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Oettli:1965:SSL

[Oet65] W. Oettli. On the solution set of a linear system with inaccurate coefficients. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1): 115–118, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Oliveira:1974:IAT

[Oli74]  F. Aleixo Oliveira.  In-

terval analysis and twopoint boundary value problems. SIAM Journal on Numerical Analysis, 11(2): 382–391, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Olson:1977:TPS

[Ols77] Andrew M. Olson. A twopoint series method for twopoint boundary value problems: Theoretical foundation. SIAM Journal on Numerical Analysis, 14(1):2–18, March 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Olver:1965:ASS

[Olv65]  F. W. J. Olver.  On

the asymptotic solutions of second-order differential equations having an irregular singularity of rank one, with an application to Whittaker functions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 225–243, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Olver:1978:NAE

[Olv78] F. W. J. Olver. A new approach to error arithmetic. SIAM Journal on Numerical Analysis, 15(2): 368–393, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Omodei:1977:NSR

[Omo77] Bernard J. Omodei. On the numerical stability of the Rayleigh–Ritz method. SIAM Journal on Numerical Analysis, 14(6):1151–1171, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Oden:1997:PDD

[OPF97] J. T. Oden, Abani Patra, and Yusheng Feng. Parallel domain decomposition solver for adaptive hp finite element methods. SIAM Journal on Numerical Analysis, 34(6): 2090–2118, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27888.

Oettli:1965:ASL

[OPW65] W. Oettli, W. Prager, and J. H. Wilkinson. Admissible solutions of linear systems with not sharply defined coefficients. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 291–299, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Ortega:1966:DDO

[OR66a] James M. Ortega and Werner C. Rheinboldt. On discretization and differentiation of operators with application to Newton’s method. SIAM Journal on Numerical Analysis, 3(1): 143–156, March 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ortega:1966:NDE

[OR66b] James M. Ortega and Maxine L. Rockoff. Nonlinear difference equations and Gauss– Seidel type iterative methods. SIAM Journal on Numerical Analysis, 3(3):497– 513, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ortega:1967:MIN

[OR67] James M. Ortega and Werner C. Rheinboldt. Monotone iterations for nonlinear equations with application to Gauss– Seidel methods. SIAM Journal on Numerical Analysis, 4 (2):171–190, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ortega:1972:GCR

[OR72] James M. Ortega and Werner C.

Rheinboldt. A general convergence result for unconstrained minimization methods. SIAM Journal on Numerical Analysis, 9(1):40–43, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Oden:1976:MFE

[OR76] J. T. Oden and J. N. Reddy.

On mixed finite element approximations. SIAM Journal on Numerical Analysis, 13 (3):393–404, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osborne:1986:RTS

[OR86] M. R. Osborne and R. D. Russell. The Riccati transformation in the solution of boundary value problems. SIAM Journal on Numerical Analysis, 23(5):1023–1033, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1990:FOI

[OR90] Stanley Osher and Leonid I. Rudin. Feature-oriented image enhancement using shock filters. SIAM Journal on Numerical Analysis, 27(4):919– 940, August 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ostermann:1993:RMP

[OR93] A. Ostermann and M. Roche. Rosenbrock methods for partial differential equations and fractional orders of convergence. SIAM Journal on Numerical Analysis, 30(4):1084– 1098, September 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ortiz:1969:TM

[Ort69]  Eduardo L. Ortiz.  The

tau method. SIAM Journal on Numerical Analysis, 6 (3):480–492, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ortega:1973:SDE

[Ort73] James M. Ortega. Stability of difference equations and convergence of iterative processes. SIAM Journal on Numerical Analysis, 10(2): 268–282, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Olver:1965:EBAb

[OS65] F. W. J. Olver and F. Stenger.

Error bounds for asymptotic solutions of secondorder differential equations having an irregular singularity of arbitrary rank. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):244–249, ???? 1965. CODEN ???? ISSN 0887-459X (print), 1095-7170 (electronic). URL http:

//links.jstor.org/sici?
sici=0887-459X(1965)2:2<
244:EBFASO>2.0.CO%3B2-N.

OSullivan:1990:CCM

[O’S90] Finbarr O’Sullivan. Convergence characteristics of methods of regularization estimators for nonlinear operator equations. SIAM Journal on Numerical Analysis, 27(6): 1635–1649, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1991:HOE

[OS91] Stanley Osher and Chi-Wang Shu. High-order essentially nonoscillatory schemes for Hamilton–Jacobi equations. SIAM Journal on Numerical Analysis, 28(4): 907–922, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Okunbor:1992:ERK

[OS92] Daniel Okunbor and Robert D. Skeel. An explicit Runge– Kutta–Nyström method is canonical if and only if its adjoint is explicit. SIAM Journal on Numerical Analysis, 29 (2):521–527, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osada:1990:CAM

[Osa90] Naoki Osada. A convergence acceleration method for some logarithmically convergent sequences. SIAM Journal on Numerical Analysis, 27 (1):178–189, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osborne:1965:SLS

[Osb65] E. E. Osborne. Smallest least squares solutions of linear equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 300–307, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Osborn:1967:AEC

[Osb67] John E. Osborn. Approximation of the eigenvalues of a class of unbounded, nonselfadjoint operators. SIAM Journal on Numerical Analysis, 4(1):45–54, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osborne:1975:SSN

[Osb75] M. R. Osborne. Some special nonlinear least squares problems. SIAM Journal on Numerical Analysis, 12(4):571– 592, September 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osborn:1976:AEN

[Osb76] John E. Osborn. Approximation of the eigenvalues of a nonselfadjoint operator arising in the study of the stability of stationary solutions of the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 13(2): 185–197, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osborne:1979:SMS

[Osb79] M. R. Osborne. The stabilized march is stable. SIAM Journal on Numerical Analysis, 16(6):923–933, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1970:MRH

[Osh70] Stanley Osher. Mesh refinements for the heat equation. SIAM Journal on Numerical Analysis, 7(2):199–205, June 1970. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1981:NSP

[Osh81] Stanley Osher. Nonlinear singular perturbation problems and one-sided difference schemes. SIAM Journal on Numerical Analysis, 18(1): 129–144, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1984:RSE

[Osh84]  Stanley Osher.  Riemann

solvers, the entropy condition, and difference approximations. SIAM Journal on Numerical Analysis, 21(2): 217–235, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osher:1985:CGM

[Osh85]  Stanley Osher.  Conver-

gence of generalized MUSCL schemes. SIAM Journal on Numerical Analysis, 22(5): 947–961, October 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ostrowski:1964:AEA

[Ost64] A. M. Ostrowski. On approximation of equations by algebraic equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):104–130, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic). See errata [Ost65].

Ostrowski:1965:EAE

[Ost65] A. M. Ostrowski. Errata: “On Approximation of Equations by Algebraic Equations” [J. Soc. Ind. Appl. Math., Ser. B Numer. anal. 1] (1964), 104–130. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):203, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic). See [Ost64].

Ostrowski:1970:TCR

[Ost70] A. M. Ostrowski. A theorem on clusters of roots of polynomial equations. SIAM Journal on Numerical Analysis, 7(4):567–570, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ostrowski:1971:SPR

[Ost71]  A. M. Ostrowski.  Some

properties of reduced polynomial equations. SIAM Journal on Numerical Analysis, 8 (4):623–638, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ostrowski:1973:PEE

[Ost73] A. M. Ostrowski. A Posteriori error estimates in iterative procedures. SIAM Journal on Numerical Analysis, 10 (2):290–298, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Oswald:1992:HCF

[Osw92] P. Oswald. Hierarchical conforming finite element methods for the biharmonic equation. SIAM Journal on Numerical Analysis, 29(6):1610– 1625, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

ONeil:1975:CTD

[OT75] T. O’Neil and J. W. Thomas. The calculation of the topological degree by quadrature. SIAM Journal on Numerical Analysis, 12(5):673– 680, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Otto:1996:MCD

[Ott96a] James S. Otto. Multigrid convergence for discretizations of singular perturbation problems with grid-aligned flow. SIAM Journal on Numerical Analysis, 33(1):399– 416, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Otto:1996:APH

[Ott96b]  Kurt Otto.  Analysis of

preconditioners for hyperbolic partial differential equations. SIAM Journal on Numerical Analysis, 33(6): 2131–2165, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26489.

Owen:1997:MCV

[Owe97] Art B. Owen. Monte Carlo variance of scrambled net quadrature. SIAM Journal on Numerical Analysis, 34 (5):1884–1910, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27746.

Pye:1973:ACH

[PA73] W. C. Pye and T. A. Atchison. An algorithm for the computation of the higher order G-transformation. SIAM Journal on Numerical Analysis, 10(1):1–7, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Padra:1997:PEE

[Pad97]  Claudio Padra.  A Pos-

teriori error estimators for nonconforming approximation of some quasi-Newtonian flows. SIAM Journal on Numerical Analysis, 34(4): 1600–1615, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27832.

Paige:1974:BMS

[Pai74] C. C. Paige. Bidiagonalization of matrices and solution of linear equations. SIAM Journal on Numerical Analysis, 11(1):197–209, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paige:1979:FNS

[Pai79] C. C. Paige. Fast numerically stable computations for generalized linear least squares problems. SIAM Journal on Numerical Analysis, 16(1): 165–171, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paige:1984:NRS

[Pai84] C. C. Paige. A note on a result of Sun Ji Guang: sensitivity of the CS and GSV decompositions. SIAM Journal on Numerical Analysis, 21(1): 186–191, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Palencia:1993:SRS

[Pal93] C. Palencia. A stability result for sectorial operators in Banach spaces. SIAM Journal on Numerical Analysis, 30 (5):1373–1384, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Palencia:1996:MNA

[Pal96] C. Palencia. Maximum norm analysis of completely discrete finite element methods for parabolic problems. SIAM Journal on Numerical Analysis, 33(4):1654– 1668, August 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25977.

Pandian:1985:CTC

[Pan85] Maharaja C. Pandian. A convergence test and componentwise error estimates for Newton type methods. SIAM Journal on Numerical Analysis, 22(4):779–791, August 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pan:1997:GSB

[Pan97] Jianhua Pan. Global superconvergence for the bilinearconstant scheme for the Stokes problem. SIAM Journal on Numerical Analysis, 34(6):2424–2430, December 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28616.

Pani:1998:GMF

[Pan98] Amiya K. Pani. An H^{1}- Galerkin mixed finite element method for parabolic partial differential equations. SIAM Journal on Numerical Analysis, 35(2):712–727, April 1998. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/28080.

Pao:1987:NMS

[Pao87] C. V. Pao. Numerical methods for semilinear parabolic equations. SIAM Journal on Numerical Analysis, 24(1): 24–35, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pao:1999:NAC

[Pao99] C. V. Pao. Numerical analysis of coupled systems of nonlinear parabolic equations. SIAM Journal on Numerical Analysis, 36(2): 393–416, April 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31316.

Parter:1965:NMG

[Par65] Seymour V. Parter. Numerical methods for generalized axially symmetric potentials. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):500–516, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Parter:1982:ESO

[Par82] Seymour V. Parter. On the eigenvalues of second order elliptic difference operators. SIAM Journal on Numerical Analysis, 19(3):518–530, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parsons:1987:GPS

[Par87a] Bradley N. Parsons. General k-part stationary iterative solutions to linear systems. SIAM Journal on Numerical Analysis, 24(1):188– 198, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parter:1987:ETG

[Par87b]  Seymour V. Parter.  On

an estimate for the threegrid MGR multigrid method. SIAM Journal on Numerical Analysis, 24(5):1032– 1045, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Park:1995:MFE

[Par95] Eun-Jae Park. Mixed finite element methods for nonlinear second-order elliptic problems. SIAM Journal on Numerical Analysis, 32(3): 865–885, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parter:1999:OGM

[Par99]  Seymour V. Parter.  On

the overlapping grid method for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 36(3):819–852, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31075.

Pasciak:1979:PCM

[Pas79] Joseph Pasciak. The penalty correction method for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 16(6):1046– 1059, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pasciak:1982:SMN

[Pas82] Joseph E. Pasciak. Spectral methods for a nonlinear initial value problem involving pseudodifferential operators. SIAM Journal on Numerical Analysis, 19(1):142– 154, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Patent:1976:EQE

[Pat76] Paul D. Patent. The effect of quadrature errors in the computation of L^{2} piecewise polynomial approximations. SIAM Journal on Numerical Analysis, 13(3): 344–361, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pehlivanov:1994:LSM

[PCL94] A. I. Pehlivanov, G. F. Carey, and R. D. Lazarov. Leastsquares mixed finite elements for second-order elliptic problems. SIAM Journal on Numerical Analysis, 31(5):1368– 1377, October 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peplow:1995:ASS

[PCW95] Andrew T. Peplow and Simon N. Chandler-Wilde. Approximate solution of second kind integral equations on infinite cylindrical surfaces. SIAM Journal on Numerical Analysis, 32(2): 594–609, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pelios:1967:RFA

[Pel67] A. Pelios. Rational function approximation as a wellconditioned matrix eigenvalue problem. SIAM Journal on Numerical Analysis, 4 (4):542–547, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pereyra:1967:ACD

[Per67a] Victor Pereyra. Accelerating the convergence of discretization algorithms. SIAM Journal on Numerical Analysis, 4 (4):508–533, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pereyra:1967:IMS

[Per67b] Victor Pereyra. Iterative methods for solving nonlinear least squares problems. SIAM Journal on Numerical Analysis, 4(1):27–36, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Percell:1976:CQC

[Per76]  Peter Percell.  On cubic

and quartic Clough–Tocher finite elements. SIAM Journal on Numerical Analysis, 13(1): 100–103, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Perthame:1990:BTS

[Per90] B. Perthame. Boltzmann type schemes for gas dynamics and the entropy property. SIAM Journal on Numerical Analysis, 27(6):1405–1421, December 1990. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Perthame:1992:SOB

[Per92] B. Perthame. Second-order Boltzmann schemes for compressible Euler equations in one and two space dimensions. SIAM Journal on Numerical Analysis, 29(1):1– 19, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Perugia:1997:FBM

[Per97] I. Perugia. A field-based mixed formulation for the two-dimensional magnetostatic problem. SIAM Journal on Numerical Analysis, 34(6): 2382–2391, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29346.

Petryshyn:1965:CPL

[Pet65] W. V. Petryshyn. Constructional proof of lax-milgram lemma and its application to non-K-P.D. abstract and differential operator equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):404–420, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Petzold:1981:ENM

[Pet81] Linda R. Petzold. An efficient numerical method for highly oscillatory ordinary differential equations. SIAM Journal on Numerical Analysis, 18 (3):455–479, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Petzold:1986:ORI

[Pet86] L. R. Petzold. Order results for implicit Runge–Kutta methods applied to differential/algebraic systems. SIAM Journal on Numerical Analysis, 23(4):837–852, August 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pettersen:1987:RCM

[Pet87] Øystein Pettersen. The random choice method: Structural stability of a conservation equation in reservoir dynamics. SIAM Journal on Numerical Analysis, 24(5): 997–1007, October 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Petkovic:1989:HLA

[Pet89] Miodrag S. Petković. On halley-like algorithms for simultaneous approximation of polynomial complex zeros. SIAM Journal on Numerical Analysis, 26(3):740–763, June 1989. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Peterson:1991:NCD

[Pet91] Todd E. Peterson. A note on the convergence of the discontinuous Galerkin method for a scalar hyperbolic equation. SIAM Journal on Numerical Analysis, 28(1):133– 140, February 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Petersson:1992:NMC

[Pet92a] N. Anders Petersson. A numerical method to calculate the two-dimensional flow around an underwater obstacle. SIAM Journal on Numerical Analysis, 29(1):20– 31, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Petras:1992:EBG

[Pet92b] Knut Petras. Error bounds for Gaussian and related quadrature and applications to R-Convex functions. SIAM Journal on Numerical Analysis, 29(2):578–585, April 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:1995:SS

[Pet95]  Jörg Peters.  C^{1}-surface

splines. SIAM Journal on Numerical Analysis, 32(2):

645–666, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Petersen:1998:GIS

[Pet98a] W. P. Petersen. A general implicit splitting for stabilizing numerical simulations of Itô stochastic differential equations. SIAM Journal on Numerical Analysis, 35 (4):1439–1451, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30397.

Petry:1998:SAT

[Pet98b] Thomas Petry. On the stability of the Abramov transfer for differential-algebraic equations of index 1. SIAM Journal on Numerical Analysis, 35 (1):201–216, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26666.

Petersson:1999:FEA

[Pet99] Joakim Petersson. A finite element analysis of optimal variable thickness sheets. SIAM Journal on Numerical Analysis, 36(6): 1759–1778, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31396.

Pflaum:1997:CCT

[Pfl97]  Christoph Pflaum.  Con-

vergence of the combination technique for secondorder elliptic differential equations. SIAM Journal on Numerical Analysis, 34(6): 2431–2455, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26029.

Pruess:1995:ANM

[PFX95] Steven Pruess, Charles T.

Fulton, and Yuantao Xie. An asymptotic numerical method for a class of singular Sturm– Liouville problems. SIAM Journal on Numerical Analysis, 32(5):1658–1676, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Prasad:1973:UCF

[PH73] J. Prasad and H. Hayashi. On the uniform convergence of Fourier–Jacobi series. SIAM Journal on Numerical Analysis, 10(1):23–27, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Phillips:1972:UCP

[Phi72] James L. Phillips. The use of collocation as a projection method for solving linear operator equations. SIAM Journal on Numerical Analysis, 9 (1):14–28, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pietrzykowski:1969:EPM

[Pie69] Tomasz Pietrzykowski. An exact potential method for constrained maxima. SIAM Journal on Numerical Analysis, 6(2):299–304, June 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Pie71].

Pietrzykowski:1971:EEP

[Pie71]  T. Pietrzykowski.  Erra-

tum: “An Exact Potential Method for Constrained Maxima” [SIAM J. Numer. Anal. 6 (1969), no. 2, 299–304]. SIAM Journal on Numerical Analysis, 8(2):481, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Pie69].

Pierre:1995:OSB

[Pie95] Roger Pierre. Optimal selection of the bubble function in the stabilization of the p1-p1 element for the Stokes problem. SIAM Journal on Numerical Analysis, 32(4):1210– 1224, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pitkaranta:1978:SDD

[Pit78] Juhani Pitkäranta. On the spatial differencing of the discrete ordinate neutron transport equation. SIAM Journal on Numerical Analysis, 15(5): 859–869, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Phillips:1990:CIE

[PK90] Timothy N. Phillips and Andreas Karageorghis. On the coefficients of integrated expansions of ultraspherical polynomials. SIAM Journal on Numerical Analysis, 27(3): 823–830, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Porsching:1987:RBM

[PL87] T. A. Porsching and M. Lin Lee. The reduced basis method for initial value problems. SIAM Journal on Numerical Analysis, 24(6):1277– 1287, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Plato:1998:MCR

[Pla98] R. Plato. The method of conjugate residuals for solving the Galerkin equations associated with symmetric positive semidefinite ill-posed problems. SIAM Journal on Numerical Analysis, 35 (4):1621–1645, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31792.

Poole:1987:MIM

[PO87] Eugene L. Poole and James M. Ortega. Multicolor ICCG methods for vector computers. SIAM Journal on Numerical Analysis, 24(6):1394– 1418, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Polak:1974:GCS

[Pol74] E. Polak. A globally converging secant method with applications to boundary value problems. SIAM Journal on Numerical Analysis, 11(3): 529–537, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Porsching:1969:JGS

[Por69] T. A. Porsching. Jacobi and Gauss–Seidel methods for nonlinear network problems. SIAM Journal on Numerical Analysis, 6(3):437– 449, September 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Porsching:1971:RCJ

[Por71] T. A. Porsching. On rates of convergence of Jacobi and Gauss–Seidel methods for Mfunctions. SIAM Journal on Numerical Analysis, 8 (3):575–582, September 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Poreda:1973:AP

[Por73] S. J. Poreda. Approximation by δ-polynomials. SIAM Journal on Numerical Analysis, 10(1):50–54, March 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Potra:1993:ILM

[Pot93] Florian A. Potra. Implementation of linear multistep methods for solving constrained equations of motion. SIAM Journal on Numerical Analysis, 30(3):774–789, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Powell:1969:LDL

[Pow69] M. J. D. Powell. The local dependence of least squares cubic splines. SIAM Journal on Numerical Analysis, 6 (3):398–413, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parlett:1973:GTP

[PP73a] B. N. Parlett and W. G.

Poole, Jr. A geometric theory for the QR, LU and power iterations. SIAM Journal on Numerical Analysis, 10 (2):389–412, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pironneau:1973:RCC

[PP73b] O. Pironneau and E. Polak. Rate of convergence of a class of methods of feasible directions. SIAM Journal on Numerical Analysis, 10(1): 161–174, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pani:1996:FEM

[PP96] Amiya K. Pani and Todd E. Peterson. Finite element methods with numerical quadrature for parabolic integrodifferential equations. SIAM Journal on Numerical Analysis, 33(3):1084– 1105, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Prenter:1976:OCE

[PR76] P. M. Prenter and R. D. Russell. Orthogonal collocation for elliptic partial differential equations. SIAM Journal on Numerical Analysis, 13 (6):923–939, December 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Passow:1977:MCS

[PR77] Eli Passow and John A.

Roulier. Monotone and convex spline interpolation. SIAM Journal on Numerical Analysis, 14(5):904–909, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Philips:1982:CFD

[PR82] Richard B. Philips and Milton E. Rose. Compact finite difference schemes for mixed initial-boundary value problems. SIAM Journal on Numerical Analysis, 19(4):698– 720, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paine:1986:CCM

[PR86a] John Paine and Robert D.

Russell. Conditioning of collocation matrices and discrete Green’s functions. SIAM Journal on Numerical Analysis, 23(2):376–392, April 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Porter:1986:NRB

[PR86b] Michael B. Porter and Edward L. Reiss. A note on the relationship between finitedifference and shooting methods for ODE eigenvalue problems. SIAM Journal on Numerical Analysis, 23(5):1034– 1039, October 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parter:1995:PLS

[PR95] Seymour V. Parter and Ernest E. Rothman. Preconditioning Legendre spectral collocation approximations to elliptic problems. SIAM Journal on Numerical Analysis, 32 (2):333–385, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:1998:AAG

[PR98] Jörg Peters and Ulrich Reif. Analysis of algorithms generalizing B-spline subdivision. SIAM Journal on Numerical Analysis, 35(2): 728–748, April 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30434.

Prasad:1970:WI

[Pra70] J. Prasad. On the weighted (0,2) interpolation. SIAM Journal on Numerical Analysis, 7(3):428–446, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Prenter:1973:CMN

[Pre73] P. M. Prenter. A collection method for the numerical solution of integral equations. SIAM Journal on Numerical Analysis, 10(4):570– 581, September 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Price:1976:CEB

[Pri76] Thomas E. Price, Jr. Cubature error bounds for a class of analytic functions. SIAM Journal on Numerical Analysis, 13(2):227–235, April 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pridor:1977:EMN

[Pri77] Adir Pridor. Estimation of moments for the numerical solution of transport problems. SIAM Journal on Numerical Analysis, 14(3):426–440, June 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Price:1979:OPN

[Pri79] Thomas E. Price, Jr. Orthogonal polynomials for nonclassical weight functions. SIAM Journal on Numerical Analysis, 16(6):999–1006, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peisker:1990:ISM

[PRS90] P. Peisker, W. Rust, and E. Stein. Iterative solution methods for plate bending problems: Multigrid and preconditioned cg algorithm.

SIAM Journal on Numerical Analysis, 27(6):1450– 1465, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pruess:1973:EES

[Pru73] Steven Pruess. Estimating the eigenvalues of Sturm– Liouville problems by approximating the differential equation. SIAM Journal on Numerical Analysis, 10(1): 55–68, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Prufer:1982:CAP

[Prü82] Michael Prüfer. A combinatorial algorithm providing alternating approximations for a zero of an M-function. SIAM Journal on Numerical Analysis, 19(3):643–652, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pryce:1984:NMR

[Pry84] J. D. Pryce. A new measure of relative error for vectors. SIAM Journal on Numerical Analysis, 21(1):202– 215, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paige:1975:SSI

[PS75] C. C. Paige and M. A. Saunders. Solution of sparse indefinite systems of linear equations. SIAM Journal on Numerical Analysis, 12(4):617– 629, September 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paige:1977:LSE

[PS77] C. C. Paige and M. A. Saunders. Least squares estimation of discrete linear dynamic systems using orthogonal transformations. SIAM Journal on Numerical Analysis, 14(2): 180–193, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parter:1980:LBI

[PS80] Seymour V. Parter and Michael Steuerwalt. On Kline and K ×K block iterative schemes for a problem arising in three-dimensional elliptic difference equations. SIAM Journal on Numerical Analysis, 17(6):823–839, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paige:1981:TGS

[PS81] C. C. Paige and M. A. Saunders. Towards a generalized singular value decomposition. SIAM Journal on Numerical Analysis, 18(3):398–405, June 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parter:1982:BIM

[PS82] Seymour V. Parter and Michael Steuerwalt. Block iterative methods for elliptic and parabolic difference equations. SIAM Journal on Numerical Analysis, 19(6):1173– 1195, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pitkaranta:1983:EEC

[PS83] Juhani Pitkäranta and L. Ridgway Scott. Error estimates for the combined spatial and angular approximations of the transport equation for slab geometry. SIAM Journal on Numerical Analysis, 20(5): 922–950, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Parter:1985:BIM

[PS85] Seymour V. Parter and Michael Steuerwalt. Block iterative methods for elliptic finite element equations. SIAM Journal on Numerical Analysis, 22(1):146–179, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pego:1988:IGS

[PS88] Robert L. Pego and Denis Serre. Instabilities in Glimm’s scheme for two systems of mixed type. SIAM Journal on Numerical Analysis, 25(5):

965–988, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:1992:SIC

[PS92] Jörg Peters and Meera Sitharam. Stability of interpolation from C^{1} cubics at the vertices of an underlying triangulation. SIAM Journal on Numerical Analysis, 29 (2):528–533, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Palczewski:1997:CRD

[PSB97] Andrzej Palczewski, Jacques Schneider, and Alexandre V. Bobylev. A consistency result for a discrete-velocity model of the Boltzmann equation. SIAM Journal on Numerical Analysis, 34(5): 1865–1883, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28900.

Papadopoulou:1989:BAI

[PSP89] E. P. Papadopoulou, Y. G.

Saridakis, and T. S. Papatheodorou. Block AOR iterative schemes for large-scale least-squares problems. SIAM Journal on Numerical Analysis, 26(3):637–660, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Powell:1979:ESH

[PT79] M. J. D. Powell and Ph. L. Toint. On the estimation of sparse Hessian matrices. SIAM Journal on Numerical Analysis, 16(6):1060– 1074, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Panier:1991:AME

[PT91] Eliane R. Panier and André L. Tits. Avoiding the maratos effect by means of a nonmonotone line search I. general constrained problems. SIAM Journal on Numerical Analysis, 28(4):1183–1195, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Papakostas:1997:HCI

[PT97] S. N. Papakostas and Ch. Tsitouras. Highly continuous interpolants for onestep ODE solvers and their application to Runge–Kutta methods. SIAM Journal on Numerical Analysis, 34 (1):22–47, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26580.

Papakostas:1996:GFE

[PTP96] S. N. Papakostas, Ch Tsitouras, and G. Papageorgiou. A general family of explicit Runge–Kutta pairs of orders 6(5). SIAM Journal on Numerical Analysis, 33(3): 917–936, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pierce:1972:HOCa

[PV72] J. G. Pierce and R. S. Varga.

Higher order convergence results for the Rayleigh–Ritz method applied to eigenvalue problems. I: Estimates relating Rayleigh–Ritz and Galerkin approximations to eigenfunctions. SIAM Journal on Numerical Analysis, 9(1): 137–151, March 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic). URL http://www.
math.kent.edu/~varga/pub/
paper_67.pdf.

Prasad:1979:LIQ

[PV79] J. Prasad and A. K. Varma.

Lacunary interpolation by quintic splines. SIAM Journal on Numerical Analysis, 16(6): 1075–1079, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peters:1970:GE

[PW70] G. Peters and J. H. Wilkinson. ax = λbx and the generalized eigenproblem. SIAM Journal on Numerical Analysis, 7(4):479–492, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Percell:1978:LRF

[PW78] Peter Percell and Mary Fanett Wheeler. A local residual finite element procedure for elliptic equations. SIAM Journal on Numerical Analysis, 15(4):705–714, August 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Percell:1980:FEC

[PW80] Peter Percell and Mary Fanett Wheeler. A C^{1} finite element collocation method for elliptic equations. SIAM Journal on Numerical Analysis, 17(5): 605–622, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pavarino:1996:PBI

[PW96] Luca F. Pavarino and Olof B. Widlund. A polylogarithmic bound for an iterative substructuring method for spectral elements in three dimensions. SIAM Journal on Numerical Analysis, 33 (4):1303–1335, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26517.

Qi:1982:NMT

[Qi82]  L. Qi.  A note on the

Moore test for nonlinear systems. SIAM Journal on Numerical Analysis, 19(4):851– 857, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Qi-nian:1999:AMD

[Qn99] Jin Qi-nian. Applications of the modified discrepancy principle to Tikhonov regularization of nonlinear illposed problems. SIAM Journal on Numerical Analysis, 36(2):475–490, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31547.

Quarteroni:1987:FSM

[Qua87] Alfio Quarteroni. Fourier spectral methods for pseudoparabolic equations. SIAM Journal on Numerical Analysis, 24(2):323–335, April 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Queck:1989:CFP

[Que89] Werner Queck. The convergence factor of preconditioned algorithms of the arrow-hurwicz type. SIAM Journal on Numerical Analysis, 26(4):1016–1030, August 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Quarteroni:1992:FEP

[QZ92] Alfio Quarteroni and Elena Zampieri. Finite element preconditioning for Legendre spectral collocation approximations to elliptic equations and systems. SIAM Journal on Numerical Analysis, 29(4): 917–936, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rabinowitz:1967:GIP

[Rab67] Philip Rabinowitz. Gaussian integration in the presence of a singularity. SIAM Journal on Numerical Analysis, 4 (2):191–201, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rachford:1968:REA

[Rac68a] H. H. Rachford, Jr. Rounding errors in alternating direction methods for parabolic problems. SIAM Journal on Numerical Analysis, 5(2): 407–421, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rachford:1968:REP

[Rac68b] H. H. Rachford, Jr. Rounding errors in parabolic problems. I: The one-space variable case. SIAM Journal on Numerical Analysis, 5(1): 156–171, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rachford:1973:TLD

[Rac73] H. H. Rachford, Jr. Two-level discrete-time Galerkin approximations for second order nonlinear parabolic partial differential equations. SIAM Journal on Numerical Analysis, 10(6):1010–1026, December 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ragozin:1985:DFS

[Rag85] David L. Ragozin. The discrete k-functional and spline smoothing of noisy data. SIAM Journal on Numerical Analysis, 22(6):1243– 1254, December 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rakotch:1976:NSL

[Rak76] E. Rakotch. Numerical solution with large matrices of Fredholm’s integral equation. SIAM Journal on Numerical Analysis, 13(1):1–7, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rall:1974:NCN

[Ral74] L. B. Rall. A note on the convergence of Newton’s method. SIAM Journal on Numerical Analysis, 11(1):34–36, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rall:1980:CET

[Ral80]  L. B. Rall.  A compar-

ison of the existence theorems of Kantorovich and Moore. SIAM Journal on Numerical Analysis, 17(1):148– 161, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ratschek:1980:CF

[Rat80]  H. Ratschek.  Centered

forms. SIAM Journal on Numerical Analysis, 17(5):656– 662, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rathsfeld:1995:NMI

[Rat95]  A. Rathsfeld.  Nyström’s

method and iterative solvers for the solution of the doublelayer potential equation over polyhedral boundaries. SIAM Journal on Numerical Analysis, 32(3):924–951, June 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rauch:1973:CTC

[Rau73] Steven W. Rauch. A convergence theory for a class of nonlinear programming problems. SIAM Journal on Numerical Analysis, 10(1):207– 228, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rauch:1985:CFE

[Rau85]  Jeffrey Rauch.  On con-

vergence of the finite element method for the wave equation. SIAM Journal on Numerical Analysis, 22(2): 245–249, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Robinson:1979:PBB

[RB79] Peter D. Robinson and Michael F. Barnsley. Pointwise bivariational bounds on solutions of Fredholm integral equations. SIAM Journal on Numerical Analysis, 16(1): 135–144, February 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rogers:1979:APT

[RBC79] Joel C. W. Rogers, Alan E.

Berger, and Melvyn Ciment. The alternating phase truncation method for numerical solution of a Stefan problem. SIAM Journal on Numerical Analysis, 16(4):563– 587, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rivlin:1966:CUA

[RC66] T. J. Rivlin and E. W. Cheney. A comparison of uniform approximations on an interval and a finite subset thereof. SIAM Journal on Numerical Analysis, 3(2): 311–320, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Russell:1978:AMS

[RC78] R. D. Russell and J. Christiansen. Adaptive mesh selection strategies for solving boundary value problems. SIAM Journal on Numerical Analysis, 15(1):59–80, February 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Richter-Dyn:1971:MIA

[RD71a] Nira Richter-Dyn. Minimal interpolation and approximation in Hilbert spaces. SIAM Journal on Numerical Analysis, 8(3):583–597, September 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Richter-Dyn:1971:PMI

[RD71b] Nira Richter-Dyn. Properties of minimal integration rules. II. SIAM Journal on Numerical Analysis, 8(3):497– 508, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Richter-Dyn:1979:BNA

[RD79] Nira Richter-Dyn. On best nonlinear approximation in sign-monotone norms and in norms induced by inner products. SIAM Journal on Numerical Analysis, 16(4):612– 622, August 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reddien:1976:AMT

[Red76] G. W. Reddien. Approximation methods for twopoint boundary value problems with nonlinear boundary conditions. SIAM Journal on Numerical Analysis, 13 (3):405–411, June 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reddien:1978:NMS

[Red78] G. W. Reddien. On Newton’s method for singular problems. SIAM Journal on Numerical Analysis, 15(5):993– 996, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reemtsen:1990:MFR

[Ree90] Rembert Reemtsen. Modifications of the first Remez algorithm. SIAM Journal on Numerical Analysis, 27(2): 507–518, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reimer:1968:BHS

[Rei68a] Manfred Reimer. Bounds for the Horner sums. SIAM Journal on Numerical Analysis, 5 (3):461–469, September 1968.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reimer:1968:FDF

[Rei68b]  Manfred Reimer.  Finite

difference forms containing derivatives of higher order. SIAM Journal on Numerical Analysis, 5(4):725– 738, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reid:1972:UCG

[Rei72] J. K. Reid. The use of conjugate gradients for systems of linear equations possessing “property A”. SIAM Journal on Numerical Analysis, 9 (2):325–332, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reinsch:1974:TES

[Rei74]  Christian Reinsch.  Two

extensions of the sardschoenberg theory of best approximation. SIAM Journal on Numerical Analysis, 11 (1):45–51, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reinhardt:1981:PEE

[Rei81] Hans-Jürgen Reinhardt. A Posteriori error estimates for the finite element solution of a singularly perturbed linear ordinary differential equation.

SIAM Journal on Numerical Analysis, 18(3):406–430, June 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reichel:1988:PCM

[Rei88] Lothar Reichel. Polynomials by conformal mapping for the Richardson iteration method for complex linear systems. SIAM Journal on Numerical Analysis, 25(6): 1359–1368, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reich:1996:SIC

[Rei96] Sebastian Reich. Symplectic integration of constrained Hamiltonian systems by composition methods. SIAM Journal on Numerical Analysis, 33(2):475–491, April 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reich:1999:BEA

[Rei99] Sebastian Reich. Backward error analysis for numerical integrators. SIAM Journal on Numerical Analysis, 36 (5):1549–1570, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32979.

Reusken:1992:MNC

[Reu92] Arnold Reusken. On maximum norm convergence of multigrid methods for twopoint boundary value problems. SIAM Journal on Numerical Analysis, 29(6):1569– 1578, December 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reusken:1994:MNC

[Reu94] Arnold Reusken. On maximum norm convergence of multigrid methods for elliptic boundary value problems. SIAM Journal on Numerical Analysis, 31(2): 378–392, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reynolds:1972:CFD

[Rey72] Albert C. Reynolds, Jr. Convergent finite difference schemes for nonlinear parabolic equations. SIAM Journal on Numerical Analysis, 9 (4):523–533, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1968:UCT

[Rhe68] Werner C. Rheinboldt. A unified convergence theory for a class of iterative processes. SIAM Journal on Numerical Analysis, 5(1):42–63, March 1968. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1978:NMC

[Rhe78] Werner C. Rheinboldt. Numerical methods for a class of finite dimensional bifurcation problems. SIAM Journal on Numerical Analysis, 15 (1):1–11, February 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1980:TMR

[Rhe80a] Werner C. Rheinboldt. On a theory of mesh-refinement processes. SIAM Journal on Numerical Analysis, 17 (6):766–778, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1980:SFN

[Rhe80b] Werner C. Rheinboldt. Solution fields of nonlinear equations and continuation methods. SIAM Journal on Numerical Analysis, 17(2): 221–237, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1982:CCB

[Rhe82] Werner C. Rheinboldt. Computation of critical boundaries on equilibrium manifolds. SIAM Journal on Numerical Analysis, 19(3): 653–669, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1993:SSP

[Rhe93] Werner C. Rheinboldt. On the sensitivity of solutions of parameterized equations. SIAM Journal on Numerical Analysis, 30(2):305–320, April 1993. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reinhardt:1999:SRD

[RHH99] Hans-Jürgen Reinhardt, Houde Han, and Dinh Nho Hào. Stability and regularization of a discrete approximation to the Cauchy problem for Laplace’s equation. SIAM Journal on Numerical Analysis, 36(3):890–905, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31695.

Rice:1966:TC

[Ric66] John R. Rice. A theory of condition. SIAM Journal on Numerical Analysis, 3 (2):287–310, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rice:1967:CCA

[Ric67] John R. Rice. Characterization of Chebyshev approximations by splines. SIAM Journal on Numerical Analysis, 4 (4):557–565, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Richter:1970:PMI

[Ric70]  Nira Richter.  Properties

of minimal integration rules. SIAM Journal on Numerical Analysis, 7(1):67–78, March 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Richter:1978:NSI

[Ric78] G. R. Richter. Numerical solution of integral equations of the first kind with nonsmooth kernels. SIAM Journal on Numerical Analysis, 15 (3):511–522, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Richter:1991:EFE

[Ric91] Gerard R. Richter. An explicit finite element method for convection-dominated steady state convection-diffusion equations. SIAM Journal on Numerical Analysis, 28(3): 744–759, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Riess:1971:NEB

[Rie71] R. D. Riess. A note on error bounds for Gauss–Chebyshev quadrature. SIAM Journal on Numerical Analysis, 8 (3):509–511, September 1971.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Riedel:1994:GEP

[Rie94] Kurt S. Riedel. Generalized epsilon-pseudospectra. SIAM Journal on Numerical Analysis, 31(4):1219–1225, August 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ringhofer:1984:CSQ

[Rin84] Christian Ringhofer. On collocation schemes for quasilinear singularly perturbed boundary value problems. SIAM Journal on Numerical Analysis, 21(5):864–882, October 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ringhofer:1990:SMN

[Rin90] Christian Ringhofer. A spectral method for the numerical simulation of quantum tunneling phenomena. SIAM Journal on Numerical Analysis, 27(1):32–50, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ringhofer:1992:SCT

[Rin92] Christian Ringhofer. A spectral collocation technique for the solution of the Wigner– Poisson problem. SIAM Journal on Numerical Analysis, 29 (3):679–700, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rippa:1992:LTT

[Rip92] Shmuel Rippa. Long and thin triangles can be good for linear interpolation. SIAM Journal on Numerical Analysis, 29 (1):257–270, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ritter:1975:MCD

[Rit75] Klaus Ritter. A method of conjugate directions for linearly constrained nonlinear programming problems. SIAM Journal on Numerical Analysis, 12(3):273–303, June 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ritov:1990:CAF

[Rit90] Y. Ritov. The convergence of an algorithm for finding the distance between a ball in a subspace and a sum of subspaces. SIAM Journal on Numerical Analysis, 27(5):1355– 1367, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rivlin:1964:PRT

[Riv64] T. J. Rivlin. A property of the ratio of trigonometric polynomials. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):131– 132, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Rivlin:1975:OSL

[Riv75] T. J. Rivlin. Optimally stable Lagrangian numerical differentiation. SIAM Journal on Numerical Analysis, 12(5): 712–725, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rivara:1984:MRP

[Riv84] María-Cecilia Rivara. Mesh refinement processes based on the generalized bisection of simplices. SIAM Journal on Numerical Analysis, 21(3): 604–613, June 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Riess:1969:EGC

[RJ69] R. D. Riess and L. W. Johnson. Estimating Gauss– Chebyshev quadrature errors. SIAM Journal on Numerical Analysis, 6(4):557– 559, December 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Riess:1974:EIF

[RJ74] R. D. Riess and L. W. Johnson. Errors in interpolating functions at the zeros of T_{n+1}(X). SIAM Journal on Numerical Analysis, 11(2): 244–253, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ramanujam:1988:MLE

[RK88] N. Ramanujam and V. M. Sunanda Kumari. Method of lines for elliptic differential equations with a small parameter. SIAM Journal on Numerical Analysis, 25(3):550–563, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rowland:1977:ECN

[RM77] J. H. Rowland and G. J. Miel.

Exit criteria for Newton– Cotes quadrature rules. SIAM Journal on Numerical Analysis, 14(6):1145–1150, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Robinson:1966:ISS

[Rob66] Stephen M. Robinson. Interpolative solution of systems of nonlinear equations. SIAM Journal on Numerical Analysis, 3(4):650–658, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Robinson:1975:STS

[Rob75] Stephen M. Robinson. Stability theory for systems of inequalities. Part I: Linear systems. SIAM Journal on Numerical Analysis, 12(5):754– 769, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Robinson:1976:STS

[Rob76] Stephen M. Robinson. Stability theory for systems of inequalities, Part II: Differentiable nonlinear systems. SIAM Journal on Numerical Analysis, 13(4):497–513, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Robinson:1979:QIR

[Rob79] Stephen M. Robinson. Quadratic interpolation is risky. SIAM Journal on Numerical Analysis, 16(3):377–379, June 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Roche:1989:IRK

[Roc89]  Michel Roche.  Implicit

Runge–Kutta methods for differential algebraic equations. SIAM Journal on Numerical Analysis, 26(4): 963–975, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rogier:1993:MNS

[Rog93]  François Rogier.  Mathe-

matical and numerical study of a magnetostatic problem around a thin shield. SIAM Journal on Numerical Analysis, 30(2):454–477, April 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rohn:1993:IIM

[Roh93]  J. Rohn.  Inverse inter-

val matrix. SIAM Journal on Numerical Analysis, 30(3): 864–870, June 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Romate:1990:LEA

[Rom90] J. E. Romate. Local error analysis of three-dimensional panel methods in terms of curvilinear surface coordinates. SIAM Journal on Numerical Analysis, 27(2): 529–542, April 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rosenberg:1967:BPM

[Ros67] Lloyd Rosenberg. Bernstein polynomials and Monte Carlo integration. SIAM Journal on Numerical Analysis, 4 (4):566–574, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rosen:1970:ASE

[Ros70a] J. B. Rosen. Approximate solution and error bounds for quasi-linear elliptic boundary value problems. SIAM Journal on Numerical Analysis, 7 (1):80–103, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rosman:1970:ERR

[Ros70b] Bernard H. Rosman. Extension of results by Rice and Schumaker on spline approximation. SIAM Journal on Numerical Analysis, 7(2): 314–316, June 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rosenzweig:1973:FDA

[Ros73]  M. B. Rosenzweig.  Fi-

nite difference approximation of the weak solution of a mildly nonlinear Dirichlet problem. SIAM Journal on Numerical Analysis, 10 (4):635–646, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rosser:1980:MFB

[Ros80] J. Barkley Rosser. Majorization formulas for a biharmonic function of two variables. SIAM Journal on Numerical Analysis, 17(2): 207–220, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rose:1981:UNT

[Ros81] Milton E. Rose. A “unified” numerical treatment of the wave equation and the Cauchy–Riemann equations. SIAM Journal on Numerical Analysis, 18(2): 372–376, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Roulier:1971:LOI

[Rou71] John Roulier. Linear operators invariant on nonnegative monotone functions. SIAM Journal on Numerical Analysis, 8(1):30–35, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Roux:1982:VSS

[Rou82] Marie-Noëlle Le Roux. Variable step size multistep methods for parabolic problems. SIAM Journal on Numerical Analysis, 19(4):725– 741, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rowland:1967:SEC

[Row67] John H. Rowland. On the structure of the error curve in Chebyshev approximation by polynomials. SIAM Journal on Numerical Analysis, 4 (3):398–405, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rowland:1969:LDP

[Row69] John H. Rowland. On the location of the deviation points in Chebyshev approximation by polynomials. SIAM Journal on Numerical Analysis, 6(1):118–126, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ratschek:1980:ACF

[RR80] H. Ratschek and J. Rokne. About the centered form. SIAM Journal on Numerical Analysis, 17(3):333–337, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rabier:1986:CCS

[RR86] P. J. Rabier and G. W. Reddien. Characterization and computation of singular points with maximum rank deficiency. SIAM Journal on Numerical Analysis, 23 (5):1040–1051, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ratschek:1987:EGO

[RR87] H. Ratschek and J. G. Rokne. Efficiency of a global optimization algorithm. SIAM Journal on Numerical Analysis, 24(5):1191–1201, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rabier:1995:NSE

[RR95] Patrick J. Rabier and Werner C. Rheinboldt. On the numerical solution of the Euler– Lagrange equations. SIAM Journal on Numerical Analysis, 32(1):318–329, February 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ranero:1996:ETD

[RR96] Ibrahim Bless Ranero and Tomás Chacón Rebello. An efficient two-dimensional vortex method with long time accuracy. SIAM Journal on Numerical Analysis, 33 (4):1425–1450, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24932.

Rohn:1998:ESL

[RR98] Jiri Rohn and Georg Rex.

Enclosing solutions of linear equations. SIAM Journal on Numerical Analysis, 35(2):524–539, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29942.

Russell:1975:NMS

[RS75] R. D. Russell and L. F.

Shampine. Numerical methods for singular boundary value problems. SIAM Journal on Numerical Analysis, 12 (1):13–36, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rabinowitz:1984:PIP

[RS84] Philip Rabinowitz and Ian H. Sloan. Product integration in the presence of a singularity. SIAM Journal on Numerical Analysis, 21(1):149– 166, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ruotsalainen:1987:SBE

[RS87] K. Ruotsalainen and J. Saranen. Some boundary element methods using Dirac’s distributions as trial functions. SIAM Journal on Numerical Analysis, 24(4):816– 827, August 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ringhofer:1989:ANM

[RS89] C. Ringhofer and C. Schmeiser. An approximate Newton method for the solution of the basic semiconductor device equations. SIAM Journal on Numerical Analysis, 26 (3):507–516, June 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Roe:1992:OPL

[RS92] P. L. Roe and D. Sidilkover. Optimum positive linear schemes for advection in two and three dimensions. SIAM Journal on Numerical Analysis, 29(6):1542–1568, December 1992. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Reddy:1995:SCC

[RS95] B. D. Reddy and J. C. Simo.

Stability and convergence of a class of enhanced strain methods. SIAM Journal on Numerical Analysis, 32(6):1705– 1728, December 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Riemenschneider:1997:MIS

[RS97a] Sherman D. Riemenschneider and Zuowei Shen. Multidimensional interpolatory subdivision schemes. SIAM Journal on Numerical Analysis, 34(6):2357–2381, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29431.

Russell:1997:SCD

[RS97b] Robert D. Russell and Weiwei Sun. Spline collocation differentiation matrices. SIAM Journal on Numerical Analysis, 34(6): 2274–2287, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27798.

Roberts:1969:PTN

[RSE69] S. M. Roberts, J. S. Shipman, and W. J. Ellis. A perturbation technique for nonlinear two-point boundary value problems. SIAM Journal on Numerical Analysis, 6 (3):347–358, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rao:1976:RAA

[RSK76] T. Mahadeva Rao, K. Subramanian, and E. V. Krishnamurthy. Residue arithmetic algorithms for exact computation of g-inverses of matrices. SIAM Journal on Numerical Analysis, 13(2): 155–171, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rouy:1992:VSA

[RT92] Elisabeth Rouy and Agnès Tourin. A viscosity solutions approach to shapefrom-shading. SIAM Journal on Numerical Analysis, 29 (3):867–884, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ryabenkii:1995:ABC

[RT95] V. S. Ryaben’kiĭ and S. V. Tsynkov. Artificial boundary conditions for the numerical solution of external viscous flow problems. SIAM Journal on Numerical Analysis, 32(5):1355–1389, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ruas:1989:ESA

[Rua89] Vitoriano Ruas. Existence and stability of asymmetric finite-element approximations in nonlinear incompressible analysis. SIAM Journal on Numerical Analysis, 26 (5):1031–1059, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rubel:1964:BPA

[Rub64] L. A. Rubel. Bounded polynomial approximation. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):133–136, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Rude:1993:FAM

[Rüd93] Ulrich Rüde. Fully adaptive multigrid methods. SIAM Journal on Numerical Analysis, 30(1):230–248, February 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ruhe:1973:ANE

[Ruh73] Axel Ruhe. Algorithms for the nonlinear eigenvalue problem. SIAM Journal on Numerical Analysis, 10(4):674– 689, September 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rulla:1996:EAI

[Rul96] Jim Rulla. Error analysis for implicit approximations to solutions to Cauchy problems. SIAM Journal on Numerical Analysis, 33(1):68–87, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rumelin:1982:NTS

[Rüm82] W. Rümelin. Numerical treatment of stochastic differential equations. SIAM Journal on Numerical Analysis, 19 (3):604–613, June 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Russell:1970:NSS

[Rus70] David L. Russell. Numerical solution of singular initial value problems. SIAM Journal on Numerical Analysis, 7(3):399–417, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Russell:1977:CCF

[Rus77] Robert D. Russell. A comparison of collocation and finite differences for two-point boundary value problems. SIAM Journal on Numerical Analysis, 14(1):19–39, March 1977. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Russell:1985:TSA

[Rus85] Thomas F. Russell. Time stepping along characteristics with incomplete iteration for a Galerkin approximation of miscible displacement in porous media. SIAM Journal on Numerical Analysis, 22(5): 970–1013, October 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rheinboldt:1974:LCU

[RV74] Werner C. Rheinboldt and James S. Vandergraft. On the local convergence of update methods. SIAM Journal on Numerical Analysis, 11 (5):1069–1085, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Renaut-Williamson:1989:FDR

[RW89a] R. A. Renaut-Williamson.

Full discretisations of u_{tt} = u_{xx} and rational approximations to cosh μz. SIAM Journal on Numerical Analysis, 26 (2):338–347, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Renaut-Williamson:1989:SRA

[RW89b] R. A. Renaut-Williamson.

Semidiscretisations of u_{tt} = u_{xx} and rational approximations to (ln z)^{2}. SIAM Journal on Numerical Analysis, 26 (2):320–337, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rosen:1995:MTA

[RW95] I. G. Rosen and Chunming Wang. A multilevel technique for the approximate solution of operator Lyapunov and algebraic Riccati equations. SIAM Journal on Numerical Analysis, 32(2): 514–541, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rulla:1996:ORC

[RW96] Jim Rulla and Noel J. Walkington. Optimal rates of convergence for degenerate parabolic problems in two dimensions. SIAM Journal on Numerical Analysis, 33(1): 56–67, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Robinson:1986:BML

[RY86] P. D. Robinson and P. K. Yuen. Bivariational methods for linear integral equations with nonsymmetric kernels. SIAM Journal on Numerical Analysis, 23(6):1230– 1240, December 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Settari:1973:GAC

[SA73] A. Settari and K. Aziz. A generalization of the additive correction methods for the iterative solution of matrix equations. SIAM Journal on Numerical Analysis, 10 (3):506–521, June 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1980:RCL

[Saa80] Y. Saad. On the rates of convergence of the Lanczos and the block Lanczos methods. SIAM Journal on Numerical Analysis, 17(5):687– 706, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1982:LBA

[Saa82]  Y. Saad.  The Lanczos

biorthogonalization algorithm and other oblique projection methods for solving large unsymmetric systems. SIAM Journal on Numerical Analysis, 19(3):485–506, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1983:ISI

[Saa83] Youcef Saad. Iterative solution of indefinite symmetric linear systems by methods using orthogonal polynomials over two disjoint intervals. SIAM Journal on Numerical Analysis, 20(4):784– 811, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1987:LSP

[Saa87] Youcef Saad. Least squares polynomials in the complex plane and their use for solving nonsymmetric linear systems. SIAM Journal on Numerical Analysis, 24(1):155– 169, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saad:1992:ASK

[Saa92] Y. Saad. Analysis of some Krylov subspace approximations to the matrix exponential operator. SIAM Journal on Numerical Analysis, 29(1): 209–228, February 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sabac:1997:OCR

[Şab97] Florin Şabac. The optimal convergence rate of monotone finite difference methods for hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 34(6): 2306–2318, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29347.

Sackett:1971:IFB

[Sac71] Gary G. Sackett. An implicit free boundary problem for the heat equation. SIAM Journal on Numerical Analysis, 8 (1):80–96, March 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sammon:1982:CES

[Sam82] Peter H. Sammon. Convergence estimates for semidiscrete parabolic equation approximations. SIAM Journal on Numerical Analysis, 19(1): 68–92, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sammon:1983:FDA

[Sam83] Peter Sammon. Fully discrete approximation methods for parabolic problems with nonsmooth initial data. SIAM Journal on Numerical Analysis, 20(3):437–470, June 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sammon:1986:NAM

[Sam86] Peter H. Sammon. Numerical approximations for a miscible displacement process in porous media. SIAM Journal on Numerical Analysis, 23 (3):508–542, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sanders:1985:MGM

[San85] Richard Sanders. The moving grid method for nonlinear hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 22(4):713– 728, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sandri:1994:FEA

[San94] Dominique Sandri. Finite element approximation of viscoelastic fluid flow: Existence of approximate solutions and error bounds. continuous approximation of the stress. SIAM Journal on Numerical Analysis, 31(2): 362–377, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sandboge:1998:QEC

[San98] Robert Sandboge. Quantitative error control for finite element methods for one-dimensional compressible flow. SIAM Journal on Numerical Analysis, 35(5): 2014–2034, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32162.

Saranen:1990:PMC

[Sar90] Jukka Saranen. Projection methods for a class of Hammerstein equations. SIAM Journal on Numerical Analysis, 27(6):1445–1449, December 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sasaki:1978:MMC

[Sas78] Tateaki Sasaki. Multidimensional Monte Carlo integration based on factorized approximation functions. SIAM Journal on Numerical Analysis, 15(5):938–952, October 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saylor:1967:NEC

[Say67] Richard Saylor. Numerical elliptic continuation. SIAM Journal on Numerical Analysis, 4(4):575–581, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saylo:1974:SOS

[Say74] Paul E. Saylo. Second order strongly implicit symmetric factorization methods for the solution of elliptic difference equations. SIAM Journal on Numerical Analysis, 11(5):894–908, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sameh:1977:STS

[SB77a] Ahmed H. Sameh and Richard P.

Brent. Solving triangular systems on a parallel computer. SIAM Journal on Numerical Analysis, 14(6):1101– 1113, December 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1977:IFP

[SB77b] A. H. Stroud and David L. Barrow. Integration formulas for the Poisson equations. SIAM Journal on Numerical Analysis, 14(5):888– 903, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Szymczak:1984:AEE

[SB84] W. G. Szymczak and I. Babuška. Adaptivity and error estimation for the finite element method applied to convection diffusion problems. SIAM Journal on Numerical Analysis, 21(5):910–954, October 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shelley:1988:OPA

[SB88] M. J. Shelley and G. R. Baker. Order-preserving approximations to successive derivatives of periodic functions by iterated splines. SIAM Journal on Numerical Analysis, 25(6): 1442–1452, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sidilkover:1993:MSS

[SB93] David Sidilkover and Achi Brandt. Multigrid solution to steady-state two-dimensional conservation laws. SIAM Journal on Numerical Analysis, 30(1):249–274, February 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1991:SGM

[SBBL91] Ralph C. Smith, Gary A.

Bogar, Kenneth L. Bowers, and John Lund. The Sinc– Galerkin method for fourthorder differential equations. SIAM Journal on Numerical Analysis, 28(3):760–788, June 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1972:PEE

[SC72] A. H. Stroud and Kwan-Wei Chen. Peano error estimates for Gauss–Laguerre quadrature formulas. SIAM Journal on Numerical Analysis, 9 (2):333–340, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schwetlick:1987:HOP

[SC87] Hubert Schwetlick and Jürgen Cleve. Higher order predictors and adaptive steplength control in path following algorithms. SIAM Journal on Numerical Analysis, 24(6):1382– 1393, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schoenberg:1965:MLD

[Sch65] I. J. Schoenberg. On monosplines of least deviation and best quadrature formulae. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):144–170, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Schoenberg:1966:MLS

[Sch66] I. J. Schoenberg. On monosplines of least square deviation and best quadrature formulae II. SIAM Journal on Numerical Analysis, 3(2): 321–328, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schechter:1968:RMC

[Sch68a] Samuel Schechter. Relaxation methods for convex problems. SIAM Journal on Numerical Analysis, 5(3):601– 612, September 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schumaker:1968:UAC

[Sch68b] Larry Schumaker. Uniform approximation by Chebyshev spline functions. II: Free knots. SIAM Journal on Numerical Analysis, 5(4):647– 656, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schultz:1969:ATM

[Sch69a] Martin H. Schultz. Approximation theory of multivariate spline functions in Sobolev spaces. SIAM Journal on Numerical Analysis, 6(4):570– 582, December 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schultz:1969:ATE

[Sch69b]  Martin H. Schultz.  L^{2}-

approximation theory of even order multivariate splines. SIAM Journal on Numerical Analysis, 6(3):467–475, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schultz:1969:MATb

[Sch69c]  Martin H. Schultz.  L^{2}-

multivariate approximation theory. SIAM Journal on Numerical Analysis, 6(2): 184–209, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schultz:1969:MATa

[Sch69d] Martin H. Schultz. L^{∞}- multivariate approximation theory. SIAM Journal on Numerical Analysis, 6(2):

161–183, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schultz:1969:RRG

[Sch69e] Martin H. Schultz. Rayleigh– Ritz–Galerkin methods for multidimensional problems. SIAM Journal on Numerical Analysis, 6(4):523–538, December 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schuurmann:1969:UUA

[Sch69f] F. J. Schuurmann. Uniqueness in the uniform approximation of a function and its derivatives. SIAM Journal on Numerical Analysis, 6(2): 305–315, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scherer:1970:BAC

[Sch70] Karl Scherer. On the best approximation of continuous functions by splines. SIAM Journal on Numerical Analysis, 7(3):418–423, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schoen:1971:FSO

[Sch71a] Kenneth Schoen. Fifth and sixth order PECE algorithms with improved stability properties. SIAM Journal on Numerical Analysis, 8(2):

244–248, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schultz:1971:EBR

[Sch71b] Martin H. Schultz. L^{2} error bounds for the Rayleigh– Ritz–Galerkin method. SIAM Journal on Numerical Analysis, 8(4):737–748, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schryer:1972:CAS

[Sch72] N. L. Schryer. Constructive approximation of solutions to linear elliptic boundary value problems. SIAM Journal on Numerical Analysis, 9(4):546–572, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schelin:1973:MCS

[Sch73] Charles W. Schelin. Monotone convergence of the sornewton iterative technique. SIAM Journal on Numerical Analysis, 10(5):933–938, October 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scherer:1974:CGL

[Sch74] Karl Scherer. Characterization of generalized Lipschitz classes by best approximation with splines. SIAM Journal on Numerical Analysis, 11(2):

283–304, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schaback:1979:SEA

[Sch79] Robert Schaback. Suboptimal exponential approximations. SIAM Journal on Numerical Analysis, 16(6):1007– 1018, December 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schreiber:1980:FEM

[Sch80] Robert Schreiber. Finite element methods of high-order accuracy for singular twopoint boundary value problems with nonsmooth solutions. SIAM Journal on Numerical Analysis, 17(4):547– 566, August 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schonhage:1982:RAR

[Sch82] Arnold Schönhage. Rational approximation to e^{−x} and related L^{2}-problems. SIAM Journal on Numerical Analysis, 19(5):1067–1080, October 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schelin:1983:CZR

[Sch83a] Charles W. Schelin. Counting zeros of real polynomials within the unit disk. SIAM Journal on Numerical Analysis, 20(5):1023–1031, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schumaker:1983:SPQ

[Sch83b] Larry L. Schumaker. On shape preserving quadratic spline interpolation. SIAM Journal on Numerical Analysis, 20(4):854–864, August 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schwartz:1983:ERE

[Sch83c] Ira Bruce Schwartz. Estimating regions of existence of unstable periodic orbits using computer-based techniques. SIAM Journal on Numerical Analysis, 20(1):106– 120, February 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schwandt:1985:KLA

[Sch85] Hartmut Schwandt. Krawczyklike algorithms for the solution of systems of nonlinear equations. SIAM Journal on Numerical Analysis, 22(4): 792–810, August 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schwandt:1986:AGC

[Sch86] Hartmut Schwandt. Almost globally convergent interval methods for discretizations of nonlinear elliptic partial differential equations. SIAM Journal on Numerical Analysis, 23(2):304–324, April 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schumaker:1989:SSF

[Sch89a] Larry L. Schumaker. On super splines and finite elements. SIAM Journal on Numerical Analysis, 26(4):997– 1005, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schwandt:1989:CRT

[Sch89b] Hartmut Schwandt. Cyclic reduction for tridiagonal systems of equations with interval coefficients on vector computers. SIAM Journal on Numerical Analysis, 26(3): 661–680, June 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schmitt:1990:AAM

[Sch90a] Bernhard A. Schmitt. An algebraic approximation for the matrix exponential in singularly perturbed boundary value problems. SIAM Journal on Numerical Analysis, 27(1):51–66, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schochet:1990:RCS

[Sch90b] S. Schochet. The rate of convergence of spectral-viscosity methods for periodic scalar conservation laws. SIAM Journal on Numerical Analysis, 27(5):1142–1159, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schecter:1993:NCS

[Sch93] Stephen Schecter. Numerical computation of saddlenode homoclinic bifurcation points. SIAM Journal on Numerical Analysis, 30(4):1155– 1178, September 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schroll:1996:CIF

[Sch96] H. J. Schroll. Convergence of implicit finite difference methods applied to nonlinear mixed systems. SIAM Journal on Numerical Analysis, 33 (3):997–1013, June 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schaeffer:1998:CDS

[Sch98a] Jack Schaeffer. Convergence of a difference scheme for the Vlasov–Poisson–Fokker– Planck system in one dimension. SIAM Journal on Numerical Analysis, 35 (3):1149–1175, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30255.

Schneider:1998:IFD

[Sch98b] Claus B. Schneider. Inversion formulas for the discretized Hilbert transform on the unit circle. SIAM Journal on Numerical Analysis, 35(1):71–77, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28514.

Scott:1975:IBC

[Sco75] Ridgway Scott. Interpolated boundary conditions in the finite element method. SIAM Journal on Numerical Analysis, 12(3):404–427, June 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scott:1981:SSS

[Sco81] David S. Scott. Solving sparse symmetric generalized eigenvalue problems without factorization. SIAM Journal on Numerical Analysis, 18(1): 102–110, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scroggs:1991:PMD

[Scr91] Jeffrey S. Scroggs. A physically motivated domain decomposition for singularly perturbed equations. SIAM Journal on Numerical Analysis, 28(1):168–178, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scuderi:1998:CMG

[Scu98] L. Scuderi. A collocation method for the generalized airfoil equation for an airfoil with a flap. SIAM Journal on Numerical Analysis, 35 (5):1725–1739, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32440.

Stroud:1971:SST

[SCWM71] A. H. Stroud, Kwan-Wei Chen, Ping-Lei Wang, and Zunkwang Mao. Some second and third degree harmonic interpolation formulas. SIAM Journal on Numerical Analysis, 8(4):681–692, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sacks-Davis:1977:SSO

[SD77]  R. Sacks-Davis.  Solution

of stiff ordinary differential equations by a second derivative method. SIAM Journal on Numerical Analysis, 14(6): 1088–1100, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Santos:1988:FEM

[SDC88] Juan Enrique Santos, Jim Douglas, Jr., and Alberto Pedro Calderón. Finite element methods for a composite model in elastodynamics. SIAM Journal on Numerical Analysis, 25(3):513–532, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schreiber:1981:FEM

[SE81] Robert Schreiber and Stanley C. Eisenstat. Finite element methods for spherically symmetric elliptic equations. SIAM Journal on Numerical Analysis, 18(3):546–558, June 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Secrest:1965:EBI

[Sec65a] Don Secrest. Error bounds for interpolation and differentiation by the use of spline functions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):440– 447, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Secrest:1965:NIA

[Sec65b] Don Secrest. Numerical integration of arbitrarily spaced data and estimation of errors. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):52– 68, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Segethova:1972:NCH

[Seg72] Jitka Segethova. Numerical construction of the Hill functions. SIAM Journal on Numerical Analysis, 9(2): 199–204, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Seidman:1996:OFB

[Sei96] Thomas I. Seidman. Optimal filtering for the backward heat equation. SIAM Journal on Numerical Analysis, 33(1): 162–170, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scherzer:1993:OPP

[SEK93] O. Scherzer, H. W. Engl, and K. Kunisch. Optimal A posteriori parameter choice for Tikhonov regularization for solving nonlinear ill-posed problems. SIAM Journal on Numerical Analysis, 30(6): 1796–1838, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Semper:1992:CFE

[Sem92] Bill Semper. Conforming finite element approximations for a fourth-order singular perturbation problem. SIAM Journal on Numerical Analysis, 29(4):1043–1058, August 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sermer:1983:GME

[Ser83] Pavol Sermer. A Galerkin method for elliptic-hyperbolic type equations. SIAM Journal on Numerical Analysis, 20 (3):471–484, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Serra:1999:HCB

[Ser99] Stefano Serra. How to choose the best iterative strategy for symmetric Toeplitz systems. SIAM Journal on Numerical Analysis, 36(4): 1078–1103, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31186.

Sewell:1966:ILC

[Sew66a] W. E. Sewell. Integrated Lipschitz conditions and approximation in the mean by interpolating polynomials. SIAM Journal on Numerical Analysis, 3(2):329–343, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sewell:1966:RMR

[Sew66b] W. E. Sewell. Remarks on the mathematical researches of J. L. Walsh. SIAM Journal on Numerical Analysis, 3 (2):344–348, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1979:ALL

[SF79] David A. Smith and William F.

Ford. Acceleration of linear and logarithmic convergence. SIAM Journal on Numerical Analysis, 16(2): 223–240, April 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schnabel:1984:TMN

[SF84] Robert B. Schnabel and Paul D. Frank. Tensor methods for nonlinear equations. SIAM Journal on Numerical Analysis, 21(5):815– 843, October 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sidi:1986:ACV

[SFS86] Avram Sidi, William F. Ford, and David A. Smith. Acceleration of convergence of vector sequences. SIAM Journal on Numerical Analysis, 23(1): 178–196, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1968:SEI

[SG68] A. H. Stroud and Edward H.

Goit, Jr. Some extensions of integration formulas. SIAM Journal on Numerical Analysis, 5(2):243–251, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stein:1966:ASR

[SH66] F. Max Stein and R. G. Huffstutler. The approximate solution of Riccati’s equation. SIAM Journal on Numerical Analysis, 3(3):425– 434, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stuart:1995:ESL

[SH95] A. M. Stuart and A. R. Humphries. The essential stability of local error control for dynamical systems. SIAM Journal on Numerical Analysis, 32(6):1940–1971, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shampine:1966:MIT

[Sha66]  L. F. Shampine.  Mono-

tone iterations and two-sided convergence. SIAM Journal on Numerical Analysis, 3 (4):607–615, December 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shampine:1968:BVP

[Sha68] L. F. Shampine. Boundary value problems for ordinary differential equations. SIAM Journal on Numerical Analysis, 5(2):219–242, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shampine:1969:BVP

[Sha69] L. F. Shampine. Boundary value problems for ordinary differential equations. II: Patch bases and monotone methods. SIAM Journal on Numerical Analysis, 6 (3):414–431, September 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shanno:1970:PSM

[Sha70] D. F. Shanno. Parameter selection for modified Newton methods for function minimization. SIAM Journal on Numerical Analysis, 7 (3):366–372, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shafer:1974:QA

[Sha74] R. E. Shafer. On quadratic approximation. SIAM Journal on Numerical Analysis, 11 (2):447–460, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shanno:1978:CNC

[Sha78] D. F. Shanno. On the convergence of a new conjugate gradient algorithm. SIAM Journal on Numerical Analysis, 15(6):1247–1257, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shampine:1985:IRK

[Sha85a] Lawrence F. Shampine. Interpolation for Runge–Kutta methods. SIAM Journal on Numerical Analysis, 22 (5):1014–1027, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shapiro:1985:OBD

[Sha85b] Alexander Shapiro. Optimal block diagonal L_{2}-scaling of matrices. SIAM Journal on Numerical Analysis, 22(1): 81–94, February 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shary:1995:OSI

[Sha95] Sergey P. Shary. On optimal solution of interval linear equations. SIAM Journal on Numerical Analysis, 32 (2):610–630, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sherman:1978:NIM

[She78] Andrew H. Sherman. On newton-iterative methods for the solution of systems of nonlinear equations. SIAM Journal on Numerical Analysis, 15(4):755–771, August 1978.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shen:1992:EEP

[She92] Jie Shen. On error estimates of projection methods for Navier–Stokes equations: First-order schemes. SIAM Journal on Numerical Analysis, 29(1):57–77, February 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shen:1995:EEP

[She95] Jie Shen. On error estimates of the penalty method for unsteady Navier–Stokes equations. SIAM Journal on Numerical Analysis, 32(2): 386–403, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shen:1996:BFD

[She96]  Jian Shen.  A block fi-

nite difference scheme for second-order elliptic problems with discontinuous coefficients. SIAM Journal on Numerical Analysis, 33(2): 686–706, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shisha:1966:IPC

[Shi66] O. Shisha. On infrapolynomials with prescribed center of gravity of their zeros. SIAM Journal on Numerical Analysis, 3(2):349–354, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Showalter:1975:PEE

[Sho75] R. E. Showalter. A Priori error estimates for approximation of parabolic boundary value problems. SIAM Journal on Numerical Analysis, 12 (3):456–463, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shreve:1973:IEE

[Shr73] David C. Shreve. Interior estimates in l^{p} for elliptic difference operators. SIAM Journal on Numerical Analysis, 10 (1):69–80, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:1996:FDP

[SHR96] Weiwei Sun, Weizhang Huang, and Robert D. Russell. Finite difference preconditioning for solving orthogonal collocation equations for boundary value problems. SIAM Journal on Numerical Analysis, 33(6): 2268–2285, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24976.

Shubert:1972:SMS

[Shu72] Bruno O. Shubert. A sequential method seeking the global maximum of a function. SIAM Journal on Numerical Analysis, 9(3):379– 388, September 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strang:1983:BS

[SI83] Gilbert Strang and Arieh Iserles. Barriers to stability. SIAM Journal on Numerical Analysis, 20(6):1251– 1257, December 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sidi:1986:CSP

[Sid86] Avram Sidi. Convergence and stability properties of minimal polynomial and reduced rank extrapolation algorithms. SIAM Journal on Numerical Analysis, 23(1): 197–209, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sidi:1997:CCS

[Sid97] Avram Sidi. A complete convergence and stability theory for a generalized Richardson extrapolation process. SIAM Journal on Numerical Analysis, 34(5):1761– 1778, October 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27858.

Sigillito:1976:PIA

[Sig76] Vincent G. Sigillito. A Priori inequalities and approximate solutions of the first boundary value problem for Δ^{2}u = f. SIAM Journal on Numerical Analysis, 13(2): 251–260, April 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Simpson:1968:AME

[Sim68] R. Bruce Simpson. Approximation of the minimizing element for a class of functionals. SIAM Journal on Numerical Analysis, 5(1):26–41, March 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Simpson:1969:RRP

[Sim69] R. B. Simpson. The Rayleigh– Ritz process for the simplest problem in the calculus of variations. SIAM Journal on Numerical Analysis, 6(2): 258–271, June 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Simpson:1971:FDM

[Sim71] R. B. Simpson. Finite difference methods for mildly nonlinear eigenvalue problems. SIAM Journal on Numerical Analysis, 8(2):190–211, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Simpson:1975:MND

[Sim75] R. B. Simpson. A method for the numerical determination of bifurcation states of nonlinear systems of equations. SIAM Journal on Numerical Analysis, 12(3):439–451, June 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sinclair:1973:DEF

[Sin73] Annette Sinclair. Determination of extremal functions in H^{p} by a Fortran program. SIAM Journal on Numerical Analysis, 10(1):137– 146, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sincovec:1977:REH

[Sin77] R. F. Sincovec. On the relative efficiency of higher order collocation methods for solving two-point boundary value problems. SIAM Journal on Numerical Analysis, 14(1): 112–123, March 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Skeel:1977:CNM

[SJ77] R. D. Skeel and L. W. Jackson. Consistency of Nordsieck methods. SIAM Journal on Numerical Analysis, 14 (5):910–924, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Skeel:1983:SVS

[SJ83] R. D. Skeel and L. W. Jackson. The stability of variable-stepsize Nordsieck methods. SIAM Journal on Numerical Analysis, 20(4): 840–853, August 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sloan:1987:LMM

[SK87] Ian H. Sloan and Philip J. Kachoyan. Lattice methods for multiple integration: theory, error analysis and examples. SIAM Journal on Numerical Analysis, 24(1):116– 128, February 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shroff:1993:SUP

[SK93] Gautam M. Shroff and Herbert B. Keller. Stabilization of unstable procedures: The recursive projection method. SIAM Journal on Numerical Analysis, 30(4):1099– 1120, September 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Skeel:1976:AFS

[Ske76] Robert Skeel. Analysis of fixed-stepsize methods. SIAM Journal on Numerical Analysis, 13(5):664–685, October 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Skeel:1981:EER

[Ske81] Robert D. Skeel. Effect of equilibration on residual size for partial pivoting. SIAM Journal on Numerical Analysis, 18(3):449–454, June 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Skeel:1982:TFP

[Ske82] Robert D. Skeel. A theoretical framework for proving accuracy results for deferred corrections. SIAM Journal on Numerical Analysis, 19(1): 171–196, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Skeel:1986:OAD

[Ske86] Robert D. Skeel. The order of accuracy for deferred corrections using uncentered end formulas. SIAM Journal on Numerical Analysis, 23 (2):393–402, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sidi:1983:PPT

[SL83a] Avram Sidi and David Levin. Prediction properties of the ttransformation. SIAM Journal on Numerical Analysis, 20 (3):589–598, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sidi:1983:ZSP

[SL83b] Avram Sidi and Doron S. Lubinsky. On the zeros of some polynomials that arise in numerical quadrature and convergence acceleration. SIAM Journal on Numerical Analysis, 20(2): 400–405, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Slepian:1968:NMD

[Sle68] David Slepian. A numerical method for determining the eigenvalues and eigenfunctions of analytic kernels. SIAM Journal on Numerical Analysis, 5(3):586– 600, September 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Slepian:1982:SCI

[Sle82]  David Slepian.  Solution

of certain integral equations with difference kernels. SIAM Journal on Numerical Analysis, 19(3):614–622, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sloan:1976:IGM

[Slo76]  Ian H. Sloan.  Iterated

Galerkin method for eigenvalue problems. SIAM Journal on Numerical Analysis, 13(5):753–760, October 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schmidt:1975:BSA

[SLW75] Eckard Schmidt, Peter Lancaster, and David Watkins. Bases of splines associated with constant coefficient differential operators. SIAM Journal on Numerical Analysis, 12(4):630–645, September 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sermer:1981:LSM

[SM81] Pavol Sermer and Rudolf Mathon. Least-squares methods for mixed-type equations. SIAM Journal on Numerical Analysis, 18(4): 705–723, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saito:1996:SAN

[SM96] Yoshihiro Saito and Taketomo Mitsui. Stability analysis of numerical schemes for stochastic differential equations. SIAM Journal on Numerical Analysis, 33(6):

2254–2267, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/22840.

Small:1988:GFM

[Sma88] R. D. Small. The generating function method of nonlinear approximation. SIAM Journal on Numerical Analysis, 25(1):235–244, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1968:CEA

[Smi68] Julius Smith. The coupled equation approach to the numerical solution of the biharmonic equation by finite differences. I. SIAM Journal on Numerical Analysis, 5(2): 323–339, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1970:CEA

[Smi70] Julius Smith. The coupled equation approach to the numerical solution of the biharmonic equation by finite differences. II. SIAM Journal on Numerical Analysis, 7(1): 104–111, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1973:ASF

[Smi73] Julius Smith. On the approximate solution of the first boundary value problem for ∇^{4}u = f. SIAM Journal on Numerical Analysis, 10 (6):967–982, December 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1981:BLA

[Smi81] Donald R. Smith. On boundary layer approximations in the numerical solution of certain stiff boundary value problems. SIAM Journal on Numerical Analysis, 18(3): 377–380, June 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smooke:1983:EEP

[Smo83] Mitchell D. Smooke. Error estimates for piecewise perturbation series solutions of the radial Schrödinger equation. SIAM Journal on Numerical Analysis, 20(2): 279–295, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sneddon:1996:SOS

[Sne96] G. E. Sneddon. Second-order spectral differentiation matrices. SIAM Journal on Numerical Analysis, 33(6): 2468–2487, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26423.

Snider:1972:IEA

[Sni72] Arthur D. Snider. An improved estimate of the accuracy of trigonometric interpolation. SIAM Journal on Numerical Analysis, 9(3):505– 508, September 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stynes:1986:UAF

[SO86] Martin Stynes and Eugene O’Riordan. A uniformly accurate finite element method for a singular perturbation problem in conservative form. SIAM Journal on Numerical Analysis, 23(2): 369–375, April 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sofroniou:1997:SRK

[SO97] M. Sofroniou and W. Oevel. Symplectic Runge–Kutta schemes. I. Order conditions. SIAM Journal on Numerical Analysis, 34(5):2063–2086, October 1997. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28102.

Sobol:1979:ESS

[Sob79a] I. M. Sobol’. Erratum: “On the Systematic Search in a Hypercube” [SIAM J. Numer. Anal. 16 (1979), no. 5, 790–793, MR 80f:65072]. SIAM Journal on Numerical Analysis, 16(6):1080, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Sob79b].

Sobol:1979:SSH

[Sob79b] I. M. Sobol’. On the systematic search in a hypercube. SIAM Journal on Numerical Analysis, 16(5):790– 793, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Sob79a].

Sonar:1998:FPO

[Son98] Thomas Sonar. On families of pointwise optimal finite volume ENO approximations. SIAM Journal on Numerical Analysis, 35(6): 2350–2369, December 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31601.

Sorensen:1980:SCC

[Sor80] D. C. Sorensen. The Qsuperlinear convergence of a collinear scaling algorithm for unconstrained optimization. SIAM Journal on Numerical Analysis, 17(1):84– 114, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sorensen:1982:NMM

[Sor82] D. C. Sorensen. Newton’s method with a model trust region modification. SIAM Journal on Numerical Analysis, 19(2):409–426, April 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sharma:1968:AHB

[SP68] A. Sharma and J. Prasad. On Abel–Hermite–Birkhoff interpolation. SIAM Journal on Numerical Analysis, 5 (4):864–881, December 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schreiber:1988:BRT

[SP88] Robert Schreiber and Beresford N. Parlett. Block reflectors: Theory and computation. SIAM Journal on Numerical Analysis, 25(1):189– 205, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). Cited in [?].

Smith:1992:ONI

[SP92] William E. Smith and David Paget. Optimal nodes for interpolatory product integration. SIAM Journal on Numerical Analysis, 29(2):

586–600, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sweldens:1994:QFA

[SP94] Wim Sweldens and Robert Piessens. Quadrature formulae and asymptotic error expansions for wavelet approximations of smooth functions. SIAM Journal on Numerical Analysis, 31(4):1240– 1264, August 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Spanier:1971:NMP

[Spa71]  Jerome Spanier.  A new

multistage procedure for systematic variance reduction in Monte Carlo. SIAM Journal on Numerical Analysis, 8 (3):548–554, September 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Spinelli:1965:PES

[Spi65] R. A. Spinelli. Poisson equation on a sphere. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3): 489–499, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Spinelli:1966:NIL

[Spi66] R. A. Spinelli. Numerical inversion of a Laplace transform. SIAM Journal on Numerical Analysis, 3(4):636– 649, December 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Spinelli:1967:PES

[Spi67] R. A. Spinelli. Poisson equation on a sphere. II. SIAM Journal on Numerical Analysis, 4(3):465–473, September 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Spijker:1971:RRE

[Spi71] M. N. Spijker. Reduction of roundoff error by splitting of difference formulas. SIAM Journal on Numerical Analysis, 8(2):345–357, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Spouge:1994:CGD

[Spo94] John L. Spouge. Computation of the gamma, digamma, and trigamma functions. SIAM Journal on Numerical Analysis, 31(3):931–944, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1965:QFM

[SS65] A. H. Stroud and D. D. Stancu. Quadrature formulas with multiple Gaussian nodes. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):129– 143, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Swarztrauber:1973:DSD

[SS73] Paul N. Swarztrauber and Roland A. Sweet. The direct solution of the discrete Poisson equation on a disk. SIAM Journal on Numerical Analysis, 10(5):900–907, October 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Soderstrom:1974:NPI

[SS74] Torsten Söderström and G. W. Stewart. On the numerical properties of an iterative method for computing the Moore–Penrose generalized inverse. SIAM Journal on Numerical Analysis, 11 (1):61–74, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Smith:1980:PIR

[SS80] William E. Smith and Ian H.

Sloan. Product-integration rules based on the zeros of Jacobi polynomials. SIAM Journal on Numerical Analysis, 17(1):1–13, February 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sloan:1982:PIP

[SS82] Ian H. Sloan and William E.

Smith. Properties of interpolatory product integration rules. SIAM Journal on Numerical Analysis, 19(2): 427–442, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stephens:1983:EUE

[SS83] A. B. Stephens and G. R. Shubin. Existence and uniqueness for an exponentially derived switching scheme. SIAM Journal on Numerical Analysis, 20(5):885–889, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sanz-Serna:1988:NMP

[SS88] J. M. Sanz-Serna. A numerical method for a partial integro-differential equation. SIAM Journal on Numerical Analysis, 25(2): 319–327, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scott:1989:EMD

[SS89] Larkin B. Scott and L. Ridgway Scott. Efficient methods for data smoothing. SIAM Journal on Numerical Analysis, 26(3):681–692, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saranen:1993:QMS

[SS93] J. Saranen and L. Schroderus. Quadrature methods for strongly elliptic equations of negative order on smooth closed curves. SIAM Journal on Numerical Analysis, 30(6): 1769–1795, December 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saranen:1995:SDM

[SS95] J. Saranen and L. Schroderus.

Some discrete methods for boundary integral equations on smooth closed curves. SIAM Journal on Numerical Analysis, 32(5):1535– 1564, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schwab:1996:OVA

[SS96a] Christoph Schwab and Manil Suri. The optimal p-version approximation of singularities on polyhedra in the boundary element method. SIAM Journal on Numerical Analysis, 33 (2):729–759, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shub:1996:CBT

[SS96b] Michael Shub and Steve Smale. Complexity of Bezout’s theorem. IV. probability of success and extensions. SIAM Journal on Numerical Analysis, 33(1):128– 148, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stenberg:1997:EAM

[SS97] Rolf Stenberg and Manil Suri. An h-p error analysis of MITC plate elements. SIAM Journal on Numerical Analysis, 34(2):544–568, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27848.

Schiff:1999:NAN

[SS99] Jeremy Schiff and S. Shnider. A natural approach to the numerical integration of Riccati differential equations. SIAM Journal on Numerical Analysis, 36(5):1392– 1413, October 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30794.

Sanz-Serna:1984:ICN

[SSA84] J. M. Sanz-Serna and L. Abia. Interpolation of the coefficients in nonlinear elliptic Galerkin procedures. SIAM Journal on Numerical Analysis, 21(1):77–83, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sanz-Serna:1991:OCC

[SSA91] J. M. Sanz-Serna and L. Abia. Order conditions for canonical Runge–Kutta schemes. SIAM Journal on Numerical Analysis, 28(4):1081– 1096, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shultz:1985:FTR

[SSB85] Gerald A. Shultz, Robert B.

Schnabel, and Richard H. Byrd. A family of trustregion-based algorithms for unconstrained minimization with strong global convergence properties. SIAM Journal on Numerical Analysis, 22(1):47–67, February 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schatz:1996:SFE

[SSW96] A. H. Schatz, I. H. Sloan, and L. B. Wahlbin. Superconvergence in finite element methods and meshes that are locally symmetric with respect to a point. SIAM Journal on Numerical Analysis, 33(2): 505–521, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saunders:1988:TCG

[SSY88] M. A. Saunders, H. D. Simon, and E. L. Yip. Two conjugategradient-type methods for unsymmetric linear equations. SIAM Journal on Numerical Analysis, 25(4): 927–940, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sweezy:1967:ESC

[ST67] W. B. Sweezy and W. J. Thron. Estimates of the speed of convergence of certain continued fractions. SIAM Journal on Numerical Analysis, 4 (2):254–270, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schumaker:1969:APH

[ST69] L. L. Schumaker and G. D. Taylor. On approximation by polynomials having restricted ranges. II. SIAM Journal on Numerical Analysis, 6(1): 31–36, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sharma:1976:CPS

[ST76] A. Sharma and J. Tzimbalario. Cardinal t-perfect Lsplines. SIAM Journal on Numerical Analysis, 13(6):915– 922, December 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saigal:1978:EAT

[ST78] R. Saigal and M. J. Todd. Efficient acceleration techniques for fixed point algorithms. SIAM Journal on Numerical Analysis, 15(5):997– 1007, October 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sloan:1986:TDI

[ST86] I. H. Sloan and V. Thomée.

Time discretization of an integro-differential equation of parabolic type. SIAM Journal on Numerical Analysis, 23(5):1052–1061, October 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sorensen:1991:OEC

[ST91] D. C. Sorensen and Ping Tak Peter Tang. On the orthogonality of eigenvectors computed by divide and conquer techniques. SIAM Journal on Numerical Analysis, 28(6):1752–1775, December 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sugiura:1992:PIQ

[ST92] Hiroshi Sugiura and Tatsuo Torii. Polynomial interpolation on quasi-equidistributed nodes on the unit disk. SIAM Journal on Numerical Analysis, 29(4):1154–1165, August 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shen:1995:NGM

[ST95] Jie Shen and Roger Temam.

Nonlinear Galerkin method using Chebyshev and Legendre polynomials. I. the onedimensional case. SIAM Journal on Numerical Analysis, 32 (1):215–234, February 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stancu:1964:RCL

[Sta64] D. D. Stancu. The remainder of certain linear approximation formulas in two variables. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):137– 163, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Starke:1991:OAD

[Sta91] Gerhard Starke. Optimal alternating direction implicit parameters for nonsymmetric systems of linear equations. SIAM Journal on Numerical Analysis, 28(5):1431– 1445, October 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Starke:1999:MBF

[Sta99]  Gerhard Starke.  Multi-

level boundary functionals for least-squares mixed finite element methods. SIAM Journal on Numerical Analysis, 36 (4):1065–1077, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32980.

Stetter:1965:SSW

[Ste65a] Hans J. Stetter. A study of strong and weak stability in discretization algorithms. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):265–280, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Steward:1965:PTS

[Ste65b] Donald V. Steward. Partitioning and tearing systems of equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 345–365, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Stenger:1968:KPE

[Ste68] Frank Stenger. Kronecker product extensions of linear operators. SIAM Journal on Numerical Analysis, 5(2): 422–435, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1970:CER

[Ste70] N. F. Stewart. Certain equivalent requirements of approximate solutions of x′ = f(t, x). SIAM Journal on Numerical Analysis, 7(2):256–270, June 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stepleman:1971:FDA

[Ste71a] Robert S. Stepleman. Finitedimensional analogues of variational problems in the plane. SIAM Journal on Numerical Analysis, 8(1):11–23, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stetter:1971:LEG

[Ste71b] Hans J. Stetter. Local estimation of the global discretization error. SIAM Journal on Numerical Analysis, 8 (3):512–523, September 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1971:EBA

[Ste71c] G. W. Stewart. Error bounds for approximate invariant subspaces, of closed linear operator. SIAM Journal on Numerical Analysis, 8(4):796– 808, December 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1972:SEP

[Ste72] G. W. Stewart. On the sensitivity of the eigenvalue problem ax = λbx. SIAM Journal on Numerical Analysis, 9 (4):669–686, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1974:CMI

[Ste74] G. W. Stewart. The convergence of multipoint iterations to multiple zeros. SIAM Journal on Numerical Analysis, 11(6):1105–1120, December 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stenger:1975:AFW

[Ste75a] Frank Stenger. An analytic function which is an approximate characteristic function. SIAM Journal on Numerical Analysis, 12(2): 239–254, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stepleman:1975:CLC

[Ste75b] Robert S. Stepleman. A characterization of local convergence for fixed point iterations in R^{1}. SIAM Journal on Numerical Analysis, 12 (6):887–894, December 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stephens:1976:CRP

[Ste76] A. B. Stephens. The convergence of the residual for projective approximations. SIAM Journal on Numerical Analysis, 13(4):607–614, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1977:PBF

[Ste77] G. W. Stewart. Perturbation bounds for the QR factorization of a matrix. SIAM Journal on Numerical Analysis, 14 (3):509–518, June 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stetter:1979:IEE

[Ste79a] Hans J. Stetter. Interpolation and error estimation in Adams PC-codes. SIAM Journal on Numerical Analysis, 16(2):311–323, April 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Steuerwalt:1979:ECN

[Ste79b] Michael Steuerwalt. The existence, computation, and number of solutions of periodic parabolic problems. SIAM Journal on Numerical Analysis, 16(3):402–420, June 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1980:EGR

[Ste80] G. W. Stewart. The efficient generation of random orthogonal matrices with an application to condition estimators. SIAM Journal on Numerical Analysis, 17(3):403– 409 (loose microfiche supplement), June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Steihaug:1983:CGM

[Ste83] Trond Steihaug. The conjugate gradient method and trust regions in large scale optimization. SIAM Journal on Numerical Analysis, 20(3): 626–637, June 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stiefel:1964:MON

[Sti64]  E. Stiefel.  Methods —

old and new — for solving the Tchebycheff approximation problem. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??): 164–176, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Stone:1968:ISI

[Sto68]  Herbert L. Stone.  Iter-

ative solution of implicit approximations of multidimensional partial differential equations. SIAM Journal on Numerical Analysis, 5 (3):530–558, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stoer:1971:NSC

[Sto71] Josef Stoer. On the numerical solution of constrained least-squares problems. SIAM Journal on Numerical Analysis, 8(2):382–411, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stoutemyer:1973:NIS

[Sto73] David R. Stoutemyer. Numerical implementation of the Schwarz alternating procedure for elliptic partial differential equations. SIAM Journal on Numerical Analysis, 10 (2):308–326, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strang:1965:UPD

[Str65a] Gilbert Strang. Unbalanced polynomials and difference methods for mixed problems. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):46– 51, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Stroud:1965:EER

[Str65b] A. H. Stroud. Error estimates for Romberg quadrature. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):480– 488, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Stroud:1966:EQE

[Str66]  A. H. Stroud.  Estimat-

ing quadrature errors for functions with low continuity. SIAM Journal on Numerical Analysis, 3(3):420– 424, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1967:IFO

[Str67a] A. H. Stroud. Integration formulas and orthogonal polynomials. SIAM Journal on Numerical Analysis, 4(3):381– 389, September 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1967:SSD

[Str67b] A. H. Stroud. Some seventh degree integration formulas for symmetric regions. SIAM Journal on Numerical Analysis, 4(1):37–44, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strang:1968:CCD

[Str68] Gilbert Strang. On the construction and comparison of difference schemes. SIAM Journal on Numerical Analysis, 5(3):506–517, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1969:FDI

[Str69a] A. H. Stroud. A fifth degree integration formula for the n-simplex. SIAM Journal on Numerical Analysis, 6 (1):90–98, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1969:IFO

[Str69b] A. H. Stroud. Integration formulas and orthogonal polynomials for two variables. SIAM Journal on Numerical Analysis, 6(2):222–229, June 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1970:IFO

[Str70] A. H. Stroud. Integration formulas and orthogonal polynomials. II. SIAM Journal on Numerical Analysis, 7(2): 271–276, June 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1972:TFD

[Str72] A. H. Stroud. Two fifth degree harmonic interpolation formulas for the n-box. SIAM Journal on Numerical Analysis, 9(3):518–521, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1973:SCF

[Str73] A. H. Stroud. Some cubature formulas for the surface of the n-sphere. SIAM Journal on Numerical Analysis, 10(4):559–569, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strand:1974:TMR

[Str74] Otto Neall Strand. Theory and methods related to the singular-function expansion and Landweber’s iteration for integral equations of the first kind. SIAM Journal on Numerical Analysis, 11 (4):798–825, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strom:1975:LN

[Str75a] Torsten Strom. On logarithmic norms. SIAM Journal on Numerical Analysis, 12(5): 741–753, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stroud:1975:SIF

[Str75b] A. H. Stroud. Some interpolation formulas for the Neumann problem for the nsphere. SIAM Journal on Numerical Analysis, 12(4):593– 604, September 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strikwerda:1984:IMS

[Str84] John C. Strikwerda. An iterative method for solving finite difference approximations to the Stokes equations. SIAM Journal on Numerical Analysis, 21(3):447–458, June 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stummel:1979:GPT

[Stu79] Friedrich Stummel. The generalized patch test. SIAM Journal on Numerical Analysis, 16(3):449–471, June 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schroll:1997:EBS

[STW97] Hans Joachim Schroll, Aslak Tveito, and Ragnar Winther. An L^{1}-error bound for a semiimplicit difference scheme applied to a stiff system of conservation laws. SIAM Journal on Numerical Analysis, 34(3):1152–1166, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26885.

Stynes:1989:AUC

[Sty89] Martin Stynes. An adaptive uniformly convergent numerical method for a semilinear singular perturbation problem. SIAM Journal on Numerical Analysis, 26(2): 442–455, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Suli:1991:CFV

[Sül91]  Endre Süli.  Convergence

of finite volume schemes for Poisson’s equation on nonuniform meshes. SIAM Journal on Numerical Analysis, 28 (5):1419–1430, October 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:1979:FEM

[Sun79] Tze Chien Sun. A finite element method for random differential equations with random coefficients. SIAM Journal on Numerical Analysis, 16(6):1019–1035, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:1983:PAG

[Sun83] Ji Guang Sun. Perturbation analysis for the generalized singular value problem. SIAM Journal on Numerical Analysis, 20(3):611–625, June 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:1996:BIA

[Sun96] W. Sun. Block iterative algorithms for solving Hermite bicubic collocation equations. SIAM Journal on Numerical Analysis, 33(2): 589–601, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:1999:SAH

[Sun99] Weiwei Sun. Spectral analysis of Hermite cubic spline collocation systems. SIAM Journal on Numerical Analysis, 36(6):1962–1975, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/32272.

Schneider:1970:CPM

[SV70] Hans Schneider and Mathukumalli Vidyasagar. Crosspositive matrices. SIAM Journal on Numerical Analysis, 7(4):508–519, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swartz:1973:NLI

[SV73] Blair K. Swartz and Richard S. Varga. A note on lacunary interpolation by splines. SIAM Journal on Numerical Analysis, 10(3):443–447, June 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sharp:1994:CIR

[SV94] P. W. Sharp and J. H. Verner. Completely imbedded Runge–Kutta pairs. SIAM Journal on Numerical Analysis, 31(4):1169–1190, August 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Saranen:1996:TCM

[SV96] J. Saranen and G. Vainkko. Trigonometric collocation methods with product integration for boundary integral equations on closed curves. SIAM Journal on Numerical Analysis, 33(4): 1577–1596, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26627.

Strand:1968:MRE

[SW68] Otto Neall Strand and Ed R.

Westwater. Minimum-RMS estimation of the numerical solution of a Fredholm integral equation of the first kind. SIAM Journal on Numerical Analysis, 5(2):287–295, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swartz:1974:RBG

[SW74a] Blair Swartz and Burton Wendroff. The relation between the Galerkin and collocation methods using smooth splines. SIAM Journal on Numerical Analysis, 11(5): 994–996, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swartz:1974:REF

[SW74b] Blair Swartz and Burton Wendroff. The relative efficiency of finite difference and finite element methods. I: Hyperbolic problems and splines. SIAM Journal on Numerical Analysis, 11(5): 979–993, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scott:1977:CSL

[SW77] Melvin R. Scott and Herman A. Watts. Computational solution of linear twopoint boundary value problems via orthonormalization. SIAM Journal on Numerical Analysis, 14(1):40–70, March 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sameh:1982:TMA

[SW82] Ahmed H. Sameh and John A. Wisniewski. A trace minimization algorithm for the generalized eigenvalue problem. SIAM Journal on Numerical Analysis, 19(6):1243– 1259, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shearer:1985:SCE

[SW85] J. M. Shearer and M. A. Wolfe. Some computable existence, uniqueness, and convergence tests for nonlinear systems. SIAM Journal on Numerical Analysis, 22(6): 1200–1207, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strikwerda:1988:EKM

[SW88] John C. Strikwerda and Bruce A. Wade. An extension of the Kreiss matrix theorem. SIAM Journal on Numerical Analysis, 25(6):1272– 1278, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Suli:1991:SMC

[SW91] Endre Süli and Antony Ware.

A spectral method of characteristics for hyperbolic problems. SIAM Journal on Numerical Analysis, 28(2): 423–445, April 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scherer:1994:CAC

[SW94a] R. Scherer and W. Wendler.

Complete algebraic characterization of A-Stable Runge– Kutta methods. SIAM Journal on Numerical Analysis, 31 (2):540–551, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Silvester:1994:FIS

[SW94b] David Silvester and Andrew Wathen. Fast iterative solution of stabilised Stokes systems Part II: Using general block preconditioners. SIAM Journal on Numerical Analysis, 31(5):1352–1367, October 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shen:1999:FAN

[SW99] Jie Shen and Shouhong Wang. A fast and accurate numerical scheme for the primitive equations of the atmosphere. SIAM Journal on Numerical Analysis, 36(3):719–737, June 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32477.

Swarztrauber:1972:NSD

[Swa72] Paul N. Swarztrauber. On the numerical solution of the Dirichlet problem for a region of general shape. SIAM Journal on Numerical Analysis, 9 (2):300–306, June 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swarztrauber:1974:DMD

[Swa74] Paul N. Swarztrauber. A direct method for the discrete solution of separable elliptic equations. SIAM Journal on Numerical Analysis, 11(6): 1136–1150, December 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swarztrauber:1979:SAD

[Swa79] Paul N. Swarztrauber. On the spectral approximation of discrete scalar and vector functions on the sphere. SIAM Journal on Numerical Analysis, 16(6):934–949, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swarztrauber:1981:AVF

[Swa81] Paul N. Swarztrauber. The approximation of vector functions and their derivatives on the sphere. SIAM Journal on Numerical Analysis, 18(2): 191–210, April 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Swartz:1988:CC

[Swa88] Blair Swartz. Conditioning collocation. SIAM Journal on Numerical Analysis, 25(1): 124–147, February 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strikwerda:1990:REB

[SWB90] John C. Strikwerda, Bruce A.

Wade, and Kenneth P. Bube. Regularity estimates up to the boundary for elliptic systems of difference equations. SIAM Journal on Numerical Analysis, 27(2):292–322, April 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Strehmel:1989:SAL

[SWC89] K. Strehmel, R. Weiner, and H. Claus. Stability analysis of linearly implicit onestep interpolation methods for stiff retarded differential equations. SIAM Journal on Numerical Analysis, 26 (5):1158–1174, October 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sweet:1974:GCR

[Swe74] Roland A. Sweet. A generalized cyclic reduction algorithm. SIAM Journal on Numerical Analysis, 11(3): 506–520, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sweet:1977:CRA

[Swe77] Roland A. Sweet. A cyclic reduction algorithm for solving block tridiagonal systems of arbitrary dimension. SIAM Journal on Numerical Analysis, 14(4):706–720, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sweby:1984:HRS

[Swe84] P. K. Sweby. High resolution schemes using flux limiters for hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 21(5): 995–1011, October 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shaw:1997:EES

[SWW97] S. Shaw, M. K. Warby, and J. R. Whiteman. Error estimates with sharp constants for a fading memory Volterra problem in linear solid viscoelasticity. SIAM Journal on Numerical Analysis, 34(3):1237–1254, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28434.

Symes:1980:NSI

[Sym80] W. W. Symes. Numerical stability in an inverse scattering problem. SIAM Journal on Numerical Analysis, 17(5): 707–732, October 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See erratum [Sym81].

Symes:1981:ENS

[Sym81] W. W. Symes. Erratum:

“Numerical Stability in an Inverse Scattering Problem” [SIAM J. Numer. Anal. 17 (1980), no. 5, 707–732, MR 81k:65135]. SIAM Journal on Numerical Analysis, 18(4): 751–752, August 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Sym80].

Shampine:1990:RCM

[SZ90] L. F. Shampine and W. Zhang.

Rate of convergence of multistep codes started by variation of order and stepsize. SIAM Journal on Numerical Analysis, 27(6):1506– 1518, December 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schmeiser:1999:CMM

[SZ99] Christian Schmeiser and Alexander Zwirchmayr. Convergence of moment methods for linear kinetic equations. SIAM Journal on Numerical Analysis, 36(1):74– 88, February 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30451.

Szyld:1988:CCI

[Szy88] Daniel B. Szyld. Criteria for combining inverse and Rayleigh quotient iteration. SIAM Journal on Numerical Analysis, 25(6):1369– 1375, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tadmor:1986:EAF

[Tad86] Eitan Tadmor. The exponential accuracy of Fourier and Chebyshev differencing methods. SIAM Journal on Numerical Analysis, 23(1):1– 10, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tadmor:1988:CTV

[Tad88] Eitan Tadmor. Convenient total variation diminishing conditions for nonlinear difference schemes. SIAM Journal on Numerical Analysis, 25 (5):1002–1014, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tadmor:1989:CSM

[Tad89]  Eitan Tadmor.  Conver-

gence of spectral methods for nonlinear conservation laws. SIAM Journal on Numerical Analysis, 26(1):30–44, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tadmor:1991:LEE

[Tad91] Eitan Tadmor. Local error estimates for discontinuous solutions of nonlinear hyperbolic equations. SIAM Journal on Numerical Analysis, 28(4):891–906, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Talay:1991:AUL

[Tal91] Denis Talay. Approximation of upper Lyapunov exponents of bilinear stochastic differential systems. SIAM Journal on Numerical Analysis, 28(4): 1141–1164, August 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tang:1998:CAO

[Tan98]  Tao Tang.  Convergence

analysis for operator-splitting methods applied to conservation laws with stiff source terms. SIAM Journal on Numerical Analysis, 35(5): 1939–1968, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30892.

Tapia:1969:WNM

[Tap69] Richard A. Tapia. The weak Newton method and boundary value problems. SIAM Journal on Numerical Analysis, 6(4):539–550, December 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tapia:1974:NMO

[Tap74a]  R. A. Tapia.  Newton’s

method for optimization problems with equality constraints. SIAM Journal on Numerical Analysis, 11(5): 874–886, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tapia:1974:NMP

[Tap74b]  R. A. Tapia.  Newton’s

method for problems with equality constraints. SIAM Journal on Numerical Analysis, 11(1):174–196, March 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tarzia:1996:NAH

[Tar96] Domingo Alberto Tarzia. Numerical analysis for the heat flux in a mixed elliptic problem to obtain a discrete steady-state two-phase Stefan problem. SIAM Journal on Numerical Analysis, 33 (4):1257–1265, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24018.

Tautenhahn:1996:EER

[Tau96]  Ulrich Tautenhahn.  Er-

ror estimates for regularization methods in Hilbert scales. SIAM Journal on Numerical Analysis, 33(6): 2120–2130, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26941.

Tavernini:1971:OSM

[Tav71] Lucio Tavernini. One-step methods for the numerical solution of Volterra functional differential equations. SIAM Journal on Numerical Analysis, 8(4):786–795, December 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tavernini:1977:FDA

[Tav77] Lucio Tavernini. Finite difference approximations for a class of semilinear Volterra evolution problems. SIAM Journal on Numerical Analysis, 14(5):931–949, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tavernini:1978:ASV

[Tav78] Lucio Tavernini. The approximate solution of Volterra differential systems with statedependent time lags. SIAM Journal on Numerical Analysis, 15(5):1039–1052, October 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Taylor:1968:APH

[Tay68] G. D. Taylor. On approximation by polynomials having restricted ranges. SIAM Journal on Numerical Analysis, 5 (2):258–268, June 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Taylor:1995:CSD

[Tay95] Mark Taylor. Cubature for the sphere and the discrete spherical harmonic transform. SIAM Journal on Numerical Analysis, 32(2): 667–670, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thiran:1982:WMS

[TD82] J. P. Thiran and P. Defert. Weak minimal H-sets for polynomials in two variables. SIAM Journal on Numerical Analysis, 19(5):1041– 1050, October 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tal-Ezer:1986:SMT

[TE86]  Hillel Tal-Ezer.  Spectral

methods in time for hyperbolic equations. SIAM Journal on Numerical Analysis, 23(1):11–26, February 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tal-Ezer:1989:SMT

[TE89]  Hillel Tal-Ezer.  Spectral

methods in time for parabolic problems. SIAM Journal on Numerical Analysis, 26(1):1– 11, February 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tai:1998:RCS

[TE98] Xue-Cheng Tai and Magne Espedal. Rate of convergence of some space decomposition methods for linear and nonlinear problems. SIAM Journal on Numerical Analysis, 35 (4):1558–1570, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29746.

Teng:1992:PMC

[Ten92] Zhen Huan Teng. Particle method and its convergence for scalar conservation laws. SIAM Journal on Numerical Analysis, 29(4):1020– 1042, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Teng:1994:AFS

[Ten94] Zhen Huan Teng. On the accuracy of fractional step methods for conservation laws in two dimensions. SIAM Journal on Numerical Analysis, 31(1):43–63, February 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tervo:1998:FSP

[Ter98]  Jouko Tervo.  An FEM

scheme of a PDE system from bioreactor theory with stability results. SIAM Journal on Numerical Analysis, 35(3):1230–1248, June 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29133.

Tewarson:1967:DMG

[Tew67] R. P. Tewarson. A direct method for generalized matrix inversion. SIAM Journal on Numerical Analysis, 4 (4):499–507, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Trefethen:1983:CFM

[TG83] Lloyd N. Trefethen and Martin H. Gutknecht. The Carathéodory–Fejér method for real rational approximation. SIAM Journal on Numerical Analysis, 20(2): 420–436, April 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tourigny:1998:NMM

[TH98] Yves Tourigny and Frank Hülsemann. A new moving mesh algorithm for the finite element solution of variational problems. SIAM Journal on Numerical Analysis, 35 (4):1416–1438, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31393.

Thatcher:1978:FEM

[Tha78] R. W. Thatcher. On the finite element method for unbounded regions. SIAM Journal on Numerical Analysis, 15 (3):466–477, June 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thomee:1967:GUS

[Tho67] Vidar Thomee. Generally unconditionally stable difference operators. SIAM Journal on Numerical Analysis, 4 (1):55–69, March 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thomee:1968:DIS

[Tho68a]  Vidar Thomee.  Discrete

interior Schauder estimates for elliptic difference operators. SIAM Journal on Numerical Analysis, 5(3):626– 645, September 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thompson:1968:SFD

[Tho68b] R. J. Thompson. On some functional differential equations: Existence of solutions and difference approximations. SIAM Journal on Numerical Analysis, 5 (3):475–487, September 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thompson:1976:CEE

[Tho76] Russell C. Thompson. Convergence and error estimates for the method of lines for certain nonlinear elliptic and elliptic-parabolic equations. SIAM Journal on Numerical Analysis, 13(1):27–43, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Twizell:1984:MMP

[TK84] E. H. Twizell and A. Q. M.

Khaliq. Multiderivative methods for periodic initial value problems. SIAM Journal on Numerical Analysis, 21(1): 111–122, February 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tian:1996:NSL

[TK96] Hong-Jiong Tian and Jiao-Xun Kuang. The numerical stability of linear multistep methods for delay differential equations with many delays. SIAM Journal on Numerical Analysis, 33(3):883–889, June 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tin:1994:IMS

[TKJ94] S.-K. Tin, N. Kopell, and C. K. R. T. Jones. Invariant manifolds and singularly perturbed boundary value problems. SIAM Journal on Numerical Analysis, 31(6):1558– 1576, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tallec:1998:NND

[TMV98] Patrick Le Tallec, Jan Mandel, and Marina Vidrascu. A Neumann–Neumann domain decomposition algorithm for solving plate and shell problems. SIAM Journal on Numerical Analysis, 35(2):836–867, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29101.

Toint:1979:SCA

[Toi79] Ph. Toint. On the superlinear convergence of an algorithm for solving a sparse minimization problem. SIAM Journal on Numerical Analysis, 16(6):1036–1045, December 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Traub:1965:PDQ

[Tra65] J. F. Traub. The principle of differentiated quotients. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(3):437–439, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Trangenstein:1977:FEM

[Tra77] John A. Trangenstein. A finite element method for the Tricomi problem in the elliptic region. SIAM Journal on Numerical Analysis, 14(6): 1066–1077, December 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tran:1996:LEE

[Tra96] Thanh Tran. Local error estimates for the Galerkin method applied to strongly elliptic integral equations on open curves. SIAM Journal on Numerical Analysis, 33 (4):1484–1493, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25140.

Trench:1972:SCD

[Tre72] William F. Trench. Stability of a class of discrete minimum variance smoothing formulas. SIAM Journal on Numerical Analysis, 9(2):307–315, June 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Trefethen:1983:CEC

[Tre83] Lloyd N. Trefethen. Circularity of the error curve and sharpness of the CF method in complex Chebyshev approximation. SIAM Journal on Numerical Analysis, 20(6): 1258–1263, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Trummer:1984:MSI

[Tru84] Manfred R. Trummer. A method for solving ill-posed linear operator equations. SIAM Journal on Numerical Analysis, 21(4):729– 737, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Trummer:1986:EIC

[Tru86] Manfred R. Trummer. An efficient implementation of a conformal mapping method based on the Szegő kernel. SIAM Journal on Numerical Analysis, 23(4):853– 872, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tam:1993:SPE

[TS93] Hon-Wah Tam and Robert D. Skeel. Stability of parallel explicit ODE methods. SIAM Journal on Numerical Analysis, 30(4):1179–1192, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tsao:1975:NIH

[Tsa75] Nai-Kuan Tsao. A note on implementing the Householder transformation. SIAM Journal on Numerical Analysis, 12(1):53–58, March 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tseng:1996:MEA

[Tse96] Ching-Yih Tseng. A multipleexchange algorithm for complex Chebyshev approximation by polynomials on the unit circle. SIAM Journal on Numerical Analysis, 33 (5):2017–2049, October 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25497.

Trefethen:1987:IPS

[TT87] Lloyd N. Trefethen and Manfred R. Trummer. An instability phenomenon in spectral methods. SIAM Journal on Numerical Analysis, 24 (5):1008–1023, October 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tang:1995:EBF

[TT95] Tao Tang and Zhen Huan Teng. Error bounds for fractional step methods for conservation laws with source terms. SIAM Journal on Numerical Analysis, 32(1):110– 127, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tadmor:1999:PEE

[TT99] Eitan Tadmor and Tao Tang.

Pointwise error estimates for scalar conservation laws with piecewise smooth solutions. SIAM Journal on Numerical Analysis, 36(6): 1739–1758, December 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33399.

Tucker:1969:SNC

[Tuc69] T. S. Tucker. Stability of nonlinear computing schemes. SIAM Journal on Numerical Analysis, 6(1):72–81, March 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Turkel:1977:CMH

[Tur77] Eli Turkel. Composite methods for hyperbolic equations. SIAM Journal on Numerical Analysis, 14(4):744– 759, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tobiska:1996:ASD

[TV96] Lutz Tobiska and Rüdiger Verfürth. Analysis of a streamline diffusion finite element method for the Stokes and Navier–Stokes equations. SIAM Journal on Numerical Analysis, 33(1):107–127, February 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Taylor:1970:CBR

[TW70] G. D. Taylor and M. J. Winter. Calculation of best restricted approximations. SIAM Journal on Numerical Analysis, 7(2):248–255, June 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thomee:1974:CEG

[TW74] Vidar Thomee and Burton Wendroff. Convergence estimates for Galerkin methods for variable coefficient initial value problems. SIAM Journal on Numerical Analysis, 11(5):1059–1068, October 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thomee:1975:GMS

[TW75] Vidar Thomee and Lars Wahlbin. On Galerkin methods in semilinear parabolic problems. SIAM Journal on Numerical Analysis, 12(3): 378–389, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tapia:1988:PNM

[TW88] R. A. Tapia and David L.

Whitley. The projected Newton method has order 1 + \sqrt{2} for the symmetric eigenvalue problem. SIAM Journal on Numerical Analysis, 25(6): 1376–1382, December 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tveito:1993:EEF

[TW93] Aslak Tveito and Ragnar Winther. An error estimate for a finite difference scheme approximating a hyperbolic system of conservation laws. SIAM Journal on Numerical Analysis, 30(2): 401–424, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Thomee:1989:SGP

[TXZ89] Vidar Thomée, Jin Chao Xu, and Nai Ying Zhang. Superconvergence of the gradient in piecewise linear finiteelement approximation to a parabolic problem. SIAM Journal on Numerical Analysis, 26(3):553–573, June 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Teng:1995:CVE

[TYZ95] Zhen Huan Teng, Lung An Ying, and Ping Wen Zhang. Convergence of the variableelliptic-vortex method for Euler equations. SIAM Journal on Numerical Analysis, 32 (3):754–774, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Urabe:1968:RED

[Ura68] Minoru Urabe. Roundoff error distribution in fixedpoint multiplication and a remark about the rounding rule. SIAM Journal on Numerical Analysis, 5(2):202–210, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Urbach:1991:CGM

[Urb91] H. P. Urbach. Convergence of the Galerkin method for two-dimensional electromagnetic problems. SIAM Journal on Numerical Analysis, 28 (3):697–710, June 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Usow:1967:ACC

[Uso67a] Karl H. Usow. On L_{1} approximation I: Computation for continuous functions and continuous dependence. SIAM Journal on Numerical Analysis, 4(1):70–88, March 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Usow:1967:AIC

[Uso67b] Karl H. Usow. On L_{1} approximation II: Computation for discrete functions and discretization effects. SIAM Journal on Numerical Analysis, 4(2):233–244, June 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Utumi:1996:OTS

[UTK96] Masaki Utumi, Ryuji Takaki, and Toshio Kawai. Optimal time step control for the numerical solution of ordinary differential equations. SIAM Journal on Numerical Analysis, 33(4): 1644–1653, August 1996.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26358.

Victory:1991:CTP

[VA91] H. D. Victory, Jr. and Edward J. Allen. The convergence theory of particle-in-cell methods for multidimensional Vlasov–Poisson systems. SIAM Journal on Numerical Analysis, 28(5):1207– 1241, October 1991. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vigo-Aguiar:1998:GPA

[VAF98] Jesus Vigo-Aguiar and José M. Ferrándiz. A general procedure for the adaptation of multistep algorithms to the integration of oscillatory problems. SIAM Journal on Numerical Analysis, 35 (4):1684–1708, August 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28676.

Vandergraft:1967:NMC

[Van67] James S. Vandergraft. Newton’s method for convex operators in partially ordered spaces. SIAM Journal on Numerical Analysis, 4(3):406– 432, September 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vandergraft:1972:APO

[Van72] James S. Vandergraft. Applications of partial orderings to the study of positive definiteness, monotonicity, and convergence of iterative methods for linear systems. SIAM Journal on Numerical Analysis, 9(1):97–104, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

VanLoan:1975:GME

[Van75] Charles F. Van Loan. A general matrix eigenvalue algorithm. SIAM Journal on Numerical Analysis, 12(6):819– 834, December 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Loan:1976:GSV

[Van76] Charles F. Van Loan. Generalizing the singular value decomposition. SIAM Journal on Numerical Analysis, 13 (1):76–83, March 1976. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Loan:1977:SME

[Van77] Charles Van Loan. The sensitivity of the matrix exponential. SIAM Journal on Numerical Analysis, 14(6):971– 981, December 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanderHouwen:1979:SRK

[van79] P. J. van der Houwen. Stabilized Runge–Kutta methods for second order differential equations without first derivatives. SIAM Journal on Numerical Analysis, 16(3): 523–537, June 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanderHouwen:1980:MSM

[van80] P. J. van der Houwen. Multistep splitting methods of high order for initial value problems. SIAM Journal on Numerical Analysis, 17(3): 410–427, June 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

VanderCruyssen:1981:SEG

[Van81a] P. Van Der Cruyssen. Stable evaluation of generalized continued fractions. SIAM Journal on Numerical Analysis, 18(5):871–881, October 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanderHouwen:1981:MNM

[van81b] P. J. van der Houwen. Modified Nyström methods for semidiscrete hyperbolic differential equations. SIAM Journal on Numerical Analysis, 18(6):1081–1097, December 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanVeldhuizen:1981:S

[van81c]  M. van Veldhuizen.  D-

stability. SIAM Journal on Numerical Analysis, 18(1): 45–64, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

vanderHouwen:1984:ISM

[van84] P. J. van der Houwen. Iterated splitting methods of high order for time-dependent partial differential equations. SIAM Journal on Numerical Analysis, 21(4):635– 656, August 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

VanLoan:1985:MWE

[Van85]  Charles Van Loan.  On

the method of weighting for equality-constrained leastsquares problems. SIAM Journal on Numerical Analysis, 22(5):851–864, October 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanPutten:1995:TDN

[van95] Maurice H. P. M. van Putten.

A two-dimensional numerical implementation of magnetohydrodynamics in divergence form. SIAM Journal on Numerical Analysis, 32(5):1504– 1518, October 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varga:1966:DMP

[Var66] Richard S. Varga. On a discrete maximum principle. SIAM Journal on Numerical Analysis, 3(2):355–359, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varga:1967:ASA

[Var67]  Richard S. Varga.  Ab-

stract: Some applications of numerical analysis to physical problems. SIAM Journal on Numerical Analysis, 4 (1):129, March 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varga:1970:MGS

[Var70] Richard S. Varga. Minimal Gerschgorin sets for partitioned matrices. SIAM Journal on Numerical Analysis, 7 (4):493–507, December 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1971:SDA

[Var71a] J. M. Varah. Stability of difference approximations to the mixed initial boundary value problems for parabolic systems. SIAM Journal on Numerical Analysis, 8(3):598– 615, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1971:SHO

[Var71b] J. M. Varah. Stability of high order accurate difference methods for parabolic equations with boundary conditions. SIAM Journal on Numerical Analysis, 8(3):569– 574, September 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1973:NSI

[Var73] J. M. Varah. On the numerical solution of ill-conditioned linear systems with applications to ill-posed problems. SIAM Journal on Numerical Analysis, 10(2): 257–267, April 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1976:ARC

[Var76] J. M. Varah. Alternate row and column elimination for solving certain linear systems. SIAM Journal on Numerical Analysis, 13(1):71–75, March 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1977:SBC

[Var77] J. M. Varah. On the stability of boundary conditions for separable difference approximations to parabolic equations. SIAM Journal on Numerical Analysis, 14(6):1114– 1125, December 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1978:SSL

[Var78] J. M. Varah. Stiffly stable linear multistep methods of extended order. SIAM Journal on Numerical Analysis, 15(6):1234–1246, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1979:STM

[Var79] J. M. Varah. On the separation of two matrices. SIAM Journal on Numerical Analysis, 16(2):216–222, April 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varah:1980:SRS

[Var80] J. M. Varah. Stability restrictions on second order, three level finite difference schemes for parabolic equations. SIAM Journal on Numerical Analysis, 17(2): 300–309, April 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Varma:1981:IRT

[Var81] A. K. Varma. An interpolatory rational trigonometric approximation. SIAM Journal on Numerical Analysis, 18(5):897–899, October 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vardi:1985:TRA

[Var85] Avi Vardi. A trust region algorithm for equality constrained minimization: convergence properties and implementation. SIAM Journal on Numerical Analysis, 22 (3):575–591, June 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vasseur:1999:KSS

[Vas99]  Alexis Vasseur.  Kinetic

semidiscretization of scalar conservation laws and convergence by using averaging lemmas. SIAM Journal on Numerical Analysis, 36(2):465–474, April 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31361.

Vatsya:1988:CCR

[Vat88] S. R. Vatsya. Convergence of conjugate residual-like methods to solve linear equations. SIAM Journal on Numerical Analysis, 25(4):957– 964, August 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vavasis:1996:SFE

[Vav96] Stephen A. Vavasis. Stable finite elements for problems with wild coefficients. SIAM Journal on Numerical Analysis, 33(3):890–916, June 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Veliov:1992:SOD

[Vel92] Vladimir Veliov. Secondorder discrete approximation to linear differential inclusions. SIAM Journal on Numerical Analysis, 29(2): 439–451, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verner:1970:QID

[Ver70] J. H. Verner. Quadratures for implicit differential equations. SIAM Journal on Numerical Analysis, 7(3):373– 385, September 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verner:1978:ERK

[Ver78] J. H. Verner. Explicit Runge– Kutta methods with estimates of the local truncation error. SIAM Journal on Numerical Analysis, 15(4): 772–790, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verner:1979:FIR

[Ver79]  J. H. Verner.  Families

of imbedded Runge–Kutta methods. SIAM Journal on Numerical Analysis, 16(5): 857–875, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verwer:1982:ARM

[Ver82] J. G. Verwer. An analysis of Rosenbrock methods for nonlinear stiff initial value problems. SIAM Journal on Numerical Analysis, 19(1):155– 170, February 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verfurth:1984:MAM

[Ver84] R. Verfürth. A multilevel algorithm for mixed problems. SIAM Journal on Numerical Analysis, 21(2): 264–271, April 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verfurth:1988:MAM

[Ver88] R. Verfürth. Multilevel algorithms for mixed problems. II. treatment of the minielement. SIAM Journal on Numerical Analysis, 25(2): 285–293, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verner:1990:CSR

[Ver90a] J. H. Verner. A contrast of some Runge–Kutta formula pairs. SIAM Journal on Numerical Analysis, 27(5):1332– 1344, October 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vertesi:1990:OLC

[Vér90b] P. Vértesi. Optimal Lebesgue constant for Lagrange interpolation. SIAM Journal on Numerical Analysis, 27 (5):1322–1331, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verner:1991:SRK

[Ver91] J. H. Verner. Some Runge– Kutta formula pairs. SIAM Journal on Numerical Analysis, 28(2):496–511, April 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verner:1993:DIH

[Ver93]  J. H. Verner.  Differen-

tiable interpolants for highorder Runge–Kutta methods. SIAM Journal on Numerical Analysis, 30(5):1446– 1466, October 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vulanovic:1993:CNA

[VF93] Relja Vulanović and Paul A. Farrell. Continuous and numerical analysis of a multiple boundary turning point problem. SIAM Journal on Numerical Analysis, 30(5):1400– 1418, October 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Victory:1986:FDM

[VG86] H. D. Victory, Jr. and Keshab Ganguly. On finitedifference methods for solving discrete-ordinates transport equations. SIAM Journal on Numerical Analysis, 23(1):78–108, February 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verschelde:1993:ACS

[VH93] Jan Verschelde and Ann Haegemans. The GBQalgorithm for constructing start systems of homotopies for polynomial systems. SIAM Journal on Numerical Analysis, 30(2): 583–594, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vozovoi:1996:AAF

[VIA96] L. Vozovoi, M. Israeli, and A. Averbuch. Analysis and application of Fourier– Gegenbauer method to stiff differential equations. SIAM Journal on Numerical Analysis, 33(5):1844–1863, October 1996. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/26359.

Victory:1980:CPD

[Vic80] H. D. Victory, Jr. Convergence properties of discreteordinates solutions for neutron transport in threedimensional media. SIAM Journal on Numerical Analysis, 17(1):71–83, February 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vila:1986:SGS

[Vil86] J. P. Vila. Simplified Godunov schemes for 2 × 2 systems of conservation laws. SIAM Journal on Numerical Analysis, 23(6):1173– 1192, December 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vila:1989:ACS

[Vil89]  J. P. Vila.  An analy-

sis of a class of secondorder accurate Godunov-type schemes. SIAM Journal on Numerical Analysis, 26(4): 830–853, August 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vincenti:1986:CCO

[Vin86] Graziano Vincenti. On the computation of the coefficients of s-orthogonal polynomials. SIAM Journal on Numerical Analysis, 23(6):1290– 1294, December 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vanaja:1990:IMF

[VK90] V. Vanaja and R. B. Kellogg. Iterative methods for a forward-backward heat equation. SIAM Journal on Numerical Analysis, 27(3): 622–635, June 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vonGolitschek:1992:ASP

[vL92] M. von Golitschek and W. A.

Light. Approximation by solutions of the planar wave equation. SIAM Journal on Numerical Analysis, 29(3): 816–830, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

VanHentenryck:1997:SPS

[VMK97] Pascal Van Hentenryck, David McAllester, and Deepak Kapur. Solving polynomial systems using a branch and prune approach. SIAM Journal on Numerical Analysis, 34(2):797–827, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28150.

Vanmaele:1997:ACV

[VMSB97] M. Vanmaele, K. W. Morton, E. Suli, and A. Borzi. Analysis of the cell vertex finite volume method for the Cauchy–Riemann equations. SIAM Journal on Numerical Analysis, 34(5): 2043–2062, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27638.

Vogel:1986:OCT

[Vog86] C. R. Vogel. Optimal choice of a truncation level for the truncated SVD solution of linear first kind integral equations when data are noisy. SIAM Journal on Numerical Analysis, 23(1):109–117, February 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Voigt:1971:OCI

[Voi71a] Robert G. Voigt. Orders of convergence for iterative procedures. SIAM Journal on Numerical Analysis, 8(2): 222–243, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Voigt:1971:RCC

[Voi71b] Robert G. Voigt. Rates of convergence for a class of iterative procedures. SIAM Journal on Numerical Analysis, 8(1):127–134, March 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Voss:1988:FOE

[Vos88] David Voss. A fifth-order exponentially fitted formula. SIAM Journal on Numerical Analysis, 25(3):670–678, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vandewalle:1993:DIM

[VP93] Stefan Vandewalle and Robert Piessens. On dynamic iteration methods for solving time-periodic differential equations. SIAM Journal on Numerical Analysis, 30(1): 286–303, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanderHouwen:1987:ERK

[vS87] P. J. van der Houwen and B. P. Sommeijer. Explicit Runge–Kutta (-Nyström) methods with reduced phase errors for computing oscillating solutions. SIAM Journal on Numerical Analysis, 24 (3):595–617, June 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanderHouwen:1989:DIR

[vS89a] P. J. van der Houwen and B. P. Sommeijer. Diagonally implicit Runge–Kutta– Nyström methods for oscillatory problems. SIAM Journal on Numerical Analysis, 26 (2):414–429, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vanderHouwen:1989:PLA

[vS89b] P. J. van der Houwen and B. P. Sommeijer. Phase-lag analysis of implicit Runge– Kutta methods. SIAM Journal on Numerical Analysis, 26 (1):214–229, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verwer:1997:SAO

[VS97] J. G. Verwer and B. P. Sommeijer. Stability analysis of an odd-even-line hopscotch method for three-dimensional advection-diffusion problems. SIAM Journal on Numerical Analysis, 34(1):376–388, February 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27697.

Varga:1980:IFM

[VSM80] R. S. Varga, E. B. Saff, and V. Mehrmann. Incomplete factorizations of matrices and connections with H-matrices. SIAM Journal on Numerical Analysis, 17 (6):787–793, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

vonPetersdorff:1997:MSK

[vSS97] Tobias von Petersdorff, Christoph Schwab, and Reinhold Schneider. Multiwavelets for second kind integral equations. SIAM Journal on Numerical Analysis, 34(6): 2212–2227, December 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27295.

Victory:1991:CAF

[VTG91] H. D. Victory, Jr., Garry Tucker, and Keshab Ganguly. The convergence analysis of fully discretized particle methods for solving Vlasov– Poisson systems. SIAM Journal on Numerical Analysis, 28(4):955–989, August 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Verschelde:1994:HEN

[VVC94] Jan Verschelde, Pierre Verlinden, and Ronald Cools. Homotopies exploiting Newton polytopes for solving sparse polynomial systems. SIAM Journal on Numerical Analysis, 31(3):915–930, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vassilevski:1995:AAM

[VW95] Panayot S. Vassilevski and Junping Wang. An application of the abstract multilevel theory to nonconforming finite element methods. SIAM Journal on Numerical Analysis, 32(1):235–248, February 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Vozovoi:1997:SAS

[VWI97] L. Vozovoi, A. Weill, and M. Israeli. Spectrally accurate solution of nonperiodic differential equations by the Fourier–Gegenbauer method. SIAM Journal on Numerical Analysis, 34 (4):1451–1471, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27881.

Wahlbin:1974:MNE

[Wah74] Lars Wahlbin. Maximum norm estimates for Friedrichs’ scheme in two dimensions. SIAM Journal on Numerical Analysis, 11(4):741–752, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wahlbin:1975:MNE

[Wah75] Lars Wahlbin. On maximum norm error estimates for Galerkin approximations to one-dimensional second order parabolic boundary value problems. SIAM Journal on Numerical Analysis, 12(2): 177–182, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wahba:1977:PAS

[Wah77] Grace Wahba. Practical approximate solutions to linear operator equations when the data are noisy. SIAM Journal on Numerical Analysis, 14 (4):651–667, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wahlbin:1980:RPS

[Wah80] Lars B. Wahlbin. A remark on parabolic smoothing and the finite element method. SIAM Journal on Numerical Analysis, 17(1):33–38, February 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wahlbin:1992:SBF

[Wah92] Lars B. Wahlbin. On superconvergence up to boundaries in finite element methods: A counterexample. SIAM Journal on Numerical Analysis, 29(4):937–946, August 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Walsh:1967:CSR

[Wal67] J. L. Walsh. On the convergence of sequences of rational functions. SIAM Journal on Numerical Analysis, 4 (2):211–221, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Walsh:1973:OCP

[Wal73] Robert T. Walsh. Optimization and comparison of partial difference methods. SIAM Journal on Numerical Analysis, 10(4):785–797, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Waldvogel:1983:PVL

[Wal83] Jörg Waldvogel. The period in the Volterra–Lotka predator-prey model. SIAM Journal on Numerical Analysis, 20(6):1264–1272, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). Dedicated to Peter Henrici on the occasion of his 60th birthday.

Walkington:1988:AWP

[Wal88] Noel J. Walkington. Acoustic wave propagation through flows with vorticity. SIAM Journal on Numerical Analysis, 25(3):533–549, June 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Walkington:1996:ACM

[Wal96] Noel J. Walkington. Algorithms for computing motion by mean curvature. SIAM Journal on Numerical Analysis, 33(6):2215– 2238, December 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26206.

Waldron:1998:ELI

[Wal98] Shayne Waldron. The error in linear interpolation at the vertices of a simplex. SIAM Journal on Numerical Analysis, 35(3): 1191–1200, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31315.

Walden:1999:FBM

[Wal99] Johan Waldén. Filter bank methods for hyperbolic PDEs. SIAM Journal on Numerical Analysis, 36(4):1183– 1233, August 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31313.

Wang:1992:CAR

[Wan92] Jun Ping Wang. Convergence analysis without regularity assumptions for multigrid algorithms based on SOR smoothing. SIAM Journal on Numerical Analysis, 29(4): 987–1001, August 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:1993:CAM

[Wan93a] Jun Ping Wang. Convergence analysis of multigrid algorithms for nonselfadjoint and indefinite elliptic problems. SIAM Journal on Numerical Analysis, 30(1):275– 285, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:1993:CAS

[Wan93b] Jun Ping Wang. Convergence analysis of the Schwarz algorithm and multilevel decomposition iterative methods. II. nonselfadjoint and indefinite elliptic problems. SIAM Journal on Numerical Analysis, 30 (4):953–970, September 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ward:1975:CSA

[War75]  Robert C. Ward.  The

combination shift QZ algorithm. SIAM Journal on Numerical Analysis, 12(6):835– 853, December 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ward:1977:NCM

[War77] Robert C. Ward. Numerical computation of the matrix exponential with accuracy estimate. SIAM Journal on Numerical Analysis, 14(4):600– 610, September 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wassyng:1982:SMR

[Was82] A. Wassyng. Solving ax = b: A method with reduced storage requirements. SIAM Journal on Numerical Analysis, 19(1):197–204, February 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Watson:1974:CBR

[Wat74] G. A. Watson. The calculation of best restricted approximations. SIAM Journal on Numerical Analysis, 11 (4):693–699, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Watson:1975:MEA

[Wat75] G. A. Watson. A multiple exchange algorithm for multivariate Chebyshev approximation. SIAM Journal on Numerical Analysis, 12(1): 46–52, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Watkins:1977:EBP

[Wat77] David S. Watkins. Error bounds for polynomial blending function methods. SIAM Journal on Numerical Analysis, 14(4):721–734, September 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Watson:1979:AGC

[Wat79] Layne T. Watson. An algorithm that is globally convergent with probability one for a class of nonlinear twopoint boundary value problems. SIAM Journal on Numerical Analysis, 16(3): 394–401, June 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Watkins:1981:DIV

[Wat81a] David S. Watkins. Determining initial values for stiff systems of ordinary differential equations. SIAM Journal on Numerical Analysis, 18(1): 13–20, February 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). In SIGNUM Meeting on Numerical Ordinary Differential Equations.

Watkins:1981:EIS

[Wat81b] David S. Watkins. Efficient initialization of stiff systems with one unknown initial condition. SIAM Journal on Numerical Analysis, 18(5):794– 800, October 1981. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wathen:1986:MIS

[Wat86]  A. J. Wathen.  Mesh-

independent spectra in the moving finite element equations. SIAM Journal on Numerical Analysis, 23(4): 797–814, August 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wissmann:1986:PSC

[WB86] Johannes W. Wissmann and Thomas Becker. Partially symmetric cubature formulas for even degrees of exactness. SIAM Journal on Numerical Analysis, 23(3):676–685, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wiegmann:1998:IIM

[WB98] Andreas Wiegmann and Kenneth P. Bube. The immersed interface method for nonlinear differential equations with discontinuous coefficients and singular sources. SIAM Journal on Numerical Analysis, 35 (1):177–200, February 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29378.

Weber:1985:MBI

[Web85] H. Weber. Multigrid bifurcation iteration. SIAM Journal on Numerical Analysis, 22 (2):262–279, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chu:1994:MRD

[wECGS94] K. w. E. Chu, W. Govaerts, and A. Spence. Matrices with rank deficiency two in eigenvalue problems and dynamical systems. SIAM Journal on Numerical Analysis, 31(2): 524–539, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weekes:1998:TWS

[Wee98] Suzanne L. Weekes. The travelling wave scheme for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 35(3):1249– 1270, June 1998. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29851.

Wegge:1966:DVN

[Weg66] Leon Wegge. On a discrete version of the Newton– Raphson method. SIAM Journal on Numerical Analysis, 3(1):134–142, March 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wegmann:1986:FNM

[Weg86] Rudolf Wegmann. On Fornberg’s numerical method for conformal mapping. SIAM Journal on Numerical Analysis, 23(6):1199–1213, December 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinstein:1969:SNE

[Wei69] Stanley E. Weinstein. Solution of nonlinear equations by iterative procedures which use approximation techniques. SIAM Journal on Numerical Analysis, 6(2):272–283, June 1969. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinstein:1971:PAF

[Wei71] Stanley E. Weinstein. Product approximations of functions of several variables. SIAM Journal on Numerical Analysis, 8(2):178–189, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinberger:1972:ONS

[Wei72a] H. F. Weinberger. On optimal numerical solution of partial differential equations. SIAM Journal on Numerical Analysis, 9(1):182–198, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinstein:1972:UAF

[Wei72b] S. E. Weinstein. Uniform approximation of functions through optimal partitioning. SIAM Journal on Numerical Analysis, 9(3):509– 517, September 1972. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinstein:1974:NSN

[Wei74a] Stanley E. Weinstein. On numerical solution of nonlinear equations through nonlinear interpolation or approximation. SIAM Journal on Numerical Analysis, 11(3): 461–471, June 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weiss:1974:AFP

[Wei74b] Richard Weiss. On the approximation of fixed points of nonlinear compact operators. SIAM Journal on Numerical Analysis, 11(3):550–553, June 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinmuller:1986:CSB

[Wei86] Ewa Weinmüller. Collocation for singular boundary value problems of second order. SIAM Journal on Numerical Analysis, 23(5):1062– 1095, October 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinmuller:1989:SSB

[Wei89] Ewa B. Weinmüller. Stability of singular boundary value problems and their discretization by finite differences. SIAM Journal on Numerical Analysis, 26(1):180– 213, February 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wei:1992:EUN

[Wei92a] Dongming Wei. Existence, uniqueness, and numerical analysis of solutions of a quasilinear parabolic problem. SIAM Journal on Numerical Analysis, 29(2): 484–497, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wei:1992:PTR

[Wei92b]  Mu Sheng Wei.  Pertur-

bation theory for the rankdeficient equality constrained least squares problem. SIAM Journal on Numerical Analysis, 29(5):1462–1481, October 1992. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinan:1993:CFM

[Wei93]  E. Weinan.  Convergence

of Fourier methods for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 30(3):650–674, June 1993. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Weideman:1994:CCE

[Wei94] J. A. C. Weideman. Computation of the complex error function. SIAM Journal on Numerical Analysis, 31 (5):1497–1518, October 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weideman:1995:ECE

[Wei95] J. A. C. Weideman. Erratum: “Computation of the complex error function” [SIAM J. Numer. Anal. 7 (1970), no. 1, 187–198]. SIAM Journal on Numerical Analysis, 32(1): 330–331, February 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [Gau70].

Welfert:1997:GPD

[Wel97] Bruno D. Welfert. Generation of pseudospectral differentiation matrices I. SIAM Journal on Numerical Analysis, 34 (4):1640–1657, August 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29554.

Wendroff:1968:WPP

[Wen68] Burton Wendroff. Well-posed problems and stable difference operators. SIAM Journal on Numerical Analysis, 5 (1):71–82, March 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Werner:1975:III

[Wer75] Helmut Werner. Interpolation and integration of initial value problems of ordinary differential equations by regular splines. SIAM Journal on Numerical Analysis, 12 (2):255–271, April 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Werner:1996:CHB

[Wer96] Bodo Werner. Computation of Hopf bifurcation with bordered matrices. SIAM Journal on Numerical Analysis, 33 (2):435–455, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weiser:1980:SEE

[WES80] A. Weiser, S. C. Eisenstat, and M. H. Schultz. On solving elliptic equations to moderate accuracy. SIAM Journal on Numerical Analysis, 17 (6):908–929, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wetton:1997:ASE

[Wet97a] Brian R. Wetton. Analysis of the spatial error for a class of finite difference methods for viscous incompressible flow. SIAM Journal on Numerical Analysis, 34(2):723–755, April 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27474.

Wetton:1997:EAC

[Wet97b] Brian R. Wetton. Error analysis for Chorin’s original fully discrete projection method and regularizations in space and time. SIAM Journal on Numerical Analysis, 34 (5):1683–1697, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28998.

Winker:1997:ATA

[WF97] Peter Winker and Kai-Tai Fang. Application of threshold-accepting to the evaluation of the discrepancy of a set of points. SIAM Journal on Numerical Analysis, 34 (5):2028–2042, October 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28607.

Wilson:1990:LTN

[WGS90] F. Wesley Wilson, R. Kent Goodrich, and Wendy Spratte. Lawson’s triangulation is nearly optimal for controlling error bounds. SIAM Journal on Numerical Analysis, 27(1): 190–197, February 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weideman:1986:SSM

[WH86] J. A. C. Weideman and B. M.

Herbst. Split-step methods for the solution of the nonlinear Schrödinger equation. SIAM Journal on Numerical Analysis, 23(3):485–507, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wu:1997:FEM

[WH97] Xiaonan Wu and Houde Han. A finite-element method for Laplace- and Helmholtz-type boundary value problems with singularities. SIAM Journal on Numerical Analysis, 34 (3):1037–1050, June 1997. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25848.

Wheeler:1973:PEE

[Whe73a] Mary Fanett Wheeler. A Priori L_{2} error estimates for Galerkin approximations to parabolic partial differential equations. SIAM Journal on Numerical Analysis, 10 (4):723–759, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wheeler:1973:EOO

[Whe73b] Mary Fanett Wheeler. L_{∞} estimates of optimal orders for Galerkin methods for one-dimensional second order parabolic and hyperbolic equations. SIAM Journal on Numerical Analysis, 10(5): 908–913, October 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wheeler:1973:OEE

[Whe73c] Mary Fanett Wheeler. An optimal L_{∞} error estimate for Galerkin approximations to solutions of two-point boundary value problems. SIAM Journal on Numerical Analysis, 10(5):914–917, October 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wheeler:1974:GPE

[Whe74] Mary Fanett Wheeler. A Galerkin procedure for estimating the flux for twopoint boundary value problems. SIAM Journal on Numerical Analysis, 11(4):764– 768, September 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wheeler:1975:GMP

[Whe75]  Mary F. Wheeler.  An

H^{−1} Galerkin method for parabolic problems in a single space variable. SIAM Journal on Numerical Analysis, 12(5): 803–817, October 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wheeler:1977:CFE

[Whe77] Mary Fanett Wheeler. A C^{0}-collocation-finite element method for two-point boundary value problems and one space dimensional parabolic problems. SIAM Journal on Numerical Analysis, 14(1): 71–90, March 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wheeler:1978:ECF

[Whe78] Mary Fanett Wheeler. An elliptic collocation-finite element method with interior penalties. SIAM Journal on Numerical Analysis, 15(1): 152–161, February 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

White:1979:SEM

[Whi79] Andrew B. White, Jr. On selection of equidistributing meshes for two-point boundary-value problems. SIAM Journal on Numerical Analysis, 16(3):472–502, June 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

White:1982:EFS

[Whi82a] R. E. White. An enthalpy formulation of the Stefan problem. SIAM Journal on Numerical Analysis, 19(6):1129– 1157, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

White:1982:NSE

[Whi82b] R. E. White. A numerical solution of the enthalpy formulation of the Stefan problem. SIAM Journal on Numerical Analysis, 19(6):1158– 1172, December 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

White:1982:NSI

[Whi82c] Andrew B. White, Jr. On the numerical solution of initial/boundary-value problems in one space dimension. SIAM Journal on Numerical Analysis, 19(4):683–697, August 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

White:1985:BAP

[Whi85] R. E. White. The binary alloy problem: existence, uniqueness, and numerical approximation. SIAM Journal on Numerical Analysis, 22(2): 205–244, April 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

White:1986:NPA

[Whi86a] R. E. White. A nonlinear parallel algorithm with application to the Stefan problem. SIAM Journal on Numerical Analysis, 23(3):639–652, June 1986. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Whitley:1986:SFR

[Whi86b] Robert Whitley. The stability of finite rank methods with applications to integral equations. SIAM Journal on Numerical Analysis, 23(1):118– 134, February 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Widder:1966:JLW

[Wid66]  D. V. Widder.  Joseph

Leonard Walsh. SIAM Journal on Numerical Analysis, 3 (2):171–172, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Widlund:1978:LMC

[Wid78] Olof Widlund. A Lanczos method for a class of nonsymmetric systems of linear equations. SIAM Journal on Numerical Analysis, 15(4): 801–812, August 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wigley:1966:CDA

[Wig66] Neil M. Wigley. On the convergence of discrete approximations to solutions of mixed boundary value problems. SIAM Journal on Numerical Analysis, 3(3):372– 382, September 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wigley:1987:SIF

[Wig87] Neil M. Wigley. Stress intensity factors and improved convergence estimates at a corner. SIAM Journal on Numerical Analysis, 24(2): 350–354, April 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wilson:1969:FCG

[Wil69] G. Wilson. Factorization of the covariance generating function of a pure moving average process. SIAM Journal on Numerical Analysis, 6 (1):1–7, March 1969. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wilson:1970:NSC

[Wil70] M. Wayne Wilson. Necessary and sufficient conditions for equidistant quadrature formula. SIAM Journal on Numerical Analysis, 7(1): 134–141, March 1970. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Williams:1972:NCA

[Wil72] Jack Williams. Numerical Chebyshev approximation in the complex plane. SIAM Journal on Numerical Analysis, 9(4):638–649, December 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Willson:1973:NPM

[Wil73] A. N. Willson, Jr. A note on pairs of matrices and matrices of monotone kind. SIAM Journal on Numerical Analysis, 10(4):618–622, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wilhelmsen:1974:NCC

[Wil74] Don R. Wilhelmsen. Nonnegative cubature on convex sets. SIAM Journal on Numerical Analysis, 11(2): 332–346, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wilhelmsen:1978:OQP

[Wil78] D. R. Wilhelmsen. Optimal quadrature for periodic analytic functions. SIAM Journal on Numerical Analysis, 15 (2):291–296, April 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Williams:1979:CCR

[Wil79]  Jack Williams.  Charac-

terization and computation of rational Chebyshev approximations in the complex plane. SIAM Journal on Numerical Analysis, 16(5):819– 827, October 1979. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wildenauer:1981:CDA

[Wil81] Peter Wildenauer. Construction of domains with all solutions, and the existence of extreme solutions. SIAM Journal on Numerical Analysis, 18(5):801–807, October 1981. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wimp:1970:DFI

[Wim70] Jet Wimp. Derivative-free iteration processes. SIAM Journal on Numerical Analysis, 7(3):329–334, September 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wimp:1977:NMA

[Wim77] Jet Wimp. New methods for accelerating the convergence of sequences arising in Laplace transform theory. SIAM Journal on Numerical Analysis, 14(2): 194–204, April 1977. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Winther:1980:SSC

[Win80] Ragnar Winther. Some superlinear convergence results for the conjugate gradient method. SIAM Journal on Numerical Analysis, 17(1): 14–17, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Winther:1982:FEM

[Win82] Ragnar Winther. A finite element method for a version of the Boussinesq equation. SIAM Journal on Numerical Analysis, 19(3):561–570, June 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wixom:1978:INC

[Wix78] James A. Wixom. Interpolation to networks of curves in E^{3}. SIAM Journal on Numerical Analysis, 15(6):1178– 1193, December 1978. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weinan:1995:PMI

[WL95] E. Weinan and Jian-Guo Liu.

Projection method. I. convergence and numerical boundary layers. SIAM Journal on Numerical Analysis, 32(4): 1017–1057, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Whitney:1967:TAR

[WM67] T. M. Whitney and R. K. Meany. Two algorithms related to the method of steepest descent. SIAM Journal on Numerical Analysis, 4(1): 109–118, March 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wollman:1996:NAO

[WO96] Stephen Wollman and Ercument Ozizmir. Numerical approximation of the onedimensional Vlasov–Poisson system with periodic boundary conditions. SIAM Journal on Numerical Analysis, 33 (4):1377–1409, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23358.

Wohlmuth:1999:HPE

[Woh99] Barbara I. Wohlmuth. Hierarchical A posteriori error estimators for mortar finite element methods with Lagrange multipliers. SIAM Journal on Numerical Analysis, 36 (5):1636–1658, October 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33051.

Wolfe:1980:MKA

[Wol80] M. A. Wolfe. A modification of Krawczyk’s algorithm.

SIAM Journal on Numerical Analysis, 17(3):376–379, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wolkenfelt:1983:RBR

[Wol83] P. H. M. Wolkenfelt. On the relation between the repetition factor and numerical stability of direct quadrature methods for second kind Volterra integral equations. SIAM Journal on Numerical Analysis, 20(5):1049– 1061, October 1983. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Womble:1989:FTM

[Wom89] David E. Womble. A fronttracking method for multiphase free boundary problems. SIAM Journal on Numerical Analysis, 26(2): 380–396, April 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wood:1971:PEI

[Woo71] Winifred L. Wood. Periodicity effects on the iterative solution of elliptic difference equations. SIAM Journal on Numerical Analysis, 8 (2):439–464, June 1971. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wozniakowski:1974:MSI

[Woz74a] H. Wozniakowski. Maximal stationary iterative methods for the solution of operator equations. SIAM Journal on Numerical Analysis, 11(5): 934–949, October 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wozniakowski:1974:REA

[Woz74b] H. Wozniakowski. Rounding error analysis for the evaluation of a polynomial and some of its derivatives. SIAM Journal on Numerical Analysis, 11(4):780–787, September 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wozniakowski:1975:GIM

[Woz75] H. Wozniakowski. Generalized information and maximal order of iteration for operator equations. SIAM Journal on Numerical Analysis, 12(1): 121–135, March 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Watanabe:1985:SDF

[WR85] Daniel S. Watanabe and Mitchell G. Roth. The stability of difference formulas for delay differential equations. SIAM Journal on Numerical Analysis, 22(1):132– 145, February 1985. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Werner:1984:CSB

[WS84] B. Werner and A. Spence.

The computation of symmetrybreaking bifurcation points. SIAM Journal on Numerical Analysis, 21(2):388–399, April 1984. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wathen:1993:FIS

[WS93] Andrew Wathen and David Silvester. Fast iterative solution of stabilised Stokes systems. Part I: Using simple diagonal preconditioners. SIAM Journal on Numerical Analysis, 30(3):630–649, June 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weideman:1988:ESO

[WT88] J. A. C. Weideman and Lloyd N. Trefethen. The eigenvalues of second-order spectral differentiation matrices. SIAM Journal on Numerical Analysis, 25(6):1279– 1298, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wu:1996:DFT

[Wu96a] Lixin Wu. Dufort–Frankeltype methods for linear and nonlinear Schrödinger equations. SIAM Journal on Numerical Analysis, 33(4): 1526–1533, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27063.

Wu:1996:USS

[Wu96b] Zi-Niu Wu. Uniqueness of steady-state solutions for difference equations on overlapping grids. SIAM Journal on Numerical Analysis, 33 (4):1336–1357, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25609.

Wuytack:1970:SRP

[Wuy70] L. Wuytack. Some remarks on a paper of D. G. Moursund. SIAM Journal on Numerical Analysis, 7(2):233–237, June 1970. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wuytack:1974:SAR

[Wuy74] Luc Wuytack. On some aspects of the rational interpolation problem. SIAM Journal on Numerical Analysis, 11 (1):52–60, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Westreich:1979:NBS

[WV79] David Westreich and Yaakov L. Varol. Numerical bifurcation at simple eigenvalues. SIAM Journal on Numerical Analysis, 16(3):538–546, June 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Werner:1983:CPO

[WW83] H. Werner and L. Wuytack.

On the continuity of the Padé operator. SIAM Journal on Numerical Analysis, 20(6): 1273–1280, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weiser:1988:CBC

[WW88] Alan Weiser and Mary Fanett Wheeler. On convergence of block-centered finite differences for elliptic problems. SIAM Journal on Numerical Analysis, 25(2): 351–375, April 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Walker:1990:LCS

[WW90] Homer F. Walker and Layne T.

Watson. Least-change secant update methods for underdetermined systems. SIAM Journal on Numerical Analysis, 27(5):1227–1262, October 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:1993:ECLa

[WW93a] Jing Hua Wang and Gerald Warnecke. On entropy consistency of large time step schemes. I. the Godunov and Glimm schemes. SIAM Journal on Numerical Analysis, 30 (5):1229–1251, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:1993:ECLb

[WW93b] Jing Hua Wang and Gerald Warnecke. On entropy consistency of large time step schemes. II. approximate Riemann solvers. SIAM Journal on Numerical Analysis, 30 (5):1252–1267, October 1993. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wynn:1964:SRD

[Wyn64] P. Wynn. On some recent developments in the theory and application of continued fractions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(??):177– 197, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic).

Wynn:1966:CSE

[Wyn66] P. Wynn. On the convergence and stability of the epsilon algorithm. SIAM Journal on Numerical Analysis, 3 (1):91–122, March 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wynn:1968:UPT

[Wyn68] Peter Wynn. Upon the Padé table derived from a Stieltjes series. SIAM Journal on Numerical Analysis, 5(4):805– 834, December 1968. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xu:1989:CBT

[Xu89] Zong Ben Xu. A computational ball test for the existence of solutions to nonlinear operator equations. SIAM Journal on Numerical Analysis, 26(1):239–248, February 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xu:1992:NCI

[Xu92] Jinchao Xu. A new class of iterative methods for nonselfadjoint or indefinite problems. SIAM Journal on Numerical Analysis, 29(2): 303–319, April 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xu:1996:TGD

[Xu96] Jinchao Xu. Two-grid discretization techniques for linear and nonlinear PDEs. SIAM Journal on Numerical Analysis, 33(5):1759– 1777, October 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23294.

Yang:1995:TDQ

[YA95] Yajun Yang and Kendall E. Atkinson. Two-dimensional quadrature for functions with a point singularity on a triangular region. SIAM Journal on Numerical Analysis, 32 (3):969–983, June 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yamamoto:1975:RRW

[Yam75]  Tetsuro Yamamoto.  The

Rayleigh–Ritz and Weinstein– Bazley methods applied to a class of ordinary differential equations of the second order. II. SIAM Journal on Numerical Analysis, 12(3): 428–438, June 1975. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yamamoto:1998:NVM

[Yam98] Nobito Yamamoto. A numerical verification method for solutions of boundary value problems with local uniqueness by Banach’s fixed-point theorem. SIAM Journal on Numerical Analysis, 35 (5):2004–2013, October 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30449.

Yan:1994:FNS

[Yan94] Yi Yan. A fast numerical solution for a second kind boundary integral equation with a logarithmic kernel. SIAM Journal on Numerical Analysis, 31(2):477–498, April 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yan:1996:AEE

[Yan96] Yin Yan. Attractors and error estimates for discretizations of incompressible Navier– Stokes equations. SIAM Journal on Numerical Analysis, 33 (4):1451–1472, August 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24809.

Yang:1999:WEI

[Yan99] Huanan Yang. On wavewise entropy inequality for high resolution schemes II: Fully discrete MUSCL schemes with exact evolution in small time. SIAM Journal on Numerical Analysis, 36(1): 1–31, February 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28149.

Yavneh:1995:MSF

[Yav95]  Irad Yavneh.  Multigrid

smoothing factors for redblack Gauss–Seidel relaxation applied to a class of elliptic operators. SIAM Journal on Numerical Analysis, 32(4): 1126–1138, August 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yeung:1996:PEC

[YB96] Wing Kwong Yeung and Kenneth P. Bube. A Priori estimates and convergence for the discrete forward and inverse problems of reflection seismology. SIAM Journal on Numerical Analysis, 33(1): 247–279, February 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yen:1993:CEM

[Yen93] Jeng Yen. Constrained equations of motion in multibody dynamics as ODEs on manifolds. SIAM Journal on Numerical Analysis, 30(2): 553–568, April 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yan:1992:OSC

[YF92] Yi Yan and Graeme Fairweather. Orthogonal spline collocation methods for some partial integrodifferential equations. SIAM Journal on Numerical Analysis, 29(3): 755–768, June 1992. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ying:1995:CVM

[Yin95] Lung An Ying. Convergence of vortex methods for threedimensional Euler equations in bounded domains. SIAM Journal on Numerical Analysis, 32(2):542–559, April 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yakowitz:1978:WMC

[YKS78] S. Yakowitz, J. E. Krimmel, and F. Szidarovszky. Weighted Monte Carlo integration. SIAM Journal on Numerical Analysis, 15(6): 1289–1300, December 1978. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Young:1972:GPC

[You72a] David M. Young. Generalizations of property A and consistent ordering. SIAM Journal on Numerical Analysis, 9 (3):454–463, September 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Young:1972:CLS

[You72b] David M. Young. On the consistency of linear stationary iterative methods. SIAM Journal on Numerical Analysis, 9(1):89–96, March 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ypma:1984:LCI

[Ypm84] T. J. Ypma. Local convergence of inexact Newton methods. SIAM Journal on Numerical Analysis, 21(3): 583–590, June 1984. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yarvin:1999:IFM

[YR99] Norman Yarvin and Vladimir Rokhlin. An improved fast multipole algorithm for potential fields on the line. SIAM Journal on Numerical Analysis, 36(2):629–666, April 1999. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32923.

Yan:1989:MGI

[YS89] Y. Yan and I. H. Sloan. Mesh grading for integral equations of the first kind with logarithmic kernel. SIAM Journal on Numerical Analysis, 26 (3):574–587, June 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yserentant:1986:MLS

[Yse86] Harry Yserentant. On the multi-level splitting of finite element spaces for indefinite elliptic boundary value problems. SIAM Journal on Numerical Analysis, 23(3):

581–595, June 1986. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ying:1994:FDC

[YZ94] Lung An Ying and Ping Wen Zhang. Fully discrete convergence estimates for vortex methods in bounded domains. SIAM Journal on Numerical Analysis, 31(2): 344–361, April 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zanna:1999:CRC

[Zan99] A. Zanna. Collocation and relaxed collocation for the Fer and the Magnus expansions. SIAM Journal on Numerical Analysis, 36 (4):1145–1182, August 1999. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32661.

Zhang:1996:NMS

[ZB96] Zhimin Zhang and Ivo Babuška. A numerical method for steady state free boundary problems. SIAM Journal on Numerical Analysis, 33(6):2184–2214, December 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/22985.

Zedek:1966:ASO

[Zed66]  Mishael Zedek.  On ap-

proximation by solutions of ordinary linear differential equations. SIAM Journal on Numerical Analysis, 3(2): 360–365, June 1966. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zemla:1995:FSD

[Zem95] Adam Zemła. On the fundamental solutions for the difference Helmholtz operator. SIAM Journal on Numerical Analysis, 32(2): 560–570, April 1995. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zennaro:1985:OSC

[Zen85]  Marino Zennaro.  One-

step collocation: uniform superconvergence, predictorcorrector method, local error estimate. SIAM Journal on Numerical Analysis, 22(6): 1135–1152, December 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zarnowski:1991:FDS

[ZH91] Roger Zarnowski and David Hoff. A finite-difference scheme for the Navier–Stokes equations of one-dimensional, isentropic, compressible flow. SIAM Journal on Numerical Analysis, 28(1):78–112, February 1991. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhao:1994:CFD

[ZH94] Jing Zhao and David Hoff.

A convergent finite-difference scheme for the Navier–Stokes equations of one-dimensional, nonisentropic, compressible flow. SIAM Journal on Numerical Analysis, 31(5):1289– 1311, October 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:1995:IEE

[Zha95] Zhen Yue Zhang. An inverse extreme eigenpair problem and its parallel iterative solution for symmetric tridiagonal matrices. SIAM Journal on Numerical Analysis, 32 (5):1620–1634, October 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:1996:TLS

[Zha96] Xuejun Zhang. Two-level Schwarz methods for the biharmonic problem discretized conforming C^{1} elements. SIAM Journal on Numerical Analysis, 33(2): 555–570, April 1996. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:1997:ASQ

[Zha97]  Zhimin Zhang.  Anal-

ysis of some quadrilateral nonconforming elements for incompressible elasticity. SIAM Journal on Numerical Analysis, 34(2): 640–663, April 1997. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28249.

Zhou:1983:LFE

[Zho83] S. Z. Zhou. The linear finite element method for a twodimensional singular boundary value problem. SIAM Journal on Numerical Analysis, 20(5):976–984, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zirilli:1982:SNS

[Zir82] Francesco Zirilli. The solution of nonlinear systems of equations by second order systems of O.D.E. and linearly implicit A-stable techniques. SIAM Journal on Numerical Analysis, 19(4):800– 815, August 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zlamal:1965:AEE

[Zla65] Milos Zlamal. Asymptotic error estimates in solving elliptic equations of the fourth order by the method of finite differences. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2): 337–344, ???? 1965. ISSN 0887-459X (print), 1095-7170 (electronic).

Zlamal:1967:DEE

[Zla67] Milos Zlamal. Discretization and error estimates for elliptic boundary value problems of the fourth order. SIAM Journal on Numerical Analysis, 4(4):626–639, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zlamal:1973:CEF

[Zla73] Milos Zlamal. Curved elements in the finite element method. I. SIAM Journal on Numerical Analysis, 10(1): 229–240, March 1973. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zlamal:1974:CEF

[Zla74] Milos Zlamal. Curved elements in the finite element method. II. SIAM Journal on Numerical Analysis, 11(2): 347–362, April 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zlatev:1980:SPS

[Zla80] Zahari Zlatev. On some pivotal strategies in Gaussian elimination by sparse technique. SIAM Journal on Numerical Analysis, 17(1):18– 30, February 1980. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zlatev:1982:UIR

[Zla82] Zahari Zlatev. Use of iterative refinement in the solution of sparse linear systems. SIAM Journal on Numerical Analysis, 19(2): 381–399, April 1982. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zemanian:1967:NED

[ZT67] A. H. Zemanian and R. P. Tewarson. The numerical evaluation of distributional transforms. SIAM Journal on Numerical Analysis, 4(2): 271–282, June 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zeng:1994:MSI

[ZW94] S. Zeng and P. Wesseling. Multigrid solution of the incompressible Navier–Stokes equations in general coordinates. SIAM Journal on Numerical Analysis, 31(6):1764– 1784, December 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zwart:1973:MSN

[Zwa73] Philip B. Zwart. Multivariate splines with nondegenerate partitions. SIAM Journal on Numerical Analysis, 10 (4):665–673, September 1973. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhu:1994:MMM

[ZWW94] Jiang Zhu, Hong Wei Wu, and Yuan Ming Wang. A mixed method for the mixed initial boundary value problems of equations of semiconductor devices. SIAM Journal on Numerical Analysis, 31 (3):731–744, June 1994. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zeng:1998:MGC

[ZZ98] Jinping Zeng and Shuzi Zhou.

On monotone and geometric convergence of Schwarz methods for two-sided obstacle problems. SIAM Journal on Numerical Analysis, 35(2):600–616, April 1998. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28892.